A Complete Bibliography of Publications in Computer Physics Communications: 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 1.25

Title word cross-reference

(O_{m}|α_{n}) [De 02]. ∗ [Tos08]. 0, 1 [DC00]. 1 [KNU00, SQ03, YRR07]. 1/2 [PS08]. 10 [GRR01]. 1s [JK01]. 2 [AH02, BJ05b, CSW02, FM00, GFP00, GMBC08, KNU00, MLF07, NHS07, PAT^{+}09, SMV01, SBCZ08, TPYV03, WCG04]. 2.0 [RS00]. 2p_{+} [JK01]. 3 [Aok01, BD00, Bal07, BFH05, CCFG05, Cha04, CBBJ02, DGV08, EL04, FDM07, FV02, GBM02, GS01a, GBD03, GBA01, GMBC08, HG02a, HBRS05, JW02, KKSR04, KMZZ05, MZB^{+}04, MNV00, NSYZ02, PCV06, QR01, QTL06, RLRR06, SJCM04, SG04b, SBB03, SG01, TAM04, WLH00, WCG04, WHL^{+}07, XON08]. 3/2 [DKC08]. 4 [CBBJ02, Tör00, WLH00]. 4fγ [KJ04]. 6 [FMD07]. 7 × 7 [LK07]. 8 [GCP^{+}02]. ^{−1} [CRUV00]. ^{2}Π [HC08]. ^{3} [MNYY00b]. ^{3}M [Eas08]. ^{4} [KM00b]. ^{(}R) [MBKJ09]. _{0.3} [JK01]. _{0.7} [JK01]. 12 [KMD^{+}02, NT05]. _{2} [CM02a, EVL00, Gha05, Gro01, HTA08, TAP01, VMMB02]. _{4} [KPL07]. _{6} [KMD^{+}02, NT05]. _{m} [MTJ02]. _{n} [MTJ02]. _{n}^{+} [GCP^{+}02]. ABC + D → AB + CD [MGG08]. α [LDZ^{+}08, RJCH00]. α → [KMH02]. B [NM03]. B_{s}^{0} → J/ψϕ [BS04a]. B_{c} [CDEW04]. B_{d}^{0} → J/ψK^{∗} [BS06b]. B_{s}^{0} → J/ψϕ [BS06b]. Bc [CWW06a, CWW06b]. β [LDZ^{+}08, MCC05, RPD^{+}05]. β = ∞ [CM03]. β^{−} [RCGC00]. BR[B → Xsy] [DGS08]. B → K^{⋆}y [Mah08b]. χ^{2} [BBBD06]. χ^{2}ϕ [Sav01]. CP (N − 1) [BPRW06]. D [DEW00, PFG06a, Sch06a, Tat07]. d = 3 [RLU00]. δf [AH03, WTH^{+}04]. e^{+}e^{−} [ABM03, AAC^{+}06, BBC^{+}01b, DDRW03, JWW00b, Por03, TA00a, TA00b]. e^{+}e^{−} → 4f [KJ04]. e^{+}e^{−} → ff̄nγ [JKW06]. e^{+}e^{−} → ffνnγ [JKW00]. e^{+}e^{−} → ππγ [SVP09]. ℓ_{1} [Lor08]. ep [Abe01, Bel01]. ϵ [GHLW03]. η [Mic07]. ηb [QWWZ09]. F [GKP^{+}06, IIK^{+}08]. F_{1} [CGM01, CG04]. F_{4} [Niu00]. F_{m}(z) [TAKN02]. ffbb [BBK^{+}07]. G [CNMC09, Kon02]. γ [JKW00, JKW06, KJ04, MKJ^{+}05]. H [MN07, BBB^{+}00, CT00]. HΨ [Mei01]. I nu(z) [Tho04a, TB87]. J [CA09, Dev05]. jj [GF01, GFG01, GF02a, GF02b, SJF07]. k [LLT^{+}02, LCHJ09, MSY07, TMN01, Yan03b]. k = 0 [FSB09]. K_{ν}(z) [TB87, Tho04a]. K_{⊥} [BCCW03]. k · p [PAD07]. l [Mic07]. $l/Mflops [FKP03]. L_{1} [Dem03]. ⟨010⟩ [LTT09]. ⟨110⟩ [LTT09]. ll [LDBG08]. m [Yan03b]. mK(m, n, k) [Yan03a]. N [ADBF03, BAD01, GBTM07, GPW^{+}09, MWA01, SvAS01, AKZ00, RF05a, RF06a, RF07, RF08, Yan03b]. n = 3 [GCP^{+}02]. ∇ · B⃗ = 0 [Cha04]. n ≈ 1000 [HB05]. ννγ [KFI^{+}01]. O(3) [CSW02]. O(α_{s}^{3}) [BKKS09]. O(N) [ÖDÇ02]. P [CWW06a, Sim09, MNYY00b]. {P, N} [LVV06]. P^{3}M [TC06]. Φ [SVP09]. ±J [HG02a, RLU00]. pp/pp̄ [TKK^{+}06]. pp/ppν [TSA^{+}03]. Q [TL09]. R [De 02, BBB^{+}00, HS03, MN01, SSB^{+}09, SNBB02, Zat06, CLFH07]. r_{12}^{k} exp(−γr_{12}^{2}) [WD04]. r_{12} exp(−γr_{12}^{2}) [SKH02a]. Σ [Kar02]. SO(5) ⊃ SO(3) [CRW09]. \sqrt{3} × \sqrt{3} [RPD^{+}05]. SU(2) [CM03]. T [ÇHM00, DEW00, BRdAHK04b, CA09, dAK01]. tanh [Wan09b]. τ [BEM^{+}02]. τ^{+}τ^{−}(nγ) [PSW00]. ττγ [PSW00]. τ → 4π [BEM^{+}02]. θ [GWK09]. × [CW02]. T l|W | [CDQF07]. U(1) [BB09a]. U(3) ⊃ SO(3) [Dra01]. U_{q}[gl(m|n)] [De 02]. W [JPS^{+}01b, JPS^{+}01a]. w = 5 [Blü09]. X [MN07, CC04]. Ξbb [CWW07]. Ξbc [CWW07]. Ξcc [CWW07]. Z [SBL^{+}04, Mic07, Pet04, RG04]. Z^{d} [HJ02].

-2 [CW02]. -a [MP03]. -aluminum [RJCH00]. -body [KNU00, ADBF03, BAD01, GPW^{+}09, SvAS01]. -centre [PAT^{+}09]. -coupled [SJF07]. -coupling [GF01, GFG01, GF02b]. -cut [CLFH07]. -D [Bal07, FMD07, MLF07]. -dim [GBD03]. -dimensional [KNU00, Tat07, Sch06a]. -direction [LTT09]. -electron [MWA01]. -factor [Kon02]. -files [BBB^{+}00]. -functions [PFG06a]. -Lagrange [LCHJ09]. -matrices [BRdAHK04b]. -matrix [dAK01, BBB^{+}00, ÇHM00, DEW00, MN01, SSB^{+}09, SNBB02, Zat06]. -method [PAD07]. -orbit [Dev05]. -pair [JPS^{+}01b, JPS^{+}01a]. -penalized [Lor08]. -pinch [Pet04, RG04]. -pinches [SBL^{+}04]. -Point [Tör00, TMN01].

-product [Tos08]. -qubit [RF05a, RF06a, RF07, RF08]. -ratio [HS03]. -ray [MKJ^{+}05]. -regime [GHLW03]. -SHAKE [GWK09]. -SiC [MCC05, RPD^{+}05]. -space [CC04]. -stability [Sim09]. -state [CBBJ02]. -theorem [CT00]. -to- [LDZ^{+}08]. -wave [CWW06a].

0 [Par04, Sha04, Vio04, Wan00]. 0-471-98495-7 [Wan00]. 0.0 [BLS01]. 0994 [Hoo04].

1 [CRPC08]. 1.0 [KRW03]. 1.02 [JS07]. 1.3 [BBPS06]. 1/f [RDSS01a]. 10-state [BKB02b]. 100 [NSYZ02]. 111 [NSYZ02]. 124 [JKW06]. 125 [Ixa01]. 126 [TA00a]. 127 [KM01b]. 128-bit [DH00]. 130 [MOS01]. 134 [AA01b]. 144 [DVL^{+}04]. 146 [Voi03]. 147 [LPR04, MSHP20]. 147/3 [MSHP20]. 150 [Ida03a]. 151 [Yos07]. 153 [Hon04]. 156 [WA07]. 161 [TND05]. 1611 [Hoo04]. 1611-0994 [Hoo04]. 162 [SM06a]. 166 [GDAG05a]. 174 [AAB^{+}07]. 175 [FNR^{+}07]. 177 [CSC^{+}08]. 178 [CGG^{+}09]. 179 [CGVA09a, Poi09]. 180 [Nat10, Ram10, Ras17, Wu10].

2 [Ano09u, ABC^{+}01, BCAD06, Hah09, HM08, Laf03, SS09a, SIE04]. 2-dimensional [SD07]. 2.0 [ABM03, BBPS07a, BCKT09, CGVA08, GS01b, HTNFBS06a]. 2.0.7 [BBPS07b]. 2.1 [HTNFBS06b, Pöt00]. 2.7 [TKK^{+}06]. 2000 [BJS00]. 2001 [DVL^{+}02, Koc02]. 2003 [Ano04b, Ano04c]. 21st [Nov02, Swe02, You02]. 2DRMP [BNFM^{+}09, SSB^{+}09].

3 [ABF^{+}01, Hah01, Laf03, Par04, Sha04, Vio04]. 3-540-43289-2 [Laf03]. 3-540-43416-X [Hoo04]. 3-540-43908-0 [Par04]. 3-540-44363-0 [Sha04]. 3-540-66783-0 [Vio04]. 3.0 [MMEH08, Sem09]. 3.1.1 [PDL04]. 3.25 [SS02a]. 31 [KS08]. 32-bit [JC01]. 36 [Tho04b]. 39 [NY08]. 3D [Wol03].

4 [SHI02, Wel01]. 47 [Tho04a]. 4B [KN07a]. 4th [ISX05]. 4th-order [ISX05].

6.2 [KRW03]. 6.3 [KRW03]. 6D [FDM07].

7 [BLS01, Wan00]. 7-0.0 [BLS01]. 77 [Dem03, Dem06].

90 [BRdAHK04a, DG08, KLM00, PS08, SS09a]. 90/95 [DG08]. 95 [vH06].

A-priori [DVG05]. a.c [KSTL03]. Ab-initio [PCCD09, MKB02, Nik03]. ABC [SCM00, TAP01]. ABCD [ATP01]. Abel [CTR00]. ABINIT [Ano09a]. absolute [SVMT00, Sus01]. absorbing [UOM01]. absorption [JK01, KV07, NI01, VAH04, Yos03, Yos07]. abstract [Por00]. ac [WGY01, iOY01]. accelerate [BK05c]. Accelerated [LBP^{+}09, WH00, FIT03, LdG^{+}07, GPW^{+}09]. Accelerating [BBD^{+}09, LSVMW08, MBKJ09]. acceleration [BDV04, KH06, NRR01, Wen01]. accelerator [KSHP02, SEE^{+}03, TIN^{+}09]. Accepting [FHF00]. Access [BNO^{+}01, Ano09s, Han00, ZC09]. accessible [BDH^{+}05]. Accord [Ano09u, Hah09]. Accounting [YvG05]. accuracy [FG04, MBR01, RMK05, SSZ01, TAKN02, TB87]. Accurate [KDW00, Mam08, SMH^{+}01, UK02a, UYK^{+}04, itVPG08, AH02, BT01, CWSH08, CD04, DWZS05, EÅU05, LTG09, Moh08, NM01a, Sim00, SR01a, VKM^{+}05, WW05, YB02b]. AcerMC [KRW03]. Acetylene [SPC^{+}05]. achievements [Mül05]. acid [CTI07]. acid-water [CTI07]. aCLIMAX [RC04]. acoustic [NN06]. Acoustical [WP06, SWP03]. acquisition [GMO03]. acting [JKKT00]. actinides [EÅ01]. action [Dür05, Elb05, HJM02]. action-computer [HJM02]. actions [KT04]. activated [RLU01]. Active [AC05a, KL01, LOY07]. activity [EYJ07, IN02]. adaptability [GGL03]. adaptation [BTS06]. adapted [Fra02, LLY07]. adapting [Jad00]. Adaptive [Bre05, CC07, CNDC09, Dys02, KNTG03, PM01, SF00, Zit09, ABRS12, BS08, DPSG06, FS08, GHPS04, HCH^{+}06, HJZ09, KTT09, KCH00, KKF^{+}04, LLCS01, LHS^{+}06, MOM^{+}00, MM05, Ros04, SJCM04, SD07, SFR05, TC06, WPL02, ZS03, Zie04, Zie05, Zie08, vdHKM08]. adaptive-resolution [ABRS12]. Addendum [Ram10, Voi03, Yos07]. adder [BDLT02]. addition [CWW06a, SI01, VEG08]. additional [VC08]. ADF95 [Str05]. ADI [BH05, Mah08a, SBD^{+}06, Zie04]. ADI-based [Zie04]. ADI-like [SBD^{+}06]. adiabatic [APV00, BBOY08, BC05, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, EKW09, KM05, MN01]. adiabatic-by-sector [MN01]. adjacency [LB04]. adjoined [CGVA09b]. adjustable [Bar03]. adjustment [bHhL07]. adjustment-stabilization [bHhL07]. admixed [BK01]. adsorbate [LJ01, YN05b]. adsorbed [WMNS09]. Adsorption [LK07, MB05a, PLL07, SPV07]. Adsorption-induced [MB05a]. advanced [CNFR01, GCD06, NK07]. advances [Ano04b]. advection [AGV00, Ida02, ISX05]. advection-diffusion [AGV00]. ADZH v2 0 [CGVA09a]. AEAA v1 0 [CGG^{+}09]. aerosol [OPB^{+}09]. affine [MP01a, PZ01]. AFMM [VCCS05]. after [ZSdD^{+}08]. Against [CJC09]. AgCo [PMH08]. agent [LNV^{+}09]. agents [BA09]. agglomerates [CGC^{+}09]. Agglomeration [WDF^{+}02]. aggregates [MDS09, Vor02]. aggregation [KP01, Vor02]. aided [ZSSA00]. air [EFG^{+}00, KMZZ05, NV09, RMVQ07]. Airy [Ixa07a]. aITALC [LR06]. Al [JK01, KNSY07b, LK07]. ALE [DMR01, DMR02]. ALEGRA [RG04]. ALEGRA-HEDP [RG04]. Alfvén [DJ04, GIME02]. Algebra [Koc02, BKKS09, BCV03, MG08b, Pee07, Pue06, Rob00, Wei04]. Algebraic [Blü04, Har00, JC07, SKF05, BS03, dSB00, IBA00, JC08b, MKS07]. algebras [AF05, dSB00]. Algorithm [WHCL07, ALV05, AH03, ABER00, Asc08, AK03, BVY05, BPRW06, BH05, BJ08, BB04b, BR09, BG09, BJ03, BSS09, BS08, Bru00a, BCCW03, CN01, CNFR01, CJK09, CKV04, ÇHM00, CHP04, CAF^{+}03, Col07, CL08b, DC00, DGAG06, DB08, Dzu09, Eas08, FMD07, FNR^{+}06, FNR^{+}07, FMN01, GFP00, GDAG05a, GDAG05b, GFS03, HSJ02, Huj05, HLB06, IVD03, JAT03, JGJ09, KKK06, KKS04, KP01, KM00a, KM01b, Kos05, LM02a, LOL06, LTG09, LZC^{+}08, LB04, Lüs05, MDS09, MY00b, MBG03, MS08a, MHR^{+}07, Mei01, MP05, Nak07, Nak08, NH09, NP01b, iNKNV08, NRR01, OCK^{+}00, OCS^{+}08, PD08, PCE^{+}08, PSP^{+}03, PSH06, PCYC02, PB09b, RD05, RK05, RY00, RLU00, ST02, iSAK^{+}08, She03, SHX02, SOYN01, SHI02, SKNV05, SWC^{+}03, SPS09, Sug01, SWP03, Tak00, TG00, TPYV03]. algorithm [TB85, Tho04b, Tót06, TL08a, TCO00, UJSW06, Wan05a, WWF08, WT01, WK02, YH02, Yan02, Yan03c, Yan03d, YZW02, YWLC04, YD07, Zha08, ZLL09, vHK00, ML03]. algorithm/simplex/spatial [WMNS09]. Algorithmic [AHS09, VPCK04, XC03]. Algorithms [BMSG01, CTSZ07, FBB01, VBFM05, AAP03, ACK05, BBD^{+}09, BOG^{+}07, CC07, CMR01, CM03, FdO09, FHW^{+}01, GGL03, JC07, KCC^{+}00, KF05a, KPF03, LPC^{+}00, LZS06, LZS08, MP01a, MRS04, Mey02, MKJ^{+}05, MHK^{+}05, Müs02a, OTY02, OMF02, OMF03, SKNV01, SSP08a, UTKF05, VAMVR08, VPP^{+}12, WL00, WC00, WCBN05, Zim02]. Aliasing [Ver04]. ALICE [Ano01a]. aligned [EMJH03b]. Alignment [LVH07, AAM^{+}01, AP05]. alkali [KB02]. alkane [RJCH00]. alkane/hydroxylated [RJCH00]. alkyl [SPV07]. all-to-all [FJC^{+}05]. allosteric [LOY07]. allotropes [AP09]. alloyed [SG05]. alloys [Bur02, SSH01]. AlN [QASF^{+}05]. alone [DGR09]. along [SGF03]. also [Var02]. ALTDSE [GNZ^{+}09]. Alternating [XZ12]. aluminum [LC08b, RJCH00]. ambipolar [WTH^{+}04]. AMBRE [GKR07]. amorphous [BH03, CCRA05, LM02b, MKB02, NI01, The05]. amphiphilic [LMS05, LNC^{+}03, NT05]. amplification [HS07, HW09, KK05]. amplified [SJHY07]. amplitude [Cha07, KLTH04, KS01]. amplitudes [BBJS09, FIJ^{+}03, Hah01, KP00]. AMR [CDQF07]. AMRA [PM01]. AMS [SM01]. analog [AP04]. analogues [IN02]. analyses [IH01]. Analysing [KS04b]. Analysis [CL03, Dom05, RDSS01a, SZ00b, ATIO06, ASJ^{+}03, AdlT03, Ano01n, Ano09t, BFMH^{+}01, BLCR05, BS06b, BBD00, BFB^{+}08, BDF^{+}08, CKK09, CGG00, CGC^{+}09, Che05, CD09a, Gre07, GME06, HDGM07, HC08, HCO01, HF06, ISSC01, JKCGJ08, JC01, KKS04, KNU00, KJ07, Krö05, LKKK07, LZ00, Liu07a, LRR^{+}09, ML03, MNYY00a, MNYY00b, MSS^{+}09, Mel01, MD00, MIM^{+}07, MKJ^{+}05, MM09, OS00b, dlRBPL09, PRSB08, Ram10, RTS01, RPY07, RSD01, RF04, SKNV01, SPS09, SR01b, SKH02b, Suz00, TCY^{+}08, TBZ12, TdRGD09, TK09, TY01, TBR07, Van05c, XC03, Yan09, ZBB^{+}06]. Analytic [Blü00, Di 02, Gut06, Ste02, BDH^{+}05, Cza06, Dra01, GSS06, KVR^{+}00, SPS09]. Analytical [Don02, PNH00, AAC^{+}06, BBC^{+}01b, BS08, DK05, Dol01, GME06, HG02b, LJ09a, RE09, WW06]. analytically [OMF03]. analyzer [SG04b, TF00]. analyzing [RC04, YH02]. anchoring [LS02]. and/or [IW01]. Anderson [CMRS02, YH02]. aneurysms [OCS^{+}08]. angle [CGG00, GW01b, ZDKG05]. Angular [Dun05, Yos07, BS04a, DK05, FIBT01, FIT03, Fri09, GFF01, GFG01, GF02a, GF02b, GSF05, GWK09, HCO01, ID09, IFF01, PFG06a, Ste02, SFSL09, ZF00]. anharmonic [TS06, dAK01]. animals [HNG05]. ANIS [GK05]. Anisotropic [HP02, dAK01, CHM^{+}09, DDD^{+}01, GMBC08, LWY01, MA09, Ots01, SF05, TBL02, Wei02a]. anisotropy [CM06, FER^{+}07b]. annealing [BH03, CEM08, FHF00, Sch06a, TL06a]. annihilation [AAC^{+}06, BBC^{+}01b, WCBN05]. Announcement [Ano01-39, Ano07-29, Ohl04, vH07]. anodic [LC08b]. anomalous [BKKS09, JKW00, JKW06, PSW00]. antenna [CLL^{+}07, LCS07, PHKL02]. antennas [Bla00]. antiferromagnet [BM06, TBL02]. antigen [Dom05]. antigen-chip [Dom05]. any [CAF^{+}03]. any-flavor [CAF^{+}03]. APACIC [KSS06, KKIS01]. APE [Ano02a, ON08, Tri01]. appearance [NY08]. Appell [CGM01, CG04]. Application [ASJ^{+}03, BD08, BJ05b, CEM08, DDEM00, EVL00, FMD07, KITK00, LFT01, NW02a, NRR01, OTY02, PKPV02, RDFF02, RMLB01, SAG^{+}02, Str00, VPK^{+}01, AC07, AAKL07, Bal07, BNO^{+}01, BFLW07, BS04b, CPS00, CSZ^{+}07, DGR09, Dup01, Dys02, HCH^{+}06, ID09, KAB^{+}00, LS09, MT01, MRF^{+}05, MS08b, NN06, OMF03, QP05, RLV^{+}08, RY00, SLC09, SEC04b, SN07, Sho04, SGL09, UTO09, Vég04, WV05, ZS07, VAH04, ZC09]. Applications [JJHvO03, LV08, Nil07a, RM05a, BP08a, BBB^{+}01, Bru04, CDH^{+}06, GC01, GLHW01, HMY^{+}02, Ida03a, Ida03b, KKKC07, KS04b, LYL07, LdG^{+}07, LHS^{+}06, MRF^{+}05, OS00b, Ram10, San00, Sta02, Suc02, WHL^{+}07, Yan03c, ZC09, KPD06]. applied [Bes02, CRS01, IF03, MKJ^{+}05, NP01b, SS02b, Tol02, TIM07, TIM08, Wal03, WGY01, WJW09]. Applying [Iwa01]. Approach [ST09, Ano09a, Bow02, CRS05, CCBL02, CGA^{+}07, CGVA08, CGVA09a, CD04, CB05, DC05a, FD03, FGF03, GBA01, HKK^{+}01, Hon04, HCIK00, Huk02, IKO00, IN02, KKKC07, Ker02, KSHP02, KM03, KM08b, LJ01, LH03, LJ09a, MC03, Man02, MT01, MGG05, MM01, NM03, Niu00, OLS^{+}01, PLL07, Pis00, PFPB^{+}09, RTS01, RM05b, RLV^{+}08, RGD^{+}01, SS09a, SG00a, SM04, SM06a, SBM09b, Sch06b, SJDC07, Ska05, Teh01, TKSR00, THC^{+}07, VF03b, VKM^{+}05, Wil02, WW06, Yak01, Yok09, YT01b, YG09, ZE00, ZPB09, SBM09a]. approaches [ABSM04, BDK^{+}06, IHAR09, PKKM02, PJSK08]. approximants [FH04, RB05]. approximate [WW06]. approximating [HKP02]. Approximation [AA08, Sch05, AT09, CSCK08, EKW09, FH04, Frü03, GSM^{+}03, Inu07, MSB09, OvSA02, OIKN02, PMG07, PCC01, Pom06, Ram10, Rob01, Roy09, SWY01, Tör00, Vak00]. approximations [DCJ07, SK08, SSA07, XD08]. APW [TKN^{+}08]. aquifer [Alf05]. ARANEA [MCLDP01]. Arbitrarily [SW09]. Arbitrary [SLMS06, BD00, BS00a, CJK09, Esi01, FKAM05, IW01, KS05, Kos05, LM02a, LL00, MK09, OKS04, XSC09, Zak06, vH06, vH07]. arbitrary-order [vH06, vH07]. arbitrary-precision [KS05]. architecture [BBB^{+}01, EFS^{+}08, EL04, GMO03, ISSC01, Oli01, PKB^{+}01, SvAS01, SIE04]. architecture/circuit [Oli01]. architectures [REAB09, TG00, VHLP09, vDGM^{+}09]. archive [FFS01]. arclength [KBV09]. area [BDH^{+}05, CLL^{+}07, SM01]. argon [HKLY07, MOC03]. argon-calculation [MOC03]. ARGON.f90 [BOPC05]. argument [TB87, Tho04a]. arguments [TB85, Tho04b]. arising [HL08b, HZGZ09, Mam08, Moh08, Ram05, SPS09]. ARKN [WYL09, WYX09]. Arnoldi [GNZ^{+}09]. aromatic [EYJ07]. array [RG04, ZA01]. articles [ME00]. artificial [LVLS02, Lik01]. ARVO [BDH^{+}05]. ASDEX [SBCZ08]. ASEP [GSM^{+}03]. ASEP/MD [GSM^{+}03]. ASOS [OPB^{+}09]. aspect [CFJ09, QG04]. aspect-ratios [CFJ09]. aspects [HL00b, JBS08]. ASPIN [LDBG08]. assembled [BNSY02]. assembler [Boy09]. assemblies [KMD^{+}02]. assembly [NT05]. assessment [SS00]. assimilation [LFT01, WLH00]. assisted [AC05a]. associated [Blü04, GBM02, KFI^{+}01]. Astrocomp [BLCR05, CBM^{+}05]. Astrocomp-grid [BLCR05]. astronomical [DDEM00, ME00]. astronomy [Ort00]. astrophysical [Mül05, SV01, VKPB09]. astrophysics [Han00, Huj05, MER^{+}00, PM01, SPS09]. astrophysics-the [Huj05]. Asymmetric [Kim07, VAH04, JC01, KJ07]. Asymmetric-dimer-symmetric-dimer [Kim07]. asymmetry [Mah08b]. ASYMPT [APV00]. asymptotics [APV00, PDA06]. asynchronous [Uhl03]. Asyrot [JC01]. AsyrotWin [JC01]. atactic [LM02b]. ATLAS [McK07]. atmosphere [LC01b]. Atmospheric [CBMS08, AdlT03, DSH02, DSH03, DSHH05, HJZL07, LHS^{+}09]. atmospheric-pressure [LHS^{+}09]. atom [Bac00, BK06a, BGH^{+}09b, CWSH08, CGG^{+}08, CGG^{+}09, Gol00, GG03, HSGBK08, Hin00, HTM01, LJ01, LDBG08, MGG05, Rap06, SNBB02, VKN07, Vie01]. atom-centered [BGH^{+}09b, HTM01]. atom-molecule [Hin00, LDBG08]. atom/ [MOC03]. ATOM/NEON [OMC00]. Atomic [HYY07, ON08, Pit05, ZF09, ASH06, BD08, BK01, BKM02, BKM05, BNSY02, Dzu09, EÅ01, FSK04, FBL00, FA00, Fro00, FTGG07, GZF04, GFG^{+}06, HGVCM^{+}02, HPC05, Hor09, Iwa01, Jia08, JHFG07, KM00b, Kon02, LHMB00, MW01, MTJ02, OS03, PFG06b, RS03, SV01, SKF05, TBR07, Wes07, Zat06]. Atomic-level [HYY07]. atomic-structure [Fro00, FTGG07]. Atomistic [KNSY07b, CCRA05, GLHW01, HOT07, Hef00, MCC05, MRF^{+}05, Müs02b, Nak07, SKNV04, Tod01, ULA^{+}02, KNSY07a]. atompaw [HTM01]. atoms [ARV02, BCP04, BM04, EÅU05, EKW09, HHM^{+}09, MOC03, Nik03, OMC00, SJP05, TNCG00, Tsa02]. attempt [MM01]. attentive [Ano01n]. attractive [NHS07, Sea02a]. attractor [ZWY04]. attractors [Lew04]. Augmented [HTM01, THM01, ADS06, BVKW02]. auralize [TKS^{+}01]. Author [Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02b, Ano02c, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano05a, Ano05b, Ano05c, Ano05d, Ano05e, Ano05f, Ano05g, Ano05h, Ano05i, Ano06a]. Authors [Ano00z]. auto [Str01a, LL08]. auto-solitary [Str01a]. auto deriv [SPF00]. autocorrelation [MSS^{+}07]. autoimmune [Dom05]. automata [BBD00, Sta00, TCO00, VK09b]. Automated [HvHHM09, VEG08, PB09b]. Automatic [Ano01n, De 02, FIJ^{+}03, Kau03, LJ08, MP01a, SMK01, Kol09, Pap01, RMMP02, Sem09, SPF00, Str05, vH06, vH07]. Automation [MA00]. Automatized [Cza06]. automaton [HS01a, KKM02, RDS02a, RDS02b]. Automodelling [DKV00]. Auxiliary [MR05, Bae03, Bae04, ZKASS05]. AV77 [CDFF05]. Avalanches [FV02]. averaged [CBKM01]. avoiding [Jen01]. Award [Ano04a, Ano04c]. Axially [SDNR05]. axis [LVH07]. Axisymmetric [SM06b, BFI^{+}00, dNKM07]. Azurin [DC05a].

B [TAP01, AAM^{+}01, AC09, FZ09, Nik03, TD03, Zat06]. B-spline [FZ09, TD03, Zat06]. B-splines [AC09, Nik03]. BaBar [ADD^{+}03, Teh01]. background [Con04, JPS^{+}01b]. backlight [CFJ09]. Bäcklund [LL08]. backscattering [WSB04]. backsubstitution [SG06]. backward [HS07, HW09, SGM^{+}09]. BAGEL [Boy09]. Baker [WS09b]. balance [BD08, CD09a, ZSdD^{+}08]. balancing [PSP^{+}03]. Ballistic [RdAGV^{+}00]. balloon [AdlT03]. ballooning [SHW01]. band [GRS06, HKK^{+}01, JC01, MS06, Nak08, PL05, PDA06, PAD07, SHX02, SYM00, WP00]. band-structure [MS06, SYM00]. bank [AL08b]. bar [CDQF07]. bar-mode [CDQF07]. Barnes [Ada04, Cza06, GKR07]. barrier [DLZ08, LLY07]. barriers [TKP06]. baryons [CWW07]. Base [CCFG05]. based [ABC^{+}03, BCC^{+}08, BDBV12, Bur02, CGIA07, CSZ^{+}07, CCRA05, DDM05, DBE^{+}04, DGR09, DHBE05, FBB01, FK00, FKP03, GMAN^{+}07, GKK^{+}08, HSJ02, IH09, ISS^{+}02, JP09, KKKC07, KH09, KSTL03, KKHL07, LM02a, LRI^{+}06, LJY07, LNV^{+}09, LZC^{+}08, LKC06, LCV06, LF02b, LM00, MY00b, Man04, MLG^{+}01, MLF07, MM05, RPD^{+}05, RGR^{+}04, ISX05, SKNV01, SKNV05, SJHY07, TS06, TYS^{+}00, UK02a, UK02b, VF03a, VEG08, VPP^{+}12, WL00, WCGL00, WDB04, WH00, Yok09, Zie04, dMBC^{+}06, dlRL09]. Baseline [HKL^{+}07a, HKL^{+}07b, HLW05]. bases [BRD04]. basis [BP08a, BC00, DD00, DO04, DO05, DSC^{+}09, FA00, GFG^{+}06, Hua09, KTT09, KTL05, MBR01, MN01, MAM04, MAM07, MSHP02, MSHP20, Pit05, RB08, SDNR05, Suz00, THM01, TS06, TKN^{+}08, UYK^{+}04, You09]. basis-set [TS06]. BAT [CKK09]. Batch [BFL^{+}01]. Bauer [OML09]. Bayesian [CKK09]. bcc [YKK07, TBL02]. BCS [BFH05, RGD^{+}01]. BCVEGPY [CDEW04, CWW06a]. BCVEGPY2.0 [CWW06a, CWW06b]. BDF [IHAR09, IVD03, VAMVR08]. be [MMMM00, VBFM05]. Beam [Bre07, PBB^{+}04, CP00, OSK04, OKS04, PPP01, QRH00, QR01, QG04, QTL06, Sch08, SBBM04, TAM04, WCG04, YRR07]. Beam-plasma [Bre07]. beams [AT09, Bar00, LdG^{+}07, Mah08a, SFF^{+}04]. bearing [KMB02]. beats [KB02]. BEEM [RdAGV^{+}00]. behavior [DR09, GDAG05a, GDAG05b, HTM^{+}08, HOI04, LWT08, Lud02, MCH02, Sat02, SKRK04, TBL02, YGT^{+}02]. behaviors [LDZ^{+}08]. behaviour [Bal01, LNLK01]. Behaviours [RDSS01b]. BEM [BP08a]. BEM/FEM/GSM [BP08a]. benchmarking [Gre04]. benchmarking-how [Gre04]. benchmarks [BMvG00, FHW^{+}01]. Beowulf [ABC^{+}03, Ano03h]. Beowulf-based [ABC^{+}03]. Beowulf-class [Ano03h]. Berlin [Hoo04, Laf03, Par04, Sha04, Vio04]. Berne [IW01]. Bessel [Tho04a, CP00, Tal09, TB87, VC08, YM03]. Best [Ano04a, Dem03, Sal02]. Bethe [Frü03]. Better [FKP03, PB09a]. between [AC05a, AC05b, Blü04, CW02, HKLY07, KTL05, KMB02, LCPC04, LJ01, MSD08, Mül02, PSK01b, RCG05, Ver00, ZSSA00]. Beyond [Gre04, SWY01, GG00]. BGK [AGJJ07, BBD00, KSC^{+}00]. Bhabha [TI01]. bi [LCHJ09, Xia01]. bi- [LCHJ09]. Bi-CGSTAB [Xia01]. bias [OD08]. big [KACB07]. BIGEBRA [AF05]. bilayers [LMS05, SDLW07]. bilinear [Ram10]. billiard [BFB^{+}08]. billiards [ISSB01]. billion [ZBB^{+}06]. billion-vertex [ZBB^{+}06]. binaries [CSS^{+}03]. Binary [dMBC^{+}06, BLCR05, GAR05, MY00a, MY00b, Mül02, Pur02, TE05, WS02]. binding [CR00, ÖDÇ02, Ver00]. bio [BMS^{+}09, DC05a]. bio-fluidic [BMS^{+}09]. bio-molecular [DC05a]. biographical [Kar01]. Biokinetical [GMAN^{+}07]. Bioler [SIE04]. Bioler-2 [SIE04]. biological [PFPB^{+}09, SIE04, YC07, ZDKG05]. biology [Rin02]. BIOMCSIM [KH01]. biomedical [LHS^{+}09]. biomembrane [Bro07]. biomembranes [LMS^{+}02]. Biomolecular [SG04a, De 07, KH01, MFVJ07, MS09]. biomolecules [SLBG09]. biophysical [FMD07, MDC09]. biosensor [RR02]. bipartite [BCHP09]. bipolar [LH03]. biquadratic [DKC08]. Birdsall [Ano04-57]. Birkhoff [CRUV00, She08]. Bisection [VPK^{+}01]. bistable [MTZ00]. Bit [WHO02, DH00, JC01]. Bit-parallel [WHO02]. black [BFI^{+}00, CGCS07, HBRS05, Leh00, RRRHD08]. blast [PPC07]. blends [BMML05]. Block [CHS09, EM08, HZGZ09, Eas08]. Block-P [Eas08]. blocked [Cha00]. BlueGene [CD09a]. Blume [DKC08]. board [TIN^{+}09, Ano05-40, Ano05-41, Ano05-42, Ano05-43, Ano05-44, Ano05-45, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano12]. body [AMP^{+}00, ADBF03, BAD01, BD08, CW00, DD01, Dzu09, EE02, GPW^{+}09, KNU00, KM05, KN07a, LZS06, LEG02, Mak01, NP01a, OMF02, Sav01, SvAS01, TND04, TND05, Var08, VPNW02, VT00c, YB02b]. Bogoliubov [MM04]. Bogoliubov-de [MM04]. Bogolyubov [BD05, DO04, DO05, DSC^{+}09, SDNR05]. Bohm [DDM07]. BOKASUN [CCGR09]. Boltzmann [Wal03, BBD00, BS00b, CT00, CK08, DPB01, DHB^{+}04, DSL09, Dys02, HL00a, HCO01, IK00, IKO00, KITK00, LNC^{+}03, Luo00, MHR^{+}07, MC08, MS05a, OCS^{+}08, PPM04, PY08, ST02, Suc02, TMTF00, TCF00, TKSR00, TdFK00, ZHC00, ZY09, vdSvdG08]. Boltzmann-like [Wal03]. BoltzTraP [MS06]. bond [BCBJ02, CYAS05, HJ02, NLC09, OK06b, ZBB^{+}06]. bond-diluted [BCBJ02]. bond-site [NLC09]. bonded [Bac02]. Bondi [Rib02]. Book [Ano00a, Bre01, Hoo04, Koc02, Laf03, Par04, Sha04, Vio04, Wan00]. Born [BS03, CCD07, OIKN02]. Bose [BGJ^{+}07, BBR03, CPS00, CC07, CCL08, CC09, LR07, Nil07a, SVS01, TQ03, TS06, ZZ09]. boson [HHH^{+}09]. bosons [ABB^{+}09, DC05b, HHW00, RM05b]. both [AA08, CL08a, Yao09]. bottom [TSA^{+}03]. botulinum [CCD07]. bounce [CBKM01]. bounce-averaged [CBKM01]. Bound [MT00, AMP^{+}00, Bac02, DSH03, GLMADB^{+}02, GPT08, LEG02, OvSA02]. bound-constraint [DSH03]. bound-state [AMP^{+}00]. boundaries [MK09, UOM01, Ver04]. Boundary [GPT08, AA08, CRS05, CLR08, CLFH07, CFKM01, CY01, CS02, CHM^{+}09, DM09, EH07, Kar02, KEM^{+}01, KSSH04, KT07, LJ09a, LC00, LÅT04, Liu07a, MK09, MNV00, MPS09, PPC07, Ram04, Ras09, Ras17, Sus01, WGL06]. boundary-layer [Ras09, Ras17]. boundary-value [LC00, Ram04]. bounds [MA06, SMZ05, ZSM05, vdEFL^{+}02]. Boussinesq [YB02a]. BOUT [DUX^{+}09, UXD^{+}09]. box [NH09]. boxes [KN07b]. bracket [KHÖ01]. brackets [UTKF05]. Bragg [MTLC01]. branched [JAT03]. BRANECODE [MFF^{+}05]. braneworld [MFF^{+}05]. breakdown [LSL07, NV09]. breaker [PPC07]. Breaking [DLZ08, ABD^{+}05, BKB02a, ISSB01]. breakup [vdSvdG08]. Breit [ZF00]. bremsstrahlung [Frü03]. Brenner [LF02b]. brief [Mar08, SMS08, Ver00]. Brillouin [Zah04, Zah05]. broad [dO02]. broadening [WCBN05]. broken [Sle00]. Brownian [DHB^{+}04, RvOvV02, SS02b, WLR^{+}08]. browsing [BBB^{+}00]. BRST [PTL04]. Bruce [Ano04-56]. brushes [LS02]. BSR [Zat06, ZF09]. BtoVVana [BS06b]. bubble [NN06, WGS00]. buffered [GS01a]. builder [ACC^{+}01, ABF^{+}01]. Building [HS01b, BBB^{+}01, FKMB09, SG04a]. built [LV08]. bulk [Mam08, Moh08]. Buneman [Ano04c]. Burgers [RE09, Zak00b]. bursting [TRGR08].

C [Koc02, KSS06, Ano09t, BCCW03, CJT06, Di 01, DH01, DM07, KS05, KKIS01, Mal00, NT05, Nat08, OGWH03, RPD^{+}05, SWS^{+}12, Tol02, iTKST01, WR01, Bre01]. C-code [Nat08]. CA3D [GS01a]. Cadabra [Pee07]. CADNA [JC08a, SJDC07]. Cahn [KEM^{+}01]. calcium [KACB07]. calcium- [KACB07]. calculate [Alf09, AC05b, BBPS07a, BBPS07b, BGH^{+}09a, CCGR09, GFG01, GLHW01, GG03, KALC08, MM01, NN09, Sar00, SV01]. calculated [MMMM00]. calculates [Hor09]. Calculating [EMJH03a, APV00, All02, BBPS02, CP00, Cip07, Cip08, Cip09, Dzu09, EKW09, Gro01, KW08, LZS06, LC01b, MS06, Mah08b, Mah09a, MFVJ07, Nat08, NY06, Por03, SHX02, SFR05, Tal09, WP00]. Calculation [Alf05, Bek06, FM03, GF01, GMBC08, Jia08, Kir06, LVLS02, LS01, OMC00, ST02, UVLRRC09, Vul03, Zah00, Zah01, AC07, AL08a, Bar03, BKM02, CCBL02, CWSH08, ÇHM00, CD04, Col07, CL08b, EVL00, FK00, FSB09, GSM^{+}03, GMAN^{+}07, Goc04, GDAG05a, GDAG05b, HHH^{+}09, HSGBK08, HHW00, IBA00, KK00, LANM^{+}01, MDS09, Man04, MR06, MWA01, MOC03, Nik03, PCCD09, Por00, Ram10, RdAGV^{+}00, SLC09, SJP05, Sea02b, Sol01, TKB^{+}04, TNCG00, TKN^{+}08, VF03a, VF03b, VS06, VT00a, Vos06, Yan09, ZSdD^{+}08, ZS08, ZDKG05]. calculations [AIOST03, AJT^{+}07, Bac00, BTI01, BH01, BK01, BKM05, BBB^{+}09b, BMG01, BD06, CN01, CD01a, CC08, Con04, Dan05a, Dan05b, DS06, Dan07, DTD^{+}02, Elm09, Fer07a, FFD00, FFG02, FTGG07, FKG00, GZF04, GIME02, GHP01, GGG01, GBTM07, GBD03, Gol00, GW01b, GRS06, HC00, HHM^{+}09, HLC08, HPC05, HTM01, IM01, IBM03, JRT00, KDW00, KLD04, Kon02, LCB^{+}00, LOCJ05, LLV^{+}01, LEG02, LR06, LZ04, Mah09b, MC03, MHGV09, MSB09, MBR01, Mei01, MN01, MAM04, MAM07, MSHP02, MSHP20, OD08, OBG09, PFG06b, PWH^{+}00, Pog05, Pue06, QASF^{+}05, RPD^{+}05, RGD^{+}01, SHW01, SIH^{+}01, SNS01, SBM02, SKNV01, SN07, SGL09, SMH^{+}01, SJ02, SR01b, SVMT00, SYM00, SNBB02, THM01, TAKN02, TND04, TND05, TYS^{+}00, VKM^{+}05, VT00b, WKP^{+}01, Wil09]. calculations [ZF09, Zha00, dSdSW08]. calculator [Bar02, DK05]. calculators [Ste02]. calibration [AAM^{+}01, HTNFBS06a, HTNFBS06b, RTS01, TNBSF04]. Campbell [WS09b]. can [BKB02b, Gre04, MMMM00]. cancer [Dom05, TdRGD09]. Candia [CCG08]. CANM [AP04]. Canon [MP04]. canonical [Bae04, FdO09, JBS08, KCH00, PRSB08, Zim05]. canonicalization [MG08b]. capabilities [BNO^{+}01]. capacitive [TC07]. capacitively [KPL07, KCR07]. captions [Ano09r]. capture [Ber03b, Car06]. capture-gamma [Car06]. capturing [Wei02a]. Car [CCFG05]. Carbon [HKK02a, AP09, CSC^{+}07, CSC^{+}08, HKK02b, KKKC07, LC08b, LF02b, NKSL05, OPO^{+}08, ÖDÇ02, PLL07, YN05b]. carbon-based [LF02b]. Carcinogenic [EYJ07]. Carlo [FNR^{+}07, JKW06, KRW03, TA00a, WA07, AW04, ABM03, ACIZ07, ASF^{+}05, AGS07, Ano03h, ABB^{+}09, Asc08, BS06a, Bae03, Bae04, BBB^{+}09a, BJ02, Bar00, BDG^{+}08, BvG02, BR09, BL00, BMML05, BHM^{+}07, BM01, BHL02, BK05b, BDYK04, BKB02a, BKB02b, Bur02, BB03, CGCS07, Che05, CGK^{+}00, Cun09, CKA^{+}09, DS01, DDD^{+}01, DGLB08, DDRW03, DH01, FNR^{+}06, FdO09, FMN01, GS01a, GPW04, GW01a, GPW^{+}09, Gra02, GOG00, GRS06, HPC05, HKLY07, HCIK00, Huk02, JKW00, Jad00, JWW00a, JWW00b, JPS^{+}01a, JPS^{+}01b, Jad03, JS06, Jun02, JBS08, KH01, KPL07, Kat02, KL06, LTA05, LF02b, MBKJ09, MRS04, MHS05, MSS^{+}09, Maz00, MSK^{+}05, MP03, MMB02, MB05a, MP06, MG09a, MABK02, MER^{+}00, MKM02, Nat08, Nil07b, OTY02, OPO^{+}08, PMA^{+}04, PSW00, Pop03, RP02, RIB01, RPD^{+}05, RS00]. Carlo [RK05, Sch04, SVP09, SLWH02, SVMT00, SSLN02, TA00b, Tak00, Tom09, TNCG00, Trö08, ULA^{+}02, Uhl03, VYK02, VPNW02, VMMB02, Wal03, WL00, WJW09, WK02, WH00, WLGX09, YC07, dS03]. carlomat [Kol09]. Carlos [Sul05]. carpets [SFSH01]. Carson [Don02]. Cartesian [CMT00, DD00, DO04, DO05, DSC^{+}09, MAM04, MAM07, SFSL09, WPL02, WD04]. Cascade [KSS06, KKIS01, BAB04, Jun02]. cascadic [DB08]. case [AIOST03, ADE^{+}02, CM02a, FGV01, HM06a, PR06, SK08, SS09b]. case-study [ADE^{+}02]. CaSPA [MDS09]. Catfish [CGCS07]. cathode [KHIL07]. Cavitation [BNS07]. CCFM [Jun02]. CCP [Ano05j]. CCP2006 [Ano07a]. CCP2007 [Ano08a]. CCP2008 [PB09a]. CCP3 [GI01]. CCP99 [BDL00]. CCSD [PKKM02]. CDF [ABF^{+}01, ABC^{+}01, BGLLW01, Sfi07]. CDiffElim [WR01]. celestial [OMF03]. Cell [De 07, MDT03, SHT08, BDYK04, CGIA07, CH09, Cle05, Dec07, EL04, FS08, HKLY07, JH09a, KSYE00, KLD04, LM02a, LC01a, PSP^{+}03, Poi08, Poi09, ISX05, SLC09, SLL01, iTKST01, UOTM03, VPCK04, YT01a, YWLC04, ZLM04, Esi01, TCY^{+}08, TDD04, VAH04, VCF^{+}04]. cell-based [CGIA07]. Cell/Monte [WJW09, KPL07]. cells [ATB^{+}01, Dom05, HFN03, TdRGD09, Zah00, Zah01]. Cellular [CGIA07, RDS02a, CD08, Jad03, JS07, KKM02, RDS02b, VK09b]. Center [SBM09a, GME06, SG00a, SM04, SM06a, Org01]. centered [BGH^{+}09b, HTM01, KBG00, MP05]. central [GZ07, ID09, KG07, KF05b, Ska05, Zie05]. central-constraint [Zie05]. central-field [KF05b]. central-upwind [GZ07, ID09]. centre [PAT^{+}09, San00]. centroids [QTMH07]. century [Nov02, Swe02, You02]. ceramics [VKN07]. certainty [PKPV02]. CESD99 [FA00]. CF [KPL07]. CFD [OS00a]. CFT [JW02]. CG [RMMP02]. CG-OPT [RMMP02]. CGSTAB [Xia01]. CH [LJ09b]. chain [FS01b, FS02, NFS01a, NFS02, NL07, PRSB08, Ryc05]. chain-molecule [NFS01a, NFS02]. Challenge [BNO^{+}01, You02]. challenges [Ano01a, EL04, Mül05]. challenging [WL08]. chamber [HFN03]. chambers [GKM^{+}00]. changes [RP02]. changing [KMR^{+}09]. channel [CGA^{+}07, CGVA08, CGVA09a, Ixa02, LVV06, Riz02]. channels [BEM^{+}02, ISS^{+}02, KACB07, MS05a, vE08]. Chaos [BFB^{+}08, Par04, LMC^{+}03, Yur02]. Chaotic [KB02, ZWY04, AOT01, ÁMRP04, ISSB01, LCPC04, LMC^{+}03, MSD08, TYSH05, kWpLwW01, WLW04]. characterisation [PB09b]. Characteristic [VPK^{+}01, Leh00]. Characteristics [KKH07, Gha05, GRS06, HKK^{+}01, SGF03]. Characterization [Lew04, TBR07]. Characterizing [GLW03, BCHP09]. Charge [Ger07, MHK02, PP02, BSTC05, BCH05, Esi01, Har02, LHC01, LHC02, MPR05, PSK01a, PSK01b, Poi08, Poi09, SGF04, UOTM03, XON08]. charged [BVKW02, DHB^{+}04, KNY05, QR01, QG04, RR02, WDF^{+}02]. charges [MP05]. Charles [Ano04-57]. Charlotte [Bur01, Hib01]. Chaste [PFPB^{+}09]. Chebyshev [ABOSPG09, Del08, Hua09, Str00, WP06]. check [SJDC07]. Chem [FKG00]. chemical [ABNÅ05, BFLW07, CRPC08, CRS01, CW01, DVL^{+}02, DVL^{+}04, HYY07, ISS^{+}02, Pin01, PJSK08, SK05, SN07, SVMT00, VEG08, WSB04, ZBB^{+}06]. chemical-physics [CRS01]. chemistry [KAB^{+}00]. chemometric [IN02]. CHEREN2 [CM06]. Cherenkov [CM06]. CHIMERA [Ano01n]. Chip [EFG^{+}00, Dom05]. chiral [CAW00, GRR01, Jan05, KSS02, KALC08]. Choice [Tak00]. cholesterics [OMY05]. Chombo [MGN07]. chromatin [LOY07]. Chromodynamics [SHT08, ABD^{+}05]. CI [ABER00, Nik03]. CIP [MS08a, MYJY01, UK02a, UK02b, UYK^{+}04]. circuit [LH03, Oli01, PPC07, Pet04, VK09a]. circuits [Bor02]. circular [DM09, VYK02]. circularly [NY07]. circulation [SVS01]. Class [Bel05, Ano03h, BBJS09, CR08, KPF03, PSH06]. classes [CTR00]. Classical [Elb05, MKB02, Rap02a, ASH06, ASF^{+}05, BFB^{+}08, CL08b, DDFI09, GF02c, HSSA01, KB02, Lüt04, MFVJ07, NP01b, Ram03, SLBG09, SS06, TBL02]. Classification [KW07, AAM^{+}01, OMF03]. Clausius [LWY01]. cleaning [JH09a, MOS00, MOS01]. clear [RTVZ08]. clear-sky [RTVZ08]. ClearSpeed [TIN^{+}09]. Clebsch [BRD04, CRW09, Dra01, KW08, RF06b]. Clifford [AF05]. close [BLCR05, HSJ02, KF05a, MPR05]. closed [BKKS09, Sim08]. closure [Ker02]. cloud [Kon01, MMTH04]. clouds [NN06, Shi09]. clues [AL08b]. Cluster [BHL02, LMM^{+}08, MSS^{+}09, Mel05, Sat02, ALV05, Ano03h, BPRW06, BOG^{+}07, BJ08, CMRS02, Gut06, KPS^{+}01, KDW00, LC01a, PKKM02, RD05, RS09, Val05, Vor02, YNS^{+}09]. cluster-cluster [Vor02]. Cluster-forming [LMM^{+}08]. cluster-surface [KPS^{+}01]. CLUSTEREASY [Fel08]. Clustering [GM00, BCCW03, Got01]. clusters [ASH06, Bac02, BBOY08, BBB^{+}04, BM01, DDFI09, Fel08, GCP^{+}02, GB05, GF02c, HGVCM^{+}02, Iwa01, JBA05, Kur02, KFB01, LNLK01, MB05b, MTJ02, NW02a, Tak03, Yur02]. CMS [Org01, ACC^{+}01, GKM^{+}00]. CMS/ECAL [Org01]. CMSapi [MRF^{+}05]. CNDO [SS09a]. CNDO/2 [SS09a]. CO [Gha05, Mah08a, VMMB02]. co-propagating [Mah08a]. coalescence [NFS01a]. coalescing [CSS^{+}03]. coarse [ASS^{+}02, BLS09a, EL06, FAiTD01, LMS^{+}02, LS09, MS09]. coarse-grained [ASS^{+}02, BLS09a, EL06, FAiTD01, MS09]. coated [CAW00]. coating [SR01a]. Coatmèlec [GMAHV^{+}09]. COBRA [SHW01]. Code [Ano04-46, LC01b, OSK04, OKS04, PMA^{+}04, RGD^{+}01, ATF^{+}09, ADBF03, Bar03, BAD01, BLCR05, BCAD06, BADC07, BMS^{+}09, BG06, BB00, BN07, BAB04, BGS^{+}04, CBKM01, CGK^{+}00, CG04, DGV08, DGS08, DTD^{+}02, DPB01, DKM07, DHBE05, EH07, Elm09, EKW09, EL04, FWP01, GMAN^{+}07, GBFS07, GS01a, GBTM07, GT04, Gre07, GZ07, GSSN00, Hah08, HSGBK08, HvDJvdM01, HTM01, ICT01, JKCGJ08, JBA^{+}07, KH09, KSYE00, KLM00, KCH00, KLTH04, KSSH04, LANM^{+}01, LKPH08, LDBG08, MS06, Man04, MGG08, MC08, Maz00, MB04, Min01, MK08, MAM07, MNV00, MDM05, Nat08, NN09, PLPS08, PBI07, PCC^{+}09, PCCD09, PM01, PSP^{+}03, PPP01, RtVR09, RJFB08, Rib02, RGR^{+}04, SHW01, Sea02b, Sea02c, SLL01, SPF00, Str05, SMK01, SSB04, Swi04, THM01, TAM04, TC06, Tol02, TDD04, Ton07]. code [UXD^{+}09, UTO09, VKPB09, VT00a, WML^{+}05, WHL^{+}07, YSM09, Zie04, Zie08, vdHKM08]. coded [HCO00, ICO03, SCO00]. codeposition [NG02]. codes [AG05, CR05, Dec07, DBE^{+}04, HDG07, HHL06, Kud09, LC01a, MCL05, PSK05, RLI07, SBM^{+}04, SJDC07, SSB^{+}09, SGF04, TCY^{+}08, WJW09, Zat06]. Coding [LS09]. coefficient [LL08, Ste02, SS02b, qX09, ZLL09]. coefficients [BRD04, BKKS09, CRW09, DK05, Dev05, DJ08, Dra01, Dy09, FGA04, Fat02, FIBT01, GF01, GFG01, GF02b, GSF05, HM06a, HB05, Kas00, KW08, RF06b, Vak00, VF03a, VF03b, Van05c, WP00, Yan03d]. coexistence [FFF01, KF05a]. coherent [SJHY07]. cohesive [KBBW02, YZD^{+}07]. coil [YD06]. coincidence [MKJ^{+}05]. cold [PCV06]. Collaboration [Ano04-46, PMA^{+}04, All01]. collaborative [dSdSW08, GI01]. collapse [HBRS05, MMTH04, SBD^{+}06]. collection [vDGM^{+}09]. Collective [AK03, LV08, BA09, YG09]. collider [BDW06]. colliders [ABM03, BBB^{+}09a, DDRW03, Kol03, Por03]. Collision [PM00, ZBB^{+}06, CL08b, FBL00, RFK08, WRMG05]. Collision-free [ZBB^{+}06]. collisional [HD04, HvDJvdM01, KA04, MV04, ST02]. collisional-radiative [HD04]. collisionally [LHMB00]. collisionless [GBC^{+}04, JBA^{+}07]. collisions [Abe01, BF04, BPP01, Che05, GG03, HSGBK08, JWW00b, Tom09, TSA^{+}03, TKK^{+}06, WM00]. collocation [LFT01, LFT03]. colloidal [All05, CMS04, KNY05, MHK02, SBD^{+}05, itVPG08, YNK05]. colloids [DHB^{+}04, FHR^{+}05, SF05]. color [AEEdR05]. colored [Gen01]. Columbus [Pit05]. Combinatorial [Flo01, DLZ08, JS08, Zim05]. combined [ASJ^{+}03, FSK04]. Combining [CL08a, DGLB08, GSM^{+}03, Hin00, SR01b]. combustion [ZLM04]. Comm [DVL^{+}04, LPR04, MSHP20, Ras17, TIM08, WA07, Yos07]. Comment [AA01b, Hon04, LHC02, Ixa01, Mar08, Ram10, WLW04]. Comments [Har02, Moh08, MA08]. Committees [Ano05j, Ano07a, Ano08a]. Common [KSS02, TBR07]. Commun [AA01b, AAB^{+}07, CSC^{+}08, CGG^{+}09, CGVA09a, Hon04, Ida03a, Ixa01, JKW06, KM01b, KS08, Nat10, Poi09, Tho04a, Tho04b, TND05, Voi03]. Communication [BFL^{+}01, TA00a, CD09a, GDC01, MP01a, SOAW08]. Communications [Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03i, Ano03j, Ano03k, Bro00, FNR^{+}07, Fij00, GDAG05a, MOS01, Ram10, SM06a, Wu10, GCD06, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b]. Communications [Ano08c, Ano09c, Ano09d, Ano09b]. community [KOS^{+}09, MOM^{+}00]. Comp [Ida03a]. compact [JS07, Jen01, SIH^{+}01]. compacting [KBBW02]. compaction [RLH^{+}09]. compacton [YB02a, Yan03a]. comparable [DCJ07]. Comparative [FHW^{+}01, BCV03]. Comparison [FS03, HKLY07, LJY07, SG06, Van05a, YZW02, BP08b, KALC08, RE09, RLI07, TBZ12, Ver00, PSK01b]. comparisons [GPW04]. Compartment [GMAN^{+}07]. COMPASS [Mar01, TLDM03]. Complete [AC05a, CK08, Zim02]. completely [JP09]. completeness [AC09]. completion [SHV^{+}01]. Complex [CIC^{+}03, KD09, LLH07, NM03, Ber02, BKM02, CDFF05, DPB01, GSS06, KM01a, LBM05, MCH02, MC08, MPK00, Mic07, MS08b, MPS09, NN09, Poi08, Poi09, RDSS01a, SK08, SHZ01, TT06, TB85, TB87, Tho04a, Tho04b, Tod01, WKP^{+}01, WRMG05, WLGX09]. Complex-scaled [NM03]. Complexity [MBC^{+}09, SSA07]. complicated [NP00]. Component [LM00, JKCGJ08, TdFK00]. components [Tót06]. composed [GBD03, HSS^{+}08]. composite [CL03, GMBC08, PKPV02]. Composition [KFB01]. compound [BAB04]. Comprehensive [SBM^{+}04, TJLR06]. compressible [Ida00, Ida03a, Ida03b, LTA05, TIM07, TIM08, dNKM07]. compression [MM05, OCK^{+}00, Pet04]. Comput [AA01b, AAB^{+}07, CSC^{+}08, CGG^{+}09, CGVA09a, DVL^{+}04, Hon04, Ixa01, JKW06, KS08, LPR04, MSHP20, Nat10, Poi09, Ras17, Tho04a, Tho04b, TND05, TIM08, Voi03, WA07, Yos07]. Computation [AS00, BMC05, GFS03, KMS09, SKH02a, WRN01, dDSFY04, BD00, BGH04, BS00a, BDT00, CNMC09, Che07, CA09, DB08, FD03, FL01, Gal00, GT01, Hon04, Ida02, Ixa07a, Kol09, KTL05, KH06, LVV09, LTG09, LJ08, MS08b, NJ00, Pap01, PPC07, PTL04, RTVZ08, SBM09b, She08, SI01, Ste01, TF04, UTKF05, VK09a, qXbL04, qX08, Yan02, Yep02, dSdSW08, dlGGS^{+}05]. Computational [Att09, BDL00, DC05a, Gou00, GI01, HKK02b, KB04, LCB^{+}00, Lan07, Mel01, MRF^{+}05, MS05b, MB05b, Nov02, OLX07, PRBD09, SWS^{+}12, SHJ07, Swe02, TdRGD09, Ano09s, BLM01, dSB00, Bor02, Bor07, Bra05, CZC00, CRS01, CMT00, CMT01, CSZ^{+}07, FS00, GGL^{+}02, GLL^{+}02, Gun02, KAB^{+}00, KB02, LNK01, LPC^{+}04, LCE^{+}09, MSK^{+}02, Min01, NP01a, OBG09, RM05a, Rin02, SG00a, SM04, SM06a, SBM09a, SI01, SAG^{+}02, Suz00, TCY^{+}08, WG01, WM00, You02, Zie08, Hoo04]. computational-task [Ano09s]. Computations [Str01a, Ada04, ABNÅ05, ABD^{+}05, BBD^{+}09, Di 01, DPSG06, FIT03, FMN01, Inu07, KMZZ05, KKF^{+}04, KM08b, Liu07b, ZE00, vDGM^{+}09]. compute [BCP04, BFLW07, CG04, Dy09, HB05, KP00, Sal02, SSP08a, Ver00]. Computer [All05, AC05a, Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03i, Ano03j, Ano03k, BA09, BR09, BDV04, CAAM08, DHMD00, Elb05, FNR^{+}07, Fij00, Goc04, GH01, GDAG05a, HMY^{+}02, JDBT06, KM01b, LPRS02, LPR04, LVLS01, LMS^{+}02, MOS01, NY06, NY08, NP00, Ram10, Rob00, SM06a, SBBM04, SAU^{+}04, TA00a, Tod01, Wu10, BCC^{+}08, BDLT02, BCG03, BG06, BL05, BKKS09, BD06, BCV03, Cha07, CRS01, Cip07, Cip08, CHP04, CPT^{+}01, DS06, Dan07, DDM07, DMD^{+}07, DJ08, DSS01, FKP03, Fra07a, GS01b, Gro01, HJM02, IOM00, JP09, JDBT09, KK04, LM02b, MTLC01, MG08b, Mas00, MVS05, Müs02b, OGKL02, Pee07, PGS02, PKKM02, Pue06, RtVR09, RDSS01a, RDS02a, RDS02b, iSHS^{+}08, SPC^{+}05, SMS^{+}00, SI01, SHI02, SHH^{+}04, SIE04, TNCG00, VT00a, Wei04, Xia01, ZSSA00]. Computer [Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09c, Ano09d, Ano09b, Koc02]. computer-generated [SI01]. computerized [FD03, GT01, Hon04, LMC^{+}03]. Computers [Esq02, Ano02a, Att09, CSS^{+}03, Mak01, Mei01, MT00, OLS^{+}01, OSK^{+}02, OCK^{+}03, Ref00, TYS^{+}00, Yos01]. computers-past [Ano02a]. Computing [BSTC05, Bre01, Bro00, CC09, FPB08, GSS06, RM05a, Sha04, Shi07, Tho01, VS01, WS09b, YM03, ZSM05, ADE^{+}02, AJ08, Ano03h, BS03, BN07, BD06, BDH^{+}05, CMRS02, CD05, COE^{+}05, CC07, CCL08, CRUV00, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, CGVA09b, CBM^{+}05, Dan09a, Dan09b, Dra01, Fel08, GBM02, GCK02, HLB06, JP09, JG02, KVR^{+}00, LbotMC01, LMC^{+}03, McK07, MA04, MA08, Nil07a, PKPV02, PMV02, Sfi07, SMZ05, SS07b, TJD09, Tri05, VPK^{+}01, ZF00, ZZ09, Vio04]. concave [Dem06]. concept [BLS01]. Concepts [San00]. concerted [Nak07]. concrete [FKMB09]. concurrent [JPS^{+}01b]. Condensate [TQ03, KALC08]. condensates [BGJ^{+}07, CC07, CCL08, CC09, LR07, Nil07a, SVS01, ZZ09]. condensation [ASJ^{+}03, CPS00, KW03]. condensed [BH03, CMR01]. condensed-matter [CMR01]. condition [MPS09, WGL06, YSM09]. Conditions [TLDM03, AAP03, AA08, CRS05, CLR08, CY01, CS02, EH07, KEM^{+}01, Kos05, KT07, LÅT04, Liu07a, MNV00, Rob00, Rob01, SGK09, WYL09, WYX09, qX08, dA08]. Condon [GME06]. conductance [KACB07, Ver00]. conducting [CAW00, QR01, QG04]. conduction [JJHvO03, TNI^{+}07]. conductivity [GMBC08]. cone [Har00]. Conference [Ano07f, Ano08d, BDL00]. confidence [AS00, Bar02, Con04]. configurable [ATB^{+}01]. configuration [AAG^{+}04, AAM^{+}01, BM04, BKM05, FFG02, RCG05, SJF07, TEP00, TNCG00]. configuration-interaction [RCG05]. configurations [BK01, BKM02, BKM05, DCNDC09, GSF05, GHIL09, SKH02b]. confined [CH09, CW01, KEM^{+}01, LC07, PS08, SA09]. confinement [ASC^{+}05, BHM^{+}07, GAR05, WML^{+}05, WSCW09]. confining [AMP^{+}00]. Conformal [RM05a, HP02]. conformation [GW01b]. Conformational [BJ05a, LH02]. congruential [DH00, WH06]. conical [BCH05, PL05]. conjectures [JW02]. conjugate [GHP01]. conjugated [vdHBP^{+}02]. conmutant [IH09]. connection [OML09]. connections [MTC07]. conquer [SKNV05]. CONQUEST [GBTM07]. Conservation [Laf03, Che07, Esi01, Poi08, Poi09, TYN02, UOTM03, ZY09]. Conservative [IIK^{+}08, ML06, Cha04, Ida02, KNTG03, TNY00, UTO09, UNK12]. conserved [GKI02, GKI04]. conserving [ACIZ07]. consideration [Kon01]. considerations [Rap06, Wen01]. consistent [BTI01, DPSG06, PHKL02, Pet04, Pit05, WH00, ZSK^{+}04]. Constant [QP05, CCD07, CW01, HM06a, Kos05, LTA05, Lei02, Mor01, SN07, Var02, qX08]. constant-pressure [Mor01]. constants [PCCD09]. constituents [GMBC08]. constitution [Bur02]. Constrained [GOG00, JPS^{+}09, bHhL07, JS06, Ros04, Zim05]. Constraint [CLR08, DSH03, GZ07, Zie05]. constraint-transport [GZ07]. constraints [GWK09, OK06b]. construct [GT01, Yan02]. Constructing [DCNDC09, FMMQ08, Hin00, Liu07b, TNY00, Sev00]. Construction [CRW09, FMG00, PJK00, Vak00, BCC^{+}06, De 02, GKR07, KLM00]. constructive [Teh01]. Contact [KBBW02, SM06b]. containing [LTG09]. contaminated [SR01a]. Contents [Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01o, Ano01p, Ano01q, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02x, Ano02p, Ano02q, Ano02r, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano04-36, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41, Ano04-42, Ano04-43, Ano04-44, Ano05-31, Ano05-32, Ano05-33, Ano05-34, Ano05-35, Ano05-36, Ano05-37, Ano05-38, Ano05-39, Ano06d, Ano07g, Ano08e, Ano09e]. context [FKMB09, LM00, MC09]. continuation [Blü00, CC07, CCL08, Cza06, KBV09]. continued [RMLB01, TB85, Tho04b]. continued-fraction [TB85, Tho04b]. continuous [AP04, BR09, HSJ02, TL06b]. continuum [FGMT02, LBM05, LH01, TP01, TD03, VPCK04, Wol03]. contour [KCH00, KK06]. contracted [AC05a, AC05b]. Contrasts [ZSSA00]. contribute [BKB02b]. Control [PK01, AAM^{+}01, DB08, FGV01, KMR^{+}09, LNV^{+}09, Nii00, SS09b, WDHE04]. controlled [BDHP08, FG04, HOI04, TL06b]. convection [EELZS04, KT05, KNT08, MZB^{+}04, Ida00]. convection-diffusion [MZB^{+}04]. convective [KKSR04, Sus01]. convective/absolute [Sus01]. Convergence [BGJ^{+}07, LOL06, KSHP02, LH01, Wen01]. converging [Maa06]. converted [WBDB04]. convex [Dem06, KH06, RLRR06]. convex/concave [Dem06]. convolution [Kas00, YZW02]. cooling [FS01b, LLLZ01]. cooperative [IN09]. Cooperativity [LLPL08]. Coordinate [BD05, SHW01]. Coordinate-space [BD05]. coordinates [CMT00, CMT01, GVMW04, JBA^{+}07, LB00, SGL09, SFSL09, XON08]. copolymer [ZM00]. CORBA [LM00]. core [BVY05, HSS^{+}08, MMTH04, NM03, ON08, PKB^{+}01, RM05b, SBD^{+}06, TND04, TND05, WLR^{+}08, AIOST03]. Coriolis [CA07]. Coriolis-coupled [CA07]. corners [Ple02]. correct [Rob00]. correcting [ZS03]. Correction [SS02b, DVL^{+}04]. corrections [FGR06, JPS^{+}01b, KLD04]. correlated [AC07, Alv09, SOS01, Zha00]. correlation [BC05, FFD00, ICO03, LCPC04, PGS02, SKH02a]. correlations [BS04a]. Corrigendum [LPR04, MSHP20, Ras17]. corrugated [YW01]. cosine [CJC09]. Cosmic [Tol02, Min01, NRR01]. COSMOCR [Min01]. cosmological [ADBF03, BADC07]. cosmology [Min01]. Cost [Got01]. Cost-effective [Got01]. Cotes [Sim08]. COULCC [Tho04b, TB85]. Coulomb [Tho04b, AMP^{+}00, BBB^{+}04, DSC06, FSB09, HJZ09, IM01, LS01, MR05, Mic07, Nob04, OIKN02, OS03, PAT^{+}09, Sar00, Sea02a, Sea02b, Sea02c, TB85]. counterpart [BSB02]. counting [CM06, Car06, Car07, RCGC00]. Coupled [BOG^{+}07, CDH^{+}06, HPC05, ASF^{+}05, AK07, Bac00, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, CA07, EMJH03b, GLL^{+}02, Gut06, Ixa02, KPL07, KCR07, LVV06, pLbL03, Mel01, PKKM02, Riz02, SJF07, SQ03, TEP00, TY01, WJW09, YT01b]. coupled-channel [CGA^{+}07, CGVA08, CGVA09a]. coupled-cluster [PKKM02]. Coupling [DTHL09, CKS00, DK05, DKC08, FIBT01, FIT03, Fri09, GF01, GFF01, GFG01, GF02a, GF02b, GSF05, IFF01, PFG06a, PHKL02, Ste02, Tót06, dlGGS^{+}05]. couplings [EH06, JKW00, JKW06, PSW00]. covariant [CMM09]. cover [HBW05]. CP [HHW00, LVV04, LPC^{+}04, LCE^{+}09]. CP-even [HHW00]. CPC [BJS00, BB09b]. CPM [LVV06]. CPsuperH [LPC^{+}04]. CPsuperH2.0 [LCE^{+}09]. CPU [FEHC01]. CR [NY08, PKKM02]. CR-39 [NY08]. CR-CCSD [PKKM02]. Crack [MCC05]. Crack-tip [MCC05]. cracks [VKN07]. Cracow [Gre07]. Crank [Sch05]. crashes [Sor02]. CRAY [ALN^{+}01, WLH00]. create [Esq04]. Critic [dlRBPL09]. Critical [CM03, JJK05, LWT08, Bal01, BBJS09, BJ08, BL00, BMML05, DGAG06, FBB01, GDAG05a, GDAG05b, KSS02, MCH02, MKM02, SS07a, Sat02, TBL02, YGT^{+}02, ZSdD^{+}08]. criticality [KF05a, SOS01]. CRModel [HvDJvdM01]. Cross [AIOST03, BS03, Cip07, Cip08, Cip09, HSGBK08, Hor09, Kol09, LDBG08, MOC03, Nik03, OMC00, Pap01, Sal03, SBM09b, Yos03, Yos07]. cross-section [Pap01, SBM09b]. Crossover [BCBJ02, ACC09, ISSB01, RCG05]. crossovers [MKM02]. crumpling [TÅT09]. cryptanalysis [ÁMRP04, WLW04]. cryptographic [ÁMRP04, kWpLwW01, WLW04]. crystal [FFK02, GOH06, GLP03, KDW00, OGG07, RIB01, SSH02, SHX02, TDY02, PCCD09]. crystalline [JGJ09, LS02, PCCD09]. crystallization [LS09]. crystallographic [CPV^{+}08]. crystallography [BH01, SHV^{+}01]. crystals [All05, BVY05, CAAM08, GLHW01, IN09, KNSY07b, KNSY07a, LPRS02, LPR04, LPC^{+}00, PKRK07, PGS02, SSPM05, SYM00, TBR07]. CTEQ [Sul05]. CTM [WLH00]. Cuba [Hah07, Hah05]. cubic [GLHW01, XZ12, Zah00, Zah05]. CUDA [LSVMW08]. cumulant [SH05]. cuprous [HSSA01]. current [Ano01a, LCV06, MMMM00, iOY01]. currents [BEM^{+}02, NY07, PHKL02, RdAGV^{+}00]. curriculum [Gou00]. curve [LSL07, ZSdD^{+}08]. curved [Den08, Vul03]. curves [APV00, BLCR05, BFL04, CGG^{+}08, CGG^{+}09, PJK00, Tam03]. curvilinear [Cha04, ID09]. cusped [WGS00]. cuspoid [KCH00]. custom [Far01]. customizable [PKB^{+}01]. cut [CLFH07]. cutoff [MHK^{+}05]. cutoff-effects [MHK^{+}05]. CWENO [KKF^{+}04]. CWO [SWS^{+}12]. cyanoadamantane [FFK02]. Cyberinfrastructure [Cho07, dSdSW08]. cycles [DDFI09, TRAdO09]. cyclotron [PPP01, WBDB04]. cylinder [CAW00, Liu07a]. cylindrical [AP05, CS07, GVMW04, HFN03, LLV^{+}01, SGF04, SGL09, SBBM04, XON08, You09].

D [Aok01, AH02, BD00, Bal07, BJ05b, BFH05, CCFG05, Cha04, CSW02, CBBJ02, DGV08, EL04, FMD07, FDM07, FM00, FV02, GFP00, GRR01, GBM02, GS01a, GBA01, GMBC08, HG02a, HBRS05, JW02, KKSR04, KMZZ05, KSSH04, MZB^{+}04, MLF07, MNV00, NHS07, NSYZ02, PCV06, QR01, QTL06, RLRR06, SJCM04, SMV01, SG04b, SBB03, SG01, SBCZ08, SQ03, TAM04, TPYV03, WLH00, WCG04, WHL^{+}07, XON08, YRR07]. D-model [NSYZ02]. D0C [NN09]. DAFT [BTS06]. Dai [Hon04]. damage [VKN07]. damping [TGB01]. DAMQT [LRR^{+}09]. DAQ [Ano01a]. Darcy [KT05, KNT08]. Dark [BBPS09, BBPS07a, BBPS07b, RRCV09]. DARWIN [AOT01, TAM04]. Data [FSBG00, PK01, Sak07, Sfi01, SEC04a, AA07, AKG02, AAM^{+}01, Ano01n, Ano09t, BBB^{+}01, BB07, BFMH^{+}01, BH08, BGLLW01, CPV^{+}08, CGG00, DDMM06, Dem03, Dem06, Dom05, EFG^{+}00, GDC01, GMO03, Han00, Hin00, KL07a, LFT01, LKKK07, LZ00, MYC09, NW02a, OK09, OCK^{+}00, OPB^{+}09, PS09, PB09b, SOAW08, SSZ01, SS09b, SC04, TKS^{+}01, TK09, WLH00, WMK09, Wen01, YWLC04, ZSdD^{+}08]. data-compression [OCK^{+}00]. database [ABNÅ05, AAM^{+}01, BCC^{+}06, BB03, TLDM03]. databases [ME00, BHNW01]. DataScan [RSD01]. dataset [HCK01]. datasets [SKNV04]. date [Fri09]. Davidson [DM07]. Dawson [Ano04b, Ano04-56, Ano04-57, Ano04-45]. DC [CHL^{+}07, RMVQ07]. DDS [WH05]. DDT [IN02]. dealiased [ICT01]. dealing [SKF05]. Debris [WH05]. Debye [BFL04, LDZ^{+}08]. Decay [Bar04, BEM^{+}02, CRS05, EH06, JPS^{+}01a, Kol03, QxW07, Ste05, TJLR06]. decaying [BAB04, Str01b]. decays [BS04a, BS06b, GPW04, MDM05, Por03]. decision [VBFM05]. decomposed [ZA01]. Decomposition [BP08a, ST09, BH07, Cha07, DTHL09, GRR01, IW01, IW02, KBG00, LM02a, LZS06, Lüs04, OMF03, SWC^{+}03, Uhl03, YWLC04]. deconfinement [KMP09]. deconfining [KSS02]. Deconvolution [KSTL03, WCBN05]. decoupling [CKS00]. Dedicated [CMR01, Tri05, Yos01]. deep [RS00]. deep-inelastic [RS00]. defect [PKRK07]. defective [PLL07]. defects [GLHW01, HHCC05, LN01, LMS05]. defined [Gal00]. deformable [Nii00]. deformation [RvOvV02, vdSvdG08]. deformations [MAM04, MAM07]. deformed [Cle05, DD00, DO04, DO05, DSC^{+}09, RGD^{+}01, SDNR05, TBR07]. degeneracy [HG02a]. degree [UNK12]. Delaunay [SMH04]. DELPHI [BCCM03]. demagnetizing [BD00]. demand [ZS08]. Demonstration [ABD^{+}05]. demonstrator [ACC^{+}01]. dense [Bun01a, CGG00, PM02, WRC^{+}04]. densities [BSTC05, BH08, CS02, GKI02, GKI04, Mam08, Moh08, dlRBPL09]. Density [BJ02, JK02, Kur02, MLF07, SG05, Alv09, Bae04, BBPS02, BBPS07a, BBPS07b, BSK^{+}03, CSW02, HHM^{+}09, HLC08, HS01a, ICO03, IBA00, KTT09, KH09, LV08, Lik01, LCV06, LRR^{+}09, MC03, MBR01, OLS^{+}01, OSK^{+}02, RLH^{+}09, RGD^{+}01, SKNV01, SKNV05, SSZ01, SMK01, TAM04, VKM^{+}05, WN01, dlGGS^{+}05]. Density-driven [MLF07]. Density-functional [Kur02, SG05, HHM^{+}09]. density-functional-theory [SKNV01]. density-matrix [WN01]. Departure [ACC09]. dependence [BS00b, MSS^{+}07, NFS02, PP02, SZ00c]. dependencies [PZ01]. dependency [ZS08]. Dependent [RPY07, BC05, BSK^{+}03, CRS05, DKMF03, DKC08, GNZ^{+}09, HTM^{+}08, HGVCM^{+}02, KFB01, MS06, MLG^{+}01, Mei01, MA09, NM01b, Nur04, PSV00, SŽ00a, TKN^{+}08, dlGGS^{+}05]. deployed [KFJ^{+}09]. deployment [Sak07]. deposition [Sch08, SLWH02]. Derivation [FAiTD01, AHS09, OMF03, SH05, WYL09]. derivative [Jam00, WGDZ04, WC05, ZWD05]. derivatives [AS03, CGVA09b, GSS06, GZDA01, KCH00, MA04, MKK05, MA08]. derive [EL06]. derives [Rob00]. described [DVL^{+}04]. Description [BJS00, BBC^{+}01a, PKB^{+}01]. description-driven [BBC^{+}01a]. Design [ABC^{+}01, BBC^{+}01a, MTLC01, MP01a, Rap06, AAKL07, CFJ09, DG08, Far01, PCA^{+}07, QRH00, RMMP02, SKNV01]. designed [KKM02, LTG09, Str05, WSCW09]. DESOLV [VBC07]. Detailed [Wro08, CD09a]. details [BDF^{+}08, PJSK08]. Detection [HHCC05, ABC^{+}03, BBPS09, CSS^{+}03, CM06, KV08, LANM^{+}01, TWY09]. detector [AAM^{+}01, BCC^{+}06, Maz00]. detectors [ABC^{+}03, NY08, PBI07, Sch08]. determinant [FA00]. Determination [YT01a, Fat02, IF03, Pet04, VEG08]. determine [BCV03, PRBD09]. Determining [ADDdM07, BKKS09, QTMH07, BVY05, BDW06, HOT07, PSH06]. Deterministic [DDM05, DVG05, LZC^{+}08]. deterministic/stochastic [DVG05]. detonations [NBPG08]. deuterium [Bat03]. developing [MRF^{+}05]. Development [HFN03, HCH^{+}06, ICO01, KKHL07, WHL^{+}07, YSM09, PFPB^{+}09, Teh01, Dan09a]. developments [HCIK00]. device [GCD06]. devices [AAG^{+}04, CBKM01, DC05a, KKKC07, LYL07, LLCS01, TDY02]. devoted [BP08a]. DFT [HTA08, PLL07]. DGLAP [CCG08, Tol02]. dHybrid [GBFS07]. diabatic [MN01]. diagonal [vH06, vH07]. diagonalization [CA09, FM00, GFS03, LB09]. diagram [NT05, TF00, TL09]. diagrams [Bar03, BT04, BCKT09, CCGR09, DKC08, FK00, Hah01, HL08a, KKK06, Ots01, TF04, VMMB02]. DIANA [TF00, TF04]. diatom [HSGBK08]. diatomic [MDT03, PJK00]. Diatomics [NW02a]. Diatomics-in-molecules [NW02a]. diblock [ZM00]. Dielectric [FER^{+}07b, Bre07, Den08, HKPL07, KM08a, LBM05, LLT^{+}02, NJ01, Zha01]. difference [BGH04, BTI01, CCL08, DR09, EMJH03b, GVMW04, GKI02, GKI04, GMAHV^{+}09, HL05, Inu07, MZB^{+}04, NN06, RLV^{+}08, Rob01, SHX02, Wan09b, KSC^{+}00]. difference-difference [GKI02, GKI04]. difference-differential [Wan09b]. differences [GLL^{+}02]. differencing [EMJH03a, dlHV08]. Different [ABSM04, BSvdDW02, Ska05, CFJ09, GMBC08, HM00, Kim03, LVLS01, PSK01b, RDSS01b, SPV07]. Differential [Hoo04, IH09, IHAR09, KA09, MGYP08, BGH04, BT01, BCV03, Che07, Don02, FMG00, GSGT03, HSGBK08, KS07, KD09, Ram05, Ras09, Ras17, VBC07, Wan09b, WR01, qXbL04, Yan02, Yan03d, YZW02]. differential-difference [BGH04]. differentiation [AA07, SPF00, Str05, VHLP09, vH06, vH07]. Diffpack [Hoo04]. Diffraction [BRdAHK04a, BSO^{+}04, BK05b, FBB01, Gro01, GSSN00, MTZ00, MP06, PCC01, Wan01, BRdAHK04b, dAK01]. diffractive [CF02]. DIFFREALWAVE [HSGBK08]. Diffuse [dA08, GLMADB^{+}02]. diffusion [AGV00, BS00b, BNSY02, CZC00, CL03, CJK09, EELZS04, GS01a, IOM00, KP01, MZB^{+}04, PC08, Ram05, RMK05, RPD^{+}05, SMSE03, TE05, VPNW02, Wei02a, WLR^{+}08]. diffusion-convection [EELZS04]. diffusion-driven [KP01]. Diffusive [BDHP08]. Digital [iSAK^{+}08, AAA^{+}00, RTVZ08]. DIII [KSSH04]. DIII-D [KSSH04]. dilepton [Abe01, Abe01]. dilute [CP00, SVS01, TS06]. diluted [BCBJ02]. dilution [CBBJ02]. dim [GBD03, GCP^{+}02, GDC01]. dimension [BGJ^{+}07, GBR^{+}09, NJ00, SBD^{+}06, TdRGD09, Vor02]. Dimensional [Ker02, AC07, Bac00, BTK^{+}02, Bre05, CSC^{+}07, CSC^{+}08, CLL^{+}07, CD01b, CTG01, CHM^{+}09, CNDC09, Eli05, ES09, FHR^{+}05, GR02, GOG00, Har02, HBMJ05, HJZL07, Huj05, HW09, IH01, IIK^{+}08, ID09, Inu07, Jad00, JKKT00, KT05, KNT08, KNTG03, KKF^{+}04, KNU00, KM01c, KK01, KA04, Krö05, LHC01, LHC02, LSL07, LCS07, LVLS02, Li03, LHS^{+}06, LKPH08, MYC09, MMTH04, MSD08, MOS00, MOS01, NYH04, Ots01, PKRK07, PD08, Pis00, QG04, RtVR09, Ram05, RM05b, RLV^{+}08, Rob00, RG04, iSHS^{+}08, Sch06a, SBJ05, SSB^{+}09, SW09, SD07, STK^{+}00, TRGR08, TMTF00, Tak03, TYN02, TZZ06, TNI^{+}07, Tat07, TY01, TYSH05, TDD04, TL09, UTO09, Var08, Ver00, Vos06, WGDZ04, WC05, WHJ06, WS09a, WTW04, XZ12, Yam00, Zak06, vHK00]. Dimensionality [Pis00, Bac02]. dimensions [BR01, BCBJ02, BSDMH05, BKKS09, Cre00, EELZS04, GBC^{+}04, JK08, LEG02, MK08, NGE^{+}04, PM00, SMH04, SGF03, SM02]. dimer [CCBL02, EVL00, Kim07, KK01]. diminishing [NRR01]. diodes [HTL^{+}03]. dipolar [TZZ06, TK08]. Dipole [DN05, SWY01]. Dirac [BBR03, BFLW07, Cun09, FZ09, GL02, GZDA01, Lüs04, MW01, MK08, MKK05, Moh07, NM01a, Pog05, SJP05, TD03, Vul03]. Direct [CC04, BBPS09, Ber03b, BDB^{+}08, FD03, Hon04, Nur04, PGS02, TG00, Wal03]. direction [LTT09, XZ12]. directions [CL08a, MSK^{+}02]. discharge [KPL07, KHIL07, KA04]. discharges [BLS09b, CS07, HKLY07, LHS^{+}09, RMVQ07, TC07, WJW09]. disconnected [Krö05]. discontinuous [Gal00, LS05]. Discoveries [TPBE04, ZSSA00]. discovery [DDEM00]. Discrete [BW01, TÅT09, Wan09b, BDK^{+}06, CNFR01, DW01, Har00, LbotMC01, Luo00, MSD08, Str00, TSI02, WHO02, YZW02]. Discrete-expansions [BW01]. Discretization [EG09, MM04, TKSR00, DPSG06, KT05, KNT08, WS09a, Zit09]. discretized [CHS09, DB08, HZGZ09, OS00b]. diseases [Dom05]. disk [WH05]. disks [SSLN02]. dislocation [Cle05]. dislocations [MS05b]. disorder [GW01a, Voi02, Voi03]. Disordered [KM05, CM02a, NH09, RLU01, Ver00]. Dispersion [QCML03, CL02, Sus01, Tam03]. dispersions [KNY05, YNK05]. dispersive [LBPS09, NN06, Ram10, Ram12, YB02a, Yan03a, Yan03b]. displaced [GME06]. displacement [Alf09, BA09, GLHW01]. display [BCD^{+}01, LYL07]. disruptions [EH03]. dissemination [KL07a]. dissipation [GIME02, KG07]. Dissipative [ZM00, AKS02, DC03, HNS01, NT05, NKV03, SVA03]. dissociation [HGVCM^{+}02]. dissociative [PNH00]. Distributed [FSBG00, GC01, AKG02, BLM01, BV00, BTS06, CDH^{+}06, Di 01, Han00, HKP02, KAB^{+}00, LbotMC01, LNV^{+}09, LM00, TG00, TYS^{+}00, WMK09, Xia01, dSdSW08]. Distribution [MK02, BDBV12, FGA04, FPB08, HTM^{+}08, HBMJ05, JH09b, KHIL07, KS84, KS08, LWY01, LC01b, OPB^{+}09, Ram10, VS06, Yan09, Yos07, YW01, vHK00]. distributions [BBOY08, Har02, JJK05, LHC01, LHC02, LV08, RF08, Sul05, Vog05, Wei02b, WHO02]. divergence [JH09a, MOS00, MOS01]. diverse [ZPB09]. divertor [KY07]. divide [SKNV05]. divide-and-conquer [SKNV05]. DL [Sea01]. DL LEED [Wan01]. DL POLY [BTS06, DHBE05, KSYE00]. DNA [CDFF05, DLZ08, Dom05, Ger07, KK05, LOY07, Zim02]. DNA-chip [Dom05]. DNS [HDG07]. documents [GM00]. Does [Nur04]. Domain [BP08a, BH07, CMT00, CMT01, Den08, DTHL09, Hei01, HL05, IW01, IW02, KBG00, LM02a, Lüs04, NFS02, NN06, Uhl03]. domains [BLM01, GS01a, NFS01a, NFS02]. doped [JK01, MB05b, NW02a]. Doppler [WCBN05]. dot [BNSY02, CLFH07, EMJH03a, EMJH03b, LLV^{+}01, LCV06, WHJ06]. dots [EMJH03b, LVLS01, MWA01, MP05, RCG05, TNCG00, Vos06, WV05, Wan00]. Double [HG02a, BMML05, BK05b, CWW07, FMN01, KFI^{+}01, LY05, RF05b, RF06b, VYK02]. double-photon [KFI^{+}01]. double-stranded [VYK02]. doubly [ACIZ07, Yan02]. doubly-periodic [Yan02]. doubly-polarized [ACIZ07]. down [CM03, TJLR06]. DP [LJ09b]. DPEMC [BK05b]. Dr [AA01b]. drag [KMB02, MY00a]. DRAGON [Tom09]. drawing [BT04, BCKT09, Cap05, HL08a]. drift [BGS^{+}04, HFN03, Jen00, Lew04, PCK00]. drift-fluid-electron [PCK00]. drift-kinetic [BGS^{+}04]. drift-wave [Jen00]. drilling [ZZH09]. driven [BBC^{+}01a, FV02, KP01, KLD04, KA05, KS04a, LLPL08, MLF07, MK05, PP02, PFPB^{+}09, RJFB08, RK05, Dan09a]. Driver [MP03]. drop [BBD00]. droplet [CMS04, KPS^{+}01, vdSvdG08]. drops [MDH04]. drunken [CN00]. dry [OML09]. DSMC [CSC^{+}04, WTW04]. DST [ADD^{+}03]. DTORH3 [GS01b]. dual [BD08, GPT08, KCR07, Liu07a, TC07]. dual-frequency [KCR07, TC07]. dual-kinetic-balance [BD08]. dual-phase-lag [Liu07a]. due [BBBR04, EG09, Voi02, Voi03]. Duffing [AA08, DWZS05, Wan06a, Wan06c, WW06, Yao09]. Duhem [FFF01]. duplexes [BSB02]. during [Dan05b, EH03, KDSB04, NG02, ZS08]. DVR3D [TKB^{+}04]. dwarf [RRRHD08]. dxhdf5 [SC04]. Dynamic [CM02b, DKC08, DGAG06, DHS00, TBL02, BSDMH05, BKB02b, DVG05, GDAG05a, GDAG05b, LNV^{+}09, MKB02, Nur04, PCA^{+}07, QCL05, SMH04, Sch05, SWC^{+}03, SHH^{+}04]. dynamic-theta [Sch05]. Dynamical [BDF^{+}08, Dan05a, DC03, LKKK07, NFS01a, NM01b, ADDdM07, BDT00, BBJ^{+}08, BFB^{+}09, BAB04, BFI^{+}00, Cun09, CKLS09, DS06, Dan07, FGF03, HK02, HOI04, LFT03, MSD08, Rob00]. Dynamics [BDHP08, KRTZ02, KMP09, NFS02, Par04, PKRK07, RvOvV02, TYS^{+}00, ZM00, ASS^{+}02, BSB02, BBB^{+}04, BTS05, BBB^{+}09b, BTK^{+}02, Bro07, BK05c, CW02, CSCK08, CDFF05, CCFG05, CLFH07, CFJ09, CCD07, CTI07, CKV04, CR00, CW00, CC00, CA07, DELG05, DC03, DD01, DHB^{+}04, DHBE05, FG04, FSK04, FS01b, FS02, GF02c, Gol00, GLW03, GW01b, HDGM07, HOT07, HL00b, HGVCM^{+}02, HM06a, HKK02a, ID09, IW01, IW02, IN09, ISH01, JAT03, KBBW02, KEL02, KCC^{+}00, Kar02, KM01a, KFJ^{+}09, KSYE00, KKSR04, KA05, KBC^{+}09, Kon01, KBG00, KS04a, KM05, LM02a, Lei02, LR07, LZS06, LZS08, LMM^{+}08, LSVMW08, LM02b, MCBR03, MFF^{+}05, Mel05, Mor01, MSH01, NT05, NH09, NKV03, iNKNV08, OSK^{+}02, OK06a, OK06b, OD07, OCK^{+}00, OMF03, OMY05, ÖDÇ02, PLPS08, PL05]. dynamics [PLS09, PZW^{+}00, PP02, PS09, QRH00, QP05, Rap02a, Rap06, Rap08, Ref00, RFK08, RRCV09, RJCH00, Ryc05, SNS01, SHZ01, SKNV05, SGL09, SLBG09, SEE^{+}03, SS02b, TT06, TNI^{+}07, Tod01, Tsa02, Val05, VKPB09, VKN07, itVPG08, WHCL07, WRMG05, WHL05, ZE00, dO02]. dynamics-whither [Rap02a]. dynamics/electronic [OLS^{+}01]. dynamos [RJFB08]. Dynasol [PZW^{+}00]. Dyson [AHS09, Maa06].

EARLINET [OPB^{+}09]. early [LAF01]. earth [KL01, Swi04]. easy [Ano09s, CD09b, Hah09, NR01, Esq04]. ECAL [Org01]. economic [SAG^{+}02]. economics [Sor02]. ECPSSR [Cip07, Cip08, Cip09, Hor09]. ed [Hoo04]. Eddy [TIM07, TIM08, DMR02, DVG05]. EDF [FPB08]. edge [ATIO06, CFJ09, PSK01a, PSK01b, SBM^{+}04, SGF04, UXD^{+}09]. edge-lit [CFJ09]. edges [Ple02]. Editorial [Pub07, Ano05-40, Ano05-41, Ano05-42, Ano05-43, Ano05-44, Ano05-45, Ano06e, Ano06f, Ano06g, Ano06h, Ano06i, Ano06j, Ano06k, Ano06l, Ano06m, Ano06n, Ano06o, Ano06p, Ano06q, Ano06r, Ano06s, Ano06t, Ano06u, Ano06v, Ano06w, Ano06x, Ano06y, Ano06z, Ano06-27, Ano07h, Ano07i, Ano07j, Ano07k, Ano07l, Ano07m, Ano07n, Ano07o, Ano07p, Ano07q, Ano07r, Ano07s, Ano07t, Ano07u, Ano07v, Ano07w, Ano07x, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano08n, Ano08o, Ano08p, Ano08q, Ano08r, Ano08s, Ano08t, Ano08u, Ano08v, Ano08w, Ano08x, Ano08y, Ano08z, Ano08-27, Ano09f, Ano09g, Ano09h, Ano09i, Ano09j, Ano09k, Ano09l, Ano09m, Ano09n, Ano09o, Ano09p, Ano09q, Ano12]. Editors [Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09c, Ano09d]. Education [Lan07, Chr00, Esq02]. educational [WCGL00]. ee4f [KJ04]. eett6f [Kol03]. Effect [Gen01, Car07, CSC^{+}07, CSC^{+}08, EMJH03b, IN02, KL01, KSTL03, LOY07, LC08a]. Effective [KTL05, Yam00, AC09, GMBC08, Got01, KP01, SWP03, Wan05a]. Effects [JTS^{+}06, BCBJ02, BCH05, CSCK08, FHR^{+}05, FFD00, GSM^{+}03, HCO01, HS01a, LLT^{+}02, LWY01, LA09, MHK^{+}05, RCGC00, SS02b, WGY01, dAK01]. Efficiency [SJ02, CM06, Car06, Car07, KNTG03, LC08b, Nat08, PWH^{+}00, SLL07]. Efficient [AJT^{+}07, BPRW06, CA09, DDMM06, FSB09, Ixa07a, JGJ09, KMZZ05, LVV09, LPC^{+}00, LR07, MMR04, MKJ^{+}05, PKKM02, SSP08a, Tak03, Vog05, WT01, Zim02, dlGGS^{+}05, vE08, BFMH^{+}01, CZC00, DWZS05, GLL^{+}02, GBD03, HLB06, KN07b, LTG09, LJ09a, MHR^{+}07, Mei01, MNH01, Müs02a, OD07, OGWH03, PKPV02, RD05, SI01, SOAW08, Sim00, Tót06, itVPG08, WKP^{+}01, WW05, YB02b, dA08]. Efficiently [ZZ09, BLM01, Bre05, KKK06]. effort [Ort00]. eigenenergies [Jam00]. eigenfunctions [CGVA09b]. eigenpairs [DM07]. eigenproblems [ALN^{+}01, RB08]. eigensolver [Bun01a]. Eigenstate [NFS01b]. eigenstates [Bac00]. eigenvalue [ABD^{+}05, GHP01, LVV07, LCHJ09, SHX02, SMZ05, ZSM05]. eigenvalues [BN07, CGVA09b, KMS09, LVV09, MM01, TYSH05]. eigenvector [Cun09]. eigenvectors [CFKM01]. eighth [Sim00]. Einstein [BGJ^{+}07, BBR03, CPS00, CC07, CCL08, CC09, DDM07, Dy09, HB05, LR07, Nil07a, SVS01, TQ03, ZZ09]. EJB [LM00]. elastic [Cha07, DHBE05, GLHW01, LV08, Nak08, PCCD09, RK05, SJP05]. elasticity [ACC09, LÅT04]. elasto [SM06b]. elasto-plastic [SM06b]. Electric [NY07, FSK04, KMR^{+}09, NT04, SGF04, WTH^{+}04]. Electrical [CTI07]. electrically [Ram12]. electro [Pis00]. electro-magnetic [Pis00]. electrochemical [HL00a]. electrochemistry [SN07]. electrodynamics [Har00, KNU00]. electrokinetic [PCF05]. Electromagnetic [CAW00, FS08, PCK00, DEW00, DW01, GFP00, HL05, JBBR01, Jen00, JTS^{+}06, KV07, LKPH08, PP09, PD08, PSP^{+}03, Poi08, Poi09, Ram10, SLMS06, UOM01, UOTM03, UTO09, VAH04, Ver04, WPL02, WRC^{+}04]. electromagnetics [FKMB09]. Electron [BRdAHK04a, BRdAHK04b, LLV^{+}01, MMMM00, NKSL05, RdAGV^{+}00, dAK01, AC07, ABSM04, Alv09, BF04, BPP01, BM04, Car06, CKV04, DC05a, EÅU05, EKW09, FPB08, Frü03, GPT08, GG03, Gut06, HPC05, KKKC07, KHIL07, KA04, Kon01, LV08, LVLS01, LVLS02, LRR^{+}09, Mam08, MCBR03, MWA01, Moh08, Nik03, dlRBPL09, PCK00, PPP01, RMLB01, RCG05, SKH02a, SBM09b, SMV01, SJHY07, SNBB02, TAM04, Ton07, Wan01, WD04, WRN01, WM00, Yak01, ZPB09, Zha00, Zha01]. electron-atom [GG03, SNBB02]. electron-capture [Car06]. electron-cyclotron [PPP01]. electron-ion [BF04, HPC05, MCBR03, SNBB02]. electron-molecule [WM00]. electron-positron [BPP01]. electron-transfer [DC05a]. electronegative [CS07]. Electronic [FW01, HP06, LTT09, LLLZ01, MWA01, SBM02, SN07, TGB01, Zha01, AJT^{+}07, BTI01, BB00, BMG01, CPV^{+}08, CTSZ07, GHP01, GBD03, HC00, HTM01, KFJ^{+}09, KLM00, LZ04, MSB09, PKSF01, QASF^{+}05, RG05, RB08, SKNV01, SMH^{+}01, THM01, TNCG00, Vos06, WKP^{+}01, YG09]. electronic-density-functional [OLS^{+}01]. electronic-structure [KLM00, MSB09, PKSF01, RB08]. electrons [EÅ01, EH03, Hor09, MK05, RLI07, SJP05, SMSE03, Sri01]. electrophoresis [KKM02, KK04]. electrophoresis-computer [KK04]. electrorheological [MS05a, SWY01, YW00]. electrostatic [AH02, BGS^{+}04, CSC^{+}07, CSC^{+}08, DTHL09, MB04, WJW09, WHL^{+}07]. electroweak [ABB^{+}09, HL08b, KP00]. element [BDK^{+}06, BLS09b, CN01, EFS^{+}08, GPT08, PKSF01, PDA06, TÅT09, Ton07, Whi00, XSC09]. element-dual [GPT08]. element/molecular [OLS^{+}01]. elementary [Fod05, Str01a]. elements [AC05a, AC05b, CN01, CGG^{+}08, CGG^{+}09, ÇHM00, GF01, GFF01, GME06, HL08b, JBBR01, KTL05, LCHJ09, LS01, OS03, PCE^{+}08, PAT^{+}09, Sar00, UTKF05, VS06, You09]. elevation [RTVZ08]. Eley [LJ01]. Eliminating [LC08a, Man02]. elimination [WR01]. ellipsoidal [LB00, WV05]. elliptic [AEB02, PKST03, Yan02, Yan03b, Yan03c]. Elman [TWY09]. ELMFIRE [SOAW08]. elongational [MDT03]. elsepa [SJP05]. Embedded [SKNV05, ASVA00, Far01, KDW00, Vég04]. Embedding [Ing01]. Emden [SPS09]. Emeraldine [CCFG05]. emergence [KOS^{+}09]. emerging [Lüt00, REAB09]. Emery [DKC08]. EMILIA [Car06]. Emission [RdAGV^{+}00, HCH^{+}06, KFI^{+}01, LC08b, RLV^{+}08, SJHY07, Yos03, Yos07]. emitted [CP00, HD04]. emitter [LC08b]. emitters [Car06]. emitting [HTL^{+}03]. empirical [SSZ01]. employing [KKF^{+}04, RMK05]. Emulator [DHMD00]. Emulsion [vdSvdG08, UVLRRC09]. EMX [AEB02]. enantiomeric [GLL^{+}02]. encounters [RRRHD08]. encryption [LMC^{+}03]. endpoint [LWT08]. energetics [OBG09]. energies [BCG03, CWSH08, EKW09, JWW00a, JPS^{+}01a, KPD06, LK07, Sea02a, SVMT00]. Energy [BBC^{+}01a, BRdAHK04a, DH01, FGV01, New07, TS08, Tol02, ATP01, BBB^{+}09a, Bes02, BFL04, BKM02, BBB^{+}09b, CCBL02, CC07, CCL08, Che05, CGA^{+}07, CGVA08, CGVA09a, Cra01, CCRA05, DVL^{+}02, DVL^{+}04, EÅ01, EVL00, FK00, Frü03, GGG01, GK05, GLL^{+}02, GPW04, GSSN00, GMO03, HP06, HG02b, HGH^{+}05, IK00, IKO00, KHIL07, KW03, LPC^{+}00, LVLS01, LLV^{+}01, LAF01, MR06, MNYY00a, MSHP02, MSHP20, MM01, Nob04, OD08, Sak07, Sch08, SEF^{+}01, SMH^{+}01, Sol01, SR01b, SKRK04, SFR05, TAP01, TZZ06, TYSH05, VS06, Wan01, WZHZ06, WLGX09, XSC09, dMBC^{+}06, BRdAHK04b, dAK01]. Engine [ON08, Vég04]. engineering [HKK^{+}01]. Engineers [Mal00, Bre01]. Engui [Hon04]. Enhanced [PM02, EHHH06, RMWH01, TGD07, TL08a, WSB04]. Enhancements [SRR^{+}00]. Enns [Koc02]. ensemble [Ber02, GCK02, HM06a, Huk02, JBS08, NSMO02, Nak08, OK06b, Zim05]. ensembles [IW02, OO05, WV04]. Entangled [KSEG05, Ryc05]. entanglement [RF06a]. entanglements [Krö05]. enumerated [SH06]. Enumeration [Jen01, BM06, SBJ05]. envelope [HS07]. envelope-kinetic [HS07]. environment [BCH05, CSZ^{+}07, GKP^{+}06, KPD06, KL07a, KW07, PDL04, iSAK^{+}08, TJD09, WR01, ZC09]. environments [PKB^{+}01, ZPB09]. enzyme [HJM02]. epitaxial [AFP02, BSvdDW02, Dan05b, NSYZ02]. epsilon [ÇHM00]. Epstein [Ram10, Yan09]. equal [PR06, Zak00a]. Equation [KD09, AA08, ATIO06, AKZ00, ASVA00, AKS01, AKS02, Bat03, BFH05, BV00, Cai09, CSCK08, CPS00, CRS09, DWZS05, DR09, DM09, DC07, DGSL09, DSL09, DKV00, Dys02, EELZS04, Fij99, Fij00, FS01a, FZ09, GNZ^{+}09, HCH^{+}06, IH09, Inu07, Ixa02, Ixa07b, KMS09, KA09, KEM^{+}01, KBV09, Kos05, LRI^{+}06, LIR^{+}06, LOCJ05, LB00, Li03, LL08, Luo00, Lüs04, MZB^{+}04, MK08, MA09, NT04, NJ01, Nur04, PC08, PAD07, PSK01a, PSK01b, PSV00, Ram05, Riz02, RLV^{+}08, ST02, SŽ00a, SG06, SR05, SW09, SGF03, Sim00, SW00a, SVA03, Sim09, SZ00c, SM02, SFSL09, Sug01, TPYV03, TQZM08, TS06, TD03, TKSR00, UNK12, UK02a, UYK^{+}04, Van05a, Vul03, WGDZ04, Wan05a, WC05, Wan06a, Wan06c, WS09a, WT01, WV05, WW06, XSC09, XZ12, Yan03d, Yao09, YB02b, Zak00a]. equation [Zak00b, Zak01, Zak06, ZY09, dlHV08]. Equations [Hoo04, IH09, IHAR09, AP04, ACK05, AHS09, AMP^{+}00, AK03, AK07, BGH04, BD05, BDP00, BT01, BDH^{+}05, BCV03, CC04, CLR08, CHS09, Che07, CJK09, CGG^{+}08, CGG^{+}09, CTG01, DKMF03, Den08, DD00, DO04, DO05, DSC^{+}09, EST00, EG09, FD03, FRdS09, Fat02, FMG00, FMMQ08, GT01, GKI02, GKI04, GSGT03, Hon04, HZGZ09, HHWH07, HHL06, IOM00, JK08, JC08b, KS07, Kas00, KKSR04, LVV04, LVV06, bLpL02, LL04, LY05, LJ09b, pLbL03, yMS01, Maa06, MM04, MOS00, MOS01, NKV03, OGWH03, PS08, PKST03, PNH00, PCV06, Ras09, RE09, Ras17, RMK05, RB00, SH05, She03, Sho04, Sho07, SPS09, SDNR05, Str00, TKP06, Tol02, UK02b, VBC07, Wan09b, qXbL04, Yan02, YB02a, Yan03a, Yan03b, Yan03c, Yan03d, YZW02]. Equations-Numerical [Hoo04]. equilibria [CHM^{+}09, KZS^{+}00, MSK^{+}05, SVMT00, ZSK^{+}04]. equilibrium [BDBV12, BL05, CD04, Col07, DGV08, Elm09, HYY07, JBA05, TCF00, ZSSA00]. equipped [Hor09]. Equivalence [CTR00, Ram10]. equivalent [EÅ01]. Eratosthenes [AA01a]. ERCS08 [Hor09]. Ergodicity [BKB02a]. ERI [REAB08, REAB09]. erosion [CMD00, WSB04]. Erratum [AAB^{+}07, CSC^{+}08, CGG^{+}09, CGVA09a, FNR^{+}07, Fij00, GDAG05a, Ida03a, JKW06, KM01b, KS08, MOS01, Nat10, Poi09, SM06a, TA00a, Tho04a, Tho04b, TND05, TIM08, WA07, Wu10]. Error [HDGM07, KL06, KTT02, Sch04, EG09, JC08a, MG09a, MG09b, SK08, vdEFL^{+}02]. errors [AW04, Inu07, WA07]. Escape [Man04]. ESPResSo [LAMH06]. essential [AFK^{+}07]. essentially [KG07]. Establishment [BB09b]. estimates [GKM^{+}00]. estimating [GBR^{+}09, JC08a]. Estimation [FGA04, BH07, CNFR01, Cra01, KMH02, KTT02, LFT03, Lik01, SSZ01, TGD06]. EtabFDC [QWWZ09]. etch [NY06]. etched [NY08]. etchers [HKPL07]. etching [MSY07, RLRR06]. ETSF IO [CPV^{+}08]. Euler [CTG01]. Eulerian [FS03, PSK05, SGF04, SSB04]. evaluating [GFF01, Mei01, PCE^{+}08]. Evaluation [FKMB09, Nob04, SM01, SG00b, ST09, WD04, BBR03, BK01, Bru04, CGM01, Dun05, GR01, GR02, GS01b, GZDA01, HS03, KS05, KNTG03, KSHP02, KCH00, Mam08, MNH01, MKK05, MN07, Moh08, MC09, OS03, Pis00, PR06, Sav01, Sch06b, SSP08b, SMH^{+}01, Sul05, VW05, WCBN05, dMBC^{+}06]. evaluations [Sal02]. evanescent [DEW01]. evaporation [DTD^{+}02, WSB04]. even [HHW00]. Event [Ano07-31, BFMH^{+}01, KRW03, KFI^{+}01, ACC^{+}01, BCD^{+}01, BBB^{+}09a, CDEW04, CB05, DDM05, DH01, JWW00b, JPS^{+}01a, Jad03, JS07, LbotMC01, LMP^{+}09, MP06, QWWZ09, RRCV09, SVP09, SEF^{+}01, SS02a, TS08, TSA^{+}03, TKK^{+}06, ABF^{+}01]. event-based [DDM05]. events [ACIZ07, BDG^{+}08, Nak07]. Evidence [BMML05, LCPC04]. Evolution [CS02, Cle05, KHIL07, SR09, AG05, BFI^{+}00, CCG08, CTSZ07, FS01a, HHL06, JS06, JPS^{+}09, KZS^{+}00, bLpL02, LL04, pLbL03, MSY07, MA06, PRBD09, SOYN01, TQ03, TGD06, Vog05, Wei02b, ZSK^{+}04]. Evolutionary [GOH06, vHLP08, AOT01, BSO^{+}04, CNFR01, Iwa01]. evolving [Bow02]. Ewald [Har02, LHC02, BTS06, FMD07, LHC01, OD07]. Exact [AC07, BM06, EELZS04, Esi01, GME06, LR07, QASF^{+}05, RM05b, SBJ05, CA09, CAF^{+}03, Dy09, HB05, LL04, pLbL03, LL08, MC03, Mil06, Mil07, Poi08, Poi09, Sch06a, SGL09, TYN02, Yan03d]. Exact-exchange [QASF^{+}05]. exactly [HNS01, TNY00]. EXAFS [TK09]. example [Teh01, Vég04]. exchange [FMN01, Gut06, PLS09, QASF^{+}05]. excitation [BCP04, DDFI09, TND04, TND05]. excitations [OS03, YG09]. excited [BCH05, CWW06b, LHMB00, MHGV09, MCBR03, NW02a, RDFF02, TA00a, TA00b]. excited-state [BCH05]. exciton [KN07b]. Excitons [vdHBP^{+}02]. excluded [BDH^{+}05]. exclusive [MP06]. execution [BLM01, REAB08, Tót06]. exercise [ZS07, ZS08]. exhibiting [KLTH04]. ExHuME [MP06]. exited [BAB04]. EXOTIC [TA00a, TA00b]. expanded [Cip09]. Expanding [HHM^{+}09, HM08, Fel08, FT08, HM06b, KTG04a]. expansion [ASF^{+}05, CRUV00, FSB09, KTT02, NFS01b, Pit05, San00, VK09b, WP06, Wei02c, Yan03c, Yan03d]. expansions [BW01, HJ02, RS09, Sea02b]. Experiment [HLW05, HKL^{+}07a, HKL^{+}07b, ADD^{+}03, ABF^{+}01, Ano03h, EFBP04, HJM02, KB02, TLDM03]. experimental [AA07, Ano01n, CHL^{+}07, ZSdD^{+}08]. experiments [DDM07, FGV01, GGQ01, Gre04, HLW05, HKL^{+}07b, SG04a, SEC04a]. expert [KS07]. Explicit [GFP00, TQ03, AKS01, De 02, FSW08, GWK09, JH09a, LPC^{+}04, LCE^{+}09, MVJ09, ON08, Van05c, itVPG08]. exploitation [ADE^{+}02]. Exploiting [MG09a, SPM00, TYS^{+}00, VHLP09, YN05a]. Exploring [MSS^{+}07, PL05, LLY07, SHH^{+}04, SIE04]. Exponential [VAMVR08, MG09a, MG09b, Ram12, VC08, dlHV08]. Exponentially [Fra07b, IVD03, ASVA00, CFMR08, KMS09, Sim00, SW00a, SVA01, SVA03, Van06, VIV01]. exponentially-fitted [ASVA00, KMS09, Sim00, SW00a, SVA01, SVA03, VIV01]. exponentials [Bun01b]. exponents [Gal00, JJK05, LCPC04, SS07a]. express [YNZ^{+}09]. Expression [TS08]. expressions [GME06, Pog05]. Extended [Huk02, Wu10, Yan02, YWYF09, Cha04, LVV06, LF02b, Nap09, NFH06, Ots01, Str01a, TNI^{+}07, Yan03c, FIT03]. Extending [BHL02, FdO09]. extensible [LAMH06, RSD01, CD01b]. Extension [ATIO06, SR01b, TV07, Dan07, DDdMS02, GWK09, KBC^{+}09, Mah09b, IW02]. extensive [EFG^{+}00]. external [BGH^{+}09a, DKC08, FHR^{+}05, FV02, JTS^{+}06, KKK06, KSS02, KDSB04, SSLN02, TV07, TL09]. externally [LLPL08]. extra [Cre00]. Extraction [HG02b, GBA01, OGG07, vHLP08]. extrapolation [dDSFY04]. extrema [Nov02]. Extreme [RRCV09, DM07, YM03]. extremely [LOCJ05]. Extremes [Sor02].

F [Sha04, RDSS01a, HD04]. F-like [HD04]. f1 [CG04]. fabrics [RGR^{+}04]. FaCE [TND05, WN01, TND04]. facility [VSBD00]. factor [DHS00, Esi01, Kon02, SKH02a, VC08]. factorisation [MA00]. factorization [AKZ00, PDA06]. factorized [PSV00]. factors [FMG00, GME06, RS03, WD04]. Faddeev [TND05, LEG02, TND04]. falling [Aok01]. Families [MKS07, De 02]. family [CJC09]. Fan [Hon04]. far [CP00]. far-field [CP00]. Farm [BFL^{+}01, BNFM^{+}09]. FARM 2DRMP [BNFM^{+}09]. farms [ABC^{+}01]. farside [Cha07]. FarSight [SEC04b]. Fast [ABRS12, BDH^{+}02, BH01, Bun01b, Bun01a, DSC06, GKK^{+}08, Ixa07b, MHS05, MS08b, RM05a, RTVZ08, Sul05, VKM^{+}05, Wei02b, WR01, YNS^{+}09, AC07, AH02, BB04b, Bru00a, CCGR09, CBMS08, CD04, EKW09, ES09, HC00, HJZ09, JK08, Kos05, Lad09, LC08a, LZC^{+}08, MP04, MG08b, MOC03, MM05, OMC00, OD08, OCK^{+}03, Oli01, PMA^{+}04, SOYN01, vHK00]. fast-switching [OD08]. Faster [DS01, HTNFBS06a, HTNFBS06b, Mas05].

FASTERD [SVP09]. faults [YKK07]. FDCSUSYDecay [QxW07]. FDTD [Ram10, MGN07, NSKS01, Ram10, Ram12, RB00, WP00, Yan09]. Fe [KEL02, KNSY07b, YKK07]. feature [MSY07]. features [HKL^{+}07b, KSS02, OS00b]. Fedosov [Tos08]. feed [TJLR06]. feed-down [TJLR06]. FEM [BP08a, RB00, WHL^{+}07]. femtosecond [Kur02]. fence [LC08a]. FERM3D [Ton07]. Fermi [BBR03, BFL^{+}01, GZDA01, MKK05, Moh07, NM01a]. fermion [ABM03, ASF^{+}05, AAC^{+}06, BCCM03, BBC^{+}01b, BPP01, CAF^{+}03, DDRW03, GGL03, JWW00b, KJ04, SW00b, Sol01, TA00a, TA00b]. fermions [BDF^{+}08, CMK^{+}03, CKLS09, DC05b, HK02, Jan05, Kol03, KALC08, MHK^{+}05, OGKL02, PS08]. ferromagnet [LTA05, RLU01, dSL02]. few [DM07, Sav01, Var08]. few-body [Sav01, Var08]. FeynArts [Hah01, HS02]. FeynEdit [HL08a]. FeynHiggs [HHH^{+}09, HHW00]. Feynman [Bek06, BT04, BCKT09, CD09b, FK00, GKR07, Hah01, HL08a, HvHHM09, KKK06, KNU00, RDFF02, Sem09, ST09, TF00, TF04]. FeynRules [CD09b]. FFL [SLL07]. FFT [BVY05, CL08a, FDM07, GBD03, MK08, iSHS^{+}08, Tak03, Yam00]. FFT-split-operator [MK08]. fgh [Sea02b]. fiber [Ger07, MTLC01, MCC05, MBC^{+}09, NSKS01]. Fickian [CL03, Ram05]. fidelity [Ker02]. fiducial [LVH07]. field [ACK05, AJ08, Bae03, Bae04, BD00, BS00a, BBB^{+}04, BK05a, BCD^{+}07, BMSG01, BKB02a, CP00, CSC^{+}07, CSC^{+}08, CGG^{+}08, CGG^{+}09, DCJ07, DKC08, DW01, DKV00, EKW09, FV02, GSM^{+}03, Goe02, GRS06, HGH^{+}05, HCH^{+}06, IH01, JTS^{+}06, JG02, KSS02, KF05b, KDSB04, KK04, KMR^{+}09, LC08b, LWY01, MR05, MSS^{+}09, MFVJ07, NSKS01, NT04, NW02b, NM01b, iNKNV08, PPC07, Pee07, PD08, Pit05, RLRR06, RMMP02, Sem09, SGF04, TL09, VCCS05, WGY01, WTH^{+}04, WRC^{+}04, Wil09, XD08, YSM09, YW01, Zha01, vHLP08]. field-theory [Pee07]. fields [ASF^{+}05, CM02b, Cre00, EL06, Fel08, FT08, FSK04, GLHW01, KDW00, Mel01, PM00, SG06, SYN01, Ver04, Yak01, ZKASS05]. FIESTA [ST09]. FiEstAS [Asc08]. Figure [Ano09r, Cap05]. filaments [LLPL08]. Files [Ano07-31, BBB^{+}00, KN07a]. filled [LPRS02, LPR04]. film [LTG09, LTT09, SLWH02]. FILMPAR [LTG09]. films [Dan05b, Mül02]. Filter [LB09, CNFR01, GKK^{+}08, SF06]. filtered [GBM02]. Filtering [LB09, AAA^{+}00, SA09, ZZH09]. filters [CSS^{+}03, KM08a]. final [BBB^{+}09a, JWW00b, SVP09]. finance [Sor02]. financial [KKH07]. find [HS01b, LJY07, Ort00, Yan03d]. Finder [CJT06, GJT03]. Finding [JS08, DM07, FD03, GKI02, GKI04, Hon04, Jam00, KKS04, bLpL02, LL04, pLbL03, RDFF02]. fine [AJ08, FMD07]. Finite [HL05, MZB^{+}04, NN06, PKSF01, RLV^{+}08, Zha00, BLS09b, BTI01, BJ08, CN01, Cha04, CCL08, DGAG06, DB08, EMJH03a, EMJH03b, EFS^{+}08, Flo01, GVMW04, GMAHV^{+}09, GDAG05a, GDAG05b, HBW05, Inu07, JBBR01, KMD^{+}02, KMP09, LLCS01, LMC^{+}03, MTJ02, MOS00, MOS01, OLS^{+}01, RP02, Rob01, SLMS06, SHX02, SS02b, Ton07, VBFD01, WHJ06, Whi00, XSC09, KSC^{+}00].

Finite-difference [HL05, NN06, BTI01, Inu07, SHX02, KSC^{+}00]. Finite-element [PKSF01, OLS^{+}01]. finite-element/molecular-dynamics/electronic-density-functional [OLS^{+}01]. finite-level [DB08]. finite-size [BJ08, DGAG06, HBW05, RP02]. finite-size-particle [VBFD01]. Finite-temperature [Zha00, KMD^{+}02]. finite-volume [Cha04, MOS00, MOS01, SLMS06]. finitely [SFSH01]. fireball [Tom09, KFJ^{+}09]. First [Ano09a, CM02a, Har01, KKKC07, LN01, RG05, Tsa02, AS03, AJT^{+}07, ABC^{+}01, ADDdM07, CR05, CBBJ02, CTI07, CGVA09b, EYJ07, FG04, GBTM07, IVD03, JPS^{+}01b, LDZ^{+}08, LA09, MCBR03, MSK^{+}05, Mor01, NKSL05, SBCZ08, WKP^{+}01, WC00, ZWD05, dSdSW08, vdHBP^{+}02, SZ04]. first-order [CBBJ02, JPS^{+}01b, LA09, ZWD05]. First-principle [Tsa02]. First-principles [Ano09a, KKKC07, AJT^{+}07, CR05, CTI07, EYJ07, FG04, GBTM07, LDZ^{+}08, MCBR03, Mor01, WKP^{+}01]. Fischer [Bur01, Hib01]. fission [DTD^{+}02, VT00a]. fit [ATP01, GKK^{+}08, TAP01]. fitted [ASVA00, CFMR08, FSW08, Fra07b, IVD03, KMS09, MKS07, PAS09, Sim00, SW00a, SVA01, SVA03, Sim08, SWFL00, Van06, VIV01, WC05, Wan06a, Wan06c, Wan06b]. Fitting [CCBL02, MYC09, Bla00, Bru00a, FGMT02, KJ07, MKJ^{+}05, Nap09, NP01b, Sim09, SF00, VAMVR08, vHLP08]. Fittino [BDW06]. five [IIK^{+}08]. five-dimensional [IIK^{+}08]. fixing [CM03, OS04]. FLAC [CGG00]. FLAPW [CMF00, FSB09]. Flash [LL07]. flat [BK05a, SLL07]. flat-plane [BK05a]. flattening [MTLC01]. Flavor [DGS08, CS02, CAF^{+}03, Lüs05, Mah09b, Mah09a]. flavors [KL01]. Flexible [Tót06, BCC^{+}06, HSS^{+}08, IW01, IW02, SWL09, SJF07]. floating [HDG07]. floating-point [HDG07]. floppy [Bac00, OBG09]. Floquet [KS04a]. Flow [PPC07, PK01, BBD00, CTG01, DVG05, DM09, GMAN^{+}07, JKKT00, JOS07, Ker02, KLD04, LTG09, MP01a, Mar08, MDT03, MC08, MLF07, NHS07, NYH04, PPM04, RSMK^{+}00, SG04b, Wal03, dNKM07]. flows [COE^{+}05, FH00, Huj05, IKO00, KITK00, KTG04a, ML06, RMVQ07, SR01a, Sus01, TFM09, TIM07, TIM08, TdFK00, TIM08]. fluctuating [DSL09, ICO01]. fluctuation [ICO03]. Fluctuations [LMS05, SSH02, HS01a, PB09b, TJLR06]. Fluid [ASC^{+}05, KF05a, CLL^{+}07, CMD00, DGV08, DUX^{+}09, FFF01, GIME02, GFP00, HKLY07, HHWH07, IKO00, ICO01, JOS07, KT07, LCS07, LCM00, LC00, LS05, MY00a, MY00b, MDT03, MC08, NYH04, PCK00, SLL07, SBCZ08, STK^{+}00, TMTF00, TK08, VKPB09, WGS00, WS02, Xia01, dNKM07]. fluid-solid [FFF01]. fluidic [BMS^{+}09]. fluids [DPB01, DGR09, HAA07, HCO01, Ida00, Ida03a, Ida03b, LNC^{+}03, SWY01, Tod01, WRMG05, YW00]. fluorescence [BG01]. fluorescent [SLL07]. flux [PCC^{+}09, Pet04]. FLY [ADBF03, BAD01, BCAD06, BADC07]. FMM [HJZ09]. FOAM [JS07, Jad00, Jad03]. Fock [BD05, DO04, DO05, DSC^{+}09, MW01, SDNR05, DD00, Dol01, GLL^{+}02, GG00, NM03, PS08, PRBD09, REAB08, SS09a]. focusing [HW09, SBBM04]. foil [BDV04]. Fokker [ABSM04, CBKM01, KA04, yMS01]. folders [BDH^{+}02]. folding [Elb05, Oka01, SSA07, WL08]. following [AAG^{+}04]. Force [TKN^{+}08, AL08a, ACC09, BK05c, CFJ09, EL06, Goe02, LZS06, MFVJ07, iNKNV08, RMMP02, SWC^{+}03, SWY01, VCCS05, YW00]. force-decomposition [SWC^{+}03]. force-field [MFVJ07, iNKNV08]. forced [SOYN01]. forces [HG02b, JKKT00, LZS08, LEG02, MK09]. forcing [AA08, Yao09]. Forest [OMF02]. Foreword [Ano01z, ME00, Sco09]. form [CRUV00, CHS09, Esi01, Lei02, Mah08a, RS03, She08, TNY00, FK00, TV07, MU06]. form-factor [Esi01]. formal [Oli01]. formalism [EE02, MM08, PTL04]. format [Ano07-31]. Formation [SCO00, BNSY02, FS01b, GB05, HOI04, KK00, MLPT08, PSK01a, PSK01b, QTL06, RRCV09, Rou01, SBD^{+}05, SHJ07, Voi02, Voi03, Yos09]. FormCalc [HS02, Hah08]. FormCalc-generated [Hah08]. formed [BSB02]. forming [GGL^{+}02, LMM^{+}08, MDH04]. forms [BKKS09, CCFG05]. formula [Inu07, MA04, MA08, Pom06, TI01, ZWD05]. formulae [Sim08]. formulas [CBMS08, NP00]. formulation [AK03, GPT08, IBA00, Leh00, YW01]. formulations [Ram12]. forsterite [LDZ^{+}08]. FORTRAN [BRdAHK04a, Hor09, KSYE00, Str05, BDH^{+}05, DG08, DGS08, Dem03, Dem06, DKM07, EH07, KLM00, MMEH08, MDM05, MA09, PS08, QRH00, Rib02, SS09a, Sar00, SPF00, TS06, vH06, vH07]. FORTRAN-90 [BRdAHK04a]. Forward [SGM^{+}09, CRS09]. Forward-backward [SGM^{+}09]. foundation [VSBD00]. Four [BCCM03, KA04, YN05a, ABM03, Bac00, BR01, BPP01, DWZS05, DDRW03, Gol00, KM00a, KM01b, MGG05, NN09, SW00b, TSA^{+}03, VT00c]. four-atom [Bac00, Gol00, MGG05]. four-body [VT00c]. Four-fermion [BCCM03, BPP01, DDRW03, SW00b]. Four-index [YN05a]. four-momentum [KM00a, KM01b]. four-point [NN09]. four-step [DWZS05]. Fourier [SVMT00, CN00, DSC06, Eli05, HC00, JP09, LC08a, MM05, NJ01, RM05a, SA09, Trö08, Wan06a, Wan06c, YZW02]. Fourier-based [MM05]. Fourth [ACK05, LJ09a, MKS07, UNK12, Van05a, Van06]. fourth-degree [UNK12]. Fourth-order [ACK05, Van05a, Van06]. FP [PWH^{+}00, TKN^{+}08]. FP- [TKN^{+}08]. FP-LAPW [PWH^{+}00]. FPGA [BCC^{+}08, EFS^{+}08]. FPGA-based [BCC^{+}08]. FPLAPW [ARV02, PAD07]. FPU [PKB^{+}01]. FracMAP [CGC^{+}09]. fractal [CGC^{+}09, GBR^{+}09, TdRGD09, Vor02]. fractal-like [CGC^{+}09]. fraction [TB85, Tho04b]. fractional [Dev05, DJ08, Sho04, Sho07]. fractions [Nap09, RMLB01]. fracture [EM08, KNSY07b]. fragmented [Tom09]. Framework [DN04, AGV00, Ano09t, BBB^{+}01, Che05, DKMF03, Dec07, DUX^{+}09, ES09, FFPW01, GPW^{+}09, ISSC01, OPB^{+}09, Tót06]. frameworks [Wel01]. Franck [GME06]. Fredholm [Str00]. Free [IH01, KSSH04, BBB^{+}09b, CHM^{+}09, HLC08, Ida00, Ida02, Ida03a, Ida03b, IKO00, OD08, SLL07, SJHY07, SR01b, SVMT00, SFR05, Wei04, ZBB^{+}06]. Free-boundary [KSSH04]. free-electron [SJHY07]. free-energy [BBB^{+}09b, IKO00]. freezing [Wil02]. Frenkel [KM05]. frequencies [FSW08, Kim03, Wan06b]. frequency [CIC^{+}03, Hei01, KCR07, LKKK07, Ram10, TC07, Wan06c]. FRET [SG04a]. FRETsg [SG04a]. Friction [CW02, KM01a, HOT07, HTM^{+}08, Müs02b, RR05, SS02b]. Frictional [KMB02, DHBE05, HKK^{+}01]. FRODO [AC05b]. Froese [Bur01, Hib01]. Front [Laf03]. frozen [NM03]. FRS [Kar01]. frustrated [Wes07]. Full [ABER00, GRS06, ADS06, BDLT02, Dür09, FS01a, FHW^{+}01, GSSN00, IIK^{+}08, Liu07b, LS05, Maz00, PAD07, PCV06, UTO09]. full- [IIK^{+}08]. full-band [PAD07]. Full-CI [ABER00]. full-electromagnetic [UTO09]. full-potential [ADS06]. full-wave [PCV06]. fullerenes [LB04]. Fully [Bac00, BSB02, ABM03, MA09, Sus01, Xia01]. function [Ada04, BFLW07, BDP00, CYAS05, CGM01, CG04, FKAM05, GT01, HG02b, ISSB01, JK02, Kar02, KHIL07, KBC^{+}09, MS08b, MNH01, New02, PGS02, RdAGV^{+}00, SK08, SA09, Str00, TGD06, TL06c, WDB04, Yan02, Yan03b, Yan03c, YW01, vdEFL^{+}02]. functional [BJ02, BSK^{+}03, CMK^{+}03, GH00, HHM^{+}09, HLC08, IBA00, KTT09, KH09, Kur02, LTG09, LCV06, Lor08, MBR01, OLS^{+}01, OSK^{+}02, SG05, SKNV01, SKNV05, SMK01, VKM^{+}05, dlGGS^{+}05]. functional/Monte [BJ02]. functionalisation [VEG08]. functionality [TV07]. functionals [HKP02]. Functions [GFS03, AS03, AA00, AA01b, AJ08, BP08a, BDBV12, BC05, BS04b, BW01, CD01a, CGA^{+}07, CGVA08, CGVA09a, DS04, FGMT02, FH04, FBB01, FPB08, FA00, GSS06, GLHW01, GZDA01, GDAG05a, GDAG05b, GME06, HTM01, Hua09, HM06b, HM08, Ixa01, JG02, KS05, Kim03, KSHP02, Kir06, KF03, KF05b, KTL05, KVR^{+}00, LdVJ06, LPC^{+}00, LKC06, LS01, Mic07, MS08b, MU06, MN07, MM09, MSHP02, MYL^{+}08, MSHP20, NM03, Nob04, PFG06a, PDM^{+}08, PAT^{+}09, RB08, Roy09, SKH02a, Sar00, Sau00, Sch06b, Sea02a, Sea02b, Sea02c, SMH^{+}01, SJF07, SFR05, TB85, TB87, Tho01, Tho04a, Tho04b, TL06b, VC08, WP00, WD04, Wei02c, YM03, Zah00, Zah01, ZBB^{+}06]. functions-calculation [GDAG05a, GDAG05b]. fusion [ASC^{+}05, BSW^{+}07, CBKM01, KMR^{+}09, OSK04, SEC04a, WML^{+}05, WSCW09]. Future [MSK^{+}02, Ano01a, Ano02a, McK07]. FV [WPL02]. FV-TD [WPL02].

g [ISH01]. G.R. [Leh00]. g permute [RLH^{+}09]. Ga [JK01, LK07, Tsa02]. GaAs [LVLS01, JK01, KFB01]. GaAs/Al [JK01]. GaGaRes [BvG02]. galaxy [RRCV09]. galaxy-sized [RRCV09]. Galerkin [FZ09, LS05, TKSR00]. Galilean [CK08, IKO00]. game [EFH^{+}07, VK09a]. GAMESS [BB00, FSBG00, KPD06, dMBC^{+}06]. GAMESS-US [KPD06]. gamma [Car06]. GaN [QASF^{+}05, Tsa02]. Ganga [Ano09s]. gap [BZ00, LLV^{+}01]. gas [BLS09b, BC00, CP00, CMD00, DHS00, GCP^{+}02, GF02c, Gut06, HCO00, HS01a, ID09, ICO03, KA04, KH06, KW03, LJ01, LNC^{+}03, MK02, NW02a, Nii00, PPC07, PCYC02, SCO00, SMV01, TNI^{+}07, Tsa02, TCO00, YB02b, Yok09, Zha01]. gas-phase [Tsa02]. gaseous [LR07]. gases [DSC06, IK00, Lon07, TS06, Wes07, WRN01, ZSSA00]. gastro [WG01]. gastro-intestinal [WG01]. gate [LLT^{+}02, LY05]. gated [KACB07]. gates [MSS00, RF05a]. GAUDI [BBB^{+}01]. Gauge [Dür05, Hei01, ALV05, ALN^{+}01, BB09a, CM03, Fod05, OS04, PM00, Tri05]. gauges [CMM09, DD01]. Gauss [AA01a, CFMR08, CP00, Del08, DSH02, DSH03, IP01, MS08b, Pom06]. GaussDal [ABNÅ05]. Gaussian [WLR^{+}08, CP00, FGMT02, FV02, Frü03, FKAM05, HKP02, MBR01, OS00b, SKH02a, SF06, VKM^{+}05, WD04, Wen01, You09]. Gaussian-core [WLR^{+}08]. Gaussian-mixture [Frü03]. Gaussian-sum [SF06]. Gaussian-type [FGMT02]. Gay [IW01]. GBL [DHS00, HS01a]. GDF [TGD06, TGD07]. Geant [Wel01]. GEANT4 [AGM^{+}00, AAB^{+}08, HFN03, YFM09]. gel [KKM02, SBD^{+}05]. gels [RvOvV02]. GeM [Che07]. GenAnneal [TL06a]. Gene [TS08]. General [CL08b, HLW05, HKL^{+}07b, ASS^{+}02, Cha00, FS00, FWP01, Gra02, HvDJvdM01, Jad00, Jad03, JJK05, KSPT04, MZB^{+}04, OK06a, SFR05, VF03a, VF03b, WTH^{+}04, ZF00, HKL^{+}07a, Pue06, Wen01]. general-purpose [ASS^{+}02, FS00, FWP01, Jad03]. Generalized [Ber02, GT01, GZDA01, LWY01, NSMO02, Alf05, BBR03, BMC05, CFJ09, CAW00, DJ08, GHP01, GMAHV^{+}09, HP02, KS05, KGM00, KHÖ01, LdVJ06, LL08, Mah08a, NM01a, NRR01, RMWH01, Sch06a, CCD07]. Generalized-ensemble [NSMO02]. generate [HHL06, KTT09, MVS05, SW00b]. generated [BB04b, Hah08, SI01]. Generating [Hah01, ON08, BRD04, BBJW05, BM04, CCRA05, Dev05, HTM01, MCLDP01, SJF07, vHK00]. Generation [CLFH07, RF05b, RF06b, AC05a, Bel05, BKM05, Boy09, CFJ09, Flo01, FIJ^{+}03, HSJ02, HvHHM09, IL07, JP09, Mas05, NP00, NYH04, Sem09, SGF04, SEF^{+}01, SMK01, TIN^{+}09, TNCG00, WHO02]. Generator [KRW03, Abe01, ABM03, ATB^{+}01, ACIZ07, AKS01, AKS02, BBB^{+}09a, BS04a, Bel01, BPP01, BvG02, CPW09, CDEW04, CWW06a, CWW06b, CWW07, GK05, Jad00, JWW00b, JPS^{+}01a, Jad03, JS07, Jun02, KFI^{+}01, Lad09, LMP^{+}09, MP06, Pap01, Pro00, QWWZ09, RS00, Sch06a, SVP09, SS02a, TA00a, TA00b, Tom09, TSA^{+}03, TKK^{+}06, CBMS08, KTBF06]. Generators [DGLB08, DH00, WH06, LBP^{+}09]. Generic [PJSK08, AAG^{+}04, Alv09, BBPS07a, BBPS07b, BBPS09, Di 01, KEM^{+}01, SDLW07]. Genetic [CB05, Bru00a, NP01b, Sug01, TL08a, WMNS09]. genetic-algorithm [WMNS09]. genetic-algorithm/simplex/spatial-grid [WMNS09]. Genetically [TL06b, TL06a]. GenMin [TL08a]. Gennes [MM04]. genomes [CHL05]. GENXICC [CWW07]. geodesic [FMMQ08, Rib02]. geomagnetic [HGH^{+}05]. geometration [KKM02]. Geometric [LWY01]. Geometrical [JS05, LANM^{+}01]. geometries [AH02, BM02b, Cha04, KEM^{+}01, LC07, MC08, SPV07]. Geometry [Sri01, BMML05, DC03, KSPT04, KS04b, Poi08, Poi09, Pop03, RJFB08, SGF04, SZ04, WTH^{+}04]. geophysics [MS05b]. George [Koc02]. GFACTOR2001 [Kon02]. GFCUBHEX [GLHW01]. Giant [ALN^{+}01]. GIAO [Dup01]. GIAO-SCF [Dup01]. GIBBS [BFL04, FFF01]. GiNaC [BD02]. Ginocchio [MS08b]. Ginzburg [BDHP08, CSCK08]. GIOD [BHNW01]. GITA [CRUV00]. given [BBJW05, KHÖ01]. GKW [PCC^{+}09]. Glass [GAR05, GGL^{+}02, BKB02a, BKB02b, CPT^{+}01, HG02a, RLU00, SH06, VKN07, YD07]. Glass-forming [GGL^{+}02]. glasses [Kat02, You02, You05]. Glassy [dO09, RR05]. GLauber [BRB09, AIOST03, BG06]. GLISSANDO [BRB09]. Global [ATP01, MTJ02, NV09, Roy09, TAP01, WTH^{+}04, AA07, BJ03, BGS^{+}04, BP08b, DR09, IIK^{+}08, JBBR01, JBA^{+}07, KPF03, LPC^{+}00, OS04, Swi04, TBZ12, TLP04, TL06b, TL08a, VPP^{+}12]. global-scale [Swi04]. globally [Maa06, BHNW01]. GLoBES [HLW05, HKL^{+}07a, HKL^{+}07b]. GLUE [RTS01]. gluon [KKK06]. gluons [KKK06, KMP09]. Gmat [CNMC09]. GMIC [OGWH03]. Go [Gra02]. Going [Jan05]. Golay [MMMM00]. gold [SPV07]. golem95 [BGH^{+}09a]. good [LCPC04]. Gordan [BRD04, CRW09, Dra01, KW08, RF06b]. Gordon [KA09]. Gourmet [Koc02]. GPUs [YNS^{+}09]. GR [TSA^{+}03, TKK^{+}06]. GRACE [BBB^{+}00, FIJ^{+}03]. GRACE/SUSY [FIJ^{+}03]. gradient [GHP01, KLD04]. gradients [BSO^{+}04, Dol01]. GRADSPH [VKPB09]. Graduate [Ano04a]. grain [Kar02, LMS^{+}02]. grained [ASS^{+}02, BLS09a, EL06, FMD07, FAiTD01, LS09, MS09]. grains [BM02b]. grammatical [TGD06]. Grand [Bae04, BNO^{+}01, JBS08, Zim05]. Grant [Kar01, Bur01]. granular [ACC09, BDK^{+}06, DHBE05, HM00, KBBW02, Rap02b]. granularity [LCB^{+}00]. GRAPE [Abe01, Mak01]. GRAPE-Dilepton [Abe01]. graph [PR06, Zim02]. graphical [AGS07, BT04, BCKT09, KJ07]. Graphics [LSVMW08, MCLDP01, CDD08, Hor09, MBKJ09]. graphs [BBJW05, HLB06, JPS^{+}01b]. grasp2K [JHFG07]. GRASP92 [FGR06, PFG06b]. Graßmann [AF05]. grating [MTLC01]. gravitating [CD05, VKPB09]. Gravitational [HBRS05, Leh00, ABC^{+}03, MMTH04]. grc [KFI^{+}01]. Green [GBM02, GLHW01, KBC^{+}09, KF03, KF05b, MNH01, RdAGV^{+}00, WP00, YW01]. Grid [HKM^{+}07, KL07a, Shi09, AAKL07, BLCR05, BS08, CSZ^{+}07, Eli08, GHPS04, IF03, Jad00, KTT09, MMTH04, MSHP02, MSHP20, OK09, Sak07, iSAK^{+}08, Sch06b, SEC04a, SFF^{+}04, WMNS09, vdHKM08, Ano09s, KKHL07, Shi07]. grid-adaptive [vdHKM08]. grid-based [CSZ^{+}07]. grid-size [BS08]. Gridless [VBFD01]. grids [CSC^{+}04, ID09, KSC^{+}00, KNT08, ISX05, SKNV01, SKNV05, SMH^{+}01, SD07, TCY^{+}08, WPL02]. Griffiths [DKC08]. GROMOS96 [BMvG00]. Gross [CPS00, DC07, MA09, TQZM08, TS06]. Ground [BH03, DC07, YN05b, BM06, CWSH08, DCNDC09, FV02, HG02a, LR07, WL08]. group [Alv09, CC04, FLO06, MI05, RF04, WN01, YT01a, Zit09]. Grouping [OGWH03]. groups [Goc04, RF05b, RF06b]. Growth [BM01, AFP02, BSvdDW02, Dan05b, MABK02, NSYZ02, RIB01]. Growth06 v2 [Dan09b]. GSA [RTS01]. GSM [BP08a]. GTC [EL04]. GTNEUT [Man04]. GTOBAS [FGMT02]. gTybalt [Wei04]. guarantees [TYN02]. Guide [BCP04, Koc02, Bar04]. guides [CLFH07]. Gustavson [CRUV00]. GUT [EH07]. GW [FER^{+}07b, GGG01, MSB09, SRR^{+}00]. gyro [PCC^{+}09]. gyro-kinetic [PCC^{+}09]. gyroid [HHCC05]. Gyrokinetic [KLD04, EL04, IIK^{+}08, Lee04, PCK00, RLI07]. gyrokinetic-ion [PCK00]. gyrokinetics [BH05]. gyrophase [SAU^{+}04]. GYutsis [VF03a].

H [EVL00, Hon04, Koc02, Laf03, Gro01, LKPH08]. H-VLPL [LKPH08]. hadron [Bar00]. Hadronic [Wel01, BEM^{+}02, CDEW04, CWW06b, CWW07, HS03]. hadronization [TSB^{+}05]. hadroproduction [CWW06a, QWWZ09]. half [HM08]. half-integer [HM08]. Hall [ADG08, CSC^{+}04, KB04]. Hall-MHD [ADG08]. Hallen [SR05]. haloes [GB05]. HAM [RE09]. Hamiltonian [HL00c, KTT02, MKS07, WZHZ06, ZF00]. Hamiltonians [MRS04, RLRR06, UCG^{+}05]. hand [BV00]. Handling [ADD^{+}03, BGLLW01, Sfi01, SS09b]. handy [YFM09]. Hankel [SSP08a, SSP08b]. Hanning [CL08a]. Hans [Hoo04]. Hansen [Vak00]. HAQMC [GPW^{+}09]. hard [MCC05, RM05b, SSLN02, TNI^{+}07]. hard-core [RM05b]. hard-point [TNI^{+}07]. hardly [Huk02]. Hardware [SEE^{+}03, ATB^{+}01, AAG^{+}04, CMR01, SI01, GPW^{+}09, LBP^{+}09]. Hardware-Accelerated [GPW^{+}09]. HARES [WKP^{+}01]. harmonic [Bek06, BFL04, Blü04, Blü09, CIC^{+}03, CHS09, DD00, DO04, DO05, DSC^{+}09, GR01, GR02, GME06, HL08b, HZGZ09, LDZ^{+}08, Maî06, PS08, SDNR05, TS06, You09]. harmonic-oscillator [DD00, DO04, DO05, DSC^{+}09]. harmonics [Bal07, CRW09, CP00, GS01b, IFF01, San00, SG00b]. Harrison [Wan00]. Hartree [BD05, DD00, DO04, DO05, DSC^{+}09, SDNR05, CWSH08, Dol01, GLL^{+}02, GG00, NM03, PS08, REAB08, SS09a]. hash [ZBB^{+}06]. HASPRNG [LBP^{+}09]. Hausdorff [WS09b]. hbook [Pöt00]. HBrowse [BBB^{+}00]. HCP [GLHW01]. HDF5 [SC04]. HDMR [KKS04]. He-like [CWSH08]. healing [MCL05]. health [SJDC07]. heart [ZS07]. Heat [TNI^{+}07, BBBR04, JJHvO03, Liu07a, RE09, SGK09]. heating [BBB^{+}04, BBBR04, DBR^{+}02]. Heavy [LMP^{+}09, Bar04, Cha07, CWW07, CS02, DGSL09, KTBF06, OSK04, TA00a, TA00b]. heavy-IoN [KTBF06, Cha07]. HEDP [RG04]. Heisenberg [TBL02, dSL02]. Heitler [Frü03]. HELAC [KP00, CPW09]. helicity [KP00, MM08]. Helium [CC08, Yos03, RY00, Yos07]. helix [YD06]. helix-coil [YD06]. help [GGL^{+}02, JS06]. HemeLB [MC08]. HENP [BNO^{+}01]. hep [AAB^{+}07, BBB^{+}01, SS07b]. hep-ph [AAB^{+}07]. hep-ph/0411186 [AAB^{+}07]. HepMC [DH01]. HERA [AAM^{+}01, BFMH^{+}01]. Hermite [FH04, LdVJ06, Pom06]. Hermitian [BFLW07]. HERWIG [KRW03, CF02]. heteroepitaxial [MABK02]. heterogeneous [Ste05]. heterogeneously [MTC07]. heterojunction [LH03]. heteropolymers [BJ05a]. heterostructures [SSPM05, Dan07]. heuristic [VF03a, VF03b]. hexadecane [VMMB02]. hexadecane-CO [VMMB02]. hexagonal [Zah01]. HF [BFH05]. HFBRAD [BD05]. HFBTHO [SDNR05]. hfodd [DSC^{+}09, DD00, DO04, DO05]. HFSZEEMAN [AJ08]. Hierarchical [The05, Ano01n, CD04, Col07, Huj05, Ort00, SKNV05, TC06]. Higgs [DKM07, DKV00, EH06, EH07, HHH^{+}09, HHW00, LPC^{+}04, LCE^{+}09]. Higgs-boson [HHH^{+}09]. Higgs-field [DKV00]. High [BBC^{+}01a, BTK^{+}02, CD05, COE^{+}05, DH01, FGV01, GK05, HJ02, KAB^{+}00, KK04, New07, SS00, Sim08, SEF^{+}01, TS08, Tol02, Bae04, BVY05, BBB^{+}09a, BADC07, BMS^{+}09, BCG03, Che05, CBM^{+}05, Cra01, DWZS05, DR09, FS01a, GPW04, GCD06, GMO03, HGH^{+}05, ISSC01, JH09a, JC01, Ker02, KKF^{+}04, LRI^{+}06, LVV09, LLT^{+}02, LKKK07, LDZ^{+}08, Mam08, MMTH04, MC08, NV09, PPM04, PPC07, PD08, PDM^{+}08, Ros04, Sak07, SBM09b, iSAK^{+}08, SLMS06, SVA01, SKRK04, TAM04, WW05, vHK00]. high- [LLT^{+}02]. high-accurate [WW05]. High-dimensional [BTK^{+}02, vHK00]. high-efficient [WW05]. high-energy [Che05, Cra01, GMO03, SKRK04]. High-energy-physics [SEF^{+}01]. high-fidelity [Ker02]. high-frequency [LKKK07]. high-intensity [PD08]. High-order [SS00, Sim08, DR09, JH09a]. high-performance [PPM04]. high-pressure [LDZ^{+}08, PDM^{+}08]. high-resolution [BVY05, KKF^{+}04, Ros04]. high-speed [GCD06]. High-temperature [HJ02, LDZ^{+}08]. Higher [SR09, LVV04, MA04, MA08, NGE^{+}04, TYSH05]. higher-energy [TYSH05]. higher-order [MA04, MA08]. highest [De 02]. Highly [CWSH08, Zak06, AT09, FBB01, KSYE00, Sau00, SŽ00a]. Highly-accurate [CWSH08]. Hilbert [KSTL03, KSHP02]. Hilbert-transform [KSTL03]. Hilliard [KEM^{+}01]. Hirvensalo [Vio04]. histogram [HF06, SR01b, VMMB02, YD06, dO02, dSL02]. histogram-reweighting [VMMB02]. histogrammed [BBBD06]. histograms [BH08]. history [MHS05, Wro08, YFM09]. HIV [CRPC08]. HIV-1 [CRPC08]. HMC [Lüs05, MSS^{+}07, UJSW06]. HMTA [GW01a]. Hogg [TSI02]. Holden [Laf03]. hole [BFI^{+}00, HBRS05, RRRHD08]. holes [CGCS07, Leh00]. Holistic [Rob01]. hollow [KHIL07]. hologram [PCA^{+}07, SI01, TIN^{+}09]. holograms [JP09]. Holographic [iSAK^{+}08]. holography [SMS^{+}00, SHI02]. HoloTrap [PCA^{+}07]. homogeneous [BBJW05, CGG^{+}08, CGG^{+}09, GBM02, Gut06, GMO03, KW03, Ste05]. homopolymer [SWL09]. homotopy [SPS09]. HONEI [vDGM^{+}09]. Hopf [AF05]. HOPPET [SR09]. hopping [Bes02]. horizon [Shi09]. HORN [SMS^{+}00, SHI02]. HORN-3 [SMS^{+}00]. HORN-4 [SHI02]. Hoshen [TG00]. Hosting [ZC09, CSZ^{+}07]. hot [TAM04]. Houches [Ano07-31, Ano09u, Hah09]. HP [WL08]. HPL [Maî06]. HPLC [NP01b]. HRMC [OPO^{+}08]. Hubbard [AC07, CA09, FM00, GOG00, KGM00, LR07, Ots01]. Huberman [TSI02]. huge [TYSH05]. human [KACB07, KN07a]. hurricane [Mas00]. Hybrid [CSC^{+}04, CKA^{+}09, KSPT04, OLS^{+}01, OSK^{+}02, OPO^{+}08, RD05, SBL^{+}04, AKS01, Cun09, GBFS07, KPF03, LKPH08, PCK00, RB00, SW00a, SvAS01, Swi04, TAM04, Tak00, TWY09, WMNS09, WBDB04]. hybrid-Darwin [TAM04]. hybridization [VPP^{+}12]. hybrids [KPF03]. HYDJET [LMP^{+}09]. hydration [DELG05]. hydro [MNV00]. hydrocarbons [EYJ07].

Hydrodynamic [MPS09, LKPH08, LLPL08, MC09, PM01, Xia01, YNK05]. Hydrodynamical [JKKT00, NBPG08]. hydrodynamics [JJHvO03, MY00b, RtVR09, Ska05, SD07, TE05, VKPB09, LMP^{+}09]. hydrogen [Bac02, CGG^{+}08, CGG^{+}09, HPC05, Jia08, KRTZ02, KF03, MOC03, PDM^{+}08, SKF05]. hydrogen-bonded [Bac02]. hydrogen-like [CGG^{+}08, CGG^{+}09, KF03, SKF05]. hydrogenic [Dy09, HB05, LS01, Sar00]. hydroxylated [RJCH00]. Hylleraas [PAT^{+}09]. Hylleraas-type [PAT^{+}09]. hyper [SNS01]. hyper-molecular [SNS01]. Hyperbolic [Laf03, BGH04, GT01, JH09a, PMG07, SS07a]. hyperbolic-function [GT01]. Hyperfine [GSF06, AJ08]. hypergeometric [CR08, CGM01, CG04, HM06b, HM08, MS08b, NP00, Wen01]. hyperplanar [BJ05b]. hypersonic [KTG04a]. hyperspherical [APV00, CGA^{+}07, CGVA08, CGVA09a]. hypervelocity [VKN07]. HypExp [HM06b, HM08]. Hysteretic [LA09].

I/O [Hah09, OCK^{+}00]. Ian [Bur01, Kar01]. IANUS [BCC^{+}08]. ICP [LCS07]. ICRF [IDS^{+}04]. ICRH [PHKL02]. Ideal [ATF^{+}09, CLR08, SIH^{+}01, IK00, Ros04, SHW01, WBC^{+}07]. identification [TdRGD09]. identity [EG09, KHÖ01]. ignition [VSBD00]. II [Ida03a, ABV02, BGLLW01, CCFG05, CMT01, CHP04, GKP^{+}06, Ida03b, IW02, PBI07, PSK01b, RF06a, THM01, Var08, Yan03b]. II.1. [ATP01]. II.2. [TAP01]. III [CvdEF^{+}05, DD00, GFF01, RF07]. ill [RMWH01]. ill-posed [RMWH01]. illumination [OKS04]. Image [MP05, BR01, DCJ07, ML03, MD00, QTMH07, RSD01, XD08]. ImageJ [DGR09]. ImageJ-based [DGR09]. images [AAA^{+}00, AEEdR05, BK05a, GBR^{+}09, HF06]. imaginary [FH04]. imaging [BSS09, LZC^{+}08]. immersed [Den08]. immersion [MIM^{+}07]. Immiscible [HCO00, BC00]. immobilization [STK^{+}00]. Impact [HDG07, MHK^{+}05, MOC03, OMC00, SM06b, VKN07]. impacts [KPS^{+}01]. impedance [DEW00, Pet04]. imperfections [BVY05]. implantation [KPS^{+}01, MIM^{+}07]. implanted [BSO^{+}04]. Implementation [BP08a, Dol01, GGQ01, KBC^{+}09, LC01a, MS08a, WRC^{+}04, ASF^{+}05, Alv09, ARV02, BB00, BTK^{+}02, BSK^{+}03, BCCW03, BVKW02, CGIA07, DDdMS02, Dup01, DHBE05, DM07, EVL00, Eli05, EFS^{+}08, GZ07, HS02, KM01c, LR07, LEG02, Maî06, MM08, MP01a, OCK^{+}03, OD07, Oli01, OBG09, PPM04, PKKM02, PCYC02, QRH00, RJFB08, RS03, SS09a, TRGR08, Tak03, TG00, WV05, WHL05, Zha08]. Implementations [VCF^{+}04, Xia01, CC04, SK08, Yam00]. Implementing [Nil07b, PLPS08, Di 01]. Implicit [JH09a, ADG08, Cha04, CCD07, DPSG06, GKI04, LBPS09, ML06, SD07, VIV01, WRC^{+}04, XZ12]. Implicit-explicit [JH09a]. implicitly [Gal00]. implosions [WSCW09]. Importance [VPNW02, ZWD05, Zim05]. importing [SC04]. improve [MC09, Pro00]. Improved [CRS05, IDS^{+}04, KM01c, RY00, SHW01, YWLC04, Cip07, Cip08, Ida00, Ida03a, Ida03b, LCE^{+}09, MHK^{+}05, Nat09, Nat10, TL08b, UOM01].

Improvement [SLL07, WLH00, BH07, Dür05]. Improvements [Col07, dMBC^{+}06]. Improving [DS04, DGLB08, GL02, PWH^{+}00, SSZ01]. impurity [MP05, WK02]. IMT [MKK05]. InAs [LVLS01]. InAs/GaAs [LVLS01]. incidence [BB07]. including [Con04, FFD00, SKH02a, WD04]. inclusion [GMBC08]. incompressible [Ida00, Ida03a, Ida03b, ZHC00]. incorporating [NW02a, SOYN01]. incorporation [ON08]. increase [SLL01]. Increasing [AA01a]. independent [ASF^{+}05, DC07, KV08, Man02, MN01, SM02, Str05, VEG08, WS09a]. Index [Ano00b, Ano00c, Ano00d, Ano00e, Ano00f, Ano00g, Ano00h, Ano00i, Ano00j, Ano00k, Ano00l, Ano00m, Ano00n, Ano00o, Ano00p, Ano00q, Ano00r, Ano00s, Ano00t, Ano00u, Ano00v, Ano00w, Ano00x, Ano00y, Ano01b, Ano01c, Ano01-27, Ano01-28, Ano01-29, Ano05d, Ano05f, Ano05h, Ano05-49, Ano05-50, Ano05-52, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01d, Ano01e, Ano01f, Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano01-35, Ano01-36, Ano01-37, Ano01-38, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02b, Ano02c, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano02-29, Ano02-30, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03-36, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Ano03-42, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i]. index [Ano04j, Ano04k, Ano04l, Ano04-48, Ano04-49, Ano04-50, Ano04-51, Ano04-52, Ano04-53, Ano04-54, Ano04-55, Ano05a, Ano05b, Ano05c, Ano05e, Ano05g, Ano05i, Ano05-46, Ano05-47, Ano05-48, Ano05-51, Ano05-53, Ano06a, Ano06-28, DGAG06, GDAG05a, GDAG05b, LVV09, MG08b, YN05a]. indexing [BFMH^{+}01]. indices [Sal02]. INDO [SS09a]. induced [AP05, Eli08, EM08, LVLS02, MB05a, MLPT08, MSH01, NM01b, NY07, RS00, Wro08, XD08]. induction [CMT00, CMT01, LZS08]. inelastic [LDBG08, MFVJ07, RS00]. inertial [ASC^{+}05, WML^{+}05, WSCW09]. inexpensive [ATB^{+}01]. inference [Vég04]. infinite [GBM02, LIR^{+}06]. Influence [CBBJ02, TMN01, SHJ07, SGK09]. Information [DDEM00, CHL05, GDC01, Han00, ME00, OGG07, Ort00, Tót08]. Infrared [NI01]. infrastructure [KKHL07]. inhomogeneities [BLCR05, NM01b, VS01]. inhomogeneous [MM04, Yak01]. Initial [BRB09, IHAR09, ASVA00, Kos05, MVS05, PAS09, Rob00, Rob01, SVA01, WW05, Wan05b, Wan06b]. Initial-State [BRB09]. initial-value [ASVA00, SVA01, WW05, Wan05b]. initialization [BDBV12]. initiated [Bar00]. initio [BTS05, MKB02, Nik03, PCCD09, SHZ01, iTKST01, YKK07, ATP01, CCBL02, HG02b, Hin00, KFJ^{+}09, MHGV09, MSH01, NW02a, RPD^{+}05, TAP01, TMN01, BGH^{+}09b]. Injection [PBB^{+}04, SJCM04]. InN [QASF^{+}05]. Input [KN07a, FGA04, Wen01]. INS [RC04]. insecticide [IN02]. Insight [YNZ^{+}09]. inspired [CPS00]. instabilities [BZ00, Lüt04]. Instability [NHS07, BH05, CDQF07, CT00, DMR01, DMR02, MDH04, MTZ00, Nur04, Sus01]. instanton [MK02, RS00]. instanton-induced [RS00]. Instructions [Ano00z]. insulator [KGM00, YH02]. integer [HM06b, HM08]. integer-valued [HM06b]. Integrability [Par04, HSSA01]. integrable [Bru04, CTR00, KW07, OMF03, qX08]. Integral [ST09, SVMT00, AA08, Alf05, BMC05, CC08, CL08b, DM09, Dun05, Dup01, EG09, Gut06, Kas00, KM05, MI05, MKK05, Moh07, MG09a, NM01a, Pis00, RDFF02, SR05, SZ00c, Str00, TAKN02, YN05a, Yao09]. Integrals [CCGR09, PR06, Sle00, ADDdM07, BD02, Bek06, BBR03, BGH^{+}09a, BW08, Cza06, Del08, Dy09, GKR07, HB05, KCH00, KK06, Mam08, MR06, Moh08, NN09, OIKN02, SKH02a, SKF05, UK02a, WD04, ZF00, dDSFY04]. integrands [IP01, Kau03]. integrate [NKV03]. Integrated [Han00, KMCS01, KPF03, LR06, ZPB09]. Integrating [VC08, SG06]. Integration [Sau00, SGF03, Asc08, BDP00, FFF01, FBB01, Fra02, Hah05, Hah07, HDGM07, Ida03a, Ida03b, JH09a, Kau03, Kim03, KK06, LIR^{+}06, LWLL07, MBR01, Mel05, MDC09, MKS07, Nur04, OMF02, PSK01a, SS00, Sch04, SW00a, SVA01, Sim08, SFR05, THC^{+}07, UJSW06, Van05b, Wu10, YWYF09, Zah04, Zah05]. integrations [Sea02c, WDB04]. integrator [Tak00, Van06]. integrators [Fra07b, SS06]. intense [BK06a, CP00, FSK04, HW09, KS04a, LCB07, Zha01]. intensity [BH01, BD06, CP00, Dan05a, Dan05b, DS06, Dan07, PD08]. Inter [BFL^{+}01, EMJH03b, GDC01]. inter-dot [EMJH03b]. Inter-Process [BFL^{+}01, GDC01]. interacting [DDD^{+}01, PS08, Tat07]. Interaction [WN01, BK06a, BKM05, FFG02, HYY07, KPD06, KTG04b, KDSB04, LCB07, LS01, yMS01, MCL05, NHS07, RCG05, Sar00, Sav01, TZZ06, TEP00, iTKST01, Vie01, WSB04, WML^{+}05, WHL05, YRR07]. Interaction-round-a-face [WN01]. Interactions [Mah08a, AH02, CF02, HL08b, HJZ09, LZS06, LLPL08, MV04, MMEH08, PHF^{+}07, PD08, SSB04, SZ04, SS05, YNK05, Zak00a]. Interactive [PCA^{+}07, WCGL00, WSCW09, BCD^{+}01, CGC^{+}09, Gre07, MPK00, PKRK07, vdB08]. Interconnected [BHNW01]. Interface [FSBG00, ZHC00, BT04, BCKT09, BB03, Den08, Hah08, Hor09, KFJ^{+}09, Liu07a, MRF^{+}05, RLU01, Tam03, WMK09, dlRL09, MCLDP01]. interfaces [Den08, GGG01, Har01, KRW03, RJCH00, YW01]. interference [KM08a]. interferometric [ABC^{+}03]. interior [DELG05]. Interlayer [BNSY02, LNK01]. intermediate [AJ08, BCG03, CRPC08, CM02a, JG02]. intermolecular [KPD06]. internal [Goc04, HCO01]. International [BDL00, BJS00]. Internet [Chr00]. internucleoside [BSB02]. Interparticle [SWY01, YW00]. Interplay [Mül02]. interpolated [CL08a]. interpolating [BS04b]. interpolation [BW01, Hin00, HDG07, PSK01b, QTMH07, Str00, UNK12, Val05]. Interpretation [HSSA01]. interpreting [RC04]. intersecting [BR01]. intersection [BCH05, PL05]. interval [Con04, FKAM05, LIR^{+}06]. intervals [AS00, Bar02]. intestinal [WG01]. intracranial [OCS^{+}08]. Intradonor [JK01]. Introducing [HLC08]. introduction [SMS08]. intrusion [TWY09]. Invar [MGPM07, MGYP08]. invariance [HP02, HL00c, SAG^{+}02]. invariant [BB04a, CK08, IKO00, PSK01a]. invariants [MGYP08, PSH06]. inverse [BV00, CRUV00]. inversion [Don02, MHK02]. investigating [TQ03].

Investigation [ACC09, BDK^{+}06, WCG04, BKB02a, CM02a, HSS^{+}08, KM01d, MB05a, MKM02, SG05]. invocation [DBE^{+}04]. involving [AA08, Ida00, Ida03a, Ida03b, LS01, Sar00, Yao09]. IoN [KTBF06, Vie01, BF04, Bar00, BSO^{+}04, Cha07, CBKM01, HPC05, JGJ09, KLD04, LdG^{+}07, LMP^{+}09, MCBR03, MOC03, MS05a, MIM^{+}07, OMC00, OSK04, OKS04, PMA^{+}04, PCK00, QTL06, Rou01, Sch08, SNBB02, WCG04, WRN01, WBDB04, OMC00, MOC03]. Ion-atom [Vie01, OMC00, MOC03]. Ion-atom/ [MOC03]. ION-ATOM/NEON [OMC00]. ion-beam [OKS04]. ion-implanted [BSO^{+}04]. ionic [GCP^{+}02, NW02a, XD08]. ionisation [BS03]. ionization [Bar00, KB02, Kur02, MOC03, OMC00, RMK05]. ionized [Lon07, WK02]. ionosphere [KKSR04, SKRK04]. ionospheric [BCD^{+}07, Eli08]. IonRock [BSO^{+}04]. ions [CWSH08, GS01a, KF03, SJP05, SHJ07, SKF05, Wro08]. IPA [PJK00]. IR [SJHY07]. irradiated [CP00]. irradiation [OSK04, RTVZ08]. Irreducible [GRR01, De 02]. Irregular [Wen01]. irreversibility [KA05]. Irreversible [Sta00, LA09]. ISBN [Hoo04, Laf03, Par04, Sha04, Vio04, Wan00]. ISICS [Cip07, Cip08, Cip09]. ISICS2008 [Cip09]. Ising [BCBJ02, BMML05, BM06, CM02b, CHP04, FV02, HBMJ05, KM01c, LTA05, NH09, SS07a]. isobaric [BFL04]. isospin [Dev05, GFG^{+}06, Mah08b]. isothermal [BFL04, TE05]. isothermal-isobaric [BFL04]. isotopes [LC01b]. Isotropic [JBS08, JOS07]. Isotropic-isotropic [JBS08]. isotropy [Koz02]. ISSN [Hoo04]. issues [Lee04, RGR^{+}04]. iterated [Sch05]. Iteration [SZ00c]. iterations [CvdEF^{+}05]. Iterative [BK06b, JBBR01, BDW06, BFLW07, CL02, FS01a, Li03, LY05, MP01a, MGG05, RB08, SMZ05, WWF08, WRC^{+}04, ZSM05]. Iteratively [DSH02, DSH03]. ITG [BGS^{+}04]. IV [CKA^{+}09, DO04, IFF01, RF08]. IVPs [FSW08]. IX [PFG06a]. Ixaru [AA01b].

J [RP02]. J4HistoryKeeper [YFM09]. J90 [WLH00]. Jacobi [KHÖ01, Yan03b]. Jacobian [Yan02, Yan03c]. JADAMILU [BN07]. Jahn [GFG^{+}06]. January [BJS00]. Japan [Sak07, Yos00]. Java [Esq04, Chr00, Esq04, KM08a, MCLDP01, RSD01]. JaxoDraw [BT04, BCKT09]. JChainsAnalyser [DGR09]. Jet [CJT06, GJT03, HFN03]. JETs [LMP^{+}09]. JetViP [Pöt00]. JetWeb [BB03]. JJGEN [SJF07]. John [Ano04b, Ano04-45]. Jones [GAR05, IW01]. Josephson [Bor02, Gen01, KSTL03, MSS00]. jumps [Ryc05]. junction [Bor02, Gen01, MSS00].

K*Grid [KKHL07, HKM^{+}07]. K- [Cip07, Cip08, Cip09]. Kac [RDFF02]. Kadomtsev [LL08]. Kallman [DC00]. Kalman [CNFR01, GKK^{+}08]. KangaROO [ADD^{+}03]. Kansa [DTHL09]. KANTBP [CGVA09a, CGA^{+}07, CGVA08]. keep [Var02]. keeping [YFM09]. Kernel [Cra01, Kos05, MP04, SHT08]. Kerr [HBRS05]. KEWPIE [BAB04]. key [MHS05, VK09b]. Keystream [ÁMRP04, WLW04]. kicked [Bow02]. Killing [dSB00]. Kinase [iTKST01]. kind [WW05, WS09a, YM03]. Kinematical [Dan05b, BD06]. Kinetic [EFBP04, HKPL07, Lon07, MABK02, RIB01, RPD^{+}05, SMH04, AGJJ07, ASC^{+}05, BD08, BSDMH05, BDBV12, BGS^{+}04, DJ04, DGV08, DKC08, HOI04, HS07, HW09, IF03, KSPT04, MCL05, MRS04, MHS05, PCC^{+}09, PSK01b, SMH^{+}01]. kinetic-fluid [DGV08]. kinetic-MHD [KSPT04]. Kinetics [AFP02, NL07, NG02, Pur02, BDHP08, KA04, Nak08, Pin01]. Kirchhoff [Pis00]. KK [JWW00b]. Klein [KA09]. KLOE [Sfi01]. KMI [HKM^{+}07]. KMI-R1 [HKM^{+}07]. knockout [BG06]. knowledge [ME00]. knowledgebase [BDG^{+}08]. Kogut [CAF^{+}03]. Kohn [SJ02, WT01, AK03, MMR04, PAD07]. Kontorova [KM05]. Kopelman [TG00]. KoralW [JPS^{+}01b]. KoralW&YFSWW3 [JPS^{+}01b]. KORALZ [JWW00a]. Korea [Cho07, SS07b]. Korringa [SJ02]. Korteweg [KD09, Zak00b, ZY09]. Kranc [HHL06]. Kronrod [AA01a]. krypton [STK^{+}00]. KSTAR [KY07]. kT [JPS^{+}09]. KtJet [BCCW03]. Kutta [Fra02, KMS09, PAS09, Van05b, ASVA00, BT01, CFMR08, MVJ09, Van05a, Van05c, VIV01, VAMVR08].

L [Cip07, Cip08, Cip09, TKN^{+}08]. L- [Cip07, Cip08, Cip09]. L1Packv2 [Lor08]. L1PMA [Dem03]. L2CXCV [Dem06]. lab [LKPH08]. LabVIEW [GCD06]. ladders [CSW02]. lag [KS01, Liu07a, Van05c]. Lagrange [LCHJ09]. Lagrangian [BGS^{+}04, CRS09, Ida02, Mel05, ML06, TYN02, TFM09, UNK12]. Laguerre [Hua09]. Lakshmanan [Par04]. lamellar [LMS05]. LAMMPS [KBC^{+}09, LS09]. lamp [SLL07]. Lanczos [BSTC05, ÇHM00, GNZ^{+}09, MA06]. Landau [BR09, CM03, KS08, BDHP08, CSCK08, KS84, LWT08, LOL06, LWLL07, OS04, PRSB08, SWL09, WL08, YD07, Zha08]. Landau-gauge-fixing [CM03]. Landau-Transition-Matrix [BR09]. landing [KPS^{+}01]. landscapes [WLGX09]. Lane [SPS09]. Langdon [Ano04-56]. Langevin [DBR^{+}02, QP05]. Langmuir [CS07]. Langtangen [Hoo04]. language [Bor07]. languages [BDK^{+}06]. LanHEP [Sem09]. lanthanides [EÅ01]. Laplace [Don02]. LAPW [PWH^{+}00]. Large [DMR02, KMD^{+}02, LM02b, RvOvV02, TIM08, ULA^{+}02, Vor02, ABD^{+}05, BVY05, BN07, CLL^{+}07, DVG05, DBE^{+}04, DM07, ES09, FTGG07, HS01b, KKK06, KTT02, MC08, MFVJ07, MDC09, MAM04, MAM07, MHK^{+}05, MM01, iNKNV08, OCK^{+}00, OGWH03, PFG06b, QG04, REAB08, RTVZ08, RCG05, Sch06b, SKNV04, SJ02, Str05, TPYV03, TYSH05, TL09, WCGL00, WV04, TIM07]. large- [TL09]. large-area [CLL^{+}07]. large-eddy [DVG05]. Large-scale [KMD^{+}02, LM02b, ABD^{+}05, BVY05, DBE^{+}04, FTGG07, MDC09, iNKNV08, OCK^{+}00, PFG06b, REAB08, RCG05, SJ02, TPYV03]. Laser [BCP04, Gha05, BBB^{+}04, BK06a, BBBR04, BDV04, CP00, DBR^{+}02, EST00, FSK04, GFP00, HD04, HS07, HW09, KDSB04, Kur02, LdG^{+}07, LKPH08, LCB07, Mah08a, MK05, MLPT08, NY07, PD08, SJHY07, SSB04, SBBM04, VSBD00, WML^{+}05, Zha01]. laser-accelerated [LdG^{+}07]. laser-atom [BK06a]. laser-driven [MK05]. laser-induced [MLPT08]. laser-matter [LCB07]. laser-plasma [VSBD00, WML^{+}05]. laser-produced [GFP00]. laser-target [PD08]. Lattice [BS00b, Cre00, EFH^{+}07, Fod05, HL00a, KS04b, MS05a, Nii00, PY08, SHT08, Suc02, vdSvdG08, ALV05, All01, ALN^{+}01, BB09a, BBD00, BC00, CYAS05, CT00, CK08, CMD00, CAF^{+}03, CMM09, CNDC09, DCNDC09, DDD^{+}01, DS04, DPB01, Di 01, DHS00, DSL09, Fel08, FT08, FKP03, FJC^{+}05, GL02, GHLW03, HvHHM09, HCO00, HCO01, HS01a, HNG05, IK00, IKO00, ICO03, JU09, JH09b, KITK00, Koz02, KK05, LNC^{+}03, Luo00, Lüs04, Lüs05, Mas05, MHR^{+}07, MC08, MSS^{+}07, MG09b, OS04, OCS^{+}08, PCF05, PPM04, PCYC02, RS09, SCO00, SS07a, SBJ05, TMTF00, TCF00, Tri05, TCO00, TdFK00, Voi02, Voi03, YB02b, Yos01, ZHC00, ZY09, ZSSA00, vHLP08, KSC^{+}00]. Lattice-BGK [KSC^{+}00]. lattice-Boltzmann [DPB01, IK00, LNC^{+}03, MC08]. Lattice-gas [Nii00, BC00, LNC^{+}03, YB02b]. LATTICEEASY [FT08]. lattices [CC09, GMAHV^{+}09, RM05b, Wes07, ZZ09]. law [MR05, Mil06, Mil07, OML09, RDSS01a]. laws [Che07, RR05, ZY09, Laf03]. layer [BBBR04, CAW00, HJZL07, KY07, LCM00, Ras09, Ras17, SSPM05, Str01b, SBCZ08]. layer-multiple-scattering [SSPM05]. layered [DW01, Liu07a]. layers [ACC09]. Laying [VSBD00]. lazy [Bru04]. LBIE [DM09]. LC [JPS^{+}01a]. LCG [BDG^{+}08, Shi07]. leading [CC04]. leakage [CJC09]. leaky [Alf05]. Least [KT04, TD03, Dem06, JC07, WWF08]. Least-squared [KT04]. LED [CFJ09]. LEED [BH01]. LEED90 [BRdAHK04a]. Legendre [Del08, SSP08b, Str00]. legs [BGH^{+}09a]. Lekner [TZZ06]. Lemaître [Rib02]. length [MSS^{+}07]. Lennard [GAR05, IW01]. Lennard-Jones [GAR05, IW01]. LEP [JWW00a]. LEP/SLC [JWW00a]. LEP2 [BCCM03, JPS^{+}01a]. LEP2/LC [JPS^{+}01a]. lepton [JWW00a]. leptonic [PAT^{+}09]. leptoquark [Bel01]. less [AW04, WA07]. level [ABB^{+}09, CPW09, Cap05, DB08, ES09, GC01, HYY07, ISSC01, LLV^{+}01, RLRR06, RGD^{+}01, SV01, TY01, Yok09, YT01b, ABF^{+}01]. levels [CCBL02, CC07, CCL08, CGA^{+}07, CGVA08, CGVA09a, EÅ01, EVL00, GZF04, LVLS01, TYS^{+}00]. LevelScheme [Cap05]. Levin [RB05]. Levin-like [RB05]. LGT [Tri01]. LHC [Shi07, ADE^{+}02, CCG08, CGCS07, GGQ01, QWWZ09]. LHCb [Ano03h]. Li [BNS07]. libraries [vDGM^{+}09]. Library [Ano04-46, BJS00, BB09b, JKW00, JKW06, MM09, PSW00, Bel05, BEM^{+}02, Boy09, DVL^{+}02, DVL^{+}04, GBM02, Hah05, Hah07, JC08a, KS05, MG09c, PMA^{+}04, Pin01, SG00a, SM04, SM06a, SBM09a, SWS^{+}12, VHLP09, BJS00, DVL^{+}04]. libration [She08]. lidar [BK06b, OPB^{+}09]. Lie [dSB00, BCV03]. life [BM02a, Teh01]. Lifshitz [HP02]. lifting [MA00]. Light [PCC01, BLCR05, BDF^{+}08, CLFH07, FWP01, GDC01, Har00, HTL^{+}03, KS04a, LPRS02, LPR04]. light-cone [Har00]. lightrays [MG09c]. Lightweight [CSZ^{+}07]. like [CGC^{+}09, CWSH08, CGG^{+}08, CGG^{+}09, HD04, KF03, MVS05, OMF02, RB05, SBD^{+}06, SKF05, Wal03]. likelihood [BDYK04, Nap09]. LILIX [Ixa02]. Limit [DDFI09, HKK02a]. Limitations [FM00]. Limiter [SZ04]. limits [KJ07, KS04a, Sor02]. LINDEN [RGD^{+}01]. Line [CDD08, JK01, Mar01]. Line-by-line [CDD08]. Linear [ADS06, BK05c, CMM09, FG04, Gao03, HHM^{+}09, RLI07, SKNV01, WC00, YG09, Bat03, BMG01, BW01, Bru00a, CN01, CIC^{+}03, CCBL02, Cha00, FGF03, GSGT03, HZGZ09, KA09, Kol03, LRI^{+}06, NP01b, SKNV05, SSP08b, Wan05b, ZA01]. linear-mixing [Bat03]. linear-rigid-rotor [CCBL02]. Linear-scaling [Gao03, HHM^{+}09, SKNV01, WC00, SKNV05]. Linearization [Ram03]. linearizations [BB04a]. Linearized [BC05, ADS06, IH09, IHAR09]. linearly [CMR01, Man02]. linearly-scaling [CMR01]. lines [HD04]. lineshape [BDM09]. link [Dür09, KT04, KSYE00]. link-cell [KSYE00]. linkages [BSB02]. linked [RS09]. linking [BDYK04]. links [HK02]. Linux [BS06a]. Liouville [CGVA09b, LVV04, LVV09]. Liouvillian [ADDdM07]. lipid [SDLW07]. Lipkin [RGD^{+}01]. liquid [All05, BNS07, CAAM08, GLP03, JBA05, LS02, LPRS02, LPR04, MSS^{+}09, MVS05, MSK^{+}05, MDH04, Mor01, MSH01, PGS02, RCGC00, RCG05, SSH01, TDY02, Yok09, Yos03, Yos07]. liquid-liquid [MSS^{+}09, Mor01]. liquids [CAAM08, GGL^{+}02, HL00b, SHZ01]. List [Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09c, Ano09d, Mas05, MDT03, YWLC04]. lists [ABRS12, BM04, SJF07]. lit [CFJ09]. literature [DDEM00]. lithium [OMC00]. little [GGL^{+}02]. load [CD09a, PSP^{+}03]. load-balancing [PSP^{+}03]. Local [MP01b, YW01, AA07, ACK05, BP08b, DM09, HTM^{+}08, JK08, LWY01, LM02b, NM01b, Ryc05, SKH02b, TLP04, TL06c, VPP^{+}12]. localised [KH09, MYL^{+}08, MSHP20, RB08]. localization [CMRS02, TIM07, TIM08]. Localized [GFS03, MSHP02, SMH^{+}01]. Locating [TL06c, LLY07, VPK^{+}01, WMNS09]. location [HS01b, QTMH07]. log [KS05]. log-derivative [Jam00]. log-sine [KS05]. Logarithmic [Dür09, Zit09]. logging [ZZH09]. logging-while-drilling [ZZH09]. logic [MSS00]. Long [HKL^{+}07a, HKL^{+}07b, PHF^{+}07, WLR^{+}08, AL08a, Cai09, CJC09, HLW05, LOCJ05, LOY07, MBG03, MRS04, RD05, SYN01, SVA01, Sim08, SS05, Tat07]. long-baseline [HLW05]. Long-range [PHF^{+}07, AL08a, CJC09, LOY07, MBG03, MRS04, RD05, SYN01, SS05, Tat07]. long-term [SVA01]. Long-time [WLR^{+}08, Sim08]. long-wave [Cai09]. loop [BD02, BGH^{+}09a, Blü00, BW08, CCGR09, FK00, KKK06, KK06, LOY07, LR06, MR06, NN09, PZ01, PR06, WMNS09, dDSFY04]. loops [PZ01]. LORES [ZDKG05]. loss [Frü03]. Lotka [Sle00]. Low [BRdAHK04a, Cun09, IL07, ZDKG05, BDBV12, CRS05, CDQF07, CS07, DS01, FH00, FS01a, GSSN00, Kat02, KG07, MSY07, MYC09, NH09, Rob00, SM06b, Wan01, BRdAHK04b, dAK01]. low- [MSY07]. low-dimensional [MYC09, Rob00]. Low-Energy [BRdAHK04a, BRdAHK04b, dAK01]. Low-lying [Cun09]. low-noise [BDBV12]. Low-pressure [IL07, CS07]. lower [WBDB04]. lowest [Kol09]. LP [Ixa07b]. LPGS [ZA01]. LPGS-decomposed [ZA01]. LQGENEP [Bel01]. LS [Car06, Car07, GF02a]. LSGP [ZA01]. LSGP- [ZA01]. LSJ [GZF04]. lsjk [KS05]. LSS [BCAD06, BADC07]. lubrication [SR01a]. LUDWIG [DPB01]. luminous [SLL07]. Lyapunov [Gal00, LCPC04]. lying [CRS05, Cun09].

M [Par04, Cip07, Cip08, Cip09, MM08]. M-shell [Cip07, Cip08, Cip09]. M.Dyna [LL00]. M3D [CBF^{+}04]. Mach [FH00]. machine [CC00, WLH00]. machines [GBD03, Tri01]. MACHO [HCK01]. macro [OML09]. macromolecular [Bin02]. macromolecules [EE02, ZDKG05]. Macroscopic [RR05, GT04, Lud02]. Macsyma [Dra01]. made [CD09b, Hah09, PKB^{+}01]. mAgic [PKB^{+}01]. mAgic-FPU [PKB^{+}01]. magnetic [ACK05, BD00, BCD^{+}07, BMSG01, BJ03, CMT00, CMT01, CGG^{+}08, CGG^{+}09, EFBP04, EKW09, FS08, GBC^{+}04, HCO01, HOI04, IH01, JTS^{+}06, JBA^{+}07, KDSB04, KMR^{+}09, KS01, NM01b, Pis00, SHW01, SG06, SYN01, SPP^{+}04, TZZ06, UTO09, VEG08, YSM09, Yak01, Zha01]. magnetically [Ram12]. magnetics [MKM02]. magnetization [SMV01]. magnetized [DGV08, HL05, JTS^{+}06, KCR07, KT07, MMTH04, Ram10, Yan09]. magneto [DGR09, HSSA01]. magneto-oscillatory [HSSA01]. magneto-rheological [DGR09]. magnetofluid [vdHKM08]. magnetohydrodynamic [DM09, Liu07b, MS08a, Pet04, SG01, YSM09]. magnetohydrodynamics [Cha04, LBPS09, Zie05]. magnetoplasma [SAU^{+}04]. magnetosphere [Swi04]. magnetron [IH01]. magnets [CM02b]. magnon [Tam03]. Magnus [PSV00, WDHE04]. Magnus-factorized [PSV00]. main [Mas00]. Making [Sal02]. Malmberg [CKV04]. man [CFH^{+}01]. management [ABNÅ05, Ano09s, BCC^{+}06, dSdSW08]. manager [Teh01]. manifestation [HSSA01]. manifold [Bac02]. manipulation [Niu00]. manipulators [MP04]. Many [KM05, BD08, BM04, CW00, DD01, Dzu09, EKW09, GFF01, Mak01, MSD08, NP01a, OMF02, SBB03, YB02b]. Many-body [KM05, BD08, CW00, DD01, Dzu09, Mak01, NP01a, OMF02, YB02b]. many-dimensional [MSD08]. many-electron [BM04, EKW09]. many-particle [GFF01, SBB03]. Maple [AF05, BS03, CD01a, DDdMS02, FIBT01, FIT03, Fri09, GFF01, GF02a, GSF05, GSF06, GI09, Har00, IFF01, bLpL02, LL04, pLbL03, PFG06a, Vul03, qXbL04]. maplet [YC07]. Mapped [ABOSPG09]. mapping [CD08]. maps [Gal00]. market [LLH07]. markets [KKH07]. Markovian [FRdS09, JS06, MVS05]. marks [LVH07]. Mars [PAD^{+}09]. martensitic [KEL02]. MAS [BDM09]. masking [UOM01]. mass [BDF^{+}08, ISS^{+}02, Jan05, JU09, PR06, UJSW06, vHLP08]. masses [CKS00, EH06, EH07, HHW00, KJ04, NN09]. massive [ABM03, Ste01]. Massively [DMD^{+}07, Jen00, BTK^{+}02, CSS^{+}03, Dec07, Dup01, GBFS07, GBD03, KCC^{+}00, MT00, NJ00, Yos01]. massless [Bek06]. Master [CCGR09, FS01a, LOCJ05, OGWH03, PR06]. MATAD [Ste01]. Matched [CSS^{+}03]. matching [CAW00, Tam03]. material [Ano09a, BD00, BDK^{+}06, Lud02, MSY07, SBM02, dNKM07]. Materials [Haf07, MRF^{+}05, CM02a, Den08, Goc04, GMBC08, Gun02, HKK^{+}01, HM00, JGJ09, KCC^{+}00, MS05b, NY06, OLS^{+}01, Ram12, YGT^{+}02, dSdSW08]. Mathematica [BC07, Bre07, Cap05, CKS00, FW01, Fer07a, GKR07, GMBC08, GHIL09, Gro01, Hah08, HL08b, HM06b, HHL06, JDBT06, JDBT09, Lor08, Maî06, MM08, PTL04, Tos08, WGDZ04, WW06]. Mathematical [Bru00b, Del03, Koc02, KTG04a, MTZ00, RR02, Suz00, GT01, KVR^{+}00, KN07a, Mas00, SZ00b, Sev00, Tho01, Yan03c]. MATHSCOUT [BC07]. Matlab [HTNFBS06a, HTNFBS06b, PBI07, TNBSF04, Tót08, Sha04]. matrices [BRdAHK04b, BN07, Bun01a, De 02, DC00, Flo01, GRR01, LB04, PFG06a, Pog05, RF08]. Matrix [BR09, Di 01, IH09, Alv09, AC05a, AC05b, BFLW07, BMG01, BBB^{+}00, CNMC09, Cha00, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, ÇHM00, DEW00, DM07, EFS^{+}08, FM03, FSB09, GF01, GFF01, GME06, HL08b, KTL05, LS01, MN01, OS03, PDA06, PAT^{+}09, Ram12, Sar00, Sch08, SSB^{+}09, SMZ05, SNBB02, TYSH05, Ton07, UTKF05, WN01, Yak01, You09, Zat06, ZSM05, dAK01, dlGGS^{+}05]. matrix-exponential [Ram12]. matrix-vector [EFS^{+}08]. MatrixExp [Pog05]. matter [BBPS07a, BBPS07b, BBPS09, BFI^{+}00, CMR01, LAMH06, LCB07, PJSK08, RRCV09]. Maximally [GFS03, MYL^{+}08]. maximally-localised [MYL^{+}08]. maximum [BDYK04]. Maxwell [MOS01, CHS09, DKMF03, Den08, HZGZ09, LCB07, MOS00, PCV06, RB00, ST02, She03]. mazer [BS00a]. mazev1 [DG08]. mazev2 [DG08]. mazev3 [DG08]. mbar [KRTZ02]. MC [FFK02, GPW04, JPS^{+}09]. MC-TESTER [GPW04]. MC3D [Wol03]. MC3D-3D [Wol03]. MCC [LSL07, MIM^{+}07]. MCDB [BDG^{+}08]. MCDF [UK02b]. MCDHF [AJ08]. MCEF [DTD^{+}02]. McGuire [Koc02]. MCHF [Fro00, FTGG07, JG02, ZF09]. MCL [KPD06]. MCNP [KN07a]. MCNP-4B [KN07a]. MD [GSM^{+}03, KPS^{+}01, LL00, PS09, TMN01]. MDGRAPE [SEE^{+}03]. MDGRAPE-2 [SEE^{+}03]. MDVRY [SLBG09]. Mean [Tam03, BKB02a, GSM^{+}03, MSS^{+}09]. means [Ano01n, BKKS09, FH04, KKSR04, Tam03]. measurements [BSO^{+}04, FKMB09, PBI07, iSAK^{+}08]. Measuring [Yur02]. Mechanical [PMH08, CN01, Dun05, HSGBK08, LLT^{+}02, OSK^{+}02, VT00b, VT00c]. mechanical/molecular [OSK^{+}02]. mechanically [RJFB08]. mechanics [HF00, JC08b, KM01d, KM03, LC00, MGG05, OMF03, Ram03, Sta00, VCCS05, WDHE04]. mechanism [Ger07, KL07a]. mechanisms [HOT07, Ste05, iTKST01]. media [Bar00, BBD00, BB04b, CGK^{+}00, JOS07, Mel01, PPM04, RK05]. medical [FFS01]. medium [CL03, NN06]. meets [MS05b, OLX07]. Mellin [Blü00, BKKS09, Blü09, Cza06, GKR07]. melting [KNSY07a, LNLK01, MVS05, MYJY01, Ste05]. melts [CW01, MPS09, Ryc05, WBC^{+}07, ZM00]. membrane [GS01a, RR02]. membranes [HJM02, SDLW07, iTKST01]. memory [BOG^{+}07, BB00, BTS06, CSCK08, CC00, TG00, TYS^{+}00, Xia01]. MEMPSODE [VPP^{+}12]. MERADGEN [ACIZ07]. mercury [SLL07]. mercury-free [SLL07]. Merging [GB05]. MERLIN [KPD06, PDL04]. MERLIN-3.1.1 [PDL04]. MERLIN/MCL [KPD06]. Mesh [BTS06, IH01, KNTG03, PM01, BFH05, DMR02, FMD07, FS08, HCH^{+}06, KKF^{+}04, LHS^{+}06, MOM^{+}00, OD07, Ros04, SJCM04, VAH04, VCF^{+}04, WTW04, Zie04, Zie08]. mesh-refining [LHS^{+}06]. meshes [MCLDP01, MOS00, MOS01]. Meshless [DM09, YNS^{+}09]. mesogenic [HSS^{+}08]. meson [CDEW04, CWW06b, DS04]. mesophase [HHCC05]. Mesoscale [JOS07, LNC^{+}03]. Mesoscopic [PCF05, YGT^{+}02, HKK^{+}01, ISSB01, iOY01, Yos00]. mesospheric [BCP04]. Metacomputing [Lüt00]. Metadata [dSdSW08]. Metal [KGM00, LY05, NP01a, TGB01, YH02]. Metal-insulator [KGM00, YH02]. metal-oxide-semiconductor [LY05]. metallic [CIC^{+}03, KPS^{+}01, PDM^{+}08, SG05]. metals [Cle05, WC00]. metamodelling [RPY07]. metastable [CRS05, vdB08]. Method [IH01, SLL01, VPK^{+}01, AP04, AA01a, ABOSPG09, Alf09, ÁMRP04, ASF^{+}05, ADS06, AKZ00, ARV02, AH02, ASVA00, Bae03, BDK^{+}06, BDW06, BLS09b, BFLW07, BK06b, BKM05, BDYK04, BTS06, BVKW02, Cai09, CMF00, CZC00, CCL08, CWSH08, CAW00, CA09, CRS09, CCRA05, DWZS05, DS01, DM09, Den08, DW01, DSH02, DSH03, DSHH05, Don02, DTHL09, DDdMS02, Dys02, DM07, EL06, Eli08, EFS^{+}08, FSW08, FNR^{+}06, FNR^{+}07, FER^{+}07b, FSB09, FZ09, GVMW04, GHP01, GT01, GMAHV^{+}09, GMAN^{+}07, GLP03, GLMADB^{+}02, GSGT03, Gre07, bHhL07, Hin00, HJZ09, Hua09, IH09, Ida02, IK00, IKO00, Ixa07b, JH09a, Jam00, JC08b, KV08, KA09, KNY05, KD09, KM01d, KM06, KN07b, KTT02, LRI^{+}06, LLY07, LLCS01, Li03, LY05, LFT01, LFT03, LKC06, LZ04, LS05, Lüs04, Maa06, Man04, MDT03, MHR^{+}07]. method [MLF07, MS05a, Mel05, MN01, MYJY01, MI05, MA06, MP01b, MABK02, NT04, NM01a, NJ01, Nik03, NFS01b, OCK^{+}03, OD07, OS04, OPO^{+}08, OGWH03, PAS09, PJK00, PMG07, PAD07, PKST03, Pis00, Pit05, PSV00, RLRR06, Ram10, Ras09, RE09, Ras17, RB08, RDFF02, RMLB01, Ros04, RMWH01, RB00, SSPM05, San00, SNS01, Sch05, SOYN01, SI01, SHZ01, Sho04, Sho07, Sim00, SW00a, SVA01, SVA03, SPS09, SJ02, SZ00c, SR01b, SM02, SRR^{+}00, SFSL09, SHH^{+}04, SS05, SFR05, TMTF00, TCF00, TQZM08, TL06b, TdFK00, UOM01, UOTM03, UK02a, UK02b, UYK^{+}04, VPCK04, Var08, WN01, WKP^{+}01, WGS00, Wal03, WGDZ04, WW05, WC05, Wan06a, Wan06c, Wan06b, Wan09b, Wan09a, WP00, WMNS09, WH00, kWpLwW01, WLW04, WTW04, Xia01, XC03, XSC09, XZ12, YNK05, YZW02, YWLC04, YD06]. method [Yok09, ZWD05, dNKM07, dO02, dlHV08]. methodologies [Bae03]. Methods [Hoo04, KNU00, AA07, AKS01, AKS02, AK07, BK01, BDP00, BT01, CFMR08, CTG01, CvdEF^{+}05, CKA^{+}09, DKMF03, DEW00, Fra02, FS01a, GSM^{+}03, GF02c, HDGM07, HGVCM^{+}02, IHAR09, KMS09, KSC^{+}00, KALC08, LVV04, LVV06, LIR^{+}06, LJY07, LZ00, LNC^{+}03, MVJ09, MPR05, MDC09, MKK05, MMMM00, MKS07, PKSF01, PKPV02, PKKM02, PSK01b, Ram03, SG06, SLMS06, SMZ05, SQ03, TBZ12, ULA^{+}02, VPCK04, Van05a, Van05c, VIV01, itVPG08, VT00c, Wan05b, WGL06, WDHE04, WYL09, WYX09, Wu10, YWYF09, YNS^{+}09, Zah04, Zah05, ZSM05, Zit09, dMBC^{+}06, vdEFL^{+}02, vHLP08, BP08a, BFI^{+}00]. Metropolis [FNR^{+}07, KM01b, DDD^{+}01, FNR^{+}06, KM00a, dS03]. mFOAM [JS07]. mFOAM-1.02 [JS07]. MgB [HTA08]. MgO [MS05b]. MHD [Ras17, ATF^{+}09, ADG08, CLR08, DM09, GIME02, GZ07, Huj05, KKSR04, KSPT04, KKF^{+}04, MMTH04, Ras09, RJFB08, Ros04, SIH^{+}01, SGK09, SPP^{+}04, Zie04, Zie08]. MHD-code [GZ07]. micelle [Car07, SCO00, Car07]. micelles [AP05]. micro [ADD^{+}03, AAA^{+}00, KHIL07, LTG09, OML09, PRSB08]. micro-canonical [PRSB08]. micro-DST [ADD^{+}03]. micro-macro [OML09]. micro-structure [AAA^{+}00, LTG09]. microcanonical [FdO09, RLU00]. microhollow [HKLY07]. micrOMEGAs [BBPS02, BBPS06, BBPS07a, BBPS07b]. micrOMEGAs 2.2 [BBPS09]. microscopic [Lud02, PY08, Rap08, VS01]. Microscopy [RdAGV^{+}00]. microstructure [CFJ09]. microstructures [Nii00]. microturbulence [Lew04]. microwave [CH09, HJZL07, NV09]. middleware [CSZ^{+}07, HKM^{+}07]. Mie [WV04]. Mika [Vio04]. Mills [MG09b]. MIMD [VT00b]. Mimetic [KT05]. MINERVA [ATF^{+}09]. MinFinder [TL08b, TL06c]. mini [HFN03]. mini-jet [HFN03]. minima [CCRA05, TL06c]. minimal [LPC^{+}04, DGS08, FIJ^{+}03, HS02]. minimisation [HP06]. minimizing [Lor08]. minimum [JPS^{+}09]. Mining [HCK01, PB09b]. miscibility [Mül02]. Mix [LL00, WSCW09]. mixed [BBD^{+}09, CHS09, Kim07, LCM00, Nap09, VKM^{+}05]. mixing [Bat03, Cun09]. Mixture [FL01, CTI07, Frü03, GAR05, TE05, WS02]. mixtures [JBS08, LL00, MY00a, MY00b, Pur02, itVPG08, VMMB02]. MJK [GG03]. mK [Yan03b]. MM [GSM^{+}03]. MMM2D [AH02]. Mn [KMD^{+}02]. MO [iTKST01, YKK07]. mobility [Mam08]. mode [AK03, CDQF07, CAW00, WBDB04]. mode-matching [CAW00]. Model [CGIA07, FK00, KOS^{+}09, RCGC00, AGJJ07, AGV00, AIOST03, AC07, AAG^{+}04, ASF^{+}05, BLS09a, BBOY08, Bat03, BBB^{+}04, BBPS07a, BBPS07b, BBPS09, BCBJ02, BBJS09, Ber03b, BG06, BCD^{+}07, BFL04, BC00, BFB^{+}08, CMRS02, CM06, CSW02, CFJ09, CBBJ02, CL02, CMD00, CS07, CHP04, DKC08, Elm09, EM08, FFK02, FGA04, FM00, FHR^{+}05, FV02, Frü03, FMN01, GGQ01, GIME02, GFP00, GLP03, GBC^{+}04, GOG00, HD04, HBMJ05, ISS^{+}02, IKO00, ICO01, ISH01, JS08, KV08, KSS02, KM01c, KITK00, KB04, KTG04a, LPC^{+}04, LR07, LBM05, LCV06, LDZ^{+}08, LMS^{+}02, LCB07, LA09, LS05, LS09, Mas00, McK07, MLF07, MK02, NHS07, NV09, NSYZ02, Nii00, Oli01, Ots01, PPM04, QTL06, QCML03, RTS01, RMK05, RDSS01a, RDS02b, RR02, RLU00, Sal02, Sch08, SDLW07, SG04a, SZ00b, Sev00]. model [SR01a, Sta02, SKF05, TRGR08, TSI02, TFM09, TWY09, TL09, VK09a, Wal03, Wen01, WCH09, WG01, WL08, YH02, YD07, YB02b, ZSdD^{+}08, ZHC00, ZY09, dO09, vdSvdG08, Dan09a, FIJ^{+}03, GMAN^{+}07, HS02]. Model-Driven [Dan09a]. model-independent [KV08]. Modeling [ABV02, CHL^{+}07, GVMW04, MY00a, MCH02, PB09a, SZ04, TDY02, AFK^{+}07, AOT01, AP05, BLS09b, BS00b, CMT00, CMT01, DVG05, Del03, GT04, HvDJvdM01, HMY^{+}02, Huj05, HHWH07, JOS07, LPRS02, LPR04, LdG^{+}07, Mar08, MTZ00, NW02b, NP01b, OCS^{+}08, PCF05, PHKL02, Pin01, Pop03, Ram10, Ram12, RDS02a, RG04, RMVQ07, SG01, Str01a, SKRK04, SBCZ08, The05, itVPG08, VSBD00, Whi00, Yep02, ZS07]. modelings [GCP^{+}02]. Modelling [TC07, AP09, Bru00b, NK07, NRDHB01, PFPB^{+}09, SBM^{+}04]. Models [Koc02, AGM^{+}00, BPRW06, BR09, BMML05, Bro07, Bru04, CYAS05, CT00, CK08, CA09, CPT^{+}01, CCRA05, FdO09, Flo01, FL01, HJ02, KMZZ05, KGM00, KM05, KACB07, LCB^{+}00, LV08, LLT^{+}02, LKKK07, Liu07b, Luo00, MKM02, RD05, RS09, Rob00, RTVZ08, SS09a, Wel01]. modern [EHHH01, TYS^{+}00]. modes [AGJJ07, BS00a, BGS^{+}04, EMJH03a]. Modifications [NP01b]. Modified [KD09, LZS06, TB87, Yan03a, Yan03b, BSB02, CFMR08, LJ09b, MBG03, Ras09, Ras17, SPS09, TL06a, WP06, kWpLwW01, Zak00a, ZLL09, Tho04a]. modify [HKK^{+}01]. modular [PKB^{+}01, Wan01]. modulated [KS01]. Module [Ano04-46, Fri09, MMR04, PMA^{+}04]. modules [PBB^{+}04, WCGL00]. Molcol [FBL00]. Moldy [Ref00]. Molecular [AP05, BBB^{+}04, BDHP08, BRdAHK04b, CDFF05, DELG05, FS01b, FS02, HOT07, KEL02, KRTZ02, OO05, PKRK07, Rap08, TNI^{+}07, TYS^{+}00, ASH06, ASS^{+}02, BSB02, BGH^{+}09b, BTS05, BBB^{+}09b, BMvG00, BK05c, CW02, CCFG05, CLFH07, CFJ09, CCD07, CTI07, CR00, CW00, CC00, DC05a, Dun05, Dup01, EVL00, FG04, FBL00, FPB08, HDGM07, HL00b, HM06a, HKK02a, IW01, IW02, IN09, ISH01, JAT03, JRT00, KCC^{+}00, KMD^{+}02, Kar02, KFJ^{+}09, KSYE00, KA05, KBC^{+}09, KBG00, KS04a, KM05, LM02a, LZS06, LZS08, LSVMW08, LL00, MBR01, MPR05, MMR04, MKB02, Mor01, MSH01, Nak08, iNKNV08, OSK^{+}02, OK06a, OK06b, OD07, OCK^{+}00, OMF03, ÖDÇ02, PLPS08, PJK00, PHF^{+}07, PLS09, PKPV02, PP02, PS09, Rap02a, Rap06, Ref00, RJCH00, RF05b, SG00a, SM04, SM06a, SBM09a, SBM09b]. molecular [SNS01, SHZ01, SKNV05, SWC^{+}03, SLBG09, SPM00, SFSL09, SEE^{+}03, SS02b, SS05, TAKN02, Tod01, TGB01, Tsa02, VCCS05, Val05, itVPG08, VPK^{+}01, WHCL07, WHL05, YWLC04, YvG05, ZE00]. Molecular-dynamics [KEL02, MSH01, OLS^{+}01, SHZ01, Tsa02]. molecule [Hin00, LCV06, LDBG08, NFS01a, NFS02, NT05, NW02a, Ton07, WM00]. molecule-doped [NW02a]. molecules [ARV02, Bac00, BBOY08, BTK^{+}02, BSS09, DN05, GLL^{+}02, HC08, HSS^{+}08, IW01, IW02, JAT03, KJ07, KLTH04, LVLS02, LRR^{+}09, MVS05, NW02a, NY07, RMLB01, ST02, SJP05, TKB^{+}04, TNCG00]. MOLED [HTL^{+}03]. Møller [ACIZ07]. MOMDIS [BG06]. Moment [BBOY08, DN05, Goc04]. momenta [FIBT01, FIT03, Fri09, GFF01, GF02a, GSF05, IFF01, PFG06a]. moments [BKKS09, ISX05, ST02]. momentum [DK05, Dun05, GFG01, GF02b, HCO01, ID09, IM01, Jia08, KM00a, KM01b, Ste02, SFSL09]. MonALISA [LNV^{+}09]. MONARC [MC01]. monitor [LNV^{+}09]. monitoring [GC01]. monolayers [SDLW07]. Monotone [CL02, Li03]. monotonic [Dem03, DB08]. Monte [FNR^{+}07, GPW^{+}09, JKW06, SVMT00, TA00a, WA07, AW04, ABM03, ACIZ07, ASF^{+}05, AGS07, Ano03h, ABB^{+}09, Asc08, BS06a, Bae03, Bae04, BBB^{+}09a, BJ02, Bar00, BDG^{+}08, BvG02, BR09, BL00, BMML05, BHM^{+}07, BM01, BHL02, BK05b, BDYK04, BKB02a, BKB02b, Bur02, BB03, CGCS07, Che05, CGK^{+}00, Cun09, CKA^{+}09, DS01, DDD^{+}01, DGLB08, DDRW03, DH01, FNR^{+}06, FdO09, FMN01, GS01a, GPW04, GW01a, Gra02, GOG00, GRS06, HPC05, HKLY07, HCIK00, Huk02, JKW00, Jad00, JWW00a, JWW00b, JPS^{+}01a, JPS^{+}01b, Jad03, JS06, Jun02, JBS08, KH01, KPL07, Kat02, KRW03, KL06, LTA05, LF02b, MBKJ09, MRS04, MHS05, MSS^{+}09, Maz00, MSK^{+}05, MP03, MMB02, MB05a, MP06, MG09a, MABK02, MER^{+}00, MKM02, Nat08, Nil07b, OTY02, OPO^{+}08, PMA^{+}04, PSW00, Pop03, RP02, RIB01]. Monte [RPD^{+}05, RS00, RK05, Sch04, SVP09, SLWH02, SSLN02, Sul05, TA00b, Tak00, Tom09, TNCG00, Trö08, ULA^{+}02, Uhl03, VYK02, VPNW02, VMMB02, Wal03, WL00, WK02, WH00, WLGX09, YC07, dS03]. Monte-Carlo [WJW09]. MontePython [Nil07b]. morphogenesis [CGIA07]. Morphological [MD00, GBA01]. morphology [BM02b, CGC^{+}09]. MOS [LLT^{+}02]. Moshinsky [UTKF05]. Mossotti [LWY01]. motility [WG01]. motion [DMR02, FRdS09, KKSR04, KLTH04, NKV03, OMF02, RLU01, Sta00, TMTF00, TGB01, Yok09]. Motion4D [MG09c]. motions [LV08]. motivated [Pee07]. moves [WL00]. moving [GSGT03, PPC07, SFF^{+}04]. MPI [BCAD06, BADC07, Gao03, MGG05]. MPI-2 [BCAD06, BADC07]. MSPHD [GSSN00]. MSSM [BDW06, BBPS02, DGS08, DKM07, HHH^{+}09, HHW00, LCE^{+}09, Mah08b, MDM05, QxW07]. much [Ort00]. MULTEM [SYM00]. Multi [DSHH05, FLO06, GIME02, Ida03a, Jad00, LbotMC01, NJ00, SQ03, TYN02, TCF00, AK07, BAD01, BBBD06, BW08, Bre05, CD01b, CR00, FH04, GFP00, Huj05, KNTG03, Li03, LLLZ01, NW02b, OSK04, ISX05, SSP08b, SIE04, THC^{+}07, Val05, WGDZ04, WC05, WMNS09, WRMG05, Xia01, Yam00, Yok09, dNKM07, BMS^{+}09, Ida03b]. multi-beam [OSK04]. multi-derivative [WGDZ04, WC05]. Multi-dimensional [Jad00, TYN02, Bre05, CD01b, Huj05, KNTG03, Li03, Yam00]. Multi-fluid [GIME02, GFP00, Xia01]. multi-loop [BW08, WMNS09]. multi-material [dNKM07]. multi-moments [ISX05]. Multi-parameter [DSHH05]. multi-particle-collision [WRMG05]. multi-phase [NW02b, Yok09]. multi-pipeline [SIE04]. multi-platform [BAD01]. multi-processing [CR00]. multi-quantum-well [LLLZ01]. multi-range [Val05]. multi-scale [THC^{+}07]. multi-sequence [SIE04]. Multi-speed [TCF00]. Multi-step [FLO06]. Multi-symplectic [SQ03, AK07]. Multi-threaded [LbotMC01]. Multi-threading [NJ00]. Multi-time-step [Ida03a, Ida03b]. multi-valued [FH04]. multi-variate [BBBD06]. multi-wavelets [SSP08b]. MULTI2D [RtVR09]. multibaric [OO05]. multibaric-multithermal [OO05]. Multibillion [Rap06]. Multibillion-atom [Rap06]. Multibondic [BJ08]. Multicanonical [Ber03a, HBMJ05, BB09a, ISH01, RD05]. multichannel [LVV07, MBG03, PB09b]. multicritical [DKC08]. Multidimensional [MM05, Asc08, BB04a, Hah05, Hah07, MYC09, TL06b, XSC09, vdHKM08]. multidomain [Hua09, PKST03]. multielectron [Kur02]. multifragmentation [DSS01]. Multigrid [MGN07, KLM00, SS05]. multigrids [OSK^{+}02]. multilane [CL02]. multilayer [HTL^{+}03, RR02]. multilayered [PP09]. Multimillion [VKN07]. multinormal [FGA04]. multiparticle [Fra07a, RFK08]. multiphase [IKO00, KITK00, TFM09, ZHC00]. multiphoton [NFS01b]. Multiple [CGK^{+}00, ABOSPG09, BDM09, CC09, FL01, FEHC01, GGQ01, GSSN00, HDGM07, KM08a, Krö05, NP00, Pet04, PCA^{+}07, SSPM05, SG04a, TYS^{+}00, UJSW06, VW05, vDGM^{+}09, vE08, GMAN^{+}07]. Multiple-Compartment [GMAN^{+}07]. multiple-CPU [FEHC01]. multiplication [BMG01, EFS^{+}08]. multiplicative [WH06]. multiplicity [VT00a]. multipliers [DH00]. multiply [Cha00]. Multipole [OIKN02, HJZ09, LM02a, OCK^{+}03, YNS^{+}09]. multiprocessor [GBD03]. Multiresolution [KCC^{+}00]. Multiscale [Bra05, All05, DC03]. multislab [dA08]. multispectral [CEM08]. multistep [Wan05b]. Multisymplectic [Cai09, Wan09a, TQZM08]. multitasking [WLH00]. multithermal [OO05]. Multithreaded [BB00, BD06, Dan09b]. multitrack [SF00]. Multivariate [RB05, Bel05, KMH02]. multiwavelets [SR05]. Muon [Kud09, GKM^{+}00]. muonium [KK00]. MUons [CBMS08]. MuPAD [AC05b]. MUPAGE [CBMS08]. MUPHY [BMS^{+}09]. MUSIC [Kud09]. MUSUN [Kud09]. mutant [CDFF05]. mutualistic [BCHP09]. MW [CHL^{+}07]. Myrinet [ACC^{+}01].

N [Tsa02]. N.H [Laf03]. n=4 [AKZ00]. Nano [HJM02, CGC^{+}09, FHR^{+}05, HTM^{+}08]. nano-agglomerates [CGC^{+}09]. nano-friction [HTM^{+}08]. Nano-scale [HJM02]. nanoclusters [PP02]. nanoconfined [TK08]. nanocrystals [CTSZ07, LNK01]. nanodevice [Hua09, LKC06]. nanodiamond [AFK^{+}07, RG05]. Nanodroplets [MMB02]. nanoflows [MS05a]. nanoparticles [BK05a, KEL02, ZLM04]. nanopattern [MLPT08]. nanopore [MB05a]. nanoscale [LY05, VKN07]. nanoscience [OLX07]. nanoscopic [BHM^{+}07]. nanostructured [KCC^{+}00, PMH08, YGT^{+}02]. nanostructures [NKSL05]. nanosystem [Ano09a]. nanotube [CSC^{+}07, CSC^{+}08, HKK02a, HKK02b, LC08b, PLL07, SGL09]. nanotubes [AFK^{+}07, KKKC07, ÖDÇ02, YN05b]. nanowire [GVMW04]. nanowires [PMH08, SG05]. narrow [DN05, LLV^{+}01]. National [Ano04-46, PMA^{+}04, Yos00, VSBD00]. natural [BSB02]. nature [CRPC08, TRAdO09]. Navier [ICT01]. Ne-like [HD04]. Near [BK05a, DEW01, JKKT00, NSKS01, Riz02, YW01]. Near-field [BK05a, NSKS01]. NearFar [Cha07]. nearside [Cha07]. nearside-farside [Cha07]. NEC [EL04]. Ned [Ano04-57]. negative [BNS07, Nob04, Rou01, Sea02a, SHJ07]. neighbor [ABRS12, MDT03, YWLC04]. neighborhood [TBR07]. neighbour [Mas05]. nematic [PGS02]. nematohydrodynamics [TDY02]. neoclassical [Lüt04, WTH^{+}04]. neon [OMC00, OMC00]. nested [Eli08, KSC^{+}00, MMTH04, PZ01, CvdEF^{+}05]. nestedness [BCHP09]. Network [FFS01, CBM^{+}05, FDM07, GC01, LV08, LFT01, LFT03, MLF07, RDSS01a, RDS02a, RDS02b, SM01, Sug01, YNZ^{+}09, ZWY04]. Networking [New07]. Networks [MTC07, BOG^{+}07, BCHP09, CD08, HSJ02, KKH07, KOS^{+}09, LLH07, Lik01, MLG^{+}01, MBC^{+}09, New02, RDSS01b, TWY09, YD07, ZBB^{+}06]. Neural [EFG^{+}00, Ano01n, LFT01, LFT03, Lik01, Sug01, TWY09, YD07]. neurotoxin [CCD07]. Neutral [PBB^{+}04, HHW00, Lon07, Man04, MOC03, OMC00]. neutralization [WCG04]. neutrino [CBMS08, GK05, HLW05, HKL^{+}07b, KL01, KFI^{+}01]. neutrinos [KL01]. neutron [CDQF07, EKW09, HBRS05, LZ00, RLV^{+}08, VT00a]. neutrons [Yos03, Yos07]. Newcomb [ATIO06]. Newton [AP04, DSH02, DSH03, Jam00, Sim08]. Newtonian [CLR08, LC00, Pue06]. NEXTCALIBUR [BPP01]. Ni [Bur02, KEL02, KMD^{+}02]. Ni-based [Bur02]. NiAl [LAF01]. Nicholson [Sch05]. NIMROD [KSSH04]. NiO [Kar02]. NIRVANA [GZ07, Zie04, Zie08]. nitride [SLC09]. NLO [CC04]. NMHDECAY [EH06]. NMR [BDM09, PCYC02]. NMscatt [MFVJ07]. NMSPEC [EH07]. NMSSM [EH06, EH07, Mah09b]. NNLO [CCG08]. no [Bar03]. nodes [GBD03, MTC07]. Nogami [RGD^{+}01]. Noise [KA05, BDBV12, Gen01, HKP02]. Noise-driven [KA05]. noises [Mil06, Mil07]. noisy [KV08]. Non [MVS05, TE05, AA08, BCD^{+}07, BL05, BFLW07, BC05, Bru00a, Cha04, CSW02, CLFH07, CL03, FRdS09, FGF03, GF02b, GBM02, GSGT03, HYY07, HSSA01, IK00, JS06, KV08, KG07, KH06, KB02, LC00, MSD08, NJ01, NP01b, NFS01b, RLRR06, RSMK^{+}00, Ram05, TNY00, UNK12, WP00, WRC^{+}04, Yao09, ZF00, ZSSA00]. non-adiabatic [BC05]. non-classical [KB02]. non-conservative [TNY00]. non-convex [RLRR06]. non-equilibrium [BL05, HYY07, ZSSA00]. non-Fickian [CL03, Ram05]. non-filtered [GBM02]. non-Hermitian [BFLW07]. non-ideal [IK00]. non-integrability [HSSA01]. non-integral [AA08, Yao09]. Non-isothermal [TE05]. non-iterative [WRC^{+}04]. non-linear [Bru00a, FGF03, GSGT03, NP01b]. Non-Markovian [MVS05, FRdS09, JS06]. non-Newtonian [LC00]. non-orthogonal [WP00, ZF00]. non-oscillatory [KG07, UNK12]. non-overlapping [CLFH07]. non-perturbative [NFS01b]. non-scalar [GF02b]. non-spherical [RSMK^{+}00]. non-staggered [Cha04]. non-trivial [MSD08]. non-uniform [BCD^{+}07, KV08, KH06, NJ01]. non-zero [CSW02]. Nonadiabatic [SK05]. nonautonomous [HL00c]. nonequilibrium [MCH02, iOY01, RMK05, Tod01]. Nonlinear [KDSB04, NYH04, WGY01, AP04, AA08, ASJ^{+}03, AH03, AOT01, AK07, BGH04, BB04b, BGS^{+}04, DWZS05, DDFI09, DKV00, EST00, FD03, FGA04, GT01, GKI02, GH01, GCD06, Hon04, KA09, bLpL02, LL04, LJ08, pLbL03, Liu07a, Lüt04, MT01, PCC^{+}09, RE09, RMWH01, SZ00c, SQ03, Wan09b, WW06, qXbL04, qX08, qX09, XZ12, Yan02, Yan03c, Yan03d, YRR07, ZLL09, dlHV08, Par04]. nonlinear-condensation [ASJ^{+}03]. nonlinearity [KLD04]. nonlinearly [YB02a, Yan03a, Yan03b]. Nonlocal [BBBR04]. Nonperturbative [Sav01]. nonrelativistic [MMR04]. nonspherical [IW02]. nonuniform [Bel05, Eli08, KV07]. nonzero [BFLW07, KJ04]. Nordsieck [BMC05]. normal [BB07, CRUV00, She08, Var02]. normalization [UCG^{+}05]. Nosé [Lei02]. notation [GZF04]. Note [Ano06-29, Ano07-30, Pub07, WYL09, Ano03-43, Ano03-44, Kar01, Koz02, qX09]. notes [BCKT09]. Novel [ZZH09, BMML05, FGF03, FH00, GBA01, HKK^{+}01, LH03, Mas05, MVS05, NJ01, PL05]. Novosibirsk [BEM^{+}02]. nozzle [CTG01]. NP [Zim02]. NP-complete [Zim02]. NPT [IW02, OK06b]. NUBEAM [PMA^{+}04]. Nuclear [MC03, New07, Bar00, BDLT02, BCG03, BFB^{+}08, DTD^{+}02, Di 02, Elm09, GFG^{+}06, HL08b, MM05, MM09, NY06, NY08, QCML03, Tom09, UTKF05]. Nucleation [SF05]. nuclei [BAB04, DSS01, RGD^{+}01]. nucleon [AIOST03, MNYY00a]. nucleon-nucleon [MNYY00a]. nucleons [Wro08]. nucleosynthesis [PCE^{+}08]. nucleus [VEG08]. nudged [Nak08]. Null [DW01, Rib02]. Null-field [DW01]. Number [DGLB08, LBP^{+}09, ATB^{+}01, DH00, FPB08, KKK06, Lad09, LCPC04, MI05, OGWH03, Pro00, Sch06a, TYSH05, WL00, WHO02, WH06]. numbers [FH00, HB05, Str05]. numer [Sea02c]. Numeric [Bre01, Ada04, BGH^{+}09b, KS05, PC08]. Numerical [AA07, AMP^{+}00, AT09, BF04, BS00a, BBD00, BCD^{+}07, BK01, BV00, BFI^{+}00, BDP00, CMS04, CSCK08, CGM01, CTG01, CvdEF^{+}05, CKA^{+}09, FRdS09, Fat02, Gal00, GR01, GR02, GHLW03, GBC^{+}04, HL00c, Hoo04, Inu07, IN09, JW02, KKSR04, KSHP02, Kon01, KM01d, KK06, LdVJ06, LdG^{+}07, LLT^{+}02, LWLL07, LC08b, LCM00, LC00, LEG02, Liu07a, MDC09, MYJY01, NRDHB01, PBB^{+}04, PHKL02, RMVQ07, SLC09, SS07a, SNS01, SJDC07, Sho07, SSP08b, Sol01, SM02, SKRK04, Sug01, TMTF00, TAKN02, TKP06, TY01, VW05, Vos06, WV05, Wil09, vdEFL^{+}02, AP04, AG05, Asc08, ASVA00, AKS01, AKS02, BDK^{+}06, BZ00, BH05, BGH^{+}09a, CCGR09, Cai09, CL08b, CRS09, DGV08, DKMF03, DGSL09, Don02, Dys02, EST00, FLO06, Fij99, Fij00, FH00, Fra02, GME06, Hah05]. numerical [Hah07, HJZL07, Huj05, HHL06, KKK06, Kau03, KL01, KCH00, KNU00, KA05, KN07b, LVV07, Lee04, LR07, LLCS01, LH03, Li03, LY05, LCB07, yMS01, MSS^{+}09, MLF07, Mil06, Mil07, Min01, MA04, MA08, MP01b, MKS07, MP05, Nur04, OCS^{+}08, PAS09, Pis00, PSK01a, PR06, PSV00, Ram05, RM05b, RS09, SMSE03, SW09, Sho04, SW00a, SVA03, Sim09, Ska05, Sus01, TKS^{+}01, TQZM08, UK02a, UK02b, Van05a, Van05b, VHLP09, WGS00, WGDZ04, WC05, Wan09a, WDB04, Wu10, YWYF09, You05, ZSK^{+}04, Zit09, vDGM^{+}09]. numerically [Tal09]. Numerov [FSW08, Sea02c]. Numerov-type [FSW08]. NumSBT [Tal09]. NVIDIA [MBKJ09]. NVT [IW02]. NWChem [KAB^{+}00, SPM00]. Nyström [Fra02, KMS09, PAS09, Van05b].

O [EVL00, Hah09, OCK^{+}00]. Object [Bre01, BHNW01, DG08, KLM00, AGV00, Che05, DM07, GGQ01, QRH00, Wil09]. Object-Oriented [Bre01, DG08, KLM00, AGV00, Che05, DM07, QRH00]. objective [KV08]. Objectivity [SM01]. Objectivity/AMS [SM01]. objects [HS01b, ICO01]. oblate [KJ07]. Obrechkoff [CWSH08, DWZS05, WW05, Wan06a, ZWD05]. observable [GG03]. observables [BDW06, HHH^{+}09, Mah09b, Mah09a]. obstacles [DEW00]. obtained [Ano04b, GZF04, Tam03, TMN01]. obtaining [KKK06, MYL^{+}08]. occurring [FK00]. ocean [NN06]. octopus [MCBR03]. ODE [WDHE04]. ODEs [CTR00, IVD03, MT01]. ODPEVP [CGVA09b]. Oedometric [OML09]. Off [KK05, ÇHM00, JC08a, KY07, Mar01, MP05, SBCZ08]. off-centered [MP05]. Off-lattice [KK05]. off-line [Mar01]. off-shell [ÇHM00]. offline [FFPW01]. offs [Oli01]. OK1 [OSK04]. OK2 [OKS04]. oligonucleotides [BSB02]. OMEGA [LANM^{+}01]. on-shell [KM00a, KM01b]. ON-SHELL2 [FK00]. One [BD02, Ker02, LKPH08, AIOST03, BS00a, BGH^{+}09a, CTG01, De 02, Dev05, Eli05, GF02b, Har02, HJZL07, Inu07, KKK06, LHC01, LHC02, LSL07, MSD08, NN09, Nik03, Ots01, Ram05, RM05b, SW09, SGF03, SBD^{+}06, SM02, TNI^{+}07, WGDZ04, WC05, Wan06a, Yos03, Yos07, Zak06]. one- [HJZL07, Nik03]. One-Dimensional [Ker02, LKPH08, CTG01, Eli05, Har02, Inu07, LHC01, LHC02, LSL07, MSD08, Ots01, Ram05, RM05b, SW09, TNI^{+}07, WGDZ04, WC05, Zak06]. one-gluon [KKK06]. One-loop [BD02, BGH^{+}09a, NN09]. one-nucleon [AIOST03]. one-parameter [De 02]. one-particle [Dev05, GF02b]. one-photon [BS00a]. one-step [WC05, Wan06a]. ONETEP [HHM^{+}09]. onia [DGSL09]. Onion [ML03]. Onion-Peeling [ML03]. online [EFG^{+}00, Gre07]. onto [Rob01]. open [AdlT03, ABNÅ05, Bae04, EHHH06, ISSB01, JP09, MSB09]. OpenDX [SC04]. opening [BJ02, Del03]. OpenMP [CC00, Goe02, MGG05]. OpenMP/MPI [MGG05]. operations [AA00, AA01b, Ixa01, RF07]. Operator [Flo01, BFLW07, CvdEF^{+}05, Cun09, CKA^{+}09, CA07, EG09, GL02, GLP03, MK08, MM01, Ram10, vdEFL^{+}02]. operator-splitting [GLP03]. operator-variational [MM01]. operators [GF02b, SFSL09].

Opportunities [Gun02]. OPT [RMMP02]. optical [ADS06, BB04b, CIC^{+}03, CC09, CFJ09, GCD06, HTNFBS06a, HTNFBS06b, MSB09, MTZ00, MBC^{+}09, NKSL05, NRDHB01, NY08, PCA^{+}07, QCML03, RG05, TNBSF04, Wes07, Whi00, YC07]. optical-properties [MSB09]. optics [SWS^{+}12, Tót08, FWP01]. Optimal [CJT06, GJT03, LFT03, SA09, VIV01, ZA01, NHS07, ZSdD^{+}08]. optimisation [BBBD06]. optimised [ASH06]. Optimization [BJ05b, Goe02, SWC^{+}03, BMSG01, Elb05, FEHC01, Iwa01, KPD06, KFJ^{+}09, KPF03, LPC^{+}00, MTJ02, OS04, PDL04, PL05, PAT^{+}09, TLP04, TL06b, TL08a, VPP^{+}12, WHCL07, WJW09, ZS03, ZS07, ZSdD^{+}08, ZS08, Zim05]. optimize [LNV^{+}09]. Optimized [BDM09, OMF02, Sch06b, SK08, FMN01, KT04, Van05a, WK02]. Optimizing [BH03, CW01, dS03]. Optimum [OD08, WMNS09]. options [TLP04]. optoelectronic [GCD06]. Orbit [BDBV12, Dev05, TEP00]. Orbit-based [BDBV12]. orbital [HLC08, KH09, Sim08, TKN^{+}08]. orbital-dependent [TKN^{+}08]. orbital-free [HLC08]. orbitals [BGH^{+}09b, FGMT02, RF05b, ZF00]. orbits [PKPV02, VPK^{+}01]. ORCO [SMSE03]. Order [GBTM07, SR09, WYX09, ACK05, AKZ00, BB04a, Blü00, CFMR08, CM02a, CBBJ02, CJK09, DR09, FMG00, HBMJ05, IVD03, JH09a, JPS^{+}01b, Kol09, LVV04, LRI^{+}06, LJ09a, LA09, MVJ09, MA04, MA08, MKS07, Poi08, Poi09, ISX05, SS00, SLMS06, Sim00, SVA01, Sim08, TB85, TB87, Tho04a, Tho04b, TK08, Van05a, Van06, WGDZ04, WHJ06, WYL09, ZWD05, vH06, vH07]. Order- [GBTM07]. order-parameter [HBMJ05]. ordered [NFS02]. ordering [JPS^{+}09, NG02]. Ordinary [IHAR09, Ram05]. organic [HTL^{+}03, MSY07]. organisation [SAU^{+}04]. organization [NYH04, Ort00, RDS02b]. organize [Ort00]. organized [SOS01]. orientable [Huj05]. orientation [CGC^{+}09, CFJ09, WMNS09]. orientation-specific [CGC^{+}09]. orientationally [NFS02]. Oriented [Bre01, AGV00, Che05, DG08, DM07, FFS01, GGQ01, KLM00, QRH00, Wil09]. origin [Riz02]. origins [CT00]. ORNL [KN07a]. ORNL-mathematical [KN07a]. orthogonal [KK01, KTT02, WP00, ZF00]. orthogonal-dimer [KK01]. Orthogonalising [IBM03]. Oscar [Ano04c]. oscillating [CM02b, DKC08, PAS09]. oscillation [HLW05, HKL^{+}07b, Ida02, NFS02, Wei02a]. oscillation-free [Ida02]. oscillations [BD06, Dan05a, Dan05b, DS06, Dan07]. oscillator [DD00, DO04, DO05, DSC^{+}09, EKW09, GME06, HL08b, HB05, MAM04, MAM07, SOYN01, SDNR05, You09]. oscillators [DDFI09, Fra02, TY01, Van05b, WYL09, Wu10, YWYF09, YT01b]. Oscillatory [BZ00, AA00, AA01b, FBB01, Fra07b, HSSA01, Ixa01, IP01, Kim03, KSHP02, KCH00, KG07, Sau00, UNK12, Van05c, WYX09]. Other [BOPC05, BC07]. OTI [Elm09]. out-of-core [BVY05]. output [BC07, Gha05]. overdamped [Gen01]. overlap [BFLW07, CvdEF^{+}05, Cun09, CKA^{+}09, CKLS09, GGL03, Jan05, KALC08, vdEFL^{+}02]. overlapping [BM02b, BDH^{+}05, CLFH07]. overload [ACC09]. Overview [BGLLW01, Oli01, KAB^{+}00]. oxidation [LAF01]. oxide [HSSA01, LY05, LC08b, RJCH00]. oxides [KKKC07]. oxygen [LN01].

P [Kar01, Eas08, SW00a, WW05, Wan05b]. P-stable [SW00a, WW05, Wan05b]. Package [KS04b, Pog05, AF05, AAG^{+}04, AGM^{+}00, BS06a, BC07, BB09a, BS06b, dSB00, BBJ^{+}08, BFB^{+}09, BDH^{+}05, CGC^{+}09, CKS00, Che07, Dem03, Dem06, DGSL09, EHHH01, EHHH06, Fer07a, FK00, Fri01, Fro00, FTGG07, GKI02, Gao03, GKI04, GI09, GDC01, GKR07, GHIL09, HKM^{+}07, HTNFBS06a, HTNFBS06b, HM06b, HHL06, Ixa02, JHFG07, KP00, KSYE00, KS84, KS08, KF05b, KVR^{+}00, bLpL02, LL04, LAMH06, pLbL03, LRR^{+}09, Lor08, LL00, MP04, MGPM07, MGYP08, MSB09, MP03, Mil06, Mil07, NFH06, Nik03, PFG06b, PZW^{+}00, PTL04, Por00, Pue06, QxW07, RMMP02, SBM09b, SBM02, Sem09, SLBG09, Ste01, SC04, Tót08, Wan01, WCH09, dlRL09, vH06, vH07]. packages [BCV03, GKP^{+}06, KPD06]. Packet [KRTZ02, BS04b, LJ01, Mei01, Sal03, ZWY04]. packets [Bow02]. packing [HSJ02, YZD^{+}07]. Padé [FH04]. Padé-approximants [FH04]. pages [Hoo04]. Painlevé [XC03, qXbL04, qX08, qX09, ZLL09]. pair [AAC^{+}06, BBC^{+}01b, JWW00a, JPS^{+}01a, JPS^{+}01b, Kol03, KFI^{+}01, Van05b]. pair-production [KFI^{+}01]. PALP [KS04b]. PANMIN [TLP04]. PANN [MNYY00a]. Papers [BDL00, Aok01]. PAPH [MNYY00b]. parabolic [BV00, Fat02]. paradigm [HHWH07]. paradigms [TYS^{+}00]. ParaGauss [MMR04]. Parallel [ATB^{+}01, BSDMH05, BMG01, BSK^{+}03, BVKW02, CR00, CW00, CC00, DN04, EVL00, Gol00, HC00, HL00b, Hef00, JKCGJ08, JRT00, LBP^{+}09, LZS08, LHS^{+}06, MP03, OPB^{+}09, QR01, RP02, RJCH00, TF04, TEP00, Uhl03, WMK09, WM00, WTW04, WHL05, WH06, ZE00, Zie08, ABC^{+}03, ABER00, ADBF03, AEB02, BAD01, BLCR05, BCAD06, BOG^{+}07, BMS^{+}09, BB00, BTK^{+}02, BJ03, CSS^{+}03, Cha00, CGIA07, CLL^{+}07, CMT00, CMT01, DMD^{+}07, Dec07, DPB01, Di 01, DUX^{+}09, Dup01, Eli05, EKW09, FMD07, FDM07, Fel08, FKP03, FEHC01, FMN01, GBFS07, GBD03, HSGBK08, HCH^{+}06, ICT01, JAT03, Jen00, JGJ09, KCC^{+}00, KAB^{+}00, KLD04, KM01c, KBG00, LCB^{+}00, LPC^{+}00, LCS07, LTG09, LR07, LLCS01, Li03, LC01a, LL00, MOM^{+}00, MC08, Mei01, MB04, MT00, Nak07, Nak08, NJ00, iNKNV08, OLS^{+}01, OSK^{+}02, OCK^{+}03]. parallel [OK06a, OK06b, OD07, ÖDÇ02, PHF^{+}07, PSP^{+}03, QRH00, Rap06, Ref00, RJFB08, SG00a, SM04, SM06a, SBM09a, iSAK^{+}08, SKNV01, SKNV05, SWC^{+}03, SBB03, SNBB02, TJD09, TRGR08, Tak03, TCY^{+}08, TG00, TC06, TLP04, Tót06, ULA^{+}02, VKPB09, VHLP09, WC00, WHO02, XON08, Yos01, Zha08, SVMT00]. parallelism [SPM00, TYS^{+}00]. Parallelizable [CA07, Xia01]. Parallelization [CMF00, FKG00, RGR^{+}04, SLWH02, WJW09, BS06a, Gao03, Goe02, MGG05]. parallelized [WHL^{+}07]. PARAMESH [MOM^{+}00]. Parameter [RPY07, AS03, Bre05, CNFR01, CGVA09b, De 02, DSHH05, HBMJ05, KKK06, KMH02, PS09, SZ00c, YM03]. Parameterization [AGM^{+}00]. parameters [Bar03, BDW06, FGA04, GMAN^{+}07, GKM^{+}00, HG02b, HM06b, HM08, IF03, LFT03, LANM^{+}01, MS08b, NY06, WV04, qX08]. parametric [CGVA09b, CBMS08]. parametrisation [GSF06]. parametrization [VCCS05]. Parametrizations [RF08]. paraxial [AT09]. parentage [Dev05, DJ08]. parity [ACIZ07, GLL^{+}02]. parity-violating [GLL^{+}02]. Parrinello [CCFG05]. Part [HTM01, Ida00, PSK01a, PSK01b, THM01]. PArthENoPE [PCE^{+}08]. Partial [Hoo04, MNYY00a, MNYY00b, FMG00, KS07, MTC07, SJP05, VBC07, qXbL04, YZW02]. Partial-wave [MNYY00a, MNYY00b, SJP05]. partially [BSTC05, LB04, Sle00]. Particle [BTS06, CPS00, CH09, KCR07, iSAK^{+}08, SWFL00, ZM00, ZLM04, ABRS12, BDYK04, BDV04, Che05, CY01, DC03, DDMM06, Dec07, Dev05, DJ08, DKM07, DEW01, EL04, FMD07, Fod05, FS08, GFF01, GF02b, GPW04, HKLY07, JH09a, JJHvO03, JS08, KLD04, LC01a, yMS01, MY00b, Man04, Mel05, MAM04, MAM07, MK09, NT05, NKV03, OD07, PCC01, PSP^{+}03, Poi08, Poi09, Pop03, Por03, RvOvV02, SLL01, SBB03, SS02b, SS05, TFM09, TC06, TE05, Tom09, UOM01, UOTM03, VPCK04, VKPB09, VBFD01, WTH^{+}04, WGL06, WRMG05, Esi01, KPL07, TCY^{+}08, TDD04, VAH04, VCF^{+}04, WJW09]. particle-based [MY00b]. particle-continuum [VPCK04]. Particle-in-cell [CH09, BDYK04, Dec07, EL04, FS08, HKLY07, JH09a, KLD04, LC01a, PSP^{+}03, Poi08, Poi09, SLL01, UOTM03, VPCK04, Esi01, KPL07, TCY^{+}08, TDD04, VAH04, VCF^{+}04, WJW09]. Particle-In-Cell/Monte [KPL07]. Particle-in-Cell/Monte-Carlo [WJW09]. Particle-inspired [CPS00]. Particles [HAA07, Bar04, CMD00, DHBE05, JKKT00, KH06, LMM^{+}08, MDM05, RSMK^{+}00, Str01a, TT06, WLR^{+}08, WV04, YZD^{+}07]. particular [AKZ00]. particulate [BC00]. partition [JK02]. Partitioning [SBB03]. PArton [KSS06, SR09, ABB^{+}09, BBB^{+}09a, CPW09, CS02, Sul05, Vog05, Wei02b, KKIS01]. Paschen [LSL07]. past [Ano02a]. Path [CC08, GOG00, MI05, SVMT00, KM05, Krö05, MG09a, RDFF02, ZE00, vE08]. Path-Integral [SVMT00, MI05, KM05]. Pathfinder [Nak07]. paths [Pet04]. Pattern [OGG07, Yan03a]. patterns [BBC^{+}01a, CLFH07, DG08, Gro01, YB02a, Par04]. Paul [Wan00]. Pauli [ZF00]. PAW [HTM01, THM01]. Pb [BNS07]. PC [FKP03, LC01a, iSAK^{+}08]. PC-based [FKP03]. PCs [Tak03]. PDE [FS00, KMH02, XC03]. PDEs [BBBD06, LJ08, LH01, PMG07, qX08, qX09, ZLL09]. PDSW [VS06]. Peaceman [Mah08a]. peak [CC09]. pedestrian [NHS07]. Peeling [ML03]. Pegasus [Vog05]. pellet [BDB^{+}08, SJCM04]. penalized [Lor08]. Penning [CBKM01, CKV04]. peptide [KPF03]. peptides [LPC^{+}00]. perception [Man02]. percolating [MDS09]. percolation [HCIK00, NLC09, Sat02]. Performance [De 07, FDM07, FEHC01, SvAS01, SHT08, BDK^{+}06, BMS^{+}09, CD05, COE^{+}05, CBM^{+}05, KAB^{+}00, MPR05, MC08, PPM04, PAT^{+}09, SBD^{+}06]. performing [CGC^{+}09, KFJ^{+}09, SGL09]. peridynamics [PLPS08]. Periodic [MNV00, AJT^{+}07, ASVA00, AK07, CY01, DWZS05, Dol01, FSW08, FBB01, HL00c, LÅT04, PKPV02, SVA01, SSLN02, THM01, VPK^{+}01, WW05, Wan05b, Wan06b, Yan02, YT01a]. peristaltic [SGK09]. permanent [DC00, FM03]. permanental [HLB06]. permanents [LB04]. Permutation [RLH^{+}09]. Permutation-reduced [RLH^{+}09]. persistency [ISSC01]. personal [Cip07, Cip08, Hib01]. perspective [Haf07]. perspectives [EL04]. PERSYS [Riz02]. perturbation [Dzu09, LRI^{+}06, LIR^{+}06, ZSdD^{+}08]. Perturbative [SR09, CS02, HS03, NFS01b]. perturbed [Fra02, Van05b, WYL09, Wu10, YWYF09]. petabyte [Ano09t]. PetaFlops [Att09]. PETAG01 [BDB^{+}08]. petawatt [KDSB04]. Petter [Hoo04]. Petviashvili [LL08]. pH [CCD07]. ph/0411186 [AAB^{+}07]. PHANTOM [BBB^{+}09a]. phantoms [KN07a]. Phase [Bur02, HBW05, KEM^{+}01, KS01, NT05, NW02b, Ots01, PRSB08, Ple02, SSLN02, Tat07, TL09, VMMB02, AGJJ07, Bae03, Bar03, Bin02, BHM^{+}07, BDHP08, CSCK08, CBBJ02, DKC08, DC05b, FHR^{+}05, GHPS04, JK02, JS05, JBS08, KM00a, KM01b, Kim07, KK01, KITK00, KK05, LJ01, Liu07a, LDZ^{+}08, LA09, MCH02, MBG03, MSS^{+}09, MSK^{+}05, Mor01, Pap01, PAS09, PP02, Pur02, RP02, RLH^{+}09, SWL09, SI01, SW00b, TMTF00, Tsa02, Van05c, YGT^{+}02, Yok09, Elm09]. Phase-field [NW02b]. phase-fitted [PAS09]. phase-lag [Van05c]. PHASE-OTI [Elm09]. phase-separating [Pur02]. phase-space [DC05b, GHPS04, Pap01]. PHASECALC [Bar03]. phases [KPS^{+}01, MVS05, RCG05]. PHEGAS [Pap01, CPW09]. phenomena [ABOSPG09, All05, BJ08, BL00, BMML05, DDM05, DDMM06, GH01, IN09, MPK00, MYJY01, Mül05, MKM02, PCF05, SAG^{+}02]. phenomenology [GHIL09, LPC^{+}04, LCE^{+}09]. PHON [Alf09]. Phonon [HKK^{+}01, Yos03, Yos07]. Phonon-band [HKK^{+}01]. phononic [SSPM05]. phonons [Alf09, Sri01]. phoretic [KH06]. phosphodiesteric [BSB02]. phospholipids [EL06]. phosphonate [BSB02]. phosphorus [Mor01]. Photo [MSH01, Ano07f, Ano08d]. Photo-induced [MSH01]. photoabsorption [GCP^{+}02]. Photoelectron [Vég04, GSSN00, Jia08]. Photofragment [ML03]. photohadronic [MER^{+}00]. photoinduced [IN09]. photon [BS00a, BvG02, CGK^{+}00, EST00, HGH^{+}05, KFI^{+}01, Nik03, VS06]. photonic [SHX02, SYM00, WP00]. photons [DDM07, LYL07]. photos [GKP^{+}06, GKP^{+}06]. Phys [AA01b, AAB^{+}07, CSC^{+}08, CGG^{+}09, CGVA09a, DVL^{+}04, Hon04, Ida03a, Ixa01, JKW06, KS08, LPR04, MSHP20, Nat10, Poi09, Ras17, Tho04a, Tho04b, TND05, TIM08, Voi03, WA07, Yos07]. Physical [Mey02, Mel01, WSB04, Yep02]. physically [RGR^{+}04]. physically-based [RGR^{+}04]. Physics [Ano02j, Ano02k, Ano02l, Ano02m, Ano02n, Ano02o, Ano03i, Ano03j, Ano03k, Ano03l, Ano03m, Ano03n, Ano03o, Ano03p, Ano03q, Ano03r, Ano03s, Ano03t, Ano03u, Ano03v, Ano03w, Ano03x, Ano03y, Ano03z, Ano03-27, Ano03-28, Ano04m, Ano04n, Ano04o, Ano04p, Ano04q, Ano04r, Ano04s, Ano04t, Ano04u, Ano04v, Ano04w, Ano04x, Ano04y, Ano04z, Ano04-27, Ano04-28, Ano04-29, Ano04-30, Ano04-31, Ano04-32, Ano04-33, Ano04-34, Ano04-35, Ano05k, Ano05l, Ano05m, Ano05n, Ano05o, Ano05p, Ano05q, Ano05r, Ano05s, Ano05t, Ano05u, Ano05v, Ano05w, Ano05x, Ano05y, Ano05z, Ano05-27, Ano05-28, Ano05-29, Ano05-30, Ano06b, Ano06c, Ano07b, Ano07c, Ano07d, Ano07e, Ano08b, Ano08c, Ano09c, Ano09d, Ano09b, BBC^{+}01a, BDL00, DH01, FNR^{+}07, FGV01]. Physics [Fij00, GDAG05a, KM01b, Lan07, MOS01, New07, Ram10, SM06a, TS08, Wu10, ABM03, ASJ^{+}03, Ano04b, Att09, BvG02, Bor07, Bra05, CRS01, CMR01, Cra01, Esq02, FS00, Fod05, GT01, GGL^{+}02, GPW04, Gou00, Gre04, GMO03, HF00, HS01b, JS08, KB02, Kud09, LVV09, LL07, Mah09b, Mah09a, MT01, MSK^{+}02, Nov02, OLX07, Pin01, RM05a, Rin02, Sak07, SEF^{+}01, SAG^{+}02, Suz00, Swe02, SC04, TA00a, TA00b, Tho01, Yan03c, BMS^{+}09]. PHYsics/scale [BMS^{+}09]. PhysicsGP [CB05]. physiological [ZS07, ZS08]. PIC [AH03, CSC^{+}04, DN04, DBR^{+}02, JBA^{+}07, LSL07, LKPH08, MIM^{+}07, SLL01, SG04b, WCG04, WHL^{+}07, YRR07]. PIC-DSMC [CSC^{+}04]. PIC-FEM [WHL^{+}07]. PIC-hydrodynamic [LKPH08]. PIC-MCC [LSL07, MIM^{+}07]. picket [LC08a]. picture [KS04a]. Piecewise [Ram04, Dem03, IH09, IHAR09, ZA01]. piecewise-linearized [IH09, IHAR09]. Piezoelectric [HM06a]. piezothermoelastic [Mel01]. pilot [AAKL07]. PIMC [Müs02a]. PIMD [Müs02a]. pinch [Pet04, RG04]. pinches [SBL^{+}04]. pioneering [Ano04b]. pipe [QR01, QG04]. pipeline [SIE04]. pipes [DM09]. Pitaevski [CPS00]. Pitaevskii [DC07, MA09, TQZM08, TS06]. pitchfork [Bal01]. pitfall [Nap09]. pits [NY06]. pixel [QTMH07]. planar [MDT03, Var02, YKK07]. Planck [ABSM04, CBKM01, KA04, yMS01]. plane [ADDdM07, BK05a, BVKW02, CR05, DEW01, GSS06, GBD03, HTA08, MMB02, MSHP02, MSHP20, THM01, VKM^{+}05]. plane-wave [CR05, MSHP02, MSHP20]. planes [MS05b]. planewave [ADS06]. plans [McK07]. Plaquette [Voi02, Voi03]. Plasma [MSY07, MIM^{+}07, Ano04b, BBBR04, Bre07, CHL^{+}07, CLL^{+}07, CH09, Del03, DBE^{+}04, DUX^{+}09, Gre04, HYY07, HKPL07, HJZL07, HL05, HHWH07, ISS^{+}02, IDS^{+}04, IL07, JKCGJ08, JTS^{+}06, KPL07, KKSR04, KCR07, KV07, Kon01, KTG04a, KTG04b, KZS^{+}00, KDSB04, KT07, Lee04, LKPH08, Liu07b, LS05, LC01a, MV04, MCL05, OLX07, PHKL02, Pin01, PSK01a, PSK01b, RLRR06, Ram10, SBM^{+}04, SGF04, SOAW08, SHJ07, SSB04, SZ04, SBBM04, STK^{+}00, TLCS04, TAM04, UXD^{+}09, VAH04, VCF^{+}04, VBFD01, VSBD00, WSB04, WML^{+}05, WRC^{+}04, Yan09, YRR07, ZSK^{+}04]. plasma-edge [SBM^{+}04]. plasma-wall [HYY07, KT07, MV04, SZ04]. plasma-wave [MCL05]. PLASMAKIN [Pin01]. plasmas [ATIO06, ATF^{+}09, ABSM04, ASC^{+}05, BF04, BDBV12, DGV08, GFP00, GBFS07, GH01, HD04, HOI04, HW09, KY07, KA04, KMR^{+}09, KSSH04, LYL07, yMS01, Mah08a, Man04, NYH04, PPP01, PCV06, SV01, SG01, TPBE04, TKP06, TDD04]. plasmastatics [Bru00b]. plasmoid [SKRK04]. plastic [SM06b]. plastically [Cle05]. plate [Var02]. platform [AAKL07, BAD01, Far01, KKHL07]. platforms [CR00]. Plato [KH09]. PLD [SMS^{+}00]. PLNoise [Mil06, Mil07]. plotting [NY06]. plugin [BBB^{+}09b]. plume [CSC^{+}04, KTG04b]. plume-to-spacecraft [KTG04b]. PLUMED [BBB^{+}09b]. plus [AIOST03, HSGBK08, LMP^{+}09]. PMCD [MP03]. PML [VAH04]. PMS [CFH^{+}01]. Podolsky [DDM07]. Point [KBG00, Tör00, AGJJ07, BCV03, DS04, HDG07, MPR05, NN09, RF05b, RF06b, Str00, TNI^{+}07, TMN01]. Point-centered [KBG00]. points [FBB01, HP02, KSS02, She08]. Poisson [KHÖ01, AS00, Con04, DHB^{+}04, Dys02, Eli05, HCH^{+}06, LdVJ06, Li03, LY05, MS05a, NT04, NJ01, QR01, QG04, TPYV03, XON08, Zie04]. polar [GVMW04, SGL09]. polarisabilities [QCL05]. polarizable [DDD^{+}01, SLBG09]. polarization [MPR05, YvG05]. polarized [ACIZ07, NY07, Vog05]. POLE [Con04]. polyalanine [YD06]. polyanions [SSH01]. polyatomic [BSS09, RMLB01]. Polychromatic [BCP04]. polycrystalline [SKH02b]. polycyclic [EYJ07]. PolydGpdCp [Ger07]. polydisperse [CGG00, SWY01, YW00]. polyelectrolytes [LH02]. polyethylene [FAiTD01, Ryc05]. polylogarithms [GR01, GR02, Maî06, VW05]. polymer [BMML05, CW01, FS01b, FS02, KMB02, KSEG05, LS02, LOY07, LS09, MB05a, Mül02, MPS09, PRSB08, ULA^{+}02, WBC^{+}07]. polymeric [Krö05]. polymerization [BJ02, ISS^{+}02, MSH01]. polymers [BLS09a, CCFG05, CNDC09, CCRA05, DCNDC09, GGL^{+}02, LC07, TiTD01, The05, VYK02, vdHBP^{+}02]. polynomial [ASF^{+}05, KTL05, KTT02, PSH06, UCG^{+}05, UNK12, Vak00]. polynomials [HLB06, KT04, Str00]. polystyrene [LM02b]. polytope [vHK00]. Polytopes [KS04b]. Pomwig [CF02]. Poor [CFH^{+}01, LH02]. PopRatio [SV01]. population [VPP^{+}12]. population-based [VPP^{+}12]. populations [SV01]. pores [BDHP08, DN05]. porous [BBD00, JOS07, NSYZ02, PPM04]. portable [BBB^{+}09b, GDC01, LL00, OCK^{+}03, Ref00, SKNV04]. portal [BLCR05]. Porting [EL04]. Pöschl [MS08b]. posed [RMWH01]. positive [FM03, LCPC04, SJP05, Sea02a]. positron [BPP01, SBM09b, WCBN05]. positrons [SJP05]. possessing [PSK01a]. possibilities [McK07]. possible [TIM07, TIM08, Var02, Vul03]. post [Pue06]. post-Newtonian [Pue06]. postprocess [BC07]. postprocessing [LB09]. potassium [KACB07, YN05b]. potential [APV00, ATP01, ADS06, Ber03b, BFLW07, CCBL02, CGG^{+}08, CGG^{+}09, CW01, DVL^{+}02, DVL^{+}04, FAiTD01, HG02b, Hin00, IK00, LRI^{+}06, LPC^{+}00, LF02b, MS08b, MAM04, MAM07, OS03, PJK00, PAT^{+}09, Riz02, SN07, SG04b, TAP01, TT06, TYSH05, WL00, XSC09, Zak06]. potentials [APV00, ASH06, AMP^{+}00, BVY05, CW00, FHR^{+}05, HSS^{+}08, IBM03, KM08b, MBG03, NW02a, ON08, OPO^{+}08, PS08, SŽ00a, Sea02a, SSLN02, TKN^{+}08, Val05, Vie01]. POTHMF [CGG^{+}09, CGG^{+}08]. POTLIB [DVL^{+}04, DVL^{+}02, DVL^{+}04]. Potts [BBJS09, BKB02a, BKB02b, CBBJ02, CGIA07, HJ02, KSS02, TL09, dO09]. power [Mil06, Mil07, NV09, RDSS01a]. power-law [Mil06, Mil07]. pp [Wan00]. PPA [TKK^{+}06]. PPA 4b [TSA^{+}03]. Practical [FJC^{+}05]. pre [Ano01n, Elm09]. pre-attentive [Ano01n]. pre-equilibrium [Elm09]. Precise [Mic07, PR06, TI01, Bru00a, CCGR09, HTNFBS06a, HTNFBS06b, KF05a, SW09, Zak06]. Precision [CCG08, BBD^{+}09, FS01a, HDG07, JWW00b, KS05, LMC^{+}03, TNBSF04]. Preconditioned [GHP01, Lüs05, Xia01]. preconditioner [HZGZ09]. preconditioners [CHS09, SBD^{+}06]. preconditioning [ADG08, GH00, UJSW06]. predator [TRAdO09]. predator-prey [TRAdO09]. predict [Gha05]. Prediction [TiTD01, BK05c, GOH06, HCH^{+}06, KPF03, SvAS01]. predictions [BL00, Bre05, CSC^{+}07, CSC^{+}08, GPW04, Oka01]. Predictive [NK07]. Preface [Ano00-27, Ano01-30, Ano04-47, AEK02, BDL00, FMP05, Gia02, GCI01, KL07b, LF02a, MG08a, PR01]. preliminary [BK01]. Prelle [DDdMS02]. preparation [Cap05]. presence [KDSB04]. present [Ano02a]. Presentation [Ano04a]. preserving [CLR08, HL00c, LB04]. pressure [BBD00, CS07, CHM^{+}09, HTA08, HJZL07, IL07, LHS^{+}09, Lei02, LDZ^{+}08, MLG^{+}01, MC09, Mor01, NV09, PDM^{+}08, QP05, Var02]. pressure-dependent [MLG^{+}01]. pressures [BNS07, KRTZ02]. prey [TRAdO09]. primordial [PCE^{+}08]. principal [HB05]. principle [RG05, Tsa02]. principles [AJT^{+}07, Ano09a, CR05, CM02a, CTI07, EYJ07, FG04, GBTM07, Har01, KKKC07, LN01, LDZ^{+}08, MCBR03, MSK^{+}05, Mor01, NKSL05, WKP^{+}01, WC00, dSdSW08, vdHBP^{+}02]. priori [DVG05, TIM07, TIM08]. Prize [Ano04-56, Ano04-57, Ano04b]. PRMAT [SNBB02]. Probability [Lik01, Man04, BH08, FPB08, FFD00, RF08, Sev00, SSZ01]. probe [CS07, NSKS01]. problem [AMP^{+}00, Bae03, Bal07, BD08, BL00, BV00, Bru00a, CRUV00, CGVA09b, FGF03, GHP01, HBW05, Huj05, JC07, JS08, KNU00, KZS^{+}00, LVV07, LC00, Man02, MMR04, MNH01, MP01b, SHV^{+}01, SZ00b, She08, SGM^{+}09, TNI^{+}07, VBFM05, Wan06b, WWF08]. problem-orientable [Huj05]. Problems [IHAR09, ASJ^{+}03, ASVA00, AKS01, BJ05b, BKM02, CFKM01, CL03, DSH03, DTHL09, FS00, FBL00, Fra07b, HCIK00, Huk02, KSTL03, LVV09, LMC^{+}03, LCHJ09, LJ09a, MT01, MVJ09, MLF07, OS00a, PAS09, Ram04, Ram05, RM05a, RMWH01, SVA01, SVA03, Sim08, Van05c, Var08, WW05, Wan05b, Wen01, WDHE04, Zim02, Zim05, dA08]. procedure [Fat02, IF03, LVV07, MVS05, MC09, PRBD09, Tam03]. procedures [FIBT01, FIT03, Fri09, GFF01, GF02a, GSF05, IFF01, PFG06a, TLP04]. Process [BFL^{+}01, PK01, BDT00, Con04, GDC01, IF03, NFS01a, QTMH07, SVP09, SOAW08, ZSdD^{+}08]. processes [AS00, ABB^{+}09, Bar00, CPW09, CZC00, Ida02, KDSB04, MER^{+}00, NFS01b, RS00, WSB04, Wen01, WCG04]. Processing [LSVMW08, AGS07, BBB^{+}01, CDD08, CR00, DDMM06, Di 01, EFG^{+}00, FEHC01, MIM^{+}07, Rap06]. processor [CGK^{+}00, De 07, MBKJ09, PKB^{+}01, REAB09, SHT08, vDGM^{+}09]. processors [BOG^{+}07, CR05, Far01, Oli01, ULA^{+}02, ZA01]. PROCRUSTES [Pue06]. produced [GFP00, HD04]. product [Kim03, Tos08, WS09b]. production [Abe01, Ano03h, AAC^{+}06, BBC^{+}01b, BvG02, CDEW04, CWW06b, CWW07, DDRW03, JWW00a, JPS^{+}01a, JPS^{+}01b, Kol03, KFI^{+}01, Por03, Tom09, TSA^{+}03]. products [GRR01, VC08]. PROFESS [HLC08]. profile [CP00, KNY05, RLRR06]. profiles [BS08, KMR^{+}09, NY06, ZDKG05]. progeny [LC01b]. Program [Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano01-35, Ano01-36, Ano01-37, Ano01-38, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano02-29, Ano02-30, Ano03-36, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Ano03-42, Ano04-48, Ano04-49, Ano04-50, Ano04-51, Ano04-52, Ano04-53, Ano04-54, Ano04-55, Ano05-46, Ano05-47, Ano05-48, Ano05-49, Ano05-50, Ano05-51, Ano05-52, Ano05-53, Ano06-28, BB09a, BJS00, BB09b, GFG01, GLHW01, KW08, APV00, AP04, AdlT03, Alf09, All02, AOT01, AJ08, AC05b, ASS^{+}02, AAC^{+}06, Bar04, BBC^{+}01b, BS03, BDW06, BBPS02, BBPS07a, BBPS07b, BCP04, BD05, BCG03, BDB^{+}08, BGH^{+}09a, BRdAHK04a, BBJW05, BM04, BKM05, BOPC05, BD06, CCGR09, CP00, CGG00, CD01a, Cha07, CRUV00, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, CGVA09b, CS02, Cip07, Cip08, Con04]. program [DS06, Dan07, Dan09a, Dan09b, DDM07, DDRW03, Dev05, DJ08, DD00, DO04, DSC^{+}09, DSS01, Dra01, Dy09, EÅ01, EÅU05, EH06, Fel08, FT08, FBL00, FFD00, FFG02, GFG^{+}06, GSM^{+}03, GS01b, Gro01, GG03, GNZ^{+}09, HHH^{+}09, HS03, HC08, HHW00, HLC08, HB05, Hor09, IW01, IW02, JWW00a, JPS^{+}01b, JK08, JU09, JG02, JC01, KTT09, KS84, KS08, Kol03, KJ04, Kol09, Kon02, Mah08b, Mah09a, MCLDP01, MFF^{+}05, MR06, MW01, MMR04, MSK^{+}05, MPK00, MFVJ07, MS09, NY06, NY08, OK06a, dlRBPL09, PS08, PKRK07, Pit05, PAT^{+}09, Por03, Ref00, RSD01, Riz02, Sar00, SV01, SCM00, SGL09, SYM00, Ste01, SDNR05, SJF07, SNBB02, TKB^{+}04, TV07, TS06, TNBSF04, TNCG00, Tót08, UCG^{+}05, VCCS05, VS06, WGDZ04, WP00, WW06, ZF00]. program [ZDKG05, CGG^{+}09, CGVA09a]. programmable [KPD06]. Programming [CB05, LL07, TS08, BDK^{+}06, BSO^{+}04, Chr00, Iwa01, MRF^{+}05, Niu00, SHH^{+}04, WMK09, ZA01, Hoo04]. Programs [BRD04, FH04, BC07, JKW00, JKW06, MA09, PSW00, Ver00]. Progress [DSL09, OS00a, NP01a]. Project [BCP04, Yos00, CFH^{+}01, BHNW01, GI01, Mak01]. projection [DTHL09, MI05, Rob01, SFSL09]. projector [BVKW02, FM00, HTM01, THM01]. projectors [RGD^{+}01]. prolate [Hua09, KJ07, LKC06]. prolates [LB09]. Prompt [Teh01, VT00a]. proof [BLS01]. proof-of-concept [BLS01]. propagating [Mah08a]. propagation [BS04b, BBBR04, EM08, FW01, HGH^{+}05, HJZL07, HL05, JC08a, JTS^{+}06, KV07, LCB07, MN01, MP01b, NN06, SLMS06, SSB^{+}09, SWP03]. Propagator [Bow02, BH07, CA07, WP06]. propagators [FJC^{+}05, Ixa07a]. propBG [CP00]. propelled [BA09]. properly [MMMM00]. properties [ADS06, Ano09a, BM01, CIC^{+}03, CTSZ07, Dup01, FKMB09, GZDA01, JRT00, KMD^{+}02, Kar02, KFB01, Lee04, LH02, MSB09, NKSL05, PMH08, PJSK08, RG05, RS09, SG00a, SM04, SM06a, SBM09a, SGK09, TiTD01, TY01, Vie01, WRN01, YC07]. property [CTI07]. proposal [Bre05]. protease [CRPC08]. Protein [DLZ08, Oka01, iTKST01, DELG05, Elb05, ISH01, LV08, MMEH08, NSMO02, SSA07, SHH^{+}04, WMNS09, WL08]. Protein-DNA [DLZ08]. protein-g [ISH01]. proteins [BDH^{+}02, DC05a, EHHH01, EHHH06, LPC^{+}00, MMEH08, SBJ05]. protocols [RDSS01b]. proton [CRS05]. protonated [GF02c]. pseudo [CCBL02, CMK^{+}03, EVL00, HDG07, IBM03, KBV09, ON08, PL05, Bru04]. pseudo-arclength [KBV09]. pseudo-dynamics [PL05]. pseudo-fermions [CMK^{+}03]. pseudo-potentials [IBM03, ON08]. pseudo-spectral [CCBL02, EVL00, HDG07]. pseudobinary [Bar03]. pseudopotentials [HP06]. pseudorandom [DH00]. Pseudospectral [LKC06, ABOSPG09, ICT01, NJ01, YZW02]. Pt [PLL07]. Public [PCE^{+}08]. Publisher [Ano01-39, Ano03-43, Ano03-44, Ano06-29, Ano07-29, Ano07-30, Ohl04, Pub07]. publishing [GDC01]. pulse [Gha05, KDSB04]. pulsed [CM02b, KHIL07]. pulses [BBBR04, BDV04, HW09, KS04a, Kur02, NY07]. PUPIL [THC^{+}07]. Pure [GF02b, CM02b, GFG01, Nap09]. purpose [ASS^{+}02, CPT^{+}01, FS00, FWP01, IOM00, Jad00, Jad03, Mak01, OK06a, iSHS^{+}08, SMS^{+}00, SHI02, SHH^{+}04, SIE04]. Pushing [Sor02]. PWBA [Cip07, Cip08, Cip09]. pwpaw [THM01]. PWT [KT07]. pyramidal [WHJ06]. pyroclastic [COE^{+}05]. PYTHIA [KRW03, SMS08, SS02a, BLS01, SEF^{+}01]. Python [Bor07, MMEH08, Nil07b, Nil07a].

Q [FKG00]. Q-Chem [FKG00]. Q2R [Sta00]. QCD [All01, BR01, CC04, CAF^{+}03, CvdEF^{+}05, CKA^{+}09, DS04, Dür09, EFH^{+}07, FKP03, FJC^{+}05, FHW^{+}01, GL02, GHLW03, GH00, HS03, HvHHM09, JS06, JPS^{+}09, JU09, KKK06, KSS02, Lüs04, Lüs05, MSS^{+}07, OS04, Vog05, Yos01]. QCD- [Vog05]. QCDd [vdEFL^{+}02]. QCDINS [RS00]. QCDOC [FMD07, FDM07]. QCMPI [TJD09]. QDENSITY [JDBT06, JDBT09]. QM [GSM^{+}03]. QM/MM [GSM^{+}03]. QMC [FM00]. Qprop [BK06a]. QQ [DGSL09]. QQ-onia [DGSL09]. Quadratic [Zah05, LLY07, ZA01]. Quadrature [Kim03, AA01a, AAP03, FKAM05, Hin00, IP01, KTT09, KD09, MKK05]. quadratures [Del08]. Quality [FGV01, KMZZ05]. Quantification [ISS^{+}02]. Quantitative [HF06]. quantities [Blü04, GFF01, GG03, MS06]. Quantized [SVS01, Har00]. quantizing [Zha01]. Quantum [AGS07, BDT00, Bes02, Bro00, DHMD00, DC05b, GPW^{+}09, JDBT06, KK01, KS04a, LJ01, Nil07b, PMV02, SYN01, SHT08, VK09a, Voi02, Voi03, Yep02, ZKASS05, ABNÅ05, ABD^{+}05, Bac02, BDLT02, BC05, BTK^{+}02, Bow02, BDM09, BNSY02, CN01, CRPC08, CA09, CHP04, CA07, DDM05, DDMM06, DMD^{+}07, DD01, DB08, Dun05, EMJH03a, EMJH03b, Fer07a, FSK04, FHR^{+}05, GCK02, GF02c, Gol00, GPT08, Haf07, HSSA01, HSGBK08, HL00b, HB05, JC07, JC08b, JDBT09, KM05, KM01d, KM03, KN07b, LVLS01, LLV^{+}01, LLT^{+}02, LCV06, LLLZ01, MSS00, Mam08, MP08, MWA01, MGG05, Mey02, MI05, MK05, MA06, Moh08, MP05, MM01, OSK^{+}02, OMF03, OGKL02, PZW^{+}00, PRBD09, PCYC02, RF05a, RF06a, RF07, RF08, Ram03, RDFF02, RS09, RCG05, SCM00, SGL09, Suc02, Suz00, TJD09, TYSH05]. quantum [TNCG00, TDD04, Tót08, VK09b, VT00b, VT00c, Vos06, WHJ06, WWF08, WDHE04, WV05, Wil09, YB02b, dSL02, RF05a, RF07, Vio04, Wan00]. quantum-number [MI05]. quark [BH07, CKS00, JWW00a, Kol03, OvSA02, TSA^{+}03]. quarkonium [DGSL09]. quarks [BDF^{+}08, KMP09]. Quarteroni [Sha04]. quartic [TY01, TYSH05, YT01b]. quartz [HM06a]. Quasi [Sch04, BFL04, HTM^{+}08, KL06, LDZ^{+}08, MT00, PSH06]. quasi-bound [MT00]. quasi-error [KL06]. quasi-harmonic [BFL04, LDZ^{+}08]. Quasi-Monte [KL06, Sch04]. quasi-polynomial [PSH06]. quasi-temperature [HTM^{+}08]. quasicrystals [Gro01]. quasilinear [FMG00, Kon01]. Quasilinearization [KM03, KM06, KM08b, MT01, KM01d, Ram04]. quasiperiodic [HL00c]. quaternary [Bur02]. quaternionic [JC07, JC08b, WWF08]. qubit [RF05a, RF06a, RF07, RF08]. QUBIT4MATLAB [Tót08]. qubits [PMV02]. quenching [RCGC00]. Quickstep [VKM^{+}05]. quiet [ZSdD^{+}08]. QWalk [MP08].

R [PKKM02, Ton07]. R-CCSD [PKKM02]. R-matrix [Ton07]. R1 [HKM^{+}07]. Racah [Fri09]. Rachford [Mah08a]. RacoonWW1.3 [DDRW03]. radar [FKMB09]. RADCAP [Ber03b]. Radial [OIKN02, WDB04, CGA^{+}07, CGVA08, CGG^{+}08, CGG^{+}09, CGVA09a, DGSL09, Dy09, HB05, KTT09, SMSE03, Sea02b, Sim09, TD03, UK02a]. Radiation [MK05, AGM^{+}00, BP08a, Leh00, Maz00, PP09, RtVR09, RMK05, SVP09, SJHY07, WSB04, WCBN05, dA08]. radiative [ACIZ07, HD04, HvDJvdM01, Huj05, TI01, WH00, Wol03, WH05]. radio [GB05]. radioactive [VS06]. radiography [KMCS01]. radiological [FFS01]. radix [Yam00]. radix-2 [Yam00]. radon [LC01b]. RAEEM [LL04]. railway [HMY^{+}02]. Raman [HS07, HW09]. ramifications [Luo00]. ramified [SFSH01]. Random [DGLB08, GG00, LBP^{+}09, SFSH01, ATB^{+}01, Bel05, BBJW05, CLFH07, CFJ09, CCRA05, FV02, HSJ02, HJ02, JH09b, Lad09, LCPC04, OTY02, Pro00, RK05, Sch06a, SYN01, SSZ01, TL06b, WHO02, WH06]. random-bond [HJ02]. range [AL08a, BHL02, CM02a, CJC09, LOY07, MBG03, MRS04, Mic07, NH09, PHF^{+}07, RD05, SYN01, SS05, Tat07, Val05]. ranged [CW00, WHL05]. ranging [MOC03, OMC00]. ranlip [Bel05]. Rao [Kas00]. Raphson [Jam00]. Raphson/log [Jam00]. Raphson/log-derivative [Jam00]. Rapid [RB08, ZZ09]. rapidity [JPS^{+}09]. rare [GCP^{+}02, GF02c, NW02a]. rare-gas [GCP^{+}02, GF02c]. ratchets [Rap02b]. Rate [MLG^{+}01, BGJ^{+}07, BBPS09, EST00, KK00, TSI02, ZS07]. Rate-based [MLG^{+}01]. RATH [bLpL02]. ratio [HS03, QG04, UVLRRC09]. rational [DR09, SK08, VC08]. ratios [BBJS09, CFJ09]. Ratip [Fri01, KF05b, NFH06]. ray [Min01, MKJ^{+}05, NRR01, Pop03, AGM^{+}00, BB07, BSO^{+}04, BG01, KMCS01, LZC^{+}08, Sal03, Vég04]. Rayleigh [DMR01, DMR02]. Rays [Tol02]. RBF [TWY09]. RBF/Elman [TWY09]. REACH [MS09]. reaction [BS00b, CZC00, CRPC08, CGA^{+}07, CGVA08, CGVA09a, DCJ07, MLG^{+}01, XD08, vE08]. reaction-diffusion [BS00b, CZC00]. reactions [Ber03b, BG06, BTS05, Elm09, HYY07, SK05, Sri01]. Reactive [RFK08, Val05, FH00, HSGBK08, LCB^{+}00, LJ01, MGG08, MN01, iNKNV08, SCM00, VT00b]. reactor [KPL07]. reactors [STK^{+}00]. ready [BAD01]. Real [KM08a, MGG08, BMSG01, Bun01a, DM07, FH04, FPB08, HSGBK08, HCO00, ICO03, MSHP02, MSHP20, MKM02, OSK^{+}02, SCO00, SKNV01, SKNV05, SMH^{+}01, Teh01, TB87, Tho04a]. real-coded [HCO00, ICO03, SCO00]. real-space [OSK^{+}02, SKNV01, SKNV05]. real-symmetric [Bun01a]. Real-time [KM08a]. realistic [CYAS05, GCP^{+}02, HKK02a, ZKASS05]. reality [TKS^{+}01]. reason [BNSY02]. reasoning [Vég04]. recall [MHS05]. Recipes [Koc02]. reciprocity [GPT08, ISSB01]. Recognition [SKH02b, DLZ08]. recoil [Nat08]. recoiling [Wro08]. recombination [PNH00]. reconfigurable [Fra07a]. reconnection [BJ03, EFBP04, FS08, GBC^{+}04, HOI04, UTO09]. reconstructing [BV00]. Reconstruction [Bat03, BG01, GGQ01, ISSC01, SF06, Teh01]. record [DH01]. recoupling [DK05, VF03a, VF03b, FIBT01]. recovery [ZS08]. rectangular [DM09]. recurrence [SHI02]. Recursion [LZ04]. recursive [KKS04, KTT02, MC03]. REDACLE [BCC^{+}06]. redox [BTS05]. REDUCE [GKI02, GKI04, UCG^{+}05, Vul03]. Reduced [Bac02, NGE^{+}04, GF01, GBC^{+}04, RLH^{+}09]. Reduced-dimensionality [Bac02]. Reducing [HKP02, Bae03]. reduction [BGJ^{+}07, Har00, MG09a, MG09b, NP00, Pis00, Wei02a]. redundancy [Man02]. reevaluation [TSI02]. reference [LRI^{+}06]. Refinement [KNTG03, PM01, FS08, HCH^{+}06, KKF^{+}04, LOL06, MOM^{+}00, Ros04, SJCM04, VAH04, VCF^{+}04, WTW04, Zie08]. refining [LHS^{+}06]. reflect [Nur04]. reflectance [SLC09]. Reflection [KV07, Ram10, WP00, Yan09]. reflections [Hib01]. reflective [CLFH07]. regaining [ZSdD^{+}08]. regarding [Ano04-56, Ano04-57]. regeneration [KL01]. regime [CMS04, GHLW03, HS03, HW09]. regimes [YM03]. region [MNYY00a]. Regional [ADE^{+}02, Org01]. registers [RF05a]. regular [BSDMH05]. Regularization [AG05, BK06b, DSHH05, RMWH01]. regularized [Cai09, DSH02, DSH03]. reinforcement [EM08]. related [ASVA00, AKS01, Lee04, SVA03]. Relating [SSA07]. relation [HJM02, KT07, LWY01, MSD08, Sus01]. relations [Blü04, Blü09, QCML03]. Relativistic [KF03, KF05b, OvSA02, AKZ00, AMP^{+}00, AT09, BD08, FFD00, FFG02, GZF04, JHFG07, Kon02, LKPH08, LS01, MMR04, MK05, ON08, PFG06b, She03, TP01, vdHKM08]. relativity [AG05, MG09c, Pue06]. relaxation [FFD00, LNK01, TCF00]. relaxing [Huk02]. RELCI [FFG02]. release [BCKT09, GKP^{+}06]. relevant [WML^{+}05]. Reliability [AAP03, AA01a]. Reliable [AA00, AA01b, Ixa01, CCRA05]. relic [BBPS02, BBPS07a, BBPS07b]. Remarks [Ano04-56, Ano04-57]. remembrance [Ano04-45]. remote [BCD^{+}01, BK06b, DSH02, DSH03, DSHH05, FGV01, SEC04b]. removal [Hor09]. Renner [HC08]. renormalization [Alv09, CC04, FLO06, MI05, WN01, Zit09]. renormalization-group [WN01]. renormalized [PKKM02]. reordering [TC06]. reorthogonalized [BSTC05]. REOS99 [FFD00]. Replica [PLS09]. Replica-exchange [PLS09]. replicas [HS01b]. Reply [AA01b, LHC02, WLW04]. represent [FA00]. representation [BDBV12, DBR^{+}02, KKK06, Mas05, SHW01]. representations [De 02, GKR07]. representing [GFG^{+}06]. repressor [CDFF05]. repressor-DNA [CDFF05]. repulsive [DKC08, LMM^{+}08, Sea02a]. required [ADE^{+}02]. research [ADE^{+}02, LAMH06, TWY09]. residue [Nat08]. Resistive [SG01, Pet04, SPP^{+}04]. Resolution [BW08, ABRS12, BVY05, BADC07, GKM^{+}00, JC01, KKF^{+}04, MMTH04, Ros04, WBDB04, ZDKG05]. resolve [YNK05]. resonance [BvG02, CLL^{+}07, HD04, KS01, LCS07, VT00c, ZPB09]. resonances [BS00a, KBV09, KM00b, MMMM00, SBM09b, TSB^{+}05]. resonant [Sal03, Wan05a]. Resonating [CYAS05]. resources [ADE^{+}02, Ano09s, BLM01, CDH^{+}06]. respect [AS03, CGVA09b]. response [AK03, FWP01, Liu07b, NP01b, NM01b, YG09, ZS07, Zha01]. resulting [Inu07, VS06]. results [All01, ACC^{+}01, HDG07, Luo00, MCL05, MP05, Yos01, You05]. retarding [SG04b]. retirement [Bur01]. retrieval [OPB^{+}09]. Reverse [OPO^{+}08]. reversibility [ISSB01]. reversible [Sta00]. Review [Bre01, Hoo04, Koc02, Laf03, Par04, Sha04, Vio04, Wan00, Ano00a]. Revised [NFH06, AH03, FFD00]. revision [Ano00z, SM04, SM06a, SBM09a]. revisited [LIR^{+}06]. REVLD [BLS09a]. Reweighting [dSL02, VMMB02]. RF [Eli08, WJW09]. rhad [HS03]. RHD [DPSG06]. RHEED [BD06, Dan05a, Dan05b, DS06, Dan07]. RHEEDGr [Dan09a]. rheological [DGR09, TiTD01]. rheology [HCO01, MDT03]. RHIC [BNO^{+}01]. Riccati [IH09, Yan03d]. Richard [Koc02]. Rideal [LJ01]. Riegeom [Por00]. Riemann [MGYP08, SWFL00]. right [BV00]. right-hand [BV00]. Rigid [EE02, CCBL02, HSS^{+}08, VPNW02]. Rigid-body [EE02]. ring [BJ02, Man02, NY07]. ring-opening [BJ02]. ring-shaped [NY07]. rings [Man02]. Risebro [Laf03]. rising [WGS00]. Rjaseekar [Par04]. RKN [Wu10, Fra07b, YWYF09]. RKN-type [Wu10, YWYF09]. RLW [Zak01]. Robust [GKM^{+}00, Tót06]. rods [JBS08]. role [AFK^{+}07, BK01, CRS01]. Rome [Org01]. ROOT [ADD^{+}03, Ano09t, WCH09]. Roothaan [MW01]. Rosen [DDM07]. Ross [Bat03]. Rostoker [SJ02]. rotating [ATF^{+}09, CC07, TQ03, Yur02, ZZ09]. rotation [Goc04, KLTH04, LVH07, MT00, PFG06a, TKB^{+}04]. rotation-vibration [TKB^{+}04]. rotation-vibrational [MT00]. rotator [Bow02]. rotor [BBOY08, CCBL02]. Rototranslational [QCL05]. round [JC08a, WN01]. round-off [JC08a]. Routes [SBD^{+}05, VEG08]. rovibrational [Bac00, CCBL02, CNMC09]. rovibrationally [LDBG08]. RT3 [HC08]. rule [FKAM05, IP01]. rules [CD09b, HvHHM09, Kim03, QCL05, Sem09]. Run [ABC^{+}01, BGLLW01]. runaway [EH03, SMSE03]. RunDec [CKS00]. Runge [VAMVR08, ASVA00, BT01, CFMR08, Fra02, KMS09, MVJ09, PAS09, Van05a, Van05b, Van05c, VIV01]. RunMC [Che05]. RunMC-an [Che05]. running [CKS00]. Runwien [dlRL09]. Ruth [OMF02]. Ruth- [OMF02]. Rydberg [KB02, NW02a]. Rydberg-excited [NW02a].

S [Par04, MM08]. Saha [RMK05]. Saleri [Sha04]. salt [CCFG05]. sample [Nap09]. samples [Nap09]. Sampling [AL08b, GW01b, Asc08, BL05, LWLL07, PM02, SR01b, TBZ12, TMN01, VPNW02, WL08, Zim05, vE08]. SANCnews [BBK^{+}07]. SANCscope [AAB^{+}07, AAB^{+}06]. SANCscope-v.1.00 [AAB^{+}07, AAB^{+}06]. Sand [PAD^{+}09, OML09, EFG^{+}00]. satisfying [YSM09]. saturations [Lüt04]. saving [FNR^{+}06, FNR^{+}07]. Savitzky [MMMM00]. SaX [MSB09]. Saxon [MAM04, MAM07]. Scalability [TCY^{+}08, PHF^{+}07, SKNV01]. Scalable [LBP^{+}09, OCK^{+}03, OCK^{+}00, SKNV04, SOAW08, BDLT02, FKP03, FS01a, LL00, iNKNV08]. scalar [Fel08, FT08, FBB01, GF02b, NN09, Sav01]. scale [ABOSPG09, ABD^{+}05, BVY05, BMS^{+}09, DBE^{+}04, EH07, FK00, FTGG07, HHM^{+}09, HJM02, KMD^{+}02, LCB^{+}00, Lee04, LOCJ05, LM02b, MC08, MDC09, iNKNV08, OCK^{+}00, PFG06b, REAB08, Rap08, RCG05, SJ02, SAG^{+}02, Swi04, TPYV03, TIM07, TIM08, THC^{+}07, ULA^{+}02, UJSW06, Vor02]. scaled [NM03]. scales [HM00]. Scaling [CR05, BJ08, BM01, BS04b, BMG01, BSvdDW02, Cha00, CMR01, DGAG06, FG04, Gao03, GDAG05a, GDAG05b, HBW05, HP02, HHM^{+}09, KF05a, SKNV01, SKNV05, SJ02, WC00]. scan [Bre05]. scanning [Bre05]. scatterers [DW01]. Scattering [DEW01, KM00b, PP09, ACIZ07, BP08a, Bel01, BCG03, BG01, CGG00, Cha07, CAW00, CY01, Di 02, DEW00, DW01, FL01, GSSN00, IBM03, LCB^{+}00, LPRS02, LPR04, LJ01, LZ00, LDBG08, MBG03, MNYY00a, MNYY00b, MGG08, MFVJ07, MNH01, MN01, PCC01, Pis00, RMLB01, RS00, SSPM05, Sal03, SJP05, SBM09b, SCM00, SNBB02, TI01, Ton07, VT00b, WK02, WV04, ZDKG05]. SCELib [SM06a, SG00a, SM04, SBM09a]. SCELib2 [SM06a, SM04]. SCELib3.0 [SBM09a]. scenario [Huj05]. scenarios [NK07]. SCF [AKG02, Cha00, CWSH08, Dup01, FKG00, REAB08, dMBC^{+}06]. schedules [ZA01]. Scheme [DK05, AEEdR05, AC05a, Cap05, CPS00, Cha04, DPSG06, Esi01, FH00, GLL^{+}02, HDG07, HL00c, HS07, KG07, LBPS09, LMC^{+}03, LHS^{+}06, yMS01, MZB^{+}04, ML06, OK06b, PC08, ISX05, Sev00, TYN02, TNY00, UTO09, UNK12, WHJ06, WS09a, WTW04, ZZH09, Zie05]. schemes [BH03, BP08b, CMT00, CMT01, ID09, Suc02, TQ03, VCF^{+}04]. Schrödinger [ACK05, AKZ00, ASVA00, AKS01, AKS02, AK07, BK06a, CJK09, CMK^{+}03, DGSL09, GH00, GNZ^{+}09, Inu07, Ixa02, Ixa07b, JK08, JC08b, KMS09, KBV09, LVV04, LVV06, LRI^{+}06, LIR^{+}06, LVV07, LB00, LY05, LCB07, Nur04, PSV00, Riz02, RLV^{+}08, SŽ00a, SW09, Sim00, SW00a, SVA03, Sim09, SM02, SFSL09, Sug01, SQ03, UK02a, UYK^{+}04, Van05a, WGDZ04, Wan05a, WC05, WS09a, WT01, XSC09, XZ12, YB02b, Zak06, dlHV08]. Schrödinger-solver [BK06a]. Schur [CD01a]. Schwarz [Lüs05]. Schwarz-preconditioned [Lüs05]. Schwinger [AHS09, ÇHM00, Maa06]. Science [MRF^{+}05, BM02a, CSZ^{+}07, Gun02, Haf07, MS05b, OLX07, Tót08, Koc02]. sciences [Han00, SBM02]. scientific [BBD^{+}09, BC07, BD06, Cap05, Dan09a, Dan09b, Esq04, MSK^{+}02, MA04, MA08, NJ00, Nil07a, SJDC07, ZC09, Sha04]. Scientists [Bre01, Mal00]. Scilab [BBJ^{+}08, BFB^{+}09]. scintillation [RCGC00]. scintillator [FWP01]. scission [RLV^{+}08]. scission-neutron [RLV^{+}08]. scope [HHM^{+}09]. scrape [KY07, SBCZ08]. scrape-off [KY07, SBCZ08]. screened [HJZ09, OS03, SJ02]. screening [MLG^{+}01]. script [HL08b]. scripts [BS06a]. SDECAY [MDM05]. SDH [MBC^{+}09]. Search [GOG00, CCRA05, Nak07, TLP04, TL06b]. searches [VPP^{+}12]. Searching [Sus01, qX08]. seawater [VS06]. Second [MVJ09, Poi08, Poi09, BB04a, CIC^{+}03, FMG00, Goc04, Hoo04, WHJ06, YM03]. Second-order [MVJ09, Poi08, Poi09, WHJ06]. secret [AEEdR05]. section [AIOST03, Pap01, SBM09b]. sections [BS03, Cip07, Cip08, Cip09, HSGBK08, Hor09, Kol09, LDBG08, MOC03, Nik03, OMC00, Sal03, Yos03, Yos07]. Sector [BBK^{+}07, ST09, MN01]. Secure [DBE^{+}04, AEEdR05, TWY09]. security [LMC^{+}03]. sedimentation [BS08]. segmental [LM02b]. segments [HSS^{+}08]. SEL [GT04]. SELECTCONF [BKM05]. Selected [BDL00, BN07, DM07, TB87]. selection [BFMH^{+}01, BKM05, CB05, Man02, TS08]. Selective [SPV07]. selenium [NI01]. Self [BTI01, MMTH04, Pet04, SOS01, BR01, BA09, BNSY02, CD05, CHL05, CGVA09b, FK00, GGG01, Jad00, Jen01, MR06, NT05, NYH04, PHKL02, Pit05, SJHY07, SBBM04, SAU^{+}04, VKPB09, WLR^{+}08, WH00, ZSK^{+}04]. self-adapting [Jad00]. self-adjoined [CGVA09b]. self-amplified [SJHY07]. self-assembled [BNSY02]. self-assembly [NT05]. self-avoiding [Jen01]. Self-consistent [BTI01, PHKL02, WH00, ZSK^{+}04]. self-consistent-field [Pit05]. self-diffusion [WLR^{+}08]. self-energy [FK00, GGG01, MR06]. self-focusing [SBBM04]. self-gravitating [CD05, VKPB09]. Self-gravitational [MMTH04]. self-intersecting [BR01]. self-organisation [SAU^{+}04]. self-organization [NYH04]. Self-organized [SOS01]. self-propelled [BA09]. self-similarity [CHL05]. semi [AAC^{+}06, ADG08, BBC^{+}01b, BFB^{+}08, BGS^{+}04, CRS09, DDEM00, FBB01, LBPS09, ML06, ON08, UNK12, Ida02, TYN02]. semi-analytical [AAC^{+}06, BBC^{+}01b]. semi-classical [BFB^{+}08]. semi-core [ON08]. semi-implicit [ADG08, LBPS09, ML06]. semi-Lagrangian [BGS^{+}04, CRS09, ML06, UNK12, Ida02, TYN02]. semi-periodic [FBB01]. semi-structured [DDEM00]. semiclassical [TDD04]. Semiconductor [Hua09, GPT08, LVLS01, LLV^{+}01, LLCS01, LVLS02, Li03, LY05, LKC06, LCV06, LLLZ01, PMV02]. semiconductors [ABV02, BSO^{+}04, dS03]. semiflexible [LLPL08, NL07]. semileptonic [HL08b]. sensing [BK06b, DSH02, DSH03, DSHH05]. Sensitivity [RTS01, FGA04, KKS04, KMZZ05, RPY07, Sal02, TBZ12]. Separability [RF06a]. separable [OvSA02]. separated [AA08, Lei02]. separating [Pur02]. separation [BDHP08, EMJH03b, JBS08, KEM^{+}01, PSK01a, PSK01b, SGF04, Var02]. separators [Nat08]. separatrix [SPP^{+}04]. sequence [SIE04]. sequences [LCPC04, SHH^{+}04, SIE04]. sequential [TLP04]. serial [Ref00]. series [AOT01, BK05c, CR08, CN00, FD03, HJ02, Hon04, LVH07, LL04, Moh07, NP00, Sea02b, Wen01, WS09b, ZS07, ZWY04]. server [FEHC01]. service [AAKL07, KKHL07, LNV^{+}09]. Services [BJS00, AAM^{+}01, AAKL07, FGV01, Han00, ISSC01]. set [Di 01, ES09, GMAN^{+}07, KTT09, Maa06, Pit05, RLRR06, Str00, TS06, UYK^{+}04, Yok09]. sets [BD08, MBR01, TKN^{+}08]. seven [NR01]. SevenOperators [HL08b]. Sewing [BG09]. SFS [MTLC01]. shadowed [SZ00b, Sev00]. SHAKE [GWK09]. shallow [ML06, Sho04, Sho07]. Sham [AK03, MMR04, PAD07, WT01]. Shannon [CHL05]. shape [BD00, LM02a, LVLS01, OKS04, RCGC00, Zah00, Zah01, ZDKG05]. shape-truncation [Zah00, Zah01]. shaped [NY07]. shapes [GMBC08, JK01]. share [AEEdR05, TSB^{+}05]. shared [BOG^{+}07, BB00, CC00]. SHAREv2 [TJLR06]. Sharing [CPV^{+}08]. sharp [BDH^{+}02]. SHdecay [Bar04]. Shear [OMY05, AP05, DJ04, RSMK^{+}00, RR05, RvOvV02]. shear-induced [AP05]. Shearingbox [GZ07]. Shearingbox-implementation [GZ07]. sheath [NT04, SHJ07]. sheets [TÅT09]. shell [Cip07, Cip08, Cip09, ÇHM00, GSF05, KM00a, KM01b, SKF05]. SHELL2 [FK00]. shielding [MCC05]. shift [Ram10]. shift-operator [Ram10]. shifts [VEG08]. Shock [Wei02a]. short [BBBR04, CW00, KS04a, NFS01a, NFS02, NH09, Ram10, WHL05]. short-ranged [CW00, WHL05]. Shortest [Krö05]. shower [Wel01]. showers [EFG^{+}00]. Si [CW02, Kim07, LTT09, NSYZ02, SPC^{+}05, Sri01]. SiC [MCC05, RPD^{+}05]. side [BV00]. Sierpinski [SFSH01]. sieve [AA01a]. Sif [LK07]. sigma [GRR01]. sign [Bae03, BFLW07, vdEFL^{+}02]. Sign-function [vdEFL^{+}02]. signals [KV08, OS00b]. silica [MKB02]. silicon [Goe02, LN01, OPO^{+}08, SLC09]. SIMD [REAB08]. SIMDized [GKK^{+}08]. similar [SHH^{+}04, SIE04]. Similarity [VBC07, CHL05]. Simple [Bro07, dO02, JK08, MC09, MPS09, TPYV03, Tod01, Vég04, ZSSA00]. simplex [BSS09, WMNS09]. simplical [Jad00]. simplifications [Pog05]. simplified [MK02]. SimScience [WCGL00]. SIMUB [BS04a]. simulate [ABSM04, DDM07, Fra07a, GS01a, JU09, KKM02, Sal03]. Simulated [FHF00, BDG^{+}08, CEM08, Sch06a, SOS01, TL06a]. Simulating [BCC^{+}08, BM02b, CCD07, CMK^{+}03, MMEH08, Mül05, OGKL02, RRRHD08, Bae04, CKV04, CHP04, DHBE05, EE02, KP01, KNY05, MSK^{+}05, NY08, PY08, SWP03]. Simulation [BDHP08, BRB09, EH03, FSK04, HK02, HGH^{+}05, HTL^{+}03, HLW05, JH09b, JBA05, KTG04b, LYL07, LHMB00, MLPT08, OSK04, PJSK08, RF05a, RF06a, RF07, RF08, TIM08, TdFK00, Var02, VK09b, WGS00, WRMG05, YZD^{+}07, ZPB09, ZM00, ACIZ07, AdlT03, All05, AH03, Aok01, ASC^{+}05, BDK^{+}06, BCCM03, Bar00, BB04b, BDB^{+}08, BR09, BTS05, BMvG00, BDF^{+}08, BS08, Bur02, BDV04, CW02, CMS04, CGC^{+}09, Che05, CGIA07, CSC^{+}07, CSC^{+}08, CSC^{+}04, CAAM08, CL02, CTI07, CLL^{+}07, DDD^{+}01, DJ04, DMR02, DDM05, DVG05, DBE^{+}04, DSC06, EFBP04, EHHH01, EHHH06, Esi01, FMD07, FWP01, FL01, FS02, HD04, HYY07, Har01, HL00a, HMY^{+}02, HL00c, HJM02, HL05, Hua09, HKL^{+}07b, HSS^{+}08, HKK02a, HS07, IIK^{+}08, IDS^{+}04, IW01, IW02, ICO01, IN09, JAT03, Jen00, JDBT06, JDBT09]. simulation [JBS08, KH01, KPL07, Kar02, KSYE00, KPS^{+}01, KKSR04, KF05a, KCR07, KBG00, KK05, Kud09, KNSY07a, KMCS01, LM02a, Lee04, LSL07, LCS07, LbotMC01, Lei02, LVLS01, LLCS01, LLT^{+}02, LH03, LZC^{+}08, LC08b, LKPH08, LAMH06, LKC06, Lon07, LMS^{+}02, LL00, LM02b, MR05, MTLC01, Maz00, MVS05, MLF07, MPK00, MMB02, Mil06, Mil07, MKB02, MIM^{+}07, MYJY01, ML06, MC01, MS09, MABK02, MSH01, NSKS01, NT05, Nak08, NN06, NYH04, OSK^{+}02, OKS04, OK06a, OK06b, ÖDÇ02, PCK00, PGS02, PP02, PS09, Poi08, Poi09, PY08, Pop03, QTL06, Rap06, RCGC00, Ref00, RJFB08, RGR^{+}04, SNS01, SM06b, SLL07, SHZ01, SKNV05, SLL01, SOAW08, SSA07, SBBM04, SAU^{+}04, STK^{+}00, Swi04, TLCS04, TMTF00, TKS^{+}01, TiTD01, Tod01, TDD04, Trö08, VYK02, VBFD01, Wal03, WTH^{+}04, WHCL07, WML^{+}05, WRC^{+}04]. simulation [Wil02, WH05, WHL05, XON08, YSM09, YD07, YNS^{+}09, YRR07, Yos00, YvG05, ZLM04, dNKM07, TIM07]. Simulational [CMT00, CMT01]. Simulations [Bin02, HM00, LHS^{+}09, LNC^{+}03, RSMK^{+}00, Wes07, Ano04b, ADBF03, ABRS12, BS06a, BA09, BF04, BSB02, BB09a, BADC07, Ber02, Ber03a, BMS^{+}09, BCD^{+}07, BDBV12, BMML05, BHM^{+}07, BL05, BM01, BGH^{+}09b, BDYK04, BKB02b, BDM09, BGS^{+}04, BK05c, CZC00, CDFF05, COE^{+}05, CDQF07, CLFH07, CDD08, CR00, CW00, CNDC09, CW01, DS01, DMR01, De 07, DGLB08, DBR^{+}02, DD01, DC05b, DUX^{+}09, Elb05, EL06, Eli08, ES09, Esq04, FFK02, Fel08, FT08, FFF01, FS08, GCP^{+}02, GBFS07, GLHW01, Gre04, GH00, GCD06, GHPS04, Haf07, HL00b, Hef00, HM06a, HBMJ05, HKLY07, HKPL07, HW09, JGJ09, KBBW02, KCC^{+}00, KMD^{+}02, Kat02, Ker02, KSPT04, KY07, KM08a, KRTZ02, KD09, KK04, KSSH04, KNSY07b, Lad09, LTA05, LBPS09, LCM00, LSVMW08, LF02b, Lud02, Lüt04, Mak01]. simulations [MV04, MFF^{+}05, MBKJ09, MHS05, MMTH04, MS08a, MRF^{+}05, MP03, MFVJ07, MDC09, MER^{+}00, MTJ02, Müs02b, NSMO02, NGE^{+}04, NKV03, NBPG08, iNKNV08, NFS01b, OLS^{+}01, OD07, Oka01, OO05, OCK^{+}00, OMF03, PHF^{+}07, PM02, PRSB08, Pet04, QRH00, RP02, RLRR06, RD05, RPD^{+}05, RJCH00, RvOvV02, SJCM04, SPC^{+}05, SEC04a, SLWH02, SWC^{+}03, SBD^{+}06, SFF^{+}04, SBL^{+}04, SvAS01, SG04b, SBB03, SPM00, SPP^{+}04, SS05, TÅT09, TMN01, THC^{+}07, Tri01, Tsa02, ULA^{+}02, Uhl03, UOM01, UOTM03, VKN07, VAH04, VCF^{+}04, WCGL00, WK02, WCG04, WBDB04, WLGX09, Xia01, YWLC04, Yos09, ZKASS05, dS03, Esq04, UVLRRC09]. Simulator [HLW05, HKL^{+}07b, CGCS07, CD01b, DMD^{+}07, GCK02, MP08, VK09a, Gha05, HKL^{+}07a]. simulators [BSW^{+}07]. simultaneous [GFS03]. sinc [WDB04]. sine [KS05]. Singer [DDdMS02]. SINGINT [Kau03]. Single [DDMM06, MAM04, MAM07, Dev05, DHS00, FH04, FK00, FS01b, FS02, GSF05, KNSY07b, KFI^{+}01, LY05, PRSB08, ISX05, SG00a, San00, SM04, SM06a, YN05b, YD06, SBM09a]. single- [DHS00, FH04, KFI^{+}01]. single-cell-based [ISX05]. Single-particle [DDMM06]. single-shell [GSF05]. single-walled [YN05b]. singlet [JC01, KJ07, Voi02, Voi03]. singlet-singlet [JC01]. singlet-triplet [KJ07]. singular [Del08, Kau03, KM08b, LC00, PNH00, Ram04, Riz02, YZW02]. singularities [BW08, GSGT03]. singularity [IM01]. sintered [KEL02]. sinusoids [CN00]. SiSe [CM02a]. site [NLC09]. sites [IW01, IW02]. Sitter [DKV00]. Six [FFK02, Bac00, BBB^{+}09a, BGH^{+}09a]. six-dimensional [Bac00]. Six-state [FFK02]. Sixth [CFMR08]. Sixth-order [CFMR08]. Size [NFS02, BJ08, BS08, Car07, DGAG06, GDAG05a, GDAG05b, HBW05, MDS09, RP02, SS02b, VBFD01, WV04]. sized [RRCV09]. sizes [MM01, MK02]. skimming [SS09b]. skin [AAA^{+}00, BBBR04]. sky [RTVZ08]. Skyrme [BD05, BFH05, DD00, DO04, DO05, DSC^{+}09, SDNR05]. slab [AH02, KV07]. slabs [JTS^{+}06]. Slavnov [PTL04]. Slavnov-Taylor1.0 [PTL04]. SLC [JWW00a]. slender [IL07]. sliding [HOT07]. slip [MS05b, SGK09]. slit [BDHP08]. slow [AAM^{+}01, Yos03, Yos07]. slowing [CM03]. slowing-down [CM03]. SM [JKW06, JKW00]. Small [CGG00, TIM07, Alf09, GF02c, MVS05, ZDKG05, TIM08]. Small-angle [CGG00]. smeared [HK02, KT04]. smearing [Dür05, Dür09]. SMMP [EHHH01, EHHH06, MMEH08]. SMMP-open-source [EHHH06]. Smoluchowski [Kos05]. smooth [FMD07, OD07]. Smoothed [BBBD06, JJHvO03, KNY05, TE05, VKPB09, BTS06]. smoothing [Dem03, Dem06]. snow [MYJY01]. social [KOS^{+}09]. Sociophysics [Sta02]. sodium [BCP04, Kur02]. soft [HSS^{+}08, KPS^{+}01, LAMH06, PJSK08, SSLN02]. soft-core [HSS^{+}08]. SOFTSUSY [All02]. Software [BG01, Org01, SMZ05, AAG^{+}04, AEB02, BBB^{+}01, BSO^{+}04, BN07, BBJ^{+}08, BFB^{+}09, CNMC09, Che07, EHHH06, Esq04, Gha05, KVR^{+}00, MP03, OPB^{+}09, PFPB^{+}09, RC04, RMMP02, SC04, Teh01, TV07, THC^{+}07, TYS^{+}00, VPP^{+}12]. softwares [LL07]. solar [KL01, RTVZ08, SLC09]. solenoidal [YSM09]. solid [BDM09, CGC^{+}09, CC08, FFF01, HKK^{+}01, HFN03, ICO01, JKKT00, KM01a, KK05, LÅT04, MMB02, Müs02b, dlRBPL09, RCG05, Ste05, WMNS09, YT01a, Yok09]. solid-state [dlRBPL09]. solidification [NW02b]. solids [ADS06, BFL04, MPK00, RR05, SBM02, THM01, YG09]. Solitary [KD09, Zak00a, Zak00b, Zak01, bLpL02, pLbL03, Str01a, YB02a, Yan03a]. Soliton [HNS01, GI09, YB02a]. solitonic [GT01]. solitons [BZ00]. solute [MLF07, NJ01]. solutes [LBM05]. Solution [BFH05, DD00, DO04, DO05, DSC^{+}09, LRI^{+}06, LIR^{+}06, Lüs04, NT04, SŽ00a, SR05, Var08, Yao09, AP04, AMP^{+}00, ASVA00, AKS01, AKS02, BD05, BTS05, BV00, CC04, CFKM01, CBF^{+}04, CRS09, DGSL09, EMJH03b, Fij99, Fij00, FS01a, FS02, GBC^{+}04, Huj05, Ixa02, JBBR01, Kas00, Kos05, LdVJ06, LVV06, Li03, LC00, MM04, MP01b, PAS09, PSK01b, RIB01, Riz02, Sho04, Sho07, Sim00, SVA03, Sim09, SDNR05, Sug01, TKP06, Van05a, WGDZ04, WC05, WDHE04, WW06, ZSK^{+}04, ZDKG05, Zie05, dA08]. solution-adaptive [Zie05]. solutions [AA08, AKZ00, AK07, BGH04, CC09, DKV00, EÅU05, EELZS04, FRdS09, GT01, GI09, HNS01, HL00c, HJZL07, bLpL02, LL04, LJ08, LJ09b, pLbL03, LL08, PAS09, Rib02, SW09, TD03, UK02a, UK02b, UYK^{+}04, VBC07, Yan02, YB02a, Yan03a, Yan03b, Yan03d, Zak06]. solvable [HNS01]. solvated [BSB02]. solvation [LBM05]. solve [CTG01, DKMF03, GNZ^{+}09, LVV07, OGWH03, PS08, TS06, Tol02]. solvent [BDH^{+}05, CCD07, GSM^{+}03, LH02, itVPG08, XD08]. solver [ADG08, AEB02, BK06a, EST00, FS00, HCH^{+}06, Ida00, Ida03a, Ida03b, KA04, MB04, MOS00, MOS01, PCV06, QR01, QG04, TP01, WPL02, WRC^{+}04, XON08, Zie04]. solvers [Bra05, FS03]. Solving [BB07, FS00, IH09, IHAR09, JS06, KEM^{+}01, LOCJ05, Maa06, SHV^{+}01, XSC09, Zim05, ACK05, AK03, BSO^{+}04, CJK09, Den08, FBL00, Fra07b, GSGT03, HHWH07, JK08, KA09, LY05, LJ09a, MZB^{+}04, MK08, MP01b, NJ01, PMG07, PAD07, PKST03, PSV00, Ras09, RE09, Ras17, She03, SZ00c, SM02, Str00, SFSL09, TPYV03, UNK12, WS09a, WYX09, XZ12, YZW02]. Some [BKM02, FGR06, JBS08, LZ00, Luo00, MA08, Bor02, BCV03, Hib01, MSD08, Roy09, Van05a, WSB04, Wen01]. sophisticated [Gre07, MM09]. sorting [REAB08, REAB09, YWLC04]. soundings [AdlT03]. source [ABNÅ05, CLL^{+}07, EHHH06, JP09, LCS07, MTLC01, MSB09, SJHY07]. sources [DDEM00, DW01]. Space [AC05a, BD05, Bre05, CC04, DC05b, DKV00, FPB08, FMMQ08, FER^{+}07b, GBFS07, GBM02, GW01b, GHPS04, Han00, ISSB01, IM01, Jia08, KM00a, KM01b, MSHP02, MSHP20, Nak08, OSK^{+}02, Pap01, PRBD09, RLH^{+}09, SKNV01, SKNV05, SW00b, SMH^{+}01, SRR^{+}00, SBB03, TKSR00, Trö08, XON08, YT01a]. spacecraft [KTG04b]. spaces [PL05, SH06]. spacetime [Rib02]. spacetimes [BFI^{+}00, Vul03]. SPAI [SBD^{+}06]. Spanish [MBC^{+}09]. Sparse [RLRR06, BN07, BMG01, Cha00, DM07, EFS^{+}08, FM03, GHP01, MYC09]. sparse-blocked [Cha00]. sparticle [EH06, EH07]. spatial [EELZS04, GKM^{+}00, HTM^{+}08, KMR^{+}09, SBD^{+}06, ZBB^{+}06]. spatial-grid [WMNS09]. Spatio [RDS02b]. Spatio-temporal [RDS02b]. spatiotemporal [GLW03]. Special [iSHS^{+}08, SMS^{+}00, SHI02, CPT^{+}01, IOM00, Mak01, SHH^{+}04, SIE04, Tho01, Van05a, Wen01]. Special-purpose [iSHS^{+}08, SMS^{+}00, SHI02, Mak01, SHH^{+}04, SIE04]. specialized [SS02a]. species [DHS00]. specific [CGC^{+}09]. spectra [All02, BB04b, BKM02, GCP^{+}02, HSSA01, Jia08, JC01, KJ07, MK05, MKJ^{+}05, MM05, MM09, Por03, RC04, RF04, TKB^{+}04, TK09, WCBN05, vHLP08]. spectral [BP08a, CCBL02, CJC09, EVL00, FS01a, HDG07, Hua09, LBPS09, PMG07, PKST03, She03, SZ00c, TPYV03]. spectrometry [ISS^{+}02]. Spectroscopic [GZF04, CDD08]. Spectroscopy [Vég04, EST00, KSTL03, MB05b, WCBN05, ZPB09]. spectrum [DKM07, GIME02, RDSS01a, Sol01, VT00a, Wan06a, Wan06c, VT00a]. Speed [GGL03, TIN^{+}09, GCD06, iSAK^{+}08, SLL01, TCF00]. Speed-up [TIN^{+}09]. SPH [JOS07, MDH04, MC09, MK09, SM06b]. SPheno [Por03]. sphere [PP09, SA09, SWFL00]. spheres [BDH^{+}05, JBS08]. spherical [Bal07, BK05a, BD05, CRW09, CMT01, MP05, OSK04, OK09, RSMK^{+}00, RJFB08, San00, Tal09, IFF01]. spherically [AG05, IW02]. spherically-symmetrical [IW02]. spheroidal [CFKM01, Hua09, Kir06, LKC06]. Spin [CY01, NH09, You02, BCC^{+}08, BDLT02, BR09, CSW02, CPT^{+}01, DKC08, FdO09, Flo01, GF02b, Goc04, HG02a, JW02, Kat02, KK01, LCV06, LDBG08, NSMO02, OTY02, PS08, PMV02, RD05, RLU00, SH06, SS06, TEP00, YD07, You05, ZPB09, GSF05]. spin- [DKC08, PS08]. spin-angular [GF02b, GSF05]. Spin-box [NH09]. spin-orbit [TEP00]. spinor [MM08]. spinor-helicity [MM08]. spins [DDD^{+}01]. spline [FZ09, NM03, TD03, Zat06]. splines [AC09, Nik03]. split [CA07, MK08]. split-operator [CA07]. splitted [Zak01]. splitting [GLP03, SG06]. splittings [AJ08, JG02]. Spontaneous [SMV01, ICO03, SJHY07]. spreading [KPS^{+}01, MMB02]. spring [EM08]. spring-block [EM08]. Springer [Hoo04, Koc02, Laf03, Par04, Sha04, Vio04]. Springer-Verlag [Hoo04, Par04, Sha04, Vio04]. sputtering [IH01, SZ00b, Sev00, WSB04]. squared [KT04]. squares [Dem06, JC07, TD03, WWF08]. SSNT [PBI07]. SSOR [GH00]. SsTools [KW07]. Stability [Van05c, ATIO06, ATF^{+}09, FGF03, SHW01, SIH^{+}01, She08, Sim09, TMN01, UVLRRC09]. Stabilization [VT00c, bHhL07, Nur04, TCF00, WZHZ06]. stabilized [BLS09b, MVJ09]. Stable [MNH01, PC08, RB00, SW00a, WW05, Wan05b]. stably [LCM00]. stack [Sch08]. stacks [LMS05]. stage [KKSR04]. stages [LAF01]. Staggered [KNT08, Cha04]. stair [Ver04]. stance [ZSdD^{+}08]. stand [DGR09]. stand-alone [DGR09]. Standalone [TP01]. Standard [FK00, FIJ^{+}03, HS02, Ano07-31, JS08, LPC^{+}04, GFF01]. standing [BB07]. Star [BCP04, EKW09, HBRS05, QTMH07, FFPW01]. star-image [QTMH07]. stars [BLCR05, CDQF07]. starting [FFF01]. State [BRB09, RPY07, AMP^{+}00, Bac02, Bat03, BM04, BKB02b, BDM09, BCH05, CBBJ02, CWSH08, DCNDC09, DC07, FFK02, FFF01, FV02, HSGBK08, HG02a, KSS02, LÅT04, LEG02, MC03, MHGV09, MHS05, dlRBPL09, PRBD09, SH06, SVP09, SJF07, Wan05a]. state-history [MHS05]. state-to-state [HSGBK08]. states [BBB^{+}09a, BM06, BH03, BJ05b, CRS05, CWW06a, CWW06b, DGV08, GFG^{+}06, GLMADB^{+}02, GPT08, HC08, JWW00b, KB02, KN07b, LJY07, LLY07, LR07, LVLS02, LHMB00, MT00, OvSA02, ON08, PJK00, RF08, RDFF02, Sav01, TNCG00, VT00c, YN05b, Zha01, ZSSA00, dO09]. states-computational [KB02]. STATFLUX [GMAN^{+}07]. Static [BKB02b, QCL05, Ver00]. statics [LMM^{+}08]. Stationary [TLCS04, Bae03, DGV08, WDHE04]. Statistical [TSB^{+}05, ASJ^{+}03, Ano09t, DSS01, GMAN^{+}07, GGL^{+}02, ISSB01, JGJ09, Nov02, PJSK08, Rin02, SAG^{+}02, Sta00, Suz00, Swe02, TY01]. Statistics [HNG05, FHF00, ISSB01, YT01b]. Status [UXD^{+}09, Ano01a, McK07]. steady [CTG01, ZSSA00]. stellarators [SIH^{+}01]. stem [Dom05]. step [BCP04, Ber03a, DWZS05, FLO06, HDGM07, Ida03a, Ida03b, IVD03, MA06, Sho07, Ver04, WGDZ04, WW05, WC05, Wan06a, WTW04, ZZH09]. steps [KV08, NR01, SSH02, Sho04]. stepsize [WDHE04]. stereo [HFN03, SGM^{+}09]. sticker [Zim02]. sticking [Tsa02]. stiff [BT01, MVJ09, PC08]. Stiffly [BT01]. still [BL00]. Stochastic [AAA^{+}00, BP08b, DD01, LH01, MRS04, ZS03, ZS07, ZSdD^{+}08, ZS08, BDP00, BM02b, BT01, DVG05, FRdS09, FM00, Lad09, TiTD01, TRAdO09, Var08, ZE00, dO09]. stock [LLH07]. Stokes [ICT01]. Storage [BNO^{+}01, Ano09t]. strain [BSO^{+}04, HTA08]. strained [Kim07]. strains [LTT09]. stranded [VYK02]. strange [Lew04]. Strategies [Tri05, Gol00, SLWH02, Ska05]. strategy [Bae03, Gra02, PL05, TLP04, VT00b, WLH00]. stratified [LCM00, TdFK00]. streams [Bru04]. strength [JK02]. strengths [EKW09, HB05]. stress [EM08, HM06a, MCC05]. stress-induced [EM08]. striation [BCD^{+}07]. strictly [SA09]. string [ABD^{+}05, GHIL09, LLY07]. strings [BR01, RK05]. STRINGVACUA [GHIL09]. strong [ACK05, CBBJ02, CKS00, KLD04, Kur02]. strongly [Alv09]. STROTAB [KJ07]. Structural [Blü09, EM08, KACB07, TBR07, Iwa01, ZBB^{+}06]. Structure [GF02c, HOI04, Mor01, AAA^{+}00, AJT^{+}07, ASH06, AJ08, BD08, BTI01, BB00, BMG01, Bro07, CPV^{+}08, DHS00, Fro00, FTGG07, FS01b, GSF06, GHP01, GOH06, GBD03, HC00, HTM01, HJM02, JHFG07, KFJ^{+}09, KLM00, KPF03, KNSY07b, LTG09, LOY07, LC08b, LB04, LZ04, MS06, MSB09, MWA01, NSKS01, New02, OKS04, Oka01, PFG06b, PKSF01, QASF^{+}05, RB08, SHV^{+}01, SG04a, SH06, SBM02, SHX02, SKNV01, SN07, SMH^{+}01, SKH02b, SYM00, THM01, Vos06, WKP^{+}01, Yos09, ZF09]. structure-preserving [LB04]. structured [DDEM00]. structures [BB07, Cle05, Flo01, GCP^{+}02, HKK02a, HKK02b, LV08, LLT^{+}02, LY05, LTT09, LF02b, LLLZ01, NM01b, OGG07, OBG09, PMV02, RRCV09, SLC09, SOAW08, Str01a, TMN01, VS01, WP00]. Student [Ano04a]. studied [Bur02, HTM^{+}08, MSH01, RvOvV02]. Studies [BS04a, BJ08, BJ03, CCG08, CSC^{+}07, CSC^{+}08, Dom05, FMD07, HKK02b, LMS^{+}02, MVS05, Min01, Rap08, iTKST01, WM00]. Study [LDZ^{+}08, PSK01a, PSK01b, RLU00, iSAK^{+}08, SGF04, SSB04, TAM04, AGJJ07, ABOSPG09, ADE^{+}02, BJ02, BZ00, Bor02, BBJ^{+}08, BFB^{+}09, BCV03, CRPC08, CH09, DELG05, DMR01, DCJ07, DC05a, DGR09, DH00, FGV01, FS01b, GAR05, GW01a, GDAG05a, GDAG05b, HOT07, HGVCM^{+}02, HM06a, HTA08, ISH01, KEL02, KL07a, KNSY07a, Kur02, LWT08, LN01, LNK01, MCL05, MCC05, PKRK07, PAT^{+}09, RIB01, RG05, RCG05, SS07a, SMSE03, SK08, SWL09, SVP09, SGM^{+}09, SHJ07, SS09b, SSLN02, TYSH05, WSCW09, YD06, YC07, YRR07, YKK07]. studying [GHIL09]. Sturm [CGVA09b, LVV04, LVV09]. sub [GS01a, QTMH07, SLC09]. sub-membrane [GS01a]. sub-pixel [QTMH07]. sub-wavelength [SLC09]. subgrid [Ker02]. sublimation [WSB04]. submonolayer [AFP02]. subroutine [Tal09]. Subroutines [WSB04]. subspace [SMZ05, ZSM05]. Successive [AS03, BB04a]. such [SSPM05]. sudden [PCC01]. SUE [CPT^{+}01]. suggestive [Niu00]. suitable [SI01]. suite [JU09, SBM^{+}04, SSB^{+}09, TKB^{+}04]. sulfur [MSH01]. sulfuric [CTI07]. sum [QCL05, Sch06b, SF06]. summation [AH02, Har02, LHC01, LHC02, MU06, TZZ06, Wen01]. sums [Bek06, Blü04, Blü09]. sunrise [CCGR09, PR06]. super [Bar04, KW07]. super-heavy [Bar04]. super-systems [KW07]. superbursts [NBPG08]. Supercomputer [Yos09, CD09a, CFH^{+}01, FMD07, FDM07]. supercomputers [BAD01, CD08, CBM^{+}05]. supercomputing [MSK^{+}02]. superconducting [KW03]. superconductivity [GOG00]. superconductors [VS01]. Superconvergence [LCHJ09]. superfield [Fer07a]. superfluid [Yos03, Yos07]. superfluorescent [MTLC01]. superheated [KNSY07a, Ste05]. SuperIso [Mah08b, Mah09b, Mah09a]. superlattice [GVMW04]. superlattices [JK01]. supernova [SBD^{+}06]. supernovae [HRN00]. superresolution [KSTL03]. Supersymmetric [DKM07, FIJ^{+}03, HS02, All02, LPC^{+}04, MDM05, Por03]. supersymmetry [Mah09a]. Suppressed [NLC09]. Surface [KNSY07a, LS02, LAF01, BVY05, BB07, BLCR05, BDHP08, BH01, BDH^{+}05, CW02, DEW01, DVL^{+}02, DVL^{+}04, EG09, HYY07, KPS^{+}01, Kim07, MKB02, NP01b, OLX07, SHV^{+}01, TCO00, WSB04, WMNS09, XSC09, ZHC00]. surface-controlled [BDHP08]. surfaces [ATP01, ABV02, BM01, BTK^{+}02, CIC^{+}03, GGG01, GBM02, GI01, Har01, HG02b, Hin00, Ida00, Ida03a, Ida03b, Ing01, KM01a, KMB02, LNK01, LTG09, MPS09, NSYZ02, NP01a, PCC01, Ple02, Rou01, SSH02, Sri01, TAP01, TGB01, Tsa02, YG09]. surrounding [LY05]. survival [SSH01]. susceptibility [VEG08]. SuSpect [DKM07]. suspended [ICO01, KH06, RSMK^{+}00]. suspension [DHB^{+}04]. suspensions [SF05, UVLRRC09, WDF^{+}02]. Susskind [CAF^{+}03]. sustainability [FKMB09]. sustained [FKP03]. SUSY [Ano09u, FIJ^{+}03, Hah09, Por03]. SusyBSG [DGS08]. SusyMath [Fer07a]. Suzuki [OMF02]. Suzuki-like [OMF02]. swap [MHR^{+}07]. Swendsen [DGAG06]. swimmers [PY08]. swimming [Rap08]. switches [Del03]. switching [OD08]. SX [EL04]. SX-6 [EL04]. Symbolic [Ada04, BGH04, KHÖ01, KH06, She08, UTKF05, Wei02c, qXbL04, CRUV00, FD03, GT01, Hon04, MU06, Niu00, Pee07, SH05, TV07, qX08, Yan02]. Symmetric [CBF^{+}04, AG05, BN07, Bun01a, CFMR08, CR00, DM07, Kim07, SŽ00a, Wan05b]. symmetrical [IW02, WS02]. Symmetries [MG09b, BCV03, Che07, MG09a]. symmetry [BD05, MMR04, RF05b, Sle00, YN05a]. Symplectic [AK07, OMF03, SS06, CFMR08, Fra07b, MKS07, SS00, SQ03, TQ03, Van06]. SYN [NR01]. Synchronization [BFL^{+}01, MTC07]. synthesis [MP01a, ZLM04]. System [BFL^{+}01, Pub07, AP04, ABNÅ05, ABC^{+}03, Ano01n, BGLLW01, Cap05, Eli05, GC01, IH01, JP09, KS07, KK01, KM00b, KTT02, KMCS01, LdVJ06, LNV^{+}09, Mar01, Mas00, MTZ00, NFS01a, NFS02, OK06b, Pee07, Pop03, Sol01, SS09b, SQ03, TKS^{+}01, TLDM03, TWY09, Ver00, Wan09a, Wei04, ZS03, ZZH09, AAB^{+}08, BNO^{+}01]. systematic [Bra05, Con04, EL06, THC^{+}07]. systems [ATP01, AC07, AJT^{+}07, ASF^{+}05, Alv09, ALN^{+}01, ADDdM07, Bae04, BBC^{+}01a, BFMH^{+}01, BB04a, BCC^{+}08, Ber02, Bin02, BHM^{+}07, BH03, BBJ^{+}08, BFB^{+}09, BVKW02, CD05, CGG00, CYAS05, DDMM06, Di 01, Dol01, DVL^{+}02, DVL^{+}04, EMJH03a, FGF03, Fra07a, GKI02, GKI04, GPW^{+}09, GSGT03, GMO03, bHhL07, HNS01, HL00a, HL00c, HZGZ09, JW02, KBBW02, KH01, KNTG03, KW07, KSEG05, Krö05, LM02a, LbotMC01, LNV^{+}09, LFT03, LANM^{+}01, LMM^{+}08, LAMH06, LM00, MCH02, MSD08, MFVJ07, MHK02, MKS07, MM01, MTJ02, NSMO02, NH09, iOY01, OTY02, OMF02, PM02, PKPV02, PCCD09, PSH06, RP02, RF05a, RF06a, RF07, RF08, RDFF02, Sle00, SOS01, SS06, TAP01, Tat07, Var08, Var02, VT00c, VPK^{+}01, WKP^{+}01, WZHZ06, WYX09, WLGX09, Zha00, ZKASS05]. Sznajd [Sta02].

T [PKKM02]. T3E [ALN^{+}01]. table [HS01b]. TADpoles [Ste01]. tailored [CR08]. tails [HBMJ05]. tangent [BGH04]. tapered [NSKS01]. target [BDB^{+}08, MOC03, OMC00, OSK04, OKS04, PD08]. targeting [vDGM^{+}09]. targets [SBM09b]. Task [CD08, Ano09s]. tauola [GKP^{+}06, BEM^{+}02, GKP^{+}06]. tauola-photos- [GKP^{+}06]. Taylor [DMR01, DMR02]. Taylor1.0 [PTL04]. TaylUR [vH06, vH07]. TD [WPL02, WT01]. TE [KV07]. TEA [Gha05]. Teaching [HF00, TPBE04]. tearing [Lüt04]. technique [Bae04, CIC^{+}03, EMJH03a, Har02, ICT01, KMH02, KA05, KK00, LHC01, LHC02, PDA06, QTMH07, Ram10, Sal03, dSL02]. techniques [AP09, Bes02, CLFH07, DSC06, GHLW03, Hei01, PBB^{+}04, PY08, Ram04, RM05a, SWC^{+}03, Tod01, TYS^{+}00]. technologies [Chr00, CBM^{+}05]. technology [CRS01, Far01, Lüt00, SMS^{+}00]. telescopes [CBMS08, GK05]. Teller [MS08b, HC08]. temperature [HTM^{+}08, HJ02, KMD^{+}02, Kar02, KLD04, Lei02, LDZ^{+}08, MTJ02, NH09, PP02, Zha00]. temperature-driven [PP02]. temperatures [DS01, FS01a, Kat02]. Temporal [RMK05, RDS02b]. tens [HHM^{+}09]. tensile [Kim07, LTT09, MDH04]. tensile-strained [Kim07]. tension [NL07, ZHC00]. Tensor [BH01, BGH^{+}09a, Bre07, GBM02, MP04, MGPM07, MGYP08, MG08b, Por00, RY00]. tensor-trick [RY00]. tensorial [HHL06]. term [SVA01]. terms [AA08, Dzu09, HTM^{+}08, KKS04, MYC09, RMWH01, Yao09]. TERS [Nat10, Nat09]. test [BJ05b, DVG05, GCP^{+}02, KHÖ01, MTJ02, OML09, PFPB^{+}09, SKNV01, qXbL04, qX09, ZLL09]. test-driven [PFPB^{+}09]. TESTER [GPW04]. testing [WL08]. tests [ABC^{+}01, BL00, BFI^{+}00, JW02, TIM07, TIM08]. tetra [HGVCM^{+}02]. tetra-atomic [HGVCM^{+}02]. tetraatomic [KLTH04, ATP01]. tetrahedral [CN01, LHS^{+}06]. tetrahedralizations [SMH04]. Tetrahedron [Zah04, Zah05]. Tetratomic [TAP01]. text [dlRL09]. text-based [dlRL09]. th [AKZ00]. their [BSB02, CGVA09b, CN00, Gro01, KCH00, LC01b, OS00b, YGT^{+}02]. them [Ort00]. theorem [CT00]. Theoretical [CS07, LNK01, Lee04, ASJ^{+}03, BL00, Bre05]. theories [ALV05, Di 02, MG09b, Tri05]. Theory [Rou01, ARV02, ALN^{+}01, BB09a, BSK^{+}03, Cha00, Cip07, Cip08, Cip09, Dzu09, Fod05, Gut06, HHM^{+}09, HLC08, JC07, KTT09, KH09, LCV06, MBR01, NP01a, OSK^{+}02, PKSF01, Pee07, QP05, RF04, Sem09, SKNV01, SKNV05, SMK01, WWF08, Wil09, dlGGS^{+}05, vHLP08, MG09c]. THERMal [KTBF06, IK00, CT00, DHS00, HS01a, LLLZ01, RP02, Sat02, TCF00, WSB04, WCH09]. Thermally [RLU01]. THERMINATOR [KTBF06]. thermionic [LLLZ01]. Thermodynamic [RS09]. Thermodynamics [BDH^{+}02, BFL04, TNI^{+}07]. Thermonuclear [HRN00, BSW^{+}07]. thermostats [BP08b]. THERMUS [WCH09]. theta [Sch05]. thickness [CAW00]. thin [BDV04, Dan05b, LTG09, LTT09, Mül02, NT04, SLWH02, TÅT09]. things [Ort00]. thiols [SPV07]. third [MKS07]. thousands [CR05, HHM^{+}09]. threaded [LbotMC01]. threading [NJ00]. Three [CSC^{+}07, CSC^{+}08, CHM^{+}09, IH01, KKF^{+}04, PKRK07, QG04, RG04, BCBJ02, BSDMH05, EELZS04, ES09, GBC^{+}04, HKPL07, JK08, KNT08, KSS02, KL01, KK01, KM00b, Krö05, LVLS02, LZS06, LHS^{+}06, LEG02, MMTH04, MOS00, MOS01, NYH04, OvSA02, PSK01b, SMH04, SBJ05, Sch06b, TRGR08, TMTF00, Tak03, TND04, TND05, TY01, Vos06, WW05, WHJ06, WTW04]. three-atomic [KM00b]. three-body [LZS06, LEG02]. Three-dimensional [CSC^{+}07, CSC^{+}08, CHM^{+}09, IH01, KKF^{+}04, PKRK07, QG04, RG04, ES09, KNT08, KK01, Krö05, LVLS02, LHS^{+}06, MMTH04, MOS00, MOS01, NYH04, SBJ05, TRGR08, TMTF00, Tak03, TY01, Vos06, WTW04]. three-grid [Sch06b]. three-quark [OvSA02]. three-step [WW05]. Threshold [FHF00, HOT07, HTM^{+}08]. thresholds [Lüt04]. Thrombosis [OCS^{+}08]. throughput [SBM09b]. thruster [CMS04, CSC^{+}04, KB04, TLCS04]. TiC [ZLM04]. tight [CR00, ÖDÇ02, Ver00]. tight-binding [CR00, Ver00]. tilt [LVH07, ZS03].

Time [FS01a, HGVCM^{+}02, Mei01, AOT01, BB04a, BMSG01, BC05, Bow02, BK05c, BSK^{+}03, CRS05, CHS09, CMT00, CMT01, DS01, DKMF03, Den08, DKC08, DC07, DB08, FNR^{+}06, FNR^{+}07, FER^{+}07b, GNZ^{+}09, HDGM07, Hei01, HL05, HZGZ09, Ida03a, Ida03b, ISSB01, JH09a, KM08a, Lee04, LOCJ05, MN01, MA06, MA09, Nak08, NM01b, NN06, Nur04, PSV00, SŽ00a, SG06, SOYN01, Sim08, SM02, SRR^{+}00, SSA07, TSI02, UJSW06, WS09a, WLR^{+}08, WTW04, ZS07, ZS08, dlGGS^{+}05, dlHV08]. Time-dependent [HGVCM^{+}02, Mei01, BC05, BSK^{+}03, CRS05, DKMF03, DKC08, GNZ^{+}09, MA09, NM01b, PSV00, SŽ00a, dlGGS^{+}05]. time-discretized [DB08]. time-domain [CMT00, CMT01, Den08, HL05, NN06]. time-ensemble [Nak08]. time-evolution [SOYN01]. time-harmonic [CHS09, HZGZ09]. time-independent [DC07, MN01, SM02, WS09a]. time-invariant [BB04a]. time-step [HDGM07, MA06, WTW04]. timelike [MG09c]. times [MSS^{+}07]. timeseries [PB09b]. timing [HDGM07]. Tinker [PLS09]. tip [CW02, MCC05, ZS03]. tip-tilt [ZS03]. TiReX [PLS09]. tissues [YC07]. Tl [LK07]. TLM [Hei01]. TMAT [dAK01]. tmLQCD [JU09]. TMOL [BRdAHK04b]. today [Shi09]. together [Gre04]. tokamak [ATIO06, ATF^{+}09, EH03, KY07, NK07, PMA^{+}04, PPP01]. tokamaks [GIME02, SJCM04, SMSE03]. Tolman [Rib02]. tomographic [BG01, LVH07]. tomography [Bal07, CEM08, GBA01]. Tool [SPF00, Str05, Ano09s, Ber03b, BBB^{+}00, CNMC09, Esq04, Gha05, GPW04, GBR^{+}09, HL08a, LPC^{+}04, LCE^{+}09, LR06, MHGV09, MCBR03, MRF^{+}05, MYL^{+}08, SMSE03, TND04, TND05, TGD06, WH05, YFM09, vdB08]. Toolkit [Mal00, SR09, BFL^{+}01, CKK09, HF00, MOM^{+}00]. tools [Di 01, MC01, Org01, SKF05, WSCW09]. top [KJ07, Kol03]. topography [ES09]. Topological [CKLS09, dlRBPL09]. topologies [RDSS01b]. topology [IW01]. TopReX [SS02a]. tops [JC01]. TORBEAM [PPP01]. torch [CHL^{+}07]. toric [KS04b]. Toroidal [Lüt04, BDBV12, GS01b, IIK^{+}08, KZS^{+}00, Liu07b, SG00b]. toroidally [ATF^{+}09]. torsion [GW01b, Vul03]. torsional [Bac02]. torus [FMD07, FDM07]. Total [MSHP02, MSHP20, BS03, NRR01]. Total-energy [MSHP02, MSHP20]. trace [KBV09]. traceability [BCC^{+}06]. tracer [Str01b]. tracing [BMSG01, PPP01, Pop03]. Track [Bla00, GGQ01, GKM^{+}00, GKK^{+}08, JGJ09, NY06, NY08, PBI07]. TRACK TEST [NY06]. TRACK VISION [NY08]. Tracking [Laf03, iSAK^{+}08, HHCC05, TCO00]. tracks [NY08, SWFL00, YFM09]. trade [Oli01]. trade-offs [Oli01]. traffic [CL02, Wal03, ZWY04]. trajectories [AGJJ07, Elb05, Nat08, ZE00]. trajectory [LJY07, MSS^{+}07]. trajectory-length [MSS^{+}07]. transactions [BD06, Dan09b]. Transcendental [MU06, Wei02c]. Transfer [Yak01, Bes02, BCH05, DC05a, GDC01, Ger07, Liu07a, RE09, SGK09, Str01b, WH00, Wol03, WH05]. Transfer-matrix [Yak01]. transferred [CHL^{+}07]. Transform [DSC06, RM05a, Don02, KA09, KSTL03, LC08a, Ras09, Ras17, SSP08b]. transformation [ASJ^{+}03, FLO06, LL08, Niu00, YN05a]. transformations [GF02a, KEL02, NP00, PZ01]. transformed [Eli05, SDNR05]. transforms [Blü00, Blü09, CR08, Dup01, HC00, KSHP02, MA00, MM05, RB05, SSP08a, Tal09, Tör00]. transistors [CSC^{+}07, CSC^{+}08, LH03, Mam08]. Transition [BR09, AGM^{+}00, BJ05b, CBBJ02, DKC08, FFD00, GAR05, HBW05, JK02, KITK00, KMP09, KT07, LJY07, LLY07, LDZ^{+}08, LA09, MSS^{+}09, Maz00, Mor01, OS03, PDA06, Tat07, Wil02, YH02, YD06]. transitions [BJ05a, BDH^{+}02, Bin02, BHM^{+}07, BKB02b, CSCK08, CM02b, FHR^{+}05, JS05, KGM00, KK01, KNSY07b, MCH02, OIKN02, PRSB08, PP02, Ple02, SWL09, SSLN02, YGT^{+}02]. Translational [TK08]. Translocation [LC07, MB05a]. Transmission [Man04, KV07, Nat08, Sch08, SYM00, WP00]. transmission/reflection [WP00]. Transport [Ano04-46, KY07, KMR^{+}09, PMA^{+}04, ABSM04, BDYK04, CMD00, CGK^{+}00, EST00, GZ07, ISSB01, KKKC07, Lee04, LLLZ01, Man04, MLF07, MBC^{+}09, NKSL05, NRR01, PAD^{+}09, PC08, Pop03, QTL06, Ros04, SYN01, SGK09, TAM04, TKP06, Vie01, WTH^{+}04, WML^{+}05, WRN01, Yak01, YNZ^{+}09, Zie05, dA08]. trap [CKV04, MA09]. trapped [RLI07]. trapping [NRDHB01, PCA^{+}07]. traps [TS06]. traveling [FD03, Hon04, LL04]. travelling [EELZS04, bLpL02, LJ08, LJ09b]. treat [GLMADB^{+}02]. treating [CA07]. Treatment [IM01, Bac02, GWK09, KL01, MMR04, PNH00, Ram05, TJLR06, WC00]. tree [ADBF03, BAD01, BCAD06, FIJ^{+}03, JKCGJ08]. treecode [AL08a, CKV04]. trends [Sch04]. tri [HGVCM^{+}02]. tri- [HGVCM^{+}02]. TRIAC [PBI07]. Trial [PDM^{+}08, PAT^{+}09]. triangular [BM06, CHS09, HZGZ09, MCLDP01, She08]. triangulation [BSDMH05]. triatomic [HC08, TT06, TKB^{+}04]. triaxial [MAM04, MAM07]. trick [RY00]. trigger [ABF^{+}01, ISSC01]. trigonometric [Sim09]. trigonometric-fitting [Sim09]. Trigonometrically [FSW08, Wan06c, Wan06b, MKS07, Sim08, WC05]. Trigonometrically-fitted [Wan06c, Wan06b, Sim08, WC05]. trimer [Bac02, RY00]. trimers [GLMADB^{+}02, OBG09]. triode [LC08b]. ‘tripleint cc’ [PAT^{+}09]. triplet [KJ07]. tritium [RCGC00]. trivial [MSD08]. tropical [Mas00]. Trotter [Inu07]. Trp [CDFF05]. truncated [ASF^{+}05, Maa06, WHJ06]. truncation [Zah00, Zah01]. Tsallis [FHF00, Sch06a]. TSIL [MR06]. tube [PCC^{+}09]. tubes [IL07]. tunable [SJHY07, Vor02]. tuning [BB03]. tunneling [MMMM00]. tunnelling [CKLS09]. turbid [CGK^{+}00]. Turbulence [Ker02, MCL05, Eli08, GLW03, HDG07, Jen00, KLD04, UXD^{+}09, YNS^{+}09]. turbulent [DVG05, Ker02, RJFB08, Str01b, TIM07, TIM08]. tweezercalib [HTNFBS06a, HTNFBS06b]. tweezers [HTNFBS06a, HTNFBS06b, TNBSF04, Whi00]. twelfth [WGDZ04]. twelfth-order [WGDZ04]. twist [CC04]. twisted [BDF^{+}08, Jan05, JU09]. twisting [DGLB08]. Two [CLL^{+}07, CNDC09, FHR^{+}05, HW09, ID09, LCS07, Moh07, STK^{+}00, TDD04, UTO09, Var08, AC07, AMP^{+}00, BCP04, BvG02, BR09, Blü00, CCGR09, CCBL02, CMK^{+}03, DS04, Dev05, DJ08, DHS00, EÅU05, EELZS04, FSW08, FK00, GR02, GBC^{+}04, GOG00, GME06, HBMJ05, HJZL07, IHAR09, IVD03, JWW00b, JK08, JKKT00, JKCGJ08, KT05, Kim03, KM01c, Krö05, LS05, Lüs05, Mah08a, MR06, Nik03, NYH04, OvSA02, PD08, PR06, RtVR09, RLV^{+}08, SKH02a, iSHS^{+}08, SSB^{+}09, SGF03, Sol01, SM02, TMTF00, TZZ06, TBZ12, TY01, TL09, TdFK00, Ver00, WGDZ04, Wan06b, WS09a, WD04, XZ12, ZY09, ZZH09]. two- [Krö05, TMTF00, TY01]. two-body [AMP^{+}00]. two-center [GME06]. two-component [JKCGJ08, TdFK00]. Two-dimensional [CLL^{+}07, CNDC09, FHR^{+}05, HW09, ID09, LCS07, STK^{+}00, TDD04, UTO09, Var08, AC07, GR02, GOG00, HBMJ05, HJZL07, JKKT00, KT05, KM01c, PD08, RtVR09, RLV^{+}08, iSHS^{+}08, SSB^{+}09, TZZ06, TL09, Ver00, WS09a, XZ12]. two-electron [EÅU05, Nik03, SKH02a, WD04]. two-fermion [JWW00b, Sol01]. two-flavor [Lüs05]. two-fluid [LS05, NYH04]. two-loop [Blü00, CCGR09, FK00, MR06]. two-particle [Dev05, DJ08]. two-phase [TMTF00]. two-photon [BvG02, Nik03]. two-point [DS04]. two-quark [OvSA02]. two-species [DHS00]. two-step [BCP04, IVD03, WGDZ04, ZZH09]. type [CFMR08, CHL^{+}07, CHP04, FSW08, FGMT02, Fra07b, PAT^{+}09, RDSS01a, SPS09, Wu10, YWYF09]. type-II [CHP04]. types [BMML05, BSvdDW02]. typical [De 02].

UCLA [DN04]. UCN [Yos07]. UHI [BF04]. UK [Wan00]. UKQCD [All01]. ultra [HGH^{+}05, Tol02]. ultra-high [HGH^{+}05]. ultracentrifugation [BS08]. ultracold [JKCGJ08]. ultradiscrete [GI09]. Ultrafast [BSS09, BCH05]. Ultrahigh [WBDB04]. ultrarelativistic [Tom09]. ultrashort [BDV04]. ultrasoft [HP06, LMM^{+}08]. ultrathin [KKKC07]. Umbrella [BL05, SR01b]. uncertainties [Con04]. Unconstrained [LÅT04]. undamped [DWZS05, Wan06a, Wan06c, WW06]. under-relaxation [TCF00]. underdense [Mah08a]. undergraduate [Chr00, Gou00]. underground [Kud09]. Understanding [Bal01, BCHP09, DSL09, Müs02b]. uniaxial [CAW00, LTT09]. unification [VPCK04]. Unified [DKMF03, Ram12, Ida00, Ida03a, Ida03b]. uniform [BCD^{+}07, JTS^{+}06, KV08, KH06, NJ01, WHO02, vHK00]. unimolecular [FS01a]. unit [YT01a]. unitary [MA06]. Units [LSVMW08, AGS07, CDD08]. Universal [BBJS09, LNLK01, Dzu09, GPW04]. universality [BBJS09, SAG^{+}02]. universe [BAD01, Fel08, FT08, Yos09]. university [HF00]. unknown [Fat02, LVH07, SSZ01]. unpolarized [Vog05]. unsteady [DM09]. Unstructured [SG04b, CSC^{+}04, LHS^{+}06, MCLDP01, MOS00, MOS01, ISX05]. up-date [Fri09]. updated [EH06]. upgrade [Dan09a, SBCZ08]. upgraded [CWW06a, CWW06b]. UPIC [Dec07]. upper [SKRK04]. upscaling [Bra05]. upwind [GZ07, ID09]. Usage [Fra07a]. Use [CMRS02, MN01, MTJ02, RF04, Swi04, BNFM^{+}09, CFJ09, FdO09, MVS05, MA04, MA08, Nap09, Sal02, SS02a, WH06]. used [BDK^{+}06, PBB^{+}04]. User [MCLDP01, Bar04, BT04, BCKT09, CGC^{+}09, Hor09]. user-interactive [CGC^{+}09]. uses [BOPC05]. USFKAD [KS07]. Using [CFJ09, GCD06, KBV09, PMG07, AGJJ07, AP04, Alf09, ASH06, AL08b, ADG08, Bar03, BS03, BDW06, Bek06, BS04a, BR09, BFL04, BH01, CN01, CLFH07, CFKM01, CA09, CMT00, CMT01, Cip07, Cip08, CC00, EMJH03a, FKMB09, FS00, GVMW04, GMAHV^{+}09, GFP00, GF02c, GSS06, Goe02, GOG00, GFS03, Haf07, Har00, HKP02, IH01, ISS^{+}02, IKO00, KPD06, KW07, KD09, LdVJ06, LLT^{+}02, LWLL07, Lik01, LSVMW08, LMS^{+}02, LNC^{+}03, Lüs04, Mah08a, MSY07, MTLC01, ML03, MBR01, Mas05, MFVJ07, MA06, MC09, MM05, ME00, MT00, Nik03, Nil07b, ON08, PL05, PLS09, PAT^{+}09, RB08, RGD^{+}01, ISX05, SNS01, iSAK^{+}08, Sch06b, SBM02, SLL07, SR05, SGF04, SOAW08, SSP08a, SSP08b, SPS09, SGL09, SM02, SDNR05, SSB04, TAM04, TIN^{+}09, TiTD01, TKSR00, TWY09, TdFK00]. using [ULA^{+}02, UVLRRC09, VKM^{+}05, itVPG08, VK09a, VK09b, Vor02, VBC07, Vul03, WPL02, WP00, WTW04, qXbL04, qX08, YWLC04, ZE00, ZF09, dDSFY04]. USPEX [GOH06]. Utilities [Fri01, NFH06]. Utilizing [BLM01, MHS05].

v [Kol03, MMEH08, DO05, FIBT01, Har00]. v.1.00 [AAB^{+}07, AAB^{+}06]. v.6.21 [BBC^{+}01b]. v1.0 [CJT06]. v1.00 [BD05]. v1.1 [BRdAHK04a, BRdAHK04b, dAK01]. V1.1.0 [vdB08]. v1.66p [SDNR05]. v1.75r [DD00]. v2.0 [Nat10, TGD07, TL08b, Nat09]. v2.08i [DO04]. v2.08k [DO05]. v2.1 [RdAGV^{+}00]. v2.3 [Mah09a]. v2.40h [DSC^{+}09]. V3.0 [Tót08, Mah09b]. vacancies [CC08]. vacua [vdB08]. vacuum [ATIO06, FS02, GHIL09, KTG04a]. valence [CYAS05]. Validation [MC01, BB03, CHL^{+}07, PS09]. Value [IHAR09, ASVA00, CFKM01, LJ09a, LC00, PAS09, Ram04, SVA01, WW05, Wan05b, Wan06b]. valued [FH04, HM06b]. vapor [JBA05, MSK^{+}05]. vapor-liquid [MSK^{+}05]. variable [CLFH07, IVD03, LL08, MBG03, SSZ01, Van05c, WTW04, qX09, Yan03d, ZLL09]. variable-coefficient [LL08, qX09, ZLL09]. variable-phase [MBG03]. variables [Str05]. Variant [RK05]. variate [BBBD06, Bel05]. Variation [IN02, NRR01]. Variational [OBG09, FMG00, GLMADB^{+}02, MM01, PAT^{+}09, SM02, Var08, WLH00, Yok09]. variety [TLP04]. various [Nii00]. varying [CAW00, Koz02]. VASIMR [IDS^{+}04]. VASP [Haf07]. VASP-a [Haf07]. Vbfnlo [ABB^{+}09]. Vector [Bal07, San00, Whi00, EFS^{+}08, EL04, Kat02, Mas05, Rap06, ULA^{+}02, WLH00]. vector-parallel [Rap06]. vectorised [KSYE00]. vehicular [Wal03]. Velocimetry [iSAK^{+}08]. Velocity [HTM^{+}08, BS08, GRS06, HOT07, Luo00, NHS07, SM06b, TKSR00]. Velocity-dependent [HTM^{+}08]. velocity-field [GRS06]. verification [UXD^{+}09]. verifying [GI09]. Verlag [Hoo04, Par04, Sha04, Vio04]. Version [Abe01, BBPS06, HHH^{+}09, Wol03, AAC^{+}06, BLS01, BFB^{+}09, BNFM^{+}09, CWW06a, CWW06b, CGVA08, CGVA09a, Cip07, Cip08, Cip09, CGK^{+}00, DS06, DD00, DO04, DSC^{+}09, EHHH06, FA00, GS01b, HTNFBS06a, HTNFBS06b, JWW00a, JPS^{+}01a, JPS^{+}01b, JS07, JC01, KRW03, KJ07, Mil07, MAM07, Nat09, Nat10, PDL04, Pit05, Pöt00, RC04, SYM00, TGD07, TL08b, vH07, BCKT09, Sem09]. versions [BD06, XD08]. versus [AA07, BDHP08, Jan05]. vertex [HBW05, SF06, ZBB^{+}06]. vertex-cover [HBW05]. vertical [LC01b]. vertically [EMJH03b]. very [OBG09, RTVZ08, WV04]. VHF [LCS07]. VHF-ICP [LCS07]. VI [ABV02, DSC^{+}09, GF02a]. via [AF05, BDH^{+}05, DGLB08, HL00c, LZS06, MCL05, Mor01, SFSL09, TCF00, TYS^{+}00, ZA01, Zim05]. vibration [KLTH04, TKB^{+}04]. vibration-rotation [KLTH04]. Vibrational [Kar02, Bac02, Bes02, MT00, RF04, SM02, VCCS05, XSC09]. VicAddress [MSY07]. video [EFH^{+}07]. viewing [Nat08]. VII [FIT03]. VIII [GSF05]. violating [GLL^{+}02]. violation [LPC^{+}04, LCE^{+}09, DGS08]. Virtual [TKS^{+}01, HF00, LKPH08]. Virtualizing [ZC09]. viscoelastic [WGS00]. viscous [MY00a, MDH04]. vision [SGM^{+}09]. Visual [GBR^{+}09, PZW^{+}00]. Visualisation [AAB^{+}08]. Visualization [AP09, OK09, AFK^{+}07, Ano09t, Bal01, iSAK^{+}08, SKNV04, SEC04b]. Visualize [Sea01, TKS^{+}01]. visualizing [MPK00]. VLab [dSdSW08]. Vlasov [AGJJ07, CRS09, DJ04, Eli05, Fij99, Fij00, FS03, GHPS04, IIK^{+}08, Jen00, LdVJ06, MV04, NGE^{+}04, PSK01a, PSK01b, PSK05, SG06, She03, SGF03, SGF04, SFF^{+}04, SSB04, UTO09, UNK12]. Vleck [Goc04]. VLIW [Far01, PKB^{+}01]. VLPL [LKPH08]. Voigt [SK08]. VOLSCAT [SBM09b]. voltage [KACB07, PPC07]. voltage-gated [KACB07]. Volterra [Sle00]. Volume [Ano05d, Ano05f, Ano05h, Ano05-35, Ano05-36, Ano05-38, Ano05-49, Ano05-50, Ano05-52, Ano01g, Ano01h, Ano01i, Ano01j, Ano01k, Ano01l, Ano01m, Ano01r, Ano01s, Ano01t, Ano01u, Ano01v, Ano01w, Ano01x, Ano01y, Ano01-31, Ano01-32, Ano01-33, Ano01-34, Ano01-35, Ano01-36, Ano01-37, Ano01-38, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano02i, Ano02s, Ano02t, Ano02u, Ano02v, Ano02w, Ano02x, Ano02y, Ano02z, Ano02-27, Ano02-28, Ano02-29, Ano02-30, Ano03a, Ano03b, Ano03c, Ano03d, Ano03e, Ano03f, Ano03g, Ano03-29, Ano03-30, Ano03-31, Ano03-32, Ano03-33, Ano03-34, Ano03-35, Ano03-36, Ano03-37, Ano03-38, Ano03-39, Ano03-40, Ano03-41, Ano03-42, Ano04d, Ano04e, Ano04f, Ano04g, Ano04h, Ano04i, Ano04j, Ano04k, Ano04l, Ano04-37, Ano04-38, Ano04-39, Ano04-40, Ano04-41]. volume [Ano04-42, Ano04-43, Ano04-44, Ano04-48, Ano04-49, Ano04-50, Ano04-51, Ano04-52, Ano04-53, Ano04-54, Ano04-55, Ano05a, Ano05b, Ano05c, Ano05e, Ano05g, Ano05i, Ano05-32, Ano05-33, Ano05-34, Ano05-37, Ano05-39, Ano05-46, Ano05-47, Ano05-48, Ano05-51, Ano05-53, Ano06a, Ano06d, Ano06-28, BDH^{+}05, Cha04, GBA01, KMP09, LTA05, LLCS01, MOS00, MOS01, SLMS06, WHJ06, dNKM07]. volume-of-fluid [dNKM07]. VORPAL [MB04]. vortex [ZZ09]. vortices [TQ03]. voxel [LZC^{+}08]. voxel-based [LZC^{+}08]. Vries [Zak00b, KD09, ZY09]. vs [HJM02, PJSK08]. Vscape [vdB08]. VTF [EFBP04].

W [RP02]. waiting [DS01]. walk [Bal01, GG00]. walks [Jen01, JH09b, MP08, SFSH01]. wall [HYY07, JKKT00, KT07, MV04, SGK09, SPP^{+}04, SZ04, SBCZ08, WGL06, WS02, HS01a, SZ04].

Wall-Limiter [SZ04]. walled [YN05b]. walls [PLL07]. Wang [BR09, DGAG06, LWT08, LOL06, LWLL07, PRSB08, SWL09, WL08, YD07, Zha08]. Wannier [GFS03, MYL^{+}08]. wannier90 [MYL^{+}08]. water [Bac02, CTI07, DN05, EVL00, JBA05, MSS^{+}09, Sho04, Sho07]. WATERWAVES [TT06]. Wave [HTM01, ISSB01, KRTZ02, THM01, AJ08, ADG08, BB07, BS04b, Bow02, BBBR04, BDP00, BVKW02, Cai09, CR05, CYAS05, CWW06a, CGA^{+}07, CGVA08, CGVA09a, DKV00, EELZS04, FW01, FPB08, FA00, GIME02, GBD03, GME06, Jen00, JTS^{+}06, JG02, Kir06, KD09, KF03, LJ01, Lew04, bLpL02, LL04, LJ08, LJ09b, LKC06, pLbL03, LS01, yMS01, MCL05, Mas00, MNYY00a, MNYY00b, Mei01, Mic07, MS08b, MP01b, MSHP02, MSHP20, NM03, PDM^{+}08, PCV06, Sal03, SJP05, SKH02a, Sar00, SLMS06, Sea02b, SWS^{+}12, SWP03, TT06, WP06, WV05, YB02b, Zak00a]. wave-packet [Sal03]. wave-particle [yMS01]. wavefields [JBBR01]. wavefunction [TP01]. wavefunctions [AC05a, AC05b, SKF05]. wavelength [SLC09]. Wavelet [TK09, MA00, OS00b]. wavelets [SSP08a, SSP08b]. wavepacket [HSGBK08, MGG08]. waves [ABC^{+}03, DJ04, DEW01, FD03, HBRS05, Hon04, IDS^{+}04, KV07, PPP01, VKM^{+}05, VAH04, WBDB04, Zak00b, Zak01]. WAVR4 [KLTH04]. way [BSW^{+}07]. weak [AJ08, FSK04, HL00c, JG02]. weakly [GLMADB^{+}02, Lon07, MV04]. wearless [HOT07, HTM^{+}08]. Web [BCD^{+}01, CBM^{+}05, DBE^{+}04, KFJ^{+}09, KKHL07]. Web-based [DBE^{+}04, KKHL07]. Web-deployed [KFJ^{+}09]. wedge [BMML05]. weight [Blü09, De 02, FKAM05, GDC01]. weighted [KOS^{+}09, MTC07, SR01b]. weights [BBJW05]. well [CHL^{+}07, LLLZ01]. well-type [CHL^{+}07]. Wells [Wan00, Mam08, Moh08]. wet [Ger07]. wetness [Nii00]. Wetting [WS02, MMB02, Mül02, Pur02]. Wheeler [KNU00]. Where [OLX07]. which [BCV03, FGA04]. while [TPBE04, ZZH09]. Whistler [ADG08]. white [RRRHD08]. whither [Rap02a]. Whittaker [Nob04, AS03]. whole [CHL05]. Wide [BCD^{+}01, KLTH04, MNYY00a, Mic07, SM01]. wide-amplitude [KLTH04]. wide-energy [MNYY00a]. widely [BLM01]. width [Zak00a]. widths [EH06]. WIEN [Gao03, dlRL09]. WIEN2k [SBM02]. Wigner [PFG06a]. Wiley [Wan00]. Wilson [BKKS09, Cun09, JU09, MHK^{+}05]. window [CJC09]. windowed [CL08a]. windows [CNDC09, HC08, Hor09, JC01]. wing [VKN07]. winners [Ano04a, Ano04b, Ano04c, Gra02]. wire [EMJH03a, RG04]. WIRED [BCD^{+}01]. Wires [Wan00, FHR^{+}05, GPT08]. within [ADS06, BD05, FSB09, FKG00, GFG^{+}06, PLPS08, SS09a, SA09]. without [BH08, GSS06, Ort00]. WKB [KM06]. Wolff [GDAG05a, GDAG05b]. wonderful [Rap02b]. Woods [MAM04, MAM07]. work [Gre04]. workflow [BCC^{+}06]. World [BCD^{+}01, PB09a, Rap02b]. worldlines [MG09c]. worldwide [Shi07, Shi07]. would [NKV03]. WPHACT [ABM03]. written [Gro01]. WTC [ZLL09]. wurtzite [GRS06, Tsa02]. WW [JKW06, JKW00]. WWW [BB03]. Wynn [ÇHM00].

X [Hoo04, AGM^{+}00, BB07, BSO^{+}04, BG01, KMCS01, LZC^{+}08, Sal03, Vég04]. X-ray [AGM^{+}00, BB07, BSO^{+}04, BG01, KMCS01, LZC^{+}08, Sal03, Vég04]. Xe [MTJ02]. XLOOPS [BD02]. xmds [CD01b]. XML [GM00]. xPerm [MG08b]. XSummer [MU06]. xylo [BSB02]. xylo/phosphodiesteric [BSB02]. xylo/phosphonate [BSB02].

yambo [MHGV09]. Yang [MG09b, OK09]. YFSWW3 [JPS^{+}01a]. yielding [RR05]. Yin [OK09]. York [Koc02]. Yukawa [HJZ09]. Yutsis [VBFM05].

Zakharov [Wan09a]. Zassenhaus [WS09b]. ZEAL [KVR^{+}00]. Zeeman [AJ08, JG02]. zero [CSW02, Pis00]. zero-dimensional [Pis00]. zeroes [JK02, JJK05]. zeros [GSS06, KVR^{+}00]. ZFITTER [AAC^{+}06, BBC^{+}01b]. ZGB [LA09]. ZnO [GRS06]. Zobrist [MHS05]. zone [Zah04, Zah05]. zones [FBB01].

References

Adam:2000:ROO

[AA00] Gh. Adam and S. Adam. Reliable operations on oscillatory functions. Computer Physics Communications, 125(1–3):127– 141, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

              1879-2944 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/S0010465599004683. See com-

ment and reply [Ixa01, AA01b].

Adam:2001:IRG

[AA01a] Gh. Adam and S. Adam. Increasing reliability of Gauss– Kronrod quadrature by Eratosthenes’ sieve method. Computer Physics Communications, 135(3):261–277, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

              2944 (electronic). URL http://www.sciencedirect.com/
              science/article/pii/S0010465500002320.

Adam:2001:RCR

[AA01b] Gh. Adam and S. Adam. Reply to the “Comment on Reliable operations on oscillatory functions” by Dr. Ixaru: [Comput. Phys. Commun. 134 (2001)]. Computer Physics Communications, 134(2):269–272, February 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0010465500001776. See [AA00, Ixa01].

Ahnert:2007:NDE

[AA07] Karsten Ahnert and Markus Abel. Numerical differentiation of experimental data: local versus global methods. Computer Physics Communications, 177(10):764–774, November 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003116.

Ahmad:2008:ASN

[AA08] Bashir Ahmad and Badra S. Alghamdi. Approximation of solutions of the nonlinear Duffing equation involving both integral and non-integral forcing terms with separated boundary conditions. Computer Physics Communications, 179(6):409–416, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508001410.

Akishin:2000:SFD

[AAA^{+}00] P. G. Akishin, E. P. Akishina, P. Akritas, I. Antoniou, J. Ioannovich, and V. V. Ivanov. Stochastic filtering of digital images of skin micro-structure. Computer Physics Communications, 126(1–2):1–11, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004762.

Andonov:2006:SV

[AAB^{+}06] A. Andonov, A. Arbuzov, D. Bardin, S. Bondarenko, P. Christova, L. Kalinovskaya, G. Nanava, and W. von Schlippe. SANCscope-v.1.00. Computer Physics Communications, 174(6):481–517, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506000038. See erratum [AAB^{+}07].

Andonov:2007:ESV

[AAB^{+}07] A. Andonov, A. Arbuzov, D. Bardin, S. Bondarenko, P. Christova, L. Kalinovskaya, G. Nanava, and W. von Schlippe. Erratum to: “SANCscope-v.1.00” [Comput. Phys. Commun. 174 (2006) 481–517, hep-ph/0411186]. Computer Physics Communications, 177(7):623–624, October 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003128. See [AAB^{+}06].

Allison:2008:GVS

[AAB^{+}08] J. Allison, M. Asai, G. Barrand, M. Donszelmann, K. Minamimoto, J. Perl, S. Tanaka, E. Tcherniaev, and J. Tinslay. The Geant4 Visualisation System. Computer Physics Communications, 178(5):331–365, March 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507004213.

Arbuzov:2006:ZSA

[AAC^{+}06] A. B. Arbuzov, M. Awramik, M. Czakon, A. Freitas, M. W. Grünewald, K. Mönig, S. Riemann, and T. Riemann. ZFIT-TER: a semi-analytical program for fermion pair production in e^{+}e^{−} annihilation, from version 6.21 to version 6.42. Computer Physics Communications, 174(9):728–758, May 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506000105.

Almeida:2004:SPC

[AAG^{+}04] N. Almeida, R. Alemany, F. Glege, J. C. da Silva, and J. Varela. A software package for the configuration of hardware devices following a generic model. Computer Physics Communications, 163(1):41–52, October 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004072.

An:2007:DAS

[AAKL07] Jooneun An, Sunil Ahn, Sungho Kim, and Jysoo Lee. The design of the application and service platform for grid pilot services. Computer Physics Communications, 177(1–2): 214, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000483.

Amorim:2001:HBD

[AAM^{+}01] A. Amorim, Vasco Amaral, Umberto Marconi, Stefan Steinbeck, António Tomé, Vincenzo Vagnoni, Helmut Wolters, and HERA-B Collaboration. The HERA-B database services: for detector configuration, calibration, alignment, slow control and data classification. Computer Physics Communications, 140(1–2):172–178, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002685.

Adam:2003:RCQ

[AAP03] Gh Adam, S. Adam, and N. M. Plakida. Reliability conditions in quadrature algorithms. Computer Physics Communications, 154(1):49–64, July 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002820.

Arnold:2009:VPL

[ABB^{+}09] K. Arnold, M. Bähr, G. Bozzi, F. Campanario, C. Englert, T. Figy, N. Greiner, C. Hackstein, V. Hankele, B. Jäger, G. Klämke, M. Kubocz, C. Oleari, S. Plätzer, S. Prestel, M. Worek, and D. Zeppenfeld. Vbfnlo: a parton level Monte Carlo for processes with electroweak bosons. Computer Physics Communications, 180(9):1661–1670, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001027.

Antos:2001:DFT

[ABC^{+}01] Jaroslav Antos, M. Breitung, Tony Chan, Paoti Chang, Yen-Chu Chen, Troy Dawson, Jim Fromm, Lisa Giacchetti, Tanya Levshina, Igor Mandrichenko, Ray Pasetes, Marilyn Schweitzer, Karen Shepelak, Miroslav Siket, Dane Skow, Stephen Wolbers, G. P. Yeh, and Ping Yeh. Design and first tests of the CDF Run 2 farms. Computer Physics Communications, 140(1–2):239–245, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002752.

Amico:2003:PBB

[ABC^{+}03] P. Amico, L. Bosi, C. Cattuto, L. Gammaitoni, F. Marchesoni, M. Punturo, F. Travasso, and H. Vocca. A parallel Beowulf-based system for the detection of gravitational waves in interferometric detectors. Computer Physics Communications, 153(2):179–189, June 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002236.

Attig:2005:DSB

[ABD^{+}05] N. Attig, G. S. Bali, Th. Düssel, Th. Lippert, H. Neff, Z. Prkaçin, and K. Schilling. Demonstration of string breaking in quantum chromodynamics by large-scale eigenvalue computations. Computer Physics Communications, 169(1–3):382– 385, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001827.

Abe:2001:GDV

[Abe01] Tetsuo Abe. GRAPE-Dilepton (Version 1.1): a generator for dilepton production in ep collisions. Computer Physics Communications, 136(1–2):126–147, May 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002460.

Ansaloni:2000:PFC

[ABER00] Roberto Ansaloni, Gian Luigi Bendazzoli, Stefano Evangelisti, and Elda Rossi. A parallel full-CI algorithm. Computer Physics Communications, 128(1–2):496–515, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005421.

Anikeev:2001:EBL

[ABF^{+}01] K. Anikeev, G. Bauer, I. Furić, D. Holmgren, A. Korn, I. Kravchenko, M. Mulhearn, P. Ngan, Ch. Paus, A. Rakitin, R. Rechenmacher, T. Shah, P. Sphicas, K. Sumorok, S. Tether, J. Tseng, and F. Wüerthwein. Event Builder and Level 3 trigger at the CDF experiment. Computer Physics Communications, 140(1–2):110–116, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002612.

Accomando:2003:WFM

[ABM03] Elena Accomando, Alessandro Ballestrero, and Ezio Maina. WPHACT 2.0: a fully massive Monte Carlo generator for four fermion physics at e^{+}e^{−} colliders. Computer Physics Communications, 150(2):166–196, February 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006859.

Alsberg:2005:GOS

[ABNÅ05] Bjørn K. Alsberg, Håvard Bjerke, Gunn M. Navestad, and Per-Olof Åstrand. GaussDal: An open source database management system for quantum chemical computations. Computer Physics Communications, 171(2):133–153, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002869.

Alexandrescu:2009:MCP

[ABOSPG09] Adrian Alexandrescu, Alfonso Bueno-Orovio, José R. Salgueiro, and Víctor M. Pérez-García. Mapped Chebyshev pseudospectral method for the study of multiple scale phenomena. Computer Physics Communications, 180(6):912–919, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004360.

Awile:2012:FNL

[ABRS12] Omar Awile, Ferit Büyükkeçeci, Sylvain Reboux, and Ivo F. Sbalzarini. Fast neighbor lists for adaptive-resolution particle simulations. Computer Physics Communications, 183(5):1073– 1081, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465512000057.

Alouani-Bibi:2004:DFP

[ABSM04] F. Alouani-Bibi, M. M. Shoucri, and J.-P. Matte. Different Fokker–Planck approaches to simulate electron transport in plasmas. Computer Physics Communications, 164(1–3): 60–66, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002577.

Ahr:2002:MSI

[ABV02] M. Ahr, M. Biehl, and T. Volkmann. Modeling (001) surfaces of II–VI semiconductors. Computer Physics Communications, 147(1–2):107–110, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002266.

Angeli:2005:CAG

[AC05a] C. Angeli and R. Cimiraglia. Computer assisted generation of the matrix elements between contracted wavefunctions in a complete active space scheme. Computer Physics Communications, 166(1):53–57, February 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005090.

Angeli:2005:FMP

[AC05b] C. Angeli and R. Cimiraglia. FRODO: a MuPAD program to calculate matrix elements between contracted wavefunctions. Computer Physics Communications, 171(1):63–68, September 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002870.

Ahn:2007:EFC

[AC07] Sul-Ah Ahn and Myoung Won Cho. Exact and fast calculation for two-dimensional correlated electron systems: application to the Hubbard model. Computer Physics Communications, 177(1–2):40, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000471.

Argenti:2009:BSE

[AC09] Luca Argenti and Renato Colle. On the B-splines effective completeness. Computer Physics Communications, 180 (9):1442–1447, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000848.

Antchev:2001:CEB

[ACC^{+}01] G. Antchev, E. Cano, S. Cittolin, S. Erhan, B. Faure, D. Gigi, J. Gutleber, C. Jacobs, F. Meijers, E. Meschi, A. Ninane, L. Orsini, L. Pollet, A. Racz, D. Samyn, N. Sinanis, W. Schleifer, and P. Sphicas. The CMS event builder demonstrator and results with Myrinet. Computer Physics Communications, 140(1–2):130–138, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002648.

Atman:2009:DEG

[ACC09] A. P. F. Atman, P. Claudin, and G. Combe. Departure from elasticity in granular layers: Investigation of a crossover overload force. Computer Physics Communications, 180(4):612– 615, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004359.

Afanasev:2007:MMC

[ACIZ07] Andrei Afanasev, Eugene Chudakov, Alexander Ilyichev, and Vladimir Zykunov. MERADGEN 1.0: Monte Carlo generator for the simulation of radiative events in parity conserving doubly-polarized Møller scattering. Computer Physics Communications, 176(3):218–231, February 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003699.

Aichinger:2005:FOA

[ACK05] M. Aichinger, S. A. Chin, and E. Krotscheck. Fourthorder algorithms for solving local Schrödinger equations in a strong magnetic field. Computer Physics Communications, 171(3):197–207, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505003115.

Adamchik:2004:SNC

[Ada04] V. S. Adamchik. Symbolic and numeric computations of the Barnes function. Computer Physics Communications, 157(3):181–190, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004983.

Antonuccio-Delogu:2003:FPT

[ADBF03] V. Antonuccio-Delogu, U. Becciani, and D. Ferro. FLY: a parallel tree N-body code for cosmological simulations. Computer Physics Communications, 155(2):159–179, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550300345X.

Adye:2003:KHM

[ADD^{+}03] T. J. Adye, A. Dorigo, R. Dubitzky, A. Forti, S. J. Gowdy, G. Hamel de Monchenault, R. G. Jacobsen, D. Kirkby, S. Kluth, E. Leonardi, A. Salnikov, and L. Wilden. Kanga-ROO: Handling the micro-DST of the BaBar experiment with ROOT. Computer Physics Communications, 150(3):197–214, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006781.

Avellar:2007:DLF

[ADDdM07] J. Avellar, L. G. S. Duarte, S. E. S. Duarte, and L. A. C. P. da Mota. Determining Liouvillian first integrals for dynamical systems in the plane. Computer Physics Communications, 177(7):584–596, October 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550700255X.

Almehed:2002:RRE

[ADE^{+}02] S. Almehed, Ch. Driouichi, P. Eerola, U. Mjörnmark, O. Smirnova, Ch. Zacharatou Jarlskog, and T. Åkesson. Regional research exploitation of the LHC: a case-study of the required computing resources. Computer Physics Communications, 145(3):341–350, June 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002746.

Arnold:2008:SIH

[ADG08] Lukas Arnold, Jürgen Dreher, and Rainer Grauer. A semiimplicit Hall-MHD solver using Whistler wave preconditioning. Computer Physics Communications, 178(8):553–557, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004845.

Alexander:2003:PSA

[AdlT03] P. Alexander and A. de la Torre. A program for the simulation and analysis of open atmospheric balloon soundings. Computer Physics Communications, 151(1):96–120, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006951.

Ambrosch-Draxl:2006:LOP

[ADS06] Claudia Ambrosch-Draxl and Jorge O. Sofo. Linear optical properties of solids within the full-potential linearized augmented planewave method. Computer Physics Communications, 175(1):1–14, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001299.

Arter:2002:PES

[AEB02] Wayne Arter, James W. Eastwood, and Nicholas J. Brealey. A parallel elliptic solver for the EMX software. Computer Physics Communications, 144(1):23–28, March 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004696.

Alvarez:2005:SSS

[AEEdR05] G. Alvarez, A. Hernández Encinas, L. Hernández Encinas, and A. Martín del Rey. A secure scheme to share secret color images. Computer Physics Communications, 173(1–2):9–16, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004224.

Attig:2002:P

[AEK02] Norbert Attig, Rüdiger Esser, and Manfred Kremer. Preface. Computer Physics Communications, 147(1–2):xi–xvi, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001947.

Ablamowicz:2005:CGH

[AF05] Rafal Ablamowicz and Bertfried Fauser. Clifford and Graßmann Hopf algebras via the BIGEBRA package for Maple. Computer Physics Communications, 170(2):115–130, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002602.

Adler:2007:ERV

[AFK^{+}07] Joan Adler, J. Fox, R. Kalish, T. Mutat, A. Sorkin, and E. Warszarski. The essential role of visualization for modeling nanotubes and nanodiamond. Computer Physics Communications, 177(1–2):19–20, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000781.

Amar:2002:KSE

[AFP02] Jacques G. Amar, Fereydoon Family, and Mihail N. Popescu. Kinetics of submonolayer epitaxial growth. Computer Physics Communications, 146(1):1–8, June 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004289.

Alcubierre:2005:RSS

[AG05] Miguel Alcubierre and José A. González. Regularization of spherically symmetric evolution codes in numerical relativity. Computer Physics Communications, 167(2):76–84, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000597.

Abbasi:2007:VMU

[AGJJ07] H. Abbasi, M. Ghadimi, M. Jenab, and N. Javaheri. Vlasov model using kinetic phase point trajectories for the study of BGK modes. Computer Physics Communications, 177(1–2): 124, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001063.

Apostolakis:2000:PMX

[AGM^{+}00] J. Apostolakis, S. Giani, M. Maire, A. V. Bagulya, and V. M. Grichine. Parameterization models for X-ray transition radiation in the GEANT4 package. Computer Physics Communications, 132(3):241–250, November 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001545.

Anderson:2007:QMC

[AGS07] Amos G. Anderson, William A. Goddard III, and Peter Schröder. Quantum Monte Carlo on graphical processing units. Computer Physics Communications, 177(3):298–306, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001993.

Abrahamyan:2000:TOO

[AGV00] Susanna Abrahamyan and José Gómez-Valdés. Towards an object-oriented framework for the advection-diffusion model. Computer Physics Communications, 132(3):224–231, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001508.

Arnold:2002:MFA

[AH02] Axel Arnold and Christian Holm. MMM2D: a fast and accurate summation method for electrostatic interactions in 2D slab geometries. Computer Physics Communications, 148(3):327–348, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005866.

Allfrey:2003:RAN

[AH03] S. J. Allfrey and R. Hatzky. A revised δf algorithm for nonlinear PIC simulation. Computer Physics Communications, 154(2):98–104, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002881.

Alkofer:2009:ADD

[AHS09] R. Alkofer, M. Q. Huber, and K. Schwenzer. Algorithmic derivation of Dyson–Schwinger equations. Computer Physics Communications, 180(6):965–976, June 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004268.

Abu-Ibrahim:2003:CSC

[AIOST03] B. Abu-Ibrahim, Y. Ogawa, Y. Suzuki, and I. Tanihata. Cross section calculations in Glauber model: I. Core plus one-nucleon case. Computer Physics Communications, 151(3):369–386, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007348.

Andersson:2008:HPC

[AJ08] Martin Andersson and Per Jönsson. HFSZEEMAN — a program for computing weak and intermediate field fine and hyperfine structure Zeeman splittings from MCDHF wave functions. Computer Physics Communications, 178(2):156–170, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700392X.

Alemany:2007:EFP

[AJT^{+}07] M. M. G. Alemany, Manish Jain, Murilo L. Tiago, Yunkai Zhou, Yousef Saad, and James R. Chelikowsky. Efficient firstprinciples calculations of the electronic structure of periodic systems. Computer Physics Communications, 177(4):339–347, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002202.

Auer:2003:CMF

[AK03] J. Auer and E. Krotscheck. Collective mode formulation of the response algorithm for solving Kohn–Sham equations. Computer Physics Communications, 151(3):265–271, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007361.

Aydin:2007:SMS

[AK07]  A. Aydin and B. Karasözen.  Symplectic and multi-

symplectic methods for coupled nonlinear Schrödinger equations with periodic solutions. Computer Physics Communications, 177(7):566–583, October 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507002500.

Alexeev:2002:DDS

[AKG02] Yuri Alexeev, Ricky A. Kendall, and Mark S. Gordon. The distributed data SCF. Computer Physics Communications, 143(1):69–82, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501004398.

Avdelas:2001:GHE

[AKS01] G. Avdelas, A. Konguetsof, and T. E. Simos. A generator of hybrid explicit methods for the numerical solution of the Schrödinger equation and related problems. Computer Physics Communications, 136(1–2):14–28, May 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002496.

Avdelas:2002:GDM

[AKS02] G. Avdelas, A. Konguetsof, and T. E. Simos. A generator of dissipative methods for the numerical solution of the Schrödinger equation. Computer Physics Communications, 148(1):59–73, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200468X.

Amirkhanov:2000:FMP

[AKZ00] I. V. Amirkhanov, S. V. Konnova, and E. P. Zhidkov. The factorization method and particular solutions of the relativistic Schrödinger equation of n th order ( n=4,6 ). Computer Physics Communications, 126(1–2):12–15, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900421X.

Ahn:2008:NTL

[AL08a] Cheol-O Ahn and Sang Hwan Lee. A new treecode for long-range force calculation. Computer Physics Communications, 178(2):121–127, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003797.

Allanach:2008:SUB

[AL08b] Benjamin C. Allanach and Christopher G. Lester. Sampling using a ‘bank’ of clues. Computer Physics Communications, 179(4):256–266, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000994.

Alford:2005:CGL

[Alf05] J. A. Alford. Calculation of the generalized leaky aquifer integral. Computer Physics Communications, 173(1–2):1–8, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004029.

Alfe:2009:PPC

[Alf09] Dario Alfè. PHON: a program to calculate phonons using the small displacement method. Computer Physics Communications, 180(12):2622–2633, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001064.

Allton:2001:RLQ

[All01] Chris Allton. Recent lattice QCD results from the UKQCD collaboration. Computer Physics Communications, 142(1– 3):168–171, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003174.

Allanach:2002:SPC

[All02] B. C. Allanach. SOFTSUSY: a program for calculating supersymmetric spectra. Computer Physics Communications, 143(3):305–331, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550100460X.

Allen:2005:CSM

[All05] Michael P. Allen. Computer simulation of multiscale phenomena in colloidal liquid crystals. Computer Physics Communications, 169(1–3):433–437, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001931.

Attig:2001:GEL

[ALN^{+}01] N. Attig, Th. Lippert, H. Neff, J. Negele, and K. Schilling. Giant eigenproblems from lattice gauge theory on CRAY T3E systems. Computer Physics Communications, 142(1– 3):196–200, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003290.

Alet:2005:CAL

[ALV05] Fabien Alet, Biagio Lucini, and Michele Vettorazzo. A cluster algorithm for lattice gauge theories. Computer Physics Communications, 169(1–3):370–373, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001797.

Alvarez:2009:DMR

[Alv09] G. Alvarez. The density matrix renormalization group for strongly correlated electron systems: a generic implementation. Computer Physics Communications, 180(9):1572–1578, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550900071X.

Amirkhanov:2000:NST

[AMP^{+}00] I. V. Amirkhanov, A. I. Machavariani, I. V. Puzynin, T. P. Puzynina, T. A. Strizh, and E. V. Zemlyanaya. Numerical solution of two-body relativistic equations for the bound-state problem with confining and Coulomb potentials. Computer Physics Communications, 126(1–2):16–21, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002234.

Alvarez:2004:KCC

[ÁMRP04] G. Álvarez, F. Montoya, M. Romera, and G. Pastor. Keystream cryptanalysis of a chaotic cryptographic method. Computer Physics Communications, 156(2):205–207, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004326. See [kWpLwW01, WLW04].

Anonymous:2000:BR

[Ano00a] Anonymous. Book review. Computer Physics Communications, 124(1):123, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004725.

Anonymous:2000:Ia

[Ano00b] Anonymous. Index. Computer Physics Communications, 124 (2–3):363–365, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599005445.

Anonymous:2000:Ib

[Ano00c] Anonymous. Index. Computer Physics Communications, 124 (2–3):366–367, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599005457.

Anonymous:2000:Ic

[Ano00d] Anonymous. Index. Computer Physics Communications, 125(1–3):316–317, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000035.

Anonymous:2000:Id

[Ano00e] Anonymous. Index. Computer Physics Communications, 125(1–3):318–319, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000047.

Anonymous:2000:Ie

[Ano00f] Anonymous. Index. Computer Physics Communications, 125 (1–3):320, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000059.

Anonymous:2000:If

[Ano00g] Anonymous. Index. Computer Physics Communications, 126(1–2):182–185, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000412.

Anonymous:2000:Ig

[Ano00h] Anonymous. Index. Computer Physics Communications, 126(3):355–357, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001284.

Anonymous:2000:Ih

[Ano00i] Anonymous. Index. Computer Physics Communications, 126(3):358–361, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001296.

Anonymous:2000:Ii

[Ano00j] Anonymous. Index. Computer Physics Communications, 126 (3):362, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001302.

Anonymous:2000:Ij

[Ano00k] Anonymous. Index. Computer Physics Communications, 127(1):173–174, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550000134X.

Anonymous:2000:Ik

[Ano00l] Anonymous. Index. Computer Physics Communications, 127 (2–3):360–362, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001417.

Anonymous:2000:Il

[Ano00m] Anonymous. Index. Computer Physics Communications, 127(2–3):363, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001429.

Anonymous:2000:Im

[Ano00n] Anonymous. Index. Computer Physics Communications, 128(1–2):527–530, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001491.

Anonymous:2000:In

[Ano00o] Anonymous. Index. Computer Physics Communications, 128 (3):639–641, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001624.

Anonymous:2000:Io

[Ano00p] Anonymous. Index. Computer Physics Communications, 128 (3):642–645, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001636.

Anonymous:2000:Ip

[Ano00q] Anonymous. Index. Computer Physics Communications, 128(3):646, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001648.

Anonymous:2000:Iq

[Ano00r] Anonymous. Index. Computer Physics Communications, 129(1–3):289–291, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001697.

Anonymous:2000:Ir

[Ano00s] Anonymous. Index. Computer Physics Communications, 130 (3):329–330, August 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001855.

Anonymous:2000:Is

[Ano00t] Anonymous. Index. Computer Physics Communications, 130(3):331, August 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001867.

Anonymous:2000:It

[Ano00u] Anonymous. Index. Computer Physics Communications, 131 (3):291–293, September 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001983.

Anonymous:2000:Iu

[Ano00v] Anonymous. Index. Computer Physics Communications, 131(3):294, September 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001995.

Anonymous:2000:Iv

[Ano00w] Anonymous. Index. Computer Physics Communications, 132 (3):317–318, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002174.

Anonymous:2000:Iw

[Ano00x] Anonymous. Index. Computer Physics Communications, 132 (3):319–320, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002186.

Anonymous:2000:Ix

[Ano00y] Anonymous. Index. Computer Physics Communications, 132(3):321, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002198.

Anonymous:2000:IAR

[Ano00z] Anonymous. Instructions to authors (1999 revision). Computer Physics Communications, 124(1):ix–xv, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599900000.

Anonymous:2000:P

[Ano00-27] Anonymous. Preface. Computer Physics Communications, 126(1–2):xi, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599005263.

Anonymous:2001:ADC

[Ano01a] Anonymous. The ALICE DAQ: current status and future challenges. Computer Physics Communications, 140(1–2):117–129, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002624.

Anonymous:2001:AIa

[Ano01b] Anonymous. Author index. Computer Physics Communications, 137(1):223–224, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002417.

Anonymous:2001:AIb

[Ano01c] Anonymous. Author index. Computer Physics Communications, 137(3):433–435, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501002831.

Anonymous:2001:AIc

[Ano01d] Anonymous. Author index. Computer Physics Communications, 139(1):151–152, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004052.

Anonymous:2001:AId

[Ano01e] Anonymous. Author index. Computer Physics Communications, 140(1–2):287–292, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004210.

Anonymous:2001:AIe

[Ano01f] Anonymous. Author index. Computer Physics Communications, 142(1–3):473–479, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004489.

Anonymous:2001:AIVa

[Ano01g] Anonymous. Author index to volume 134. Computer Physics Communications, 134(3):429–431, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001394.

Anonymous:2001:AIVb

[Ano01h] Anonymous. Author index to volume 135. Computer Physics Communications, 135(3):380–382, April 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001941.

Anonymous:2001:AIVc

[Ano01i] Anonymous. Author index to volume 136. Computer Physics Communications, 136(3):353–355, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002272.

Anonymous:2001:AIVd

[Ano01j] Anonymous. Author index to volume 138. Computer Physics Communications, 138(3):306–308, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003940.

Anonymous:2001:AIVe

[Ano01k] Anonymous. Author index to volume 139. Computer Physics Communications, 139(3):331–333, October 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004088.

Anonymous:2001:AIVf

[Ano01l] Anonymous. Author index to volume 140. Computer Physics Communications, 140(3):517–523, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004258.

Anonymous:2001:AIVg

[Ano01m] Anonymous. Author index to volume 141. Computer Physics Communications, 141(3):430–432, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004520.

Anonymous:2001:AAC

[Ano01n] Anonymous. Automatic analysis of CHIMERA experimental data by means of a hierarchical pre-attentive neural system. Computer Physics Communications, 140(1–2):13–20, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550100251X.

Anonymous:2001:Ca

[Ano01o] Anonymous. Contents. Computer Physics Communications, 139(1):xiii–xiv, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004064.

Anonymous:2001:Cb

[Ano01p] Anonymous. Contents. Computer Physics Communications, 140(1–2):xiii–xv, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004234.

Anonymous:2001:Cc

[Ano01q] Anonymous. Contents. Computer Physics Communications, 142(1–3):xiii–xviii, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004507.

Anonymous:2001:CVa

[Ano01r] Anonymous. Contents to volume 134. Computer Physics Communications, 134(3):426–428, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001382.

Anonymous:2001:CVb

[Ano01s] Anonymous. Contents to volume 135. Computer Physics Communications, 135(3):377–379, April 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100193X.

Anonymous:2001:CVc

[Ano01t] Anonymous. Contents to volume 136. Computer Physics Communications, 136(3):351–352, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002260.

Anonymous:2001:CVd

[Ano01u] Anonymous. Contents to volume 137. Computer Physics Communications, 137(3):429–431, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550100282X.

Anonymous:2001:CVe

[Ano01v] Anonymous. Contents to volume 138. Computer Physics Communications, 138(3):303–305, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003939.

Anonymous:2001:CVf

[Ano01w] Anonymous. Contents to volume 139. Computer Physics Communications, 139(3):327–329, October 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004076.

Anonymous:2001:CVg

[Ano01x] Anonymous. Contents to volume 140. Computer Physics Communications, 140(3):513–516, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004246.

Anonymous:2001:CVh

[Ano01y] Anonymous. Contents to volume 141. Computer Physics Communications, 141(3):427–429, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004519.

Anonymous:2001:F

[Ano01z] Anonymous. Foreword. Computer Physics Communications, 140(1–2):xi–xii, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501004222.

Anonymous:2001:Ia

[Ano01-27] Anonymous. Index. Computer Physics Communications, 133 (2–3):407–408, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002204.

Anonymous:2001:Ib

[Ano01-28] Anonymous. Index. Computer Physics Communications, 133 (2–3):409–410, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002216.

Anonymous:2001:Ic

[Ano01-29] Anonymous. Index. Computer Physics Communications, 133(2–3):411, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002228.

Anonymous:2001:P

[Ano01-30] Anonymous. Preface. Computer Physics Communications, 142 (1–3):xi–xii, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003034.

Anonymous:2001:PIVa

[Ano01-31] Anonymous. Program index to volume 134. Computer Physics Communications, 134(3):432, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001400.

Anonymous:2001:PIVb

[Ano01-32] Anonymous. Program index to volume 135. Computer Physics Communications, 135(3):383–384, April 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001953.

Anonymous:2001:PIVc

[Ano01-33] Anonymous. Program index to volume 136. Computer Physics Communications, 136(3):356–357, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002284.

Anonymous:2001:PIVd

[Ano01-34] Anonymous. Program index to volume 137. Computer Physics Communications, 137(3):437–438, July 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002843.

Anonymous:2001:PIVe

[Ano01-35] Anonymous. Program index to volume 138. Computer Physics Communications, 138(3):309, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003952.

Anonymous:2001:PIVf

[Ano01-36] Anonymous. Program index to volume 139. Computer Physics Communications, 139(3):335–336, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100409X.

Anonymous:2001:PIVg

[Ano01-37] Anonymous. Program index to volume 140. Computer Physics Communications, 140(3):524, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100426X.

Anonymous:2001:PIVh

[Ano01-38] Anonymous. Program index to volume 141. Computer Physics Communications, 141(3):433–434, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004532.

Anonymous:2001:PA

[Ano01-39] Anonymous. Publisher’s announcement. Computer Physics Communications, 135(1):ix, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001424.

Anonymous:2002:ACP

[Ano02a] Anonymous. APE computers-past, present and future. Computer Physics Communications, 147(1–2):402–409, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003144.

Anonymous:2002:AIa

[Ano02b] Anonymous. Author index. Computer Physics Communications, 145(1):203–204, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004034.

Anonymous:2002:AIb

[Ano02c] Anonymous. Author index. Computer Physics Communications, 146(1):137–138, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502005295.

Anonymous:2002:AIVa

[Ano02d] Anonymous. Author index to volume 143. Computer Physics Communications, 143(3):335–337, March 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001650.

Anonymous:2002:AIVb

[Ano02e] Anonymous. Author index to volume 144. Computer Physics Communications, 144(3):350–352, April 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004101.

Anonymous:2002:AIVc

[Ano02f] Anonymous. Author index to volume 145. Computer Physics Communications, 145(3):398–400, June 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005039.

Anonymous:2002:AIVd

[Ano02g] Anonymous. Author index to volume 146. Computer Physics Communications, 146(3):349–352, July 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005763.

Anonymous:2002:AIVe

[Ano02h] Anonymous. Author index to volume 147. Computer Physics Communications, 147(3):870–883, September 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006392.

Anonymous:2002:AIVf

[Ano02i] Anonymous. Author index to volume 148. Computer Physics Communications, 148(3):356–358, November 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007221.

Anonymous:2002:CPCa

[Ano02j] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 147(3):iii–v, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006379.

Anonymous:2002:CPCb

[Ano02k] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 148(1):v–vii, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006549.

Anonymous:2002:CPCc

[Ano02l] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 148(2):iii–v, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006641.

Anonymous:2002:CPCd

[Ano02m] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 148(3):iii–v, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007208.

Anonymous:2002:CPCe

[Ano02n] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 149(1):v–vii, November 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007464.

Anonymous:2002:CPCf

[Ano02o] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 149(2):iii–v, December 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200766X.

Anonymous:2002:Ca

[Ano02p] Anonymous. Contents. Computer Physics Communications, 145(1):xiii–xiv, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004502.

Anonymous:2002:Cb

[Ano02q] Anonymous. Contents. Computer Physics Communications, 146(1):xiii–xiv, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502005362.

Anonymous:2002:Cc

[Ano02r] Anonymous. Contents. Computer Physics Communications, 147(1–2):xvii–xxvi, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550200615X.

Anonymous:2002:CVa

[Ano02s] Anonymous. Contents to volume 143. Computer Physics Communications, 143(3):333–334, March 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001649.

Anonymous:2002:CVb

[Ano02t] Anonymous. Contents to volume 144. Computer Physics Communications, 144(3):347–349, April 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004095.

Anonymous:2002:CVc

[Ano02u] Anonymous. Contents to volume 145. Computer Physics Communications, 145(3):395–397, June 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502005027.

Anonymous:2002:CVd

[Ano02v] Anonymous. Contents to volume 146. Computer Physics Communications, 146(3):345–348, July 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005751.

Anonymous:2002:CVe

[Ano02w] Anonymous. Contents to volume 147. Computer Physics Communications, 147(3):859–869, September 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006380.

Anonymous:2002:CVf

[Ano02x] Anonymous. Contents to volume 148. Computer Physics Communications, 148(3):353–355, November 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200721X.

Anonymous:2002:PIVa

[Ano02y] Anonymous. Program index to volume 143. Computer Physics Communications, 143(3):338, March 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001662.

Anonymous:2002:PIVb

[Ano02z] Anonymous. Program index to volume 144. Computer Physics Communications, 144(3):353–354, April 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004113.

Anonymous:2002:PIVc

[Ano02-27] Anonymous. Program index to volume 145. Computer Physics Communications, 145(3):401, June 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005040.

Anonymous:2002:PIVd

[Ano02-28] Anonymous. Program index to volume 146. Computer Physics Communications, 146(3):353, July 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005775.

Anonymous:2002:PIVe

[Ano02-29] Anonymous. Program index to volume 147. Computer Physics Communications, 147(3):884, September 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006409.

Anonymous:2002:PIVf

[Ano02-30] Anonymous. Program index to volume 148. Computer Physics Communications, 148(3):359, November 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007233.

Anonymous:2003:AIVa

[Ano03a] Anonymous. Author index to volume 149. Computer Physics Communications, 149(3):197–198, January 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200783X.

Anonymous:2003:AIVb

[Ano03b] Anonymous. Author index to volume 150. Computer Physics Communications, 150(3):326–328, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008585.

Anonymous:2003:AIVc

[Ano03c] Anonymous. Author index to volume 151. Computer Physics Communications, 151(3):389–391, April 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503001723.

Anonymous:2003:AIVd

[Ano03d] Anonymous. Author index to volume 152. Computer Physics Communications, 152(3):343–344, May 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002340.

Anonymous:2003:AIVe

[Ano03e] Anonymous. Author index to volume 153. Computer Physics Communications, 153(3):512–515, July 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003084.

Anonymous:2003:AIVf

[Ano03f] Anonymous. Author index to volume 154. Computer Physics Communications, 154(3):237–238, August 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003801.

Anonymous:2003:AIVg

[Ano03g] Anonymous. Author index to volume 155. Computer Physics Communications, 155(3):263–265, November 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004491.

Anonymous:2003:BCC

[Ano03h] Anonymous. A Beowulf-class computing cluster for the Monte Carlo production of the LHCb experiment. Computer Physics Communications, 150(2):129–139, February 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200677X.

Anonymous:2003:CPCa

[Ano03i] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 149(3):iii–v, January 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007816.

Anonymous:2003:CPCb

[Ano03j] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 150(1):v–vii, January 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502008093.

Anonymous:2003:CPCc

[Ano03k] Anonymous. Computer physics communications — list of editors. Computer Physics Communications, 150(2):iii–v, February 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502008330.

Anonymous:2003:CPCd

[Ano03l] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 150(3):iii–v, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502008561.

Anonymous:2003:CPCe

[Ano03m] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 151(1):v–vii, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502008871.

Anonymous:2003:CPCf

[Ano03n] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 151(2):iii–v, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503001449.

Anonymous:2003:CPCg

[Ano03o] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 151(3):iii–v, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550300170X.

Anonymous:2003:CPCh

[Ano03p] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 152(1):v–vii, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503001887.

Anonymous:2003:CPCi

[Ano03q] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 152(2):iii–v, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002169.

Anonymous:2003:CPCj

[Ano03r] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 152(3):iii–v, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002327.

Anonymous:2003:CPCk

[Ano03s] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 153(1):v–vii, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002558.

Anonymous:2003:CPCl

[Ano03t] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 153(2):iii–v, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002686.

Anonymous:2003:CPCm

[Ano03u] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 153(3):v–vii, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003060.

Anonymous:2003:CPCn

[Ano03v] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 154(1):v–vii, July 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003369.

Anonymous:2003:CPCo

[Ano03w] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 154(2):iii–v, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550300359X.

Anonymous:2003:CPCp

[Ano03x] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 154(3):iii–v, August 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503003783.

Anonymous:2003:CPCq

[Ano03y] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 155(1):v–vii, September 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004077.

Anonymous:2003:CPCr

[Ano03z] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 155(2):iii–v, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004193.

Anonymous:2003:CPCs

[Ano03-27] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 155(3):iii–v, November 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004478.

Anonymous:2003:CPCt

[Ano03-28] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 156(1):v–vii, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004764.

Anonymous:2003:CVa

[Ano03-29] Anonymous. Contents to volume 149. Computer Physics Communications, 149(3):195–196, January 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502007828.

Anonymous:2003:CVb

[Ano03-30] Anonymous. Contents to volume 150. Computer Physics Communications, 150(3):324–325, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008573.

Anonymous:2003:CVc

[Ano03-31] Anonymous. Contents to volume 151. Computer Physics Communications, 151(3):387–388, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503001711.

Anonymous:2003:CVd

[Ano03-32] Anonymous. Contents to volume 152. Computer Physics Communications, 152(3):341–342, May 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002339.

Anonymous:2003:CVe

[Ano03-33] Anonymous. Contents to volume 153. Computer Physics Communications, 153(3):509–511, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503003072.

Anonymous:2003:CVf

[Ano03-34] Anonymous. Contents to volume 154. Computer Physics Communications, 154(3):235–236, August 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003795.

Anonymous:2003:CVg

[Ano03-35] Anonymous. Contents to volume 155. Computer Physics Communications, 155(3):261–262, November 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550300448X.

Anonymous:2003:PIVa

[Ano03-36] Anonymous. Program index to volume 149. Computer Physics Communications, 149(3):199, January 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007841.

Anonymous:2003:PIVb

[Ano03-37] Anonymous. Program index to volume 150. Computer Physics Communications, 150(3):329, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008597.

Anonymous:2003:PIVc

[Ano03-38] Anonymous. Program index to volume 151. Computer Physics Communications, 151(3):392–393, April 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503001735.

Anonymous:2003:PIVd

[Ano03-39] Anonymous. Program index to volume 152. Computer Physics Communications, 152(3):345, May 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002352.

Anonymous:2003:PIVe

[Ano03-40] Anonymous. Program index to volume 153. Computer Physics Communications, 153(3):516–517, July 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003096.

Anonymous:2003:PIVf

[Ano03-41] Anonymous. Program index to volume 154. Computer Physics Communications, 154(3):239, August 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003813.

Anonymous:2003:PIVg

[Ano03-42] Anonymous. Program index to volume 155. Computer Physics Communications, 155(3):266, November 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004508.

Anonymous:2003:PNa

[Ano03-43] Anonymous. Publisher’s note. Computer Physics Communications, 150(3):vii, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502008706.

Anonymous:2003:PNb

[Ano03-44] Anonymous. Publisher’s note. Computer Physics Communications, 155(2):vii, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004296.

Anonymous:2004:BGS

[Ano04a] Anonymous. 2003 best graduate student presentation award winners. Computer Physics Communications, 164(1–3):xxiii– xxiv, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003388.

Anonymous:2004:JDP

[Ano04b] Anonymous. 2003 John Dawson Prize winners for “pioneering advances to plasma physics obtained through simulations”. Computer Physics Communications, 164(1–3):xiv–xv, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003340.

Anonymous:2004:OBA

[Ano04c] Anonymous. 2003 Oscar Buneman Award winners. Computer Physics Communications, 164(1–3):xxi–xxii, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003376.

Anonymous:2004:AIVa

[Ano04d] Anonymous. Author index to volume 156. Computer Physics Communications, 156(3):326–328, January 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503005307.

Anonymous:2004:AIVb

[Ano04e] Anonymous. Author index to volume 157. Computer Physics Communications, 157(3):256–257, March 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000438.

Anonymous:2004:AIVc

[Ano04f] Anonymous. Author index to volume 158. Computer Physics Communications, 158(3):198–199, April 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001195.

Anonymous:2004:AIVd

[Ano04g] Anonymous. Author index to volume 159. Computer Physics Communications, 159(3):247–248, June 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001791.

Anonymous:2004:AIVe

[Ano04h] Anonymous. Author index to volume 160. Computer Physics Communications, 160(3):245–246, July 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002395.

Anonymous:2004:AIVf

[Ano04i] Anonymous. Author index to volume 161. Computer Physics Communications, 161(3):188–189, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003595.

Anonymous:2004:AIVg

[Ano04j] Anonymous. Author index to volume 162. Computer Physics Communications, 162(3):220–221, October 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004230.

Anonymous:2004:AIVh

[Ano04k] Anonymous. Author index to volume 163. Computer Physics Communications, 163(3):205–206, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004862.

Anonymous:2004:AIVi

[Ano04l] Anonymous. Author index to volume 164. Computer Physics Communications, 164(1–3):483–491, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550400517X.

Anonymous:2004:CPCa

[Ano04m] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 156(2):iii–v, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503005058.

Anonymous:2004:CPCb

[Ano04n] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 156(3):iii–v, January 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503005289.

Anonymous:2004:CPCc

[Ano04o] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 157(1):v–vii, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000116.

Anonymous:2004:CPCd

[Ano04p] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 157(2):iii–v, February 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000281.

Anonymous:2004:CPCe

[Ano04q] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 157(3):iii–v, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504000414.

Anonymous:2004:CPCf

[Ano04r] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 158(1):v–vii, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550400061X.

Anonymous:2004:CPCg

[Ano04s] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 158(2):iii–vi, April 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504000888.

Anonymous:2004:CPCh

[Ano04t] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 158(3):iii–vi, April 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001171.

Anonymous:2004:CPCi

[Ano04u] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 159(1):v–viii, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001377.

Anonymous:2004:CPCj

[Ano04v] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 159(2):iii–vi, May 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001535.

Anonymous:2004:CPCk

[Ano04w] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 159(3):iii–vi, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001778.

Anonymous:2004:CPCl

[Ano04x] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 160(1):v–viii, June 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001912.

Anonymous:2004:CPCm

[Ano04y] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 160(2):iii–vi, July 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002218.

Anonymous:2004:CPCn

[Ano04z] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 160(3):iii–vi, July 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002371.

Anonymous:2004:CPCo

[Ano04-27] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 161(1–2):v–viii, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003443.

Anonymous:2004:CPCp

[Ano04-28] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 161(3):iii–vi, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003698.

Anonymous:2004:CPCq

[Ano04-29] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 162(1):v– viii, September 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003807.

Anonymous:2004:CPCr

[Ano04-30] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 162(2):iii– vi, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003960.

Anonymous:2004:CPCs

[Ano04-31] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 162(3):iii–vi, October 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004217.

Anonymous:2004:CPCt

[Ano04-32] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 163(1):v–viii, October 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004461.

Anonymous:2004:CPCu

[Ano04-33] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 163(2):iii–vi, November 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004692.

Anonymous:2004:CPCv

[Ano04-34] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 163(3):iii–vi, November 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004849.

Anonymous:2004:CPCw

[Ano04-35] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 164(1–3): v–viii, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504005168.

Anonymous:2004:C

[Ano04-36] Anonymous. Contents. Computer Physics Communications, 164(1–3):xxv–xxx, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005259.

Anonymous:2004:CVa

[Ano04-37] Anonymous. Contents to volume 156. Computer Physics Communications, 156(3):323–325, January 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503005290.

Anonymous:2004:CVb

[Ano04-38] Anonymous. Contents to volume 157. Computer Physics Communications, 157(3):254–255, March 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000426.

Anonymous:2004:CVc

[Ano04-39] Anonymous. Contents to volume 158. Computer Physics Communications, 158(3):196–197, April 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001183.

Anonymous:2004:CVd

[Ano04-40] Anonymous. Contents to volume 159. Computer Physics Communications, 159(3):245–246, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550400178X.

Anonymous:2004:CVe

[Ano04-41] Anonymous. Contents to volume 160. Computer Physics Communications, 160(3):243–244, July 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002383.

Anonymous:2004:CVf

[Ano04-42] Anonymous. Contents to volume 161. Computer Physics Communications, 161(3):186–187, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003583.

Anonymous:2004:CVg

[Ano04-43] Anonymous. Contents to volume 162. Computer Physics Communications, 162(3):218–219, October 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004229.

Anonymous:2004:CVh

[Ano04-44] Anonymous. Contents to volume 163. Computer Physics Communications, 163(3):203–204, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004850.

Anonymous:2004:RJD

[Ano04-45] Anonymous. In remembrance of John Dawson. Computer Physics Communications, 164(1–3):xii–xiii, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003339.

Anonymous:2004:NTC

[Ano04-46] Anonymous. The national transport code collaboration module library. Computer Physics Communications, 164(1–3): 108–113, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002668.

Anonymous:2004:P

[Ano04-47] Anonymous. Preface. Computer Physics Communications, 164(1–3):xi, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003327.

Anonymous:2004:PIVa

[Ano04-48] Anonymous. Program index to volume 156. Computer Physics Communications, 156(3):329, January 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503005319.

Anonymous:2004:PIVb

[Ano04-49] Anonymous. Program index to volume 157. Computer Physics Communications, 157(3):258–259, March 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400044X.

Anonymous:2004:PIVc

[Ano04-50] Anonymous. Program index to volume 158. Computer Physics Communications, 158(3):200–201, April 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001201.

Anonymous:2004:PIVd

[Ano04-51] Anonymous. Program index to volume 159. Computer Physics Communications, 159(3):249–250, June 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001808.

Anonymous:2004:PIVe

[Ano04-52] Anonymous. Program index to volume 160. Computer Physics Communications, 160(3):247–248, July 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002401.

Anonymous:2004:PIVf

[Ano04-53] Anonymous. Program index to volume 161. Computer Physics Communications, 161(3):190, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003601.

Anonymous:2004:PIVg

[Ano04-54] Anonymous. Program index to volume 162. Computer Physics Communications, 162(3):222–223, October 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004242.

Anonymous:2004:PIVh

[Ano04-55] Anonymous. Program index to volume 163. Computer Physics Communications, 163(3):207–208, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004874.

Anonymous:2004:RBL

[Ano04-56] Anonymous. Remarks by bruce langdon regarding the dawson prize. Computer Physics Communications, 164(1–3): xix–xx, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003364.

Anonymous:2004:RCK

[Ano04-57] Anonymous. Remarks by charles K. (ned) birdsall regarding the dawson prize. Computer Physics Communications, 164(1–3):xvi–xviii, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003352.

Anonymous:2005:AIVa

[Ano05a] Anonymous. Author index to volume 165. Computer Physics Communications, 165(3):273–274, February 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000081.

Anonymous:2005:AIVb

[Ano05b] Anonymous. Author index to volume 166. Computer Physics Communications, 166(3):243–244, March 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000494.

Anonymous:2005:AIVc

[Ano05c] Anonymous. Author index to volume 167. Computer Physics Communications, 167(3):255–257, May 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002183.

Anonymous:2005:AIVd

[Ano05d] Anonymous. Author index to volume 168. Computer Physics Communications, 168(3):235–236, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003012.

Anonymous:2005:AIVe

[Ano05e] Anonymous. Author index to volume 169. Computer Physics Communications, 169(1–3):467–475, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003292.

Anonymous:2005:AIVf

[Ano05f] Anonymous. Author index to volume 170. Computer Physics Communications, 170(3):301–302, August 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500384X.

Anonymous:2005:AIVg

[Ano05g] Anonymous. Author index to volume 171. Computer Physics Communications, 171(3):221–222, October 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500456X.

Anonymous:2005:AIVh

[Ano05h] Anonymous. Author index to volume 172. Computer Physics Communications, 172(3):217–218, November 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005242.

Anonymous:2005:AIVi

[Ano05i] Anonymous. Author index to volume 173. Computer Physics Communications, 173(3):195–196, December 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005588.

Anonymous:2005:CC

[Ano05j] Anonymous. CCP 2004 committees. Computer Physics Communications, 169(1–3):xiii–xiv, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000998.

Anonymous:2005:CPCa

[Ano05k] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 165(1):v–viii, January 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504005387.

Anonymous:2005:CPCb

[Ano05l] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 165(2):iii–vi, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504005508.

Anonymous:2005:CPCc

[Ano05m] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 165(3):iii–vi, February 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000068.

Anonymous:2005:CPCd

[Ano05n] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 166(1):v–viii, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000226.

Anonymous:2005:CPCe

[Ano05o] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 166(2):iii–vi, March 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000354.

Anonymous:2005:CPCf

[Ano05p] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 166(3):iii–vi, March 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000470.

Anonymous:2005:CPCg

[Ano05q] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 167(1):v–viii, April 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000901.

Anonymous:2005:CPCh

[Ano05r] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 167(2):iii–vi, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002092.

Anonymous:2005:CPCi

[Ano05s] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 167(3):iii–vi, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500216X.

Anonymous:2005:CPCj

[Ano05t] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 168(1):v–viii, May 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002432.

Anonymous:2005:CPCk

[Ano05u] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 169(1–3):v–viii, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003280.

Anonymous:2005:CPCl

[Ano05v] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 170(1):v–viii, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003462.

Anonymous:2005:CPCm

[Ano05w] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 170(2):iii–vi, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003656.

Anonymous:2005:CPCn

[Ano05x] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 170(3):iii–vi, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003826.

Anonymous:2005:CPCo

[Ano05y] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 171(1):v– viii, September 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550500408X.

Anonymous:2005:CPCp

[Ano05z] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 171(2):iii– vi, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550500439X.

Anonymous:2005:CPCq

[Ano05-27] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 171(3):iii–vi, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004546.

Anonymous:2005:CPCr

[Ano05-28] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 172(1):v–viii, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004820.

Anonymous:2005:CPCs

[Ano05-29] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 172(2):iii–vi, November 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004959.

Anonymous:2005:CPCt

[Ano05-30] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 173(1–2):v–viii, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505005400.

Anonymous:2005:C

[Ano05-31] Anonymous. Contents. Computer Physics Communications, 169(1–3):xv–xxi, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550500336X.

Anonymous:2005:CVa

[Ano05-32] Anonymous. Contents to volume 165. Computer Physics Communications, 165(3):271–272, February 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500007X.

Anonymous:2005:CVb

[Ano05-33] Anonymous. Contents to volume 166. Computer Physics Communications, 166(3):241–242, March 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000482.

Anonymous:2005:CVc

[Ano05-34] Anonymous. Contents to volume 167. Computer Physics Communications, 167(3):253–254, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002171.

Anonymous:2005:CVd

[Ano05-35] Anonymous. Contents to volume 168. Computer Physics Communications, 168(3):233–234, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003000.

Anonymous:2005:CVe

[Ano05-36] Anonymous. Contents to volume 170. Computer Physics Communications, 170(3):299–300, August 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003838.

Anonymous:2005:CVf

[Ano05-37] Anonymous. Contents to volume 171. Computer Physics Communications, 171(3):219–220, October 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004558.

Anonymous:2005:CVg

[Ano05-38] Anonymous. Contents to volume 172. Computer Physics Communications, 172(3):215–216, November 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005230.

Anonymous:2005:CVh

[Ano05-39] Anonymous. Contents to volume 173. Computer Physics Communications, 173(3):193–194, December 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005576.

Anonymous:2005:EBa

[Ano05-40] Anonymous. Editorial Board. Computer Physics Communications, 171(3):??, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505004510.

Anonymous:2005:EBb

[Ano05-41] Anonymous. Editorial Board. Computer Physics Communications, 172(1):??, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505004789.

Anonymous:2005:EBc

[Ano05-42] Anonymous. Editorial Board. Computer Physics Communications, 172(2):??, November 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505004935.

Anonymous:2005:EBd

[Ano05-43] Anonymous. Editorial Board. Computer Physics Communications, 172(3):??, November 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005205.

Anonymous:2005:EBe

[Ano05-44] Anonymous. Editorial Board. Computer Physics Communications, 173(1–2):??, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005369.

Anonymous:2005:EBf

[Ano05-45] Anonymous. Editorial Board. Computer Physics Communications, 173(3):??, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005540.

Anonymous:2005:PIVa

[Ano05-46] Anonymous. Program index to volume 165. Computer Physics Communications, 165(3):275, February 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000093.

Anonymous:2005:PIVb

[Ano05-47] Anonymous. Program index to volume 166. Computer Physics Communications, 166(3):245–246, March 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000500.

Anonymous:2005:PIVc

[Ano05-48] Anonymous. Program index to volume 167. Computer Physics Communications, 167(3):258, May 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002195.

Anonymous:2005:PIVd

[Ano05-49] Anonymous. Program index to volume 168. Computer Physics Communications, 168(3):237, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003024.

Anonymous:2005:PIVe

[Ano05-50] Anonymous. Program index to volume 170. Computer Physics Communications, 170(3):303–304, August 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003851.

Anonymous:2005:PIVf

[Ano05-51] Anonymous. Program index to volume 171. Computer Physics Communications, 171(3):223–224, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004571.

Anonymous:2005:PIVg

[Ano05-52] Anonymous. Program index to volume 172. Computer Physics Communications, 172(3):219, November 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005254.

Anonymous:2005:PIVh

[Ano05-53] Anonymous. Program index to volume 173. Computer Physics Communications, 173(3):197–198, December 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500559X.

Anonymous:2006:AIV

[Ano06a] Anonymous. Author index to volume 174. Computer Physics Communications, 174(12):1011–1015, June 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001780.

Anonymous:2006:CPCa

[Ano06b] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 174(1):v–vii, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550500603X.

Anonymous:2006:CPCb

[Ano06c] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 175(1):v–vii, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002025.

Anonymous:2006:CV

[Ano06d] Anonymous. Contents to volume 174. Computer Physics Communications, 174(12):1005–1010, June 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001779.

Anonymous:2006:EBa

[Ano06e] Anonymous. Editorial Board. Computer Physics Communications, 174(1):??, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505005990.

Anonymous:2006:EBb

[Ano06f] Anonymous. Editorial Board. Computer Physics Communications, 174(2):??, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505006132.

Anonymous:2006:EBc

[Ano06g] Anonymous. Editorial Board. Computer Physics Communications, 174(3):??, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000208.

Anonymous:2006:EBd

[Ano06h] Anonymous. Editorial Board. Computer Physics Communications, 174(4):??, February 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000427.

Anonymous:2006:EBe

[Ano06i] Anonymous. Editorial Board. Computer Physics Communications, 174(5):??, March 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000580.

Anonymous:2006:EBf

[Ano06j] Anonymous. Editorial Board. Computer Physics Communications, 174(6):??, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000762.

Anonymous:2006:EBg

[Ano06k] Anonymous. Editorial Board. Computer Physics Communications, 174(7):??, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000865.

Anonymous:2006:EBh

[Ano06l] Anonymous. Editorial Board. Computer Physics Communications, 174(8):??, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001160.

Anonymous:2006:EBi

[Ano06m] Anonymous. Editorial Board. Computer Physics Communications, 174(9):??, May 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001330.

Anonymous:2006:EBj

[Ano06n] Anonymous. Editorial Board. Computer Physics Communications, 174(10):??, May 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001470.

Anonymous:2006:EBk

[Ano06o] Anonymous. Editorial Board. Computer Physics Communications, 174(11):??, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001627.

Anonymous:2006:EBl

[Ano06p] Anonymous. Editorial Board. Computer Physics Communications, 174(12):??, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001743.

Anonymous:2006:EBm

[Ano06q] Anonymous. Editorial Board. Computer Physics Communications, 175(1):??, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001986.

Anonymous:2006:EBn

[Ano06r] Anonymous. Editorial Board. Computer Physics Communications, 175(2):??, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002190.

Anonymous:2006:EBo

[Ano06s] Anonymous. Editorial Board. Computer Physics Communications, 175(3):??, August 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002347.

Anonymous:2006:EBp

[Ano06t] Anonymous. Editorial Board. Computer Physics Communications, 175(4):??, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002487.

Anonymous:2006:EBq

[Ano06u] Anonymous. Editorial Board. Computer Physics Communications, 175(5):??, September 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002785.

Anonymous:2006:EBr

[Ano06v] Anonymous. Editorial Board. Computer Physics Communications, 175(6):??, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550600302X.

Anonymous:2006:EBs

[Ano06w] Anonymous. Editorial Board. Computer Physics Communications, 175(7):??, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506003262.

Anonymous:2006:EBt

[Ano06x] Anonymous. Editorial Board. Computer Physics Communications, 175(8):??, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506003420.

Anonymous:2006:EBu

[Ano06y] Anonymous. Editorial Board. Computer Physics Communications, 175(9):??, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506003547.

Anonymous:2006:EBv

[Ano06z] Anonymous. Editorial Board. Computer Physics Communications, 175(10):??, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003742.

Anonymous:2006:EBw

[Ano06-27] Anonymous. Editorial Board. Computer Physics Communications, 175(11–12):??, December 1–15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003845.

Anonymous:2006:PIV

[Ano06-28] Anonymous. Program index to volume 174. Computer Physics Communications, 174(12):1016–1019, June 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001792.

Anonymous:2006:PN

[Ano06-29] Anonymous. Publisher’s note. Computer Physics Communications, 174(1):ix, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505004698.

Anonymous:2007:CC

[Ano07a] Anonymous. CCP2006 committees. Computer Physics Communications, 177(1–2):xiii–xiv, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001543.

Anonymous:2007:CPCa

[Ano07b] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 176(1):v–vii, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506004085.

Anonymous:2007:CPCb

[Ano07c] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 176(6):iii–v, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001506.

Anonymous:2007:CPCc

[Ano07d] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 176(7):iii–v, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001725.

Anonymous:2007:CPCd

[Ano07e] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 177(1–2):v–vii, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507002627.

Anonymous:2007:CP

[Ano07f] Anonymous. Conference photo. Computer Physics Communications, 177(1–2):xv, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001555.

Anonymous:2007:C

[Ano07g] Anonymous. Contents. Computer Physics Communications, 177(1–2):xvii–xxiii, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002664.

Anonymous:2007:EBa

[Ano07h] Anonymous. Editorial Board. Computer Physics Communications, 176(1):??, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506004048.

Anonymous:2007:EBb

[Ano07i] Anonymous. Editorial Board. Computer Physics Communications, 176(2):??, January 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550600422X.

Anonymous:2007:EBc

[Ano07j] Anonymous. Editorial Board. Computer Physics Communications, 176(3):??, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000033.

Anonymous:2007:EBd

[Ano07k] Anonymous. Editorial Board. Computer Physics Communications, 176(4):??, February 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000161.

Anonymous:2007:EBe

[Ano07l] Anonymous. Editorial Board. Computer Physics Communications, 176(5):??, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000355.

Anonymous:2007:EBf

[Ano07m] Anonymous. Editorial Board. Computer Physics Communications, 176(6):??, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001476.

Anonymous:2007:EBg

[Ano07n] Anonymous. Editorial Board. Computer Physics Communications, 176(7):??, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001695.

Anonymous:2007:EBh

[Ano07o] Anonymous. Editorial Board. Computer Physics Communications, 176(8):??, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001889.

Anonymous:2007:EBi

[Ano07p] Anonymous. Editorial Board. Computer Physics Communications, 176(9–10):??, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002081.

Anonymous:2007:EBj

[Ano07q] Anonymous. Editorial Board. Computer Physics Communications, 176(11–12):??, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002305.

Anonymous:2007:EBk

[Ano07r] Anonymous. Editorial Board. Computer Physics Communications, 177(1–2):??, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002585.

Anonymous:2007:EBl

[Ano07s] Anonymous. Editorial Board. Computer Physics Communications, 177(3):??, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507003177.

Anonymous:2007:EBm

[Ano07t] Anonymous. Editorial Board. Computer Physics Communications, 177(4):??, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507003293.

Anonymous:2007:EBn

[Ano07u] Anonymous. Editorial Board. Computer Physics Communications, 177(5):??, September 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003384.

Anonymous:2007:EBo

[Ano07v] Anonymous. Editorial Board. Computer Physics Communications, 177(6):??, September 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003554.

Anonymous:2007:EBp

[Ano07w] Anonymous. Editorial Board. Computer Physics Communications, 177(7):??, October 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507003864.

Anonymous:2007:EBq

[Ano07x] Anonymous. Editorial Board. Computer Physics Communications, 177(8):??, October 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507003980.

Anonymous:2007:EBr

[Ano07y] Anonymous. Editorial Board. Computer Physics Communications, 177(9):??, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507004092.

Anonymous:2007:EBs

[Ano07z] Anonymous. Editorial Board. Computer Physics Communications, 177(10):??, November 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507004286.

Anonymous:2007:EBt

[Ano07-27] Anonymous. Editorial Board. Computer Physics Communications, 177(11):??, December 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507004389.

Anonymous:2007:EBu

[Ano07-28] Anonymous. Editorial Board. Computer Physics Communications, 177(12):??, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550700450X.

Anonymous:2007:PA

[Ano07-29] Anonymous. Publisher’s announcement. Computer Physics Communications, 176(4):??, February 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507000136.

Anonymous:2007:PN

[Ano07-30] Anonymous. Publisher’s note. Computer Physics Communications, 177(6):??, September 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003499.

Anonymous:2007:SFH

[Ano07-31] Anonymous. A standard format for les houches event files. Computer Physics Communications, 176(4):300–304, February 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506004164.

Anonymous:2008:CC

[Ano08a] Anonymous. CCP2007 committees. Computer Physics Communications, 179(1–3):xiv, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000659.

Anonymous:2008:CPCa

[Ano08b] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 178(1):v–vii, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700478X.

Anonymous:2008:CPCb

[Ano08c] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 179(1–3):v–vii, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002038.

Anonymous:2008:CP

[Ano08d] Anonymous. Conference photo. Computer Physics Communications, 179(1–3):xiii, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000660.

Anonymous:2008:C

[Ano08e] Anonymous. Contents. Computer Physics Communications, 179(1–3):xv–xvii, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508002051.

Anonymous:2008:EBa

[Ano08f] Anonymous. Editorial Board. Computer Physics Communications, 178(1):??, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507004742.

Anonymous:2008:EBb

[Ano08g] Anonymous. Editorial Board. Computer Physics Communications, 178(2):??, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507004961.

Anonymous:2008:EBc

[Ano08h] Anonymous. Editorial Board. Computer Physics Communications, 178(3):??, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550700505X.

Anonymous:2008:EBd

[Ano08i] Anonymous. Editorial Board. Computer Physics Communications, 178(4):??, February 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000520.

Anonymous:2008:EBe

[Ano08j] Anonymous. Editorial Board. Computer Physics Communications, 178(5):??, March 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000696.

Anonymous:2008:EBf

[Ano08k] Anonymous. Editorial Board. Computer Physics Communications, 178(6):??, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000829.

Anonymous:2008:EBg

[Ano08l] Anonymous. Editorial Board. Computer Physics Communications, 178(7):??, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550800101X.

Anonymous:2008:EBh

[Ano08m] Anonymous. Editorial Board. Computer Physics Communications, 178(8):??, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001173.

Anonymous:2008:EBi

[Ano08n] Anonymous. Editorial Board. Computer Physics Communications, 178(9):??, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001264.

Anonymous:2008:EBj

[Ano08o] Anonymous. Editorial Board. Computer Physics Communications, 178(10):??, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001549.

Anonymous:2008:EBk

[Ano08p] Anonymous. Editorial Board. Computer Physics Communications, 178(11):??, June 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001689.

Anonymous:2008:EBl

[Ano08q] Anonymous. Editorial Board. Computer Physics Communications, 178(12):??, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001835.

Anonymous:2008:EBm

[Ano08r] Anonymous. Editorial Board. Computer Physics Communications, 179(1–3):??, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001999.

Anonymous:2008:EBn

[Ano08s] Anonymous. Editorial Board. Computer Physics Communications, 179(4):??, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508002373.

Anonymous:2008:EBo

[Ano08t] Anonymous. Editorial Board. Computer Physics Communications, 179(5):??, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002580.

Anonymous:2008:EBp

[Ano08u] Anonymous. Editorial Board. Computer Physics Communications, 179(6):??, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002774.

Anonymous:2008:EBq

[Ano08v] Anonymous. Editorial Board. Computer Physics Communications, 179(7):??, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508002919.

Anonymous:2008:EBr

[Ano08w] Anonymous. Editorial Board. Computer Physics Communications, 179(8):??, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003068.

Anonymous:2008:EBs

[Ano08x] Anonymous. Editorial Board. Computer Physics Communications, 179(9):??, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003226.

Anonymous:2008:EBt

[Ano08y] Anonymous. Editorial Board. Computer Physics Communications, 179(10):??, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003378.

Anonymous:2008:EBu

[Ano08z] Anonymous. Editorial Board. Computer Physics Communications, 179(11):??, December 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003470.

Anonymous:2008:EBv

[Ano08-27] Anonymous. Editorial Board. Computer Physics Communications, 179(12):??, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003561.

Anonymous:2009:AFP

[Ano09a] Anonymous. ABINIT: First-principles approach to material and nanosystem properties. Computer Physics Communications, 180(12):2582–2615, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509002276.

Anonymous:2009:CPCc

[Ano09b] Anonymous. Computer Physics Communications. Computer Physics Communications, 180(12):??, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509003130.

Anonymous:2009:CPCa

[Ano09c] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 180(1):v–vii, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004062.

Anonymous:2009:CPCb

[Ano09d] Anonymous. Computer Physics Communications — list of editors. Computer Physics Communications, 180(3):iii–v, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900054X.

Anonymous:2009:C

[Ano09e] Anonymous. Contents. Computer Physics Communications, 180(12):ix–x, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550900335X.

Anonymous:2009:EBa

[Ano09f] Anonymous. Editorial Board. Computer Physics Communications, 180(1):??, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004025.

Anonymous:2009:EBb

[Ano09g] Anonymous. Editorial Board. Computer Physics Communications, 180(2):??, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000101.

Anonymous:2009:EBc

[Ano09h] Anonymous. Editorial Board. Computer Physics Communications, 180(3):??, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000514.

Anonymous:2009:EBd

[Ano09i] Anonymous. Editorial Board. Computer Physics Communications, 180(4):??, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000770.

Anonymous:2009:EBe

[Ano09j] Anonymous. Editorial Board. Computer Physics Communications, 180(5):??, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000903.

Anonymous:2009:EBf

[Ano09k] Anonymous. Editorial Board. Computer Physics Communications, 180(6):??, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001350.

Anonymous:2009:EBg

[Ano09l] Anonymous. Editorial Board. Computer Physics Communications, 180(7):??, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001702.

Anonymous:2009:EBh

[Ano09m] Anonymous. Editorial Board. Computer Physics Communications, 180(8):??, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002161.

Anonymous:2009:EBi

[Ano09n] Anonymous. Editorial Board. Computer Physics Communications, 180(9):??, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002446.

Anonymous:2009:EBj

[Ano09o] Anonymous. Editorial Board. Computer Physics Communications, 180(10):??, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002641.

Anonymous:2009:EBk

[Ano09p] Anonymous. Editorial Board. Computer Physics Communications, 180(11):??, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002896.

Anonymous:2009:EBl

[Ano09q] Anonymous. Editorial Board. Computer Physics Communications, 180(12):??, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509003300.

Anonymous:2009:FC

[Ano09r] Anonymous. Figure captions. Computer Physics Communications, 180(12):iv, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550900321X.

Anonymous:2009:GTC

[Ano09s] Anonymous. Ganga: a tool for computational-task management and easy access to Grid resources. Computer Physics Communications, 180(11):2303–2316, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001970.

Anonymous:2009:RCF

[Ano09t] Anonymous. ROOT — a C++ framework for petabyte data storage, statistical analysis and visualization. Computer Physics Communications, 180(12):2499–2512, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002550.

Anonymous:2009:SHA

[Ano09u] Anonymous. SUSY Les Houches Accord 2. Computer Physics Communications, 180(1):8–25, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002737.

Anonymous:2012:EB

[Ano12] Anonymous. Editorial Board. Computer Physics Communications, 183(5):??, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465512000409.

Aoki:2001:SFP

[Aok01] Takayuki Aoki. 3D simulation for falling papers. Computer Physics Communications, 142(1–3):326–329, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003617.

Alvarez:2001:DEP

[AOT01] A. Alvarez, A. Orfila, and J. Tintore. DARWIN: An evolutionary program for nonlinear modeling of chaotic time series. Computer Physics Communications, 136(3):334–349, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001540.

Abrashkevich:2004:CPN

[AP04] Alexander Abrashkevich and I. V. Puzynin. CANM, a program for numerical solution of a system of nonlinear equations using the continuous analog of Newton’s method. Computer Physics Communications, 156(2):154–170, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004612.

Arya:2005:MMS

[AP05] Gaurav Arya and Athanassios Z. Panagiotopoulos. Molecular modeling of shear-induced alignment of cylindrical micelles. Computer Physics Communications, 169(1–3):262–266, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001578.

Adler:2009:VTM

[AP09] Joan Adler and Polina Pine. Visualization techniques for modelling carbon allotropes. Computer Physics Communications, 180(4):580–582, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004323.

Abrashkevich:2000:APC

[APV00] A. G. Abrashkevich, I. V. Puzynin, and S. I. Vinitsky. ASYMPT: a program for calculating asymptotics of hyperspherical potential curves and adiabatic potentials. Computer Physics Communications, 125(1–3):259–281, March 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004750.

Aray:2002:IAM

[ARV02] Yosslen Aray, Jesus Rodriguez, and David Vega. An implementation of the atoms in molecules theory to the FPLAPW method. Computer Physics Communications, 143(3):199–212, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004283.

Aguilar-Saavedra:2000:CCI

[AS00] J. A. Aguilar-Saavedra. Computation of confidence intervals for Poisson processes. Computer Physics Communications, 130(1–2):190–203, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000357.

Abad:2003:SDW

[AS03] J. Abad and J. Sesma. Successive derivatives of Whittaker functions with respect to the first parameter. Computer Physics Communications, 156(1):13–21, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003692.

Atzeni:2005:FKS

[ASC^{+}05] S. Atzeni, A. Schiavi, F. Califano, F. Cattani, F. Cornolti, D. Del Sarto, T. V. Liseykina, A. Macchi, and F. Pegoraro. Fluid and kinetic simulation of inertial confinement fusion plasmas. Computer Physics Communications, 169(1–3):153–159, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001335.

Ascasibar:2008:FSM

[Asc08] Yago Ascasibar. FiEstAS sampling — a Monte Carlo algorithm for multidimensional numerical integration. Computer Physics Communications, 179(12):881–887, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002683.

Alvarez:2005:TPE

[ASF^{+}05] G. Alvarez, C. Sen, N. Furukawa, Y. Motome, and E. Dagotto. The truncated polynomial expansion Monte Carlo method for fermion systems coupled to classical fields: a model independent implementation. Computer Physics Communications, 168(1):32–45, May 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505000780.

Ali:2006:SAM

[ASH06] M. M. Ali, Roger Smith, and S. Hobday. The structure of atomic and molecular clusters, optimised using classical potentials. Computer Physics Communications, 175(7):451–464, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002323.

Aksenov:2003:ACN

[ASJ^{+}03] Sergej V. Aksenov, Michael A. Savageau, Ulrich D. Jentschura, Jens Becher, Gerhard Soff, and Peter J. Mohr. Application of the combined nonlinear-condensation transformation to problems in statistical analysis and theoretical physics.

Computer Physics Communications, 150(1):1–20, January 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006276.

Aoyagi:2002:GPC

[ASS^{+}02] Takeshi Aoyagi, Fumio Sawa, Tatsuya Shoji, Hiroo Fukunaga, Jun ichi Takimoto, and Masao Doi. A generalpurpose coarse-grained molecular dynamics program. Computer Physics Communications, 145(2):267–279, May 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002710.

Avdelas:2000:EEF

[ASVA00] G. Avdelas, T. E. Simos, and J. Vigo-Aguiar. An embedded exponentially-fitted Runge–Kutta method for the numerical solution of the Schrödinger equation and related periodic initial-value problems. Computer Physics Communications, 131(1–2):52–67, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000801.

Assous:2009:NPA

[AT09] Franck Assous and Felix Tsipis. Numerical paraxial approximation for highly relativistic beams. Computer Physics Communications, 180(7):1086–1097, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508004530.

Ackermann:2001:PRN

[ATB^{+}01] J. Ackermann, U. Tangen, B. Bödekker, J. Breyer, E. Stoll, and J. S. McCaskill. Parallel random number generator for inexpensive configurable hardware cells. Computer Physics Communications, 140(3):293–302, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002429.

Aiba:2009:MIM

[ATF^{+}09] N. Aiba, S. Tokuda, M. Furukawa, P. B. Snyder, and M. S. Chu. MINERVA: Ideal MHD stability code for toroidally rotating tokamak plasmas. Computer Physics Communications, 180(8):1282–1304, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000654.

Aiba:2006:ENE

[ATIO06] Nobuyuki Aiba, Shinji Tokuda, Tomoko Ishizawa, and Masao Okamoto. Extension of the Newcomb equation into the vacuum for the stability analysis of tokamak edge plasmas. Computer Physics Communications, 175(4):269–289, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001950.

Aguado:2001:GFI

[ATP01] Alfredo Aguado, César Tablero, and Miguel Paniagua. Global fit of ab initio potential energy surfaces: II.1. Tetraatomic systems ABCD. Computer Physics Communications, 134(1):97–109, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001818.

Attig:2009:CPP

[Att09] Norbert Attig. Computational physics with petaflops computers. Computer Physics Communications, 180(4):555–558, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004499.

ALPHA:2004:MCE

[AW04] ALPHA Collaboration and Ulli Wolff. Monte Carlo errors with less errors. Computer Physics Communications, 156 (2):143–153, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004673.

See erratum [WA07].

Baglietto:2009:CSC

[BA09] Gabriel Baglietto and Ezequiel V. Albano. Computer simulations of the collective displacement of self-propelled agents. Computer Physics Communications, 180(4):527–531, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004438.

Bouriquet:2004:KDC

[BAB04] Bertrand Bouriquet, Yasuhisa Abe, and David Boilley. KEW-PIE: a dynamical cascade code for decaying exited compound nuclei. Computer Physics Communications, 159(1):1–18, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550400075X.

Bacic:2000:FCS

[Bac00] Zlatko Bacić. Fully coupled six-dimensional calculations of rovibrational eigenstates of floppy four-atom molecules. Computer Physics Communications, 128(1–2):46–54, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005172.

Bacic:2002:RDQ

[Bac02] Zlatko Bacić. Reduced-dimensionality quantum bound state treatment of hydrogen-bonded clusters: torsional vibrational manifold of the water trimer. Computer Physics Communications, 145(1):184–193, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502001534.

Becciani:2001:YRF

[BAD01] U. Becciani and V. Antonuccio-Delogu. Are you ready to FLY in the universe? A multi-platform N-body tree code for parallel supercomputers. Computer Physics Communications, 136(1–2):54–63, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002538.

Becciani:2007:FMH

[BADC07] U. Becciani, V. Antonuccio-Delogu, and M. Comparato. FLY: MPI-2 high resolution code for LSS cosmological simulations. Computer Physics Communications, 176(3):211–217, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506003687.

Baeurle:2003:SPA

[Bae03] S. A. Baeurle. The stationary phase auxiliary field Monte Carlo method: a new strategy for reducing the sign problem of auxiliary field methodologies. Computer Physics Communications, 154(2):111–120, August 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002844.

Baeurle:2004:GCA

[Bae04] S. A. Baeurle. Grand canonical auxiliary field Monte Carlo: a new technique for simulating open systems at high density. Computer Physics Communications, 157(3):201–206, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503005216.

Ball:2001:UCB

[Bal01] R. Ball. Understanding critical behaviour through visualization: a walk around the pitchfork. Computer Physics Communications, 142(1–3):71–75, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003228.

Balandin:2007:VSH

[Bal07] A. L. Balandin. Vector spherical harmonics application to 3-D tomography problem. Computer Physics Communications, 176(7):457–464, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506004310.

Barashenkov:2000:MCS

[Bar00] V. S. Barashenkov. Monte Carlo simulation of ionization and nuclear processes initiated by hadron and ion beams in media. Computer Physics Communications, 126(1–2):28–31, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004178.

Barlow:2002:CCI

[Bar02] Roger Barlow. A calculator for confidence intervals. Computer Physics Communications, 149(2):97–102, December 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200588X.

Barlow:2003:PCC

[Bar03] D. A. Barlow. PHASECALC: a code for the calculation of pseudobinary phase diagrams using no adjustable parameters. Computer Physics Communications, 153(2):227–232, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503001589.

Barbot:2004:DSH

[Bar04] C. Barbot. Decay of super-heavy particles: user guide of the SHdecay program. Computer Physics Communications, 157(1):63–86, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004697.

Bates:2003:RRL

[Bat03] Jason W. Bates. Reconstruction of Ross’ linear-mixing model for the equation of state of deuterium. Computer Physics Communications, 151(2):149–170, March 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006987.

Bolding:2000:MSM

[BB00] Barry Bolding and Kim Baldridge. Multithreaded shared memory parallel implementation of the electronic structure code GAMESS. Computer Physics Communications, 128(1–2):55– 66, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000679.

Butterworth:2003:JWI

[BB03] J. M. Butterworth and S. Butterworth. JetWeb: a WWW interface and database for Monte Carlo tuning and validation. Computer Physics Communications, 153(2):164–178, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550300208X.

Belkhouche:2004:SLS

[BB04a] Fethi Belkhouche and Boumediene Belkhouche. Successive linearizations of second order multidimensional timeinvariant systems. Computer Physics Communications, 162 (2):79–88, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003303.

Beskrovnyy:2004:NAF

[BB04b] Vladislav Beskrovnyy and Pascal Baldi. A new algorithm for fast simulation of optical spectra generated in a nonlinear media. Computer Physics Communications, 160(3):204–212, July 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002073.

Basham:2007:SSS

[BB07] Mark Basham and Roger A. Bennett. Solving surface structures from normal incidence X-ray standing wave data. Computer Physics Communications, 177(5):459–466, September 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700241X.

Bazavov:2009:PPM

[BB09a] Alexei Bazavov and Bernd A. Berg. Program package for multicanonical simulations of U(1) lattice gauge theory. Computer Physics Communications, 180(11):2339–2347, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002148.

Burke:2009:ECP

[BB09b] Phil Burke and Val Burke. Establishment of CPC and the program library. Computer Physics Communications, 180 (12):vii–viii, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509003245.

Busby:2000:HGT

[BBB^{+}00] D. W. Busby, P. G. Burke, V. M. Burke, C. J. Noble, N. S. Scott, and I. T. A. Spence. HBrowse: a GRACE tool for browsing R-matrix H-files. Computer Physics Communications, 131(3):202–224, September 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001223.

Barrand:2001:GSA

[BBB^{+}01] G. Barrand, I. Belyaev, P. Binko, M. Cattaneo, R. Chytracek, G. Corti, M. Frank, G. Gracia, J. Harvey, E. van Herwijnen, P. Maley, P. Mato, S. Probst, and F. Ranjard. GAUDI — A software architecture and framework for building HEP data processing applications. Computer Physics Communications, 140(1–2):45–55, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002545.

Batishchev:2004:MDM

[BBB^{+}04] O. V. Batishchev, A. A. Batishcheva, V. Yu. Bychenkov, C. M. Albukrek, A. V. Brantov, and W. Rozmus. Molecular dynamics model for heating of Coulomb clusters by laser field. Computer Physics Communications, 164(1–3):53–59, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002565.

Ballestrero:2009:PMC

[BBB^{+}09a] Alessandro Ballestrero, Aissa Belhouari, Giuseppe Bevilacqua, Vladimir Kashkan, and Ezio Maina. PHANTOM: a Monte Carlo event generator for six parton final states at high energy colliders. Computer Physics Communications, 180(3):401– 417, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003342.

Bonomi:2009:PPP

[BBB^{+}09b] Massimiliano Bonomi, Davide Branduardi, Giovanni Bussi, Carlo Camilloni, Davide Provasi, Paolo Raiteri, Davide Donadio, Fabrizio Marinelli, Fabio Pietrucci, Ricardo A. Broglia, and Michele Parrinello. PLUMED: a portable plugin for free-energy calculations with molecular dynamics. Computer Physics Communications, 180(10):1961–1972, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900157X.

Beddall:2006:SMV

[BBBD06] A. Beddall, A. Beddall, A. Bingül, and Y. Durmaz. Smoothed multi-variate histogrammed PDEs, and χ^{2} optimisation. Computer Physics Communications, 175(11–12):700–707, December 1–15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002992.

Brantov:2004:NHW

[BBBR04] A. V. Brantov, V. Yu. Bychenkov, O. V. Batishchev, and W. Rozmus. Nonlocal heat wave propagation due to skin layer plasma heating by short laser pulses. Computer Physics Communications, 164(1–3):67–72, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002589.

Baker:2001:DPD

[BBC^{+}01a] Nigel Baker, Alain Bazan, Guy Chevenier, Florida Estrella, Zsolt Kovacs, Jean-Marie Le Goff, Richard Mc-Clatchey, and Peter Martin. Design patterns for descriptiondriven systems in high energy physics. Computer Physics Communications, 140(1–2):1–12, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002508.

Bardin:2001:ZVS

[BBC^{+}01b] D. Bardin, M. Bilenky, P. Christova, M. Jack, L. Kalinovskaya, A. Olchevski, S. Riemann, and T. Riemann. ZFIT-TER v.6.21: a semi-analytical program for fermion pair production in e^{+}e^{−} annihilation. Computer Physics Communications, 133(2–3):229–395, January 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001521.

Bernsdorf:2000:NAP

[BBD00] J. Bernsdorf, G. Brenner, and F. Durst. Numerical analysis of the pressure drop in porous media flow with lattice Boltzmann (BGK) automata. Computer Physics Communications, 129(1–3):247–255, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001119.

Baboulin:2009:ASC

[BBD^{+}09] Marc Baboulin, Alfredo Buttari, Jack Dongarra, Jakub Kurzak, Julie Langou, Julien Langou, Piotr Luszczek, and Stanimire Tomov. Accelerating scientific computations with mixed precision algorithms. Computer Physics Communications, 180(12):2526–2533, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003846.

Bordeianu:2008:SSP

[BBJ^{+}08] C. C. Bordeianu, C. Besliu, Al. Jipa, D. Felea, and I. V. Grossu. Scilab software package for the study of dynamical systems. Computer Physics Communications, 178(10):788–793, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000052.

Berche:2009:URC

[BBJS09] B. Berche, P. Butera, W. Janke, and L. Shchur. Universal ratios of critical amplitudes in the Potts model universality class. Computer Physics Communications, 180(4):493– 496, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004402.

Bogacz:2005:PGH

[BBJW05] L. Bogacz, Z. Burda, W. Janke, and B. Waclaw. A program generating homogeneous random graphs with given weights. Computer Physics Communications, 173(3):162–174, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004285.

Bardin:2007:SS

[BBK^{+}07] D. Bardin, S. Bondarenko, L. Kalinovskaya, G. Nanava, L. Rumyantsev, and W. von Schlippe. SANCnews: Sector ffbb. Computer Physics Communications, 177(9):738–756, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003062.

Ballentine:2008:MDC

[BBOY08] G. E. Ballentine, G. F. Bertsch, N. Onishi, and K. Yabana. Moment distributions of clusters and molecules in the adiabatic rotor model. Computer Physics Communications, 178(1):48–51, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700313X.

Belanger:2002:MPC

[BBPS02] G. Bélanger, F. Boudjema, A. Pukhov, and A. Semenov. micrOMEGAs: a program for calculating the relic density in the MSSM. Computer Physics Communications, 149(2):103–120, December 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005969.

Belanger:2006:MV

[BBPS06] G. Bélanger, F. Boudjema, A. Pukhov, and A. Semenov. micrOMEGAs: Version 1.3. Computer Physics Communications, 174(7):577–604, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506000026.

Belanger:2007:MPCa

[BBPS07a] G. Bélanger, F. Boudjema, A. Pukhov, and A. Semenov. micrOMEGAs 2.0: a program to calculate the relic density of dark matter in a generic model. Computer Physics Communications, 176(5):367–382, March 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506004188.

Belanger:2007:MPCb

[BBPS07b] G. Bélanger, F. Boudjema, A. Pukhov, and A. Semenov. micrOMEGAs 2.0.7: a program to calculate the relic density of dark matter in a generic model. Computer Physics Communications, 177(11):894–895, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003682.

Belanger:2009:DMD

[BBPS09] G. Bélanger, F. Boudjema, A. Pukhov, and A. Semenov. Dark matter direct detection rate in a generic model with micrOMEGAs 2.2. Computer Physics Communications, 180 (5):747–767, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550800386X.

Bhagat:2003:EGB

[BBR03] Vikram Bhagat, Ranjan Bhattacharya, and Dhiranjan Roy. On the evaluation of generalized Bose–Einstein and Fermi– Dirac integrals. Computer Physics Communications, 155(1): 7–20, September 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002947.

Boghosian:2000:PBI

[BC00] Bruce M. Boghosian and Peter V. Coveney. A particulate basis for an immiscible lattice-gas model. Computer Physics Communications, 129(1–3):46–55, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000916.

Bonella:2005:LTD

[BC05] Sara Bonella and David Coker. Linearized, time-dependent, non-adiabatic quantum correlation functions. Computer Physics Communications, 169(1–3):267–273, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500158X.

Barnett:2007:MMP

[BC07] Michael P. Barnett and Joseph F. Capitani. The MATH-SCOUT Mathematica package to postprocess the output of other scientific programs. Computer Physics Communications, 177(12):944–950, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003670.

Becciani:2006:FMP

[BCAD06] U. Becciani, M. Comparato, and V. Antonuccio-Delogu. FLY MPI-2: a parallel tree code for LSS. Computer Physics Communications, 174(7):605–606, April 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000713.

Berche:2002:CEB

[BCBJ02] P. E. Berche, C. Chatelain, B. Berche, and W. Janke. Crossover effects in the bond-diluted Ising model in three dimensions. Computer Physics Communications, 147(1–2):427–430, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502003193.

Barone:2006:RFD

[BCC^{+}06] Luciano M. Barone, Simone Cau, Francesca Cavallari, Silvia Costantini, Ioan Dafinei, Marcella Diemoz, Remo Moro, Giovanni Organtini, Riccardo Paramatti, and Fabio Pellegrino. REDACLE: a flexible database for traceability and workflow management for detector construction. Computer Physics Communications, 174(11):869–875, June 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001056.

Belletti:2008:SSS

[BCC^{+}08] F. Belletti, M. Cotallo, A. Cruz, L. A. Fernández, A. Gordillo, A. Maiorano, F. Mantovani, E. Marinari, V. Martín-Mayor, A. Muñoz-Sudupe, D. Navarro, S. Pérez-Gaviro, J. J. Ruiz-Lorenzo, S. F. Schifano, D. Sciretti, A. Tarancón, R. Tripiccione, and J. L. Velasco. Simulating spin systems on IANUS, an FPGA-based computer. Computer Physics Communications, 178(3):208–216, February 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004183.

Ballestrero:2003:FFS

[BCCM03] A. Ballestrero, R. Chierici, F. Cossutti, and E. Migliore. Four-fermion simulation at LEP2 in DELPHI. Computer Physics Communications, 152(2):175–190, May 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008226.

Butterworth:2003:KCI

[BCCW03] J. M. Butterworth, J. P. Couchman, B. E. Cox, and B. M. Waugh. KtJet: a C++ implementation of the K_{⊥} clustering algorithm. Computer Physics Communications, 153(1):85–96, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503001565.

Ballaminut:2001:WWW

[BCD^{+}01] A. Ballaminut, C. Colonello, M. Dönszelmann, E. van Herwijnen, D. Köper, J. Korhonen, M. Litmaath, J. Perl, A. Theodorou, D. Whiteson, and E. Wolff. WIRED — world wide Web interactive remote event display. Computer Physics Communications, 140(1–2):266–273, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002776.

Besse:2007:NSI

[BCD^{+}07] Christophe Besse, Jean Claudel, Pierre Degond, Fabrice Deluzet, Gérard Gallice, and Christian Tessieras. Numerical simulations of the ionospheric striation model in a non-uniform magnetic field. Computer Physics Communications, 176(2):75–90, January 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003171.

Bertulani:2003:CPN

[BCG03] C. A. Bertulani, C. M. Campbell, and T. Glasmacher. A computer program for nuclear scattering at intermediate and high energies. Computer Physics Communications, 152(3):317–340, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200824X.

Burghardt:2005:UES

[BCH05] Irene Burghardt, Lorenz S. Cederbaum, and James T. Hynes. Ultrafast excited-state charge transfer at a conical intersection: effects of an environment. Computer Physics Communications, 169(1–3):95–98, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001190.

Burgos:2009:UCN

[BCHP09] Enrique Burgos, Horacio Ceva, Laura Hernández, and R. P. J. Perazzo. Understanding and characterizing nestedness in mutualistic bipartite networks. Computer Physics Communications, 180(4):532–535, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003792.

Binosi:2009:JGU

[BCKT09] D. Binosi, J. Collins, C. Kaufhold, and L. Theussl. JaxoDraw: a graphical user interface for drawing Feynman diagrams. Version 2.0 release notes. Computer Physics Communications, 180(9):1709–1715, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000757.

Bellanger:2004:PCT

[BCP04] Véronique Bellanger, Arnaud Courcelle, and Alain Petit. A program to compute the two-step excitation of mesospheric sodium atoms for the polychromatic laser guide star project. Computer Physics Communications, 162(2):143–150, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003534.

Butcher:2003:CSS

[BCV03] J. Butcher, J. Carminati, and K. T. Vu. A comparative study of some computer algebra packages which determine the Lie point symmetries of differential equations. Computer Physics Communications, 155(2):92–114, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003485.

Bagneres:2000:CDF

[BD00] Anne Bagnérés and Sophie Durbiano. 3D computation of the demagnetizing field in a magnetic material of arbitrary shape. Computer Physics Communications, 130(1–2):54–74, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000436.

Bauer:2002:OLI

[BD02] C. Bauer and H. S. Do. One-loop integrals with XLOOPS– GiNaC. Computer Physics Communications, 144(2):154–168, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001583.

Bennaceur:2005:CSS

[BD05] K. Bennaceur and J. Dobaczewski. Coordinate-space solution of the Skyrme–Hartree–Fock–Bogolyubov equations within spherical symmetry. The program HFBRAD (v1.00). Computer Physics Communications, 168(2):96–122, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002304.

Brzuszek:2006:MTS

[BD06] Marcin Brzuszek and Andrzej Daniluk. Multithreaded transactions in scientific computing: New versions of a computer program for kinematical calculations of RHEED intensity oscillations. Computer Physics Communications, 175(10): 678–681, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002979.

Beloy:2008:ADK

[BD08] Kyle Beloy and Andrei Derevianko. Application of the dual-kinetic-balance sets in the relativistic many-body problem of atomic structure. Computer Physics Communications, 179(5):310–319, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508001148.

Bezshyyko:2008:PPD

[BDB^{+}08] O. Bezshyyko, A. Dolinskii, K. Bezshyyko, I. Kadenko, R. Yermolenko, and V. Ziemann. PETAG01: a program for the direct simulation of a pellet target. Computer Physics Communications, 178(2):144–155, January 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003918.

Bierwage:2012:OBR

[BDBV12] A. Bierwage, C. Di Troia, S. Briguglio, and G. Vlad. Orbitbased representation of equilibrium distribution functions for low-noise initialization of kinetic simulations of toroidal plasmas. Computer Physics Communications, 183(5):1107–1123, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465512000288.

Boucaud:2008:DTM

[BDF^{+}08] Ph. Boucaud, P. Dimopoulos, F. Farchioni, R. Frezzotti, V. Gimenez, G. Herdoiza, K. Jansen, V. Lubicz, C. Michael, G. Münster, D. Palao, G. C. Rossi, L. Scorzato, A. Shindler, S. Simula, T. Sudmann, C. Urbach, and U. Wenger. Dynamical twisted mass fermions with light quarks: simulation and analysis details. Computer Physics Communications, 179(10):695–715, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002270.

Belov:2008:LMK

[BDG^{+}08] S. Belov, L. Dudko, E. Galkin, A. Gusev, W. Pokorski, and A. Sherstnev. LCG MCDB — a knowledgebase of Monte–Carlo simulated events. Computer Physics Communications, 178(3):222–229, February 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003931.

Bakk:2002:TPF

[BDH^{+}02] Audun Bakk, Paul G. Dommersnes, Alex Hansen, Johan S. Høye, Kim Sneppen, and Mogens H. Jensen. Thermodynamics of proteins: Fast folders and sharp transitions. Computer Physics Communications, 147(1–2):307–312, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200293X.

Busa:2005:AFP

[BDH^{+}05] Ján Busa, Jozef Dzurina, Edik Hayryan, Shura Hayryan, Chin-Kun Hu, Ján Plávka, Imrich Pokorný, Jaroslav Skrivánek, and Ming-Chya Wu. ARVO: a Fortran package for computing the solvent accessible surface area and the excluded volume of overlapping spheres via analytic equations. Computer Physics Communications, 165(1):59–96, January 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004151.

Binder:2008:SSC

[BDHP08] Kurt Binder, Subir K. Das, Jürgen Horbach, and Sanjay Puri. Simulation of surface-controlled phase separation in slit pores: Diffusive Ginzburg–Landau kinetics versus molecular dynamics. Computer Physics Communications, 179(1–3):1– 7, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000064.

Balevicius:2006:IPP

[BDK^{+}06] R. Balevicius, A. Dziugys, R. Kacianauskas, A. Maknickas, and K. Vislavicius. Investigation of performance of programming approaches and languages used for numerical simulation of granular material by the discrete element method. Computer Physics Communications, 175(6):404–415, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002141.

Boghosian:2000:PSP

[BDL00] B. Boghosian, G. D. Doolen, and D. P. Landau. Preface to selected papers from the international conference on computational physics CCP99. Computer Physics Communications, 127(1):xi–xii, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000576.

Berman:2002:QFA

[BDLT02] G. P. Berman, G. D. Doolen, G. V. López, and V. I. Tsifrinovich. A quantum full adder for a scalable nuclear spin quantum computer. Computer Physics Communications, 146(3):324–330, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550200423X.

Brouwer:2009:OMQ

[BDM09] William J. Brouwer, Michael C. Davis, and Karl T. Mueller. Optimized multiple quantum MAS lineshape simulations in solid state NMR. Computer Physics Communications, 180 (10):1973–1982, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001581.

Breuer:2000:NIM

[BDP00] Heinz-Peter Breuer, Uwe Dorner, and Francesco Petruccione. Numerical integration methods for stochastic wave function equations. Computer Physics Communications, 132 (1–2):30–43, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001351.

Berman:2000:QCD

[BDT00] G. P. Berman, G. D. Doolen, and V. I. Tsifrinovich. Quantum computation as a dynamical process. Computer Physics Communications, 127(1):91–99, May 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000031X.

Bychenkov:2004:CSP

[BDV04] V. Yu. Bychenkov, G. I. Dudnikova, and V. A. Vshivkov. Computer simulation of particle acceleration in thin foil by ultrashort laser pulses. Computer Physics Communications, 164(1–3):459–463, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003194.

Bechtle:2006:FPD

[BDW06] P. Bechtle, K. Desch, and P. Wienemann. Fittino, a program for determining MSSM parameters from collider observables using an iterative method. Computer Physics Communications, 174(1):47–70, January 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004728.

Bowers:2004:MLM

[BDYK04] Kevin J. Bowers, Barbara G. Devolder, Lin Yin, and Thomas J. T. Kwan. A maximum likelihood method for linking particle-in-cell and Monte–Carlo transport simulations. Computer Physics Communications, 164(1–3):311–317, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002978.

Bekavac:2006:CMF

[Bek06] Stefan Bekavac. Calculation of massless Feynman integrals using harmonic sums. Computer Physics Communications, 175(3):180–195, August 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506001123.

Bellagamba:2001:LLG

[Bel01] Lorenzo Bellagamba. LQGENEP: a leptoquark generator for ep scattering. Computer Physics Communications, 141 (1):83–97, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501002958.

Beliakov:2005:CLR

[Bel05]  Gleb Beliakov.  Class library ranlip for multivariate

nonuniform random variate generation. Computer Physics Communications, 170(1):93–108, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002596.

Bondar:2002:NHC

[BEM^{+}02] A. E. Bondar, S. I. Eidelman, A. I. Milstein, T. Pierzchala, N. I. Root, Z. Was, and M. Worek. Novosibirsk hadronic currents for τ → 4π channels of τ decay library TAUOLA. Computer Physics Communications, 146(2):139–153, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200262X.

Berg:2002:GES

[Ber02] Bernd A. Berg. Generalized ensemble simulations for complex systems. Computer Physics Communications, 147(1–2):52–57, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002035.

Berg:2003:MSS

[Ber03a] Bernd A. Berg. Multicanonical simulations step by step. Computer Physics Communications, 153(3):397–406, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002455.

Bertulani:2003:RPM

[Ber03b] C. A. Bertulani. RADCAP: a potential model tool for direct capture reactions. Computer Physics Communications, 156(1):123–141, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503004417.

Beswick:2002:QHT

[Bes02] J. A. Beswick. Quantum hopping techniques applied to vibrational energy transfer. Computer Physics Communications, 145(1):64–77, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502001479.

Balakin:2004:NSE

[BF04] A. A. Balakin and G. M. Fraiman. Numerical simulations of electron-ion collisions in UHI plasmas. Computer Physics Communications, 164(1–3):46–52, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002553.

Bordeianu:2008:CAS

[BFB^{+}08] C. C. Bordeianu, D. Felea, C. Besliu, A. Jipa, and I. V. Grossu. Chaos analysis of a semi-classical nuclear billiard model. Computer Physics Communications, 179(1–3):199– 201, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000416.

Bordeianu:2009:NVS

[BFB^{+}09] C. C. Bordeianu, D. Felea, C. Besliu, Al. Jipa, and I. V. Grossu. A new version of Scilab software package for the study of dynamical systems. Computer Physics Communications, 180(11):2398–2399, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509002264.

Bonche:2005:SSH

[BFH05] P. Bonche, H. Flocard, and P. H. Heenen. Solution of the Skyrme HF + BCS equation on a 3D mesh. Computer Physics Communications, 171(1):49–62, September 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002821.

Brandt:2000:NEM

[BFI^{+}00] S. Brandt, J. A. Font, J. Ma Ibáñez, J. Massó, and E. Seidel. Numerical evolution of matter in dynamical axisymmetric black hole spacetimes: I. Methods and tests. Computer Physics Communications, 124(2–3):169–196, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004476.

Breitung:2001:FBS

[BFL^{+}01] M. Breitung, J. Fromm, T. Levshina, I. Mandrichenko, and M. Schweitzer. Farm batch system and Fermi inter-process communication and synchronization toolkit. Computer Physics Communications, 140(1–2):253–265, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002788.

Blanco:2004:GII

[BFL04] M. A. Blanco, E. Francisco, and V. Luaña. GIBBS: isothermal-isobaric thermodynamics of solids from energy curves using a quasi-harmonic Debye model. Computer Physics Communications, 158(1):57–72, March 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503005472.

Bloch:2007:IMC

[BFLW07] J. Bloch, A. Frommer, B. Lang, and T. Wettig. An iterative method to compute the sign function of a non-Hermitian matrix and its application to the overlap Dirac operator at nonzero chemical potential. Computer Physics Communications, 177(12):933–943, December 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003633.

Bauerdick:2001:EIS

[BFMH^{+}01] L. A. T. Bauerdick, Adrian Fox-Murphy, Tobias Haas, Stefan Stonjek, and Enrico Tassi. Event indexing systems for efficient selection and analysis of HERA data. Computer Physics Communications, 137(2):236–246, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100162X.

Brunetti:2001:SXR

[BG01] A. Brunetti and B. Golosio. Software for X-ray fluorescence and scattering tomographic reconstruction. Computer Physics Communications, 141(3):412–425, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004192.

Bertulani:2006:MGM

[BG06] C. A. Bertulani and A. Gade. MOMDIS: a Glauber model computer code for knockout reactions. Computer Physics Communications, 175(5):372–380, September 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001937.

Booth:2009:SA

[BG09] T. E. Booth and J. E. Gubernatis. Sewing algorithm. Computer Physics Communications, 180(4):509–511, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003871.

Baldwin:2004:SCH

[BGH04] D. Baldwin, Ü. Göktas, and W. Hereman. Symbolic computation of hyperbolic tangent solutions for nonlinear differential-difference equations. Computer Physics Communications, 162(3):203–217, October 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504003728.

Binoth:2009:GNP

[BGH^{+}09a] T. Binoth, J.-Ph. Guillet, G. Heinrich, E. Pilon, and T. Reiter. golem95: a numerical program to calculate oneloop tensor integrals with up to six external legs. Computer Physics Communications, 180(11):2317–2330, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002057.

Blum:2009:IMS

[BGH^{+}09b] Volker Blum, Ralf Gehrke, Felix Hanke, Paula Havu, Ville Havu, Xinguo Ren, Karsten Reuter, and Matthias Scheffler. Ab initio molecular simulations with numeric atom-centered orbitals. Computer Physics Communications, 180(11):2175–2196, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002033.

Bao:2007:CRD

[BGJ^{+}07] Weizhu Bao, Yunyi Ge, Dieter Jaksch, Peter A. Markowich, and Rada M. Weishäupl. Convergence rate of dimension reduction in Bose–Einstein condensates. Computer Physics Communications, 177(11):832–850, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003256.

Buckley-Geer:2001:OCR

[BGLLW01] L. Buckley-Geer, S. Lammel, M. Leininger, and T. Watts. Overview of the CDF Run II data handling system. Computer Physics Communications, 140(1–2):198–208, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002715.

Brunetti:2004:SLC

[BGS^{+}04] M. Brunetti, V. Grandgirard, O. Sauter, J. Vaclavik, and L. Villard. A semi-Lagrangian code for nonlinear global simulations of electrostatic drift-kinetic ITG modes. Computer Physics Communications, 163(1):1–21, October 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004126.

Blum:2001:FLI

[BH01] V. Blum and K. Heinz. Fast LEED intensity calculations for surface crystallography using tensor LEED. Computer Physics Communications, 134(3):392–425, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002095.

Blaudeck:2003:GSC

[BH03] P. Blaudeck and K. H. Hoffmann. Ground states for condensed amorphous systems: Optimizing annealing schemes. Computer Physics Communications, 150(3):293–299, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006884.

Belli:2005:NIA

[BH05] E. A. Belli and G. W. Hammett. A numerical instability in an ADI algorithm for gyrokinetics. Computer Physics Communications, 172(2):119–132, November 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500398X.

Burch:2007:DDI

[BH07] Tommy Burch and Christian Hagen. Domain decomposition improvement of quark propagator estimation. Computer Physics Communications, 176(2):137–145, January 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506003407.

Berg:2008:DPD

[BH08] Bernd A. Berg and Robert C. Harris. From data to probability densities without histograms. Computer Physics Communications, 179(6):443–448, September 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001458.

Han:2007:ASM

[bHhL07] Wen biao Han and Xin hao Liao. The adjustment-stabilization method for constrained systems. Computer Physics Communications, 177(6):500–505, September 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002494.

Blote:2002:CMC

[BHL02] H. W. J. Blöte, J. R. Heringa, and E. Luijten. Cluster Monte Carlo: Extending the range. Computer Physics Communications, 147(1–2):58–63, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002047.

Binder:2007:MCS

[BHM^{+}07] Kurt Binder, Jürgen Horbach, Andrey Milchev, Marcus Müller, and Richard Vink. Monte Carlo simulations of phase transitions of systems in nanoscopic confinement. Computer Physics Communications, 177(1–2):140–145, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000938.

Bunn:2001:GPG

[BHNW01] Julian J. Bunn, Koen Holtman, Harvey B. Newman, and Richard P. Wilkinson. The GIOD Project — Globally Interconnected Object Databases. Computer Physics Communications, 140(1–2):162–171, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002673.

Binder:2002:SPT

[Bin02] Kurt Binder. Simulations of phase transitions in macromolecular systems. Computer Physics Communications, 147 (1–2):22–33, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502001996.

Ballone:2002:DFM

[BJ02] P. Ballone and R. O. Jones. Density functional/Monte Carlo study of ring-opening polymerization. Computer Physics Communications, 147(1–2):325–330, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002965.

Breslau:2003:PAG

[BJ03] J. A. Breslau and S. C. Jardin. A parallel algorithm for global magnetic reconnection studies. Computer Physics Communications, 151(1):8–24, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502006938.

Bachmann:2005:CTH

[BJ05a] Michael Bachmann and Wolfhard Janke. Conformational transitions of heteropolymers. Computer Physics Communications, 169(1–3):111–113, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001232.

Bligaard:2005:OHT

[BJ05b] Thomas Bligaard and Hannes Jónsson. Optimization of hyperplanar transition states: Application to 2D test problems. Computer Physics Communications, 169(1–3):284–288, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550500161X.

Berg:2008:MCA

[BJ08] Bernd A. Berg and Wolfhard Janke. Multibondic cluster algorithm for finite-size scaling studies of critical phenomena. Computer Physics Communications, 179(1–3):21–25, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000428.

Burke:2000:CIP

[BJS00] P. G. Burke, C. Jackson, and N. S. Scott. CPC International Program Library: Description of library services January 2000. Computer Physics Communications, 124(1):xvi–xxvi, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599900012.

Bogdanovich:2001:NMP

[BK01] P. Bogdanovich and R. Karpuskienė. Numerical methods of the preliminary evaluation of the role of admixed configurations in atomic calculations. Computer Physics Communications, 134(3):321–334, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002149.

Bazhan:2005:NFF

[BK05a] W. Bazhan and K. Kolwas. Near-field flat-plane images of spherical nanoparticles. Computer Physics Communications, 165(3):191–198, February 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504004618.

Boonekamp:2005:DMC

[BK05b] M. Boonekamp and T. Kúcs. DPEMC: a Monte Carlo for double diffraction. Computer Physics Communications, 167(3):217–228, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505000767.

Brutovsky:2005:LPF

[BK05c] Branislav Brutovsky and Gerald R. Kneller. Linear prediction of force time series to accelerate molecular dynamics simulations. Computer Physics Communications, 169(1–3):339–342, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001736.

Bauer:2006:QSS

[BK06a] Dieter Bauer and Peter Koval. Qprop: a Schrödingersolver for intense laser-atom interaction. Computer Physics Communications, 174(5):396–421, March 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005825.

Bockmann:2006:IRM

[BK06b] Christine Böckmann and Andreas Kirsche. Iterative regularization method for lidar remote sensing. Computer Physics Communications, 174(8):607–615, April 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000117.

Brangian:2002:EBM

[BKB02a] C. Brangian, W. Kob, and K. Binder. Ergodicity breaking in a mean field Potts glass: a Monte Carlo investigation. Computer Physics Communications, 147(1–2):154–157, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002357.

Brangian:2002:SDG

[BKB02b] C. Brangian, W. Kob, and K. Binder. Static and dynamic glass transitions in the 10-state Potts glass: What can Monte Carlo simulations contribute? Computer Physics Communications, 146(1):9–15, June 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004290.

Blumlein:2009:DCF

[BKKS09] Johannes Blümlein, Manuel Kauers, Sebastian Klein, and Carsten Schneider. Determining the closed forms of the O(α_{s}^{3}) anomalous dimensions and Wilson coefficients from Mellin moments by means of computer algebra. Computer Physics Communications, 180(11):2143–2165, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002008.

Bogdanovich:2002:SPC

[BKM02] P. Bogdanovich, R. Karpuskienė, and A. Momkauskaitė. Some problems of calculation of energy spectra of complex atomic configurations. Computer Physics Communications, 143(2):174–180, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004465.

Bogdanovich:2005:PGS

[BKM05] P. Bogdanovich, R. Karpuskiene, and A. Momkauskaite. A program of generation and selection of configurations for the configuration interaction method in atomic calculations SELECT-CONF. Computer Physics Communications, 172(2):133–143, November 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003954.

Binder:2000:MCT

[BL00] K. Binder and E. Luijten. Monte Carlo tests of theoretical predictions for critical phenomena: still a problem? Computer Physics Communications, 127(1):126–130, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000151.

Blaak:2005:USN

[BL05] Ronald Blaak and Hartmut Löwen. Umbrella sampling in non-equilibrium computer simulations. Computer Physics Communications, 169(1–3):64–68, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001141.

Blankenbecler:2000:TFA

[Bla00] Richard Blankenbecler. Track fitting and antennas. Computer Physics Communications, 124(1):28–39, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004300.

Becciani:2005:PCA

[BLCR05] U. Becciani, A. F. Lanza, A. Costa, and M. Rodonò. A parallel code for the analysis of light curves of close binary stars with surface inhomogeneities through the astrocompgrid portal. Computer Physics Communications, 170(1):11–30, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002341.

Basney:2001:UWD

[BLM01] Jim Basney, Miron Livny, and Paolo Mazzanti. Utilizing widely distributed computational resources efficiently with execution domains. Computer Physics Communications, 140(1–2):246–252, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002764.

Li:2002:RMP

[bLpL02] Zhi bin Li and Yin ping Liu. RATH: a Maple package for finding travelling solitary wave solutions to nonlinear evolution equations. Computer Physics Communications, 148(2):256–266, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005593.

Bertini:2001:PVP

[BLS01] Marc Bertini, Leif Lönnblad, and Torbjörn Sjöstrand. Pythia version 7-0.0 — a proof-of-concept version. Computer Physics Communications, 134(3):365–391, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000206X.

Bailey:2009:RCG

[BLS09a] A. G. Bailey, C. P. Lowe, and A. P. Sutton. REVLD: a coarse-grained model for polymers. Computer Physics Communications, 180(4):594–599, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550900040X.

Becker:2009:SFE

[BLS09b] M. M. Becker, D. Loffhagen, and W. Schmidt. A stabilized finite element method for modeling of gas discharges. Computer Physics Communications, 180(8):1230–1241, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000447.

Blumlein:2000:ACM

[Blü00] Johannes Blümlein. Analytic continuation of Mellin transforms up to two-loop order. Computer Physics Communications, 133(1):76–104, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001569.

Blumlein:2004:ARB

[Blü04] Johannes Blümlein. Algebraic relations between harmonic sums and associated quantities. Computer Physics Communications, 159(1):19–54, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503005411.

Blumlein:2009:SRH

[Blü09] Johannes Blümlein. Structural relations of harmonic sums and Mellin transforms up to weight w = 5. Computer Physics Communications, 180(11):2218–2249, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002124.

Blackman:2001:GCS

[BM01] J. A. Blackman and P. A. Mulheran. Growth of clusters on surfaces: Monte Carlo simulations and scaling properties. Computer Physics Communications, 137(1):195–205, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001795.

Banavar:2002:SL

[BM02a] Jayanth R. Banavar and Amos Maritan. The science of life. Computer Physics Communications, 146(1):129–130, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004460.

Brodatzki:2002:SSG

[BM02b] Uwe Brodatzki and Klaus Mecke. Simulating stochastic geometries: morphology of overlapping grains. Computer Physics Communications, 147(1–2):218–221, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002497.

Bogdanovich:2004:PGC

[BM04] P. Bogdanovich and A. Momkauskaitė. A program for generating configuration state lists in many-electron atoms. Computer Physics Communications, 157(3):217–225, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503005198.

Blakeley:2006:EEG

[BM06] N. D. Blakeley and R. P. Millane. Exact enumeration of the ground states of a triangular Ising antiferromagnet. Computer Physics Communications, 174(3):198–201, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005758.

Bustamante:2005:CGN

[BMC05] M. G. Bustamante, J. E. Miraglia, and F. D. Colavecchia. Computation of a generalized Nordsieck integral. Computer Physics Communications, 171(1):40–48, September 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002857.

Bowler:2001:PSM

[BMG01] D. R. Bowler, T. Miyazaki, and M. J. Gillan. Parallel sparse matrix multiplication for linear scaling electronic structure calculations. Computer Physics Communications, 137(2):255–273, June 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001643.

Binder:2005:MCS

[BMML05] Kurt Binder, M. Müller, A. Milchev, and D. P. Landau. Monte Carlo simulations of Ising models and polymer blends in double wedge geometry: Evidence for novel types of critical phenomena. Computer Physics Communications, 169(1–3):226–229, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001499.

Bernaschi:2009:MPM

[BMS^{+}09] M. Bernaschi, S. Melchionna, S. Succi, M. Fyta, E. Kaxiras, and J. K. Sircar. MUPHY: a parallel MUlti PHYsics/ scale code for high performance bio-fluidic simulations. Computer Physics Communications, 180(9):1495–1502, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001118.

Blackwell:2001:ART

[BMSG01] B. D. Blackwell, B. F. McMillan, A. C. Searle, and H. J. Gardner. Algorithms for real time magnetic field tracing and optimization. Computer Physics Communications, 142(1–3):243–247, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003265.

Bonvin:2000:GBM

[BMvG00] Alexandre M. J. J. Bonvin, Alan E. Mark, and Wilfred F. van Gunsteren. The GROMOS96 benchmarks for molecular simulation. Computer Physics Communications, 128(3):550– 557, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005408.

Bollhofer:2007:JSC

[BN07]  Matthias Bollhöfer and Yvan Notay.  JADAMILU: a

software code for computing selected eigenvalues of large sparse symmetric matrices. Computer Physics Communications, 177(12):951–964, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003700.

Burke:2009:FVF

[BNFM^{+}09] V. M. Burke, C. J. Noble, V. Faro-Maza, A. Maniopoulou, and N. S. Scott. FARM 2DRMP: a version of FARM for use with 2DRMP. Computer Physics Communications, 180 (12):2450–2451, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002379.

Bernardo:2001:NCH

[BNO^{+}01] L. Bernardo, H. Nordberg, D. Olson, A. Shoshani, A. Sim, A. Vaniachine, D. Zimmerman, B. Gibbard, R. Porter, T. Wenaus, and D. Malon. New capabilities in the HENP Grand Challenge Storage Access System and its application at RHIC.

Computer Physics Communications, 140(1–2):179–188, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002697.

Bazhirov:2007:CLP

[BNS07] Timur T. Bazhirov, Genri E. Norman, and Vladimir V. Stegailov. Cavitation in liquid Pb and Li under negative pressures. Computer Physics Communications, 177(1–2):41, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001403.

Brunev:2002:IAD

[BNSY02] D. V. Brunev, I. G. Neizvestny, N. L. Shwartz, and Z. Sh. Yanovitskaya. Interlayer atomic diffusion as the reason for self-assembled quantum dot formation. Computer Physics Communications, 147(1–2):255–258, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002588.

Bentz:2007:CCA

[BOG^{+}07] Jonathan L. Bentz, Ryan M. Olson, Mark S. Gordon, Michael W. Schmidt, and Ricky A. Kendall. Coupled cluster algorithms for networks of shared memory parallel processors. Computer Physics Communications, 176(9–10):589– 600, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001816.

Bradley:2005:OUP

[BOPC05] J. Bradley, S. F. C. O’Rourke, R. T. Pedlow, and D. S. F. Crothers. Other uses of the program ‘ARGON.f90’. Computer Physics Communications, 172(1):42–44, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505003978.

Borcherds:2002:CSS

[Bor02] P. H. Borcherds. A computational study of some Josephson junction circuits. Computer Physics Communications, 147(1–2):87–90, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002229.

Borcherds:2007:PLC

[Bor07] P. H. Borcherds. Python: a language for computational physics. Computer Physics Communications, 177(1–2):199– 201, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000732.

Bowman:2002:PAT

[Bow02] Gary E. Bowman. Propagator approach for time evolving wave packets in the quantum kicked rotator. Computer Physics Communications, 143(2):181–186, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004477.

Boyle:2009:BAG

[Boy09] Peter A. Boyle. The BAGEL assembler generation library. Computer Physics Communications, 180(12):2739–2748, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002616.

Barka:2008:ISB

[BP08a] André Barka and Clément Picard. Implementation of spectral basis functions in BEM/FEM/GSM Domain Decomposition Methods devoted to scattering and radiation applications. Computer Physics Communications, 178(6):438–448, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004602.

Bussi:2008:STC

[BP08b] Giovanni Bussi and Michele Parrinello. Stochastic thermostats: comparison of local and global schemes. Computer Physics Communications, 179(1–3):26–29, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000106.

Berends:2001:NFF

[BPP01] F. A. Berends, C. G. Papadopoulos, and R. Pittau. NEXTCAL-IBUR — A four-fermion generator for electron-positron collisions. Computer Physics Communications, 136(1–2):148–172, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002472.

Beard:2006:ECA

[BPRW06] B. B. Beard, M. Pepe, S. Riederer, and U.-J. Wiese. Efficient cluster algorithm for CP (N − 1) models. Computer Physics Communications, 175(10):629–634, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002633.

Bacus:2001:ISI

[BR01] Bob Bacus and V. G. J. Rodgers. The image of selfintersecting QCD strings in four dimensions. Computer Physics Communications, 136(1–2):37–53, May 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002526.

Bhar:2009:CST

[BR09] Shyamal Bhar and Soumen Kumar Roy. Computer simulation of two continuous spin models using wang–landautransition-matrix Monte Carlo algorithm. Computer Physics Communications, 180(5):699–707, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003883.

Brandt:2005:MSS

[Bra05] A. Brandt. Multiscale solvers and systematic upscaling in computational physics. Computer Physics Communications, 169(1–3):438–441, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001943.

Broniowski:2009:GGI

[BRB09] Wojciech Broniowski, Maciej Rybczyński, and Piotr Bozek. GLISSANDO: GLauber Initial-State Simulation AND mOre. … Computer Physics Communications, 180(1):69–83, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002713.

Bahri:2004:PGC

[BRD04] C. Bahri, D. J. Rowe, and J. P. Draayer. Programs for generating Clebsch–Gordan coefficients of SU(3) in SU(2) and SO(3) bases. Computer Physics Communications, 159(2):121–143, May 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000839.

Blanco-Rey:2004:FLE

[BRdAHK04a] Maria Blanco-Rey, Pedro de Andres, Georg Held, and David A. King. A FORTRAN-90 low-energy electron diffraction program (LEED90 v1.1). Computer Physics Communications, 161(3):151–165, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002310.

Blanco-Rey:2004:MML

[BRdAHK04b] Maria Blanco-Rey, Pedro de Andres, Georg Held, and David A. King. Molecular t-matrices for Low-Energy Electron Diffraction (TMOL v1.1). Computer Physics Communications, 161(3):166–178, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002322.

Brealey:2001:BRB

[Bre01] Nicholas Brealey. Book review: C++ and Object-Oriented Numeric Computing for Scientists and Engineers. Computer Physics Communications, 137(3):427–428, July 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002430.

Brein:2005:ASP

[Bre05] Oliver Brein. Adaptive scanning — a proposal how to scan theoretical predictions over a multi-dimensional parameter space efficiently. Computer Physics Communications, 170(1):42–48, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002572.

Bret:2007:BPD

[Bre07] A. Bret. Beam-plasma dielectric tensor with Mathematica. Computer Physics Communications, 176(5):362–366, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506004140.

Brooks:2000:QCC

[Bro00] M. Brooks. Quantum computing and communications. Computer Physics Communications, 124(2–3):357–358, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005007.

Brown:2007:SMB

[Bro07] Frank L. H. Brown. Simple models for biomembrane structure and dynamics. Computer Physics Communications, 177 (1–2):172–175, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000896.

Brunetti:2000:FPG

[Bru00a] Antonio Brunetti. A fast and precise genetic algorithm for a non-linear fitting problem. Computer Physics Communications, 124(2–3):204–211, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004543.

Brushlinsky:2000:MMP

[Bru00b] K. V. Brushlinsky. Mathematical modelling in plasmastatics. Computer Physics Communications, 126(1–2):37–40, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002441.

Brunelli:2004:PAS

[Bru04] J. C. Brunelli. PSEUDO: applications of streams and lazy evaluation to integrable models. Computer Physics Communications, 163(1):22–40, October 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400414X.

Bastin:2000:NCO

[BS00a] T. Bastin and E. Solano. Numerical computation of onephoton mazer resonances for arbitrary field modes. Computer Physics Communications, 124(2–3):197–203, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559900452X.

Blaak:2000:LDR

[BS00b] Ronald Blaak and Peter M. A. Sloot. Lattice dependence of reaction-diffusion in lattice Boltzmann modeling. Computer Physics Communications, 129(1–3):256–266, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001120.

Bartlett:2003:BTI

[BS03] Philip L. Bartlett and Andris T. Stelbovics. Born total ionisation cross sections: An algebraic computing program using Maple. Computer Physics Communications, 154(2):159–174, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002868.

Belkov:2004:SAC

[BS04a] A. Bel’kov and S. Shulga. Studies of angular correlations in the decays B_{s}^{0} → J/ψϕ by using the SIMUB generator. Computer Physics Communications, 156(3):221–240, January 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550300465X.

Borisov:2004:AIS

[BS04b] A. G. Borisov and S. V. Shabanov. An application of the interpolating scaling functions to wave packet propagation. Computer Physics Communications, 160(1):1–7, June 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001456.

Badal:2006:PLS

[BS06a] Andreu Badal and Josep Sempau. A package of Linux scripts for the parallelization of Monte Carlo simulations. Computer Physics Communications, 175(6):440–450, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550600230X.

Belkov:2006:BPA

[BS06b] A. Bel’kov and S. Shulga. BtoVVana: the package for analysis of B_{s}^{0} → J/ψϕ and B_{d}^{0} → J/ψK^{∗} decays. Computer Physics Communications, 174(6):464–480, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005928.

Brown:2008:NAG

[BS08] Patrick H. Brown and Peter Schuck. A new adaptive gridsize algorithm for the simulation of sedimentation velocity profiles in analytical ultracentrifugation. Computer Physics Communications, 178(2):105–120, January 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003785.

Barvik:2002:FSM

[BSB02] Ivan Barvík, Josef Stepánek, and Jiří Bok. Fully solvated molecular dynamics simulations of duplexes formed by modified oligonucleotides with xylo/phosphodiesteric and xylo/phosphonate internucleoside linkages and their natural counterpart. Computer Physics Communications, 147 (1–2):158–161, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002369.

Beyer:2005:PDK

[BSDMH05] Tilo Beyer, Gernot Schaller, Andreas Deutsch, and Michael Meyer-Hermann. Parallel dynamic and kinetic regular triangulation in three dimensions. Computer Physics Communications, 172(2):86–108, November 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505003991.

Burdick:2003:PIT

[BSK^{+}03] W. Russell Burdick, Yousef Saad, Leeor Kronik, Igor Vasiliev, Manish Jain, and James R. Chelikowsky. Parallel implementation of time-dependent density functional theory. Computer Physics Communications, 156(1):22–42, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004132.

Bleicher:2004:ISS

[BSO^{+}04] Lucas Bleicher, José Marcos Sasaki, Renata Villela Orloski, Lisandro Pavie Cardoso, Marcelo Assaoka Hayashi, and Jacobus Willibrordus Swart. IonRock: software for solving strain gradients of ion-implanted semiconductors by Xray diffraction measurements and evolutionary programming. Computer Physics Communications, 160(2):158–165, July 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001316.

Brichta:2009:UIP

[BSS09] Jean-Paul Brichta, Aden N. Seaman, and Joseph H. Sanderson. Ultrafast imaging of polyatomic molecules with simplex algorithm. Computer Physics Communications, 180(2):197–200, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003147.

Bekas:2005:CCD

[BSTC05] Constantine Bekas, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Computing charge densities with partially reorthogonalized Lanczos. Computer Physics Communications, 171(3):175–186, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505003103.

Brendel:2002:DTS

[BSvdDW02] L. Brendel, A. Schindler, M. von den Driesch, and D. E. Wolf. Different types of scaling in epitaxial growth. Computer Physics Communications, 147(1–2):111–114, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002278.

Becoulet:2007:WTT

[BSW^{+}07] A. Bécoulet, P. Strand, H. Wilson, M. Romanelli, L.-G. Eriksson, and The contributors to the European Task Force on Integrated Modelling Activity. The way towards thermonuclear fusion simulators. Computer Physics Communications, 177(1–2): 55–59, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001348.

Burrage:2001:SAR

[BT01] Kevin Burrage and Tianhai Tian. Stiffly accurate Runge– Kutta methods for stiff stochastic differential equations. Computer Physics Communications, 142(1–3):186–190, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003241.

Binosi:2004:JGU

[BT04] D. Binosi and L. Theußl. JaxoDraw: a graphical user interface for drawing Feynman diagrams. Computer Physics Communications, 161(1–2):76–86, August 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002115.

Benham:2001:SCF

[BTI01] S. P. Benham, J. M. Thijssen, and J. E. Inglesfield. Selfconsistent finite-difference electronic structure calculations. Computer Physics Communications, 136(1–2):64–76, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002551.

Borowski:2002:HDQ

[BTK^{+}02] S. Borowski, S. Thiel, T. Klüner, H.-J. Freund, R. Tisma, and H. Lederer. High-dimensional quantum dynamics of molecules on surfaces: a massively parallel implementation. Computer Physics Communications, 143(2):162–173, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004453.

Blumberger:2005:IMD

[BTS05] Jochen Blumberger, Yoshitaka Tateyama, and Michiel Sprik. Ab initio molecular dynamics simulation of redox reactions in solution. Computer Physics Communications, 169(1–3):256– 261, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001566.

Bush:2006:DDD

[BTS06] I. J. Bush, I. T. Todorov, and W. Smith. A DAFT DL POLY distributed memory adaptation of the Smoothed Particle Mesh Ewald method. Computer Physics Communications, 175(5):323–329, September 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001901.

Bunge:2001:FED

[Bun01a] Carlos F. Bunge. Fast eigensolver for dense real-symmetric matrices. Computer Physics Communications, 138(1):92–100, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002156.

Bunge:2001:FE

[Bun01b] Carlos F. Bunge. Fast exponentials. Computer Physics Communications, 139(3):243–245, October 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002168.

Burke:2001:RCF

[Bur01] Phil Burke. On the retirement of Charlotte Froese Fischer and Ian Grant. Computer Physics Communications, 138(1):1, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002442.

Bursik:2002:PCN

[Bur02] Jiří Buršík. Phase constitution of Ni-based quaternary alloys studied by Monte Carlo simulation. Computer Physics Communications, 147(1–2):162–165, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002370.

Borukhov:2000:NSI

[BV00] V. T. Borukhov and P. N. Vabishchevich. Numerical solution of the inverse problem of reconstructing a distributed right-hand side of a parabolic equation. Computer Physics Communications, 126(1–2):32–36, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004166.

Berends:2002:GMC

[BvG02] F. A. Berends and R. van Gulik. GaGaRes: a Monte Carlo generator for resonance production in two-photon physics. Computer Physics Communications, 144(1):82–103, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004660.

Bylaska:2002:PIP

[BVKW02] Eric J. Bylaska, Marat Valiev, Ryoichi Kawai, and John H. Weare. Parallel implementation of the projector augmented plane wave method for charged systems. Computer Physics Communications, 143(1):11–28, February 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004131.

Bakhos:2005:CHR

[BVY05] M. Bakhos, A. P. Vincent, and D. A. Yuen. An out-of-core high-resolution FFT algorithm for determining largescale imperfections of surface potentials in crystals. Computer Physics Communications, 168(3):198–208, June 15, 2005.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002560.

Brock:2001:DEL

[BW01] Jerry S. Brock and J. Renae Wiseman. Discrete-expansions for linear interpolation functions. Computer Physics Communications, 142(1–3):206–213, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003344.

Bogner:2008:RSM

[BW08] Christian Bogner and Stefan Weinzierl. Resolution of singularities for multi-loop integrals. Computer Physics Communications, 178(8):596–610, April 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004729.

Barashenkov:2000:OIG

[BZ00] I. V. Barashenkov and E. V. Zemlyanaya. Oscillatory instabilities of gap solitons: a numerical study. Computer Physics Communications, 126(1–2):22–27, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002416.

Cvitas:2007:PSO

[CA07] Marko T. Cvitas and Stuart C. Althorpe. Parallelizable splitoperator propagator for treating coriolis-coupled quantum dynamics. Computer Physics Communications, 177(4):357–361, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002275.

Cho:2009:EEC

[CA09] Myoung Won Cho and Sul-Ah Ahn. Efficient and exact computation of Hubbard and t–J models using quantum diagonalization method. Computer Physics Communications, 180(4):549– 554, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003810.

Cheung:2008:CSL

[CAAM08] David L. Cheung, Lucian Anton, Michael P. Allen, and Andrew J. Masters. Computer simulation of liquids and liquid crystals. Computer Physics Communications, 179(1–3):61– 65, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000167.

Collaboration:2003:EAF

[CAF^{+}03] JLQCD Collaboration, S. Aoki, M. Fukugita, S. Hashimoto, K.-I. Ishikawa, N. Ishizuka, Y. Iwasaki, K. Kanaya, T. Kaneko, Y. Kuramashi, M. Okawa, N. Tsutsui, A. Ukawa, N. Yamada, and T. Yoshié. An exact algorithm for any-flavor lattice QCD with Kogut–Susskind fermion. Computer Physics Communications, 155(3):183–208, November 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003539.

Cai:2009:MNM

[Cai09] Jiaxiang Cai. Multisymplectic numerical method for the regularized long-wave equation. Computer Physics Communications, 180(10):1821–1831, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001556.

Caprio:2005:LLS

[Cap05] M. A. Caprio. LevelScheme: a level scheme drawing and scientific figure preparation system for Mathematica. Computer Physics Communications, 171(2):107–118, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002900.

Carles:2006:ELC

[Car06] A. Grau Carles. EMILIA, the LS counting efficiency for electron-capture and capture-gamma emitters. Computer Physics Communications, 174(1):35–46, January 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004674.

Carles:2007:MMS

[Car07] A. Grau Carles. MICELLE, the micelle size effect on the LS counting efficiency. Computer Physics Communications, 176(4):305–317, February 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550600395X.

Cheng:2000:ESC

[CAW00] Dajun Cheng, Yahia M. M. Antar, and Gang Wang. Electromagnetic scattering by a conducting cylinder coated with a uniaxial chiral layer of varying thickness: generalized modematching method. Computer Physics Communications, 125 (1–3):90–104, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004609.

Cranmer:2005:PGP

[CB05] Kyle Cranmer and R. Sean Bowman. PhysicsGP: a genetic programming approach to event selection. Computer Physics Communications, 167(3):165–176, May 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500069X.

Chatelain:2002:IDS

[CBBJ02] C. Chatelain, P.-E. Berche, B. Berche, and W. Janke. Influence of dilution on the strong first-order phase transition of the 3D 4-state Potts model. Computer Physics Communications, 147(1–2):431–434, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200320X.

Chen:2004:SSM

[CBF^{+}04] J. Chen, J. Breslau, G. Fu, S. Jardin, and W. Park. Symmetric solution in M3D. Computer Physics Communications, 164(1–3):468–471, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003212.

Chacon:2001:BAI

[CBKM01] L. Chacón, D. C. Barnes, D. A. Knoll, and G. H. Miley. A bounce-averaged ion Fokker–Planck code for Penning fusion devices. Computer Physics Communications, 134(2):183–208, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002009.

Costa:2005:AWT

[CBM^{+}05] A. Costa, U. Becciani, P. Miocchi, V. Antonuccio, R. Capuzzo Dolcetta, P. Di Matteo, and V. Rosato. Astrocomp: Web technologies for high performance computing on a network of supercomputers. Computer Physics Communications, 166(1):17–25, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504005041.

Carminati:2008:AMP

[CBMS08] G. Carminati, M. Bazzotti, A. Margiotta, and M. Spurio. Atmospheric MUons from PArametric formulas: a fast GEnerator for neutrino telescopes (MUPAGE). Computer Physics Communications, 179(12):915–923, December 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800266X.

Couturier:2000:PMD

[CC00] Raphaël Couturier and Christophe Chipot. Parallel molecular dynamics using OpenMP on a shared memory machine. Computer Physics Communications, 124(1):49–59, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004324.

Cafarella:2004:DSR

[CC04] Alessandro Cafarella and Claudio Corianò. Direct solution of renormalization group equations of QCD in xspace: NLO implementations at leading twist. Computer Physics Communications, 160(3):213–242, July 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002036.

Chang:2007:ACA

[CC07] S.-L. Chang and C.-S. Chien. Adaptive continuation algorithms for computing energy levels of rotating Bose–Einstein condensates. Computer Physics Communications, 177(9):707–719, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003086.

Clark:2008:PIC

[CC08] Bryan K. Clark and David M. Ceperley. Path integral calculations of vacancies in solid helium. Computer Physics Communications, 179(1–3):82–88, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000209.

Chang:2009:CMP

[CC09] S.-L. Chang and C.-S. Chien. Computing multiple peak solutions for Bose–Einstein condensates in optical lattices. Computer Physics Communications, 180(6):926–947, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004384.

Castillo-Chara:2002:FIP

[CCBL02] J. Castillo-Chará, John W. Bevan, and Robert R. Lucchese. Fitting of an ab initio potential of two linearrigid-rotor dimer and the calculation of rovibrational energy levels by the pseudo-spectral approach. Computer Physics Communications, 145(1):48–63, May 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001467.

Chen:2007:SBN

[CCD07] Yongzhi Chen, Xin Chen, and Yuefan Deng. Simulating botulinum neurotoxin with constant pH molecular dynamics in Generalized Born implicit solvent. Computer Physics Communications, 177(1–2):210–213, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000501.

Cavazzoni:2005:BSF

[CCFG05] Carlo Cavazzoni, Renato Colle, Riccardo Farchioni, and Giuseppe Grosso. Base and salt 3D forms of Emeraldine II polymers by Car–Parrinello molecular dynamics. Computer Physics Communications, 169(1–3):135–138, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001293.

Cafarella:2008:PSN

[CCG08] Alessandro Cafarella, Claudio Corianò, and Marco Guzzi. Precision studies of the NNLO DGLAP evolution at the LHC with candia. Computer Physics Communications, 179(9):665–684, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002257.

Caffo:2009:BFP

[CCGR09] Michele Caffo, Henryk Czyz, Michal Gunia, and Ettore Remiddi. BOKASUN: a fast and precise numerical program to calculate the master integrals of the two-loop sunrise diagrams. Computer Physics Communications, 180(3):427–430, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003706.

Chang:2008:FDC

[CCL08] S.-L. Chang, C.-S. Chien, and Z.-C. Li. A finite difference continuation method for computing energy levels of Bose–Einstein condensates. Computer Physics Communications, 179(4):208– 226, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000957.

Curco:2005:MGR

[CCRA05] David Curcó, Jordi Casanovas, Marc Roca, and Carlos Alemán. A method for generating reliable atomistic models of amorphous polymers based on a random search of energy minima. Computer Physics Communications, 169(1–3):335–338, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001724.

Carvalho:2001:MPC

[CD01a] M. J. Carvalho and S. D’Agostino. A Maple program for calculations with Schur functions. Computer Physics Communications, 141(2):282–295, November 30, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004027.

Collecutt:2001:XEM

[CD01b] G. Collecutt and P. D. Drummond. xmds: eXtensible multi-dimensional simulator. Computer Physics Communications, 142(1–3):219–223, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003095.

Colonna:2004:HAF

[CD04] G. Colonna and A. D’Angola. A hierarchical approach for fast and accurate equilibrium calculation. Computer Physics Communications, 163(3):177–190, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004333.

Capuzzo-Dolcetta:2005:HPC

[CD05] Roberto Capuzzo-Dolcetta. High performance computing for self-gravitating systems. Computer Physics Communications, 169(1–3):365–369, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001785.

Chen:2008:TMS

[CD08] Yongzhi Chen and Yuefan Deng. Task mapping on supercomputers with cellular networks. Computer Physics Communications, 179(7):479–485, October 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001513.

Chen:2009:DAC

[CD09a] Yongzhi Chen and Yuefan Deng. A detailed analysis of communication load balance on BlueGene supercomputer.

Computer Physics Communications, 180(8):1251–1258, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000460.

Christensen:2009:FFR

[CD09b] Neil D. Christensen and Claude Duhr. FeynRules — Feynman rules made easy. Computer Physics Communications, 180(9):1614–1641, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000733.

Collange:2008:LLS

[CDD08] Sylvain Collange, Marc Daumas, and David Defour. Lineby-line spectroscopic simulations on graphics processing units. Computer Physics Communications, 178(2):135–143, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003773.

Chang:2004:BEG

[CDEW04] Chao-Hsi Chang, Chafik Driouichi, Paula Eerola, and Xing-Gang Wu. BCVEGPY: an event generator for hadronic production of the B_{c} meson. Computer Physics Communications, 159(3):192–224, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504000992.

Cavallari:2005:MDS

[CDFF05] Manuela Cavallari, Alessandro Desideri, Mattia Falconi, and Mauro Ferrario. Molecular dynamics simulations of the trp repressor-DNA complex and the AV77 mutant. Computer Physics Communications, 169(1–3):130–134, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001281.

Coveney:2006:CAD

[CDH^{+}06] P. V. Coveney, G. De Fabritiis, M. J. Harvey, S. M. Pickles, and A. R. Porter. Coupled applications on distributed resources.

Computer Physics Communications, 175(6):389–396, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002116.

Cerda-Duran:2007:ASL

[CDQF07] Pablo Cerdá-Durán, Vicent Quilis, and José A. Font. AMR simulations of the low T l|W | bar-mode instability of neutron stars. Computer Physics Communications, 177(3):288–297, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002068.

Cai:2008:ASA

[CEM08] Weiwei Cai, David J. Ewing, and Lin Ma. Application of simulated annealing for multispectral tomography. Computer Physics Communications, 179(4):250–255, August 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000982.

Cox:2002:PHD

[CF02] Brian Cox and Jeff Forshaw. Pomwig: Herwig for diffractive interactions. Computer Physics Communications, 144 (1):104–110, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501004672.

Csikor:2001:PPP

[CFH^{+}01] F. Csikor, Z. Fodor, P. Hegedüs, V. K. Horváth, S. D. Katz, and A. Piróth. The PMS project: Poor man’s supercomputer. Computer Physics Communications, 134(2):139–149, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000182X.

Chang:2009:UGM

[CFJ09] Jee-Gong Chang, Yu-Bin Fang, and Shin-Pon Ju. Using a generalized molecular dynamics force model for random microstructure generation of different aspect-ratios and orientation for use in the optical design of an LED edge-lit backlight. Computer Physics Communications, 180(8):1259–1270, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000629.

Charalambopoulos:2001:SBV

[CFKM01] A. Charalambopoulos, D. I. Fotiadis, D. Kourounis, and C. V. Massalas. On the solution of boundary value problems using spheroidal eigenvectors. Computer Physics Communications, 139(2):153–171, September 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002065.

Calvo:2008:SOS

[CFMR08] M. Calvo, J. M. Franco, J. I. Montijano, and L. Rández. Sixth-order symmetric and symplectic exponentially fitted modified Runge–Kutta methods of Gauss type. Computer Physics Communications, 178(10):732–744, May 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000489.

Colavecchia:2004:FCC

[CG04] F. D. Colavecchia and G. Gasaneo. f1: a code to compute Appell’s F_{1} hypergeometric function. Computer Physics Communications, 157(1):32–38, February 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004909.

Chuluunbaatar:2007:KPC

[CGA^{+}07] O. Chuluunbaatar, A. A. Gusev, A. G. Abrashkevich, A. Amaya-Tapia, M. S. Kaschiev, S. Y. Larsen, and S. I. Vinitsky. KANTBP: a program for computing energy levels, reaction matrix and radial wave functions in the coupledchannel hyperspherical adiabatic approach. Computer Physics Communications, 177(8):649–675, October 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002986.

Chakrabarty:2009:FUI

[CGC^{+}09] Rajan K. Chakrabarty, Mark A. Garro, Shammah Chancellor, Christopher Herald, and Hans Moosmüller. FracMAP: a user-interactive package for performing simulation and orientation-specific morphology analysis of fractal-like solid nano-agglomerates. Computer Physics Communications, 180 (8):1376–1381, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000411.

Cavaglia:2007:CMC

[CGCS07] M. Cavaglià, R. Godang, L. Cremaldi, and D. Summers. Catfish: a Monte Carlo simulator for black holes at the LHC. Computer Physics Communications, 177(6):506–517, September 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002524.

Carsughi:2000:SAS

[CGG00] Flavio Carsughi, Achille Giacometti, and Domenico Gazzillo. Small-angle scattering data analysis for dense polydisperse systems: the FLAC program. Computer Physics Communications, 133(1):66–75, December 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002022.

Chuluunbaatar:2008:PPC

[CGG^{+}08] O. Chuluunbaatar, A. A. Gusev, V. P. Gerdt, V. A. Rostovtsev, S. I. Vinitsky, A. G. Abrashkevich, M. S. Kaschiev, and V. V. Serov. POTHMF: a program for computing potential curves and matrix elements of the coupled adiabatic radial equations for a hydrogen-like atom in a homogeneous magnetic field. Computer Physics Communications, 178(4):301–330, February 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004195. See erra-

tum [CGG^{+}09].

Chuluunbaatar:2009:EPAa

[CGG^{+}09] O. Chuluunbaatar, A. A. Gusev, V. P. Gerdt, V. A. Rostovtsev, S. I. Vinitsky, A. G. Abrashkevich, M. S. Kaschiev, and V. V. Serov. Erratum to: Program AEAA v1 0,“POTHMF: a program for computing potential curves and matrix elements of the coupled adiabatic radial equations for a hydrogen-like atom in a homogeneous magnetic field” [Comput. Phys. Commun. 178 (2008) 301]. Computer Physics Communications, 180(6):1011, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004256. See [CGG^{+}08].

Chen:2007:PIC

[CGIA07] Nan Chen, James A. Glazier, Jesús A. Izaguirre, and Mark S. Alber. A parallel implementation of the cellular Potts model for simulation of cell-based morphogenesis. Computer Physics Communications, 176(11–12):670–681, June 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002044.

Colasanti:2000:MPV

[CGK^{+}00] Alberto Colasanti, Giovanni Guida, Annamaria Kisslinger, Raffaele Liuzzi, Maria Quarto, Patrizia Riccio, Giuseppe Roberti, and Fulvia Villani. Multiple processor version of a Monte Carlo code for photon transport in turbid media. Computer Physics Communications, 132(1–2):84–93, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001387.

Colavecchia:2001:NEA

[CGM01] F. D. Colavecchia, G. Gasaneo, and J. E. Miraglia. Numerical evaluation of Appell’s F_{1} hypergeometric function. Computer Physics Communications, 138(1):29–43, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001862.

Chuluunbaatar:2008:KNV

[CGVA08] O. Chuluunbaatar, A. A. Gusev, S. I. Vinitsky, and A. G. Abrashkevich. KANTBP 2.0: New version of a program for computing energy levels, reaction matrix and radial wave functions in the coupled-channel hyperspherical adiabatic approach. Computer Physics Communications, 179(9):685–693, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002269. See erra-

tum [CGVA09a].

Chuluunbaatar:2009:EPAb

[CGVA09a] O. Chuluunbaatar, A. A. Gusev, S. I. Vinitsky, and A. G. Abrashkevich. Erratum to: Program ADZH v2 0, KANTBP 2.0: New version of a program for computing energy levels, reaction matrix and radial wave functions in the coupled-channel hyperspherical adiabatic approach” [Comput. Phys. Commun. 179 (2008) 685]. Computer Physics Communications, 180(6):1012, June 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004244. See [CGVA08].

Chuluunbaatar:2009:OPC

[CGVA09b] O. Chuluunbaatar, A. A. Gusev, S. I. Vinitsky, and A. G. Abrashkevich. ODPEVP: a program for computing eigenvalues and eigenfunctions and their first derivatives with respect to the parameter of the parametric self-adjoined Sturm–Liouville problem. Computer Physics Communications, 180(8):1358–1375, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001283.

Chuang:2009:PCS

[CH09] Y. C. Chuang and J. Y. Hsu. Particle-in-cell study of microwave confined plasma. Computer Physics Communications, 180(4):656–658, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000307.

Challacombe:2000:GPS

[Cha00] Matt Challacombe. A general parallel sparse-blocked matrix multiply for linear scaling SCF theory. Computer Physics Communications, 128(1–2):93–107, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000746.

Chacon:2004:NSC

[Cha04] L. Chacón. A non-staggered, conservative, ∇ · B⃗ = 0, finite-volume scheme for 3D implicit extended magnetohydrodynamics in curvilinear geometries. Computer Physics Communications, 163(3):143–171, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004369.

Cha:2007:NCP

[Cha07] Moon Hoe Cha. NearFar: a computer program for nearsidefarside decomposition of heavy-ion elastic scattering amplitude. Computer Physics Communications, 176(4):318–325, February 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506003961.

Chekanov:2005:ROO

[Che05] S. Chekanov. RunMC-an object-oriented analysis framework for Monte Carlo simulation of high-energy particle collisions. Computer Physics Communications, 173(3):175–185, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004297.

Cheviakov:2007:GSP

[Che07] Alexei F. Cheviakov. GeM software package for computation of symmetries and conservation laws of differential equations. Computer Physics Communications, 176(1):48–61, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600316X.

Chen:2005:SIS

[CHL05] Ta-Yuan Chen, Li-Ching Hsieh, and Hoong-Chien Lee. Shannon information and self-similarity in whole genomes. Computer Physics Communications, 169(1–3):218–221, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001475.

Chau:2007:MEV

[CHL^{+}07] S. W. Chau, K. L. Hsu, D. L. Lin, J. S. Chen, and C. C. Tzeng. Modeling and experimental validation of a 1.2 MW DC transferred well-type plasma torch. Computer Physics Communications, 177(1–2):114–117, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001269.

Cizek:2000:SLA

[ÇHM00] M. Çižek, J. Horáček, and H.-D. Meyer. Schwinger– Lanczos algorithm for calculation of off-shell T -matrix elements and Wynn’s epsilon algorithm. Computer Physics Communications, 131(1–2):41–51, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000076X.

Cooper:2009:TDA

[CHM^{+}09] W. A. Cooper, S. P. Hirshman, P. Merkel, J. P. Graves, J. Kisslinger, H. F. G. Wobig, Y. Narushima, S. Okamura, and K. Y. Watanabe. Three-dimensional anisotropic pressure free boundary equilibria. Computer Physics Communications, 180(9):1524–1533, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001179.

Cho:2007:CK

[Cho07] Kihyeon Cho. Cyberinfrastructure in korea. Computer Physics Communications, 177(1–2):247–248, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000586.

Cole:2004:ASI

[CHP04] J. H. Cole, L. C. L. Hollenberg, and S. Prawer. An algorithm for simulating the Ising model on a type-II quantum computer. Computer Physics Communications, 161(1–2):18–26, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002085.

Christian:2000:JPI

[Chr00] Wolfgang Christian. Java programming and Internet technologies for undergraduate education. Computer Physics Communications, 127(1):16–22, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000187.

Cheng:2009:BTP

[CHS09] Guang-Hui Cheng, Ting-Zhu Huang, and Shu-Qian Shen. Block triangular preconditioners for the discretized timeharmonic Maxwell equations in mixed form. Computer Physics Communications, 180(2):192–196, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003044.

Calmels:2003:CFT

[CIC^{+}03] L. Calmels, J. E. Inglesfield, S. Crampin, E. Arola, and Th. Rasing. Complex frequency technique for linear and second harmonic optical properties of metallic surfaces. Computer Physics Communications, 151(3):251–264, April 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008500.

Cipolla:2007:IVI

[Cip07] Sam J. Cipolla. An improved version of ISICS: a program for calculating K-, L- and M-shell cross sections from PWBA and ECPSSR theory using a personal computer. Computer Physics Communications, 176(2):157–159, January 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506003195.

Cipolla:2008:IVI

[Cip08] Sam J. Cipolla. An improved version of ISICS: a program for calculating K-, L- and M-shell cross sections from PWBA and ECPSSR theory using a personal computer. Computer Physics Communications, 179(8):616, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002233.

Cipolla:2009:IEV

[Cip09] Sam J. Cipolla. ISICS2008: An expanded version of ISICS for calculating K-, L-, and M-shell cross sections from PWBA and ECPSSR theory. Computer Physics Communications, 180(9):1716–1717, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001337.

Chen:2009:ALR

[CJC09] Kui Fu Chen, Jing Tao Jiang, and Stephen Crowsen. Against the long-range spectral leakage of the cosine window family. Computer Physics Communications, 180(6):904–911, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004311.

Chin:2009:AOD

[CJK09] S. A. Chin, S. Janecek, and E. Krotscheck. An arbitrary order diffusion algorithm for solving Schrödinger equations. Computer Physics Communications, 180(9):1700–1708, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001131.

Chumakov:2006:OJF

[CJT06] S. Chumakov, E. Jankowski, and F. V. Tkachov. Optimal jet finder (v1.0 C++). Computer Physics Communications, 175(7):487–496, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002463.

Chikatamarla:2008:CGI

[CK08] Shyam S. Chikatamarla and Iliya V. Karlin. Complete Galilean invariant lattice Boltzmann models. Computer Physics Communications, 179(1–3):140–143, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000301.

Cundy:2009:NMQ

[CKA^{+}09] N. Cundy, S. Krieg, G. Arnold, A. Frommer, Th. Lippert, and K. Schilling. Numerical methods for the QCD overlap operator IV: Hybrid Monte Carlo. Computer Physics Communications, 180(1):26–54, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002877.

Caldwell:2009:BBA

[CKK09] Allen Caldwell, Daniel Kollár, and Kevin Kröninger. BAT — the Bayesian analysis toolkit. Computer Physics Communications, 180(11):2197–2209, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509002045.

Cundy:2009:TTD

[CKLS09] Nigel Cundy, Stefan Krieg, Thomas Lippert, and Andreas Schäfer. Topological tunnelling with dynamical overlap fermions. Computer Physics Communications, 180(2):201–208, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003160.

Chetyrkin:2000:RMP

[CKS00] K. G. Chetyrkin, J. H. Kühn, and M. Steinhauser. Run-Dec: a Mathematica package for running and decoupling of the strong coupling and quark masses. Computer Physics Communications, 133(1):43–65, December 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001557.

Christlieb:2004:TAS

[CKV04] A. J. Christlieb, R. Krasny, and J. P. Verboncoeur. A treecode algorithm for simulating electron dynamics in a Penning–Malmberg trap. Computer Physics Communications, 164(1–3):306–310, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002966.

Cho:2002:MIS

[CL02] H.-J. Cho and Shih-Ching Lo. Monotone iterative simulation of the multilane traffic dispersion model. Computer Physics Communications, 148(3):291–298, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502005581.

Chen:2003:ANF

[CL03] Han-Taw Chen and Kuo-Chi Liu. Analysis of non-Fickian diffusion problems in a composite medium. Computer Physics Communications, 150(1):31–42, January 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006793.

Chen:2008:CHW

[CL08a] Kui Fu Chen and Yan Feng Li. Combining the hanning windowed interpolated FFT in both directions. Computer Physics Communications, 178(12):924–928, June 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000933.

Colonna:2008:GNA

[CL08b] G. Colonna and A. Laricchiuta. General numerical algorithm for classical collision integral calculation. Computer Physics Communications, 178(11):809–816, June 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000623.

Cleri:2005:EDC

[Cle05] Fabrizio Cleri. Evolution of dislocation cell structures in plastically deformed metals. Computer Physics Communications, 169(1–3):44–49, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505001098.

Chang:2007:GRN

[CLFH07] Jee-Gong Chang, Cheng-Tai Lee, Yu-Bin Fang, and Chi-Chuan Hwang. Generation of random non-overlapping dot patterns for light guides using molecular dynamics simulations with variable r-cut and reflective boundary techniques. Computer Physics Communications, 177(11):851–862, December 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003463.

Choi:2007:TDF

[CLL^{+}07] Sunmi Choi, Sunghee Lee, Jaekoo Lee, Kyungnam Kim, and Geunyoung Yeom. Two-dimensional fluid simulation of largearea plasma source with parallel resonance antenna. Computer Physics Communications, 177(1–2):126, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001130.

Cecere:2008:CPB

[CLR08] Mariana Cécere, Luis Lehner, and Oscar Reula. Constraint preserving boundary conditions for the ideal Newtonian MHD equations. Computer Physics Communications, 179(8):545– 554, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001665.

Celino:2002:FPI

[CM02a] Massimo Celino and Carlo Massobrio. First principles investigation of the intermediate range order in disordered materials: the case of SiSe_{2}. Computer Physics Communications, 147(1–2):166–169, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002382.

Chakrabarti:2002:DTP

[CM02b] Bikas K. Chakrabarti and Arkajyoti Misra. Dynamic transitions in pure Ising magnets under pulsed and oscillating fields. Computer Physics Communications, 147(1–2):120–125, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002291.

Cucchieri:2003:CSL

[CM03] Attilio Cucchieri and Tereza Mendes. Critical slowing-down in SU(2) landau-gauge-fixing algorithms at β = ∞. Computer Physics Communications, 154(1):1–48, July 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002790.

Carles:2006:CCC

[CM06] A. Grau Carles and A. Grau Malonda. CHEREN2, the Cherenkov counting efficiency by an anisotropy detection model. Computer Physics Communications, 174(1):30–34, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004662.

Chopard:2000:LGM

[CMD00] Bastien Chopard, Alexandre Masselot, and Alexandre Dupuis. A lattice gas model for erosion and particles transport in a fluid. Computer Physics Communications, 129(1–3):167– 176, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001041.

Canning:2000:PFM

[CMF00] A. Canning, W. Mannstadt, and A. J. Freeman. Parallelization of the FLAPW method. Computer Physics Communications, 130(3):233–243, August 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000120X.

Collaboration:2003:SSF

[CMK^{+}03] Alpha Collaboration, Michele Della Morte, Francesco Knechtli, Juri Rolf, Rainer Sommer, Ines Wetzorke, and Ulli Wolff. Simulating the Schrödinger functional with two pseudo-fermions. Computer Physics Communications, 156(1):62–72, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004363.

Cucchieri:2009:LCG

[CMM09] Attilio Cucchieri, Axel Maas, and Tereza Mendes. Linear covariant gauges on the lattice. Computer Physics Communications, 180(2):215–225, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003184.

Cleri:2001:DHL

[CMR01] Fabrizio Cleri, Alessandro Marongiu, and Vittorio Rosato. Dedicated hardware for linearly-scaling algorithms in condensedmatter physics. Computer Physics Communications, 139 (1):20–33, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501002247.

Cain:2002:UCC

[CMRS02] Philipp Cain, Frank Milde, Rudolf A. Römer, and Michael Schreiber. Use of cluster computing for the Anderson model of localization. Computer Physics Communications, 147(1– 2):246–250, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002564.

Carretero:2004:NSC

[CMS04] Jorge Carretero and Manuel Martínez-Sánchez. Numerical simulation of a colloidal thruster in the droplet regime. Computer Physics Communications, 164(1–3):202–208, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002814.

Chou:2000:SMT

[CMT00] Wenchien Chou, Ryoji Matsumoto, and Fumiko Tajima. Simulational modeling of time-domain magnetic induction using parallel computational schemes: (I) in the Cartesian coordinates. Computer Physics Communications, 131(1–2):26–40, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000709.

Chou:2001:SMT

[CMT01] Wenchien Chou, Ryoji Matsumoto, and Fumiko Tajima. Simulational modeling of time-domain magnetic induction using parallel computational schemes: (II) in the spherical coordinates. Computer Physics Communications, 138(2):175–191, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002107.

Czyz:2000:DST

[CN00] Janusz Czyz and Dorota Nieścioruk. On “drunken sinusoids” and their Fourier series. Computer Physics Communications, 126(1–2):51–56, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004154.

Cadilhe:2001:FEA

[CN01] M. W. Cadilhe and J. J. Soares Neto. A finite element algorithm using linear tetrahedral elements for quantum mechanical calculations. Computer Physics Communications, 136(3):183– 197, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002393.

Cunha-Netto:2009:TDL

[CNDC09] A. G. Cunha-Netto, Ronald Dickman, and A. A. Caparica. Two-dimensional lattice polymers: Adaptive windows simulations. Computer Physics Communications, 180(4):583–586, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004335.

Chan:2001:AEA

[CNFR01] Zeke S. H. Chan, H. W. Ngan, Y. F. Fung, and A. B. Rad. An advanced evolutionary algorithm for parameter estimation of the discrete Kalman filter. Computer Physics Communications, 142(1–3):248–254, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003320.

Castro:2009:GST

[CNMC09] M. E. Castro, A. Niño, and C. Muñoz-Caro. Gmat. A software tool for the computation of the rovibrational G matrix. Computer Physics Communications, 180(7):1183–1187, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000204.

Cavazzoni:2005:HPC

[COE^{+}05] Carlo Cavazzoni, Tomaso Esposti Ongaro, Giovanni Erbacci, Augusto Neri, and Giovanni Macedonio. High performance computing simulations of pyroclastic flows. Computer Physics Communications, 169(1–3):454–456, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001979.

Colonna:2007:IHA

[Col07] Gianpiero Colonna. Improvements of hierarchical algorithm for equilibrium calculation. Computer Physics Communications, 177(6):493–499, September 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002457.

Conrad:2004:PCI

[Con04] Jan Conrad. A program for confidence interval calculations for a Poisson process with background including systematic uncertainties: POLE 1.0. Computer Physics Communications, 158(2):117–123, April 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504000037.

Caron:2000:PPC

[CP00] C. F. R. Caron and R. M. Potvliege. propBG: a program for calculating the intensity and the far-field profile of harmonics emitted by a dilute gas irradiated by an intense Gaussian or Bessel–Gauss laser beam. Computer Physics Communications, 126(3):269–293, April 11, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004932.

Cerimele:2000:PIS

[CPS00] M. M. Cerimele, F. Pistella, and S. Succi. Particleinspired scheme for the Gross–Pitaevski equation: An application to Bose–Einstein condensation. Computer Physics Communications, 129(1–3):82–90, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000953.

Cruz:2001:SSP

[CPT^{+}01] A. Cruz, J. Pech, A. Tarancón, P. Téllez, C. L. Ullod, and C. Ungil. SUE: a special purpose computer for spin glass models. Computer Physics Communications, 133(2–3):165–176, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001703.

Caliste:2008:SES

[CPV^{+}08] D. Caliste, Y. Pouillon, M. J. Verstraete, V. Olevano, and X. Gonze. Sharing electronic structure and crystallographic data with ETSF IO. Computer Physics Communications, 179(10):748–758, November 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001963.

Cafarella:2009:HPG

[CPW09] Alessandro Cafarella, Costas G. Papadopoulos, and Małgorzata Worek. Helac–Phegas: a generator for all parton level processes. Computer Physics Communications, 180(10):1941– 1955, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001465.

Colombo:2000:PTB

[CR00] Luciano Colombo and Mario Rosati. Parallel tight-binding molecular dynamics simulations on symmetric multi-processing platforms. Computer Physics Communications, 128(1–2):108– 117, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000795.

Canning:2005:SFP

[CR05] A. Canning and D. Raczkowski. Scaling first-principles plane-wave codes to thousands of processors. Computer Physics Communications, 169(1–3):449–453, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001967.

Chatterjee:2008:CNT

[CR08] Sanjukta Chatterjee and Dhiranjan Roy. A class of new transforms tailored for the hypergeometric series. Computer Physics Communications, 179(8):555–561, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550800177X.

Cranmer:2001:KEH

[Cra01] Kyle Cranmer. Kernel estimation in high-energy physics. Computer Physics Communications, 136(3):198–207, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002435.

Creutz:2000:LFE

[Cre00] Michael Creutz. Lattice fields and extra dimensions. Computer Physics Communications, 127(1):37–42, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000175.

Carnevale:2008:NRI

[CRPC08] V. Carnevale, S. Raugei, S. Piana, and P. Carloni. On the nature of the reaction intermediate in the HIV-1 protease: a quantum chemical study. Computer Physics Communications, 179(1–3):120–123, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550800026X.

Chillemi:2001:RCT

[CRS01] G. Chillemi, M. Rosati, and N. Sanna. The role of computer technology in applied computational chemical-physics. Computer Physics Communications, 139(1):1–19, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002235.

Carjan:2005:IBC

[CRS05] Nicolae Carjan, Margarit Rizea, and Dan Strottman. Improved boundary conditions for the decay of low lying metastable proton states in a time-dependent approach. Computer Physics Communications, 173(1–2):41–60, December 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500425X.

Crouseilles:2009:FSL

[CRS09] Nicolas Crouseilles, Thomas Respaud, and Eric Sonnendrücker. A forward semi-Lagrangian method for the numerical solution of the Vlasov equation. Computer Physics Communications, 180(10):1730–1745, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001301.

Chekanov:2000:GSC

[CRUV00] N. A. Chekanov, V. A. Rostovtsev, Y. Uwano, and S. I. Vinitsky. GITA ^{−1}: a symbolic computing program for an inverse problem of the Birkhoff–Gustavson normal form expansion. Computer Physics Communications, 126(1–2):47– 50, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004919.

Caprio:2009:CSH

[CRW09] M. A. Caprio, D. J. Rowe, and T. A. Welsh. Construction of SO(5) ⊃ SO(3) spherical harmonics and Clebsch–Gordan coefficients. Computer Physics Communications, 180(7):1150– 1163, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550900006X.

Chuvakin:2002:EPP

[CS02] A. Chuvakin and J. Smith. Evolution program for parton densities with perturbative heavy flavor boundary conditions. Computer Physics Communications, 143(3):257–286, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004556.

Chung:2007:TMC

[CS07] T. H. Chung and Y. M. Shin. Theoretical model of cylindrical Langmuir probe for low-pressure electronegative discharges. Computer Physics Communications, 177(1–2):127, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700104X.

Cheng:2004:HPD

[CSC^{+}04] Shannon Cheng, Mark Santi, Murat Celik, Manuel Martinez-Sanchez, and Jaime Peraire. Hybrid PIC-DSMC simulation of a Hall thruster plume on unstructured grids. Computer Physics Communications, 164(1–3):73–79, December 1– 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002590.

Chen:2007:TDS

[CSC^{+}07] P.-Y. Chen, Y.-L. Shao, K.-W. Cheng, K.-H. Hsu, J.-S. Wu, and J.-P. Ju. Three-dimensional simulation studies on electrostatic predictions for carbon nanotube field effect transistors. Computer Physics Communications, 177(9):683–688, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003013. See erra-

tum [CSC^{+}08].

Chen:2008:ETD

[CSC^{+}08] P.-Y. Chen, Y.-L. Shao, K.-W. Cheng, K.-H. Hsu, J.-S. Wu, and J.-P. Yu. Erratum to “Three-dimensional simulation studies on electrostatic predictions for carbon nanotube field effect transistors” [Comput. Phys. Commun. 177 (2007) 683– 688]. Computer Physics Communications, 178(10):794, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000465. See [CSC^{+}07].

Cassol-Seewald:2008:NAG

[CSCK08] N. C. Cassol-Seewald, M. I. M. Copetti, and G. Krein. Numerical approximation of the Ginzburg–Landau equation with memory effects in the dynamics of phase transitions. Computer Physics Communications, 179(5):297–309, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001100.

Calzavarini:2003:MFC

[CSS^{+}03] E. Calzavarini, L. Sartori, F. Schifano, R. Tripiccione, and A. Viceré. Matched filters for coalescing binaries detection on massively parallel computers. Computer Physics Communications, 152(3):295–306, May 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008482.

Chandrasekharan:2002:SLM

[CSW02] S. Chandrasekharan, B. Scarlet, and U.-J. Wiese. From spin ladders to the 2D O(3) model at non-zero density. Computer Physics Communications, 147(1–2):388–393, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003119.

Coveney:2007:AHE

[CSZ^{+}07] P. V. Coveney, R. S. Saksena, S. J. Zasada, M. McKeown, and S. Pickles. The application hosting environment: Lightweight middleware for grid-based computational science. Computer Physics Communications, 176(6):406–418, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506004206.

Chen:2000:TOI

[CT00] Hudong Chen and Chris Teixeira. H-theorem and origins of instability in thermal lattice Boltzmann models. Computer Physics Communications, 129(1–3):21–31, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000898.

Colonna:2001:NMS

[CTG01] G. Colonna, M. Tuttafesta, and D. Giordano. Numerical methods to solve Euler equations in one-dimensional steady nozzle flow. Computer Physics Communications, 138(3):213–221, August 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002119.

Choe:2007:EPS

[CTI07] Yoong-Kee Choe, Eiji Tsuchida, and Tamio Ikeshoji. Electrical property of a sulfuric acid-water mixture from the first-principles molecular dynamics simulation. Computer Physics Communications, 177(1–2):38–39, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001427.

Cheb-Terrab:2000:AOE

[CTR00] E. S. Cheb-Terrab and A. D. Roche. Abel ODEs: Equivalence and integrable classes. Computer Physics Communications, 130(1–2):204–231, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000424.

Chelikowsky:2007:AEE

[CTSZ07] James R. Chelikowsky, Murilo L. Tiago, Yousef Saad, and Yunkai Zhou. Algorithms for the evolution of electronic properties in nanocrystals. Computer Physics Communications, 177(1–2):1–5, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000458.

Cundy:2009:LLW

[Cun09] Nigel Cundy. Low-lying Wilson Dirac operator eigenvector mixing in dynamical overlap hybrid Monte Carlo. Computer Physics Communications, 180(2):180–191, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003020.

Cundy:2005:NMQ

[CvdEF^{+}05] N. Cundy, J. van den Eshof, A. Frommer, S. Krieg, Th. Lippert, and K. Schäfer. Numerical methods for the QCD overlap operator: III. Nested iterations. Computer Physics Communications, 165(3):221–242, February 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005028.

Cornwell:2000:PMD

[CW00] C. F. Cornwell and L. T. Wille. Parallel molecular dynamics simulations for short-ranged many-body potentials. Computer Physics Communications, 128(1–2):477–491, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000515.

Czezowski:2001:OSC

[CW01] Adam Czezowski and Cliff Woodward. Optimizing simulations of confined polymer melts at constant chemical potential. Computer Physics Communications, 142(1–3):117–122, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003447.

Cai:2002:FBS

[CW02] Jun Cai and Jian-Sheng Wang. Friction between Si tip and (001)-2×1 surface: a molecular dynamics simulation. Computer Physics Communications, 147(1–2):145–148, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002825.

Chen:2008:HAG

[CWSH08] Jiaqi Chen, Zhongcheng Wang, Hezhu Shao, and Hailing Hao. Highly-accurate ground state energies of the He atom and the He-like ions by Hartree SCF calculation with Obrechkoff method. Computer Physics Communications, 179(7):486–491, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001525.

Chang:2006:BUV

[CWW06a] Chao-Hsi Chang, Jian-Xiong Wang, and Xing-Gang Wu. BCVEGPY2.0: An upgraded version of the generator BCVEGPY with the addition of hadroproduction of the P - wave Bc states. Computer Physics Communications, 174 (3):241–251, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505005163.

Chang:2006:UVG

[CWW06b] Chao-Hsi Chang, Jian-Xiong Wang, and Xing-Gang Wu. An upgraded version of the generator BCVEGPY2.0 for hadronic production of Bc meson and its excited states. Computer Physics Communications, 175(9):624–627, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002906.

Chang:2007:GGH

[CWW07] Chao-Hsi Chang, Jian-Xiong Wang, and Xing-Gang Wu. GENXICC: a generator for hadronic production of the double heavy baryons Ξcc, Ξbc, and Ξbb. Computer Physics Communications, 177(5):467–478, September 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002512.

Chu:2001:SSP

[CY01] L. L. Chu and K. W. Yu. Spin scattering of a particle for periodic boundary conditions. Computer Physics Communications, 142(1–3):155–159, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003575.

Casula:2005:RVB

[CYAS05] Michele Casula, Seiji Yunoki, Claudio Attaccalite, and Sandro Sorella. Resonating valence bond wave function: from lattice models to realistic systems. Computer Physics Communications, 169(1–3):386–393, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001839.

Czakon:2006:AAC

[Cza06] M. Czakon. Automatized analytic continuation of Mellin– Barnes integrals. Computer Physics Communications, 175 (8):559–571, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002657.

Cao:2000:ECM

[CZC00] Lilun Cao, Jianping Zhu, and Pasquale Cinnella. An efficient computational method for simulations of reaction-diffusion processes. Computer Physics Communications, 126(1–2):41–46, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559900226X.

deAbreu:2008:DCE

[dA08] Marcos Pimenta de Abreu. Diffuse conditions for efficient solution of multislab radiation transport problems. Computer Physics Communications, 179(6):380–384, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550800146X.

deAndres:2001:AAE

[dAK01] P. L. de Andres and D. A. King. Anisotropic and anharmonic effects through the t-matrix for Low-Energy Electron Diffraction (TMAT V1.1). Computer Physics Communications, 138(3):281–301, August 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002399.

Daniluk:2005:DCR

[Dan05a] Andrzej Daniluk. Dynamical calculations for RHEED intensity oscillations. Computer Physics Communications, 166 (2):123–140, March 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504005594.

Daniluk:2005:KCR

[Dan05b] Andrzej Daniluk. Kinematical calculations of RHEED intensity oscillations during the growth of thin epitaxial films. Computer Physics Communications, 170(3):265–286, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500281X.

Daniluk:2007:ECP

[Dan07] Andrzej Daniluk. An extension of the computer program for dynamical calculations of RHEED intensity oscillations.

Heterostructures. Computer Physics Communications, 176 (1):70–73, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506003249.

Daniluk:2009:MDD

[Dan09a] Andrzej Daniluk. Model-Driven Development for scientific computing. An upgrade of the RHEEDGr program. Computer Physics Communications, 180(11):2394–2397, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002136.

Daniluk:2009:MTS

[Dan09b] Andrzej Daniluk. Multithreaded transactions in scientific computing. the Growth06 v2 program. Computer Physics Communications, 180(7):1219–1220, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000393.

Ditz:2008:CMT

[DB08] P. Ditz and A. Borzi‘. A cascadic monotonic time-discretized algorithm for finite-level quantum control computation. Computer Physics Communications, 178(5):393–399, March 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004237.

Dimitrov:2004:SWB

[DBE^{+}04] D. A. Dimitrov, R. Busby, J. Exby, D. L. Bruhwiler, and J. R. Cary. Secure Web-based invocation of largescale plasma simulation codes. Computer Physics Communications, 164(1–3):464–467, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003200.

Detering:2002:LRL

[DBR^{+}02] F. Detering, V. Yu. Bychenkov, W. Rozmus, R. Sydora, and C. E. Capjack. Langevin representation of laser heating in PIC simulations. Computer Physics Communications, 143(1):48–53, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004325.

Delic:2000:PMK

[DC00] George Delic and Gordon G. Cash. The permanent of 0, 1 matrices and Kallman’s algorithm. Computer Physics Communications, 124(2–3):315–329, February 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004488.

DeFabritiis:2003:DGM

[DC03] G. De Fabritiis and P. V. Coveney. Dynamical geometry for multiscale dissipative particle dynamics. Computer Physics Communications, 153(2):209–226, June 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503001991.

DiFelice:2005:CAS

[DC05a] Rosa Di Felice and Stefano Corni. Computational approach to study electron-transfer proteins: Azurin for bio-molecular devices. Computer Physics Communications, 169(1–3):9–13, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001013.

Drummond:2005:QPS

[DC05b] P. D. Drummond and J. F. Corney. Quantum phasespace simulations of fermions and bosons. Computer Physics Communications, 169(1–3):412–415, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001888.

Dion:2007:GST

[DC07] Claude M. Dion and Eric Cancès. Ground state of the timeindependent Gross–Pitaevskii equation. Computer Physics Communications, 177(10):787–798, November 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003074.

Deng:2007:CSI

[DCJ07] Shaozhong Deng, Wei Cai, and Donald Jacobs. A comparable study of image approximations to the reaction field. Computer Physics Communications, 177(9):689–699, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003049.

DaSilva:2009:CGS

[DCNDC09] Cláudio José DaSilva, A. G. Cunha-Netto, Ronald Dickman, and A. A. Caparica. Constructing ground state configurations of lattice polymers. Computer Physics Communications, 180(4):590–593, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550800413X.

Dobaczewski:2000:SSH

[DD00] J. Dobaczewski and J. Dudek. Solution of the Skyrme– Hartree–Fock equations in the Cartesian deformed harmonicoscillator basis. (III) HFODD (v1.75r): a new version of the program. Computer Physics Communications, 131(1– 2):164–186, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001211.

Deuar:2001:SGQ

[DD01] P. Deuar and P. D. Drummond. Stochastic gauges in quantum dynamics for many-body simulations. Computer Physics Communications, 142(1–3):442–445, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003848.

Danese:2001:MCM

[DDD^{+}01] G. Danese, I. De Lotto, A. De Marchi, F. Leporati, T. Bellini, M. Buscaglia, and F. Mantegazza. Monte Carlo Metropolis simulation of interacting anisotropic polarizable spins on a lattice. Computer Physics Communications, 134(1):47–57, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001909.

Duarte:2002:EPS

[DDdMS02] L. G. S. Duarte, S. E. S. Duarte, L. A. C. P. da Mota, and J. E. F. Skea. An extension of the Prelle–Singer method and a Maple implementation. Computer Physics Communications, 144(1):46–62, March 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004623.

Dkaki:2000:IDS

[DDEM00] Taoufiq Dkaki, Bernard Dousset, Daniel Egret, and Josiane Mothe. Information discovery from semi-structured sources — application to astronomical literature. Computer Physics Communications, 127(2–3):198–206, May 10, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599005093.

DeLauro:2009:LCN

[DDFI09] E. De Lauro, S. De Martino, M. Falanga, and L. Gr. Ixaru. Limit cycles in nonlinear excitation of clusters of classical oscillators. Computer Physics Communications, 180(10): 1832–1838, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001568.

DeRaedt:2005:DEB

[DDM05] K. De Raedt, H. De Raedt, and K. Michielsen. Deterministic event-based simulation of quantum phenomena. Computer Physics Communications, 171(1):19–39, September 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002845.

DeRaedt:2007:CPS

[DDM07] K. De Raedt, H. De Raedt, and K. Michielsen. A computer program to simulate Einstein–Podolsky–Rosen– Bohm experiments with photons. Computer Physics Communications, 176(11–12):642–651, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001853.

DeRaedt:2006:EDP

[DDMM06] H. De Raedt, K. De Raedt, K. Michielsen, and S. Miyashita. Efficient data processing and quantum phenomena: Singleparticle systems. Computer Physics Communications, 174 (10):803–817, May 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000373.

Denner:2003:RMC

[DDRW03] A. Denner, S. Dittmaier, M. Roth, and D. Wackeroth. RacoonWW1.3: a Monte Carlo program for four-fermion production at e^{+}e^{−} colliders. Computer Physics Communications, 153(3):462–507, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002054.

deDoncker:2004:CLI

[dDSFY04] Elise de Doncker, Yoshimitsu Shimizu, Junpei Fujimoto, and Fukuko Yuasa. Computation of loop integrals using extrapolation. Computer Physics Communications, 159(3):145–156, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000815.

DeWit:2002:ACE

[De 02] David De Wit. Automatic construction of explicit R matrices for the one-parameter families of irreducible typical highest weight (O_{m}|α_{n}) representations of U_{q}[gl(m|n)]. Computer Physics Communications, 145(2):205–255, May 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004635.

DeFabritiis:2007:PCP

[De 07] G. De Fabritiis. Performance of the cell processor for biomolecular simulations. Computer Physics Communications, 176 (11–12):660–664, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001981.

Decyk:2007:UFM

[Dec07] Viktor K. Decyk. UPIC: a framework for massively parallel particle-in-cell codes. Computer Physics Communications, 177(1–2):95–97, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001452.

Deluzet:2003:MMP

[Del03] Fabrice Deluzet. Mathematical modeling of plasma opening switches. Computer Physics Communications, 152(1):34–54, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007968.

Deloff:2008:GLC

[Del08] A. Deloff. Gauss–Legendre and Chebyshev quadratures for singular integrals. Computer Physics Communications, 179(12):908–914, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002567.

Damjanovic:2005:MDS

[DELG05] Ana Damjanović, Bertrand García-Moreno E., Eaton E. Lattman, and Angel E. García. Molecular dynamics study of hydration of the protein interior. Computer Physics Communications, 169(1–3):126–129, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500127X.

Demetriou:2003:LFP

[Dem03] I. C. Demetriou. L1PMA: a Fortran 77 package for best L_{1} piecewise monotonic data smoothing. Computer Physics Communications, 151(3):315–338, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502007397.

Demetriou:2006:LFP

[Dem06] I. C. Demetriou. L2CXCV: a Fortran 77 package for least squares convex/concave data smoothing. Computer Physics Communications, 174(8):643–668, April 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000075.

Deng:2008:IIM

[Den08] Shaozhong Deng. On the immersed interface method for solving time-domain Maxwell’s equations in materials with curved dielectric interfaces. Computer Physics Communications, 179(11):791–800, December 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002440.

Deveikis:2005:PGO

[Dev05] A. Deveikis. A program for generating one-particle and two-particle coefficients of fractional parentage for the single j-orbit with isospin. Computer Physics Communications, 173(3):186–192, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505004315.

Doicu:2000:MME

[DEW00] A. Doicu, Yu. A. Eremin, and T. Wriedt. T - and D-matrix methods for electromagnetic scattering by impedance obstacles. Computer Physics Communications, 124(1):19–27, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004269.

Doicu:2001:SEW

[DEW01] A. Doicu, Yu. Eremin, and T. Wriedt. Scattering of evanescent waves by a particle on or near a plane surface. Computer Physics Communications, 134(1):1–10, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000179X.

Decyk:2008:OOD

[DG08] Viktor K. Decyk and Henry J. Gardner. Object-oriented design patterns in Fortran 90/95: mazev1, mazev2 and mazev3. Computer Physics Communications, 178(8):611–620, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004857.

Dilaver:2006:DCI

[DGAG06] M. Dilaver, S. Gündüç, M. Aydin, and Y. Gündüç. Dynamic critical index of the Swendsen–Wang algorithm by dynamic finite-size scaling. Computer Physics Communications, 175(8):553–558, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002736.

Deng:2008:IRN

[DGLB08] Lih-Yuan Deng, Rui Guo, Dennis K. J. Lin, and Fengshan Bai. Improving random number generators in the Monte Carlo simulations via twisting and combining. Computer Physics Communications, 178(6):401–408, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700447X.

Dominguez-Garcia:2009:JIB

[DGR09] P. Domínguez-García and M. A. Rubio. JChainsAnalyser: an ImageJ-based stand-alone application for the study of magneto-rheological fluids. Computer Physics Communications, 180(10):1956–1960, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001271.

Degrassi:2008:SFC

[DGS08] G. Degrassi, P. Gambino, and P. Slavich. SusyBSG: a Fortran code for BR[B → Xsy] in the MSSM with Minimal Flavor Violation. Computer Physics Communications, 179(10):759–771, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002312.

Domenech-Garret:2009:QOP

[DGSL09] Juan-Luis Domenech-Garret and Miguel-Angel Sanchis-Lozano. QQ-onia package: a numerical solution to the Schrödinger radial equation for heavy quarkonium. Computer Physics Communications, 180(5):768–778, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003895.

DeBartolo:2008:KFN

[DGV08] R. De Bartolo, A. Greco, and P. Veltri. A 3D kinetic-fluid numerical code for stationary equilibrium states in magnetized plasmas. Computer Physics Communications, 178(9):647–664, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004912.

Dyadkin:2000:SBM

[DH00] Iosif G. Dyadkin and Kenneth G. Hamilton. A study of 128-bit multipliers for congruential pseudorandom number generators. Computer Physics Communications, 125(1–3): 239–258, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://cpc.

cs.qub.ac.uk/summaries/ADLK;   http://www.elsevier.
com/gej-ng//10/15/40/55/25/42/abstract.html;   http:
//www.sciencedirect.com/science/article/pii/S0010465599004671.

Dobbs:2001:HCM

[DH01] Matt Dobbs and Jørgen Beck Hansen. The HepMC C++ Monte Carlo event record for high energy physics. Computer Physics Communications, 134(1):41–46, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001892.

Dobnikar:2004:PBB

[DHB^{+}04] J. Dobnikar, D. Halozan, M. Brumen, H.-H. von Grünberg, and R. Rzehak. Poisson–Boltzmann Brownian dynamics of charged colloids in suspension. Computer Physics Communications, 159(2):73–92, May 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504000530.

Dutt:2005:IGD

[DHBE05] Meenakshi Dutt, Bruno Hancock, Craig Bentham, and James Elliott. An implementation of granular dynamics for simulating frictional elastic particles based on the DL POLY code. Computer Physics Communications, 166(1):26–44, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005053.

DeRaedt:2000:QCE

[DHMD00] Hans De Raedt, Anthony H. Hams, Kristel Michielsen, and Koen De Raedt. Quantum computer emulator. Computer Physics Communications, 132(1–2):1–20, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001326.

Dubbeldam:2000:DSF

[DHS00] D. Dubbeldam, A. G. Hoekstra, and P. M. A. Sloot. Dynamic structure factor in single- and two-species thermal GBL lattice gas. Computer Physics Communications, 129(1–3):13– 20, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000886.

DiPierro:2001:MDP

[Di 01] Massimo Di Pierro. Matrix distributed processing: a set of C++ tools for implementing generic lattice computations on parallel systems. Computer Physics Communications, 141(1):98–148, November 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002971.

DiMarzio:2002:ANS

[Di 02] Frank Di Marzio. Analytic nuclear scattering theories. Computer Physics Communications, 143(3):241–256, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004593.

Dannert:2004:VSK

[DJ04] Tilman Dannert and Frank Jenko. Vlasov simulation of kinetic shear Alfvén waves. Computer Physics Communications, 163(2):67–78, November 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504004357.

Deveikis:2008:CPT

[DJ08] A. Deveikis and A. Juodagalvis. A computer program for twoparticle generalized coefficients of fractional parentage. Computer Physics Communications, 179(8):607–613, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002208.

Deveikis:2005:ASC

[DK05] A. Deveikis and A. Kuznecovas. The analytical scheme calculator for angular momentum coupling and recoupling coefficients. Computer Physics Communications, 172(1):60–67, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505003589.

Deviren:2008:DPT

[DKC08] Bayram Deviren, Mustafa Keskin, and Osman Canko. Dynamic phase transition and multicritical dynamic phase diagrams of the kinetic spin-3/2 Blume–Emery–Griffiths model with repulsive biquadratic coupling under a timedependent oscillating external field. Computer Physics Communications, 178(6):420–437, March 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004596.

Djouadi:2007:SFC

[DKM07] Abdelhak Djouadi, Jean-Loïc Kneur, and Gilbert Moultaka. SuSpect: a Fortran code for the supersymmetric and Higgs particle spectrum in the MSSM. Computer Physics Communications, 176(6):426–455, March 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550600419X.

DeRaedt:2003:UFN

[DKMF03] H. De Raedt, J. S. Kole, K. F. L. Michielsen, and M. T. Figge. Unified framework for numerical methods to solve the time-dependent Maxwell equations. Computer Physics Communications, 156(1):43–61, December 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004144.

Dyshko:2000:ASH

[DKV00] A. L. Dyshko, N. B. Konyukhova, and N. A. Voronov. Automodelling solutions of the Higgs-field nonlinear wave equation in the de Sitter space. Computer Physics Communications, 126(1–2):57–62, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002428.

delaGrandmaison:2005:ECC

[dlGGS^{+}05] Emmanuel Lorin de la Grandmaison, Shivaraju B. Gowda, Yousef Saad, Murilo L. Tiago, and James R. Chelikowsky. Efficient computation of the coupling matrix in timedependent density functional theory. Computer Physics Communications, 167(1):7–22, April 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005648.

delaHoz:2008:ETD

[dlHV08] F. de la Hoz and F. Vadillo. An exponential time differencing method for the nonlinear Schrödinger equation. Computer Physics Communications, 179(7):449–456, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001483.

Otero-de-la-Roza:2009:CNP

[dlRBPL09] A. Otero de-la Roza, M. A. Blanco, A. Martín Pendás, and Víctor Luaña. Critic: a new program for the topological analysis of solid-state electron densities. Computer Physics Communications, 180(1):157–166, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002865.

delaRoza:2009:RTB

[dlRL09] A. Otero de la Roza and Víctor Luaña. Runwien: a text-based interface for the WIEN package. Computer Physics Communications, 180(5):800–812, May 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003937.

Deremble:2008:PDR

[DLZ08] Cyril Deremble, Richard Lavery, and Krystyna Zakrzewska. Protein-DNA recognition: Breaking the combinatorial barrier. Computer Physics Communications, 179(1–3):112–119, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000258.

Dziubak:2007:OOC

[DM07] Tomasz Dziubak and Jacek Matulewski. An object-oriented C++ implementation of Davidson method for finding a few selected extreme eigenpairs of a large, sparse, real, symmetric matrix. Computer Physics Communications, 177(8):676–682, October 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003037.

Dehghan:2009:MLB

[DM09] Mehdi Dehghan and Davoud Mirzaei. Meshless local boundary integral equation (LBIE) method for the unsteady magnetohydrodynamic (MHD) flow in rectangular and circular pipes. Computer Physics Communications, 180(9):1458–1466, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001039.

deMatos:2006:BSG

[dMBC^{+}06] Jonathan de Matos, Eduardo Bortolato, Alexandre Camilo, Jr., João A. Martini, Ronaldo A. L. Gonçalves, and Paulo S. L. de Souza. Binary SCF: GAMESS improvements for energy evaluation based on SCF methods. Computer Physics Communications, 174(1):1–16, January 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004686.

DeRaedt:2007:MPQ

[DMD^{+}07] K. De Raedt, K. Michielsen, H. De Raedt, B. Trieu, G. Arnold, M. Richter, Th. Lippert, H. Watanabe, and N. Ito. Massively parallel quantum computer simulator. Computer Physics Communications, 176(2):121–136, January 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003390.

Darlington:2001:SAS

[DMR01] Rebecca M. Darlington, Thomas L. McAbee, and Garry Rodrigue. A study of ALE simulations of Rayleigh–Taylor instability. Computer Physics Communications, 135(1):58–73, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002162.

Darlington:2002:LES

[DMR02] Rebecca M. Darlington, Thomas L. McAbee, and Garry Rodrigue. Large eddy simulation and ALE mesh motion in Rayleigh–Taylor instability simulation. Computer Physics Communications, 144(3):261–276, April 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002631.

Decyk:2004:UPP

[DN04] Viktor K. Decyk and Charles D. Norton. UCLA parallel PIC framework. Computer Physics Communications, 164(1–3):80–85, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002607.

Dellago:2005:DMW

[DN05] Christoph Dellago and Mor M. Naor. Dipole moment of water molecules in narrow pores. Computer Physics Communications, 169(1–3):36–39, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001074.

deNiem:2007:VFM

[dNKM07] D. de Niem, E. Kührt, and U. Motschmann. A volumeof-fluid method for simulation of compressible axisymmetric multi-material flow. Computer Physics Communications, 176(3):170–190, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550600364X.

deOliveira:2002:SDB

[dO02] Paulo Murilo Castro de Oliveira. Simple dynamics for broad histogram method. Computer Physics Communications, 147(1–2):410–413, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003156.

Dobaczewski:2004:SSH

[DO04] J. Dobaczewski and P. Olbratowski. Solution of the Skyrme– Hartree–Fock–Bogolyubov equations in the Cartesian deformed harmonic-oscillator basis. (IV) HFODD (v2.08i): a new version of the program. Computer Physics Communications, 158(3):158–191, April 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504000797.

Dobaczewski:2005:SSH

[DO05] J. Dobaczewski and P. Olbratowski. Solution of the Skyrme– Hartree–Fock–Bogolyubov equations in the Cartesian deformed harmonic-oscillator basis. (V) HFODD (v2.08k). Computer Physics Communications, 167(3):214–216, May 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000779.

deOliveira:2009:GSS

[dO09] Mário J. de Oliveira. Glassy states in the stochastic Potts model. Computer Physics Communications, 180(4):480–484, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000228.

Doll:2001:IAH

[Dol01] K. Doll. Implementation of analytical Hartree–Fock gradients for periodic systems. Computer Physics Communications, 137(1):74–88, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501001722.

Domany:2005:ADC

[Dom05] Eytan Domany. Analysis of DNA-chip and antigen-chip data: studies of cancer, stem cells and autoimmune diseases. Computer Physics Communications, 169(1–3):183–187, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001384.

Donolato:2002:ANI

[Don02] C. Donolato. Analytical and numerical inversion of the Laplace–Carson transform by a differential method. Computer Physics Communications, 145(2):298–309, May 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002813.

Desplat:2001:LPL

[DPB01] Jean-Christophe Desplat, Ignacio Pagonabarraga, and Peter Bladon. LUDWIG: a parallel lattice-Boltzmann code for complex fluids. Computer Physics Communications, 134 (3):273–290, March 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002058.

Dorfi:2006:TMC

[DPSG06] E. A. Dorfi, H. Pikall, A. Stökl, and A. Gautschy. Towards a more consistent discretization scheme for adaptive implicit RHD computations. Computer Physics Communications, 174(10):771–782, May 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506000154.

Dehghan:2009:GBH

[DR09] Mehdi Dehghan and Narges Rastegar. On the global behavior of a high-order rational difference equation. Computer Physics Communications, 180(6):873–878, June 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004232.

Draeger:2001:MPC

[Dra01] Thomas Draeger. A Macsyma program for computing analytic Clebsch–Gordan coefficients of U(3) ⊃ SO(3). Computer Physics Communications, 139(3):246–262, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002053.

Dall:2001:FMC

[DS01] Jesper Dall and Paolo Sibani. Faster Monte Carlo simulations at low temperatures. The waiting time method. Computer Physics Communications, 141(2):260–267, November 30, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100412X.

daSilva:2003:OMM

[dS03]  Cesar R. S. da Silva.  Optimizing Metropolis Monte

Carlo simulations of semiconductors. Computer Physics Communications, 153(3):392–396, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550300225X.

DeGrand:2004:IMT

[DS04] Thomas DeGrand and Stefan Schaefer. Improving meson two-point functions in lattice QCD. Computer Physics Communications, 159(3):185–191, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504000803.

Daniluk:2006:NVC

[DS06] Andrzej Daniluk and Kazimierz Skrobas. A new version of a computer program for dynamical calculations of RHEED intensity oscillations. Computer Physics Communications, 174(1):83–85, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005096.

Bernardes:2000:KAC

[dSB00] Esmerindo de Sousa Bernardes. Killing — an algebraic computational package for Lie algebras. Computer Physics Communications, 130(1–2):137–175, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000060.

Duncan:2006:FFT

[DSC06] A. Duncan, R. D. Sedgewick, and R. D. Coalson. Fast Fourier transform simulation techniques for Coulomb gases. Computer Physics Communications, 175(2):73–77, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001044.

Dobaczewski:2009:SSH

[DSC^{+}09] J. Dobaczewski, W. Satula, B. G. Carlsson, J. Engel, P. Olbratowski, P. Powalowski, M. Sadziak, J. Sarich, N. Schunck, A. Staszczak, M. Stoitsov, M. Zalewski, and H. Zduńczuk. Solution of the Skyrme–Hartree–Fock–Bogolyubov equations in the Cartesian deformed harmonic-oscillator basis.: (VI) hfodd (v2.40h): a new version of the program. Computer Physics Communications, 180(11):2361–2391, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002598.

daSilveira:2008:MMD

[dSdSW08] Pedro R. C. da Silveira, Cesar R. S. da Silva, and Renata M. Wentzcovitch. Metadata management for distributed first principles calculations in VLab — a collaborative cyberinfrastructure for materials computation. Computer Physics Communications, 178(3):186–198, February 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003955.

Doicu:2002:IRG

[DSH02] Adrian Doicu, Franz Schreier, and Michael Hess. Iteratively regularized Gauss–Newton method for atmospheric remote sensing. Computer Physics Communications, 148(2):214–226, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005556.

Doicu:2003:IRG

[DSH03] Adrian Doicu, Franz Schreier, and Michael Hess. Iteratively regularized Gauss–Newton method for bound-constraint problems in atmospheric remote sensing. Computer Physics Communications, 153(1):59–65, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503001383.

Doicu:2005:MPR

[DSHH05] Adrian Doicu, Franz Schreier, Siegfried Hilgers, and Michael Hess. Multi-parameter regularization method for atmospheric remote sensing. Computer Physics Communications, 165(1):1– 9, January 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004382.

deSouza:2002:RHT

[dSL02] Adauto J. F. de Souza and M. L. Lyra. Reweighting histogram technique for the quantum Heisenberg ferromagnet. Computer Physics Communications, 146(1):16–23, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004307.

Dunweg:2009:PUF

[DSL09] Burkhard Dünweg, Ulf D. Schiller, and Anthony J. C. Ladd. Progress in the understanding of the fluctuating lattice Boltzmann equation. Computer Physics Communications, 180 (4):605–608, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000290.

Donangelo:2001:CPS

[DSS01] R. Donangelo, K. Sneppen, and S. R. Souza. A computer program for statistical multifragmentation of nuclei. Computer Physics Communications, 140(3):405–411, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002090.

Deppman:2002:MCN

[DTD^{+}02] A. Deppman, O. A. P. Tavares, S. B. Duarte, E. C. de Oliveira, J. D. T. Arruda-Neto, S. R. de Pina, V. P. Likhachev, O. Rodriguez, J. Mesa, and M. Gonçalves. The MCEF code for nuclear evaporation and fission calculations. Computer Physics Communications, 145(3):385–394, June 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002783.

Duan:2009:CPD

[DTHL09] Y. Duan, P. F. Tang, T. Z. Huang, and S. J. Lai. Coupling projection domain decomposition method and Kansa’s method in electrostatic problems. Computer Physics Communications, 180(2):209–214, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003172.

Dunlap:2005:AMM

[Dun05] Brett I. Dunlap. Angular momentum in molecular quantum mechanical integral evaluation. Computer Physics Communications, 165(1):18–36, January 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004400.

Dupuis:2001:NIT

[Dup01] M. Dupuis. New integral transforms for molecular properties and application to a massively parallel GIAO-SCF implementation. Computer Physics Communications, 134(2):150–166, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001958.

Durr:2005:GAI

[Dür05] Stephan Dürr. Gauge action improvement and smearing. Computer Physics Communications, 172(3):163–186, November 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004017.

Durr:2009:LLS

[Dür09] Stephan Dürr. Logarithmic link smearing for full QCD. Computer Physics Communications, 180(8):1338–1357, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900068X.

Dudson:2009:BFP

[DUX^{+}09] B. D. Dudson, M. V. Umansky, X. Q. Xu, P. B. Snyder, and H. R. Wilson. BOUT++: a framework for parallel plasma fluid simulations. Computer Physics Communications, 180(9):1467–1480, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001040.

DeStefano:2005:PDT

[DVG05] Giuliano De Stefano, Oleg V. Vasilyev, and Daniel E. Goldstein. A-priori dynamic test for deterministic/stochastic modeling in large-eddy simulation of turbulent flow. Computer Physics Communications, 169(1–3):210–213, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500144X.

Duchovic:2002:PPE

[DVL^{+}02] Ronald J. Duchovic, Yuri L. Volobuev, Gillian C. Lynch, Donald G. Truhlar, Thomas C. Allison, Albert F. Wagner, Bruce C. Garrett, and Jose C. Corchado. POTLIB 2001: a potential energy surface library for chemical systems. Computer Physics Communications, 144(2):169–187, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004374.

Duchovic:2004:CPL

[DVL^{+}04] Ronald J. Duchovic, Yuri L. Volobuev, Gillian C. Lynch, Donald G. Truhlar, Thomas C. Allison, Albert F. Wagner, Bruce C. Garrett, and Jose C. Corchado. A correction to the POTLIB Library described in “POTLIB 2001: a potential energy surface library for chemical systems”: [Comput. Phys. Comm. 144 (2002) 169–187]. Computer Physics Communications, 156(3):319–322, January 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550300434X.

Doicu:2001:NFM

[DW01] Adrian Doicu and Thomas Wriedt. Null-field method with discrete sources to electromagnetic scattering from layered scatterers. Computer Physics Communications, 138(2):136–142, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002028.

Dai:2005:NHE

[DWZS05] Yongming Dai, Zhongcheng Wang, Deying Zhao, and Xiaolong Song. A new high efficient and high accurate Obrechkoff fourstep method for the periodic nonlinear undamped Duffing’s equation. Computer Physics Communications, 165(2):110–126, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004138.

Dy:2009:PCE

[Dy09] Hoang-Binh Dy. A program to compute exact hydrogenic radial integrals and Einstein coefficients. Computer Physics Communications, 180(10):2020, October 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001842.

Dyshlovenko:2002:ANM

[Dys02] P. E. Dyshlovenko. Adaptive numerical method for Poisson– Boltzmann equation and its application. Computer Physics Communications, 147(1–2):335–338, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002989.

Dzuba:2009:UAC

[Dzu09] V. A. Dzuba. An universal algorithm of calculating terms of atomic many-body perturbation theory. Computer Physics Communications, 180(3):392–400, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003731.

Edvardsson:2001:APE

[EÅ01] Sverker Edvardsson and Daniel Åberg. An atomic program for energy levels of equivalent electrons: lanthanides and actinides. Computer Physics Communications, 133(2–3):396–406, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001715.

Eastwood:2008:BPA

[Eas08] J. W. Eastwood. The Block-P^{3}M algorithm. Computer Physics Communications, 179(1–3):46–50, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000143.

Edvardsson:2005:PAS

[EÅU05] Sverker Edvardsson, Daniel Åberg, and Per Uddholm. A program for accurate solutions of two-electron atoms. Computer Physics Communications, 165(3):260–270, February 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005077.

Ejtehadi:2002:RBF

[EE02] M. R. Ejtehadi and R. Everaers. Rigid-body formalism for simulating macromolecules. Computer Physics Communications, 147(1–2):339–341, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002990.

Elwakil:2004:ETW

[EELZS04] S. A. Elwakil, S. K. El-Labany, M. A. Zahran, and R. Sabry. Exact travelling wave solutions for a diffusion-convection equation in two and three spatial dimensions. Computer Physics Communications, 158(2):113–116, April 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000529.

Egedal:2004:KSV

[EFBP04] J. Egedal, W. Fox, E. Belonohy, and M. Porkolab. Kinetic simulation of the VTF magnetic reconnection experiment. Computer Physics Communications, 164(1–3):29–33, December 1– 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002528.

Eppler:2000:NCS

[EFG^{+}00] W. Eppler, T. Fischer, H. Gemmeke, A. Chilingarian, and A. Vardanyan. Neural chip SAND in online data processing of extensive air showers. Computer Physics Communications, 126(1–2):63–66, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004233.

Egri:2007:LQV

[EFH^{+}07] Győző I. Egri, Zoltán Fodor, Christian Hoelbling, Sándor D. Katz, Dániel Nógrádi, and Kálmán K. Szabó. Lattice QCD as a video game. Computer Physics Communications, 177(8):631–639, October 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003025.

Elkurdi:2008:FAI

[EFS^{+}08] Yousef Elkurdi, David Fernández, Evgueni Souleimanov, Dennis Giannacopoulos, and Warren J. Gross. FPGA architecture and implementation of sparse matrix-vector multiplication for the finite element method. Computer Physics Communications, 178(8):558–570, April 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004869.

Ergul:2009:DED

[EG09] Özgür Ergül and Levent Gürel. Discretization error due to the identity operator in surface integral equations. Computer Physics Communications, 180(10):1746–1752, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001313.

Eriksson:2003:SRE

[EH03] L.-G. Eriksson and P. Helander. Simulation of runaway electrons during tokamak disruptions. Computer Physics Communications, 154(3):175–196, August 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002935.

Ellwanger:2006:NUP

[EH06] Ulrich Ellwanger and Cyril Hugonie. NMHDECAY 2.1: An updated program for sparticle masses, Higgs masses, couplings and decay widths in the NMSSM. Computer Physics Communications, 175(4):290–303, August 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001883.

Ellwanger:2007:NFC

[EH07] Ulrich Ellwanger and Cyril Hugonie. NMSPEC: a Fortran code for the sparticle and Higgs masses in the NMSSM with GUT scale boundary conditions. Computer Physics Communications, 177(4):399–407, August 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002263.

Eisenmenger:2001:SMP

[EHHH01] Frank Eisenmenger, Ulrich H. E. Hansmann, Shura Hayryan, and Chin-Kun Hu. [SMMP] A modern package for simulation of proteins. Computer Physics Communications, 138(2):192–212, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001977.

Eisenmenger:2006:EVS

[EHHH06] Frank Eisenmenger, Ulrich H. E. Hansmann, Shura Hayryan, and Chin-Kun Hu. An enhanced version of SMMP-open-source software package for simulation of proteins. Computer Physics Communications, 174(5):422–429, March 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005850.

Engel:2009:FPC

[EKW09] D. Engel, M. Klews, and G. Wunner. A fast parallel code for calculating energies and oscillator strengths of many-electron atoms at neutron star magnetic field strengths in adiabatic approximation. Computer Physics Communications, 180(2):302–311, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003007.

Ethier:2004:PGP

[EL04] S. Ethier and Z. Lin. Porting the 3D gyrokinetic particlein-cell code GTC to the NEC SX-6 vector architecture: perspectives and challenges. Computer Physics Communications, 164(1–3):456–458, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003182.

Elezgaray:2006:SMD

[EL06] J. Elezgaray and M. Laguerre. A systematic method to derive force fields for coarse-grained simulations of phospholipids. Computer Physics Communications, 175(4):264–268, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506001585.

Elber:2005:CSP

[Elb05] Ron Elber. Computer simulations of protein folding: Classical trajectories by optimization of action. Computer Physics Communications, 169(1–3):277–283, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001608.

Eliasson:2005:PIO

[Eli05] Bengt Eliasson. The parallel implementation of the onedimensional Fourier transformed Vlasov–Poisson system. Computer Physics Communications, 170(2):205–230, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002614.

Eliasson:2008:NNG

[Eli08] B. Eliasson. A nonuniform nested grid method for simulations of RF induced ionospheric turbulence. Computer Physics Communications, 178(1):8–14, January 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003657.

Elmaghraby:2009:POP

[Elm09] Elsayed K. Elmaghraby. PHASE-OTI: a pre-equilibrium model code for nuclear reactions calculations. Computer Physics Communications, 180(9):1694–1699, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001155.

Esleta:2008:SRS

[EM08] G. A. Esleta and C. Monterola. Structural reinforcement in a spring-block model of stress-induced fracture propagation. Computer Physics Communications, 178(9):635–646, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004900.

El-Moghraby:2003:CMQ

[EMJH03a] D. El-Moghraby, R. G. Johnson, and P. Harrison. Calculating modes of quantum wire and dot systems using a finite differencing technique. Computer Physics Communications, 150(3):235–246, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006902.

El-Moghraby:2003:EID

[EMJH03b] D. El-Moghraby, R. G. Johnson, and P. Harrison. The effect of inter-dot separation on the finite difference solution of vertically aligned coupled quantum dots. Computer Physics Communications, 155(3):236–243, November 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003679.

Ertl:2009:FLS

[ES09] Otmar Ertl and Siegfried Selberherr. A fast level set framework for large three-dimensional topography simulations. Computer Physics Communications, 180(8):1242–1250, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000459.

Esirkepov:2001:ECC

[Esi01] T. Zh. Esirkepov. Exact charge conservation scheme for Particle-In-Cell simulation with an arbitrary form-factor. Computer Physics Communications, 135(2):144–153, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002289.

Esquembre:2002:CPE

[Esq02] Francisco Esquembre. Computers in physics education. Computer Physics Communications, 147(1–2):13–18, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502001972.

Esquembre:2004:EJS

[Esq04] Francisco Esquembre. Easy Java Simulations: a software tool to create scientific simulations in Java. Computer Physics Communications, 156(2):199–204, January 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004405.

Ehlert:2000:NSR

[EST00] J. Ehlert, H. Stiel, and K. Teuchner. A numerical solver for rate equations and photon transport equations in nonlinear laser spectroscopy. Computer Physics Communications, 124(2–3):330–339, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004506.

Eggert:2000:PIP

[EVL00] Philip Eggert, Alexandra Viel, and Claude Leforestier. Parallel implementation of a pseudo-spectral calculation of molecular energy levels: Application to the water dimer (H_{2}O)_{2}. Computer Physics Communications, 128(1–2):315–325, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005184.

Efimov:2007:CAP

[EYJ07] Oleg Efimov, Young-Gui Yoon, and Seung-Hoon Jhi. Carcinogenic activity of polycyclic aromatic hydrocarbons in firstprinciples. Computer Physics Communications, 177(1–2):50, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001397.

Fritzsche:2000:CNV

[FA00] S. Fritzsche and J. Anton. CESD99 — A new version to represent atomic wave functions in a determinant basis. Computer Physics Communications, 124(2–3):353–355, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004555.

Fukunaga:2001:DCG

[FAiTD01] Hiroo Fukunaga, Takeshi Aoyagi, Jun ichi Takimoto, and Masao Doi. Derivation of coarse-grained potential for polyethylene. Computer Physics Communications, 142(1–3):224– 226, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003113.

Faraboschi:2001:DTP

[Far01] Paolo Faraboschi. The design of a technology platform for custom VLIW embedded processors. Computer Physics Communications, 139(1):104–108, September 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002338.

Fatullayev:2002:NPD

[Fat02] Afet Golayoglu Fatullayev. Numerical procedure for the determination of an unknown coefficients in parabolic equations. Computer Physics Communications, 144(1):29–33, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004702.

Ferre-Borrull:2001:IHO

[FBB01] Josep Ferré-Borrull and Salvador Bosch. On the integration of highly oscillatory functions: Algorithms based on critical points and semi-periodic zones in scalar diffraction. Computer Physics Communications, 134(1):11–22, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001806.

Flower:2000:MPS

[FBL00] D. R. Flower, G. Bourhis, and J.-M. Launay. Molcol: a program for solving atomic and molecular collision problems. Computer Physics Communications, 131(3):187–201, September 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001181.

Fan:2003:DAC

[FD03] Engui Fan and H. H. Dai. A direct approach with computerized symbolic computation for finding a series of traveling waves to nonlinear equations. Computer Physics Communications, 153(1):17–30, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502008731. See [Hon04].

Fang:2007:PFN

[FDM07] Bin Fang, Yuefan Deng, and Glenn Martyna. Performance of the 3D FFT on the 6D network torus QC-DOC parallel supercomputer. Computer Physics Communications, 176(8):531–538, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507000276.

Fiore:2009:EUC

[FdO09] Carlos E. Fiore and Mário J. de Oliveira. Extending the use of canonical and microcanonical Monte Carlo algorithms to spin models. Computer Physics Communications, 180(9):1434–1441, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000836.

Fung:2001:POP

[FEHC01] Y. F. Fung, M. F. Ercan, T. K. Ho, and W. L. Cheung. Performance optimization for parallel processing on a multiple-CPU server. Computer Physics Communications, 142(1–3):191–195, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003277.

Felder:2008:CPL

[Fel08] Gary Felder. CLUSTEREASY: a program for lattice simulations of scalar fields in an expanding universe on parallel computing clusters. Computer Physics Communications, 179(8):604–606, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002221.

Ferrari:2007:SMP

[Fer07a] A. F. Ferrari. SusyMath: a Mathematica package for quantum superfield calculations. Computer Physics Communications, 176(5):334–346, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003973.

Freysoldt:2007:DAG

[FER^{+}07b] Christoph Freysoldt, Philipp Eggert, Patrick Rinke, Arno Schindlmayr, R. W. Godby, and Matthias Scheffler. Dielectric anisotropy in the GW space–time method. Computer Physics Communications, 176(1):1–13, January 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003134.

Fritzsche:2000:RRP

[FFD00] S. Fritzsche, C. Froese Fischer, and C. Z. Dong. REOS99: a revised program for transition probability calculations including relativistic, correlation, and relaxation effects. Computer Physics Communications, 124(2–3):340–352, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004531.

Fernandes:2001:SSS

[FFF01] Fernando M. S. Silva Fernandes, Rui P. S. Fartaria, and Filomena F. M. Freitas. The starting state in simulations of the fluid-solid coexistence by Gibbs–Duhem integration. Computer Physics Communications, 141(3):403–411, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004350.

Fritzsche:2002:RPR

[FFG02] S. Fritzsche, C. Froese Fischer, and G. Gaigalas. RELCI: a program for relativistic configuration interaction calculations. Computer Physics Communications, 148(1):103–123, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004630.

Fabianski:2002:SSM

[FFK02] Robert Fabiański, Lucyna Firlej, and Bogdan Kuchta. Six-state model of cyanoadamantane crystal. MC simulations. Computer Physics Communications, 147(1–2):174–177, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002837.

Fine:2001:SOF

[FFPW01] V. Fine, Y. Fisyak, V. Perevoztchikov, and T. Wenaus. The STAR offline framework. Computer Physics Communications, 140(1–2):76–81, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002570.

Ferraris:2001:NOR

[FFS01] M. Ferraris, P. Frixione, and S. Squarcia. Network oriented radiological and medical archive. Computer Physics Communications, 140(1–2):226–232, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002739.

Fattebert:2004:LSF

[FG04] Jean-Luc Fattebert and François Gygi. Linear scaling firstprinciples molecular dynamics with controlled accuracy. Computer Physics Communications, 162(1):24–36, September 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003261.

Fang:2004:ESC

[FGA04] Shoufan Fang, George Z. Gertner, and Alan A. Anderson. Estimation of sensitivity coefficients of nonlinear model input parameters which have a multinormal distribution. Computer Physics Communications, 157(1):9–16, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004880.

Filho:2003:NAS

[FGF03] Tarcísio M. Rocha Filho, Iram M. Gléria, and Annibal Figueiredo. A novel approach for the stability problem in non-linear dynamical systems. Computer Physics Communications, 155(1):21–30, September 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002959.

Faure:2002:GFC

[FGMT02] Alexandre Faure, Jimena D. Gorfinkiel, Lesley A. Morgan, and Jonathan Tennyson. GTOBAS: fitting continuum functions with Gaussian-type orbitals. Computer Physics Communications, 144(2):224–241, April 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001418.

FroeseFischer:2006:SCG

[FGR06] Charlotte Froese Fischer, Gediminas Gaigalas, and Yuri Ralchenko. Some corrections to GRASP92. Computer Physics Communications, 175(11–12):738–744, December 1– 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506003377.

Ferrari:2001:QSR

[FGV01] T. Ferrari, A. Ghiselli, and C. Vistoli. Quality of services for remote control in high energy physics experiments: a case study. Computer Physics Communications, 140(1–2):209–218, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002727.

Filippova:2000:NNS

[FH00] Olga Filippova and Dieter Haenel. A novel numerical scheme for reactive flows at low Mach numbers. Computer Physics Communications, 129(1–3):267–274, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001132.

Feil:2004:PAR

[FH04] T. M. Feil and H. H. H. Homeier. Programs for the approximation of real and imaginary single- and multi-valued functions by means of Hermite–Padé-approximants. Computer Physics Communications, 158(2):124–135, April 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000785.

Fachat:2000:SAT

[FHF00] André Fachat, Karl Heinz Hoffmann, and Astrid Franz. Simulated annealing with threshold accepting or Tsallis statistics. Computer Physics Communications, 132(3):232–240, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001533.

Franzrahe:2005:TDM

[FHR^{+}05] K. Franzrahe, P. Henseler, A. Ricci, W. Strepp, S. Sengupta, M. Dreher, Chr. Kircher, M. Lohrer, W. Quester, K. Binder, and P. Nielaba. Two-dimensional model colloids and nano wires: phase transitions, effects of external potentials and quantum effects. Computer Physics Communications, 169(1–3):197– 202, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001414.

Frezzotti:2001:CBF

[FHW^{+}01] Roberto Frezzotti, Martin Hasenbusch, Ulli Wolff, Jochen Heitger, and Karl Jansen. Comparative benchmarks of full QCD algorithms. Computer Physics Communications, 136(1–2):1– 13, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002423.

Fritzsche:2001:MPC

[FIBT01] S. Fritzsche, T. Inghoff, T. Bastug, and M. Tomaselli. Maple procedures for the coupling of angular momenta. V. Recoupling coefficients. Computer Physics Communications, 139(3):314–326, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002193.

Fijalkow:1999:NSV

[Fij99] Eric Fijalkow. A numerical solution to the Vlasov equation. Computer Physics Communications, 116(2–3):319–328, February 1999. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465598001465. See erra-

tum [Fij00].

Fijalkow:2000:ENS

[Fij00] Eric Fijalkow. Erratum to “A numerical solution to the Vlasov equation” [computer physics communications 116 (1999) 319– 328]. Computer Physics Communications, 124(2–3):359, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004336. See [?].

Fujimoto:2003:GSA

[FIJ^{+}03] J. Fujimoto, T. Ishikawa, M. Jimbo, T. Kaneko, K. Kato, S. Kawabata, T. Kon, M. Kuroda, Y. Kurihara, Y. Shimizu, and H. Tanaka. GRACE/SUSY: Automatic generation of tree amplitudes in the Minimal Supersymmetric Standard Model. Computer Physics Communications, 153(1):106–134, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503001590.

Fritzsche:2003:MPC

[FIT03] S. Fritzsche, T. Inghoff, and M. Tomaselli. Maple procedures for the coupling of angular momenta. VII. Extended and accelerated computations. Computer Physics Communications, 153(3):424–444, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002273.

Foley:2005:PAA

[FJC^{+}05] Justin Foley, K. Jimmy Juge, Alan Ó Cais, Mike Peardon, Sinéad M. Ryan, Jon-Ivar Skullerud, and TrinLat Collaboration. Practical all-to-all propagators for lattice QCD. Computer Physics Communications, 172(3):145–162, November 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505004005.

Fleischer:2000:SFB

[FK00] J. Fleischer and M. Yu. Kalmykov. ON-SHELL2: FORM based package for the calculation of two-loop self-energy single scale Feynman diagrams occurring in the standard model. Computer Physics Communications, 128(3):531–549, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005329.

Fukuda:2005:GQR

[FKAM05] H. Fukuda, M. Katuya, E. O. Alt, and A. V. Matveenko. Gaussian quadrature rule for arbitrary weight function and interval. Computer Physics Communications, 167(2):143–150, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000676.

Furlani:2000:PSC

[FKG00] Thomas R. Furlani, Jing Kong, and Peter M. W. Gill. Parallelization of SCF calculations within Q-Chem. Computer Physics Communications, 128(1–2):170–177, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550000059X.

Ferrieres:2009:ECE

[FKMB09] X. Ferrieres, G. Klysz, P. Mazet, and J. P. Balayssac. Evaluation of the concrete electromagnetics properties by using radar measurements in a context of building sustainability. Computer Physics Communications, 180(8):1277–1281, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000642.

Fodor:2003:BTM

[FKP03] Zoltán Fodor, Sándor D. Katz, and Gábor Papp. Better than $l/mflops sustained: a scalable PC-based parallel computer for lattice QCD. Computer Physics Communications, 152(2):121– 134, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007762.

Fruhwirth:2001:MMM

[FL01] R. Frühwirth and M. Liendl. Mixture models of multiple scattering: computation and simulation. Computer Physics Communications, 141(2):230–246, November 30, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004039.

Florek:2001:OMG

[Flo01] Wojciech Florek. Operator matrices generation: Combinatorial structures in finite spin models. Computer Physics Communications, 138(3):264–280, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002405.

Ferreira:2006:MST

[FLO06] João V. B. Ferreira, Valter L. Líbero, and Luiz N. Oliveira. Multi-step transformation in numerical renormalization group. Computer Physics Communications, 174(11):862–868, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506000555.

Fettes:2000:LSD

[FM00] Werner Fettes and Ingo Morgenstern. Limitations of the stochastic diagonalization and the projector QMC in the 2D Hubbard model. Computer Physics Communications, 124 (2–3):148–168, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004403.

Forbert:2003:CPS

[FM03] Harald Forbert and Dominik Marx. Calculation of the permanent of a sparse positive matrix. Computer Physics Communications, 150(3):267–273, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006835.

Fang:2007:FGP

[FMD07] Bin Fang, Glenn Martyna, and Yuefan Deng. A fine grained parallel smooth particle mesh Ewald algorithm for biophysical simulation studies: Application to the 6-D torus QCDOC supercomputer. Computer Physics Communications, 177(4):362–377, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507002445.

Filippov:2000:CVF

[FMG00] V. M. Filippov, S. P. Mikailovna, and Yake Gondo. Construction of variational factors for quasilinear second order partial differential equations. Computer Physics Communications, 126(1–2):67–71, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004245.

Fredericks:2008:CSG

[FMMQ08] E. Fredericks, F. M. Mahomed, E. Momoniat, and A. Qadir. Constructing a space from the geodesic equations. Computer Physics Communications, 179(6):438–442, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001380.

Furukawa:2001:MCA

[FMN01] Nobuo Furukawa, Yukitoshi Motome, and Hisaho Nakata. Monte Carlo algorithm for the double exchange model optimized for parallel computations. Computer Physics Communications, 142(1–3):410–413, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003769.

Ferrario:2005:P

[FMP05] Mauro Ferrario, Simone Melchionna, and Carlo Pierleoni. Preface. Computer Physics Communications, 169(1–3):xi–xii, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000986.

Fartaria:2006:TSA

[FNR^{+}06] Rui P. S. Fartaria, Rodrigo S. Neves, Pedro C. R. Rodrigues, Filomena F. M. Freitas, and Fernando M. S. Silva Fernandes. A time saving algorithm for the Monte Carlo method of Metropolis. Computer Physics Communications, 175(2):116– 121, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506001147. See erra-

tum [FNR^{+}07].

Fartaria:2007:ETS

[FNR^{+}07] Rui P. S. Fartaria, Rodrigo S. Neves, Pedro C. R. Rodrigues, Filomena F. M. Freitas, and Fernando M. S. Silva Fernandes. Erratum to “A time saving algorithm for the Monte Carlo method of Metropolis” [Computer Physics Communications 175 (2006) 116–121]. Computer Physics Communications, 176(3):250, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506004176. See [FNR^{+}06].

Fodor:2005:LGT

[Fod05] Zoltán Fodor. Lattice gauge theory and elementary particle physics. Computer Physics Communications, 169(1–3):347– 352, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550500175X.

Francisco:2008:ECE

[FPB08] E. Francisco, A. Martín Pendás, and M. A. Blanco. EDF: Computing electron number probability distribution functions in real space from molecular wave functions. Computer Physics Communications, 178(8):621–634, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004882.

Franco:2002:RKN

[Fra02] J. M. Franco. Runge–Kutta–Nyström methods adapted to the numerical integration of perturbed oscillators. Computer Physics Communications, 147(3):770–787, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004605.

Fragner:2007:URC

[Fra07a] Heinrich Fragner. Usage of a reconfigurable computer to simulate multiparticle systems. Computer Physics Communications, 176(5):327–333, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003985.

Franco:2007:EFS

[Fra07b] J. M. Franco. Exponentially fitted symplectic integrators of RKN type for solving oscillatory problems. Computer Physics Communications, 177(6):479–492, September 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002433.

Farias:2009:NSN

[FRdS09] R. L. S. Farias, Rudnei O. Ramos, and L. A. da Silva. Numerical solutions for non-Markovian stochastic equations of motion. Computer Physics Communications, 180(4):574–579, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004220.

Fritzsche:2001:URP

[Fri01] S. Fritzsche. Utilities for the ratip package. Computer Physics Communications, 141(1):163–174, November 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004003.

Fritzsche:2009:MPC

[Fri09] S. Fritzsche. Maple procedures for the coupling of angular momenta. An up-date of the Racah module. Computer Physics Communications, 180(10):2021–2023, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001969.

FroeseFischer:2000:MAS

[Fro00] Charlotte Froese Fischer. The MCHF atomic-structure package. Computer Physics Communications, 128(3):635–636, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000096.

Fruhwirth:2003:GMA

[Frü03] R. Frühwirth. A Gaussian-mixture approximation of the Bethe–Heitler model of electron energy loss by bremsstrahlung. Computer Physics Communications, 154(2):131–142, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002923.

Fitzgerald:2000:SPC

[FS00] Rosa M. Fitzgerald and Granville Sewell. Solving problems in computational physics using a general-purpose PDE solver. Computer Physics Communications, 124(2–3):132–138, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559900435X.

Frankcombe:2001:TEU

[FS01a] Terry J. Frankcombe and Sean C. Smith. Time evolution in the unimolecular master equation at low temperatures: full spectral solution with scalable iterative methods and high precision. Computer Physics Communications, 141(1):39–54, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002983.

Fujiwara:2001:MDS

[FS01b] Susumu Fujiwara and Tetsuya Sato. Molecular dynamics study of structure formation of a single polymer chain by cooling. Computer Physics Communications, 142(1–3):123–126, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003496.

Fujiwara:2002:MDS

[FS02] Susumu Fujiwara and Tetsuya Sato. Molecular dynamics simulation of a single polymer chain in vacuum and in solution. Computer Physics Communications, 147(1–2):342–345, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502003004.

Filbet:2003:CEV

[FS03] F. Filbet and E. Sonnendrücker. Comparison of Eulerian Vlasov solvers. Computer Physics Communications, 150(3): 247–266, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550200694X.

Fujimoto:2008:EPC

[FS08] Keizo Fujimoto and Richard D. Sydora. Electromagnetic particle-in-cell simulations on magnetic reconnection with adaptive mesh refinement. Computer Physics Communications, 178(12):915–923, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000908.

Friedrich:2009:ECC

[FSB09] Christoph Friedrich, Arno Schindlmayr, and Stefan Blügel. Efficient calculation of the Coulomb matrix and its expansion around k = 0 within the FLAPW method. Computer Physics Communications, 180(3):347–359, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003664.

Fletcher:2000:DDI

[FSBG00] Graham D. Fletcher, Michael W. Schmidt, Brett M. Bode, and Mark S. Gordon. The distributed data interface in GAMESS. Computer Physics Communications, 128(1–2):190–200, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000734.

Fischer:2004:SAQ

[FSK04] R. Fischer, A. Staudt, and C. H. Keitel. Simulation of atomic quantum dynamics in combined intense laser and weak electric fields. Computer Physics Communications, 157(2):139–146, February 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503005174.

Fang:2008:TFE

[FSW08] Yonglei Fang, Yongzhong Song, and Xinyuan Wu. Trigonometrically fitted explicit Numerov-type method for periodic IVPs with two frequencies. Computer Physics Communications, 179(11):801–811, December 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002488.

Felder:2008:LPL

[FT08] Gary Felder and Igor Tkachev. LATTICEEASY: a program for lattice simulations of scalar fields in an expanding universe. Computer Physics Communications, 178(12):929–932, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550800091X.

FroeseFischer:2007:MAS

[FTGG07] Charlotte Froese Fischer, Georgio Tachiev, Gediminas Gaigalas, and Michel R. Godefroid. An MCHF atomic-structure package for large-scale calculations. Computer Physics Communications, 176(8):559–579, April 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000446.

Frontera:2002:AGS

[FV02] Carlos Frontera and Eduard Vives. ”avalanches” in the ground state of the 3D Gaussian random field Ising model driven by an external field. Computer Physics Communications, 147(1–2):455–458, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003284.

Falloon:2001:EWP

[FW01] P. E. Falloon and J. B. Wang. Electronic wave propagation with Mathematica. Computer Physics Communications, 134(2):167–182, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550000196X.

Frlez:2001:OGP

[FWP01] E. Frlez, B. K. Wright, and D. Pocanić. optics: generalpurpose scintillator light response simulation code. Computer Physics Communications, 134(1):110–135, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001934.

FroeseFischer:2009:BSG

[FZ09] Charlotte Froese Fischer and Oleg Zatsarinny. A Bspline Galerkin method for the Dirac equation. Computer Physics Communications, 180(6):879–886, June 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800427X.

Galvanetto:2000:NCL

[Gal00] Ugo Galvanetto. Numerical computation of Lyapunov exponents in discontinuous maps implicitly defined. Computer Physics Communications, 131(1–2):1–9, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000552.

Gao:2003:LSP

[Gao03] Shiwu Gao. Linear-scaling parallelization of the WIEN package with MPI. Computer Physics Communications, 153(2):190– 198, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002248.

Gallo:2005:GTC

[GAR05] Paola Gallo, Andrea Attili, and Mauro Rovere. Glass transition in confinement: a Lennard-Jones binary mixture study. Computer Physics Communications, 169(1–3):214–217, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001451.

Gheller:2005:MCF

[GB05] Claudio Gheller and Gianfranco Brunetti. Merging clusters and the formation of radio haloes. Computer Physics Communications, 169(1–3):378–381, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001815.

Golosio:2001:NMA

[GBA01] B. Golosio, A. Brunetti, and S. R. Amendolia. A novel morphological approach to volume extraction in 3D tomography. Computer Physics Communications, 141(2):217–224, November 30, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003988.

Grasso:2004:NSR

[GBC^{+}04] D. Grasso, D. Borgogno, F. Califano, D. Farina, F. Pegoraro, and F. Porcelli. Numerical solution of a reduced model of collisionless magnetic reconnection in two and three dimensions. Computer Physics Communications, 164(1–3):23–28, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002516.

Goedecker:2003:EDF

[GBD03] Stefan Goedecker, Mireille Boulet, and Thierry Deutsch. An efficient 3-dim FFT for plane wave electronic structure calculations on massively parallel machines composed of multiprocessor nodes. Computer Physics Communications, 154 (2):105–110, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550300287X.

Gargate:2007:DMP

[GBFS07] L. Gargaté, R. Bingham, R. A. Fonseca, and L. O. Silva. dHybrid: a massively parallel code for hybrid simulations of space plasmas. Computer Physics Communications, 176(6):419–425, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506004309.

Gay-Balmaz:2002:LCF

[GBM02] Philippe Gay-Balmaz and Olivier J. F. Martin. A library for computing the filtered and non-filtered 3D Green’s tensor associated with infinite homogeneous space and surfaces. Computer Physics Communications, 144(1):111–120, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004714.

Grossu:2009:VTE

[GBR^{+}09] I. V. Grossu, C. Besliu, M. V. Rusu, Al. Jipa, C. C. Bordeianu, and D. Felea. Visual tool for estimating the fractal dimension of images. Computer Physics Communications, 180(10):1999–2001, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001635.

Gillan:2007:OFP

[GBTM07] M. J. Gillan, D. R. Bowler, A. S. Torralba, and T. Miyazaki. Order-N first-principles calculations with the CONQUEST code. Computer Physics Communications, 177(1–2):14–18, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001658.

Ghiselli:2001:DAM

[GC01] A. Ghiselli and The MONARC Collaboration. Distributed applications monitoring at system and network level. Computer Physics Communications, 140(1–2):219–225, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002806.

Gutierrez-Castrejon:2006:ULA

[GCD06] R. Gutiérrez-Castrejón and M. Duelk. Using LabVIEW for advanced nonlinear optoelectronic device simulations in high-speed optical communications. Computer Physics Communications, 174(6):431–440, March 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005886.

Group:2001:P

[GCI01] CCP3 Working Group, Simon Crampin, and John Inglesfield. Preface. Computer Physics Communications, 137(1):xi, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001667.

George:2002:SEQ

[GCK02] Jeffrey A. George, Michael E. Colvin, and V. V. Krishnan. A simulator for ensemble quantum computing. Computer Physics Communications, 144(3):277–283, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002667.

Galindez:2002:DMR

[GCP^{+}02] J. Galindez, F. Calvo, P. Paska, D. Hrivnak, R. Kalus, and F. X. Gadéa. DIM modelings for realistic simulations of ionic rare-gas clusters, test on structures and photoabsorption spectra of Ar_{n}^{+} (n = 3–8 ). Computer Physics Communications, 145(1):126–140, May 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001509.

Gunduc:2005:ESD

[GDAG05a] Semra Gündüç, Mehmet Dilaver, Meral Aydin, and Yigit Gündüç. Erratum to “A study of dynamic finite size scaling behavior of the scaling functions-calculation of dynamic critical index of Wolff algorithm” [Computer Physics Communications 166 (2005) 1–7]. Computer Physics Communications, 172(3):214, November 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505004340. See [GDAG05b].

Gunduc:2005:SDF

[GDAG05b] Semra Gündüç, Mehmet Dilaver, Meral Aydin, and Yigit Gündüç. A study of dynamic finite size scaling behavior of the scaling functions-calculation of dynamic critical index of Wolff algorithm. Computer Physics Communications, 166(1):1–7, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004631. See erra-

tum [GDAG05a].

Gaspar:2001:DPL

[GDC01] C. Gaspar, M. Dönszelmann, and Ph. Charpentier. DIM, a portable, light weight package for information publishing, data transfer and inter-process communication. Computer Physics Communications, 140(1–2):102–109, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002600.

Genchev:2001:ECN

[Gen01] Z. D. Genchev. Effect of colored noise on an overdamped Josephson junction. Computer Physics Communications, 135(1):35–39, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500002113.

Gervasio:2007:CTM

[Ger07] Francesco Luigi Gervasio. Charge transfer mechanism in a polydgpdcp fiber and in wet DNA. Computer Physics Communications, 177(1–2):27–29, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001099.

Gaigalas:2001:CRC

[GF01] Gediminas Gaigalas and Stephan Fritzsche. Calculation of reduced coefficients and matrix elements in jj-coupling. Computer Physics Communications, 134(1):86–96, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001764.

Gaigalas:2002:MPC

[GF02a] G. Gaigalas and S. Fritzsche. Maple procedures for the coupling of angular momenta. VI. LS–jj transformations. Computer Physics Communications, 149(1):39–60, November 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200591X.

Gaigalas:2002:PSA

[GF02b] G. Gaigalas and S. Fritzsche. Pure spin-angular momentum coefficients for non-scalar one-particle operators in jj-coupling. Computer Physics Communications, 148(3):349–351, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005891.

Gianturco:2002:SDS

[GF02c]  F. A. Gianturco and Francesco Filippone.  Structure

and dynamics of small protonated rare-gas clusters using quantum and classical methods. Computer Physics Communications, 145(1):78–96, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502001480.

Gaigalas:2001:MPC

[GFF01] G. Gaigalas, S. Fritzsche, and B. Fricke. Maple procedures for the coupling of angular momenta. III. Standard quantities for evaluating many-particle matrix elements. Computer Physics Communications, 135(2):219–237, April 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002344.

Gaigalas:2001:PCP

[GFG01] Gediminas Gaigalas, Stephan Fritzsche, and Ian P. Grant. Program to calculate pure angular momentum coefficients in jjcoupling. Computer Physics Communications, 139(3):263–278, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002132.

Gaigalas:2006:JPR

[GFG^{+}06] G. Gaigalas, S. Fritzsche, E. Gaidamauskas, G. Kirsanskas, and T. Zalandauskas. Jahn — a program for representing atomic and nuclear states within an isospin basis. Computer Physics Communications, 175(1):52–66, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001093.

Garcia:2000:EEA

[GFP00] S. García, F. Fuentes, and C. Paz. Explicit electromagnetic algorithm for 2D using a multi-fluid model in laser-produced plasmas. Computer Physics Communications, 127(2–3):271– 285, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005214.

Gygi:2003:CML

[GFS03] François Gygi, Jean-Luc Fattebert, and Eric Schwegler. Computation of maximally localized Wannier functions using a simultaneous diagonalization algorithm. Computer Physics Communications, 155(1):1–6, September 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003151.

Gubernatis:2000:RWB

[GG00] J. E. Gubernatis and M. Guerrero. Random walk beyond Hartree–Fock. Computer Physics Communications, 128(1– 2):201–209, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599005287.

Grum-Grzhimailo:2003:MPC

[GG03] A. N. Grum-Grzhimailo. MJK: a program to calculate observable quantities in electron-atom collisions. Computer Physics Communications, 152(1):101–119, April 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200797X.

Garcia-Gonzalez:2001:GSE

[GGG01] P. García-González and R. W. Godby. GW self-energy calculations for surfaces and interfaces. Computer Physics Communications, 137(1):108–122, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001746.

Glotzer:2002:GFL

[GGL^{+}02] S. C. Glotzer, Y. Gebremichael, N. Lacevic, T. B. Schrøder, and F. W. Starr. Glass-forming liquids and polymers: with a little help from computational statistical physics. Computer Physics Communications, 146(1):24–29, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004319.

Gavai:2003:SAO

[GGL03] Rajiv V. Gavai, Sourendu Gupta, and Robert Lacaze. Speed and adaptability of overlap fermion algorithms. Computer Physics Communications, 154(2):143–158, August 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002972.

Gaines:2001:IOO

[GGQ01] Irwin Gaines, Saul Gonzalez, and Sijin Qian. Implementation of an object oriented track reconstruction model into multiple LHC experiments. Computer Physics Communications, 140(1–2):21–30, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002521.

Guagnelli:2000:SPS

[GH00] Marco Guagnelli and Jochen Heitger. SSOR preconditioning in simulations of the QCD Schrödinger functional. Computer Physics Communications, 130(1–2):12–21, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000126.

Group:2001:CSN

[GH01] The Complexity Simulation Group and Takaya Hayashi. Computer simulation of nonlinear phenomena in plasmas. Computer Physics Communications, 142(1–3):49–54, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003149.

Ghani:2005:TCL

[Gha05] B. Abdul Ghani. TEA CO_{2} Laser Simulator: a software tool to predict the output pulse characteristics of TEA CO_{2} laser. Computer Physics Communications, 171(2):93–106, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002894.

Gray:2009:SMP

[GHIL09] James Gray, Yang-Hui He, Anton Ilderton, and André Lukas. STRINGVACUA. A Mathematica package for studying vacuum configurations in string phenomenology. Computer Physics Communications, 180(1):107–119, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002841.

Giusti:2003:NTL

[GHLW03] L. Giusti, C. Hoelbling, M. Lüscher, and H. Wittig. Numerical techniques for lattice QCD in the ϵ-regime. Computer Physics Communications, 153(1):31–51, June 1, 2003.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008743.

Gan:2001:PCG

[GHP01] C. K. Gan, P. D. Haynes, and M. C. Payne. Preconditioned conjugate gradient method for the sparse generalized eigenvalue problem in electronic structure calculations. Computer Physics Communications, 134(1):33–40, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001880.

Gutnic:2004:VSA

[GHPS04] M. Gutnic, M. Haefele, I. Paun, and E. Sonnendrücker. Vlasov simulations on an adaptive phase-space grid. Computer Physics Communications, 164(1–3):214–219, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002838.

Grimley:2001:CCP

[GI01] T. B. Grimley and J. E. Inglesfield. The Collaborative Computational Project on surfaces: CCP3. Computer Physics Communications, 137(1):1–3, June 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001679.

Gao:2009:MPV

[GI09] Min Gao and Masaaki Ito. A Maple package for verifying ultradiscrete soliton solutions. Computer Physics Communications, 180(7):1196–1205, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550900023X.

Gianturco:2002:P

[Gia02] Franco A. Gianturco. Preface. Computer Physics Communications, 145(1):xi–xii, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502001443.

Galkin:2002:MFM

[GIME02] S. A. Galkin, A. A. Ivanov, S. Yu. Medvedev, and A. G. Elfimov. Multi-fluid MHD model and calculations of Alfvén wave spectrum and dissipation in tokamaks. Computer Physics Communications, 143(1):29–47, February 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004143.

Grigoriev:2003:OJF

[GJT03] D. Yu. Grigoriev, E. Jankowski, and F. V. Tkachov. Optimal jet finder. Computer Physics Communications, 155(1): 42–64, September 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002911.

Gazizov:2005:AHE

[GK05] A. Gazizov and M. Kowalski. ANIS: High energy neutrino generator for neutrino telescopes. Computer Physics Communications, 172(3):203–213, November 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004194.

Gao:2002:RPF

[GKI02] Min Gao, Yasuyuki Kato, and Masaaki Ito. A REDUCE package for finding conserved densities of systems of nonlinear difference-difference equations. Computer Physics Communications, 148(2):242–255, October 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005568.

Gao:2004:RPF

[GKI04] Min Gao, Yasuyuki Kato, and Masaaki Ito. A REDUCE package for finding conserved densities of systems of implicit difference-difference equations. Computer Physics Communications, 160(1):69–89, June 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504001122.

Gorbunov:2008:FSK

[GKK^{+}08] S. Gorbunov, U. Kebschull, I. Kisel, V. Lindenstruth, and W. F. J. Müller. Fast SIMDized Kalman filter based track fit. Computer Physics Communications, 178(5):374–383, March 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004262.

Golutvin:2000:RET

[GKM^{+}00] I. A. Golutvin, Y. T. Kiriouchine, S. A. Movchan, G. A. Ososkov, V. V. Palichik, and E. A. Tikhonenko. Robust estimates of track parameters and spatial resolution for CMS muon chambers. Computer Physics Communications, 126(1–2):72– 76, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002490.

Golonka:2006:TPE

[GKP^{+}06] P. Golonka, B. Kersevan, T. Pierzchala, E. Richter-Was, Z. Was, and M. Worek. The tauola-photos-F environment for the TAUOLA and PHOTOS packages, release II. Computer Physics Communications, 174(10):818–835, May 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506000403.

Gluza:2007:AMP

[GKR07] J. Gluza, K. Kajda, and T. Riemann. AMBRE — a Mathematica package for the construction of Mellin–Barnes representations for Feynman integrals. Computer Physics Communications, 177(11):879–893, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003505.

Gattringer:2002:IDO

[GL02]  Christof Gattringer and C. B. Lang.  Improving the

Dirac operator in lattice QCD. Computer Physics Communications, 147(1–2):398–401, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003132.

Golubov:2001:GPC

[GLHW01] S. I. Golubov, Xiangli Liu, Hanchen Huang, and C. H. Woo. GFCUBHEX: Program to calculate elastic Green’s functions and displacement fields for applications in atomistic simulations of defects in cubic and HCP crystals. Computer Physics Communications, 137(2):312–324, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001576.

Gnani:2002:ECH

[GLL^{+}02] M. Gnani, P. Lazzeretti, A. Ligabue, U. Pincelli, and A. Soncini. An efficient coupled Hartree–Fock computational scheme for parity-violating energy differences in enantiomeric molecules. Computer Physics Communications, 144(2):130– 140, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200156X.

Gonzalez-Lezana:2002:VMT

[GLMADB^{+}02] T. González-Lezana, S. Miret-Artés, G. Delgado-Barrio, P. Villarreal, J. Rubayo-Soneira, I. Baccarelli, F. Paesani, and F. A. Gianturco. A variational method to treat diffuse states in weakly bound trimers. Computer Physics Communications, 145(1):156–183, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502001522.

Glowinski:2003:OSM

[GLP03] R. Glowinski, P. Lin, and X.-B. Pan. An operatorsplitting method for a liquid crystal model. Computer Physics Communications, 152(3):242–252, May 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008238.

Guan:2003:CSD

[GLW03] Shuguang Guan, C.-H. Lai, and G. W. Wei. Characterizing the spatiotemporal dynamics of turbulence. Computer Physics Communications, 155(2):77–91, October 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002960.

Guillaume:2000:CXD

[GM00] Damien Guillaume and Fionn Murtagh. Clustering of XML documents. Computer Physics Communications, 127(2–3):215– 227, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005111.

Garcia-March:2009:GFD

[GMAHV^{+}09] Miguel A. García-March, Miguel Arevalillo-Herráez, Francisco R. Villatoro, Fernando Giménez, and Pedro Fernández de Córdoba. A generalized finite difference method using Coatmèlec lattices. Computer Physics Communications, 180 (7):1125–1133, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000319.

Garcia:2007:SCS

[GMAN^{+}07] F. Garcia, J. Mesa, J. D. T. Arruda-Neto, O. Helene, V. Vanin, F. Milian, A. Deppman, T. E. Rodrigues, and O. Rodriguez. The STATFLUX code: a statistical method for calculation of flow and set of parameters, based on the Multiple-Compartment Biokinetical Model. Computer Physics Communications, 176(5):347–361, March 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003924.

Gomez-Munoz:2008:CEC

[GMBC08] José Luis Gómez-Muñoz and Julián Bravo-Castillero. Calculation of effective conductivity of 2D and 3D composite materials with anisotropic constituents and different inclusion shapes in Mathematica. Computer Physics Communications, 179(4):275–287, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000799.

Guseinov:2006:EAE

[GME06] I. I. Guseinov, B. A. Mamedov, and A. S. Ekenoglu. Exact analytical expressions and numerical analysis of twocenter Franck–Condon factors and matrix elements over displaced harmonic oscillator wave functions. Computer Physics Communications, 175(3):226–231, August 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001603.

Gutleber:2003:THA

[GMO03] Johannes Gutleber, Steven Murray, and Luciano Orsini. Towards a homogeneous architecture for high-energy physics data acquisition systems. Computer Physics Communications, 153(2):155–163, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503001619.

Guan:2009:AAL

[GNZ^{+}09] Xiaoxu Guan, C. J. Noble, O. Zatsarinny, K. Bartschat, and B. I. Schneider. ALTDSE: An Arnoldi–Lanczos program to solve the time-dependent Schrödinger equation. Computer Physics Communications, 180(12):2401–2409, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001003.

Goc:2004:CCV

[Goc04] Roman Goc. Computer calculation of the Van Vleck second moment for materials with internal rotation of spin groups. Computer Physics Communications, 162(2):102–112, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003315.

Goedecker:2002:OPF

[Goe02] Stefan Goedecker. Optimization and parallelization of a force field for silicon using OpenMP. Computer Physics Communications, 148(1):124–135, October 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004666.

Guerrero:2000:SST

[GOG00] M. Guerrero, G. Ortiz, and J. E. Gubernatis. Search for superconductivity in the two-dimensional Hubbard model using the constrained path Monte Carlo. Computer Physics Communications, 127(1):143–149, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000242.

Glass:2006:UEC

[GOH06] Colin W. Glass, Artem R. Oganov, and Nikolaus Hansen. USPEX — evolutionary crystal structure prediction. Computer Physics Communications, 175(11–12):713–720, December 1–15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002931.

Goldfield:2000:PSF

[Gol00] E. M. Goldfield. Parallel strategies for four-atom quantum dynamics calculations. Computer Physics Communications, 128(1–2):178–189, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000631.

Gottlieb:2001:CEC

[Got01] Steven Gottlieb. Cost-effective clustering. Computer Physics Communications, 142(1–3):43–48, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003125.

Gould:2000:CPU

[Gou00] Harvey Gould. Computational physics and the undergraduate curriculum. Computer Physics Communications, 127(1):6–10, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000345.

Gospavic:2008:BED

[GPT08] Radovan Gospavic, Viktor Popov, and Goran Todorovic. Boundary element-dual reciprocity formulation for bound electron states in semiconductor quantum wires. Computer Physics Communications, 178(5):366–373, March 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004250.

Golonka:2004:MTU

[GPW04] P. Golonka, T. Pierzchala, and Z. Was. MC-TESTER: a universal tool for comparisons of Monte Carlo predictions for particle decays in high energy physics. Computer Physics Communications, 157(1):39–62, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004661.

Gothandaraman:2009:HAQ

[GPW^{+}09] Akila Gothandaraman, Gregory D. Peterson, G. Lee Warren, Robert J. Hinde, and Robert J. Harrison. A Hardware-Accelerated Quantum Monte Carlo framework (HAQMC) for N-body systems. Computer Physics Communications, 180 (12):2563–2573, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002070.

Gehrmann:2001:NEH

[GR01] T. Gehrmann and E. Remiddi. Numerical evaluation of harmonic polylogarithms. Computer Physics Communications, 141(2):296–312, November 30, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004118.

Gehrmann:2002:NET

[GR02] T. Gehrmann and E. Remiddi. Numerical evaluation of two-dimensional harmonic polylogarithms. Computer Physics Communications, 144(2):200–223, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550200139X.

Grassberger:2002:GWG

[Gra02] Peter Grassberger. Go with the winners: a general Monte Carlo strategy. Computer Physics Communications, 147(1–2):64–70, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002059.

Greenwald:2004:BBH

[Gre04] Martin Greenwald. Beyond benchmarking-how experiments and simulations can work together in plasma physics. Computer Physics Communications, 164(1–3):1–8, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002486.

Grebosz:2007:CCI

[Gre07] Jerzy Grebosz. The Cracow code — an interactive method of sophisticated online analysis. Computer Physics Communications, 176(4):251–265, February 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003675.

Grodzicky:2001:CPW

[Gro01] Roman Grodzicky. A computer program written in Mathematica for calculating H_{2} quasicrystals and their diffraction patterns. Computer Physics Communications, 136(3):236–249, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002484.

Gates:2001:IDP

[GRR01] S. James Gates, Jr., B. Radak, and V. G. J. Rodgers. Irreducible decomposition of products of 10D chiral sigma matrices. Computer Physics Communications, 136(1–2):173–181, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002514.

Guo:2006:FBM

[GRS06] Baozeng Guo, Umberto Ravaioli, and Maritin Staedele. Full band Monte Carlo calculations of velocity-field characteristics of wurtzite ZnO. Computer Physics Communications, 175(7):482–486, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002694.

Gil:2001:CMC

[GS01a] Amparo Gil and Javier Segura. CA3D: a Monte Carlo code to simulate 3D buffered diffusion of ions in sub-membrane domains. Computer Physics Communications, 136(3):269–293, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001448.

Gil:2001:DNV

[GS01b] Amparo Gil and Javier Segura. DTORH3 2.0: a new version of a computer program for the evaluation of toroidal harmonics. Computer Physics Communications, 139(2):186–191, September 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001886.

Gaigalas:2005:MPC

[GSF05] G. Gaigalas, O. Scharf, and S. Fritzsche. Maple procedures for the coupling of angular momenta. VIII. Spinangular coefficients for single-shell configurations. Computer Physics Communications, 166(2):141–169, March 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005582.

Gaigalas:2006:HSP

[GSF06] G. Gaigalas, O. Scharf, and S. Fritzsche. Hyperfine structure parametrisation in Maple. Computer Physics Communications, 174(3):202–221, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005709.

Gousheh:2003:MSS

[GSGT03] S. S. Gousheh, H. R. Sepangi, and K. Ghafoori-Tabrizi. A method for solving systems of non-linear differential equations with moving singularities. Computer Physics Communications, 149(3):135–141, January 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502006306.

Galvan:2003:AMP

[GSM^{+}03] I. Fdez Galván, M. L. Sánchez, M. E. Martín, F. J. Olivares del Valle, and M. A. Aguilar. ASEP/MD: a program for the calculation of solvent effects combining QM/ MM methods and the mean field approximation. Computer Physics Communications, 155(3):244–259, November 1– 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003515.

Gillan:2006:CZA

[GSS06] C. J. Gillan, A. Schuchinsky, and I. Spence. Computing zeros of analytic functions in the complex plane without using derivatives. Computer Physics Communications, 175(4):304–313, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506001949.

Gunnella:2000:MFM

[GSSN00] R. Gunnella, F. Solal, D. Sébilleau, and C. R. Natoli. MSPHD: a full multiple scattering code for low energy photoelectron diffraction. Computer Physics Communications, 132(3):251–266, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001454.

Gao:2001:GHF

[GT01] Yi-Tian Gao and Bo Tian. Generalized hyperbolic-function method with computerized symbolic computation to construct the solitonic solutions to nonlinear equations of mathematical physics. Computer Physics Communications, 133 (2–3):158–164, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001685.

Glasser:2004:SMM

[GT04] A. H. Glasser and X. Z. Tang. The SEL macroscopic modeling code. Computer Physics Communications, 164(1–3): 237–243, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002863.

Gunton:2002:OCS

[Gun02] J. D. Gunton. Opportunities for computational science in materials science. Computer Physics Communications, 146(1):131, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200454X.

Gutle:2006:AEI

[Gut06] Claudine Gutlé. Analytic exchange integral for coupled cluster theory in the homogeneous electron gas. Computer Physics Communications, 174(10):836–845, May 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000129.

Galeriu:2004:MNS

[GVMW04] C. Galeriu, L. C. Lew Yan Voon, R. Melnik, and M. Willatzen. Modeling a nanowire superlattice using the finite difference method in cylindrical polar coordinates. Computer Physics Communications, 157(2):147–159, February 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004934.

Goossens:2001:MCS

[GW01a] D. J. Goossens and T. R. Welberry. Monte Carlo study of disorder in HMTA. Computer Physics Communications, 142(1–3):387–390, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003897.

Guntert:2001:SCS

[GW01b] Peter Güntert and Kurt Wüthrich. Sampling of conformation space in torsion angle dynamics calculations. Computer Physics Communications, 138(2):155–169, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002041.

Gonnet:2009:SES

[GWK09] Pedro Gonnet, Jens H. Walther, and Petros Koumoutsakos. θ-SHAKE: An extension to SHAKE for the explicit treatment of angular constraints. Computer Physics Communications, 180(3):360–364, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003676.

Gressel:2007:SIC

[GZ07] O. Gressel and U. Ziegler. Shearingbox-implementation for the central-upwind, constraint-transport MHD-code NIRVANA. Computer Physics Communications, 176(11–12):652–659, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001865.

Gong:2001:GFD

[GZDA01] Zhigang Gong, Ladislav Zejda, Werner Däppen, and Josep M. Aparicio. Generalized Fermi–Dirac functions and derivatives: properties and evaluation. Computer Physics Communications, 136(3):294–309, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550100145X.

Gaigalas:2004:SLN

[GZF04] G. Gaigalas, T. Zalandauskas, and S. Fritzsche. Spectroscopic LSJ notation for atomic levels obtained from relativistic calculations. Computer Physics Communications, 157(3):239–253, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503005186.

Herrmann:2007:PF

[HAA07] Hans J. Herrmann, José S. Andrade, Jr., and Ascânio D. Araújo. Particles in fluids. Computer Physics Communications, 177(1–2):158–161, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000884.

Hafner:2007:MSU

[Haf07] J. Hafner. Materials simulations using VASP-a quantum perspective to materials science. Computer Physics Communications, 177(1–2):6–13, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550700080X.

Hahn:2001:GFD

[Hah01] Thomas Hahn. Generating Feynman diagrams and amplitudes with FeynArts 3. Computer Physics Communications, 140(3):418–431, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002909.

Hahn:2005:CLM

[Hah05] T. Hahn. Cuba — a library for multidimensional numerical integration. Computer Physics Communications, 168(2):78–95, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000792.

Hahn:2007:CLM

[Hah07] Thomas Hahn. Cuba — a library for multidimensional numerical integration. Computer Physics Communications, 176 (11–12):712–713, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002056.

Hahn:2008:MIF

[Hah08] T. Hahn. A Mathematica interface for FormCalc-generated code. Computer Physics Communications, 178(3):217–221, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004201.

Hahn:2009:SHA

[Hah09] T. Hahn. SUSY Les Houches Accord 2 I/O made easy. Computer Physics Communications, 180(9):1681–1693, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001106.

Hanisch:2000:IAD

[Han00] R. J. Hanisch. Integrated access to distributed data and information services in astrophysics and the space sciences. Computer Physics Communications, 127(2–3):177–187, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559900507X.

Harris:2000:ARD

[Har00] Frank E. Harris. Algebraic reduction in discrete light-cone quantized electrodynamics using Maple V. Computer Physics Communications, 132(1–2):21–29, October 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001338.

Harrison:2001:FPS

[Har01] N. M. Harrison. First principles simulation of surfaces and interfaces. Computer Physics Communications, 137(1):59–73, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001710.

Harris:2002:CES

[Har02] Frank E. Harris. Comments on “Ewald summation technique for one-dimensional charge distributions”. Computer Physics Communications, 146(2):271–273, July 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004526. See [LHC01, LHC02].

Hoang-Binh:2005:PCE

[HB05] D. Hoang-Binh. A program to compute exact hydrogenic radial integrals, oscillator strengths, and Einstein coefficients, for principal quantum numbers up to n ≈ 1000. Computer Physics Communications, 166(3):191–196, March 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005661.

Hilfer:2005:MST

[HBMJ05] Rudolf Hilfer, Bibudhananda Biswal, Hans-Georg Mattutis, and Wolfhard Janke. Multicanonical simulations of the tails of the order-parameter distribution of the two-dimensional Ising model. Computer Physics Communications, 169(1–3):230–233, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001505.

Hawke:2005:GWC

[HBRS05] Ian Hawke, Luca Baiotti, Luciano Rezzolla, and Erik Schnetter. Gravitational waves from the 3D collapse of a neutron star to a Kerr black hole. Computer Physics Communications, 169(1–3):374–377, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001803.

Hartmann:2005:PTF

[HBW05] Alexander K. Hartmann, Wolfgang Barthel, and Martin Weigt. Phase transition and finite-size scaling in the vertex-cover problem. Computer Physics Communications, 169(1–3):234–237, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001517.

Haynes:2000:PFF

[HC00] Peter D. Haynes and Michel Côté. Parallel fast Fourier transforms for electronic structure calculations. Computer Physics Communications, 130(1–2):130–136, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000497.

He:2008:RWP

[HC08] Sheng-Gui He and Dennis J. Clouthier. RT3: a Windows program for the Renner–Teller analysis of ^{2}Π states of triatomic molecules. Computer Physics Communications, 178(9):676– 684, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004894.

Hsu:2006:DPP

[HCH^{+}06] K.-H. Hsu, P.-Y. Chen, C.-T. Hung, L.-H. Chen, and J.-S. Wu. Development of a parallel Poisson’s equation solver with adaptive mesh refinement and its application in field emission prediction. Computer Physics Communications, 174(12):948–960, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506000567.

Hu:2000:RDM

[HCIK00] Chin-Kun Hu, Jau-Ann Chen, N. Sh. Izmailian, and P. Kleban. Recent developments in the Monte Carlo approach to percolation problems. Computer Physics Communications, 126(1–2): 77–81, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599002386.

Hegland:2001:MMD

[HCK01] Markus Hegland, William Clarke, and Margaret Kahn. Mining the MACHO dataset. Computer Physics Communications, 142(1–3):22–28, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003071.

Hashimoto:2000:IRC

[HCO00] Yasuhiro Hashimoto, Yu Chen, and Hirotada Ohashi. Immiscible real-coded lattice gas. Computer Physics Communications, 129(1–3):56–62, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000928.

Hirabayashi:2001:LBA

[HCO01] Miki Hirabayashi, Yu Chen, and Hirotada Ohashi. A lattice Boltzmann analysis of the effects of the internal angular momentum on rheology of magnetic fluids. Computer Physics Communications, 142(1–3):148–150, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003551.

Hajiyev:2004:CRM

[HD04] E. Hajiyev and A. Demir. A collisional-radiative model for simulation of ne-like and F-like resonance lines emitted from laser produced plasmas. Computer Physics Communications, 164(1–3):86–90, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002619.

Homann:2007:IFP

[HDG07] Holger Homann, Jürgen Dreher, and Rainer Grauer. Impact of the floating-point precision and interpolation scheme on the results of DNS of turbulence by pseudo-spectral codes. Computer Physics Communications, 177(7):560–565, October 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507002998.

Han:2007:ETA

[HDGM07] Guowen Han, Yuefan Deng, James Glimm, and Glenn Martyna. Error and timing analysis of multiple time-step integration methods for molecular dynamics. Computer Physics Communications, 176(4):271–291, February 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003729.

Heffelfinger:2000:PAS

[Hef00] Grant S. Heffelfinger. Parallel atomistic simulations. Computer Physics Communications, 128(1–2):219–237, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000503.

Hein:2001:GTT

[Hei01] Steffen Hein. Gauge techniques in time and frequency domain TLM. Computer Physics Communications, 136(1–2):77–89, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001436.

Heermann:2000:TPV

[HF00] Dieter W. Heermann and Thomas T. Fuhrmann. Teaching physics in the virtual university: the mechanics toolkit. Computer Physics Communications, 127(1):11–15, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000333.

Holub:2006:QHA

[HF06] Oliver Holub and Sérgio T. Ferreira. Quantitative histogram analysis of images. Computer Physics Communications, 175(9):620–623, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002712.

Hoshina:2003:DGS

[HFN03] Kotoyo Hoshina, Keisuke Fujii, and Osamu Nitoh. Development of a Geant4 solid for stereo mini-jet cells in a cylindrical drift chamber. Computer Physics Communications, 153(3):373–391, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002066.

Hatano:2002:DDG

[HG02a] Naomichi Hatano and James E. Gubernatis. Double degeneracy in the ground state of the 3D ±J spin glass. Computer Physics Communications, 147(1–2):414–418, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003168.

Hayes:2002:EAP

[HG02b] J. M. Hayes and J. C. Greer. Extraction of analytical potential function parameters from ab initio potential energy surfaces and analytical forces. Computer Physics Communications, 147(3):803–825, September 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004563.

Homola:2005:SUH

[HGH^{+}05] P. Homola, D. Góra, D. Heck, H. Klages, J. Pekala, M. Risse, B. Wilczyńska, and H. Wilczyński. Simulation of ultra-high energy photon propagation in the geomagnetic field. Computer Physics Communications, 173(1–2):71–90, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505004212.

Hernandez:2002:TDM

[HGVCM^{+}02] M. I. Hernández, A. García-Vela, J. Campos-Martínez, O. Roncero, P. Villarreal, and G. Delgado-Barrio. Time-dependent methods to study the dissociation dynamics of tri- and tetraatomic clusters. Computer Physics Communications, 145 (1):97–125, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502001492.

Harting:2005:DTD

[HHCC05] Jens Harting, Matthew J. Harvey, Jonathan Chin, and Peter V. Coveney. Detection and tracking of defects in the gyroid mesophase. Computer Physics Communications, 165(2):97–109, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504004953.

Hahn:2009:FPC

[HHH^{+}09] T. Hahn, S. Heinemeyer, W. Hollik, H. Rzehak, and G. Weiglein. FeynHiggs: a program for the calculation of MSSM Higgs-boson observables — version 2.6.5. Computer Physics Communications, 180(8):1426–1427, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000691.

Husa:2006:KMP

[HHL06] Sascha Husa, Ian Hinder, and Christiane Lechner. Kranc: a Mathematica package to generate numerical codes for tensorial evolution equations. Computer Physics Communications, 174(12):983–1004, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506001020.

Hine:2009:LSD

[HHM^{+}09] N. D. M. Hine, P. D. Haynes, A. A. Mostofi, C.-K. Skylaris, and M. C. Payne. Linear-scaling density-functional theory with tens of thousands of atoms: Expanding the scope and scale of calculations with ONETEP. Computer Physics Communications, 180(7):1041–1053, July 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004414.

Heinemeyer:2000:FPC

[HHW00] S. Heinemeyer, W. Hollik, and G. Weiglein. FeynHiggs: a program for the calculation of the masses of the neutral CP-even Higgs bosons in the MSSM. Computer Physics Communications, 124(1):76–89, January 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599003641.

Hung:2007:NPS

[HHWH07] C.-T. Hung, M.-H. Hu, J.-S. Wu, and F.-N. Hwang. A new paradigm for solving plasma fluid modeling equations. Computer Physics Communications, 177(1–2):138–139, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700094X.

Hibbert:2001:CFF

[Hib01] A. Hibbert. Charlotte Froese Fischer — some personal reflections. Computer Physics Communications, 138(1):2–9, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002454.

Hinde:2000:CAM

[Hin00] Robert J. Hinde. Constructing atom-molecule potential surfaces from ab initio data: a method combining quadrature and interpolation. Computer Physics Communications, 130(1–2): 1–11, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005238.

Hellmund:2002:HTS

[HJ02] Meik Hellmund and Wolfhard Janke. High-temperature series expansions for random-bond Potts models on Z^{d}. Computer Physics Communications, 147(1–2):435–438, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003211.

Hoyrup:2002:NSS

[HJM02] Pernille Høyrup, Kent Jørgensen, and Ole G. Mouritsen. Nano-scale structure in membranes in relation to enzyme action-computer simulation vs. experiment. Computer Physics Communications, 147(1–2):313–320, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002941.

Huang:2009:FYA

[HJZ09] Jingfang Huang, Jun Jia, and Bo Zhang. FMM–Yukawa: An adaptive fast multipole method for screened Coulomb interactions. Computer Physics Communications, 180(11):2331–2338, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002082.

Hu:2007:PMA

[HJZL07] Xi-Wei Hu, Zhong-He Jiang, Shu Zhang, and Minghai Liu. The propagation of a microwave in an atmospheric pressure plasma layer: one- and two-dimensional numerical solutions. Computer Physics Communications, 177(1–2):76–77, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001178.

Hasenfratz:2002:SDF

[HK02] Anna Hasenfratz and Francesco Knechtli. Simulation of dynamical fermions with smeared links. Computer Physics Communications, 148(1):81–86, October 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004708.

Hayashi:2001:PBE

[HKK^{+}01] Keiji Hayashi, Takuo Kanayama, Seiji Kawakita, Syo Sunouchi, Mitsuhiro Takimoto, Yasunori Kawamura, and Akifumi Maeda. Phonon-band engineering as a novel approach to modify frictional characteristics of mesoscopic solid materials. Computer Physics Communications, 142(1–3):238–242, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003198.

Huhtala:2002:CNS

[HKK02a] Maria Huhtala, Antti Kuronen, and Kimmo Kaski. Carbon nanotube structures: molecular dynamics simulation at realistic limit. Computer Physics Communications, 146(1):30–37, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004320.

Huhtala:2002:CSC

[HKK02b] Maria Huhtala, Antti Kuronen, and Kimmo Kaski. Computational studies of carbon nanotube structures. Computer Physics Communications, 147(1–2):91–96, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002230.

Huber:2007:GGL

[HKL^{+}07a] Patrick Huber, Joachim Kopp, Manfred Lindner, Mark Rolinec, and Walter Winter. GLoBES: General Long Baseline Experiment Simulator. Computer Physics Communications, 177(5):439–440, September 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507002470.

Huber:2007:NFS

[HKL^{+}07b] Patrick Huber, Joachim Kopp, Manfred Lindner, Mark Rolinec, and Walter Winter. New features in the simulation of neutrino oscillation experiments with GLoBES 3.0: (general long baseline experiment simulator). Computer Physics Communications, 177(5):432–438, September 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002421.

Hong:2007:CBP

[HKLY07] Y. J. Hong, G. J. Kim, S. H. Lee, and M. Yoon. Comparison between particle-in-cell Monte–Carlo and fluid simulations of argon microhollow discharges. Computer Physics Communications, 177(1–2):128, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001026.

Hahm:2007:KRG

[HKM^{+}07] Jaegyoon Hahm, Oh-Kyoung Kwon, Hun Joo Myung, Hyoungwoo Park, and Jongsuk Ruth Lee. KMI-R1: a grid middleware package for K*grid. Computer Physics Communications, 177(1–2):215, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000495.

Hoffman:2002:RGN

[HKP02] David K. Hoffman, Donald J. Kouri, and Eli Pollak. Reducing Gaussian noise using distributed approximating functionals. Computer Physics Communications, 147(3):759–769, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004575.

Hong:2007:KPS

[HKPL07] Y. J. Hong, H. S. Ko, G. Y. Park, and J. K. Lee. Kinetic plasma simulations for three dielectric etchers. Computer Physics Communications, 177(1–2):122–123, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001191.

He:2000:LBS

[HL00a] Xiaoyi He and Ning Li. Lattice Boltzmann simulation of electrochemical systems. Computer Physics Communications, 129(1–3):158–166, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550000103X.

Hedman:2000:PAQ

[HL00b] Fredrik Hedman and Aatto Laaksonen. Parallel aspects of quantum molecular dynamics simulations of liquids. Computer Physics Communications, 128(1–2):284–294, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000825.

Hong:2000:NSP

[HL00c] Jialin Hong and Ying Liu. Numerical simulation of periodic and quasiperiodic solutions for nonautonomous Hamiltonian systems via the scheme preserving weak invariance.

Computer Physics Communications, 131(1–2):86–94, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000850.

Huang:2005:FDT

[HL05] Sh. J. Huang and F. Li. Finite-difference time-domain simulation of electromagnetic propagation in magnetized plasma. Computer Physics Communications, 166(1):45–52, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005065.

Hahn:2008:FTD

[HL08a] T. Hahn and P. Lang. FeynEdit — a tool for drawing Feynman diagrams. Computer Physics Communications, 179(12):931–935, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002749.

Haxton:2008:SMS

[HL08b] Wick Haxton and Cecilia Lunardini. SevenOperators, a Mathematica script for harmonic oscillator nuclear matrix elements arising in semileptonic electroweak interactions. Computer Physics Communications, 179(5):345–358, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001112.

Huo:2006:EAC

[HLB06] Yan Huo, Heng Liang, and Fengshan Bai. An efficient algorithm for computing permanental polynomials of graphs. Computer Physics Communications, 175(3):196–203, August 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001275.

Ho:2008:IPN

[HLC08] Gregory S. Ho, Vincent L. Lignères, and Emily A. Carter. Introducing PROFESS: a new program for orbital-free density functional theory calculations. Computer Physics Communications, 179(11):839–854, December 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002476.

Huber:2005:SLB

[HLW05] P. Huber, M. Lindner, and W. Winter. Simulation of long-baseline neutrino oscillation experiments with GLoBES: (general long baseline experiment simulator). Computer Physics Communications, 167(3):195–202, May 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000743.

Herrmann:2000:SGM

[HM00] H. J. Herrmann and M. Müller. Simulations of granular materials on different scales. Computer Physics Communications, 127(1):120–125, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550000014X.

Herzbach:2006:PCM

[HM06a] Daniel Herzbach and Martin H. Müser. Piezoelectric coefficients by molecular dynamics simulations in the constant stress ensemble: a case study of quartz. Computer Physics Communications, 174(1):17–23, January 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004704.

Huber:2006:HMP

[HM06b] T. Huber and D. Maître. HypExp, a Mathematica package for expanding hypergeometric functions around integer-valued parameters. Computer Physics Communications, 175(2):122–144, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506001135.

Huber:2008:HEH

[HM08] Tobias Huber and Daniel Maître. HypExp 2, expanding hypergeometric functions about half-integer parameters. Computer Physics Communications, 178(10):755–776, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000027.

Ho:2002:CSM

[HMY^{+}02] T. K. Ho, B. H. Mao, Z. Z. Yuan, H. D. Liu, and Y. F. Fung. Computer simulation and modeling in railway applications. Computer Physics Communications, 143(1):1–10, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004106.

Hsu:2005:SLA

[HNG05] Hsiao-Ping Hsu, Walder Nadler, and Peter Grassberger. Statistics of lattice animals. Computer Physics Communications, 169(1–3):114–116, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001244.

Hasebe:2001:SSE

[HNS01] Katsuya Hasebe, Akihiro Nakayama, and Yuki Sugiyama. Soliton solutions of exactly solvable dissipative systems. Computer Physics Communications, 142(1–3):259–262, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100337X.

Horiuchi:2004:SFD

[HOI04] Ritoku Horiuchi, H. Ohtani, and A. Ishizawa. Structure formation and dynamical behavior of kinetic plasmas controlled by magnetic reconnection. Computer Physics Communications, 164(1–3):17–22, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002504.

Hong:2004:CDA

[Hon04] Woo-Pyo Hong. Comment on “A direct approach with computerized symbolic computation for finding a series of traveling waves to nonlinear equations” by Engui Fan and H. H. Dai: [Comput. Phys. Commun. 153 (2003) 17]. Computer Physics Communications, 158(3):192–195, April 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504000542. See [FD03].

Hoover:2004:BRB

[Hoo04] William G. Hoover. Book review: Computational Partial Differential Equations-Numerical Methods and Diffpack Programming: Hans Petter Langtangen, second ed., Springer-Verlag, Berlin, 2003, 855 pages, ISBN 3-540-43416-X, ISSN 1611-0994. Computer Physics Communications, 161(3):181–182, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001481.

Horvat:2009:EFP

[Hor09] Vladimir Horvat. ERCS08: a FORTRAN program equipped with a Windows graphics user interface that calculates ECPSSR cross sections for the removal of atomic electrons. Computer Physics Communications, 180(6):995–1003, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004517.

Hayashi:2007:MDS

[HOT07] Keiji Hayashi, Noriyoshi Omote, and Daisuke Tanaka. Molecular dynamics study of atomistic mechanisms of wearless friction determining threshold sliding velocity. Computer Physics Communications, 177(1–2):51, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001373.

Henkel:2002:ASG

[HP02] Malte Henkel and Michel Pleimling. Anisotropic scaling and generalized conformal invariance at Lifshitz points. Computer Physics Communications, 147(1–2):419–422, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200317X.

Hasnip:2006:EEM

[HP06] P. J. Hasnip and C. J. Pickard. Electronic energy minimisation with ultrasoft pseudopotentials. Computer Physics Communications, 174(1):24–29, January 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004716.

Holzmann:2005:CEI

[HPC05] Markus Holzmann, Carlo Pierleoni, and David M. Ceperley. Coupled electron-ion Monte Carlo calculations of atomic hydrogen. Computer Physics Communications, 169(1–3):421–425, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001906.

Hillebrandt:2000:TS

[HRN00] Wolfgang Hillebrandt, Martin Reinecke, and Jens C. Niemeyer. Thermonuclear supernovae. Computer Physics Communications, 127(1):53–58, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000217.

Hoekstra:2001:WED

[HS01a] A. G. Hoekstra and P. M. A. Sloot. Wall effects on density fluctuations in the GBL thermal lattice gas automaton. Computer Physics Communications, 142(1–3):151–154, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003563.

Holtman:2001:BLL

[HS01b] Koen Holtman and Heinz Stockinger. Building a large location table to find replicas of physics objects. Computer Physics Communications, 140(1–2):146–152, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100265X.

Hahn:2002:IMS

[HS02] Thomas Hahn and Christian Schappacher. The implementation of the Minimal Supersymmetric Standard Model in FeynArts and FormCalc. Computer Physics Communications, 143(1):54–68, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004362.

Harlander:2003:RPE

[HS03] Robert V. Harlander and Matthias Steinhauser. rhad: a program for the evaluation of the hadronic R-ratio in the perturbative regime of QCD. Computer Physics Communications, 153(2):244–274, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002042.

Hur:2007:NEK

[HS07] Min Sup Hur and Hyyong Suk. New envelope-kinetic scheme for the simulation of Raman backward laser amplification. Computer Physics Communications, 177(1–2):68–69, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001142.

Hankel:2008:DPR

[HSGBK08] Marlies Hankel, Sean C. Smith, Stephen K. Gray, and Gabriel G. Balint-Kurti. DIFFREALWAVE: a parallel real wavepacket code for the quantum mechanical calculation of reactive state-to-state differential cross sections in atom plus diatom collisions. Computer Physics Communications, 179(8):569–578, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508001938.

Hannemann:2002:RCP

[HSJ02] A. Hannemann, J. C. Schön, and M. Jansen. A random close packing based algorithm for the generation of continuous random networks. Computer Physics Communications, 144(3):284–296, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002680.

Hughes:2008:ISC

[HSS^{+}08] Zak E. Hughes, Lorna M. Stimson, Henk Slim, Juho S. Lintuvuori, Jaroslav M. Ilnytskyi, and Mark R. Wilson. An investigation of soft-core potentials for the simulation of mesogenic molecules and molecules composed of rigid and flexible segments. Computer Physics Communications, 178(10):724–731, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000477.

Hammura:2001:IMO

[HSSA01] Kiyotaka Hammura, Kazuo Sakai, Miho Seyama, and Yoshinobu Aoyagi. Interpretation of magneto-oscillatory spectra in cuprous oxide as quantum manifestation of classical non-integrability. Computer Physics Communications, 142(1–3):311–315, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003423.

Hosseini:2008:MUP

[HTA08] M. Hosseini, A. Tavana, and M. Akhavan. MgB_{2} under pressure and plane strain: a DFT study. Computer Physics Communications, 179(6):385–390, September 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001471.

Houili:2003:MSM

[HTL^{+}03] H. Houili, E. Tutis, H. Lütjens, M. N. Bussac, and L. Zuppiroli. MOLED: Simulation of multilayer organic light emitting diodes. Computer Physics Communications, 156(1):108–122, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503004351.

Holzwarth:2001:PAW

[HTM01] N. A. W. Holzwarth, A. R. Tackett, and G. E. Matthews. A Projector Augmented Wave (PAW) code for electronic structure calculations, Part I: atompaw for generating atomcentered functions. Computer Physics Communications, 135 (3):329–347, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002447.

Hayashi:2008:VDT

[HTM^{+}08] Keiji Hayashi, Daisuke Tanaka, Tomoki Maruyama, Harumi Araki, Daisuke Matsumura, and Misato Kaneko. Velocitydependent threshold behavior of wearless nano-friction as studied in terms of spatial distribution of the local quasitemperature. Computer Physics Communications, 179(1–3): 98–101, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000222.

Hansen:2006:TFVa

[HTNFBS06a] Poul Martin Hansen, Iva Marija Tolić-Nørrelykke, Henrik Flyvbjerg, and Kirstine Berg-Sørensen. tweezercalib 2.0: Faster version of Matlab package for precise calibration of optical tweezers. Computer Physics Communications, 174(6):518–520, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505006247.

Hansen:2006:TFVb

[HTNFBS06b] Poul Martin Hansen, Iva Marija Tolic-Nørrelykke, Henrik Flyvbjerg, and Kirstine Berg-Sørensen. tweezercalib 2.1: Faster version of Matlab package for precise calibration of optical tweezers. Computer Physics Communications, 175(8):572–573, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600289X.

Huang:2009:SNS

[Hua09] Chia-Chien Huang. Semiconductor nanodevice simulation by multidomain spectral method with Chebyshev, prolate spheroidal and Laguerre basis functions. Computer Physics Communications, 180(3):375–383, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800369X.

Hujeirat:2005:PON

[Huj05] A. Hujeirat. A problem-orientable numerical algorithm for modeling multi-dimensional radiative MHD flows in astrophysics-the hierarchical solution scenario. Computer Physics Communications, 168(1):1–24, May 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000640.

Hukushima:2002:EEM

[Huk02] Koji Hukushima. Extended ensemble Monte Carlo approach to hardly relaxing problems. Computer Physics Communications, 147(1–2):77–82, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002072.

Hartgers:2001:CGC

[HvDJvdM01] A. Hartgers, J. van Dijk, J. Jonkers, and J. A. M. van der Mullen. CRModel: a general collisional radiative modeling code. Computer Physics Communications, 135(2):199–218, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002319.

Hart:2009:AGL

[HvHHM09] A. Hart, G. M. von Hippel, R. R. Horgan, and E. H. Müller. Automated generation of lattice QCD Feynman rules. Computer Physics Communications, 180(12):2698–2716, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001325.

Hur:2009:TDS

[HW09] Min Sup Hur and Jonathan S. Wurtele. Two-dimensional simulations of the amplification and focusing of intense laser pulses in the kinetic regime of Raman backward amplification in plasmas. Computer Physics Communications, 180(4):651– 655, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000289.

Hamaguchi:2007:ALS

[HYY07] Satoshi Hamaguchi, Masashi Yamashiro, and Hideaki Yamada. Atomic-level simulation of non-equilibrium surface chemical reactions under plasma-wall interaction. Computer Physics Communications, 177(1–2):108–109, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001075.

Huang:2009:NBT

[HZGZ09] Ting-Zhu Huang, Li-Tao Zhang, Tong-Xiang Gu, and Xian-Yu Zuo. New block triangular preconditioner for linear systems arising from the discretized time-harmonic Maxwell equations. Computer Physics Communications, 180(10):1853–1859, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001660.

Ismail-Beigi:2000:NAF

[IBA00] Sohrab Ismail-Beigi and T. A. Arias. New algebraic formulation of density functional calculation. Computer Physics Communications, 128(1–2):1–45, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000722.

Ivanov:2003:OPP

[IBM03] I. A. Ivanov, M. W. J. Bromley, and J. Mitroy. Orthogonalising pseudo-potentials in scattering calculations. Computer Physics Communications, 152(1):9–24, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007579.

Inoue:2001:DSM

[ICO01] Yasuhiro Inoue, Yu Chen, and Hirotada Ohashi. Development of a simulation model for solid objects suspended in a fluctuating fluid. Computer Physics Communications, 142(1–3):114–116, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003435.

Inoue:2003:DCS

[ICO03] Yasuhiro Inoue, Yu Chen, and Hirotada Ohashi. On the density correlation of the spontaneous fluctuation in a real-coded lattice gas. Computer Physics Communications, 153(1):66–70, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503001541.

Iovieno:2001:NTP

[ICT01] Michele Iovieno, Carlo Cavazzoni, and Daniela Tordella. A new technique for a parallel dealiased pseudospectral Navier–Stokes code. Computer Physics Communications, 141(3):365–374, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004337.

Illenseer:2009:TDC

[ID09] Tobias F. Illenseer and Wolfgang J. Duschl. Two-dimensional central-upwind schemes for curvilinear grids and application to gas dynamics with angular momentum. Computer Physics Communications, 180(11):2283–2302, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002367.

Ida:2000:IUS

[Ida00] Masato Ida. An improved unified solver for compressible and incompressible fluids involving free surfaces. Part I. Convection. Computer Physics Communications, 132(1–2):44–65, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001363.

Ida:2002:CSL

[Ida02] Masato Ida. A conservative semi-Lagrangian method for oscillation-free computation of advection processes. Computer Physics Communications, 143(2):142–154, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004416.

Ida:2003:EIU

[Ida03a] Masato Ida. Erratum to “An improved unified solver for compressible and incompressible fluids involving free surfaces. II. Multi-time-step integration and applications”: [Comp. Phys. Commun. 150 (2003) 300–322]. Computer Physics Communications, 150(3):323, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008159. See [Ida03b].

Ida:2003:IUS

[Ida03b] Masato Ida. An improved unified solver for compressible and incompressible fluids involving free surfaces. II. Multitime-step integration and applications. Computer Physics Communications, 150(3):300–322, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006914. See erratum [Ida03a].

Ilin:2004:ISI

[IDS^{+}04] Andrew V. Ilin, Franklin R. Chang Díaz, Jared P. Squire, Alfonso G. Tarditi, and Mark D. Carter. Improved simulation of the ICRF waves in the VASIMR plasma. Computer Physics Communications, 164(1–3):251–257, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002887.

Isotani:2003:GPA

[IF03] Sadao Isotani and Américo Tsuneo Fujii. A grid procedure applied to the determination of parameters of a kinetic process. Computer Physics Communications, 151(1):1–7, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006926.

Inghoff:2001:MPC

[IFF01] T. Inghoff, S. Fritzsche, and B. Fricke. Maple procedures for the coupling of angular momenta. IV. Spherical harmonics. Computer Physics Communications, 139(3):297–313, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002181.

Ido:2001:TDM

[IH01] Shunji Ido and Ryusuke Hirose. Three-dimensional magnetic field analyses on the magnetron sputtering system by using free mesh method. Computer Physics Communications, 142(1–3):338–342, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003678.

Ibanez:2009:SDM

[IH09] Javier Ibáñez and Vicente Hernández. Solving differential matrix Riccati equations by a piecewise-linearized method based on the conmutant equation. Computer Physics Communications, 180(11):2103–2114, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001933.

Ibanez:2009:SIV

[IHAR09] J. Ibáñez, V. Hernández, E. Arias, and P. A. Ruiz. Solving initial value problems for ordinary differential equations by two approaches: BDF and piecewise-linearized methods. Computer Physics Communications, 180(5):712–723, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003925.

Idomura:2008:CGG

[IIK^{+}08] Yasuhiro Idomura, Masato Ida, Takuma Kano, Nobuyuki Aiba, and Shinji Tokuda. Conservative global gyrokinetic toroidal full- f five-dimensional Vlasov simulation. Computer Physics Communications, 179(6):391–403, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001409.

Ihle:2000:TLB

[IK00] T. Ihle and D. M. Kroll. Thermal lattice-Boltzmann method for non-ideal gases with potential energy. Computer Physics Communications, 129(1–3):1–12, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000874.

Inamuro:2000:GIM

[IKO00] Takaji Inamuro, Nobuharu Konishi, and Fumimaru Ogino. A Galilean invariant model of the lattice Boltzmann method for multiphase fluid flows using free-energy approach. Computer Physics Communications, 129(1–3):32–45, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000904.

Iza:2007:LPP

[IL07] F. Iza and J. K. Lee. Low-pressure plasma generation inside slender tubes. Computer Physics Communications, 177(1–2): 72–75, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001166.

Ivanov:2001:TCS

[IM01] I. A. Ivanov and J. Mitroy. Treatment of the Coulomb singularity in momentum space calculations. Computer Physics Communications, 134(3):317–320, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002101.

Itoh:2002:VEI

[IN02] S. Itoh and U. Nagashima. Variation effect on the insecticide activity of DDT analogues. A chemometric approach. Computer Physics Communications, 147(1–2):182–185, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002849.

Ishida:2009:NSD

[IN09]  Kunio Ishida and Keiichiro Nasu.  Numerical simula-

tion on the dynamics of photoinduced cooperative phenomena in molecular crystals. Computer Physics Communications, 180(9):1489–1494, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550900109X.

Inglesfield:2001:ES

[Ing01] J. E. Inglesfield. Embedding at surfaces. Computer Physics Communications, 137(1):89–107, June 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001734.

Nomura:2008:SPA

[iNKNV08] Ken ichi Nomura, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. A scalable parallel algorithm for large-scale reactive force-field molecular dynamics simulations. Computer Physics Communications, 178(2):73–87, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003748.

Inui:2007:NER

[Inu07] Norio Inui. Numerical errors resulting from finite-difference approximation in computations of a one-dimensional Schrödinger equation with the Trotter formula. Computer Physics Communications, 177(10):781–786, November 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550700327X.

Ishikawa:2000:SPC

[IOM00] T. Ishikawa, Y. Ohtsuka, and M. Matsunami. A special purpose computer for diffusion equations. Computer Physics Communications, 131(1–2):68–77, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000813.

Ohe:2001:ACN

[iOY01] Jun ichiro Ohe and Kousuke Yakubo. AC current in nonequilibrium mesoscopic systems. Computer Physics Communications, 142(1–3):436–441, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003800.

Ixaru:2001:GQR

[IP01] L. Gr. Ixaru and B. Paternoster. A Gauss quadrature rule for oscillatory integrands. Computer Physics Communications, 133(2–3):177–188, January 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001739.

Satake:2008:SHS

[iSAK^{+}08] Shin ichi Satake, Takafumi Anraku, Hiroyuki Kanamori, Tomoaki Kunugi, Kazuho Sato, and Tomoyoshi Ito. Study on high speed parallel algorithm using PC grid environment for visualization measurements by digital holographic particle tracking velocimetry. Computer Physics Communications, 178(1):1–7, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003645.

Isobe:2001:MMD

[ISH01] M. Isobe, H. Shimizu, and Y. Hiwatari. A multicanonical molecular dynamics study for a model protein-g. Computer Physics Communications, 142(1–3):144–147, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100354X.

Satake:2008:SPC

[iSHS^{+}08] Shin ichi Satake, Yoshiaki Hiroi, Yuya Suzuki, Nobuyuki Masuda, and Tomoyoshi Ito. Special-purpose computer for two-dimensional FFT. Computer Physics Communications, 179(6):404–408, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508001355.

Ihrig:2002:QCP

[ISS^{+}02] Dieter F. Ihrig, Jörg Stockhaus, Frank Scheide, Oliver Winkelhake, and Oliver Streuber. Quantification of channels of plasma polymerization using a chemical model based on mass spectrometry. Computer Physics Communications, 147(1–2):186–189, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002412.

Ishio:2001:WFS

[ISSB01] Hiromu Ishio, Alexander I. Saichev, Almas F. Sadreev, and Karl-Fredrik Berggren. Wave function statistics for mesoscopic transport through chaotic open billiards: time reversibility, space reciprocity breaking and statistical crossover. Computer Physics Communications, 142(1–3):64–70, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003204.

Innocente:2001:CSA

[ISSC01] V. Innocente, L. Silvestris, D. Stickland, and CMS Software Group. CMS software architecture: Software framework, services and persistency in high level trigger, reconstruction and analysis. Computer Physics Communications, 140(1–2):31–44, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002533.

S:2005:OSC

[ISX05] S. Ii, M. Shimuta, and F. Xiao. A 4th-order and single-cell-based advection scheme on unstructured grids using multimoments. Computer Physics Communications, 173(1–2):17–33, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004236.

Tsuda:2001:IMS

[iTKST01] Ken ichiro Tsuda, Hiroki Kaneko, Jiro Shimada, and Toshikazu Takada. Ab initio MO studies of interaction mechanisms of protein kinase C with cell membranes. Computer Physics Communications, 142(1–3):140–143, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003538.

Veld:2008:AEM

[itVPG08] Pieter J. in ’t Veld, Steven J. Plimpton, and Gary S. Grest. Accurate and efficient methods for modeling colloidal mixtures in an explicit solvent using molecular dynamics. Computer Physics Communications, 179(5):320–329, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550800115X.

Ixaru:2003:EFV

[IVD03] L. Gr. Ixaru, G. Vanden Berghe, and H. De Meyer. Exponentially fitted variable two-step BDF algorithm for first order ODEs. Computer Physics Communications, 150(2):116–128, February 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006768.

Ilnytskyi:2001:DDM

[IW01] Jaroslav Ilnytskyi and Mark R. Wilson. A domain decomposition molecular dynamics program for the simulation of flexible molecules with an arbitrary topology of Lennard-Jones and/or Gay–Berne sites. Computer Physics Communications, 134(1):23–32, February 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001879.

Ilnytskyi:2002:DDM

[IW02] Jaroslav M. Ilnytskyi and Mark R. Wilson. A domain decomposition molecular dynamics program for the simulation of flexible molecules of spherically-symmetrical and nonspherical sites. II. Extension to NVT and NPT ensembles. Computer Physics Communications, 148(1):43–58, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004678.

Iwamatsu:2001:AEP

[Iwa01] Masao Iwamatsu. Applying evolutionary programming to structural optimization of atomic clusters. Computer Physics Communications, 142(1–3):214–218, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004490.

Ixaru:2001:CRO

[Ixa01] L. Gr. Ixaru. A comment on “Reliable operations on oscillatory functions”: [Comput. Phys. Commun. 125 (2000) 127]. Computer Physics Communications, 134(2):267–268, February 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001788. See [AA00,

AA01b].

Ixaru:2002:LPS

[Ixa02] L. Gr. Ixaru. LILIX — a package for the solution of the coupled channel Schrödinger equation. Computer Physics Communications, 147(3):834–852, September 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004599.

Ixaru:2007:ECA

[Ixa07a] L. Gr. Ixaru. Efficient computation of the Airy propagators. Computer Physics Communications, 176(11–12):637– 641, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001798.

Ixaru:2007:FLM

[Ixa07b] L. Gr. Ixaru. Fast LP method for the Schrödinger equation. Computer Physics Communications, 177(12):897–907, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003517.

Jadach:2000:FMD

[Jad00] S. Jadach. Foam: Multi-dimensional general purpose Monte Carlo generator with self-adapting simplical grid. Computer Physics Communications, 130(3):244–259, August 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000473.

Jadach:2003:FGP

[Jad03] S. Jadach. Foam: a general-purpose cellular Monte Carlo event generator. Computer Physics Communications, 152(1):55–100, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007555.

Jamieson:2000:NRL

[Jam00] M. J. Jamieson. Newton–Raphson/log-derivative method for finding eigenenergies. Computer Physics Communications, 125(1–3):193–195, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004865.

Jansen:2005:GCT

[Jan05] Karl Jansen. Going chiral: twisted mass versus overlap fermions. Computer Physics Communications, 169(1–3):362– 364, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001773.

Jabbarzadeh:2003:PAM

[JAT03] A. Jabbarzadeh, J. D. Atkinson, and R. I. Tanner. A parallel algorithm for molecular dynamics simulation of branched molecules. Computer Physics Communications, 150(2):65–84, February 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006318.

Johansson:2005:SWV

[JBA05] Erik Johansson, Kim Bolton, and Peter Ahlström. Simulation of water vapor clusters in equilibrium with liquid water. Computer Physics Communications, 169(1–3):54–56, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001116.

Jolliet:2007:GCP

[JBA^{+}07] S. Jolliet, A. Bottino, P. Angelino, R. Hatzky, T. M. Tran, B. F. Mcmillan, O. Sauter, K. Appert, Y. Idomura, and L. Villard. A global collisionless PIC code in magnetic coordinates. Computer Physics Communications, 177(5):409–425, September 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002251.

Jaun:2001:ISG

[JBBR01] A. Jaun, K. Blomqvist, A. Bondeson, and T. Rylander. Iterative solution of global electromagnetic wavefields with finite elements. Computer Physics Communications, 135(1):74–81, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000223X.

Jungblut:2008:IIP

[JBS08] S. Jungblut, K. Binder, and T. Schilling. Isotropic-isotropic phase separation in mixtures of rods and spheres: Some aspects of Monte Carlo simulation in the grand canonical ensemble. Computer Physics Communications, 179(1–3):13–16, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000088.

Judge:2001:ABW

[JC01] R. H. Judge and Dennis J. Clouthier. AsyrotWin: a 32bit Windows version of Asyrot, a program for the analysis of high resolution singlet-singlet band spectra of asymmetric tops. Computer Physics Communications, 135(3):293–311, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002332.

Jiang:2007:AAL

[JC07] Tongsong Jiang and Li Chen. Algebraic algorithms for least squares problem in quaternionic quantum theory. Computer Physics Communications, 176(7):481–485, April 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000306.

Jezequel:2008:CLE

[JC08a] Fabienne Jézéquel and Jean-Marie Chesneaux. CADNA: a library for estimating round-off error propagation. Computer Physics Communications, 178(12):933–955, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000775.

Jiang:2008:AMS

[JC08b] Tongsong Jiang and Li Chen. An algebraic method for Schrödinger equations in quaternionic quantum mechanics. Computer Physics Communications, 178(11):795–799, June 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000490.

Julia-Diaz:2006:QMQ

[JDBT06] Bruno Juliá-Díaz, Joseph M. Burdis, and Frank Tabakin. QDENSITY — a Mathematica quantum computer simulation. Computer Physics Communications, 174(11):914–934, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506000543.

Julia-Diaz:2009:QMQ

[JDBT09] Bruno Juliá-Díaz, Joseph M. Burdis, and Frank Tabakin. QDENSITY — a Mathematica quantum computer simulation. Computer Physics Communications, 180(3):474, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003457.

Jenko:2000:MPV

[Jen00] Frank Jenko. Massively parallel Vlasov simulation of electromagnetic drift-wave turbulence. Computer Physics Communications, 125(1–3):196–209, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004890.

Jensen:2001:ECS

[Jen01] Iwan Jensen. Enumeration of compact self-avoiding walks. Computer Physics Communications, 142(1–3):109–113, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550100340X.

Jonsson:2002:PCW

[JG02]  Per Jönsson and Stefan Gustafsson.  A program for

computing weak and intermediate field Zeeman splittings from MCHF wave functions. Computer Physics Communications, 144(2):188–199, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004611.

Jeon:2009:EPA

[JGJ09] Byoungseon Jeon and Niels Grønbech-Jensen. Efficient parallel algorithm for statistical ion track simulations in crystalline materials. Computer Physics Communications, 180(2):231–237, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003202.

Jacobs:2009:IET

[JH09a] G. B. Jacobs and J. S. Hesthaven. Implicit-explicit time integration of a high-order particle-in-cell method with hyperbolic divergence cleaning. Computer Physics Communications, 180(10):1760–1767, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001490.

Jiang:2009:SDR

[JH09b] J. G. Jiang and Y. N. Huang. Simulation of distribution of random walks on a lattice. Computer Physics Communications, 180(2):177–179, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003019.

Jonsson:2007:GRA

[JHFG07] P. Jönsson, X. He, C. Froese Fischer, and I. P. Grant. The grasp2K relativistic atomic structure package. Computer Physics Communications, 177(7):597–622, October 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003001.

Jiang:2008:CAH

[Jia08] T. F. Jiang. Calculation of atomic hydrogen and its photoelectron spectra in momentum space. Computer Physics Communications, 178(8):571–577, April 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004870.

Jeong:2003:SPH

[JJHvO03] J. H. Jeong, M. S. Jhon, J. S. Halow, and J. van Osdol. Smoothed particle hydrodynamics: Applications to heat conduction. Computer Physics Communications, 153(1):71–84, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503001553.

Janke:2005:CEG

[JJK05] Wolfhard Janke, Des Johnston, and Ralph Kenna. Critical exponents from general distributions of zeroes. Computer Physics Communications, 169(1–3):457–461, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001980.

Jiang:2001:IAL

[JK01] P. Jiang and W. C. Kok. Intradonor 1s-2p_{+} absorption line shapes of doped GaAs/Al_{0.3}Ga_{0.7}As superlattices. Computer Physics Communications, 142(1–3):300–304, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003654.

Janke:2002:DPF

[JK02] Wolfhard Janke and Ralph Kenna. Density of partition function zeroes and phase transition strength. Computer Physics Communications, 147(1–2):443–446, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003235.

Janecek:2008:FSP

[JK08] S. Janecek and E. Krotscheck. A fast and simple program for solving local Schrödinger equations in two and three dimensions. Computer Physics Communications, 178(11):835–842, June 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000453.

Jeon:2008:PTC

[JKCGJ08] Byoungseon Jeon, Joel D. Kress, Lee A. Collins, and Niels Grønbech-Jensen. Parallel TREE code for twocomponent ultracold plasma analysis. Computer Physics Communications, 178(4):272–279, February 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507004171.

Jasberg:2000:HFA

[JKKT00] A. Jäsberg, A. Koponen, M. Kataja, and J. Timonen. Hydrodynamical forces acting on particles in a two-dimensional flow near a solid wall. Computer Physics Communications, 129(1–3):196–206, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001077.

Jacholkowska:2000:LSA

[JKW00] A. Jacholkowska, J. Kalinowski, and Z. Was. Library of SM and anomalous WW γ couplings for the e^{+}e^{−} → ffνnγ Monte Carlo programs. Computer Physics Communications, 124(2–3):238–242, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004385. See erratum [JKW06].

Jacholkowska:2006:ELS

[JKW06] A. Jacholkowska, J. Kalinowski, and Z. Was. Erratum to “Library of SM and anomalous WW γ couplings for the e^{+}e^{−} → ff̄nγ Monte Carlo programs” [Comput. Phys. Commun. 124 (2000) 238]. Computer Physics Communications, 174(3):253, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005746. See [JKW00].

Jiang:2007:MSM

[JOS07] Fangming Jiang, Mónica S. A. Oliveira, and Antonio C. M. Sousa. Mesoscale SPH modeling of fluid flow in isotropic porous media. Computer Physics Communications, 176(7):471–480, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700029X.

Jackin:2009:COS

[JP09] B. J. Jackin and P. K. Palanisamy. A completely open source based computing system for computer generation of Fourier holograms. Computer Physics Communications, 180 (10):1882–1887, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001246.

Jadach:2001:MCE

[JPS^{+}01a] S. Jadach, W. Placzek, M. Skrzypek, B. F. L. Ward, and Z. Was. The Monte Carlo event generator YFSWW3 version 1.16 for W -pair production and decay at LEP2/LC energies. Computer Physics Communications, 140(3):432–474, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002880.

Jadach:2001:MCP

[JPS^{+}01b] S. Jadach, W. Placzek, M. Skrzypek, B. F. L. Ward, and Z. Was. The Monte Carlo program KoralW version 1.51 and the concurrent Monte Carlo KoralW&YFSWW3 with all background graphs and first-order corrections to W -pair production. Computer Physics Communications, 140(3):475–512, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550100296X.

Jadach:2009:CMQ

[JPS^{+}09] S. Jadach, W. Placzek, M. Skrzypek, P. Stephens, and Z. Was. Constrained MC for QCD evolution with rapidity ordering and minimum kT. Computer Physics Communications, 180(5):675– 698, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003822.

Jonsson:2000:PCM

[JRT00] Dan Jonsson, Kenneth Ruud, and Peter R. Taylor. Parallel calculations of molecular properties. Computer Physics Communications, 128(1–2):412–433, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000680.

Janke:2005:GPT

[JS05] Wolfhard Janke and Adriaan M. J. Schakel. Geometrical phase transitions. Computer Physics Communications, 169(1–3):222– 225, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001487.

Jadach:2006:SCM

[JS06] S. Jadach and M. Skrzypek. Solving constrained Markovian evolution in QCD with the help of the non-Markovian Monte Carlo. Computer Physics Communications, 175(8):511–527, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002451.

Jadach:2007:MCV

[JS07] S. Jadach and P. Sawicki. mFOAM-1.02: a compact version of the cellular event generator FOAM. Computer Physics Communications, 177(5):441–458, September 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002287.

Jureit:2008:FSM

[JS08] Jan-Hendrik Jureit and Christoph A. Stephan. Finding the standard model of particle physics a combinatorial problem. Computer Physics Communications, 178(3):230–247, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003967.

Jin:2006:EEM

[JTS^{+}06] Fanya Jin, Honghui Tong, Zhongbing Shi, Deli Tang, and Paul K. Chu. Effects of external magnetic field on propagation of electromagnetic wave in uniform magnetized plasma slabs. Computer Physics Communications, 175(8):545–552, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002724.

Jansen:2009:TPS

[JU09] Karl Jansen and Carsten Urbach. tmLQCD: a program suite to simulate Wilson twisted mass lattice QCD. Computer Physics Communications, 180(12):2717–2738, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001647.

Jung:2002:CMC

[Jun02] H. Jung. The CCFM Monte Carlo generator Cascade. Computer Physics Communications, 143(1):100–111, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004386.

Janke:2002:NTC

[JW02] Wolfhard Janke and Martin Weigel. Numerical tests of CFT conjectures for 3D spin systems. Computer Physics Communications, 147(1–2):382–387, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003107.

Jadach:2000:MCP

[JWW00a] S. Jadach, B. F. L. Ward, and Z. Was. The Monte Carlo program KORALZ, for the lepton or quark pair production at LEP/SLC energies from version 4.0 to version 4.04. Computer Physics Communications, 124(2–3):233–237, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004373.

Jadach:2000:PMC

[JWW00b] S. Jadach, B. F. L. Ward, and Z. Was. The precision Monte Carlo event generator KK for two-fermion final states in e^{+}e^{−} collisions. Computer Physics Communications, 130(3):260–325, August 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000485.

Kolobov:2004:FDF

[KA04] Vladimir Kolobov and Robert Arslanbekov. Four dimensional Fokker–Planck solver for electron kinetics in collisional gas discharge plasmas. Computer Physics Communications, 164(1–3):195–201, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002802.

Komatsu:2005:NDN

[KA05] Nobuyoshi Komatsu and Takashi Abe. Noise-driven numerical irreversibility in molecular dynamics technique. Computer Physics Communications, 171(3):187–196, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505003127.

Kanth:2009:DTM

[KA09] A. S. V. Ravi Kanth and K. Aruna. Differential transform method for solving the linear and nonlinear Klein–Gordon equation. Computer Physics Communications, 180(5):708– 711, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003901.

Kendall:2000:HPC

[KAB^{+}00] Ricky A. Kendall, Edoardo Aprà, David E. Bernholdt, Eric J. Bylaska, Michel Dupuis, George I. Fann, Robert J. Harrison, Jialin Ju, Jeffrey A. Nichols, Jarek Nieplocha, T. P. Straatsma, Theresa L. Windus, and Adrian T. Wong. High performance computational chemistry: An overview of NWChem a distributed parallel application. Computer Physics Communications, 128(1–2):260–283, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000655.

Kranjc:2007:SMH

[KACB07] Agata Kranjc, Claudio Anselmi, Paolo Carloni, and Frank E. Blaney. Structural models of human big conductance calcium- and voltage-gated potassium channels. Computer Physics Communications, 177(1–2):21–26, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000793.

Krieg:2008:CMC

[KALC08] Stefan Krieg, Norbert Attig, Thomas Lippert, and Nigel Cundy. A comparison of methods to calculate the chiral condensate with overlap fermions. Computer Physics Communications, 179(1–3):181–183, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000374.

Karwowski:2001:IPG

[Kar01] Jacek Karwowski. Ian P. Grant, FRS — a biographical note. Computer Physics Communications, 138(1):10–17, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002466.

Karakasidis:2002:VPN

[Kar02] Theodoros E. Karakasidis. Vibrational properties of a Σ 5(310)[001] NiO grain boundary as a function of temperature: a molecular dynamics simulation. Computer Physics Communications, 147(1–2):198–201, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003831.

Kastner:2000:RCS

[Kas00] S. O. Kastner. Rao coefficients for the solution of convolution integral equations. Computer Physics Communications, 127(2–3):286–289, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599005226.

Katzgraber:2002:MCS

[Kat02] Helmut G. Katzgraber. Monte Carlo simulations of vector spin glasses at low temperatures. Computer Physics Communications, 147(1–2):439–442, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003223.

Kauer:2003:SAN

[Kau03]  Nikolas Kauer.  SINGINT: Automatic numerical inte-

gration of singular integrands. Computer Physics Communications, 153(2):233–243, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503001607.

Krug:2002:CIN

[KB02] Andreas Krug and Andreas Buchleitner. Chaotic ionization of non-classical alkali Rydberg states-computational physics beats experiment. Computer Physics Communications, 147(1–2):394–397, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003120.

Koo:2004:CMH

[KB04] Justin W. Koo and Iain D. Boyd. Computational model of a Hall thruster. Computer Physics Communications, 164(1– 3):442–447, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003169.

Kadau:2002:CDS

[KBBW02] Dirk Kadau, Guido Bartels, Lothar Brendel, and Dietrich E. Wolf. Contact dynamics simulations of compacting cohesive granular systems. Computer Physics Communications, 147(1–2):190–193, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002424.

Kong:2009:IGF

[KBC^{+}09] Ling Ti Kong, Guido Bartels, Carlos Campañá, Colin Denniston, and Martin H. Müser. Implementation of Green’s function molecular dynamics: An extension to LAMMPS. Computer Physics Communications, 180(6):1004–1010, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004529.

Koradi:2000:PCD

[KBG00] R. Koradi, M. Billeter, and P. Güntert. Point-centered domain decomposition for parallel molecular dynamics simulation. Computer Physics Communications, 124(2–3):139–147, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004361.

Klosiewicz:2009:UPA

[KBV09] P. Klosiewicz, J. Broeckhove, and W. Vanroose. Using pseudo-arclength continuation to trace the resonances of the Schrödinger equation. Computer Physics Communications, 180(4):545–548, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003949.

Kalia:2000:MAM

[KCC^{+}00] Rajiv K. Kalia, Timothy J. Campbell, Alok Chatterjee, Aiichiro Nakano, Priya Vashishta, and Shuji Ogata. Multiresolution algorithms for massively parallel molecular dynamics simulations of nanostructured materials. Computer Physics Communications, 128(1–2):245–259, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599005391.

Kirk:2000:ACC

[KCH00] N. P. Kirk, J. N. L. Connor, and C. A. Hobbs. An adaptive contour code for the numerical evaluation of the oscillatory cuspoid canonical integrals and their derivatives. Computer Physics Communications, 132(1–2):142–165, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001260.

Kim:2007:PSM

[KCR07] Daeho Kim, Hyonu Chang, and Chang-Mo Ryu. Particle simulation of a magnetized dual-frequency capacitively coupled plasma. Computer Physics Communications, 177(1–2): 125, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001051.

Korkmaz:2009:SWS

[KD09] Alper Korkmaz and Idris Dag. Solitary wave simulations of complex modified Korteweg–de Vries equation using differential quadrature method. Computer Physics Communications, 180(9):1516–1523, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001167.

Krasovitskii:2004:NPD

[KDSB04] V. B. Krasovitskii, V. G. Dorofeenko, V. I. Sotnikov, and B. S. Bauer. Nonlinear processes during the interaction of a petawatt laser pulse with a plasma in the presence of an external magnetic field. Computer Physics Communications, 164(1–3):103–107, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002656.

Klintenberg:2000:ACF

[KDW00] M. Klintenberg, S. E. Derenzo, and M. J. Weber. Accurate crystal fields for embedded cluster calculations. Computer Physics Communications, 131(1–2):120–128, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000710.

Kadau:2002:MDS

[KEL02] K. Kadau, P. Entel, and P. S. Lomdahl. Moleculardynamics study of martensitic transformations in sintered Fe– Ni nanoparticles. Computer Physics Communications, 147 (1–2):126–129, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002308.

Kenzler:2001:PSC

[KEM^{+}01] R. Kenzler, F. Eurich, P. Maass, B. Rinn, J. Schropp, E. Bohl, and W. Dieterich. Phase separation in confined geometries: Solving the Cahn–Hilliard equation with generic boundary conditions. Computer Physics Communications, 133(2–3):139–157, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001594.

Kerstein:2002:ODT

[Ker02] Alan R. Kerstein. One-dimensional turbulence: a new approach to high-fidelity subgrid closure of turbulent flow simulations. Computer Physics Communications, 148(1):1–16, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005520.

Koval:2003:RWG

[KF03] Peter Koval and Stephan Fritzsche. Relativistic wave and Green’s functions for hydrogen-like ions. Computer Physics Communications, 152(2):191–207, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502007981.

Kim:2005:FCC

[KF05a] Young C. Kim and Michael E. Fisher. Fluid coexistence close to criticality: scaling algorithms for precise simulation. Computer Physics Communications, 169(1–3):295–300, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001633.

Koval:2005:RCF

[KF05b] Peter Koval and Stephan Fritzsche. Relativistic central-field Green’s functions for the ratip package. Computer Physics Communications, 172(3):187–202, November 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004200.

Kwong:2001:CDP

[KFB01] H. H. Kwong, Y. P. Feng, and T. B. Boo. Composition dependent properties of GaAs clusters. Computer Physics Communications, 142(1–3):290–294, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003484.

Kurihara:2001:GEG

[KFI^{+}01] Y. Kurihara, J. Fujimoto, T. Ishikawa, Y. Shimizu, and T. Munehisa. grc ννγ: Event generator for the singleand double-photon emission associated with neutrino pairproduction. Computer Physics Communications, 136(3):250– 268, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000254X.

Keith:2009:WDI

[KFJ^{+}09] J. Brandon Keith, Jacob R. Fennick, Chad E. Junkermeier, Daniel R. Nelson, and James P. Lewis. A Web-deployed interface for performing ab initio molecular dynamics, optimization, and electronic structure in Fireball. Computer Physics Communications, 180(3):418–426, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003354.

Kissmann:2007:LDE

[KG07] R. Kissmann and R. Grauer. A low dissipation essentially non-oscillatory central scheme. Computer Physics Communications, 176(8):522–530, April 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000252.

Koch:2000:MIT

[KGM00] Erik Koch, Olle Gunnarsson, and Richard M. Martin. Metalinsulator transitions in generalized Hubbard models. Computer Physics Communications, 127(1):137–142, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000230.

Kamberaj:2001:MCS

[KH01] H. Kamberaj and V. Helms. Monte Carlo simulation of biomolecular systems with BIOMCSIM. Computer Physics Communications, 141(3):375–402, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004349.

Kroger:2006:SCP

[KH06] Martin Kröger and Markus Hütter. Symbolic computation of the phoretic acceleration of convex particles suspended in a non-uniform gas. Computer Physics Communications, 175(10):650–664, November 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003006.

Kenny:2009:PLO

[KH09] S. D. Kenny and A. P. Horsfield. Plato: a localised orbital based density functional theory code. Computer Physics Communications, 180(12):2616–2621, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002562.

Kim:2007:EEE

[KHIL07] G. J. Kim, Y. J. Hong, F. Iza, and J. K. Lee. Evolution of the electron energy distribution function in pulsed micro hollow cathode discharge. Computer Physics Communications, 177(1– 2):131, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001002.

Kroger:2001:STJ

[KHÖ01] Martin Kröger, Markus Hütter, and Hans Christian Öttinger. Symbolic test of the Jacobi identity for given generalized ‘poisson’ bracket. Computer Physics Communications, 137(2):325– 340, June 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001618.

Kim:2003:QRI

[Kim03] Kyung Joong Kim. Quadrature rules for the integration of the product of two oscillatory functions with different frequencies. Computer Physics Communications, 153(2):135–144, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002017.

Kim:2007:ADS

[Kim07] Yong-Sung Kim. Asymmetric-dimer-symmetric-dimer mixed phase on tensile-strained Si(001) surface. Computer Physics Communications, 177(1–2):42, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465507001385.

Kirby:2006:CSW

[Kir06] P. Kirby. Calculation of spheroidal wave functions. Computer Physics Communications, 175(7):465–472, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002621.

Kono:2000:ALB

[KITK00] Koji Kono, Tatsumi Ishizuka, Hiroshi Tsuda, and Atsushi Kurosawa. Application of lattice Boltzmann model to multiphase flows with phase transition. Computer Physics Communications, 129(1–3):110–120, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000989.

Kolodziej:2004:EPN

[KJ04] Karol Kolodziej and Fred Jegerlehner. ee4fγ — a program for e^{+}e^{−} → 4f, 4fγ with nonzero fermion masses. Computer Physics Communications, 159(2):106–120, May 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001031.

Kodet:2007:NGV

[KJ07] John Kodet and Richard H. Judge. A new graphical version of STROTAB: The analysis and fitting of singlet-triplet spectra of asymmetric top molecules in the prolate or oblate limits. Computer Physics Communications, 176(9–10):601– 616, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001646.

Kosarev:2000:NCT

[KK00] E. L. Kosarev and E. P. Krasnoperov. A new calculation technique of muonium formation rate. Computer Physics Communications, 126(1–2):93–100, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004257.

Koga:2001:QPT

[KK01] Akihisa Koga and Norio Kawakami. Quantum phase transitions for the three-dimensional orthogonal-dimer spin system. Computer Physics Communications, 142(1–3):172–175, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003186.

Krawczyk:2004:HFE

[KK04] M. J. Krawczyk and K. Kulakowski. High field electrophoresiscomputer simulations. Computer Physics Communications, 163(3):172–176, November 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504004102.

Krawczyk:2005:LSS

[KK05] M. J. Krawczyk and K. Kulakowski. Off-lattice simulation of the solid phase DNA amplification. Computer Physics Communications, 170(2):131–136, August 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002626.

Kurihara:2006:NCI

[KK06]  Y. Kurihara and T. Kaneko.  Numerical contour in-

tegration for loop integrals. Computer Physics Communications, 174(7):530–539, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550500593X.

Kleimann:2004:TDM

[KKF^{+}04] Jens Kleimann, Andreas Kopp, Horst Fichtner, Rainer Grauer, and Kai Germaschewski. Three-dimensional MHD highresolution computations with CWENO employing adaptive mesh refinement. Computer Physics Communications, 158 (1):47–56, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503005460.

Kim:2007:CNF

[KKH07] Kyungsik Kim, Soo Yong Kim, and Deock-Ho Ha. Characteristics of networks in financial markets. Computer Physics Communications, 177(1–2):184–185, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000823.

Kwak:2007:DWB

[KKHL07] Jae-Hyuck Kwak, Sangwan Kim, Young Chul Hwang, and Pillwoo Lee. Development of Web-based Grid service platform for K*Grid infrastructure. Computer Physics Communications, 177(1–2):217, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000549.

Kuhn:2001:APC

[KKIS01] R. Kuhn, F. Krauss, B. Iványi, and G. Soff. APACIC++ 1.0 A PArton cascade in C++. Computer Physics Communications, 134(2):223–266, February 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002010.

Kapoyannis:2006:NAE

[KKK06] A. S. Kapoyannis, A. I. Karanikas, and C. N. Ktorides. A numerical algorithm for efficiently obtaining a Feynman parameter representation of one-gluon loop QCD Feynman diagrams for a large number of external gluons. Computer Physics Communications, 174(8):631–642, April 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000063.

Kang:2007:FPA

[KKKC07] Yong-Ju Kang, Joongoo Kang, Yong-Hoon Kim, and Kee Joo Chang. First-principles approach to the electron transport and applications for devices based on carbon nanotubes and ultrathin oxides. Computer Physics Communications, 177(1–2):30– 33, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001439.

Krawczyk:2002:NCA

[KKM02] M. J. Krawczyk, K. Kulakowski, and A. Z. Maksymowicz. New cellular automaton designed to simulate geometration in gel electrophoresis. Computer Physics Communications, 147(1–2):354–357, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550200303X.

Kaya:2004:RAF

[KKS04] H. Kaya, M. Kaplan, and H. Saygin. A recursive algorithm for finding HDMR terms for sensitivity analysis. Computer Physics Communications, 158(2):106–112, April 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504000232.

Kholodov:2004:NSC

[KKSR04] Y. A. Kholodov, A. S. Kholodov, E. L. Stupitzki, and A. Y. Repin. Numerical simulation of the convective plasma dynamics stage at the ionosphere motion by means of 3D MHD equations. Computer Physics Communications, 164(1–3):91–97, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002620.

Kim:2001:ERE

[KL01] Jai Sam Kim and Kyungsoo Lee. The earth regeneration effect of solar neutrinos: a numerical treatment with three active neutrino flavors. Computer Physics Communications, 135(2):176– 189, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000237X.

Kleiss:2006:EMC

[KL06] Ronald Kleiss and Achilleas Lazopoulos. Error in Monte Carlo, quasi-error in quasi-Monte Carlo. Computer Physics Communications, 175(2):93–115, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506001019.

Kim:2007:SDD

[KL07a] Hyung Jinn Kim and Jong Suk Ruth Lee. A study on a data dissemination mechanism for grid environment. Computer Physics Communications, 177(1–2):216, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000768.

Kim:2007:P

[KL07b] Jai Sam Kim and Jae Koo Lee. Preface. Computer Physics Communications, 177(1–2):xi–xii, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001531.

Kniep:2004:GPC

[KLD04] James C. Kniep, Jean-Noel G. Leboeuf, and Viktor K. Decyk. Gyrokinetic particle-in-cell calculations of ion temperature gradient driven turbulence with parallel nonlinearity and strong flow corrections. Computer Physics Communications, 164(1–3):98–102, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002644.

Kim:2000:OOC

[KLM00] Yong-Hoon Kim, In-Ho Lee, and Richard M. Martin. Objectoriented construction of a multigrid electronic-structure code with Fortran 90. Computer Physics Communications, 131 (1–2):10–25, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000564.

Kozin:2004:NVR

[KLTH04] Igor N. Kozin, Mark M. Law, Jonathan Tennyson, and Jeremy M. Hutson. New vibration-rotation code for tetraatomic molecules exhibiting wide-amplitude motion: WAVR4. Computer Physics Communications, 163(2):117–131, November 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004096.

Kharraziha:2000:MAS

[KM00a] Hamid Kharraziha and Stefano Moretti. The Metropolis algorithm for on-shell four-momentum phase space. Computer Physics Communications, 127(2–3):242–260, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599005044. See erratum [KM01b].

Kolganova:2000:SRT

[KM00b] E. A. Kolganova and A. K. Motovilov. Scattering and resonances in the ^{4}He three-atomic system. Computer Physics Communications, 126(1–2):88–92, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004749.

Kawaguchi:2001:FCD

[KM01a] Takaaki Kawaguchi and Hiroshi Matsukawa. Friction and complex dynamics at solid surfaces. Computer Physics Communications, 142(1–3):100–104, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003368.

Kharraziha:2001:EMA

[KM01b] Hamid Kharraziha and Stefano Moretti. Erratum to: “The Metropolis algorithm for on-shell four-momentum phase space”: [Computer Physics Commun. 127 (2000) 242–260]. Computer Physics Communications, 134(1):136–138, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001831. See [KM00a].

Knoll:2001:IPI

[KM01c] P. Knoll and S. Mirzaei. Improved parallel implementation of the two-dimensional Ising model. Computer Physics Communications, 141(3):313–321, December 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004167.

Krivec:2001:NIQ

[KM01d] R. Krivec and V. B. Mandelzweig. Numerical investigation of quasilinearization method in quantum mechanics. Computer Physics Communications, 138(1):69–79, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001916.

Krivec:2003:QAQ

[KM03] R. Krivec and V. B. Mandelzweig. Quasilinearization approach to quantum mechanics. Computer Physics Communications, 152(2):165–174, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502008214.

Krajewski:2005:MBQ

[KM05] Florian R. Krajewski and Martin H. Müser. Many-body quantum dynamics by adiabatic path-integral molecular dynamics: Disordered Frenkel Kontorova models. Computer Physics Communications, 169(1–3):426–429, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001918.

Krivec:2006:QMW

[KM06] R. Krivec and V. B. Mandelzweig. Quasilinearization method and WKB. Computer Physics Communications, 174(2):119– 126, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505005102.

Kireev:2008:RTJ

[KM08a] Alexandre N. Kireev and Olivier J. F. Martin. Real-time Java simulations of multiple interference dielectric filters. Computer Physics Communications, 179(12):903–907, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002531.

Krivec:2008:QAC

[KM08b] R. Krivec and V. B. Mandelzweig. Quasilinearization approach to computations with singular potentials. Computer Physics Communications, 179(12):865–867, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002543.

Kreer:2002:FDB

[KMB02] T. Kreer, M. H. Müser, and K. Binder. Frictional drag between polymer bearing surfaces. Computer Physics Communications, 147(1–2):358–361, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003041.

Kwan:2001:ISS

[KMCS01] Thomas J. T. Kwan, Allen R. Mathews, Peggy J. Christenson, and Charles M. Snell. Integrated system simulation in X-ray radiography. Computer Physics Communications, 142(1–3):263–269, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003381.

Kamieniarz:2002:LSS

[KMD^{+}02] G. Kamieniarz, R. Matysiak, A. Caramico D’Auria, F. Esposito, and C. Benelli. Large-scale simulations of the finitetemperature properties of the molecular assemblies Mn_{6} and Ni_{12}. Computer Physics Communications, 147(1–2):194–197, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002436.

Knuteson:2002:PNM

[KMH02] B. Knuteson, H. E. Miettinen, and L. Holmström. α → PDE: a new multivariate technique for parameter estimation. Computer Physics Communications, 145(3):351–356, June 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002795.

Krein:2009:DDT

[KMP09] G. Krein, J. M. Machado, and A. O. Pereira. Dynamics of the deconfinement transition of quarks and gluons in a finite volume. Computer Physics Communications, 180(4):564– 573, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000058.

Kroetz:2009:TCF

[KMR^{+}09] T. Kroetz, F. A. Marcus, M. Roberto, I. L. Caldas, R. L. Viana, and E. C. da Silva. Transport control in fusion plasmas by changing electric and magnetic field spatial profiles. Computer Physics Communications, 180(4):642–650, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004426.

Kalogiratou:2009:CES

[KMS09] Z. Kalogiratou, Th. Monovasilis, and T. E. Simos. Computation of the eigenvalues of the Schrödinger equation by exponentially-fitted Runge–Kutta–Nyström methods. Computer Physics Communications, 180(2):167–176, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002993.

Kioutsioukis:2005:ESC

[KMZZ05] Ioannis Kioutsioukis, Dimitrios Melas, Christos Zerefos, and Ioannis Ziomas. Efficient sensitivity computations in 3D air quality models. Computer Physics Communications, 167 (1):23–33, April 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550400565X.

Krstic:2007:IFO

[KN07a] D. Krstić and D. Nikezić. Input files with ORNL-mathematical phantoms of the human body for MCNP-4B. Computer Physics Communications, 176(1):33–37, January 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002943.

Kubota:2007:ENM

[KN07b] Y. Kubota and K. Nobusada. An efficient numerical method for exciton states in quantum boxes. Computer Physics Communications, 177(1–2):43, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001294.

Kuksin:2007:SMS

[KNSY07a] A. Yu. Kuksin, G. E. Norman, V. V. Stegailov, and A. V. Yanilkin. Surface melting of superheated crystals. Atomistic simulation study. Computer Physics Communications, 177(1–2):34–37, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001440.

Kuksin:2007:ASS

[KNSY07b] Alexey Kuksin, Genri Norman, Vladimir Stegailov, and Alexey Yanilkin. Atomistic simulations of structure transitions and fracture in Fe and Al single crystals. Computer Physics Communications, 177(1–2):48, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001300.

Karasozen:2008:SGD

[KNT08] B. Karasözen, A. D. Nemtsev, and V. G. Tsybulin. Staggered grids discretization in three-dimensional Darcy convection. Computer Physics Communications, 178(12):885–893, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000787.

Keppens:2003:AMR

[KNTG03] R. Keppens, M. Nool, G. Tóth, and J. P. Goedbloed. Adaptive mesh refinement for conservative systems: multi-dimensional efficiency evaluation. Computer Physics Communications, 153(3):317–339, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503001395.

Klimenko:2000:MNA

[KNU00] S. V. Klimenko, I. N. Nikitin, and W. F. Urazmetov. Methods of numerical analysis of 1-dimensional 2-body problem in Wheeler–Feynman electrodynamics. Computer Physics Communications, 126(1–2):82–87, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599005330.

Kim:2005:SPM

[KNY05] Kang Kim, Yasuya Nakayama, and Ryoichi Yamamoto. A smoothed profile method for simulating charged colloidal dispersions. Computer Physics Communications, 169(1–3):104– 106, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001219.

Kocbach:2002:BRB

[Koc02] L. Kocbach. Book review: Computer Algebra Recipes: a Gourmet’s Guide to the Mathematical Models of Science: Richard H. Enns and George C. McGuire, Springer, New York, 2001. Computer Physics Communications, 144(3):343–346, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001935.

Kolodziej:2003:EVP

[Kol03] Karol Kolodziej. eett6f v. 1.0: a program for top quark pair production and decay into 6 fermions at linear colliders. Computer Physics Communications, 151(3):339–353, April 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007385.

Kolodziej:2009:CPA

[Kol09] Karol Kolodziej. carlomat: a program for automatic computation of lowest order cross sections. Computer Physics Communications, 180(9):1671–1680, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001076.

Kontar:2001:NCQ

[Kon01] Eduard P. Kontar. Numerical consideration of quasilinear electron cloud dynamics in plasma. Computer Physics Communications, 138(3):222–233, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002144.

Kondo:2002:GPR

[Kon02] T. Kondo. GFACTOR2001: a program for relativistic atomic g-factor calculations. Computer Physics Communications, 146(2):261–270, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002643.

Kostoglou:2005:CKS

[Kos05] Margaritis Kostoglou. On the constant kernel Smoluchowski equation: fast algorithm for solution with arbitrary initial conditions. Computer Physics Communications, 173(1–2):34–40, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004248.

Kumpula:2009:MCE

[KOS^{+}09] J. M. Kumpula, J.-P. Onnela, J. Saramäki, J. Kertész, and K. Kaski. Model of community emergence in weighted social networks. Computer Physics Communications, 180(4):517– 522, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004347.

Kozlowski:2002:NVL

[Koz02] Wlodzimierz Kozlowski. A note on varying lattice isotropy. Computer Physics Communications, 143(2):155–161, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004441.

Kanaki:2000:HPC

[KP00] Aggeliki Kanaki and Costas G. Papadopoulos. HELAC: a package to compute electroweak helicity amplitudes. Computer Physics Communications, 132(3):306–315, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550000151X.

Kempf:2001:EAS

[KP01] Sascha Kempf and Susanne Pfalzner. An effective algorithm for simulating diffusion-driven aggregation. Computer Physics Communications, 137(2):225–235, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001503.

Kalatzis:2006:POE

[KPD06] Fanis G. Kalatzis, Dimitrios G. Papageorgiou, and Ioannis N. Demetropoulos. A programmable optimization environment using the GAMESS-US and MERLIN/MCL packages. Applications on intermolecular interaction energies. Computer Physics Communications, 175(5):359–371, September 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001925.

Klepeis:2003:NCH

[KPF03] J. L. Klepeis, M. J. Pieja, and C. A. Floudas. A new class of hybrid global optimization algorithms for peptide structure prediction: integrated hybrids. Computer Physics Communications, 151(2):121–140, March 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200735X.

Kang:2007:PCM

[KPL07] Jonghyuk Kang, Ganyoung Park, and Jaekoo Lee. Particle-In-Cell/Monte Carlo simulation of CF_{4} discharge in capacitively coupled plasma reactor. Computer Physics Communications, 177(1–2):129–130, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001014.

Kholmurodov:2001:MSC

[KPS^{+}01] Kholmirzo Kholmurodov, Igor Puzynin, William Smith, Kenji Yasuoka, and Toshikazu Ebisuzaki. MD simulation of cluster-surface impacts for metallic phases: soft landing, droplet spreading and implantation. Computer Physics Communications, 141(1):1–16, November 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002922.

Kroger:2005:SMD

[Krö05] Martin Kröger. Shortest multiple disconnected path for the analysis of entanglements in two- and three-dimensional polymeric systems. Computer Physics Communications, 168(3): 209–232, June 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505002808.

Knaup:2002:WPM

[KRTZ02] M. Knaup, P.-G. Reinhard, C. Toepffer, and G. Zwicknagel. Wave packet molecular dynamics simulations of hydrogen at mbar pressures. Computer Physics Communications, 147(1–2):202–204, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550200245X.

Kersevan:2003:MCE

[KRW03] B. P. Kersevan and E. Richter-Was. The Monte Carlo Event Generator AcerMC version 1.0 with interfaces to PYTHIA 6.2 and HERWIG 6.3. Computer Physics Communications, 149(3):142–194, January 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502005921.

Kolbig:1984:PPL

[KS84] K. S. Kölbig and B. Schorr. A program package for the Landau distribution. Computer Physics Communications, 31(1):97– 111, January 1984. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0010465584900857. See erratum

[KS08].

Krupski:2001:PLA

[KS01] Marcin Krupski and Pawel B. Sczaniecki. Phase lag in the amplitude modulated magnetic resonance. Computer Physics Communications, 138(2):170–174, August 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002077.

Korolkov:2004:QMD

[KS04a] Mikhail V. Korolkov and Burkhard Schmidt. Quantum molecular dynamics driven by short and intense light pulses: Towards the limits of the Floquet picture. Computer Physics Communications, 161(1–2):1–17, August 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002061.

Kreuzer:2004:PPA

[KS04b] Maximilian Kreuzer and Harald Skarke. PALP: a package for analysing lattice polytopes with applications to toric geometry. Computer Physics Communications, 157(1):87–106, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004910.

Kalmykov:2005:LCL

[KS05] M. Yu. Kalmykov and A. Sheplyakov. lsjk — a C++ library for arbitrary-precision numeric evaluation of the generalized log-sine functions. Computer Physics Communications, 172(1):45–59, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505003139.

Kadamani:2007:UES

[KS07] S. Kadamani and A. D. Snider. USFKAD: An expert system for partial differential equations. Computer Physics Communications, 176(1):62–69, January 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003638.

Kolbig:2008:EPP

[KS08] K. S. Kölbig and B. Schorr. Erratum to: “A program package for the Landau distribution” [Comput. Phys. Commun. 31 (1984) 97–111]. Computer Physics Communications, 178(12): 972, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001136. See [KS84].

Kandhai:2000:FDL

[KSC^{+}00] D. Kandhai, W. Soll, S. Chen, A. Hoekstra, and P. Sloot. Finite-Difference Lattice-BGK methods on nested grids. Computer Physics Communications, 129(1–3):100–109, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000977.

Kremer:2005:EPS

[KSEG05] Kurt Kremer, Sathish K. Sukumaran, Ralf Everaers, and Gary S. Grest. Entangled polymer systems. Computer Physics Communications, 169(1–3):75–81, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001165.

King:2002:NEH

[KSHP02] Frederick W. King, Gregory J. Smethells, Geir T. Helleloid, and Paul J. Pelzl. Numerical evaluation of Hilbert transforms for oscillatory functions: a convergence accelerator approach. Computer Physics Communications, 145(2):256–266, May 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001558.

Kim:2004:HKM

[KSPT04] Charlson C. Kim, Carl R. Sovinec, Scott E. Parker, and The NIMROD Team. Hybrid kinetic-MHD simulations in general geometry. Computer Physics Communications, 164(1–3):448–455, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003170.

Karsch:2002:CFD

[KSS02] F. Karsch, Ch. Schmidt, and S. Stickan. Common features of deconfining and chiral critical points in QCD and the three state Potts model in an external field. Computer Physics Communications, 147(1–2):451–454, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003272.

Krauss:2006:APC

[KSS06] F. Krauss, A. Schälicke, and G. Soff. APACIC++ 2.0. A PArton Cascade In C++. Computer Physics Communications, 174(11):876–902, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506000099.

Kruger:2004:FBS

[KSSH04] S. E. Kruger, C. R. Sovinec, D. D. Schnack, and E. D. Held. Free-boundary simulations of DIII-D plasmas with the NIMROD code. Computer Physics Communications, 164(1– 3):34–39, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550400253X.

Kosarev:2003:DPS

[KSTL03] E. L. Kosarev, A. Ya. Shul’man, M. A. Tarasov, and T. Lindström. Deconvolution problems and superresolution in hilberttransform spectroscopy based on a.c. Josephson effect. Computer Physics Communications, 151(2):171–186, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007014.

Kholmurodov:2000:HVL

[KSYE00] Kholmirzo Kholmurodov, William Smith, Kenji Yasuoka, and Toshikazu Ebisuzaki. A highly vectorised “link-cell” FOR-TRAN code for the DL POLY molecular dynamics simulation package. Computer Physics Communications, 125(1–3):167– 192, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004853.

Katz:2004:LSO

[KT04] S. D. Katz and B. C. Tóth. Least-squared optimized polynomials for smeared link actions. Computer Physics Communications, 158(3):137–149, April 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000736.

Karasozen:2005:MDT

[KT05] B. Karasözen and V. G. Tsybulin. Mimetic discretization of two-dimensional Darcy convection. Computer Physics Communications, 167(3):203–213, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505000652.

Kuhn:2007:MPW

[KT07] S. Kuhn, D. D. Tskhakaya, and D. Tskhakaya (Jr. ). The magnetized plasma-wall transition (PWT) and its relation to fluid boundary conditions. Computer Physics Communications, 177(1–2):80–83, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001154.

Kisiel:2006:TTH

[KTBF06] Adam Kisiel, Tomasz Taluć, Wojciech Broniowski, and Wojciech Florkowski. THERMINATOR: THERMal heavy-IoN generATOR. Computer Physics Communications, 174(8): 669–687, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000087.

Korsun:2004:MMH

[KTG04a] A. G. Korsun, E. M. Tverdokhlebova, and F. F. Gabdullin. Mathematical model of hypersonic plasma flows expanding in vacuum. Computer Physics Communications, 164(1–3):434–441, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504003157.

Korsun:2004:SPP

[KTG04b] A. G. Korsun, E. M. Tverdokhlebova, and F. F. Gabdullin. Simulation of plasma plume-to-spacecraft interaction. Computer Physics Communications, 164(1–3):353–364, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003030.

Kozin:2005:ECM

[KTL05] Igor N. Kozin, Jonathan Tennyson, and Mark M. Law. Effective computation of matrix elements between polynomial basis functions. Computer Physics Communications, 165(1):10–14, January 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004394.

Kunishima:2002:EER

[KTT02] Wataru Kunishima, Tetsuji Tokihiro, and Hiroshi Tanaka. Error estimation of recursive orthogonal polynomial expansion method for large Hamiltonian system. Computer Physics Communications, 148(2):171–181, October 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005507.

Kakhiani:2009:PGB

[KTT09] Khatuna Kakhiani, Kakha Tsereteli, and Paata Tsereteli. A program to generate a basis set adaptive radial quadrature grid for density functional theory. Computer Physics Communications, 180(2):256–268, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003330.

Kudryavtsev:2009:MSC

[Kud09] V. A. Kudryavtsev. Muon simulation codes MUSIC and MUSUN for underground physics. Computer Physics Communications, 180(3):339–346, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0010465508003640.

Kurkina:2002:DFS

[Kur02] L. I. Kurkina. Density-functional study of multielectron ionization of sodium clusters by strong femtosecond laser pulses. Computer Physics Communications, 147(1–2):205–208, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002461.

Kim:2007:RAT

[KV07] H. C. Kim and J. P. Verboncoeur. Reflection, absorption and transmission of TE electromagnetic waves propagation in a nonuniform plasma slab. Computer Physics Communications, 177(1–2):118–121, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001257.

Kalafut:2008:OMI

[KV08] Bennett Kalafut and Koen Visscher. An objective, modelindependent method for detection of non-uniform steps in noisy signals. Computer Physics Communications, 179(10): 716–723, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508002282.

Kravanja:2000:ZMS

[KVR^{+}00] P. Kravanja, M. Van Barel, O. Ragos, M. N. Vrahatis, and F. A. Zafiropoulos. ZEAL: a mathematical software package for computing zeros of analytic functions. Computer Physics Communications, 124(2–3):212–232, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004294.

Krogh:2003:CEH

[KW03] J. W. Krogh and M. Wierzbowska. The condensation energy of the homogeneous superconducting gas. Computer Physics Communications, 150(1):53–63, January 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006148.

Kiselev:2007:CIS

[KW07] A. V. Kiselev and T. Wolf. Classification of integrable supersystems using the sstools environment. Computer Physics Communications, 177(3):315–328, August 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002391.

Kuhn:2008:PCS

[KW08] Markus Kuhn and Hans Walliser. Program for calculating SU(4) Clebsch–Gordan coefficients. Computer Physics Communications, 179(10):733–740, November 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002300.

Wong:2001:MCC

[kWpLwW01] Wai kit Wong, Lap piu Lee, and Kwok wo Wong. A modified chaotic cryptographic method. Computer Physics Communications, 138(3):234–236, August 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100220X. See comment and reply

[ÁMRP04, WLW04].

Kim:2007:TSS

[KY07] S. S. Kim and S. W. Yoon. Transport simulations for the scrape-off layer and divertor plasmas in KSTAR tokamak. Computer Physics Communications, 177(1–2):70–71, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001129.

Kostomarov:2000:PET

[KZS^{+}00] D. P. Kostomarov, F. S. Zaitsev, A. G. Shishkin, D. C. Robinson, M. R. O’Brien, and M. Gryaznevich. The problem of evolution of toroidal plasma equilibria. Computer Physics Communications, 126(1–2):101–106, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002283.

Loscar:2009:HEF

[LA09] Ernesto S. Loscar and Ezequiel V. Albano. Hysteretic effects in the first-order irreversible phase transition of the ZGB model. Computer Physics Communications, 180(4):488–492, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004165.

Ladd:2009:FRN

[Lad09] Anthony J. C. Ladd. A fast random number generator for stochastic simulations. Computer Physics Communications, 180(11):2140–2142, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001994.

Lozovoi:2001:SEE

[LAF01] A. Y. Lozovoi, A. Alavi, and M. W. Finnis. Surface energy and the early stages of oxidation of nial(110). Computer Physics Communications, 137(1):174–194, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001783.

Laforest:2003:BRB

[Laf03] M. Laforest. Book review: Front Tracking for Hyperbolic Conservation Laws: by H. Holden and N.H. Risebro (Springer, Berlin, 2002) ISBN 3-540-43289-2. Computer Physics Communications, 155(2):181–182, October 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503003461.

Limbach:2006:EES

[LAMH06] H. J. Limbach, A. Arnold, B. A. Mann, and C. Holm. ESPResSo — an extensible simulation package for research on soft matter systems. Computer Physics Communications, 174(9):704–727, May 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550500576X.

Landau:2007:CPE

[Lan07] Rubin H. Landau. Computational physics education; why, what and how. Computer Physics Communications, 177 (1–2):191–194, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000719.

Likhachev:2001:OCC

[LANM^{+}01] V. P. Likhachev, J. D. T. Arruda-Neto, M. N. Martins, F. Garcia, J. Mesa, O. Rodriguez, V. B. Shostak, I. Evseev, S. A.

Paschuk, H. R. Schelin, and V. V. Denyak. OMEGA code: calculation of the geometrical parameters of detection systems. Computer Physics Communications, 135(2):190–198, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002307.

Linna:2004:UPB

[LÅT04] R. P. Linna, J. A. Åström, and J. Timonen. Unconstrained periodic boundary conditions for solid state elasticity. Computer Physics Communications, 158(1):26–38, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503005435.

Levitina:2000:SEE

[LB00] Tatiana Levitina and Erkki J. Brändas. On the Schrödinger equation in ellipsoidal coordinates. Computer Physics Communications, 126(1–2):107–113, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046559900449X.

Liang:2004:PSP

[LB04] Heng Liang and Fengshan Bai. A partially structure-preserving algorithm for the permanents of adjacency matrices of fullerenes. Computer Physics Communications, 163(2):79–84, November 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004163.

Levitina:2009:FDF

[LB09] T. Levitina and E. J. Brändas. Filter diagonalization: Filtering and postprocessing with prolates. Computer Physics Communications, 180(9):1448–1457, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000861.

Levy:2005:DCM

[LBM05] Nicolas Levy, Daniel Borgis, and Massimo Marchi. A dielectric continuum model of solvation for complex solutes. Computer Physics Communications, 169(1–3):69–74, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001153.

Legrand:2001:MTD

[LbotMC01] Iosif Legrand and On behalf of the MONARC Collaboration. Multi-threaded, discrete event simulation of distributed computing systems. Computer Physics Communications, 140(1–2):274–285, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002818.

Lee:2009:HHA

[LBP^{+}09] JunKyu Lee, Yu Bi, Gregory D. Peterson, Robert J. Hinde, and Robert J. Harrison. HASPRNG: Hardware Accelerated Scalable Parallel Random Number Generators. Computer Physics Communications, 180(12):2574–2581, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002100.

Laveder:2009:SIS

[LBPS09] D. Laveder, D. Borgogno, T. Passot, and P. L. Sulem. On a semi-implicit scheme for spectral simulations of dispersive magnetohydrodynamics. Computer Physics Communications, 180(10):1860–1869, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001672.

Lima:2000:NSS

[LC00] P. M. Lima and M. Carpentier. Numerical solution of a singular boundary-value problem in non-Newtonian fluid mechanics. Computer Physics Communications, 126(1–2):114–120, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004221.

Lu:2001:IPP

[LC01a] Quan Ming Lu and Dong Sheng Cai. Implementation of parallel plasma particle-in-cell codes on PC cluster. Computer Physics Communications, 135(1):93–104, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002277.

Lupu:2001:CCV

[LC01b] A. Lupu and V. Cuculeanu. Code for calculating the vertical distribution of radon isotopes and their progeny in the atmosphere. Computer Physics Communications, 141 (1):149–162, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501002995.

Luijten:2007:TPC

[LC07] Erik Luijten and Angelo Cacciuto. Translocation of polymers out of confined geometries. Computer Physics Communications, 177(1–2):150–153, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000902.

Li:2008:EPF

[LC08a] Yan Feng Li and Kui Fu Chen. Eliminating the picket fence effect of the fast Fourier transform. Computer Physics Communications, 178(7):486–491, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507004663.

Li:2008:NSF

[LC08b] Yiming Li and Hui-Wen Cheng. Numerical simulation of field emission efficiency of anodic aluminum oxide carbon nanotube field emitter in the triode structure. Computer Physics Communications, 179(1–3):107–111, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000246.

Lagana:2000:CGP

[LCB^{+}00] Antonio Laganà, Stefano Crocchianti, Alessandro Bolloni, Valentina Piermarini, Ranieri Baraglia, Renato Ferrini, and Domenico Laforenza. Computational granularity and parallel models to scale up reactive scattering calculations. Computer Physics Communications, 128(1–2):295–314, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005160.

Lorin:2007:NMS

[LCB07] E. Lorin, S. Chelkowski, and A. Bandrauk. A numerical Maxwell–Schrödinger model for intense laser-matter interaction and propagation. Computer Physics Communications, 177(12):908–932, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003529.

Lee:2009:CIC

[LCE^{+}09] J. S. Lee, M. Carena, J. Ellis, A. Pilaftsis, and C. E. M. Wagner. CPsuperH2.0: An improved computational tool for Higgs phenomenology in the MSSM with explicit CP violation. Computer Physics Communications, 180(2):312–331, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003032.

Li:2009:SBL

[LCHJ09] Z.-C. Li, C.-S. Chien, H.-T. Huang, and B.-W. Jeng. Superconvergence of bi-k-Lagrange elements for eigenvalue problems. Computer Physics Communications, 180(11):2268–2282, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002355.

Lignieres:2000:NSM

[LCM00] François Lignières, Francesco Califano, and André Mangeney. Numerical simulations of a mixed layer in a stably stratified fluid. Computer Physics Communications, 132(3):197–223, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001272.

Lee:2004:ECB

[LCPC04] Po-Han Lee, Yi Chen, Soo-Chang Pei, and Yih-Yuh Chen. Evidence of the correlation between positive Lyapunov exponents and good chaotic random number sequences. Computer Physics Communications, 160(3):187–203, July 15, 2004.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001663.

Lee:2007:TDF

[LCS07] Sunghee Lee, Sunmi Choi, and Jongwon Shon. Twodimensional fluid simulation of VHF-ICP source with parallel resonance antenna. Computer Physics Communications, 177(1–2):133, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000987.

Liu:2006:MSQ

[LCV06] Jinn-Liang Liu, Jen-Hao Chen, and O. Voskoboynikov. A model for semiconductor quantum dot molecule based on the current spin density functional theory. Computer Physics Communications, 175(9):575–582, November 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002748.

Lopez-Duran:2008:ASS

[LDBG08] D. López-Durán, E. Bodo, and F. A. Gianturco. AS-PIN: An a ll spin scattering code for atom-molecule rovibrationally inelastic cross sections. Computer Physics Communications, 179(11):821–838, December 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002464.

Lefebvre:2007:NMA

[LdG^{+}07] E. Lefebvre, E. d’Humières, L. Gremillet, S. Fritzler, and V. Malka. Numerical modeling and applications of laseraccelerated ion beams. Computer Physics Communications, 177(1–2):60–63, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001245.

LeBourdiec:2006:NSV

[LdVJ06] S. Le Bourdiec, F. de Vuyst, and L. Jacquet. Numerical solution of the Vlasov–Poisson system using generalized Hermite functions. Computer Physics Communications, 175(8):528–544, October 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002670.

Liu:2008:SHP

[LDZ^{+}08] Lei Liu, Jianguo Du, Jijun Zhao, Hong Liu, Di Wu, and Fuliang Zhao. Study of high-pressure and high-temperature behaviors and α-to-β phase transition of forsterite by firstprinciples and quasi-harmonic Debye model. Computer Physics Communications, 179(6):417–423, September 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001422.

Lee:2004:TNP

[Lee04] W. W. Lee. Theoretical and numerical properties of a gyrokinetic plasma: issues related to transport time scale simulation. Computer Physics Communications, 164(1–3): 244–250, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002875.

Liu:2002:NIT

[LEG02] H. Liu, Ch. Elster, and W. Glöckle. Numerical implementation of three-body forces in bound state Faddeev calculations in three dimensions. Computer Physics Communications, 147(1–2):170–173, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002394.

Lehner:2000:GRB

[Leh00] Luis Lehner. Gravitational radiation, black holes and the characteristic formulation of G.R. Computer Physics Communications, 127(1):43–52, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000199.

Leimkuhler:2002:SFN

[Lei02] Ben Leimkuhler. A separated form of Nosé dynamics for constant temperature and pressure simulation. Computer Physics Communications, 148(2):206–213, October 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005544.

Lewandowski:2004:CSA

[Lew04] J. L. V. Lewandowski. Characterization of strange attractors in drift wave microturbulence. Computer Physics Communications, 164(1–3):114–117, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400267X.

Landau:2002:P

[LF02a] David P. Landau and Fereydoon Family. Preface. Computer Physics Communications, 146(1):x–xi, June 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004277.

Los:2002:MCS

[LF02b] J. H. Los and A. Fasolino. Monte Carlo simulations of carbonbased structures based on an extended brenner potential. Computer Physics Communications, 147(1–2):178–181, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002400.

Liaqat:2001:ANN

[LFT01] Ali Liaqat, Makoto Fukuhara, and Tatsuoki Takeda. Application of neural network collocation method to data assimilation. Computer Physics Communications, 141(3):350–364, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004313.

Liaqat:2003:OEP

[LFT03] Ali Liaqat, Makoto Fukuhara, and Tatsuoki Takeda. Optimal estimation of parameters of dynamical systems by neural network collocation method. Computer Physics Communications, 150(3):215–234, February 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200680X.

Lythe:2001:SPC

[LH01] Grant Lythe and Salman Habib. Stochastic PDEs: convergence to the continuum? Computer Physics Communications, 142(1–3):29–35, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003083.

Limbach:2002:CPP

[LH02] Hans Jörg Limbach and Christian Holm. Conformational properties of poor solvent polyelectrolytes. Computer Physics Communications, 147(1–2):321–324, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002953.

Li:2003:NNA

[LH03] Yiming Li and Kuen-Yu Huang. A novel numerical approach to heterojunction bipolar transistors circuit simulation. Computer Physics Communications, 152(3):307–316, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502008494.

Langridge:2001:EST

[LHC01] D. J. Langridge, J. F. Hart, and S. Crampin. Ewald summation technique for one-dimensional charge distributions. Computer Physics Communications, 134(1):78–85, February 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001922. See comments

and reply [Har02, LHC02].

Langridge:2002:RCE

[LHC02] D. J. Langridge, J. F. Hart, and S. Crampin. Reply to comment on “Ewald summation technique for one-dimensional charge distributions”. Computer Physics Communications, 146(2):274–275, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004538.

See [LHC01, Har02].

Loveall:2000:SCE

[LHMB00] D. Loveall, M. B. Hamley, B. J. Miller, and K. Bartschat. Simulation of collisionally excited atomic states. Computer Physics Communications, 124(1):90–103, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004270.

Lian:2006:PAM

[LHS^{+}06] Y.-Y. Lian, K.-H. Hsu, Y.-L. Shao, Y.-M. Lee, Y.-W. Jeng, and J.-S. Wu. Parallel adaptive mesh-refining scheme on a three-dimensional unstructured tetrahedral mesh and its applications. Computer Physics Communications, 175(11–12): 721–737, December 1–15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506001962.

Lee:2009:SBA

[LHS^{+}09] Seung Min Lee, Yong Jun Hong, Young Sik Seo, Felipe Iza, Gyoo Cheon Kim, and Jae Koo Lee. Simulations of biomedical atmospheric-pressure discharges. Computer Physics Communications, 180(4):636–641, April 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000083.

Li:2003:PMI

[Li03] Yiming Li. A parallel monotone iterative method for the numerical solution of multi-dimensional semiconductor Poisson equation. Computer Physics Communications, 153(3):359–372, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002030.

Likas:2001:PDE

[Lik01] Aristidis Likas. Probability density estimation using artificial neural networks. Computer Physics Communications, 135(2):167–175, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002356.

Ledoux:2006:SSEb

[LIR^{+}06] V. Ledoux, L. Gr. Ixaru, M. Rizea, M. Van Daele, and G. Vanden Berghe. Solution of the Schrödinger equation over an infinite integration interval by perturbation methods, revisited. Computer Physics Communications, 175(9):612–619, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002761.

Liu:2007:NAD

[Liu07a] Kuo-Chi Liu. Numerical analysis of dual-phase-lag heat transfer in a layered cylinder with nonlinear interface boundary conditions. Computer Physics Communications, 177(3):307–314, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002020.

Liu:2007:CPR

[Liu07b] Yueqiang Liu. Constructing plasma response models from full toroidal magnetohydrodynamic computations. Computer Physics Communications, 176(3):161–169, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003626.

Lemoine:2001:QWP

[LJ01] Didier Lemoine and Bret Jackson. Quantum wave packet approach to the Eley–Rideal reactive scattering between a gas phase atom and an adsorbate. Computer Physics Communications, 137(3):415–426, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001655.

Liang:2008:ACT

[LJ08] Songxin Liang and David J. Jeffrey. Automatic computation of the travelling wave solutions to nonlinear PDEs. Computer Physics Communications, 178(9):700–712, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004948.

Liang:2009:EAA

[LJ09a] Songxin Liang and David J. Jeffrey. An efficient analytical approach for solving fourth order boundary value problems. Computer Physics Communications, 180(11):2034–2040, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001878.

Liang:2009:NTW

[LJ09b] Songxin Liang and David J. Jeffrey. New travelling wave solutions to modified CH and DP equations. Computer Physics Communications, 180(9):1429–1433, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000708.

Lee:2007:CTB

[LJY07] Geunjung Lee, Jinwoo Jeong, and Young-Gui Yoon. Comparison of trajectory based methods to find transition states. Computer Physics Communications, 177(1–2):52–53, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550700166X.

Lee:2007:AEA

[LK07] Geunsik Lee and Jai Sam Kim. Adsorption energies of Al, Ga, In, Tl on Sif(111)-7 × 7. Computer Physics Communications, 177(1–2):44, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001282.

Lin:2006:PMB

[LKC06] Wenbin Lin, Narayan Kovvali, and Lawrence Carin. Pseudospectral method based on prolate spheroidal wave functions for semiconductor nanodevice simulation. Computer Physics Communications, 175(2):78–85, July 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001068.

Lim:2007:DMH

[LKKK07] Gyuchang Lim, Soo Yoo Kim, Ji-Hyun Kang, and Kyungsik Kim. Dynamical models of high-frequency data analysis. Computer Physics Communications, 177(1–2):187, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000677.

Liljo:2008:ODE

[LKPH08] Jalo Liljo, Anupam Karmakar, A. Pukhov, and M. Hochbruck. One-dimensional electromagnetic relativistic PIC-hydrodynamic hybrid simulation code H-VLPL (hybrid virtual laser plasma lab). Computer Physics Communications, 179(6):371–379, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001379.

Lyubartsev:2000:MDM

[LL00] Alexander P. Lyubartsev and Aatto Laaksonen. M.dyna mix — a scalable portable parallel MD simulation package for arbitrary molecular mixtures. Computer Physics Communications, 128(3):565–589, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599005299.

Li:2004:RMP

[LL04] Zhi-Bin Li and Yin-Ping Liu. RAEEM: a Maple package for finding a series of exact traveling wave solutions for nonlinear evolution equations. Computer Physics Communications, 163(3):191–201, November 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004370.

Lee:2007:PPS

[LL07] Koo-Chul Lee and Julian Lee. Programming physics softwares in Flash. Computer Physics Communications, 177(1–2):195– 198, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000720.

Liu:2008:ABT

[LL08] Jian-Guo Liu and Ye-Zhou Li. Auto–Bäcklund transformation and exact solutions of the generalized variable-coefficient Kadomtsev–Petviashvili equation. Computer Physics Communications, 179(10):724–732, November 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002324.

Li:2001:NPA

[LLCS01] Yiming Li, Jinn-Liang Liu, Tien-Sheng Chao, and S. M. Sze. A new parallel adaptive finite volume method for the numerical simulation of semiconductor devices. Computer Physics Communications, 142(1–3):285–289, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003472.

Lee:2007:CNS

[LLH07] Kyoung Eun Lee, Jae Woo Lee, and Byoung Hee Hong. Complex networks in a stock market. Computer Physics Communications, 177(1–2):186, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000665.

Lough:2001:ETT

[LLLZ01] B. Lough, S. P. Lee, R. A. Lewis, and C. Zhang. Electronic thermal transport and thermionic cooling in semiconductor multi-quantum-well structures. Computer Physics Communications, 142(1–3):274–280, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003459.

Llopis:2008:CHI

[LLPL08] I. Llopis, M. Cosentino Lagomarsino, I. Pagonabarraga, and C. P. Lowe. Cooperativity and hydrodynamic interactions in externally driven semiflexible filaments. Computer Physics Communications, 179(1–3):150–154, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000325.

Li:2002:NSQ

[LLT^{+}02] Yiming Li, Jam-Wem Lee, Ting-Wei Tang, Tien-Sheng Chao, Tan-Fu Lei, and S. M. Sze. Numerical simulation of quantum effects in high-k gate dielectric MOS structures using quantum mechanical models. Computer Physics Communications, 147(1–2):214–217, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002485.

Li:2001:EEL

[LLV^{+}01] Yiming Li, Jinn-Liang Liu, O. Voskoboynikov, C. P. Lee, and S. M. Sze. Electron energy level calculations for cylindrical narrow gap semiconductor quantum dot. Computer Physics Communications, 140(3):399–404, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002910.

Lee:2007:QSA

[LLY07] Daewon Lee, Jaewook Lee, and Young-Gui Yoon. A quadratic string adapted barrier exploring method for locating transition states. Computer Physics Communications, 177(1–2): 218, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000537.

Lunney:2000:CBD

[LM00] T. F. Lunney and A. J. McCaughey. Component based distributed systems — CORBA and EJB in context. Computer Physics Communications, 127(2–3):207–214, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046559900510X.

Lakshminarasimhulu:2002:CMB

[LM02a] Pasupulati Lakshminarasimhulu and Jeffry D. Madura. A cell multipole based domain decomposition algorithm for molecular dynamics simulation of systems of arbitrary shape. Computer Physics Communications, 144(2):141–153, April 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502001571.

Lyulin:2002:LSC

[LM02b] Alexey V. Lyulin and M. A. J. Michels. Large-scale computer simulation of local segmental dynamics in amorphous atactic polystyrene. Computer Physics Communications, 147(1–2):298–302, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002916.

Li:2003:SCE

[LMC^{+}03] Shujun Li, Xuanqin Mou, Yuanlong Cai, Zhen Ji, and Jihong Zhang. On the security of a chaotic encryption scheme: problems with computerized chaos in finite computing precision. Computer Physics Communications, 153(1):52–58, June 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502008755.

Likos:2008:CFS

[LMM^{+}08] Christos N. Likos, Bianca M. Mladek, Angel J. Moreno, Dieter Gottwald, and Gerhard Kahl. Cluster-forming systems of ultrasoft repulsive particles: statics and dynamics. Computer Physics Communications, 179(1–3):71–76, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000180.

Lokhtin:2009:HIE

[LMP^{+}09] I. P. Lokhtin, L. V. Malinina, S. V. Petrushanko, A. M. Snigirev, I. Arsene, and K. Tywoniuk. Heavy ion event generator HYDJET++ (HYDrodynamics plus JETs). Computer Physics Communications, 180(5):779–799, May 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003913.

Lopez:2002:CSS

[LMS^{+}02] Carlos F. Lopez, Preston B. Moore, John C. Shelley, Mee Y. Shelley, and Michael L. Klein. Computer simulation studies of biomembranes using a coarse grain model. Computer Physics Communications, 147(1–2):1–6, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001959.

Loison:2005:FDL

[LMS05] Claire Loison, Michel Mareschal, and Friederike Schmid. Fluctuations and defects in lamellar stacks of amphiphilic bilayers. Computer Physics Communications, 169(1–3):99–103, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001207.

Lee:2001:FPS

[LN01] Young Joo Lee and R. M. Nieminen. First principles study of oxygen defects in silicon. Computer Physics Communications, 142(1–3):305–310, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003666.

Love:2003:SAF

[LNC^{+}03] P. J. Love, M. Nekovee, P. V. Coveney, J. Chin, N. González-Segredo, and J. M. R. Martin. Simulations of amphiphilic fluids using mesoscale lattice-Boltzmann and lattice-gas methods. Computer Physics Communications, 153(3):340–358, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002005.

Lee:2001:TCS

[LNK01] Young Joo Lee, R. M. Nieminen, and S. Kim. Theoretical and computational study of the interlayer relaxation of surfaces and nanocrystals. Computer Physics Communications, 142(1–3):414–417, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003782.

Lee:2001:UMB

[LNLK01] Young Joo Lee, R. M. Nieminen, Eok-Kyun Lee, and S. Kim. Universal melting behaviour of clusters. Computer Physics Communications, 142(1–3):201–205, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003319.

Legrand:2009:MAB

[LNV^{+}09] I. Legrand, H. Newman, R. Voicu, C. Cirstoiu, C. Grigoras, C. Dobre, A. Muraru, A. Costan, M. Dediu, and C. Stratan. MonALISA: An agent based, dynamic service system to monitor, control and optimize distributed systems. Computer Physics Communications, 180(12):2472–2498, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002410.

Lee:2005:SME

[LOCJ05] Hwee Kuan Lee, Yutaka Okabe, X. Cheng, and M. B. A. Jalil. Solving the master equation for extremely long time scale calculations. Computer Physics Communications, 168(3):159–164, June 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002274.

Lee:2006:CRW

[LOL06] Hwee Kuan Lee, Yutaka Okabe, and D. P. Landau. Convergence and refinement of the Wang–Landau algorithm. Computer Physics Communications, 175(1):36–40, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001081.

Longo:2007:KSN

[Lon07] S. Longo. Kinetic simulation of neutral and weakly ionized gases. Computer Physics Communications, 177(1–2):102– 105, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001233.

Loris:2008:LMP

[Lor08] Ignace Loris. L1Packv2: a Mathematica package for minimizing an ℓ_{1}-penalized functional. Computer Physics Communications, 179(12):895–902, December 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800252X.

Li:2007:DAP

[LOY07] Ming Li and Zhong-Can Ou-Yang. DNA as active polymer: long-range allosteric effect and chromatin loop structure. Computer Physics Communications, 177(1–2):176–179, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507000926.

Lee:2000:EPA

[LPC^{+}00] Jooyoung Lee, Jarosław Pillardy, Cezary Czaplewski, Yelena Arnautova, Daniel R. Ripoll, Adam Liwo, Kenneth D. Gibson, Ryszard J. Wawak, and Harold A. Scheraga. Efficient parallel algorithms in global optimization of potential energy functions for peptides, proteins, and crystals. Computer Physics Communications, 128(1–2):399–411, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005159.

Lee:2004:CCT

[LPC^{+}04] J. S. Lee, A. Pilaftsis, M. Carena, S. Y. Choi, M. Drees, J. Ellis, and C. E. M. Wagner. CPsuperH: a computational tool for Higgs phenomenology in the minimal supersymmetric standard model with explicit CP violation. Computer Physics Communications, 156(3):283–317, January 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004636.

Lednei:2004:CCM

[LPR04] M. F. Lednei, I. P. Pinkevich, and V. Yu Reshetnyak. Corrigendum to: “Computer modeling of light scattering in filled liquid crystals”: [Comput. Phys. Comm. 147 (2002) 362–365]. Computer Physics Communications, 156(3):318, January 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004582. See [LPRS02].

Lednei:2002:CML

[LPRS02] M. F. Lednei, I. P. Pinkevich, V. Yu. Reshetnyak, and T. J. Sluckin. Computer modeling of light scattering in filled liquid crystals. Computer Physics Communications, 147(1– 2):362–365, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502003053.

See corrigendum [LPR04].

Lorca:2006:ITL

[LR06] Alejandro Lorca and Tord Riemann. An integrated tool for loop calculations: aITALC. Computer Physics Communications, 174(1):71–82, January 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550500473X.

Leung:2007:EPI

[LR07] Mary Ann E. Leung and William P. Reinhardt. Efficient parallel implementation of Bose Hubbard model: Exact numerical ground states and dynamics of gaseous Bose–Einstein condensates. Computer Physics Communications, 177(4):348–356, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002238.

Ledoux:2006:SSEa

[LRI^{+}06] V. Ledoux, M. Rizea, L. Ixaru, G. Vanden Berghe, and M. Van Daele. Solution of the Schrödinger equation by a high order perturbation method based on a linear reference potential. Computer Physics Communications, 175(6):424–439, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600261X.

Lopez:2009:DPA

[LRR^{+}09] Rafael López, Jaime Fernández Rico, Guillermo Ramírez, Ignacio Ema, and David Zorrilla. DAMQT: a package for the analysis of electron density in molecules. Computer Physics Communications, 180(9):1654–1660, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000873.

Lugosi:2001:CME

[LS01] L. Lugosi and L. Sarkadi. Calculation of the matrix elements of the Coulomb interaction involving relativistic hydrogenic wave functions. Computer Physics Communications, 141(1):73–82, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002946.

Lange:2002:SAL

[LS02] Harald Lange and Friederike Schmid. Surface anchoring on liquid crystalline polymer brushes. Computer Physics Communications, 147(1–2):276–281, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002606.

Loverich:2005:DGM

[LS05] J. Loverich and U. Shumlak. A discontinuous Galerkin method for the full two-fluid plasma model. Computer Physics Communications, 169(1–3):251–255, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001554.

Luo:2009:CCG

[LS09]  Chuanfu Luo and Jens-Uwe Sommer.  Coding coarse

grained polymer model for LAMMPS and its application to polymer crystallization. Computer Physics Communications, 180(8):1382–1391, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000423.

Lee:2007:PBC

[LSL07] S. M. Lee, Y. S. Seo, and J. K. Lee. Paschen breakdown curve by one-dimensional PIC-MCC simulation. Computer Physics Communications, 177(1–2):132, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000999.

Liu:2008:AMD

[LSVMW08] Weiguo Liu, Bertil Schmidt, Gerrit Voss, and Wolfgang Müller-Wittig. Accelerating molecular dynamics simulations using graphics processing units with CUDA. Computer Physics Communications, 179(9):634–641, November 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002191.

Landau:2005:MCS

[LTA05] D. P. Landau, F. Tavazza, and J. Adler. Monte Carlo simulations of a compressible Ising ferromagnet at constant volume. Computer Physics Communications, 169(1–3):149–152, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001323.

Lee:2009:FPA

[LTG09] Y. C. Lee, H. M. Thompson, and P. H. Gaskell. FILM-PAR: a parallel algorithm designed for the efficient and accurate computation of thin film flow on functional surfaces containing micro-structure. Computer Physics Communications, 180(12):2634–2649, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001921.

Lin:2009:ESS

[LTT09] J.-Y. Lin, Y.-H. Tang, and M.-H. Tsai. Electronic structures of the Si(001) thin film under ⟨110⟩- and ⟨010⟩-direction uniaxial tensile strains. Computer Physics Communications, 180 (4):659–663, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004177.

Luding:2002:MSM

[Lud02] Stefan Luding. From microscopic simulations to macroscopic material behavior. Computer Physics Communications, 147(1–2):134–140, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002321.

Luo:2000:SRR

[Luo00] Li-Shi Luo. Some recent results on discrete velocity models and ramifications for lattice Boltzmann equation. Computer Physics Communications, 129(1–3):63–74, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000093X.

Luscher:2004:SDE

[Lüs04] Martin Lüscher. Solution of the Dirac equation in lattice QCD using a domain decomposition method. Computer Physics Communications, 156(3):209–220, January 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004867.

Luscher:2005:SPH

[Lüs05] Martin Lüscher. Schwarz-preconditioned HMC algorithm for two-flavor lattice QCD. Computer Physics Communications, 165(3):199–220, February 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005016.

Luthi:2000:MET

[Lüt00] Hans Peter Lüthi. Metacomputing, an emerging technology? Computer Physics Communications, 128(1–2):326–332, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005196.

Lutjens:2004:TSN

[Lüt04] H. Lütjens. Toroidal simulations of nonlinear thresholds and saturations of classical and neoclassical tearing instabilities. Computer Physics Communications, 164(1–3):286– 290, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002930.

Leherte:2008:CMP

[LV08] Laurence Leherte and Daniel P. Vercauteren. Collective motions in protein structures: Applications of elastic network models built from electron density distributions. Computer Physics Communications, 179(1–3):171–180, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000362.

Levine:2007:AFM

[LVH07] Zachary H. Levine, Alex Volkovitsky, and Howard K. Hung. Alignment of fiducial marks in a tomographic tilt series with an unknown rotation axis. Computer Physics Communications, 176(11–12):694–700, June 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550700183X.

Li:2001:CSE

[LVLS01] Yiming Li, O. Voskoboynikov, C. P. Lee, and S. M. Sze. Computer simulation of electron energy levels for different shape inas/gaas semiconductor quantum dots. Computer Physics Communications, 141(1):66–72, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003976.

Li:2002:CIE

[LVLS02] Yiming Li, O. Voskoboynikov, C. P. Lee, and S. M. Sze. Calculation of induced electron states in three-dimensional semiconductor artificial molecules. Computer Physics Communications, 147(1–2):209–213, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002473.

Ledoux:2004:CMH

[LVV04] V. Ledoux, M. Van Daele, and G. Vanden Berghe. CP methods of higher order for Sturm–Liouville and Schrödinger equations. Computer Physics Communications, 162(3):151–165, October 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003704.

Ledoux:2006:CME

[LVV06] V. Ledoux, M. Van Daele, and G. Vanden Berghe. CPM{P, N} methods extended for the solution of coupled channel Schrödinger equations. Computer Physics Communications, 174(5):357–370, March 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505005801.

Ledoux:2007:NPS

[LVV07] V. Ledoux, M. Van Daele, and G. Vanden Berghe. A numerical procedure to solve the multichannel Schrödinger eigenvalue problem. Computer Physics Communications, 176(3):191–199, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506003651.

Ledoux:2009:ECH

[LVV09] V. Ledoux, M. Van Daele, and G. Vanden Berghe. Efficient computation of high index Sturm–Liouville eigenvalues for problems in physics. Computer Physics Communications, 180(2):241–250, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003305.

Li:2007:NIU

[LWLL07] Y. W. Li, T. Wüst, D. P. Landau, and H. Q. Lin. Numerical integration using Wang–Landau sampling. Computer Physics Communications, 177(6):524–529, September 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507002950.

Landau:2008:CEB

[LWT08] D. P. Landau, Fugao Wang, and Shan-Ho Tsai. Critical endpoint behavior: a Wang–Landau study. Computer Physics Communications, 179(1–3):8–12, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000076.

Lo:2001:GAE

[LWY01] C. K. Lo, Jones T. K. Wan, and K. W. Yu. Geometric anisotropic effects on local field distribution: Generalized Clausius–Mossotti relation. Computer Physics Communications, 142(1–3):453–456, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003885.

Li:2005:NIM

[LY05] Yiming Li and Shao-Ming Yu. A numerical iterative method for solving Schrödinger and Poisson equations in nanoscale single, double and surrounding gate metal-oxide-semiconductor structures. Computer Physics Communications, 169(1–3):309–312, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001669.

Lee:2007:SPP

[LYL07] Hae June Lee, Hyun Jin Yoon, and Jae Koo Lee. Simulation of photons from plasmas for the applications to display devices. Computer Physics Communications, 177(1–2):106– 107, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001105.

Litvinenko:2000:SMN

[LZ00] E. I. Litvinenko and E. P. Zhidkov. Some methods of neutron scattering data analysis. Computer Physics Communications, 127(2–3):229–241, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004920.

Lorin:2004:RME

[LZ04] E. Lorin and G. Zérah. Recursion method for electronic structure calculations. Computer Physics Communications, 158(1):39–46, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503005459.

Li:2008:FAV

[LZC^{+}08] Ning Li, Hua-Xia Zhao, Sang-Hyun Cho, Jung-Gil Choi, and Myoung-Hee Kim. A fast algorithm for voxel-based deterministic simulation of X-ray imaging. Computer Physics Communications, 178(7):518–523, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550700464X.

Li:2006:MFD

[LZS06] Jianhui Li, Zhongwu Zhou, and Richard J. Sadus. Modified force decomposition algorithms for calculating three-body interactions via molecular dynamics. Computer Physics Communications, 175(11–12):683–691, December 1–15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002967.

Li:2008:PAM

[LZS08] Jianhui Li, Zhongwu Zhou, and Richard J. Sadus. Parallel algorithms for molecular dynamics with induction forces. Computer Physics Communications, 178(5):384–392, March 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004365.

Maslen:2000:ALF

[MA00] M. Maslen and P. Abbott. Automation of the lifting factorisation of wavelet transforms. Computer Physics Communications, 127(2–3):309–326, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004518.

Mohankumar:2004:UHO

[MA04] N. Mohankumar and Scott M. Auerbach. On the use of higherorder formula for numerical derivatives in scientific computing. Computer Physics Communications, 161(3):109–118, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001304.

Mohankumar:2006:TSB

[MA06] N. Mohankumar and Scott M. Auerbach. On time-step bounds in unitary quantum evolution using the Lanczos method. Computer Physics Communications, 175(7):473–481, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002682.

Mohankumar:2008:SCU

[MA08] N. Mohankumar and Scott M. Auerbach. Some comments on the use of higher-order formula for numerical derivatives in scientific computing. Computer Physics Communications, 179(11):773–776, December 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002336.

Muruganandam:2009:FPT

[MA09] P. Muruganandam and S. K. Adhikari. Fortran programs for the time-dependent Gross–Pitaevskii equation in a fully anisotropic trap. Computer Physics Communications, 180(10):1888–1912, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550900126X.

Maas:2006:SST

[Maa06] Axel Maas. Solving a set of truncated Dyson–Schwinger equations with a globally converging method. Computer Physics Communications, 175(3):167–179, August 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001111.

Much:2002:KMC

[MABK02] F. Much, M. Ahr, M. Biehl, and W. Kinzel. A kinetic Monte Carlo method for the simulation of heteroepitaxial growth. Computer Physics Communications, 147(1–2):226–229, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002515.

Mahdy:2008:ITC

[Mah08a] A. I. Mahdy. Interactions of two co-propagating laser beams in underdense plasmas using a generalized Peaceman–Rachford ADI form. Computer Physics Communications, 178(4):249– 271, February 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507004079.

Mahmoudi:2008:SPC

[Mah08b] F. Mahmoudi. SuperIso: a program for calculating the isospin asymmetry of B → K^{⋆}y in the MSSM. Computer Physics Communications, 178(10):745–754, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507005140.

Mahmoudi:2009:SVP

[Mah09a] F. Mahmoudi. SuperIso v2.3: a program for calculating flavor physics observables in supersymmetry. Computer Physics Communications, 180(9):1579–1613, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000721.

Mahmoudi:2009:SVF

[Mah09b] F. Mahmoudi. SuperIso v3.0, flavor physics observables calculations: extension to NMSSM. Computer Physics Communications, 180(9):1718–1719, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001489.

Maitre:2006:HMI

[Maî06] D. Maître. HPL, a Mathematica implementation of the harmonic polylogarithms. Computer Physics Communications, 174(3):222–240, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505005771.

Makino:2001:GPS

[Mak01] Junichiro Makino. GRAPE Project: special-purpose computers for many-body simulations. Computer Physics Communications, 139(1):45–54, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002296.

Maley:2000:CTE

[Mal00] David Maley. C++ toolkit for engineers and scientists. Computer Physics Communications, 128(3):637–638, June 2000.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000084.

Mohammed-Azizi:2004:SPC

[MAM04] B. Mohammed-Azizi and D. E. Medjadi. Single particle calculations for a Woods–Saxon potential with triaxial deformations, and large Cartesian oscillator basis. Computer Physics Communications, 156(3):241–282, January 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004648.

Mohammed-Azizi:2007:SPC

[MAM07] B. Mohammed-Azizi and D. E. Medjadi. Single particle calculations for a Woods–Saxon potential with triaxial deformations, and large Cartesian oscillator basis (new version code). Computer Physics Communications, 176(9–10):634–635, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001671.

Mamedov:2008:AEI

[Mam08] B. A. Mamedov. Accurate evaluation of integrals arising from the bulk electron densities in quantum wells of high electron mobility transistors. Computer Physics Communications, 178(9):673–675, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000040.

Mancini:2002:REA

[Man02] G. Mancini. A redundancy eliminating approach to linearly independent rings selection in the ring perception problem. Computer Physics Communications, 143(2):187–197, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004544.

Mandrekas:2004:GCC

[Man04] John Mandrekas. GTNEUT: a code for the calculation of neutral particle transport in plasmas based on the transmission and escape probability method. Computer Physics Communications, 161(1–2):36–64, August 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400205X.

Martin:2001:CLS

[Mar01] A. Martin. The COMPASS off-line system. Computer Physics Communications, 140(1–2):82–91, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002582.

Marro:2008:BCM

[Mar08] J. Marro. A brief comment on the modeling of flow. Computer Physics Communications, 179(1–3):144–149, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000313.

Masjukov:2000:MCM

[Mas00] Vladimir V. Masjukov. A mathematical and computer model of the main tropical hurricane wave system. Computer Physics Communications, 126(1–2):121–125, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004798.

Mason:2005:FNL

[Mas05] D. R. Mason. Faster neighbour list generation using a novel lattice vector representation. Computer Physics Communications, 170(1):31–41, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002328.

Mazziotta:2000:MCC

[Maz00] M. N. Mazziotta. A Monte Carlo code for full simulation of a transition radiation detector. Computer Physics Communications, 132(1–2):110–123, October 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001466.

Messmer:2004:PES

[MB04] Peter Messmer and David L. Bruhwiler. A parallel electrostatic solver for the VORPAL code. Computer Physics Communications, 164(1–3):118–121, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002681.

Milchev:2005:AIP

[MB05a] Andrey Milchev and Kurt Binder. Adsorption-induced polymer translocation through a nanopore: a Monte Carlo investigation. Computer Physics Communications, 169(1–3):107–110, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001220.

Moroni:2005:CSD

[MB05b] Saverio Moroni and Stefano Baroni. Computational spectroscopy of doped he clusters. Computer Physics Communications, 169(1–3):404–407, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001864.

Mouronte:2009:CSO

[MBC^{+}09] M. L. Mouronte, R. M. Benito, J. P. Cárdenas, A. Santiago, V. Feliú, P. van Wijngaarden, and L. G. Moyano. Complexity in Spanish optical fiber and SDH transport networks. Computer Physics Communications, 180(4):523–526, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000022.

Martinazzo:2003:MVP

[MBG03] R. Martinazzo, E. Bodo, and F. A. Gianturco. A modified variable-phase algorithm for multichannel scattering with long-range potentials. Computer Physics Communications, 151(2):187–198, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502007373.

Martinsen:2009:AMC

[MBKJ09] Paul Martinsen, Johannes Blaschke, Rainer Künnemeyer, and Robert Jordan. Accelerating Monte Carlo simulations with an NVIDIA ^{(}R) graphics processor. Computer Physics Communications, 180(10):1983–1989, October 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001593.

Martin:2001:IAM

[MBR01] Jan M. L. Martin, Charles W. Bauschlicher, Jr., and Alessandra Ricca. On the integration accuracy in molecular density functional theory calculations using Gaussian basis sets. Computer Physics Communications, 133(2–3):189–201, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001740.

Morita:2001:VMS

[MC01] Youhei Morita and The MONARC Collaboration. Validation of the MONARC simulation tools. Computer Physics Communications, 140(1–2):153–161, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002661.

Mainegra:2003:NSD

[MC03] E. Mainegra and R. Capote. Nuclear state density calculations: An exact recursive approach. Computer Physics Communications, 150(1):43–52, January 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006136.

Mazzeo:2008:HHP

[MC08] M. D. Mazzeo and P. V. Coveney. HemeLB: a high performance parallel lattice-Boltzmann code for large scale fluid flow in complex geometries. Computer Physics Communications, 178(12):894–914, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000805.

Molteni:2009:SPI

[MC09] Diego Molteni and Andrea Colagrossi. A simple procedure to improve the pressure evaluation in hydrodynamic context using the SPH. Computer Physics Communications, 180(6):861– 872, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004219.

Marques:2003:OFP

[MCBR03] Miguel A. L. Marques, Alberto Castro, George F. Bertsch, and Angel Rubio. octopus: a first-principles tool for excited electron-ion dynamics. Computer Physics Communications, 151(1):60–78, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502006860.

Mattoni:2005:CTS

[MCC05] Alessandro Mattoni, Luciano Colombo, and Fabrizio Cleri. Crack-tip stress shielding by a hard fiber in β-SiC: an atomistic study. Computer Physics Communications, 169(1–3):40– 43, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001086.

Marro:2002:MNP

[MCH02] J. Marro, J. M. Cortes, and P. I. Hurtado. Modeling nonequilibrium phase transitions and critical behavior in complex systems. Computer Physics Communications, 147(1–2):115–119, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200228X.

McKee:2007:ACM

[McK07] Shawn McKee. The ATLAS computing model: status, plans and future possibilities. Computer Physics Communications, 177(1–2):231–234, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507000513.

Marchetto:2005:THP

[MCL05] Chiara Marchetto, Francesco Califano, and Maurizio Lontano. Turbulence healing via plasma-wave interaction: the results of a study via kinetic codes. Computer Physics Communications, 169(1–3):207–209, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001438.

Marchand:2001:APG

[MCLDP01] Richard Marchand, Mathieu Charbonneau-Lefort, Mathieu Dumberry, and Benoit Pronovost. ARANEA, a program for generating unstructured triangular meshes with a Java Graphics User Interface. Computer Physics Communications, 139(2):172–185, September 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001874.

Michielsen:2000:MIA

[MD00] K. Michielsen and H. De Raedt. Morphological image analysis. Computer Physics Communications, 132(1–2):94–103, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001399.

Milotti:2009:NIM

[MDC09] Edoardo Milotti, Alessio Del Fabbro, and Roberto Chignola. Numerical integration methods for large-scale biophysical simulations. Computer Physics Communications, 180 (11):2166–2174, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509002021.

Melean:2004:STI

[MDH04] Yasmin Meleán, Leonardo Di G. Sigalotti, and Anwar Hasmy. On the SPH tensile instability in forming viscous liquid drops. Computer Physics Communications, 157(3):191–200, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550300523X.

Muhlleitner:2005:SFC

[MDM05] M. Mühlleitner, A. Djouadi, and Y. Mambrini. SDECAY: a Fortran code for the decays of the supersymmetric particles in the MSSM. Computer Physics Communications, 168(1):46–70, May 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000822.

Magee:2009:CAC

[MDS09] James E. Magee, Helen Dutton, and Flor R. Siperstein. CaSPA — an algorithm for calculation of the size of percolating aggregates. Computer Physics Communications, 180(9):1503–1510, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550900112X.

Matin:2003:CNL

[MDT03] M. L. Matin, P. J. Daivis, and B. D. Todd. Cell neighbor list method for planar elongational flow: rheology of a diatomic fluid. Computer Physics Communications, 151(1):35– 46, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006999.

Murtagh:2000:FAI

[ME00] Fionn Murtagh and Daniel Egret. Foreword to articles on “from information to knowledge using astronomical databases”. Computer Physics Communications, 127(2–3):175–176, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005068.

Meijer:2001:TDW

[Mei01] Anthony J. H. M. Meijer. Time-dependent wave packet calculations on parallel computers: a new and efficient algorithm for evaluating HΨ. Computer Physics Communications, 141(3):330–341, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004180.

Melnik:2001:CAC

[Mel01] R. V. N. Melnik. Computational analysis of coupled physical fields in piezothermoelastic media. Computer Physics Communications, 142(1–3):231–237, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003150.

Melheim:2005:CIM

[Mel05] Jens A. Melheim. Cluster integration method in Lagrangian particle dynamics. Computer Physics Communications, 171(3):155–161, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002924.

Mucke:2000:MCS

[MER^{+}00] A. Mücke, Ralph Engel, J. P. Rachen, R. J. Protheroe, and Todor Stanev. Monte Carlo simulations of photohadronic processes in astrophysics. Computer Physics Communications, 124(2–3):290–314, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004464.

Meyer:2002:PQA

[Mey02] David A. Meyer. Physical quantum algorithms. Computer Physics Communications, 146(3):295–301, July 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004204.

Martin:2005:BPS

[MFF^{+}05] Johannes U. Martin, Gary N. Felder, Andrei V. Frolov, Lev Kofman, and Marco Peloso. BRANECODE: a program for simulations of braneworld dynamics. Computer Physics Communications, 171(1):69–78, September 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002882.

Merzel:2007:NPC

[MFVJ07] Franci Merzel, Fabien Fontaine-Vive, and Mark R. Johnson. NMscatt: a program for calculating inelastic scattering from large biomolecular systems using classical forcefield simulations. Computer Physics Communications, 177 (6):530–538, September 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002469.

Mareschal:2008:P

[MG08a] Michel Mareschal and Paul Geerlings. Preface. Computer Physics Communications, 179(1–3):xi–xii, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000647.

Martin-Garcia:2008:XFI

[MG08b] José M. Martín-García. xPerm: fast index canonicalization for tensor computer algebra. Computer Physics Communications, 179(8):597–603, October 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800221X.

Morte:2009:ESE

[MG09a] Michele Della Morte and Leonardo Giusti. Exploiting symmetries for exponential error reduction in path integral Monte Carlo. Computer Physics Communications, 180(6):813–818, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003688.

Morte:2009:SEE

[MG09b] Michele Della Morte and Leonardo Giusti. Symmetries and exponential error reduction in Yang–Mills theories on the lattice. Computer Physics Communications, 180(6):819–826, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001052.

Muller:2009:MLL

[MG09c] Thomas Müller and Frank Grave. Motion4D — a library for lightrays and timelike worldlines in the Theory of Relativity. Computer Physics Communications, 180(11):2355–2360, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002343.

Medvedev:2005:OMA

[MGG05] Dmitry M. Medvedev, Evelyn M. Goldfield, and Stephen K. Gray. An OpenMP/MPI approach to the parallelization of iterative four-atom quantum mechanics. Computer Physics Communications, 166(2):94–108, March 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005260.

Mayneris:2008:RWC

[MGG08] Jordi Mayneris, Miguel González, and Stephen K. Gray. Real wavepacket code for ABC +D → AB +CD reactive scattering. Computer Physics Communications, 179(10):741–747, November 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003159.

Meglicki:2007:MFC

[MGN07] Zdzislaw Meglicki, Stephen K. Gray, and Boyana Norris. Multigrid FDTD with chombo. Computer Physics Communications, 176(2):109–120, January 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003237.

Martin-Garcia:2007:ITP

[MGPM07] J. M. Martín-García, R. Portugal, and L. R. U. Manssur. The invar tensor package. Computer Physics Communications, 177(8):640–648, October 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507002561.

Martin-Garcia:2008:ITP

[MGYP08] J. M. Martín-García, D. Yllanes, and R. Portugal. The invar tensor package: Differential invariants of Riemann. Computer Physics Communications, 179(8):586–590, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550800180X.

Marini:2009:YIT

[MHGV09] Andrea Marini, Conor Hogan, Myrta Grüning, and Daniele Varsano. yambo: An ab initio tool for excited state calculations. Computer Physics Communications, 180(8):1392–1403, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000472.

Messina:2002:CIC

[MHK02] René Messina, Christian Holm, and Kurt Kremer. Charge inversion in colloidal systems. Computer Physics Communications, 147(1–2):282–285, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002886.

Morte:2005:ILC

[MHK^{+}05] Michele Della Morte, Roland Hoffmann, Francesco Knechtli, Ulli Wolff, and ALPHA Collaboration. Impact of large cutoff-effects on algorithms for improved Wilson fermions. Computer Physics Communications, 165(1):49–58, January 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005004.

Mattila:2007:ESA

[MHR^{+}07] Keijo Mattila, Jari Hyväluoma, Tuomo Rossi, Mats Aspnäs, and Jan Westerholm. An efficient swap algorithm for the lattice Boltzmann method. Computer Physics Communications, 176(3):200–210, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003663.

Mason:2005:FRS

[MHS05] D. R. Mason, T. S. Hudson, and A. P. Sutton. Fast recall of state-history in kinetic Monte Carlo simulations utilizing the zobrist key. Computer Physics Communications, 165(1):37–48, January 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550400462X.

Mizusaki:2005:PIR

[MI05] Takahiro Mizusaki and Masatoshi Imada. Path-integral renormalization group method with quantum-number projection. Computer Physics Communications, 169(1–3):408–411, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001876.

Michel:2007:PCW

[Mic07] N. Michel. Precise Coulomb wave functions for a wide range of complex l, η and z. Computer Physics Communications, 176(3):232–249, February 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003717.

Milotti:2006:PPE

[Mil06] Edoardo Milotti. PLNoise: a package for exact numerical simulation of power-law noises. Computer Physics Communications, 175(3):212–225, August 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001597.

Milotti:2007:NVP

[Mil07] Edoardo Milotti. New version of PLNoise: a package for exact numerical simulation of power-law noises. Computer Physics Communications, 177(4):391–398, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700224X.

Miyagawa:2007:PAP

[MIM^{+}07] Y. Miyagawa, M. Ikeyama, S. Miyagawa, M. Tanaka, and H. Nakadate. Plasma analysis for the plasma immersion ion implantation processing by a PIC-MCC simulation. Computer Physics Communications, 177(1–2):84–87, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001221.

Miniati:2001:CNC

[Min01] Francesco Miniati. COSMOCR: a numerical code for cosmic ray studies in computational cosmology. Computer Physics Communications, 141(1):17–38, November 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002934.

Munster:2002:DIS

[MK02] Gernot Münster and Christel Kamp. Distribution of instanton sizes in a simplified instanton gas model. Computer Physics Communications, 147(1–2):423–426, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003181.

Mocken:2005:RSL

[MK05] Guido R. Mocken and Christoph H. Keitel. Radiation spectra of laser-driven quantum relativistic electrons. Computer Physics Communications, 166(3):171–190, March 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005612.

Mocken:2008:FSO

[MK08] Guido R. Mocken and Christoph H. Keitel. FFT-split-operator code for solving the Dirac equation in 2+1 dimensions. Computer Physics Communications, 178(11):868–882, June 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000611.

Monaghan:2009:SPB

[MK09] J. J. Monaghan and J. B. Kajtar. SPH particle boundary forces for arbitrary boundaries. Computer Physics Communications, 180(10):1811–1820, October 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001544.

Mischler:2002:CIM

[MKB02] C. Mischler, W. Kob, and K. Binder. Classical and abinitio molecular dynamic simulation of an amorphous silica surface. Computer Physics Communications, 147(1–2):222–225, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002503.

Morhac:2005:EFA

[MKJ^{+}05] M. Morhác, J. Kliman, M. Jandel, L. Krupa, V. Matousek, J. H. Hamilton, and A. V. Ramayya. Efficient fitting algorithms applied to analysis of coincidence γ-ray spectra. Computer Physics Communications, 172(1):19–41, October 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505003607.

Mohankumar:2005:EFD

[MKK05] N. Mohankumar, T. Kannan, and S. Kanmani. On the evaluation of Fermi–Dirac integral and its derivatives by IMT and DE quadrature methods. Computer Physics Communications, 168(2):71–77, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505000639.

Murtazaev:2002:MCI

[MKM02] A. K. Murtazaev, I. K. Kamilov, and M. A. Magomedov. Monte Carlo investigation of critical phenomena in models of real magnetics with crossovers. Computer Physics Communications, 147(1–2):447–450, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003247.

Monovasilis:2007:FTF

[MKS07] Th. Monovasilis, Z. Kalogiratou, and T. E. Simos. Families of third and fourth algebraic order trigonometrically fitted symplectic methods for the numerical integration of Hamiltonian systems. Computer Physics Communications, 177(10):757–763, November 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003104.

Manzhos:2003:PIA

[ML03] Sergei Manzhos and Hans-Peter Loock. Photofragment image analysis using the Onion-Peeling Algorithm. Computer Physics Communications, 154(1):76–87, July 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002777.

Mohammadian:2006:CSI

[ML06] A. Mohammadian and D. Y. Le Roux. Conservative semiimplicit semi-Lagrangian scheme for simulation of shallow flows. Computer Physics Communications, 174(2):99–108, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004765.

Meca:2007:DDF

[MLF07] Antonio Soto Meca, Francisco Alhama López, and Carlos González Fernández. Density-driven flow and solute transport problems. A 2-D numerical model based on the network simulation method. Computer Physics Communications, 177(9):720–728, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003098.

Matheu:2001:RBS

[MLG^{+}01] David M. Matheu, Thomas A. Lada II, William H. Green, Anthony M. Dean, and Jeffrey M. Grenda. Rate-based screening of pressure-dependent reaction networks. Computer Physics Communications, 138(3):237–249, August 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002375.

Monville:2008:SLI

[MLPT08] M. E. Monville, J. D. Lamb, J. C. G. Pereira, and J. A. Truter. Simulation of laser-induced nanopattern formation. Computer Physics Communications, 178(1):40–47, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003724.

Munehisa:2001:ACE

[MM01] T. Munehisa and Y. Munehisa. An attempt to calculate energy eigenvalues in quantum systems of large sizes (operator-variational approach). Computer Physics Communications, 142(1–3):176–180, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003216.

McPeake:2004:DSI

[MM04] Dermot McPeake and J. F. McCann. Discretization and solution of the inhomogeneous bogoliubov-de Gennes equations. Computer Physics Communications, 161(3):119–128, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002152.

Morhac:2005:MNS

[MM05] M. Morhác and V. Matousek. Multidimensional nuclear spectra compression using fast adaptive Fourier-based transforms. Computer Physics Communications, 165(2):127–138, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004965.

Maitre:2008:SMI

[MM08] D. Maître and P. Mastrolia. S@M, a Mathematica implementation of the spinor-helicity formalism. Computer Physics Communications, 179(7):501–534, October 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001793.

Morhac:2009:LSF

[MM09] Miroslav Morhác and Vladislav Matousek. Library of sophisticated functions for analysis of nuclear spectra. Computer Physics Communications, 180(10):1913–1940, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001477.

Milchev:2002:NSP

[MMB02] Andrey Milchev, Alexander Milchev, and Kurt Binder. Nanodroplets on a solid plane: wetting and spreading in a Monte Carlo simulation. Computer Physics Communications, 146 (1):38–53, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004332.

Meinke:2008:SVS

[MMEH08] Jan H. Meinke, Sandipan Mohanty, Frank Eisenmenger, and Ulrich H. E. Hansmann. SMMP v. 3.0 — simulating proteins and protein interactions in Python and Fortran. Computer Physics Communications, 178(6):459–470, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004614.

Molinas-Mata:2000:ERS

[MMMM00] Pau Molinàs-Mata and Patrici Molinàs-Mata. Electron resonances and Savitzky–Golay methods: How can the tunneling current be properly calculated? Computer Physics Communications, 132(1–2):104–109, October 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001430.

Matveev:2004:EST

[MMR04] Alexei V. Matveev, Markus Mayer, and Notker Rösch. Efficient symmetry treatment for the nonrelativistic and relativistic molecular Kohn–Sham problem. the symmetry module of the program paragauss. Computer Physics Communications, 160(2):91–119, July 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550400102X.

Matsumoto:2004:SGC

[MMTH04] Tomoaki Matsumoto, Masahiro N. Machida, Kohji Tomisaka, and Tomoyuki Hanawa. Self-gravitational collapse of a magnetized cloud core: high resolution simulations with threedimensional MHD nested grid. Computer Physics Communications, 164(1–3):229–236, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002851.

Milnikov:2001:UDB

[MN01] Gennady V. Mil’nikov and Hiroki Nakamura. Use of diabatic basis in the adiabatic-by-sector R-matrix propagation method in time-independent reactive scattering calculations. Computer Physics Communications, 140(3):381–398, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002879.

Mohankumar:2007:EF

[MN07]  N. Mohankumar and A. Natarajan.  On the evalua-

tion of H and X functions. Computer Physics Communications, 176(4):266–270, February 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003705.

Milnikov:2001:SEE

[MNH01] Gennady V. Mil’nikov, Hiroki Nakamura, and Jiri Horácek. Stable and efficient evaluation of Green’s function in scattering problem. Computer Physics Communications, 135(3):278–292, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500002381.

Morgan:2000:PBC

[MNV00] David L. Morgan, Jr., J. Robert Neely, and Harry C. Vantine. Periodic boundary conditions in a 3D hydro code. Computer Physics Communications, 126(1–2):131–136, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004208.

Matsuda:2000:PPWa

[MNYY00a] Masanori Matsuda, Junichi Nagata, Hiro Yoshino, and Yutaka Yoshino. PANN: Partial-wave analysis of nucleonnucleon scattering in wide-energy region. Computer Physics Communications, 131(3):225–263, September 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001235.

Matsuda:2000:PPWb

[MNYY00b] Masanori Matsuda, Junichi Nagata, Hiro Yoshino, and Yutaka Yoshino. PAPH: Partial-wave analysis of p–^{3}He scattering. Computer Physics Communications, 131(3):264–288, September 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001259.

McSherry:2003:IAA

[MOC03] D. M. McSherry, S. F. C. O’Rourke, and D. S. F. Crothers. Ion-atom/ argon-calculation of ionization cross sections by fast ion impact for neutral target atoms ranging from hydrogen to argon. Computer Physics Communications, 155(2):144–158, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503003424.

Mohankumar:2007:TNS

[Moh07] N. Mohankumar. Two new series for the Fermi–Dirac integral. Computer Physics Communications, 176(11–12):665– 669, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002007.

Mohankumar:2008:CAE

[Moh08] N. Mohankumar. Comments on the accurate evaluation of integrals arising from the bulk electron densities in quantum wells. Computer Physics Communications, 179(11):819–820, December 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002506.

MacNeice:2000:PPA

[MOM^{+}00] Peter MacNeice, Kevin M. Olson, Clark Mobarry, Rosalinda de Fainchtein, and Charles Packer. PARAMESH: a parallel adaptive mesh refinement community toolkit. Computer Physics Communications, 126(3):330–354, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005019.

Morishita:2001:SLP

[Mor01] Tetsuya Morishita. Structure of liquid phosphorus: a liquid-liquid phase transition via constant-pressure firstprinciples molecular dynamics. Computer Physics Communications, 142(1–3):356–360, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003745.

Munz:2000:TDF

[MOS00] C.-D. Munz, P. Ommes, and R. Schneider. A threedimensional finite-volume solver for the Maxwell equations with divergence cleaning on unstructured meshes. Computer Physics Communications, 130(1–2):83–117, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550000045X. See erratum [MOS01].

Munz:2001:ETD

[MOS01] C.-D. Munz, P. Ommes, and R. Schneider. Erratum to: “A three-dimensional finite-volume solver for the Maxwell equations with divergence cleaning on unstructured meshes” [Computer Physics Communications 130 (2000) 83–117]. Computer Physics Communications, 135(2):260, April 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001412. See [MOS00].

Marongiu:2001:AIA

[MP01a] Alessandro Marongiu and Paolo Palazzari. Automatic implementation of affine iterative algorithms: Design flow and communication synthesis. Computer Physics Communications, 139(1):109–131, September 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100234X.

Moiseenko:2001:LSM

[MP01b] V. E. Moiseenko and V. V. Pilipenko. Local solution method for numerical solving of the wave propagation problem. Computer Physics Communications, 141(3):342–349, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004209.

Mendes:2003:PMC

[MP03] B. Mendes and A. Pereira. Parallel Monte Carlo driver (PMCD)-a software package for Monte Carlo simulations in parallel. Computer Physics Communications, 151(1):89–95, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006896.

Manssur:2004:CPF

[MP04] L. R. U. Manssur and R. Portugal. The canon package: a fast kernel for tensor manipulators. Computer Physics Communications, 157(2):173–180, February 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004946.

Movilla:2005:ICS

[MP05] J. L. Movilla and J. Planelles. Image charges in spherical quantum dots with an off-centered impurity: algorithm and numerical results. Computer Physics Communications, 170(2):144–152, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002651.

Monk:2006:EMC

[MP06] J. Monk and A. Pilkington. ExHuME 1.3: a Monte Carlo event generator for exclusive diffraction. Computer Physics Communications, 175(3):232–239, August 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001913.

Marquezino:2008:QSQ

[MP08] F. L. Marquezino and R. Portugal. The QWalk simulator of quantum walks. Computer Physics Communications, 179(5):359–369, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508001124.

Merimaa:2000:ISP

[MPK00] J. Merimaa, L. F. Perondi, and K. Kaski. An interactive simulation program for visualizing complex phenomena in solids. Computer Physics Communications, 124(1):60–75, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599002647.

Masia:2005:PMP

[MPR05] Marco Masia, Michael Probst, and Rossend Rey. On the performance of molecular polarization methods close to a point charge. Computer Physics Communications, 169(1–3):331–334, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001712.

Muller:2009:HBC

[MPS09] M. Müller, C. Pastorino, and J. Servantie. Hydrodynamic boundary condition of polymer melts at simple and complex surfaces. Computer Physics Communications, 180(4):600–604, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004189.

Maggs:2005:AFS

[MR05] A. C. Maggs and J. Rottler. Auxiliary field simulation and Coulomb’s law. Computer Physics Communications, 169 (1–3):160–165, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505001347.

Martin:2006:TPC

[MR06] Stephen P. Martin and David G. Robertson. TSIL: a program for the calculation of two-loop self-energy integrals. Computer Physics Communications, 174(2):133–151, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505004741.

Meloni:2005:CMS

[MRF^{+}05] Simone Meloni, Mario Rosati, Alessandro Federico, Luca Ferraro, Alessandro Mattoni, and Luciano Colombo. Computational materials science application programming interface (CMSapi): a tool for developing applications for atomistic simulations. Computer Physics Communications, 169(1–3):462– 466, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001992.

Mason:2004:SKM

[MRS04] D. R. Mason, R. E. Rudd, and A. P. Sutton. Stochastic kinetic Monte Carlo algorithms for long-range Hamiltonians. Computer Physics Communications, 160(2):140–157, July 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001651.

Melchionna:2005:LBP

[MS05a] Simone Melchionna and Sauro Succi. Lattice Boltzmann– Poisson method for electrorheological nanoflows in ion channels. Computer Physics Communications, 169(1–3):203–206, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001426.

Miranda:2005:CMS

[MS05b] Caetano R. Miranda and Sandro Scandolo. Computational materials science meets geophysics: dislocations and slip planes of MgO. Computer Physics Communications, 169(1–3):24–27, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001049.

Madsen:2006:BCC

[MS06] Georg K. H. Madsen and David J. Singh. BoltzTraP. A code for calculating band-structure dependent quantities. Computer Physics Communications, 175(1):67–71, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001305.

Matsumoto:2008:ICA

[MS08a] Yosuke Matsumoto and Kanako Seki. Implementation of the CIP algorithm to magnetohydrodynamic simulations. Computer Physics Communications, 179(5):289–296, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001094.

Michel:2008:FCG

[MS08b] N. Michel and M. V. Stoitsov. Fast computation of the Gauss hypergeometric function with all its parameters complex with application to the Pöschl–Teller–Ginocchio potential wave functions. Computer Physics Communications, 178(7):535– 551, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004699.

Moritsugu:2009:RPC

[MS09] Kei Moritsugu and Jeremy C. Smith. REACH: a program for coarse-grained biomolecular simulation. Computer Physics Communications, 180(7):1188–1195, July 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000216.

Martin-Samos:2009:SOS

[MSB09] Layla Martin-Samos and Giovanni Bussi. SaX: An open source package for electronic-structure and optical-properties calculations in the GW approximation. Computer Physics Communications, 180(8):1416–1425, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000496.

Mazrooei-Sebdani:2008:NTR

[MSD08] Reza Mazrooei-Sebdani and Mehdi Dehghan. A nontrivial relation between some many-dimensional chaotic discrete dynamical systems and some one-dimensional chaotic discrete dynamical systems. Computer Physics Communications, 179(9):628–633, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002178.

Munejiri:2001:PIP

[MSH01] Shuji Munejiri, Fuyuki Shimojo, and Kozo Hoshino. Photoinduced polymerization in liquid sulfur studied by an ab initio molecular-dynamics simulation. Computer Physics Communications, 142(1–3):131–135, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003514.

Mostofi:2002:TEC

[MSHP02] Arash A. Mostofi, Chris-Kriton Skylaris, Peter D. Haynes, and Mike C. Payne. Total-energy calculations on a real space grid with localized functions and a plane-wave basis. Computer Physics Communications, 147(3):788–802, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004617. See corrigendum

[MSHP20].

Mostofi:2020:CTE

[MSHP20] Arash A. Mostofi, Chris-Kriton Skylaris, Peter D. Haynes, and Mike C. Payne. Corrigendum to “Total-energy calculations on a real space grid with localised functions and a plane-wave basis” [Comput. Phys. Comm. 147/3 (2002) 788–802]. Computer Physics Communications, 252(??):Article 107232, July 2020. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465520300680. See [MSHP02].

McCurdy:2002:FDS

[MSK^{+}02] C. William McCurdy, Horst D. Simon, William T. C. Kramer, Robert F. Lucas, William E. Johnston, and David H. Bailey. Future directions in scientific supercomputing for computational physics. Computer Physics Communications, 147(1–2):34–39, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200200X.

McGrath:2005:TMC

[MSK^{+}05] Matthew J. McGrath, J. Ilja Siepmann, I-Feng W. Kuo, Christopher J. Mundy, Joost VandeVondele, Michiel Sprik, Jürg Hutter, Fawzi Mohamed, Matthias Krack, and Michele Parrinello. Toward a Monte Carlo program for simulating vapor-liquid phase equilibria from first principles. Computer Physics Communications, 169(1–3):289–294, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001621.

Makhlin:2000:JJQ

[MSS00] Yuriy Makhlin, Gerd Schön, and Alexander Shnirman. Josephson junction quantum logic gates. Computer Physics Communications, 127(1):156–164, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000254.

Meyer:2007:EHT

[MSS^{+}07] Harvey Meyer, Hubert Simma, Rainer Sommer, Michele Della Morte, Oliver Witzel, Ulli Wolff, and Alpha Collaboration. Exploring the HMC trajectory-length dependence of autocorrelation times in lattice QCD. Computer Physics Communications, 176(2):91–97, January 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003201.

Mazza:2009:CMC

[MSS^{+}09] Marco G. Mazza, Kevin Stokely, Elena G. Strekalova, H. Eugene Stanley, and Giancarlo Franzese. Cluster Monte Carlo and numerical mean field analysis for the water liquid-liquid phase transition. Computer Physics Communications, 180(4):497– 502, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000344.

Makabe:2007:PEF

[MSY07] T. Makabe, T. Shimada, and T. Yagisawa. Plasma etching and feature evolution of organic low- k material by using vicaddress. Computer Physics Communications, 177(1–2):64– 67, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700118X.

Mussa:2000:BQB

[MT00] Hamse Y. Mussa and Jonathan Tennyson. Bound and quasibound rotation-vibrational states using massively parallel computers. Computer Physics Communications, 128(1–2):434–445, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000588.

Mandelzweig:2001:QAN

[MT01] V. B. Mandelzweig and F. Tabakin. Quasilinearization approach to nonlinear problems in physics with application to nonlinear ODEs. Computer Physics Communications, 141(2):268–281, November 30, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004155.

Marro:2007:NHW

[MTC07] J. Marro, Joaquín J. Torres, and Jesús M. Cortés. Networks with heterogeneously weighted connections and partial synchronization of nodes. Computer Physics Communications, 177(1–2):180–183, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000872.

Munro:2002:GOF

[MTJ02] Lindsey J. Munro, Arnold Tharrington, and Kenneth D. Jordan. Global optimization and finite temperature simulations of atomic clusters: Use of Xe_{n}Ar_{m} clusters as test systems. Computer Physics Communications, 145(1):1–23, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004301.

Man:2001:DFB

[MTLC01] W. S. Man, H. Y. Tam, S. Y. Liu, and K. C. Chan. Design of fiber Bragg grating for flattening of superfluorescent fiber source (SFS) using computer simulation. Computer Physics Communications, 142(1–3):270–273, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003411.

Miliukova:2000:MMD

[MTZ00] Olga Yu. Miliukova, Vyacheslav A. Trofimov, and Irina G. Zakharova. Mathematical modeling of diffraction instability in optical bistable system. Computer Physics Communications, 126(1–2):126–130, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004774.

Moch:2006:XTF

[MU06] S. Moch and P. Uwer. XSummer — transcendental functions and symbolic summation in Form. Computer Physics Communications, 174(9):759–770, May 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506000397.

Muller:2002:IBW

[Mül02] Marcus Müller. Interplay between wetting and miscibility in thin binary polymer films. Computer Physics Communications, 147(1–2):292–297, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002904.

Muller:2005:SAP

[Mül05] Ewald Müller. Simulating astrophysical phenomena: challenges and achievements. Computer Physics Communications, 169(1–3):353–361, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505001761.

Muser:2002:NEA

[Müs02a] M. H. Müser. On new efficient algorithms for PIMC and PIMD. Computer Physics Communications, 147(1–2):83–86, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002217.

Muser:2002:TAU

[Müs02b] Martin H. Müser. Towards an atomistic understanding of solid friction by computer simulations. Computer Physics Communications, 146(1):54–62, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502004344.

Manfredi:2004:VSP

[MV04] G. Manfredi and F. Valsaque. Vlasov simulations of plasmawall interactions in a weakly collisional plasma. Computer Physics Communications, 164(1–3):262–268, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002905.

Martin-Vaquero:2009:SOS

[MVJ09] J. Martín-Vaquero and B. Janssen. Second-order stabilized explicit Runge–Kutta methods for stiff problems. Computer Physics Communications, 180(10):1802–1810, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001532.

McBride:2005:NMM

[MVS05] Carl McBride, Carlos Vega, and Eduardo Sanz. Non-Markovian melting: a novel procedure to generate initial liquid like phases for small molecules for use in computer simulation studies. Computer Physics Communications, 170(2):137–143, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500264X.

Matsuoka:2001:ADF

[MW01] Osamu Matsuoka and Yoshihiro Watanabe. An atomic Dirac–Fock–Roothaan program. Computer Physics Communications, 139(2):218–234, September 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001989.

McCarthy:2001:ESC

[MWA01] S. A. McCarthy, J. B. Wang, and P. C. Abbott. Electronic structure calculation for N-electron quantum dots. Computer Physics Communications, 141(1):175–204, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004015.

Malevanets:2000:MVD

[MY00a] A. Malevanets and J. M. Yeomans. Modeling viscous drag in binary fluid mixtures. Computer Physics Communications, 127(1):105–112, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000308.

Malevanets:2000:PBA

[MY00b] A. Malevanets and J. M. Yeomans. A particle-based algorithm for the hydrodynamics of binary fluid mixtures. Computer Physics Communications, 129(1–3):282–288, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001156.

Manzhos:2009:FSM

[MYC09] Sergei Manzhos, Koichi Yamashita, and Tucker Carrington, Jr. Fitting sparse multidimensional data with lowdimensional terms. Computer Physics Communications, 180 (10):2002–2012, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001817.

Mizoe:2001:NSS

[MYJY01] H. Mizoe, Seong Y. Yoon, M. Josho, and T. Yabe. Numerical simulation on snow melting phenomena by CIP method. Computer Physics Communications, 135(2):154–166, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002290.

Mostofi:2008:WTO

[MYL^{+}08] Arash A. Mostofi, Jonathan R. Yates, Young-Su Lee, Ivo Souza, David Vanderbilt, and Nicola Marzari. wannier90: a tool for obtaining maximally-localised Wannier functions. Computer Physics Communications, 178(9):685–699, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004936.

McTaggart:2004:FDS

[MZB^{+}04] N. McTaggart, R. Zagórski, X. Bonnin, A. Runov, and R. Schneider. Finite difference scheme for solving general 3D convection-diffusion equation. Computer Physics Communications, 164(1–3):318–329, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400298X.

Nakano:2007:PPS

[Nak07] Aiichiro Nakano. Pathfinder: a parallel search algorithm for concerted atomistic events. Computer Physics Communications, 176(4):292–299, February 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003948.

Nakano:2008:STE

[Nak08] Aiichiro Nakano. A space–time-ensemble parallel nudged elastic band algorithm for molecular kinetics simulation. Computer Physics Communications, 178(4):280–289, February 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004225.

Nappi:2009:PUE

[Nap09] A. Nappi. A pitfall in the use of extended likelihood for fitting fractions of pure samples in a mixed sample. Computer Physics Communications, 180(2):269–275, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003652.

Nath:2008:MCC

[Nat08] S. Nath. A Monte Carlo C-code for calculating transmission efficiency of recoil separators and viewing residue trajectories. Computer Physics Communications, 179(7):492–500, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001641.

Nath:2009:TVI

[Nat09] S. Nath. TERS v2.0: An improved version of TERS. Computer Physics Communications, 180(11):2392–2393, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002069. See erratum [?].

Nath:2010:ETV

[Nat10] S. Nath. Erratum to “TERS v2.0: An improved version of TERS” [Comput. Phys. Commun. 180 (2009) 2392]. Computer Physics Communications, 181(9):1659, September 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046551000202X. See [Nat09].

Noel:2008:HSD

[NBPG08] Claire Noël, Yves Busegnies, Miltiadis V. Papalexandris, and Stephane Goriely. Hydrodynamical simulations of detonations in superbursts. Computer Physics Communications, 179(1–3):190–193, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000398.

Newman:2002:SFN

[New02] M. E. J. Newman. The structure and function of networks. Computer Physics Communications, 147(1–2):40–45, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002011.

Newman:2007:NHE

[New07] Harvey B. Newman. Networking for high energy and nuclear physics. Computer Physics Communications, 177(1–2):224– 230, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000525.

Nikkinen:2006:REU

[NFH06] J. Nikkinen, S. Fritzsche, and S. Heinäsmäki. Revised and extended utilities for the ratip package. Computer Physics Communications, 175(5):348–358, September 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001895.

Nakamura:2001:DPC

[NFS01a] Hiroaki Nakamura, Susumu Fujiwara, and Tetsuya Sato. Dynamical process of coalescence of domains in a short chain-molecule system. Computer Physics Communications, 142(1–3):127–130, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003502.

Nurhuda:2001:EEM

[NFS01b] M. Nurhuda, F. H. M. Faisal, and U. Schwengelbeck. Eigenstate expansion method for simulations of non-perturbative multiphoton processes. Computer Physics Communications, 134(3):291–306, March 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002071.

Nakamura:2002:DOO

[NFS02] Hiroaki Nakamura, Susumu Fujiwara, and Tetsuya Sato. Dynamics of orientationally ordered domains in a short chainmolecule system: Size dependence of domain oscillation. Computer Physics Communications, 147(1–2):346–349, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003016.

Ni:2002:KOD

[NG02] Jun Ni and Binglin Gu. Kinetics of ordering during codeposition. Computer Physics Communications, 147(1–2):234–237, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002539.

Newman:2004:RVS

[NGE^{+}04] D. L. Newman, M. V. Goldman, R. E. Ergun, L. Andersson, and N. Sen. Reduced Vlasov simulations in higher dimensions. Computer Physics Communications, 164(1–3): 122–127, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002693.

Nemnes:2009:SBA

[NH09] G. A. Nemnes and K. H. Hoffmann. Spin-box algorithm for low temperature dynamics of short range disordered Ising spin systems. Computer Physics Communications, 180(7):1098– 1103, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004542.

Nakayama:2007:IPF

[NHS07] Akihiro Nakayama, Katsuya Hasebe, and Yuki Sugiyama. Instability of pedestrian flow in 2D optimal velocity model with attractive interaction. Computer Physics Communications, 177(1–2):162–163, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000811.

Nakamura:2001:IAA

[NI01] Kazuma Nakamura and Atushi Ikawa. Infrared absorption in amorphous selenium. Computer Physics Communications, 142(1–3):295–299, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003642.

Niimura:2000:LGM

[Nii00] Hiroaki Niimura. Lattice-gas model with wetness control for various deformable microstructures. Computer Physics Communications, 129(1–3):145–157, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001028.

Nikolopoulos:2003:PIC

[Nik03] L. A. A. Nikolopoulos. A package for the ab-initio calculation of one- and two-photon cross sections of two-electron atoms, using a CI B-splines method. Computer Physics Communications, 150(2):140–165, February 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006847.

Nilsen:2007:PSC

[Nil07a] Jon K. Nilsen. Python in scientific computing: Applications to Bose–Einstein condensates. Computer Physics Communications, 177(1–2):45, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507001312.

Nilsen:2007:MIQ

[Nil07b] Jon Kristian Nilsen. MontePython: Implementing quantum Monte Carlo using Python. Computer Physics Communications, 177(10):799–814, November 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003141.

Niukkanen:2000:TSN

[Niu00] A. W. Niukkanen. A transformation of F_{4} suggestive of a new approach to symbolic manipulation programming. Computer Physics Communications, 126(1–2):137–140, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900538X.

Nielsen:2000:MTN

[NJ00] Ida M. B. Nielsen and Curtis L. Janssen. Multi-threading: a new dimension to massively parallel scientific computation. Computer Physics Communications, 128(1–2):238–244, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550000062X.

Nielsen:2001:NPF

[NJ01] Ida M. B. Nielsen and Curtis L. Janssen. A novel pseudospectral Fourier method for solving Poisson’s equation for a solute in a non-uniform dielectric. Computer Physics Communications, 136(1–2):29–36, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500002502.

Na:2007:PMA

[NK07] Yong-Su Na and J. Y. Kim. Predictive modelling of advanced tokamak scenarios. Computer Physics Communications, 177(1– 2):134, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000963.

Neaton:2005:ETO

[NKSL05] J. B. Neaton, K. H. Khoo, C. D. Spataru, and Steven G. Louie. Electron transport and optical properties of carbon nanostructures from first principles. Computer Physics Communications, 169(1–3):1–8, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001001.

Nikunen:2003:HWY

[NKV03] P. Nikunen, M. Karttunen, and I. Vattulainen. How would you integrate the equations of motion in dissipative particle dynamics simulations? Computer Physics Communications, 153(3):407–423, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002029.

Nam:2007:KSC

[NL07] Gimoon Nam and Nam-Kyung Lee. Kinetics of a semiflexible chain under tension. Computer Physics Communications, 177(1–2):188, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000689.

Nduwayo:2009:SBS

[NLC09] L. Nduwayo, R. Lindebaum, and N. Chetty. Suppressed bondsite percolation. Computer Physics Communications, 180 (4):503–508, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000034.

Natarajan:2001:AMG

[NM01a] A. Natarajan and N. Mohankumar. An accurate method for the generalized Fermi–Dirac integral. Computer Physics Communications, 137(3):361–365, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001539.

Nishino:2001:DRL

[NM01b] Masamichi Nishino and Seiji Miyashita. Dynamical response of local magnetic structures induced by inhomogeneities to time-dependent magnetic field. Computer Physics Communications, 142(1–3):382–386, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003873.

Nikolopoulos:2003:CSH

[NM03] L. A. A. Nikolopoulos and L. B. Madsen. Complex-scaled Hartree–Fock wave functions for the frozen core: a B spline approach. Computer Physics Communications, 151(1):47–59, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006823.

Norton:2006:FDT

[NN06] Guy V. Norton and Jorge C. Novarini. Finite-difference time-domain simulation of acoustic propagation in dispersive medium: An application to bubble clouds in the ocean. Computer Physics Communications, 174(12):961–965, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506000725.

Nhung:2009:DCC

[NN09] Dao Thi Nhung and Le Duc Ninh. D0C: a code to calculate scalar one-loop four-point integrals with complex masses. Computer Physics Communications, 180(11):2258–2267, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900232X.

Noble:2004:ENE

[Nob04] C. J. Noble. Evaluation of negative energy Coulomb (whittaker) functions. Computer Physics Communications, 159 (1):55–62, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504000025.

Novotny:2002:CSP

[Nov02] M. A. Novotny. Computational statistical physics: 21st century extrema. Computer Physics Communications, 146(1):132–133, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004472.

Niukkanen:2000:CGC

[NP00] A. W. Niukkanen and O. S. Paramonova. Computer generation of complicated transformations and reduction formulas for multiple hypergeometric series. Computer Physics Communications, 126(1–2):141–148, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599005147.

Nekovee:2001:RPC

[NP01a] Maziar Nekovee and J. M. Pitarke. Recent progress in the computational many-body theory of metal surfaces. Computer Physics Communications, 137(1):123–142, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001758.

Nikitas:2001:MCG

[NP01b] P. Nikitas and A. Papageorgiou. Modifications of the classical genetic algorithm for non-linear fitting applied to response surface modeling in HPLC. Computer Physics Communications, 141(2):225–229, November 30, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100399X.

Nieminen:2001:HSS

[NR01] T. A. Nieminen and J. E. R. Ross. How to SYN in seven easy steps. Computer Physics Communications, 142(1–3): 160–163, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003587.

Nieminen:2001:NMO

[NRDHB01] T. A. Nieminen, H. Rubinsztein-Dunlop, N. R. Heckenberg, and A. I. Bishop. Numerical modelling of optical trapping.

Computer Physics Communications, 142(1–3):468–471, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003915.

Nutaro:2001:AGT

[NRR01] T. Nutaro, S. Riyavong, and D. Ruffolo. Application of a generalized total variation diminishing algorithm to cosmic ray transport and acceleration. Computer Physics Communications, 134(2):209–222, February 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002034.

Nakamura:2001:FST

[NSKS01] Hiroaki Nakamura, Toshiharu Saiki, Hirotomo Kambe, and Keiji Sawada. FDTD simulation of tapered structure of near-field fiber probe. Computer Physics Communications, 142(1–3):464–467, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003927.

Nagasima:2002:GES

[NSMO02] Takehiro Nagasima, Yuji Sugita, Ayori Mitsutake, and Yuko Okamoto. Generalized-ensemble simulations of spin systems and protein systems. Computer Physics Communications, 146(1):69–76, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004368.

Neizvestny:2002:MEG

[NSYZ02] I. G. Neizvestny, N. L. Shwartz, Z. Sh. Yanovitskaya, and A. V. Zverev. 3D-model of epitaxial growth on porous 111 and 100 Si surfaces. Computer Physics Communications, 147(1–2):272–275, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002874.

Nanbu:2004:SMP

[NT04] Kenichi Nanbu and Lizhu Tong. Solution method of the Poisson equation for the electric field with a thin sheath. Computer Physics Communications, 164(1–3):428–433, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003145.

Nakamura:2005:PDS

[NT05] Hiroaki Nakamura and Yuichi Tamura. Phase diagram for self-assembly of amphiphilic molecule C_{12}E_{6} by dissipative particle dynamics simulation. Computer Physics Communications, 169(1–3):139–143, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500130X.

Nurhuda:2004:DDS

[Nur04] M. Nurhuda. Does the dynamic stabilization reflect the numerical instability of direct integration of time dependent Schrödinger equation? Computer Physics Communications, 162(1):1–8, September 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002127.

Nam:2009:GMH

[NV09] Sang Ki Nam and John P. Verboncoeur. Global model for high power microwave breakdown at high pressure in air. Computer Physics Communications, 180(4):628–635, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550800430X.

Naumkin:2002:DMP

[NW02a] F. Y. Naumkin and D. J. Wales. Diatomics-in-molecules potentials incorporating ab initio data: Application to ionic, Rydberg-excited, and molecule-doped rare gas clusters. Computer Physics Communications, 145(1):141–155, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001510.

Nestler:2002:PFM

[NW02b] Britta Nestler and Adam A. Wheeler. Phase-field modeling of multi-phase solidification. Computer Physics Communications, 147(1–2):230–233, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002527.

Nikezic:2006:CPT

[NY06] D. Nikezic and K. N. Yu. Computer program TRACK TEST for calculating parameters and plotting profiles for etch pits in nuclear track materials. Computer Physics Communications, 174(2):160–165, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005187.

Nobusada:2007:ECR

[NY07] Katsuyuki Nobusada and Kazuhiro Yabana. Electric currents in ring-shaped molecules induced by circularly polarized laser pulses. Computer Physics Communications, 177(1–2): 54, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001361.

Nikezic:2008:CPT

[NY08] D. Nikezic and K. N. Yu. Computer program TRACK VISION for simulating optical appearance of etched tracks in CR-39 nuclear track detectors. Computer Physics Communications, 178(8):591–595, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507004717.

Numata:2004:NTD

[NYH04] Ryusuke Numata, Zensho Yoshida, and Takaya Hayashi. Nonlinear three-dimensional simulation for self-organization and flow generation in two-fluid plasmas. Computer Physics Communications, 164(1–3):291–296, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002942.

Orlandini:2009:VCS

[OBG09] S. Orlandini, I. Baccarelli, and F. A. Gianturco. Variational calculations of structures and energetics in very floppy trimers: a new computational implementation. Computer Physics Communications, 180(3):384–391, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800372X.

Omeltchenko:2000:SLS

[OCK^{+}00] Andrey Omeltchenko, Timothy J. Campbell, Rajiv K. Kalia, Xinlian Liu, Aiichiro Nakano, and Priya Vashishta. Scalable I/O of large-scale molecular dynamics simulations: a data-compression algorithm. Computer Physics Communications, 131(1–2):78–85, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000837.

Ogata:2003:SPI

[OCK^{+}03] Shuji Ogata, Timothy J. Campbell, Rajiv K. Kalia, Aiichiro Nakano, Priya Vashishta, and Satyavani Vemparala. Scalable and portable implementation of the fast multipole method on parallel computers. Computer Physics Communications, 153(3):445–461, July 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002467.

Ouared:2008:TMI

[OCS^{+}08] R. Ouared, B. Chopard, B. Stahl, D. A. Rüfenacht, H. Yilmaz, and G. Courbebaisse. Thrombosis modeling in intracranial aneurysms: a lattice Boltzmann numerical algorithm. Computer Physics Communications, 179(1–3):128–131, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000283.

Oh:2007:EPI

[OD07] Kwang Jin Oh and Yuefan Deng. An efficient parallel implementation of the smooth particle mesh Ewald method for molecular dynamics simulations. Computer Physics Communications, 177(5):426–431, September 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002408.

Oberhofer:2008:OBF

[OD08] Harald Oberhofer and Christoph Dellago. Optimum bias for fast-switching free energy calculations. Computer Physics Communications, 179(1–3):41–45, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000131.

Ozdogan:2002:PTB

[ÖDÇ02] Cem Özdogan, Gülay Dereli, and Tahir Çagin. O(N) parallel tight binding molecular dynamics simulation of carbon nanotubes. Computer Physics Communications, 148(2):188–205, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005532.

Okuyan:2007:PIE

[OGG07] Erhan Okuyan, Ugur Güdükbay, and Oguz Gülseren. Pattern information extraction from crystal structures. Computer Physics Communications, 176(7):486–506, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000240.

Ortiz:2002:SFQ

[OGKL02] G. Ortiz, J. E. Gubernatis, E. Knill, and R. Laflamme. Simulating fermions on a quantum computer. Computer Physics Communications, 146(3):302–316, July 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004216.

Ovcharenko:2003:GGM

[OGWH03] A. M. Ovcharenko, S. I. Golubov, C. H. Woo, and Hanchen Huang. GMIC++: Grouping method in C++: an efficient method to solve large number of Master equations. Computer Physics Communications, 152(2):208–226, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502008160.

Ohl:2004:PA

[Ohl04] Thorsten Ohl. Publisher’s announcement. Computer Physics Communications, 163(1):ix, October 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504004564.

Ohsaki:2002:RIC

[OIKN02] A. Ohsaki, A. Igarashi, T. Kai, and S. Nakazaki. Radial integrals in the Coulomb–Born approximation: Multipole transitions. Computer Physics Communications, 147 (3):826–833, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004587.

Oh:2006:GPP

[OK06a] Kwang Jin Oh and Michael L. Klein. A general purpose parallel molecular dynamics simulation program. Computer Physics Communications, 174(7):560–568, April 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505006235.

Oh:2006:PMD

[OK06b] Kwang Jin Oh and Michael L. Klein. A parallel molecular dynamics simulation scheme for a molecular system with bond constraints in NPT ensemble. Computer Physics Communications, 174(4):263–269, February 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005837.

Ohno:2009:VSD

[OK09] Nobuaki Ohno and Akira Kageyama. Visualization of spherical data by Yin–Yang grid. Computer Physics Communications, 180(9):1534–1538, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001180.

Okamoto:2001:PFS

[Oka01] Yuko Okamoto. Protein folding simulations and structure predictions. Computer Physics Communications, 142(1–3):

55–63, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003162.

Ogoyski:2004:COSb

[OKS04] A. I. Ogoyski, S. Kawata, and T. Someya. Code OK2 — a simulation code of ion-beam illumination on an arbitrary shape and structure target. Computer Physics Communications, 161(3):143–150, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002139.

Olivieri:2001:OFM

[Oli01] Mauro Olivieri. Overview on a formal model of architecture/circuit trade-offs for the implementation of fast processors. Computer Physics Communications, 139(1):144–150, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002363.

Ogata:2001:HFE

[OLS^{+}01] Shuji Ogata, Elefterios Lidorikis, Fuyuki Shimojo, Aiichiro Nakano, Priya Vashishta, and Rajiv K. Kalia. Hybrid finiteelement/molecular-dynamics/electronic-density-functional approach to materials simulations on parallel computers. Computer Physics Communications, 138(2):143–154, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100203X.

Ostrikov:2007:CPN

[OLX07] K. Ostrikov, I. Levchenko, and S. Xu. Computational plasma nanoscience: Where plasma physics meets surface science. Computer Physics Communications, 177(1–2):110–113, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001270.

ORourke:2000:IAN

[OMC00] S. F. C. O’Rourke, D. M. McSherry, and D. S. F. Crothers. ION-ATOM/NEON — calculation of ionization cross sections by fast ion impact for neutral target atoms ranging from lithium to neon. Computer Physics Communications, 131(1–2):129–141, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000758.

Omelyan:2002:OFR

[OMF02] I. P. Omelyan, I. M. Mryglod, and R. Folk. Optimized Forest–Ruth- and Suzuki-like algorithms for integration of motion in many-body systems. Computer Physics Communications, 146(2):188–202, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502004514.

Omelyan:2003:SAI

[OMF03] I. P. Omelyan, I. M. Mryglod, and R. Folk. Symplectic analytically integrable decomposition algorithms: classification, derivation, and application to molecular dynamics, quantum and celestial mechanics simulations. Computer Physics Communications, 151(3):272–314, April 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007543.

Oquendo:2009:OTB

[OML09] W. F. Oquendo, J. D. Muñoz, and A. Lizcano. Oedometric test, Bauer’s law and the micro-macro connection for a dry sand. Computer Physics Communications, 180(4):616– 620, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000046.

Orlandini:2005:SDC

[OMY05] Enzo Orlandini, Davide Marenduzzo, and Julia M. Yeomans. Shear dynamics in cholesterics. Computer Physics Communications, 169(1–3):122–125, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001268.

Oliveira:2008:GRP

[ON08] Micael J. T. Oliveira and Fernando Nogueira. Generating relativistic pseudo-potentials with explicit incorporation of semicore states using APE, the atomic pseudo-potentials engine. Computer Physics Communications, 178(7):524–534, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004651.

Okumura:2005:MSM

[OO05] Hisashi Okumura and Yuko Okamoto. Molecular simulations in the multibaric-multithermal ensembles. Computer Physics Communications, 169(1–3):317–321, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001682.

Osterloh:2009:PSR

[OPB^{+}09] Lukas Osterloh, Carlos Pérez, David Böhme, José Maria Baldasano, Christine Böckmann, Lars Schneidenbach, and David Vicente. Parallel software for retrieval of aerosol distribution from LIDAR data in the framework of EARLINET– ASOS. Computer Physics Communications, 180(11):2095– 2102, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550900191X.

Opletal:2008:HHR

[OPO^{+}08] G. Opletal, T. C. Petersen, B. O’Malley, I. K. Snook, D. G. McCulloch, and I. Yarovsky. HRMC: Hybrid reverse Monte Carlo method with silicon and carbon potentials. Computer Physics Communications, 178(10):777–787, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000039.

Organtini:2001:STR

[Org01] Giovanni Organtini. Software tools at the Rome CMS/ ECAL Regional Center. Computer Physics Communications, 140(1–2):92–101, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002594.

Ortiz:2000:HOI

[Ort00] Patricio F. Ortiz. The hierarchical organization of information in astronomy or “how to organize things to find them without much effort”. Computer Physics Communications, 127(2–3):188–197, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599005081.

Orszag:2000:CPP

[OS00a] Steven A. Orszag and I. Staroselsky. CFD: Progress and problems. Computer Physics Communications, 127(1):165–171, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000321.

Ososkov:2000:GWF

[OS00b] G. Ososkov and A. Shitov. Gaussian wavelet features and their applications for analysis of discretized signals. Computer Physics Communications, 126(1–2):149–157, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599002271.

Orban:2003:ETM

[OS03] A. Orbán and B. Sulik. The evaluation of transition matrix elements for atomic excitations by a screened Coulomb potential. Computer Physics Communications, 151(2):199–215, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006963.

Oliveira:2004:GOM

[OS04] O. Oliveira and P. J. Silva. A global optimization method for Landau gauge fixing in lattice QCD. Computer Physics Communications, 158(2):73–88, April 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503005423.

Ogata:2002:HQM

[OSK^{+}02] Shuji Ogata, Fuyuki Shimojo, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Hybrid quantum mechanical/molecular dynamics simulation on parallel computers: density functional theory on real-space multigrids. Computer Physics Communications, 149(1):30–38, November 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006124.

Ogoyski:2004:COSa

[OSK04] A. I. Ogoyski, T. Someya, and S. Kawata. Code OK1 — simulation of multi-beam irradiation on a spherical target in heavy ion fusion. Computer Physics Communications, 157(2):160–172, February 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004922.

Otsuka:2001:PDO

[Ots01] Hiromi Otsuka. Phase diagrams of the one-dimensional anisotropic extended Hubbard model. Computer Physics Communications, 142(1–3):181–185, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100323X.

Okabe:2002:ANM

[OTY02] Yutaka Okabe, Yusuke Tomita, and Chiaki Yamaguchi. Application of new Monte Carlo algorithms to random spin systems. Computer Physics Communications, 146(1):63–68, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004356.

Oettel:2002:RTQ

[OvSA02] M. Oettel, L. von Smekal, and R. Alkofer. Relativistic threequark bound states in separable two-quark approximation. Computer Physics Communications, 144(1):63–81, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004659.

Persson:2007:FBF

[PAD07] Clas Persson and Claudia Ambrosch-Draxl. A full-band FPLAPW + k · p-method for solving the Kohn–Sham equation. Computer Physics Communications, 177(3):280–287, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001841.

Parteli:2009:STM

[PAD^{+}09] Eric J. R. Parteli, Murilo P. Almeida, Orencio Durán, José S. Andrade, Jr., and Hans J. Herrmann. Sand transport on Mars. Computer Physics Communications, 180(4):609–611, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004281.

Papadopoulos:2001:PPS

[Pap01] Costas G. Papadopoulos. PHEGAS: a phase-space generator for automatic cross-section computation. Computer Physics Communications, 137(2):247–254, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001631.

Parmananda:2004:BRB

[Par04] P. Parmananda. Book review: Nonlinear Dynamics Integrability, Chaos and Patterns: M. Lakshmanan and S. Rjaseekar, Springer-Verlag, Berlin, 2002, ISBN 3-540-43908-0. Computer Physics Communications, 161(3):179–180, August 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504001468.

Papadopoulos:2009:PFR

[PAS09] D. F. Papadopoulos, Z. A. Anastassi, and T. E. Simos. A phase-fitted Runge–Kutta–Nyström method for the numerical solution of initial value problems with oscillating solutions. Computer Physics Communications, 180(10):1839–1846, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001623.

Plummer:2009:TTP

[PAT^{+}09] M. Plummer, E. A. G. Armour, A. C. Todd, C. P. Franklin, and J. N. Cooper. ‘tt ‘tripleint cc’: a program for 2-centre variational leptonic Coulomb potential matrix elements using Hylleraas-type trial functions, with a performance optimization study. Computer Physics Communications, 180(12):2410–2423, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509002288.

Plascak:2009:CMB

[PB09a] J. A. Plascak and A. T. Bernardes. CCP2008 — modeling a better world. Computer Physics Communications, 180(4): v, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000368.

Pretty:2009:DMA

[PB09b] D. G. Pretty and B. D. Blackwell. A data mining algorithm for automated characterisation of fluctuations in multichannel timeseries. Computer Physics Communications, 180 (10):1768–1776, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001507.

Pankin:2004:NTU

[PBB^{+}04] Alexei Pankin, Glenn Bateman, Robert Budny, Arnold Kritz, Douglas McCune, Alexei Polevoi, and Irina Voitsekhovitch. Numerical techniques used in neutral beam injection modules. Computer Physics Communications, 164(1–3):421–427, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003133.

Patiris:2007:TIM

[PBI07] D. L. Patiris, K. Blekas, and K. G. Ioannides. TRIAC II. A Matlab code for track measurements from SSNT detectors. Computer Physics Communications, 177(3):329–338, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002196.

Pereverzev:2008:SNS

[PC08] G. V. Pereverzev and G. Corrigan. Stable numeric scheme for diffusion equation with a stiff transport. Computer Physics Communications, 179(8):579–585, October 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800218X.

Pleguezuelos:2007:HIH

[PCA^{+}07] E. Pleguezuelos, A. Carnicer, J. Andilla, E. Martín-Badosa, and M. Montes-Usategui. HoloTrap: Interactive hologram design for multiple dynamic optical trapping. Computer Physics Communications, 176(11–12):701–709, June 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002032.

Petrella:2001:LPS

[PCC01] G. Petrella, L. Cassidei, and F. Ciriaco. Light particle scattering diffraction from surfaces in the sudden approximation. Computer Physics Communications, 138(2):128–135, August 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002016.

Peeters:2009:NGK

[PCC^{+}09] A. G. Peeters, Y. Camenen, F. J. Casson, W. A. Hornsby, A. P. Snodin, D. Strintzi, and G. Szepesi. The nonlinear gyro-kinetic flux tube code GKW. Computer Physics Communications, 180(12):2650–2672, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002112.

Perger:2009:ICE

[PCCD09] W. F. Perger, J. Criswell, B. Civalleri, and R. Dovesi. Ab-initio calculation of elastic constants of crystalline systems with the CRYSTAL code. Computer Physics Communications, 180(10):1753–1759, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001453.

Pisanti:2008:PPA

[PCE^{+}08] O. Pisanti, A. Cirillo, S. Esposito, F. Iocco, G. Mangano, G. Miele, and P. D. Serpico. PArthENoPE: Public algorithm evaluating the nucleosynthesis of primordial elements. Computer Physics Communications, 178(12):956–971, June 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000921.

Pagonabarraga:2005:MLM

[PCF05] Ignacio Pagonabarraga, Fabrizio Capuani, and Daan Frenkel. Mesoscopic lattice modeling of electrokinetic phenomena. Computer Physics Communications, 169(1–3):192–196, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001402.

Parker:2000:EGI

[PCK00] Scott E. Parker, Yang Chen, and Charlson C. Kim. Electromagnetic gyrokinetic-ion drift-fluid-electron hybrid simulation. Computer Physics Communications, 127(1):59–70, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000278.

Popovich:2006:FWS

[PCV06] P. Popovich, W. A. Cooper, and L. Villard. A full-wave solver of the Maxwell’s equations in 3D cold plasmas. Computer Physics Communications, 175(4):250–263, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001457.

Pravia:2002:TNI

[PCYC02] Marco A. Pravia, Zhiying Chen, Jeffrey Yepez, and David G. Cory. Towards a NMR implementation of a quantum lattice gas algorithm. Computer Physics Communications, 146(3):339– 344, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004253.

Petrov:2008:TDE

[PD08] G. M. Petrov and J. Davis. A two-dimensional electromagnetic field algorithm for high-intensity laser-target interactions. Computer Physics Communications, 179(12):868–880, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002555.

Perrey-Debain:2006:BFT

[PDA06] Emmanuel Perrey-Debain and I. David Abrahams. A band factorization technique for transition matrix element asymptotics. Computer Physics Communications, 175(5):315–322, September 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002128.

Papageorgiou:2004:MNV

[PDL04] D. G. Papageorgiou, I. N. Demetropoulos, and I. E. Lagaris. MERLIN-3.1.1. A new version of the Merlin optimization environment. Computer Physics Communications, 159(1):70–71, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000761.

Pierleoni:2008:TWF

[PDM^{+}08] Carlo Pierleoni, Kris T. Delaney, Miguel A. Morales, David M. Ceperley, and Markus Holzmann. Trial wave functions for high-pressure metallic hydrogen. Computer Physics Communications, 179(1–3):89–97, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000210.

Peeters:2007:CFT

[Pee07] Kasper Peeters. Cadabra: a field-theory motivated symbolic computer algebra system. Computer Physics Communications, 176(8):550–558, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507000318.

Peterson:2004:SCI

[Pet04] Kyle Peterson. Self consistent impedance determination in multiple circuit paths for resistive magnetohydrodynamic zpinch flux compression simulations. Computer Physics Communications, 164(1–3):414–420, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003121.

Pagaran:2006:MPC

[PFG06a] J. Pagaran, S. Fritzsche, and G. Gaigalas. Maple procedures for the coupling of angular momenta. IX. Wigner Dfunctions and rotation matrices. Computer Physics Communications, 174(8):616–630, April 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550600004X.

Parpia:2006:GPL

[PFG06b] F. A. Parpia, C. Froese Fischer, and I. P. Grant. GRASP92: a package for large-scale relativistic atomic structure calculations. Computer Physics Communications, 175(11–12): 745–747, December 1–15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506003389.

Pitt-Francis:2009:CTD

[PFPB^{+}09] Joe Pitt-Francis, Pras Pathmanathan, Miguel O. Bernabeu, Rafel Bordas, Jonathan Cooper, Alexander G. Fletcher, Gary R. Mirams, Philip Murray, James M. Osborne, Alex Walter, S. Jon Chapman, Alan Garny, Ingeborg M. M. van Leeuwen, Philip K. Maini, Blanca Rodríguez, Sarah L. Waters, Jonathan P. Whiteley, Helen M. Byrne, and David J. Gavaghan. Chaste: a test-driven approach to software development for biological modelling. Computer Physics Communications, 180(12):2452–2471, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002604.

Phuong:2002:DCF

[PGS02] Nguyen Hoang Phuong, Guido Germano, and Friederike Schmid. The direct correlation function in nematic liquid crystals from computer simulation. Computer Physics Communications, 147(1–2):350–353, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502003028.

Patra:2007:LRI

[PHF^{+}07] Michael Patra, Marja T. Hyvönen, Emma Falck, Mohsen Sabouri-Ghomi, Ilpo Vattulainen, and Mikko Karttunen. Longrange interactions and parallel scalability in molecular simulations. Computer Physics Communications, 176(1):14–22, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506003146.

Pecoul:2002:NMC

[PHKL02] S. Pécoul, S. Heuraux, R. Koch, and G. Leclert. Numerical modeling of the coupling of an ICRH antenna with a plasma with self-consistent antenna currents. Computer Physics Communications, 146(2):166–187, July 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002692.

Pinhao:2001:PCK

[Pin01] Nuno R. Pinhão. PLASMAKIN: a chemical kinetics library for plasma physics modeling. Computer Physics Communications, 135(1):105–131, March 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002265.

Pisani:2000:DRN

[Pis00] Lorenzo Pisani. Dimensionality reduction in the numerical evaluation of electro-magnetic scattering: from the Kirchhoff’s integral method to a zero-dimensional approach. Computer Physics Communications, 125(1–3):83–89, March 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004610.

Pitzer:2005:ASC

[Pit05] Russell M. Pitzer. Atomic self-consistent-field program by the basis set expansion method: Columbus version. Computer Physics Communications, 170(3):239–264, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002638.

Pashov:2000:CPC

[PJK00] A. Pashov, W. Jastrzebski, and P. Kowalczyk. Construction of potential curves for diatomic molecular states by the IPA method. Computer Physics Communications, 128(3):622– 634, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000102.

Praprotnik:2008:SAS

[PJSK08] Matej Praprotnik, Christoph Junghans, Luigi Delle Site, and Kurt Kremer. Simulation approaches to soft matter: Generic statistical properties vs. chemical details. Computer Physics Communications, 179(1–3):51–60, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000155.

Pasqualucci:2001:PDF

[PK01] E. Pasqualucci and KLOE collaboration. Process and data flow control in KLOE. Computer Physics Communications, 140(1–2):139–145, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002636.

Paolucci:2001:MFM

[PKB^{+}01] Pier S. Paolucci, Philippe Kajfasz, Philippe Bonnot, Bernard Candaele, Daniel Maufroid, Elena Pastorelli, Andrea Ricciardi, Yves Fusella, and Eugenio Guarino. mAgic-FPU and MADE: a customizable VLIW core and the modular VLIW processor architecture description environments. Computer Physics Communications, 139(1):132–143, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002351.

Piecuch:2002:ECI

[PKKM02] Piotr Piecuch, Stanislaw A. Kucharski, Karol Kowalski, and Monika Musial. Efficient computer implementation of the renormalized coupled-cluster methods: The R-CCSD[T], R-CCSD(T), CR-CCSD[T], and CR-CCSD(T) approaches. Computer Physics Communications, 149(2):71–96, December 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502005982.

Perdiou:2002:AEC

[PKPV02] A. E. Perdiou, V. S. Kalantonis, E. A. Perdios, and M. N. Vrahatis. Application of efficient composite methods for computing with certainty periodic orbits in molecular systems. Computer Physics Communications, 148(2):227–235, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200560X.

Patriarca:2007:TDI

[PKRK07] M. Patriarca, A. Kuronen, M. Robles, and K. Kaski. Threedimensional interactive molecular dynamics program for the study of defect dynamics in crystals. Computer Physics Communications, 176(1):38–47, January 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003158.

Pask:2001:FEM

[PKSF01] J. E. Pask, B. M. Klein, P. A. Sterne, and C. Y. Fong. Finite-element methods in electronic-structure theory. Computer Physics Communications, 135(1):1–34, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002125.

Pfeiffer:2003:MSM

[PKST03] Harald P. Pfeiffer, Lawrence E. Kidder, Mark A. Scheel, and Saul A. Teukolsky. A multidomain spectral method for solving elliptic equations. Computer Physics Communications, 152(3):253–273, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502008470.

Passerone:2005:ECI

[PL05] Daniele Passerone and Teodoro Laino. Exploring conical intersection spaces using pseudo-dynamics and band optimization: a novel strategy. Computer Physics Communications, 169(1–3):305–308, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001657.

Liu:2003:MPF

[pLbL03] Yin ping Liu and Zhi bin Li. A Maple package for finding exact solitary wave solutions of coupled nonlinear evolution equations. Computer Physics Communications, 155(1):65–76, September 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503003473.

Pleimling:2002:PTS

[Ple02] Michel Pleimling. Phase transitions at surfaces, edges, and corners. Computer Physics Communications, 147(1–2):101–106, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002254.

Park:2007:APD

[PLL07] Yongjin Park, Rob J. W. E. Lahaye, and Young Hee Lee. Adsorption of pt on defective carbon nanotube walls: a DFT approach. Computer Physics Communications, 177(1–2):46, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001336.

Parks:2008:IPW

[PLPS08] Michael L. Parks, Richard B. Lehoucq, Steven J. Plimpton, and Stewart A. Silling. Implementing peridynamics within a molecular dynamics code. Computer Physics Communications, 179(11):777–783, December 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002348.

Penev:2009:TRE

[PLS09] Evgeni S. Penev, Sotiria Lampoudi, and Joan-Emma Shea. TiReX: Replica-exchange molecular dynamics using Tinker. Computer Physics Communications, 180(10):2013–2019, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001866.

Poschl:2000:CSG

[PM00] W. Pöschl and B. Müller. Collision of SU(2) gauge fields in 3+1 dimensions. Computer Physics Communications, 125(1–3):282–315, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004841.

Plewa:2001:AAM

[PM01] T. Plewa and E. Müller. AMRA: An adaptive mesh refinement hydrodynamic code for astrophysics. Computer Physics Communications, 138(2):101–127, August 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001990.

Paul:2002:ESS

[PM02] W. Paul and M. Müller. Enhanced sampling in simulations of dense systems. Computer Physics Communications, 146(1):113–117, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004423.

Pankin:2004:TMC

[PMA^{+}04] Alexei Pankin, Douglas McCune, Robert Andre, Glenn Bateman, and Arnold Kritz. The tokamak Monte Carlo fast ion module NUBEAM in the national transport code collaboration library. Computer Physics Communications, 159(3):157–184, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001109.

Pedram:2007:USM

[PMG07] P. Pedram, M. Mirzaei, and S. S. Gousheh. Using spectral method as an approximation for solving hyperbolic PDEs. Computer Physics Communications, 176(9–10):581–588, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550700032X.

Pisov:2008:MPA

[PMH08] Stoyan Pisov, Oksana Melikhova, and Marc Hou. Mechanical properties of agco nanostructured nanowires. Computer Physics Communications, 179(1–3):102–106, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000234.

Privman:2002:QCS

[PMV02] Vladimir Privman, Dima Mozyrsky, and Israel D. Vagner. Quantum computing with spin qubits in semiconductor structures. Computer Physics Communications, 146(3):331–338, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004241.

Pichl:2000:ATS

[PNH00] Lukás Pichl, Hiroki Nakamura, and Jiří Horácek. Analytical treatment of singular equations in dissociative recombination. Computer Physics Communications, 124(1):1–18, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004191.

Poghosyan:2005:PCS

[Pog05] V. A. Poghosyan. Package for calculations and simplifications of expressions with Dirac matrices (matrixexp). Computer Physics Communications, 170(3):287–295, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505003577.

Pointon:2008:SOE

[Poi08] T. D. Pointon. Second-order, exact charge conservation for electromagnetic particle-in-cell simulation in complex geometry. Computer Physics Communications, 179(8):535–544, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001653;   http:/
/www.sciencedirect.com/science/article/pii/S0010465509001684.

See erratum [Poi09].

Pointon:2009:ESO

[Poi09] T. D. Pointon. Erratum to “Second-order, exact charge conservation for electromagnetic particle-in-cell simulation in complex geometry” [Comput. Phys. Commun. 179 (2008) 535– 544]. Computer Physics Communications, 180(10):2024, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001684. See [Poi08].

Pomorski:2006:GHA

[Pom06] Krzysztof Pomorski. Gauss–Hermite approximation formula. Computer Physics Communications, 174(3):181–186, February 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505005175.

Popescu:2003:GMS

[Pop03] Lucretiu M. Popescu. A geometry modeling system for ray tracing or particle transport Monte Carlo simulation. Computer Physics Communications, 150(1):21–30, January 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200629X.

Portugal:2000:RPA

[Por00] R. Portugal. The riegeom package: abstract tensor calculation. Computer Physics Communications, 126(3):261–268, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004804.

Porod:2003:SPC

[Por03] W. Porod. SPheno, a program for calculating supersymmetric spectra, SUSY particle decays and SUSY particle production at e^{+}e^{−} colliders. Computer Physics Communications, 153(2):275–315, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503002224.

Potter:2000:JHV

[Pöt00] B. Pötter. JetViP 2.1: the hbook version. Computer Physics Communications, 133(1):105–118, December 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001582.

Pisov:2002:CDT

[PP02] S. Pisov and A. Proykova. Charge dependence of temperaturedriven phase transitions of molecular nanoclusters: Molecular dynamics simulation. Computer Physics Communications, 147(1–2):238–241, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002540.

Pena:2009:SER

[PP09] O. Peña and U. Pal. Scattering of electromagnetic radiation by a multilayered sphere. Computer Physics Communications, 180(11):2348–2354, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509002306.

Park:2007:FFC

[PPC07] Seong-Kwan Park, Kyong-Yop Park, and Hi Jun Choe. Flow field computation for the high voltage gas blast circuit breaker with the moving boundary. Computer Physics Communications, 177(9):729–737, November 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003530.

Pan:2004:HPL

[PPM04] Chongxun Pan, Jan F. Prins, and Cass T. Miller. A highperformance lattice Boltzmann implementation to model flow in porous media. Computer Physics Communications, 158 (2):89–105, April 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504000049.

Poli:2001:TBT

[PPP01] E. Poli, A. G. Peeters, and G. V. Pereverzev. TORBEAM, a beam tracing code for electron-cyclotron waves in tokamak plasmas. Computer Physics Communications, 136(1–2):90– 104, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001461.

Palazzari:2001:P

[PR01] Paolo Palazzari and Vittorio Rosato. Preface. Computer Physics Communications, 139(1):xi–xii, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002223.

Pozzorini:2006:PNE

[PR06] S. Pozzorini and E. Remiddi. Precise numerical evaluation of the two loop sunrise graph Master integrals in the equal mass case. Computer Physics Communications, 175(5):381–387, September 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506002153.

Portes:2009:CPD

[PRBD09] D. Portes, Jr., H. Rodrigues, B. Baseia, and S. B. Duarte. Computational procedure to determine quantum state evolution in Fock space. Computer Physics Communications, 180(2):226–230, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003196.

Proykova:2000:HIR

[Pro00] Ana Proykova. How to improve a random number generator. Computer Physics Communications, 124(2–3):125–131, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004348.

Paul:2008:PTS

[PRSB08] W. Paul, F. Rampf, T. Strauch, and K. Binder. Phase transitions in a single polymer chain: a micro-canonical analysis of Wang–Landau simulations. Computer Physics Communications, 179(1–3):17–20, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550800009X.

Pal:2008:FPS

[PS08] Hridis Kumar Pal and Alok Shukla. A Fortran 90 program to solve the Hartree–Fock equations for interacting spin-1/2 fermions confined in harmonic potentials. Computer Physics Communications, 179(4):267–274, August 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001446.

Poghosyan:2009:NPV

[PS09] Armen H. Poghosyan and Aram A. Shahinyan. A new parameter for validation molecular dynamics simulation (MD) data. Computer Physics Communications, 180(2):238–240, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003299.

Pongracz:2006:ADC

[PSH06] Barna Pongrácz, Gábor Szederkényi, and Katalin M. Hangos. An algorithm for determining a class of invariants in quasi-polynomial systems. Computer Physics Communications, 175(3):204–211, August 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506001287.

Pohn:2001:SFCa

[PSK01a] E. Pohn, M. Shoucri, and G. Kamelander. Study of the formation of a charge separation at a plasma edge. Part I. The numerical integration of a Vlasov equation possessing an invariant. Computer Physics Communications, 137(3):380–395, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001965.

Pohn:2001:SFCb

[PSK01b] E. Pohn, M. Shoucri, and G. Kamelander. Study of the formation of a charge separation at a plasma edge. Part II. Comparison between three different interpolation methods for the solution of the kinetic Vlasov equation. Computer Physics Communications, 137(3):396–404, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001849.

Pohn:2005:EVC

[PSK05] E. Pohn, M. Shoucri, and G. Kamelander. Eulerian Vlasov codes. Computer Physics Communications, 166(2):81–93, March 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504005107.

Plimpton:2003:LBA

[PSP^{+}03] Steven J. Plimpton, David B. Seidel, Michael F. Pasik, Rebecca S. Coats, and Gary R. Montry. A load-balancing algorithm for a parallel electromagnetic particle-in-cell code. Computer Physics Communications, 152(3):227–241, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007956.

Puzynin:2000:MFM

[PSV00] I. V. Puzynin, A. V. Selin, and S. I. Vinitsky. Magnusfactorized method for numerical solving the time-dependent Schrödinger equation. Computer Physics Communications, 126(1–2):158–161, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002258.

Paul:2000:LAC

[PSW00] T. Paul, J. Swain, and Z. Was. Library of anomalous ττγ couplings for τ^{+}τ^{−}(nγ) Monte Carlo programs. Computer Physics Communications, 124(2–3):243–246, February 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004397.

Picariello:2004:STM

[PTL04] Marco Picariello and Emilio Torrente-Lujan. Slavnov-Taylor1.0: a Mathematica package for computation in BRST formalism. Computer Physics Communications, 156(2):171– 179, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004624.

Publisher:2007:PNN

[Pub07] The Publisher. Publisher’s note — new editorial system. Computer Physics Communications, 176(1):ix, January 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506003936.

Puetzfeld:2006:PCA

[Pue06] Dirk Puetzfeld. PROCRUSTES: a computer algebra package for post-Newtonian calculations in General Relativity. Computer Physics Communications, 175(7):497–508, October 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002669.

Puri:2002:KWP

[Pur02] Sanjay Puri. Kinetics of wetting for phase-separating binary mixtures. Computer Physics Communications, 147(1– 2):286–291, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002898.

Petersen:2000:IEF

[PWH^{+}00] Max Petersen, Frank Wagner, Lars Hufnagel, Matthias Scheffler, Peter Blaha, and Karlheinz Schwarz. Improving the efficiency of FP-LAPW calculations. Computer Physics Communications, 126(3):294–309, April 11, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004956.

Pooley:2008:LBS

[PY08] C. M. Pooley and J. M. Yeomans. Lattice Boltzmann simulation techniques for simulating microscopic swimmers. Computer Physics Communications, 179(1–3):159–164, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000349.

Popp:2001:LTN

[PZ01] Andreas Popp and Karl-Heinz Zimmermann. On loop transformations of nested loops with affine dependencies. Computer Physics Communications, 139(1):90–103, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002326.

Peng:2000:DVQ

[PZW^{+}00] Tong Peng, Dong H. Zhang, Dun-You Wang, Yi-Min Li, and J. Z. H. Zhang. Dynasol: a visual quantum dynamics package. Computer Physics Communications, 128(1–2):492–495, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559900541X.

Qteish:2005:EEC

[QASF^{+}05] A. Qteish, A. I. Al-Sharif, M. Fuchs, M. Scheffler, S. Boeck, and J. Neugebauer. Exact-exchange calculations of the electronic structure of AlN, GaN and InN. Computer Physics Communications, 169(1–3):28–31, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001050.

Quinet:2005:RSR

[QCL05] Olivier Quinet, Benoît Champagne, and Paolo Lazzeretti. Rototranslational sum rules for static and dynamic polarisabilities. Computer Physics Communications, 173(3):131–139, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004339.

Quesada:2003:DRN

[QCML03] J. M. Quesada, R. Capote, A. Molina, and M. Lozano. Dispersion relations in the nuclear optical model. Computer Physics Communications, 153(1):97–105, June 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503001577.

Qiang:2004:TDP

[QG04] Ji Qiang and Robert L. Gluckstern. Three-dimensional Poisson solver for a charged beam with large aspect ratio in a conducting pipe. Computer Physics Communications, 160(2):120–128, July 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001298.

Quigley:2005:CPL

[QP05] D. Quigley and M. I. J. Probert. Constant pressure Langevin dynamics: theory and application. Computer Physics Communications, 169(1–3):322–325, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001694.

Qiang:2001:PPS

[QR01] Ji Qiang and Robert D. Ryne. Parallel 3D Poisson solver for a charged beam in a conducting pipe. Computer Physics Communications, 138(1):18–28, July 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001850.

Qiang:2000:FIO

[QRH00] Ji Qiang, Robert D. Ryne, and Salman Habib. Fortran implementation of object-oriented design in parallel beam dynamics simulations. Computer Physics Communications, 133(1):18–33, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000165X.

Qiang:2006:MIB

[QTL06] J. Qiang, D. Todd, and D. Leitner. A 3D model for ion beam formation and transport simulation. Computer Physics Communications, 175(6):416–423, September 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002165.

Quine:2007:DSI

[QTMH07] Brendan M. Quine, Valery Tarasyuk, Henok Mebrahtu, and Richard Hornsey. Determining star-image location: a new sub-pixel interpolation technique to process image centroids. Computer Physics Communications, 177(9):700–706, November 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003050.

Qiao:2009:EEG

[QWWZ09] Cong-Feng Qiao, Jian Wang, Jian Xiong Wang, and Yangheng Zheng. EtabFDC: An ηb event generator in hadroproduction at LHC. Computer Physics Communications, 180(1):61–68, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002701.

Xu:2008:SPI

[qX08] Gui qiong Xu. Searching for Painlevé integrable conditions of nonlinear PDEs with constant parameters using symbolic computation. Computer Physics Communications, 178(7):505–517, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004687.

Xu:2009:NPT

[qX09] Gui qiong Xu. A note on the Painlevé test for nonlinear variable-coefficient PDEs. Computer Physics Communications, 180(7):1137–1144, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000356.

Xu:2004:SCP

[qXbL04] Gui qiong Xu and Zhi bin Li. Symbolic computation of the Painlevé test for nonlinear partial differential equations using Maple. Computer Physics Communications, 161(1–2):65–75, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002103.

Qi:2007:FMD

[QxW07] Wei Qi and Jian xiong Wang. FDCSUSYDecay: An MSSM decay package. Computer Physics Communications, 177 (4):378–390, August 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507002214.

Ramos:2003:LMC

[Ram03] J. I. Ramos. Linearization methods in classical and quantum mechanics. Computer Physics Communications, 153(2):199– 208, June 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002261.

Ramos:2004:PQT

[Ram04] J. I. Ramos. Piecewise quasilinearization techniques for singular boundary-value problems. Computer Physics Communications, 158(1):12–25, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503005228.

Ramos:2005:NTO

[Ram05] J. I. Ramos. On the numerical treatment of an ordinary differential equation arising in one-dimensional non-Fickian diffusion problems. Computer Physics Communications, 170(3):231–238, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002687.

Ramadan:2010:AFA

[Ram10] Omar Ramadan. Addendum to: “A FDTD analysis on magnetized plasma of Epstein distribution and reflection calculation” [Computer Physics Communications 180 (1) (2009) 55–60]: a short comment on the equivalence of the shiftoperator FDTD method and the bilinear frequency approximation technique for modeling dispersive electromagnetic applications. Computer Physics Communications, 181(7):1275–1276, July 2010. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465510000822. See [Yan09].

Ramadan:2012:UME

[Ram12] Omar Ramadan. Unified matrix-exponential FDTD formulations for modeling electrically and magnetically dispersive materials. Computer Physics Communications, 183(5):1101– 1106, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465512000276.

Rapaport:2002:CMD

[Rap02a] D. C. Rapaport. Classical molecular dynamics-whither? Computer Physics Communications, 146(1):134, June 15, 2002.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004484.

Rapaport:2002:WWG

[Rap02b] D. C. Rapaport. The wonderful world of granular ratchets. Computer Physics Communications, 147(1–2):141–144, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002333.

Rapaport:2006:MAM

[Rap06] D. C. Rapaport. Multibillion-atom molecular dynamics simulation: Design considerations for vector-parallel processing. Computer Physics Communications, 174(7):521–529, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505006260.

Rapaport:2008:MDS

[Rap08] D. C. Rapaport. Molecular dynamics studies of swimming at the microscopic scale. Computer Physics Communications, 179(1–3):155–158, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508000337.

Rashidi:2009:MDT

[Ras09] M. M. Rashidi. The modified differential transform method for solving MHD boundary-layer equations. Computer Physics Communications, 180(11):2210–2217, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002094. See corrigendum [Ras17].

Rashidi:2017:CMD

[Ras17] M. M. Rashidi. Corrigendum to “The modified differential transform method for solving MHD boundary-layer equations” [comput. phys. comm. 180 (2009) 2210–2217]. Computer Physics Communications, 212(??):285, March 2017. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465516303435. See [Ras09].

Rylander:2000:SFF

[RB00] Thomas Rylander and Anders Bondeson. Stable FEM– FDTD hybrid method for Maxwell’s equations. Computer Physics Communications, 125(1–3):75–82, March 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004634.

Roy:2005:MAL

[RB05] Dhiranjan Roy and Ranjan Bhattacharya. Multivariate approximants with levin-like transforms. Computer Physics Communications, 172(1):1–18, October 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003590.

Rayson:2008:RIM

[RB08] M. J. Rayson and P. R. Briddon. Rapid iterative method for electronic-structure eigenproblems using localised basis functions. Computer Physics Communications, 178(2):128–134, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003803.

Ramirez-Cuesta:2004:ANV

[RC04] A. J. Ramirez-Cuesta. aCLIMAX 4.0.1, the new version of the software for analyzing and interpreting INS spectra. Computer Physics Communications, 157(3):226–238, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503005204.

Rontani:2005:CBL

[RCG05] Massimo Rontani, Carlo Cavazzoni, and Guido Goldoni. The crossover between liquid and solid electron phases in quantum dots: a large-scale configuration-interaction study. Computer Physics Communications, 169(1–3):430–432, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500192X.

Ravat:2000:MSQ

[RCGC00] B. Ravat, H. Chapelotte, M. Grivet, and A. Chambaudet. Model and simulation of quenching effects on the shape of a tritium β^{−} in a liquid scintillation counting. Computer Physics Communications, 127(2–3):261–267, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005056.

Reynal:2005:HMC

[RD05] S. Reynal and H. T. Diep. Hybrid multicanonical cluster algorithm for efficient simulations of long-range spin models. Computer Physics Communications, 169(1–3):243–246, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001530.

Reuter:2000:GFC

[RdAGV^{+}00] K. Reuter, P. L. de Andres, F. J. Garcia-Vidal, D. Sestovic, F. Flores, and K. Heinz. Green’s function calculation of ballistic electron emission microscopy currents (BEEM v2.1). Computer Physics Communications, 127(2–3):327–342, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005020.

Rejcek:2002:AFK

[RDFF02] J. M. Rejcek, S. Datta, N. G. Fazleev, and J. L. Fry. Application of the Feynman–Kac path integral method in finding excited states of quantum systems. Computer Physics Communications, 146(2):154–165, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002679.

Ren:2002:CAM

[RDS02a] Zhiliang Ren, Zhidong Deng, and Zengqi Sun. Cellular automaton modeling of computer network. Computer Physics Communications, 144(3):243–251, April 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501004684.

Ren:2002:STO

[RDS02b] Zhiliang Ren, Zhidong Deng, and Zengqi Sun. Spatio-temporal organization of a cellular automaton model for computer network. Computer Physics Communications, 144(3):310–316, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002722.

Ren:2001:APS

[RDSS01a] Zhiliang Ren, Zhidong Deng, Dianxun Shuai, and Zengqi Sun. Analysis of power spectrum and 1/f type power law in a complex computer network model. Computer Physics Communications, 136(3):225–235, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001606.

Ren:2001:BND

[RDSS01b] Zhiliang Ren, Zhidong Deng, Zengqi Sun, and Dianxun Shuai. Behaviours of networks with different topologies and protocols. Computer Physics Communications, 141(2):247–259, November 30, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004040.

Rashidi:2009:NAM

[RE09] M. M. Rashidi and E. Erfani. New analytical method for solving Burgers’ and nonlinear heat transfer equations and comparison with HAM. Computer Physics Communications, 180(9):1539–1544, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001192.

Ramdas:2008:ESS

[REAB08] Tirath Ramdas, Gregory K. Egan, David Abramson, and Kim K. Baldridge. On ERI sorting for SIMD execution of large-scale Hartree–Fock SCF. Computer Physics Communications, 178(11):817–834, June 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000635.

Ramdas:2009:ESE

[REAB09] Tirath Ramdas, Gregory K. Egan, David Abramson, and Kim K. Baldridge. ERI sorting for emerging processor architectures. Computer Physics Communications, 180(8):1221–1229, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000435.

Refson:2000:MPM

[Ref00] Keith Refson. Moldy: a portable molecular dynamics simulation program for serial and parallel computers. Computer Physics Communications, 126(3):310–329, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004968.

Rykhlinskaya:2004:UGT

[RF04] K. Rykhlinskaya and S. Fritzsche. Use of group theory for the analysis of vibrational spectra. Computer Physics Communications, 162(2):124–142, September 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003522.

Radtke:2005:SQQ

[RF05a] T. Radtke and S. Fritzsche. Simulation of n-qubit quantum systems. I. Quantum registers and quantum gates. Computer Physics Communications, 173(1–2):91–113, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505004261.

Rykhlinskaya:2005:GMS

[RF05b] K. Rykhlinskaya and S. Fritzsche. Generation of molecular symmetry orbitals for the point and double groups. Computer Physics Communications, 171(2):119–132, September 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500295X.

Radtke:2006:SQQ

[RF06a] T. Radtke and S. Fritzsche. Simulation of n-qubit quantum systems. II. Separability and entanglement. Computer Physics Communications, 175(2):145–166, July 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001445.

Rykhlinskaya:2006:GCG

[RF06b] K. Rykhlinskaya and S. Fritzsche. Generation of Clebsch– Gordan coefficients for the point and double groups. Computer Physics Communications, 174(11):903–913, June 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000531.

Radtke:2007:SQQ

[RF07] T. Radtke and S. Fritzsche. Simulation of n-qubit quantum systems. III. Quantum operations. Computer Physics Communications, 176(9–10):617–633, May 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507001804.

Radtke:2008:SQQ

[RF08] T. Radtke and S. Fritzsche. Simulation of n-qubit quantum systems. IV. Parametrizations of quantum states, matrices and probability distributions. Computer Physics Communications, 179(9):647–664, November 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002245.

Rohlf:2008:RMC

[RFK08] Katrin Rohlf, Simon Fraser, and Raymond Kapral. Reactive multiparticle collision dynamics. Computer Physics Communications, 179(1–3):132–139, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000295.

Robinson:2004:TDP

[RG04] Allen C. Robinson and Christopher J. Garasi. Threedimensional z-pinch wire array modeling with ALEGRA-HEDP. Computer Physics Communications, 164(1–3):408– 413, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550400311X.

Raty:2005:FPS

[RG05] Jean-Yves Raty and Giulia Galli. First principle study of nanodiamond optical and electronic properties. Computer Physics Communications, 169(1–3):14–19, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001025.

Rodriguez:2001:LCL

[RGD^{+}01] O. Rodríguez, F. García, H. Dias, J. Mesa, J. D. T. Arruda-Neto, E. Garrote, and F. Guzmán. LINDEN: Code for level density calculations of deformed nuclei using Lipkin– Nogami projectors in the BCS approach. Computer Physics Communications, 137(3):405–414, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001564.

Romero:2004:PIC

[RGR^{+}04] Sergio Romero, Eladio Gutiérrez, Luis F. Romero, Oscar Plata, and Emilio L. Zapata. Parallelization issues of a code for physically-based simulation of fabrics. Computer Physics Communications, 162(3):188–202, October 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400373X.

Rak:2001:KMC

[RIB01] Miroslawa Rak, Marek Izdebski, and Andrzej Brozi. Kinetic Monte Carlo study of crystal growth from solution. Computer Physics Communications, 138(3):250–263, August 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002387.

Ribeiro:2002:FCN

[Rib02] Marcelo B. Ribeiro. A Fortran code for null geodesic solutions in the Lemaître–Tolman–Bondi spacetime. Computer Physics Communications, 148(2):236–241, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502005611.

Rintoul:2002:NBC

[Rin02] Mark D. Rintoul. The new biology and computational statistical physics. Computer Physics Communications, 146(1):77–83, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200437X.

Rizea:2002:PPS

[Riz02] M. Rizea. PERSYS — a program for the solution near the origin of coupled channel Schrödinger equation with singular potential. Computer Physics Communications, 143(1):83–99, February 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004295.

Roy:2000:PMD

[RJCH00] S. Roy, R. Y. Jin, V. Chaudhary, and W. L. Hase. Parallel molecular dynamics simulations of alkane/hydroxylated α-aluminum oxide interfaces. Computer Physics Communications, 128(1–2):210–218, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000400.

Reuter:2008:PIM

[RJFB08] K. Reuter, F. Jenko, C. B. Forest, and R. A. Bayliss. A parallel implementation of an MHD code for the simulation of mechanically driven, turbulent dynamos in spherical geometry. Computer Physics Communications, 179(4):245–249, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000970.

Rosso:2005:VMC

[RK05] Alberto Rosso and Werner Krauth. Variant Monte Carlo algorithm for driven elastic strings in random media. Computer Physics Communications, 169(1–3):188–191, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001396.

Reinhard:2009:GPR

[RLH^{+}09] F. Reinhard, O. F. Lange, J. S. Hub, J. Haas, and H. Grubmüller. g permute: Permutation-reduced phase space density compaction. Computer Physics Communications, 180 (3):455–458, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003743.

Rewoldt:2007:LCG

[RLI07] G. Rewoldt, Z. Lin, and Y. Idomura. Linear comparison of gyrokinetic codes with trapped electrons. Computer Physics Communications, 177(10):775–780, November 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003268.

Radjenovic:2006:SFL

[RLRR06] Branislav Radjenović, Jae Koo Lee, and Marija Radmilović-Radjenović. Sparse field level set method for non-convex Hamiltonians in 3D plasma etching profile simulations. Computer Physics Communications, 174(2):127–132, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005151.

Ruiz-Lorenzo:2000:SMA

[RLU00] J. J. Ruiz-Lorenzo and C. L. Ullod. Study of a microcanonical algorithm on the ±J spin glass model in d = 3. Computer Physics Communications, 125(1–3):210–220, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004907.

Roters:2001:TAI

[RLU01] L. Roters, S. Lübeck, and K. D. Usadel. Thermally activated interface motion in a disordered ferromagnet. Computer Physics Communications, 142(1–3):361–363, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003757.

Rizea:2008:FDA

[RLV^{+}08] M. Rizea, V. Ledoux, M. Van Daele, G. Vanden Berghe, and N. Carjan. Finite difference approach for the two-dimensional Schrödinger equation with application to scission-neutron emission. Computer Physics Communications, 179(7):466–478, October 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001501.

Raynolds:2005:ACC

[RM05a] James E. Raynolds and Lenore R. Mullin. Applications of conformal computing techniques to problems in computational physics: the fast Fourier transform. Computer Physics Communications, 170(1):1–10, July 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002286.

Rigol:2005:ENA

[RM05b] Marcos Rigol and Alejandro Muramatsu. Exact numerical approach to hard-core bosons on one-dimensional lattices. Computer Physics Communications, 169(1–3):416–420, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550500189X.

Rauenzahn:2005:TAN

[RMK05] R. M. Rauenzahn, V. A. Mousseau, and D. A. Knoll. Temporal accuracy of the nonequilibrium radiation diffusion equations employing a saha ionization model. Computer Physics Communications, 172(2):109–118, November 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003966.

Ribeiro:2001:AMC

[RMLB01] E. M. S. Ribeiro, L. E. Machado, M.-T. Lee, and L. M. Brescansin. Application of the method of continued fractions to electron scattering by polyatomic molecules. Computer Physics Communications, 136(1–2):117–125, May 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001515.

Reith:2002:COS

[RMMP02] Dirk Reith, Hendrik Meyer, and Florian Müller-Plathe. CG-OPT: a software package for automatic force field design. Computer Physics Communications, 148(3):299–313, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005623.

Rogier:2007:NMD

[RMVQ07] François Rogier, Jean-Charles Mateo-Velez, and Géraldine Quinio. Numerical modeling of DC discharges in air flows. Computer Physics Communications, 177(1–2):78–79, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001117.

Roths:2001:GRM

[RMWH01] T. Roths, M. Marth, J. Weese, and J. Honerkamp. A generalized regularization method for nonlinear ill-posed problems enhanced for nonlinear regularization terms. Computer Physics Communications, 139(3):279–296, October 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100217X.

Roberts:2000:CAD

[Rob00] A. J. Roberts. Computer algebra derives correct initial conditions for low-dimensional dynamical models. Computer Physics Communications, 126(3):187–206, April 11, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004944.

Roberts:2001:HPI

[Rob01] A. J. Roberts. Holistic projection of initial conditions onto a finite difference approximation. Computer Physics Communications, 142(1–3):316–321, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003599.

Rossmanith:2004:HRC

[Ros04] James A. Rossmanith. A high-resolution constrained transport method with adaptive mesh refinement for ideal MHD. Computer Physics Communications, 164(1–3):128–133, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550400270X.

Rous:2001:TNI

[Rou01] P. J. Rous. Theory of negative ion formation at surfaces. Computer Physics Communications, 137(1):33–58, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001837.

Roy:2009:GAS

[Roy09] Dhiranjan Roy. Global approximation for some functions. Computer Physics Communications, 180(8):1315–1337, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000678.

Radev:2002:PJW

[RP02] R. Radev and A. Proykova. Parallel J–W Monte Carlo simulations of thermal phase changes in finite-size systems. Computer Physics Communications, 147(1–2):242–245, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002552.

Righi:2005:KMC

[RPD^{+}05] M. C. Righi, C. A. Pignedoli, R. Di Felice, C. M. Bertoni, and A. Catellani. Kinetic Monte Carlo simulations of C diffusion on \sqrt{3} × \sqrt{3} β-SiC(111) based on ab initio calculations. Computer Physics Communications, 169(1–3):50–53, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001104.

Ratto:2007:SDP

[RPY07] Marco Ratto, Andrea Pagano, and Peter Young. State dependent parameter metamodelling and sensitivity analysis.

Computer Physics Communications, 177(11):863–876, December 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003475.

Rossokhaty:2002:MMB

[RR02] V. Rossokhaty and N. Rossokhata. Mathematical model of a biosensor with multilayer charged membrane. Computer Physics Communications, 147(1–2):366–369, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502003065.

Rottler:2005:MFL

[RR05] Jörg Rottler and Mark O. Robbins. Macroscopic friction laws and shear yielding of glassy solids. Computer Physics Communications, 169(1–3):177–182, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001372.

Rosa:2009:EED

[RRCV09] Reinaldo R. Rosa, Fernando M. Ramos, Cesar A. Caretta, and Haroldo F. Campos Velho. Extreme event dynamics in the formation of galaxy-sized dark matter structures. Computer Physics Communications, 180(4):621–624, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004001.

Rosswog:2008:SBH

[RRRHD08] Stephan Rosswog, Enrico Ramirez-Ruiz, W. Raphael Hix, and M. Dan. Simulating black hole white dwarf encounters. Computer Physics Communications, 179(1–3):184–189, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000386.

Ringwald:2000:QMC

[RS00] A. Ringwald and F. Schrempp. QCDINS 2.0- A Monte Carlo generator for instanton-induced processes in deep-inelastic scattering. Computer Physics Communications, 132(3):267– 305, November 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550000148X.

Rios:2003:IAF

[RS03] Cibrán Santamarina Rios and Juan José Saborido Silva. An implementation of atomic form factors. Computer Physics Communications, 151(1):79–88, March 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502006872.

Rigol:2009:TPQ

[RS09] Marcos Rigol and Rajiv R. P. Singh. Thermodynamic properties of quantum lattice models from numerical linked cluster expansions. Computer Physics Communications, 180(4):540– 544, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003780.

Ritley:2001:DEP

[RSD01] K. A. Ritley, M. Schlestein, and H. Dosch. DataScan: An extensible program for image analysis in Java. Computer Physics Communications, 137(2):300–311, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001552.

Raiskinmaki:2000:SNS

[RSMK^{+}00] P. Raiskinmäki, A. Shakib-Manesh, A. Koponen, A. Jäsberg, M. Kataja, and J. Timonen. Simulations of non-spherical particles suspended in a shear flow. Computer Physics Communications, 129(1–3):185–195, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001065.

Ratto:2001:SAM

[RTS01] M. Ratto, S. Tarantola, and A. Saltelli. Sensitivity analysis in model calibration: GSA–GLUE approach. Computer Physics Communications, 136(3):212–224, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100159X.

Ramis:2009:MCC

[RtVR09] R. Ramis, J. Meyer ter Vehn, and J. Ramírez. MULTI2D — a computer code for two-dimensional radiation hydrodynamics. Computer Physics Communications, 180(6):977–994, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004505.

Romero:2008:FCS

[RTVZ08] L. F. Romero, Siham Tabik, Jesús M. Vías, and Emilio L. Zapata. Fast clear-sky solar irradiation computation for very large digital elevation models. Computer Physics Communications, 178(11):800–808, June 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000507.

Rzepiela:2002:LSD

[RvOvV02] Anna A. Rzepiela, Joost H. J. van Opheusden, and Ton van Vliet. Large shear deformation of particle gels studied by Brownian dynamics simulations. Computer Physics Communications, 147(1–2):303–306, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002928.

Roudnev:2000:ITT

[RY00] V. Roudnev and S. Yakovlev. Improved tensor-trick algorithm: application to helium trimer. Computer Physics Communications, 126(1–2):162–164, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000023.

Ryckaert:2005:LJE

[Ryc05] Jean-Paul Ryckaert. From local jumps to entangled chain dynamics in polyethylene melts. Computer Physics Communications, 169(1–3):89–94, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505001189.

Soler:2009:OFF

[SA09] José M. Soler and Eduardo Anglada. Optimal Fourier filtering of a function that is strictly confined within a sphere. Computer Physics Communications, 180(7):1134–1136, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000332.

Stanley:2002:ACS

[SAG^{+}02] H. E. Stanley, L. A. N. Amaral, P. Gopikrishnan, V. Plerou, and M. A. Salinger. Application of computational statistical physics to scale invariance and universality in economic phenomena. Computer Physics Communications, 146(1):84–92, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004381.

Sakamoto:2007:DGD

[Sak07] Hiroshi Sakamoto. Data grid deployment for high energy physics in japan. Computer Physics Communications, 177 (1–2):239–242, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000562.

Saltelli:2002:MBU

[Sal02] Andrea Saltelli. Making best use of model evaluations to compute sensitivity indices. Computer Physics Communications, 145(2):280–297, May 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002801.

Salek:2003:WPT

[Sal03] Pawel Salek. A wave-packet technique to simulate resonant X-ray scattering cross sections. Computer Physics Communications, 150(2):85–98, February 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200632X.

Sanna:2000:VSH

[San00] N. Sanna. Vector spherical harmonics: Concepts and applications to the single centre expansion method. Computer Physics Communications, 132(1–2):66–83, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001375.

Sarkadi:2000:FPC

[Sar00] L. Sarkadi. A Fortran program to calculate the matrix elements of the Coulomb interaction involving hydrogenic wave functions. Computer Physics Communications, 133(1):119–127, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001600.

Satz:2002:CPT

[Sat02] Helmut Satz. Cluster percolation and thermal critical behavior. Computer Physics Communications, 147(1–2):46–51, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002023.

Sauter:2000:IHO

[Sau00] Thilo Sauter. Integration of highly oscillatory functions. Computer Physics Communications, 125(1–3):119–126, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004658.

Subbarao:2004:CSG

[SAU^{+}04] D. Subbarao, Manish Arora, R. Uma, M. V. R. Phanendra, and Krishna K. Singh. Computer simulation of gyrophase self-organisation in a magnetoplasma. Computer Physics Communications, 164(1–3):477–481, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003236.

Savkli:2001:NEF

[Sav01]  Çetin Savkli. Nonperturbative evaluation of the few-body

states for scalar χ^{2}ϕ interaction. Computer Physics Communications, 135(3):312–328, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550000240X.

Stijnman:2003:PSP

[SBB03] M. A. Stijnman, R. H. Bisseling, and G. T. Barkema. Partitioning 3D space for parallel many-particle simulations. Computer Physics Communications, 149(3):121–134, January 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006288.

Subbarao:2004:CSC

[SBBM04] D. Subbarao, Karuna Batra, Manik Bali, and Sugata Mitra. Computer simulation of cylindrical laser beam selffocusing in a plasma. Computer Physics Communications, 164(1–3):472–476, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504003224.

Subba:2008:FMA

[SBCZ08] F. Subba, X. Bonnin, D. Coster, and R. Zanino. 2D fluid modeling of the ASDEX upgrade scrape-off layer up to the first wall. Computer Physics Communications, 179(1–3):194– 198, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000404.

Sciortino:2005:RCG

[SBD^{+}05] Francesco Sciortino, Sergey V. Buldyrev, Cristiano De Michele, Giuseppe Foffi, Neda Ghofraniha, Emilia La Nave, Angel Moreno, Stefano Mossa, Ivan Saika-Voivod, Piero Tartaglia, and Emanuela Zaccarelli. Routes to colloidal gel formation. Computer Physics Communications, 169(1–3):166–171, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001359.

Smolarski:2006:PSA

[SBD^{+}06] Dennis C. Smolarski, Ramesh Balakrishnan, Eduardo F. D’Azevedo, John W. Fettig, Bronson Messer, Anthony Mezzacappa, Faisal Saied, Paul E. Saylor, and F. Douglas Swesty. On the performance of SPAI and ADI-like preconditioners for core collapse supernova simulations in one spatial dimension.

Computer Physics Communications, 175(5):330–338, September 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002177.

Schiemann:2005:EET

[SBJ05] Reinhard Schiemann, Michael Bachmann, and Wolfhard Janke. Exact enumeration of three-dimensional lattice proteins. Computer Physics Communications, 166(1):8–16, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550400503X.

Sotnikov:2004:HSP

[SBL^{+}04] V. I. Sotnikov, B. S. Bauer, J. N. Leboeuf, P. Hellinger, P. Trávnícek, and V. Fiala. Hybrid simulations of Zpinches. Computer Physics Communications, 164(1–3):150– 155, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002735.

Schwarz:2002:ESC

[SBM02] K. Schwarz, P. Blaha, and G. K. H. Madsen. Electronic structure calculations of solids using the WIEN2k package for material sciences. Computer Physics Communications, 147(1–2):71–76, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002060.

Schneider:2004:CSC

[SBM^{+}04] Ralf Schneider, Xavier Bonnin, Neil McTaggart, Alex Runov, Matthias Borchardt, Jörg Riemann, Andreas Mutzke, Konstantin Matyash, Henry Leyh, Manoj Warrier, David Coster, Wolfgang Eckstein, and Renate Dohmen. Comprehensive suite of codes for plasma-edge modelling. Computer Physics Communications, 164(1–3):9–16, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002498.

Sanna:2009:SNR

[SBM09a] N. Sanna, I. Baccarelli, and G. Morelli. SCELib3.0: The new revision of SCELib, the parallel computational library of molecular properties in the Single Center Approach. Computer Physics Communications, 180(12):2544–2549, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900229X.

Sanna:2009:VPE

[SBM09b] N. Sanna, I. Baccarelli, and G. Morelli. The VOLSCAT package for electron and positron scattering of molecular targets: a new high throughput approach to crosssection and resonances computation. Computer Physics Communications, 180(12):2550–2562, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002331.

Szczesniak:2004:DSP

[SC04] Ireneusz Szcześniak and John R. Cary. dxhdf5: a software package for importing HDF5 physics data into OpenDX. Computer Physics Communications, 164(1–3):365–369, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003042.

Schlier:2004:ETQ

[Sch04] Ch. Schlier. Error trends in quasi-Monte Carlo integration. Computer Physics Communications, 159(2):93–105, May 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504000827.

Schanze:2005:ACN

[Sch05] Thomas Schanze. Approximation of the Crank–Nicholson method by the iterated dynamic-theta method. Computer Physics Communications, 165(1):15–17, January 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504004783.

Schanze:2006:EDT

[Sch06a] Thomas Schanze. An exact D-dimensional Tsallis random number generator for generalized simulated annealing. Computer Physics Communications, 175(11–12):708–712, December 1–15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002980.

Schreier:2006:OEL

[Sch06b] Franz Schreier. Optimized evaluation of a large sum of functions using a three-grid approach. Computer Physics Communications, 174(10):783–792, May 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000178.

Schiavi:2008:TMM

[Sch08] A. Schiavi. A transmission matrix model for ion beam energy deposition in stack detectors. Computer Physics Communications, 178(1):35–39, January 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003694.

Skouteris:2000:AQR

[SCM00] D. Skouteris, J. F. Castillo, and D. E. Manolopoulos. ABC: a quantum reactive scattering program. Computer Physics Communications, 133(1):128–135, December 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001673.

Sakai:2000:FMR

[SCO00] Tomonori Sakai, Yu Chen, and Hirotada Ohashi. Formation of micelle in the real-coded lattice gas. Computer Physics Communications, 129(1–3):75–81, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000941.

Scott:2009:F

[Sco09] N. Stanley Scott. Foreword. Computer Physics Communications, 180(12):v, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509003129.

Stokl:2007:DIH

[SD07] A. Stökl and E. A. Dorfi. 2-dimensional implicit hydrodynamics on adaptive grids. Computer Physics Communications, 177(11):815–831, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003153.

Schmid:2007:GML

[SDLW07] Friederike Schmid, Dominik Düchs, Olaf Lenz, and Beate West. A generic model for lipid monolayers, bilayers, and membranes. Computer Physics Communications, 177(1–2):168– 171, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000860.

Stoitsov:2005:ADS

[SDNR05] M. V. Stoitsov, J. Dobaczewski, W. Nazarewicz, and P. Ring. Axially deformed solution of the Skyrme–Hartree–Fock– Bogolyubov equations using the transformed harmonic oscillator basis. The program HFBTHO (v1.66p). Computer Physics Communications, 167(1):43–63, April 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000305.

Searle:2001:DV

[Sea01] B. G. Searle. DL visualize. Computer Physics Communications, 137(1):25–32, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501001709.

Seaton:2002:CFA

[Sea02a] M. J. Seaton. Coulomb functions for attractive and repulsive potentials and for positive and negative energies. Computer Physics Communications, 146(2):225–249, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002758.

Seaton:2002:FCC

[Sea02b] M. J. Seaton. fgh, a code for the calculation of Coulomb radial wave functions from series expansions. Computer Physics Communications, 146(2):250–253, July 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200276X.

Seaton:2002:NCN

[Sea02c] M. J. Seaton. numer, a code for Numerov integrations of Coulomb functions. Computer Physics Communications, 146(2):254–260, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002771.

Shasharina:2004:DGF

[SEC04a] Svetlana Shasharina, Robert Eger, and John Cary. Data grid for fusion simulations and experiments. Computer Physics Communications, 164(1–3):134–137, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002711.

Shasharina:2004:FAR

[SEC04b] Svetlana Shasharina, Robert Eger, and John Cary. FarSight: application for remote visualization. Computer Physics Communications, 164(1–3):402–407, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003108.

Susukita:2003:HAM

[SEE^{+}03] Ryutaro Susukita, Toshikazu Ebisuzaki, Bruce G. Elmegreen, Hideaki Furusawa, Kenya Kato, Atsushi Kawai, Yoshinao Kobayashi, Takahiro Koishi, Geoffrey D. McNiven, Tetsu Narumi, and Kenji Yasuoka. Hardware accelerator for molecular dynamics: MDGRAPE-2. Computer Physics Communications, 155(2):115–131, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503003497.

Sjostrand:2001:HEP

[SEF^{+}01] Torbjörn Sjöstrand, Patrik Edén, Christer Friberg, Leif Lönnblad, Gabriela Miu, Stephen Mrenna, and Emanuel Norrbin. High-energy-physics event generation with Pythia 6.1. Computer Physics Communications, 135(2):238–259, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002368.

Semenov:2009:LPA

[Sem09] A. V. Semenov. LanHEP — a package for the automatic generation of Feynman rules in field theory. Version 3.0. Computer Physics Communications, 180(3):431–454, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003718.

Sevastianov:2000:PSC

[Sev00] L. A. Sevastianov. The probability scheme of constructing the mathematical model of shadowed sputtering. Computer Physics Communications, 130(1–2):41–46, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000369.

Strandlie:2000:AMF

[SF00] A. Strandlie and R. Frühwirth. Adaptive multitrack fitting. Computer Physics Communications, 133(1):34–42, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001661.

Schilling:2005:NSA

[SF05] Tanja Schilling and Daan Frenkel. Nucleation in suspensions of anisotropic colloids. Computer Physics Communications, 169(1–3):117–121, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001256.

Speer:2006:GSF

[SF06] T. Speer and R. Frühwirth. A Gaussian-sum filter for vertex reconstruction. Computer Physics Communications, 174 (12):935–947, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000749.

Sonnendrucker:2004:VSB

[SFF^{+}04] E. Sonnendrücker, F. Filbet, A. Friedman, E. Oudet, and J.-L. Vay. Vlasov simulations of beams with a moving grid. Computer Physics Communications, 164(1–3):390–395, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550400308X.

Sfiligoi:2001:DHK

[Sfi01] I. Sfiligoi. Data handling in KLOE. Computer Physics Communications, 140(1–2):189–197, October 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002703.

Sfiligoi:2007:CC

[Sfi07] Igor Sfiligoi. CDF computing. Computer Physics Communications, 177(1–2):235–238, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000550.

Swendsen:2005:AIM

[SFR05] Robert H. Swendsen, Marc Fasnacht, and John M. Rosenberg. The adaptive integration method for calculating general free energy functions. Computer Physics Communications, 169(1–3):274–276, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001591.

Seeger:2001:RWF

[SFSH01] S. Seeger, A. Franz, C. Schulzky, and K. H. Hoffmann. Random walks on finitely ramified Sierpinski carpets. Computer Physics Communications, 134(3):307–316, March 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002083.

Suarez:2009:MSM

[SFSL09] J. Suarez, S. C. Farantos, S. Stamatiadis, and L. Lathouwers. A method for solving the molecular Schrödinger equation in Cartesian coordinates via angular momentum projection operators. Computer Physics Communications, 180(11):2025–2033, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001854.

Sanna:2000:SPC

[SG00a] N. Sanna and F. A. Gianturco. SCELib: a parallel computational library of molecular properties in the single center approach. Computer Physics Communications, 128(1–2):139–169, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000783.

Segura:2000:ETH

[SG00b] J. Segura and A. Gil. Evaluation of toroidal harmonics. Computer Physics Communications, 124(1):104–122, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004282.

Storer:2001:RMM

[SG01] Robin Storer and Henry Gardner. Resistive magnetohydrodynamic modeling for 3D plasmas. Computer Physics Communications, 142(1–3):343–348, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100371X.

Schroder:2004:FBS

[SG04a] G. F. Schröder and H. Grubmüller. FRETsg: Biomolecular structure model building from multiple FRET experiments. Computer Physics Communications, 158(3):150–157, April 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504000748.

Spirkin:2004:UPS

[SG04b] Anton Spirkin and Nikolaos A. Gatsonis. Unstructured 3D PIC simulations of the flow in a retarding potential analyzer. Computer Physics Communications, 164(1–3):383–389, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003078.

Schreiber:2005:DFI

[SG05] Michael Schreiber and Sibylle Gemming. Density-functional investigation of alloyed metallic nanowires. Computer Physics Communications, 169(1–3):57–59, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001128.

Schmitz:2006:CTS

[SG06] H. Schmitz and R. Grauer. Comparison of time splitting and backsubstitution methods for integrating Vlasov’s equation with magnetic fields. Computer Physics Communications, 175(2):86–92, July 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550600110X.

Shoucri:2003:IVE

[SGF03] M. Shoucri, H. Gerhauser, and K.-H Finken. Integration of the Vlasov equation along characteristics in one and two dimensions. Computer Physics Communications, 154(1):65–75, July 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002819.

Shoucri:2004:SGC

[SGF04] M. Shoucri, H. Gerhauser, and K.-H. Finken. Study of the generation of a charge separation and electric field at a plasma edge using Eulerian Vlasov codes in cylindrical geometry. Computer Physics Communications, 164(1–3):138–149, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002723.

Srinivas:2009:ISC

[SGK09] S. Srinivas, R. Gayathri, and M. Kothandapani. The influence of slip conditions, wall properties and heat transfer on MHD peristaltic transport. Computer Physics Communications, 180(11):2115–2122, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001945.

Skouteris:2009:PPE

[SGL09] Dimitris Skouteris, Osvaldo Gervasi, and Antonio Laganà. A program for performing exact quantum dynamics calculations using cylindrical polar coordinates: a nanotube application. Computer Physics Communications, 180(3):459–465, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003755.

SilvadeMenezes:2009:FBS

[SGM^{+}09] Fortunato Silva de Menezes, Davi Geiger, Ricardo Martins Abreu e Silva, George Distler, and Arefin Hu. Forwardbackward study of the stereo vision problem. Computer Physics Communications, 180(4):664–668, April 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000320.

Seeger:2005:SDC

[SH05] Steffen Seeger and Karl Heinz Hoffmann. On symbolic derivation of the cumulant equations. Computer Physics Communications, 168(3):165–176, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500233X.

Schubert:2006:SES

[SH06] Sven Schubert and Karl Heinz Hoffmann. The structure of enumerated spin glass state spaces. Computer Physics Communications, 174(3):191–197, February 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005710.

Shacham:2004:BRB

[Sha04] Mordechai Shacham. Book review: Scientific Computing with MATLAB: A. Quarteroni, F. Saleri, Springer-Verlag, Berlin, 2003, ISBN 3-540-44363-0. Computer Physics Communications, 161(3):183–185, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400147X.

Shebalin:2003:SAS

[She03] John V. Shebalin. A spectral algorithm for solving the relativistic Vlasov–Maxwell equations. Computer Physics Communications, 156(1):86–94, December 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004387.

Shevchenko:2008:SCB

[She08] I. I. Shevchenko. Symbolic computation of the Birkhoff normal form in the problem of stability of the triangular libration points. Computer Physics Communications, 178(9):665–672, May 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004924.

Sugie:2004:SPCa

[SHH^{+}04] Takashige Sugie, Keiji Horita, Satoru Hosono, Tomoyoshi Ito, and Toshikazu Ebisuzaki. A special-purpose computer for exploring similar protein sequences by the dynamic programming method. Computer Physics Communications, 157(1):1–8, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004879.

Shimobaba:2002:SPC

[SHI02] Tomoyoshi Shimobaba, Sinsuke Hishinuma, and Tomoyoshi Ito. Special-purpose computer for holography HORN-4 with recurrence algorithm. Computer Physics Communications, 148(2):160–170, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502004733.

Shiers:2007:WLC

[Shi07] Jamie Shiers. The Worldwide LHC Computing Grid (worldwide LCG). Computer Physics Communications, 177(1– 2):219–223, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550700077X.

Shiers:2009:GTC

[Shi09] Jamie Shiers. Grid today, clouds on the horizon. Computer Physics Communications, 180(4):559–563, April 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004153.

Simek:2007:CSN

[SHJ07] J. Simek, R. Hrach, and P. Jelinek. Computational study of negative ions influence on plasma sheath formation. Computer Physics Communications, 177(1–2):137, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000951.

Shoucri:2004:AFS

[Sho04] M. Shoucri. The application of a fractional steps method for the numerical solution of the shallow water equations. Computer Physics Communications, 164(1–3):396–401, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504003091.

Shoucri:2007:NSS

[Sho07] Magdi Shoucri. Numerical solution of the shallow water equations with a fractional step method. Computer Physics Communications, 176(1):23–32, January 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003183.

Spray:2008:PLQ

[SHT08] J. Spray, J. Hill, and A. Trew. Performance of a lattice quantum chromodynamics kernel on the cell processor. Computer Physics Communications, 179(9):642–646, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002294.

Saldin:2001:SSC

[SHV^{+}01] D. K. Saldin, R. Harder, H. Vogler, W. Moritz, and I. K. Robinson. Solving the structure completion problem in surface crystallography. Computer Physics Communications, 137(1):12–24, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501001692.

Sanchez:2001:IMC

[SHW01] R. Sanchez, S. P. Hirshman, and H. V. Wong. Improved magnetic coordinate representation for ideal ballooning stability calculations with the COBRA code. Computer Physics Communications, 135(1):82–92, March 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002253.

Shen:2002:FDE

[SHX02] Linfang Shen, Sailing He, and Sanshui Xiao. A finitedifference eigenvalue algorithm for calculating the band structure of a photonic crystal. Computer Physics Communications, 143(3):213–221, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501004568.

Shimojo:2001:IMD

[SHZ01] Fuyuki Shimojo, Kozo Hoshino, and Y. Zempo. Ab initio molecular-dynamics simulation method for complex liquids. Computer Physics Communications, 142(1–3):364–367, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003770.

Shimobaba:2001:ECM

[SI01] Tomoyoshi Shimobaba and Tomoyoshi Ito. An efficient computational method suitable for hardware of computer-generated hologram with phase computation by addition. Computer Physics Communications, 138(1):44–52, July 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001898.

Sugie:2004:SPCb

[SIE04] Takashige Sugie, Tomoyoshi Ito, and Toshikazu Ebisuzaki. A special-purpose computer for exploring similar biological sequences: Bioler-2 with multi-pipeline and multi-sequence architecture. Computer Physics Communications, 162(1):37– 50, September 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504003273.

Sanchez:2001:IMS

[SIH^{+}01] R. Sanchez, M. Yu. Isaev, S. P. Hirshman, W. A. Cooper, G. Y. Fu, J. A. Jimenez, L. P. Ku, M. I. Mikhailov, D. A. Monticello, A. H. Reiman, and A. A. Subbotin. Ideal MHD stability calculations for compact stellarators. Computer Physics Communications, 141(1):55–65, November 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003964.

Simos:2000:AEO

[Sim00] T. E. Simos. An accurate eighth order exponentially-fitted method for the efficient solution of the Schrödinger equation. Computer Physics Communications, 125(1–3):21–59, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004592.

Simos:2008:HOC

[Sim08] T. E. Simos. High-order closed Newton–Cotes trigonometricallyfitted formulae for long-time integration of orbital problems. Computer Physics Communications, 178(3):199–207, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004067.

Simos:2009:STF

[Sim09] T. E. Simos. P -stability, trigonometric-fitting and the numerical solution of the radial Schrödinger equation. Computer Physics Communications, 180(7):1072–1085, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004463.

Smirnov:2002:ESS

[SJ02] A. V. Smirnov and D. D. Johnson. Efficiency and scaling of the screened Korringa–Kohn–Rostoker method for large-scale calculations. Computer Physics Communications, 148(1):74–80, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004691.

Samtaney:2004:AMR

[SJCM04] R. Samtaney, S. C. Jardin, P. Colella, and D. F. Martin. 3D adaptive mesh refinement simulations of pellet injection in tokamaks. Computer Physics Communications, 164(1–3):220–228, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400284X.

Scott:2007:NHC

[SJDC07] N. S. Scott, F. Jézéquel, C. Denis, and J.-M. Chesneaux. Numerical ‘health check’ for scientific codes: the CADNA approach. Computer Physics Communications, 176(8):507–521, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000331.

Sturesson:2007:JFP

[SJF07] L. Sturesson, P. Jönsson, and C. Froese Fischer. JJ-GEN: a flexible program for generating lists of jj-coupled configuration state functions. Computer Physics Communications, 177(6):539–550, September 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002548.

Shin:2007:TCI

[SJHY07] S. H. Shin, D. Y. Jeong, J. E. Han, and M. Yoon. A tunable coherent IR radiation source based on a selfamplified spontaneous emission free-electron laser. Computer Physics Communications, 177(1–2):136, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507001038.

Salvat:2005:EDP

[SJP05] Francesc Salvat, Aleksander Jablonski, and Cedric J. Powell. elsepa — Dirac partial-wave calculation of elastic scattering of electrons and positrons by atoms, positive ions and molecules. Computer Physics Communications, 165(2):157–190, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004795.

Sergi:2005:NCR

[SK05] Alessandro Sergi and Raymond Kapral. Nonadiabatic chemical reactions. Computer Physics Communications, 169(1–3):400– 403, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001852.

Schreier:2008:OIR

[SK08] Franz Schreier and Dieter Kohlert. Optimized implementations of rational approximations — a case study on the Voigt and complex error function. Computer Physics Communications, 179(7):457–465, October 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001495.

Skandera:2005:DSC

[Ska05] Dan Skandera. Different strategies of the central approach to the numerical hydrodynamics. Computer Physics Communications, 169(1–3):238–242, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001529.

Surzhykov:2005:ATD

[SKF05] Andrey Surzhykov, Peter Koval, and Stephan Fritzsche. Algebraic tools for dealing with the atomic shell model. I. Wavefunctions and integrals for hydrogen-like ions. Computer Physics Communications, 165(2):139–156, January 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004643.

Samson:2002:CTE

[SKH02a] Claire C. M. Samson, Wim Klopper, and Trygve Helgaker. Computation of two-electron Gaussian integrals for wave functions including the correlation factor r_{12} exp(−γr_{12}^{2}). Computer Physics Communications, 149(1):1–10, November 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502005908.

Stankovic:2002:RAL

[SKH02b] Igor Stankovic, Martin Kröger, and Siegfried Hess. Recognition and analysis of local structure in polycrystalline configurations. Computer Physics Communications, 145(3):371–384, June 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002655.

Shimojo:2001:LSD

[SKNV01] Fuyuki Shimojo, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Linear-scaling density-functional-theory calculations of electronic structure based on real-space grids: design, analysis, and scalability test of parallel algorithms. Computer Physics Communications, 140(3):303–314, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002478.

Sharma:2004:SPV

[SKNV04] Ashish Sharma, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Scalable and portable visualization of large atomistic datasets. Computer Physics Communications, 163(1): 53–64, October 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504004084.

Shimojo:2005:EDC

[SKNV05] Fuyuki Shimojo, Rajiv K. Kalia, Aiichiro Nakano, and Priya Vashishta. Embedded divide-and-conquer algorithm on hierarchical real-space grids: parallel molecular dynamics simulation based on linear-scaling density functional theory. Computer Physics Communications, 167(3):151–164, May 1, 2005.

CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000688.

Stupitsky:2004:NMB

[SKRK04] E. L. Stupitsky, A. S. Kholodov, A. Y. Repin, and Y. A. Kholodov. Numerical modeling of behavior of high-energy plasmoid in upper ionosphere. Computer Physics Communications, 164(1–3):258–261, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002899.

Souaille:2009:MPC

[SLBG09] M. Souaille, H. Loirat, D. Borgis, and M. P. Gaigeot. MD-VRY: a polarizable classical molecular dynamics package for biomolecules. Computer Physics Communications, 180(2): 276–301, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508002890.

Sahoo:2009:NCR

[SLC09] Kartika Chandra Sahoo, Yiming Li, and Edward Yi Chang. Numerical calculation of the reflectance of sub-wavelength structures on silicon nitride for solar cell application. Computer Physics Communications, 180(10):1721–1729, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001258.

Slepnyov:2000:ILV

[Sle00] S. K. Slepnyov. Integrals of Lotka–Volterra systems with partially broken symmetry. Computer Physics Communications, 126(1–2):165–167, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599002295.

Shon:2001:MIS

[SLL01] C. H. Shon, H. J. Lee, and J. K. Lee. Method to increase the simulation speed of particle-in-cell (PIC) code. Computer Physics Communications, 141(3):322–329, December 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004179.

Seo:2007:ILE

[SLL07] Y. S. Seo, S. M. Lee, and J. K. Lee. Improvement of luminous efficiency in a mercury-free flat fluorescent lamp (FFL) by using fluid simulation. Computer Physics Communications, 177(1–2): 135, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000975.

Schwartzkopff:2006:AHO

[SLMS06] T. Schwartzkopff, F. Lörcher, C.-D. Munz, and R. Schneider. Arbitrary high order finite-volume methods for electromagnetic wave propagation. Computer Physics Communications, 174(9):689–703, May 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506000130.

Shu:2002:PSM

[SLWH02] J. W. Shu, Q. Lu, W. O. Wong, and Hanchen Huang. Parallelization strategies for Monte Carlo simulations of thin film deposition. Computer Physics Communications, 144(1):34–45, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001388.

Sato:2001:EOA

[SM01] H. Sato and Y. Morita. Evaluation of objectivity/AMS on the wide area network. Computer Physics Communications, 140(1–2):233–238, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002740.

Stare:2002:NSV

[SM02] Jernej Stare and Janez Mavri. Numerical solving of the vibrational time-independent Schrödinger equation in one and two dimensions using the variational method. Computer Physics Communications, 143(3):222–240, March 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100457X.

Sanna:2004:SNR

[SM04] N. Sanna and G. Morelli. SCELib2: the new revision of SCELib, the parallel computational library of molecular properties in the single center approach. Computer Physics Communications, 162(1):51–78, September 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550400325X. See erratum [SM06a].

Sanna:2006:ESN

[SM06a] N. Sanna and G. Morelli. Erratum to “SCELib2: the new revision of SCELib, the parallel computational library of molecular properties in the single center approach” [Computer Physics Communications 162 (2004) 51–78]. Computer Physics Communications, 175(7):509, October 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002645. See [SM04].

Seo:2006:ASS

[SM06b] Songwon Seo and Oakkey Min. Axisymmetric SPH simulation of elasto-plastic contact in the low velocity impact. Computer Physics Communications, 175(9):583–603, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506002700.

Skylaris:2001:AKE

[SMH^{+}01] Chris-Kriton Skylaris, Arash A. Mostofi, Peter D. Haynes, Chris J. Pickard, and Mike C. Payne. Accurate kinetic energy evaluation in electronic structure calculations with localized functions on real space grids. Computer Physics Communications, 140(3):315–322, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550100248X.

Schaller:2004:KDD

[SMH04] Gernot Schaller and Michael Meyer-Hermann. Kinetic and dynamic Delaunay tetrahedralizations in three dimensions. Computer Physics Communications, 162(1):9–23, September 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003248.

Strange:2001:ACG

[SMK01] R. Strange, F. R. Manby, and P. J. Knowles. Automatic code generation in density functional theory. Computer Physics Communications, 136(3):310–318, May 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001485.

Shimobaba:2000:SPC

[SMS^{+}00] Tomoyoshi Shimobaba, Nobuyuki Masuda, Takashige Sugie, Satoru Hosono, Shinobu Tsukui, and Tomoyoshi Ito. Specialpurpose computer for holography HORN-3 with PLD technology. Computer Physics Communications, 130(1–2):75–82, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000448.

Sjostrand:2008:BIP

[SMS08] Torbjörn Sjöstrand, Stephen Mrenna, and Peter Skands. A brief introduction to PYTHIA 8.1. Computer Physics Communications, 178(11):852–867, June 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000441.

Sanchez:2003:ONT

[SMSE03] R. Sánchez, J. R. Martín-Solís, and B. Esposito. ORCO: a numerical tool to study the radial diffusion of runaway electrons in tokamaks. Computer Physics Communications, 156(1):95–107, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503004399.

Senatore:2001:SME

[SMV01] Gaetano Senatore, S. Moroni, and D. Varsano. Spontaneous magnetization of the 2D electron gas. Computer Physics Communications, 142(1–3):406–409, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003721.

Shepard:2005:SCE

[SMZ05] Ron Shepard, Michael Minkoff, and Yunkai Zhou. Software for computing eigenvalue bounds for iterative subspace matrix methods. Computer Physics Communications, 170(1):109–114, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002547.

Shiratori:2007:ESC

[SN07] K. Shiratori and K. Nobusada. Electronic structure calculations at constant chemical potential toward the application to electrochemistry. Computer Physics Communications, 177(1– 2):47, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001324.

Sunderland:2002:PMP

[SNBB02] A. G. Sunderland, C. J. Noble, V. M. Burke, and P. G. Burke. A parallel R-matrix program PRMAT for electronatom and electron-ion scattering calculations. Computer Physics Communications, 145(3):311–340, June 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001406.

Sanz-Navarro:2001:NCU

[SNS01] C. F. Sanz-Navarro and Roger Smith. Numerical calculations using the hyper-molecular dynamics simulation method. Computer Physics Communications, 137(1):206–221, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001801.

Signell:2008:SPS

[SOAW08] Artur Signell, Francisco Ogando, Mats Aspnäs, and Jan Westerholm. Scalable plasma simulation with ELMFIRE using efficient data structures for process communication. Computer Physics Communications, 179(5):330–338, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001343.

Soloveva:2001:NCE

[Sol01] T. M. Solov’eva. Numerical calculation of the energy spectrum of a two-fermion system. Computer Physics Communications, 136(3):208–211, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500002459.

Sornette:2002:PLE

[Sor02] D. Sornette. Pushing the limits: Extremes and crashes in finance and economics. Computer Physics Communications, 147(1–2):19–21, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502001984.

Sloot:2001:SOC

[SOS01] P. M. A. Sloot, B. J. Overeinder, and A. Schoneveld. Selforganized criticality in simulated correlated systems. Computer Physics Communications, 142(1–3):76–81, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003253.

Shima:2001:FOM

[SOYN01] H. Shima, H. Obuse, K. Yakubo, and T. Nakayama. The forced oscillator method incorporating the fast time-evolution algorithm. Computer Physics Communications, 142(1–3): 418–423, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003794.

Sbraccia:2005:ASC

[SPC^{+}05] C. Sbraccia, C. A. Pignedoli, A. Catellani, R. Di Felice, P. L. Silvestrelli, F. Toigo, F. Ancilotto, and C. M. Bertoni. Acetylene on Si(111) from computer simulations. Computer Physics Communications, 169(1–3):32–35, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001062.

Stamatiadis:2000:ATA

[SPF00] S. Stamatiadis, R. Prosmiti, and S. C. Farantos. auto deriv: Tool for automatic differentiation of a Fortran code. Computer Physics Communications, 127(2–3):343–355, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599005135.

Straatsma:2000:NEP

[SPM00] T. P. Straatsma, M. Philippopoulos, and J. A. McCammon. NWChem: Exploiting parallelism in molecular simulations. Computer Physics Communications, 128(1–2):377–385, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000540.

Strauss:2004:MSR

[SPP^{+}04] H. R. Strauss, A. Pletzer, W. Park, S. Jardin, J. Breslau, and L. Sugiyama. MHD simulations with resistive wall and magnetic separatrix. Computer Physics Communications, 164(1–3):40–45, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002541.

Singh:2009:AAL

[SPS09] Om P. Singh, Rajesh K. Pandey, and Vineet K. Singh. An analytic algorithm of Lane–Emden type equations arising in astrophysics using modified homotopy analysis method. Computer Physics Communications, 180(7):1116–1124, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000277.

Schapotschnikow:2007:SAA

[SPV07] Philipp Schapotschnikow, René Pool, and Thijs J. H. Vlugt. Selective adsorption of alkyl thiols on gold in different geometries. Computer Physics Communications, 177(1–2):154– 157, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000859.

Sun:2003:MSM

[SQ03] Jian-Qiang Sun and Meng-Zhao Qin. Multi-symplectic methods for the coupled 1D nonlinear Schrödinger system. Computer Physics Communications, 155(3):221–235, November 1– 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503002856.

Simpson:2001:ALM

[SR01a] M. E. Simpson and A. J. Roberts. An accurate lubrication model of contaminated coating flows. Computer Physics Communications, 142(1–3):330–332, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003629.

Souaille:2001:EWH

[SR01b] Marc Souaille and Benoît Roux. Extension to the weighted histogram analysis method: combining umbrella sampling with free energy calculations. Computer Physics Communications, 135(1):40–57, March 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500002150.

Shamsi:2005:SHI

[SR05] M. Shamsi and M. Razzaghi. Solution of Hallen’s integral equation using multiwavelets. Computer Physics Communications, 168(3):187–197, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002298.

Salam:2009:HOP

[SR09] G. P. Salam and J. Rojo. A higher order perturbative parton evolution toolkit (HOPPET). Computer Physics Communications, 180(1):120–156, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002853.

Srivastava:2001:GEP

[Sri01] G. P. Srivastava. Geometry, electrons, phonons and reactions on Si(001) surfaces. Computer Physics Communications, 137(1):143–162, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550100176X.

Steinbeck:2000:EGS

[SRR^{+}00] L. Steinbeck, A. Rubio, L. Reining, M. Torrent, I. D. White, and R. W. Godby. Enhancements to the GW space–time method. Computer Physics Communications, 125(1–3):105– 118, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559900466X.

Schlier:2000:HOS

[SS00] Ch. Schlier and A. Seiter. High-order symplectic integration: an assessment. Computer Physics Communications, 130(1–2):176–189, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000114.

Slabospitsky:2002:TSE

[SS02a] S. R. Slabospitsky and L. Sonnenschein. TopReX 3.25 — a specialized event generator for use with PYTHIA. Computer Physics Communications, 148(1):87–102, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550200471X.

Sutmann:2002:CFS

[SS02b] G. Sutmann and B. Steffen. Correction of finite size effects in molecular dynamics applied to the friction coefficient of a Brownian particle. Computer Physics Communications, 147(1–2):374–377, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003089.

Sutmann:2005:PPP

[SS05] Godehard Sutmann and Bernhard Steffen. A particle– particle particle–multigrid method for long-range interactions in molecular simulations. Computer Physics Communications, 169(1–3):343–346, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001748.

Steinigeweg:2006:SIC

[SS06] Robin Steinigeweg and Heinz-Jürgen Schmidt. Symplectic integrators for classical spin systems. Computer Physics Communications, 174(11):853–861, June 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000385.

Sakaniwa:2007:NSC

[SS07a] Yasunori Sakaniwa and Hiroyuki Shima. Numerical study on critical exponents of hyperbolic Ising lattice. Computer Physics Communications, 177(1–2):189, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507000690.

Son:2007:HCK

[SS07b] Dongchul Son and Jun-Suhk Suh. HEP computing in korea. Computer Physics Communications, 177(1–2):243–246, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507000574.

Sahu:2009:FIH

[SS09a] Sridhar Sahu and Alok Shukla. Fortran 90 implementation of the Hartree–Fock approach within the CNDO/2 and INDO models. Computer Physics Communications, 180(5):724–734, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003834.

Stevanato:2009:CSH

[SS09b] Elisa Stevanato and Roberto Stroili. A case study for the handling of data in a skimming control system. Computer Physics Communications, 180(7):1109–1115, July 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000265.

Steinhofel:2007:RTC

[SSA07] K. Steinhöfel, A. Skaliotis, and A. A. Albrecht. Relating time complexity of protein folding simulation to approximations of folding time. Computer Physics Communications, 176(7):465– 470, April 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000264.

Strozzi:2004:SLP

[SSB04] D. J. Strozzi, M. M. Shoucri, and A. Bers. Study of laser plasma interactions using an Eulerian Vlasov code. Computer Physics Communications, 164(1–3):156–159, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002747.

Scott:2009:STD

[SSB^{+}09] N. S. Scott, M. P. Scott, P. G. Burke, T. Stitt, V. Faro-Maza, C. Denis, and A. Maniopoulou. 2DRMP: a suite of twodimensional R-matrix propagation codes. Computer Physics Communications, 180(12):2424–2449, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509002380.

Senda:2001:SPL

[SSH01] Y. Senda, F. Shimojo, and K. Hoshino. A survival of the polyanions in liquid alloys. Computer Physics Communications, 142(1–3):401–405, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003708.

Selke:2002:FSC

[SSH02] W. Selke, F. Szalma, and J. S. Hager. Fluctuations of steps on crystal surfaces. Computer Physics Communications, 147(1–2):251–254, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002576.

Strepp:2002:PTH

[SSLN02] W. Strepp, S. Sengupta, M. Lohrer, and P. Nielaba. Phase transitions of hard and soft disks in external periodic potentials: a Monte Carlo study. Computer Physics Communications, 147(1–2):370–373, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003077.

Singh:2008:EAC

[SSP08a] Vineet K. Singh, Om P. Singh, and Rajesh K. Pandey. Efficient algorithms to compute Hankel transforms using wavelets. Computer Physics Communications, 179(11):812–818, December 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550800249X.

Singh:2008:NEH

[SSP08b] Vineet K. Singh, Om P. Singh, and Rajesh K. Pandey. Numerical evaluation of the Hankel transform by using linear Legendre multi-wavelets. Computer Physics Communications, 179(6):424–429, September 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001434.

Sainidou:2005:LMS

[SSPM05] R. Sainidou, N. Stefanou, I. E. Psarobas, and A. Modinos. A layer-multiple-scattering method for phononic crystals and heterostructures of such. Computer Physics Communications, 166(3):197–240, March 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005600.

Soloviev:2001:IAE

[SSZ01] A. G. Soloviev, A. N. Sosnin, and E. P. Zhidkov. Improving the accuracy of estimation of unknown random variable probability density over empirical data. Computer Physics Communications, 137(3):366–379, July 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001588.

Sabbane:2002:CAC

[ST02] Mohamed Sabbane and Mohamed Tij. Calculation algorithm for the collisional moments of the Boltzmann equation for Maxwell molecules. Computer Physics Communications, 149(1):19–29, November 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502005957.

Smirnov:2009:FIE

[ST09] A. V. Smirnov and M. N. Tentyukov. Feynman Integral Evaluation by a Sector decomposiTion Approach (FI-ESTA). Computer Physics Communications, 180(5):735–746, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003858.

Stauffer:2000:ISM

[Sta00] Dietrich Stauffer. Irreversible statistical mechanics from reversible motion: Q2R automata. Computer Physics Communications, 127(1):113–119, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000205.

Stauffer:2002:SSM

[Sta02] Dietrich Stauffer. Sociophysics: the Sznajd model and its applications. Computer Physics Communications, 146(1):93–98, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004393.

Steinhauser:2001:MPP

[Ste01] Matthias Steinhauser. MATAD: a program package for the computation of MAssive TADpoles. Computer Physics Communications, 134(3):335–364, March 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500002046.

Stevenson:2002:AAM

[Ste02] P. D. Stevenson. Analytic angular momentum coupling coefficient calculators. Computer Physics Communications, 147 (3):853–858, September 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004629.

Stegailov:2005:HHM

[Ste05] Vladimir Stegailov. Homogeneous and heterogeneous mechanisms of superheated solid melting and decay. Computer Physics Communications, 169(1–3):247–250, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001542.

Suetomi:2000:TDF

[STK^{+}00] E. Suetomi, M. Tanaka, S. Kamiya, S. Hayashi, S. Ikeda, H. Sugawara, and Y. Sakai. Two-dimensional fluid simulation of plasma reactors for the immobilization of krypton. Computer Physics Communications, 125(1–3):60–74, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004622.

Streltsov:2000:ACL

[Str00] I. P. Streltsov. Application of Chebyshev and Legendre polynomials on discrete point set to function interpolation and solving Fredholm integral equations. Computer Physics Communications, 126(1–2):178–181, April 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599005202.

Strunin:2001:CAS

[Str01a] Dmitry V. Strunin. Computations of auto-solitary structures modeling extended elementary particles. Computer Physics Communications, 142(1–3):95–99, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003332.

Strunin:2001:TTD

[Str01b] Dmitry V. Strunin. A transfer of a tracer in decaying turbulent layer. Computer Physics Communications, 142(1–3): 333–337, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003630.

Straka:2005:ATA

[Str05] Christian W. Straka. ADF95: Tool for automatic differentiation of a FORTRAN code designed for large numbers of independent variables. Computer Physics Communications, 168(2):123–139, June 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505000810.

Succi:2002:LBS

[Suc02] Sauro Succi. Lattice Boltzmann schemes for quantum applications. Computer Physics Communications, 146(3):317–323, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004228.

Sugawara:2001:NSS

[Sug01] M. Sugawara. Numerical solution of the Schrödinger equation by neural network and genetic algorithm. Computer Physics Communications, 140(3):366–380, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002867.

Sullivan:2005:FEC

[Sul05] Zack Sullivan. Fast evaluation of CTEQ parton distributions in Monte Carlos. Computer Physics Communications, 168(1):25– 31, May 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000809.

Suslov:2001:SCA

[Sus01] Sergey A. Suslov. Searching convective/absolute instability boundary for flows with fully numerical dispersion relation. Computer Physics Communications, 142(1–3):322–325, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003605.

Suzuki:2000:MBC

[Suz00] Masuo Suzuki. Mathematical basis of computational statistical physics and quantum analysis. Computer Physics Communications, 127(1):32–36, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000266.

Silva:2001:PPC

[SV01] A. I. Silva and S. M. Viegas. PopRatio: a program to calculate atomic level populations in astrophysical plasmas. Computer Physics Communications, 136(3):319–333, May 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001497.

Simos:2001:EFH

[SVA01] T. E. Simos and Jesus Vigo-Aguiar. An exponentiallyfitted high order method for long-term integration of periodic initial-value problems. Computer Physics Communications, 140(3):358–365, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002855.

Simos:2003:DEF

[SVA03] T. E. Simos and Jesus Vigo-Aguiar. A dissipative exponentiallyfitted method for the numerical solution of the Schrödinger equation and related problems. Computer Physics Communications, 152(3):274–294, May 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502008718.

Spinnato:2001:PPB

[SvAS01] P. F. Spinnato, G. D. van Albada, and P. M. A. Sloot. Performance prediction of N-body simulations on a hybrid architecture. Computer Physics Communications, 139(1):34– 44, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501002259.

Srinivasan:2000:PFP

[SVMT00] Jay Srinivasan, Yuri L. Volobuev, Steven L. Mielke, and Donald G. Truhlar. Parallel Fourier Path-Integral Monte Carlo calculations of absolute free energies and chemical equilibria. Computer Physics Communications, 128(1–2):446–464, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000527.

Shekhovtsova:2009:FMC

[SVP09] O. Shekhovtsova, G. Venanzoni, and G. Pancheri. FAS-TERD: a Monte Carlo event generator for the study of final state radiation in the process e^{+}e^{−} → ππγ at DAΦNE. Computer Physics Communications, 180(7):1206–1218, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000381.

Simula:2001:QCD

[SVS01] T. P. Simula, S. M. M. Virtanen, and M. M. Salomaa. Quantized circulation in dilute Bose–Einstein condensates. Computer Physics Communications, 142(1–3):396–400, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003691.

Simos:2000:PSH

[SW00a] T. E. Simos and P. S. Williams. A P-stable hybrid exponentially-fitted method for the numerical integration of the Schrödinger equation. Computer Physics Communications, 131(1–2):109–119, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001247.

Skrzypek:2000:HGF

[SW00b] M. Skrzypek and Z. Was. How to generate four-fermion phase space. Computer Physics Communications, 125(1–3):8– 20, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004579.

Shao:2009:APN

[SW09] Hezhu Shao and Zhongcheng Wang. Arbitrarily precise numerical solutions of the one-dimensional Schrödinger equation. Computer Physics Communications, 180(1):1–7, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002725.

Shu:2003:OTP

[SWC^{+}03] Ji Wu Shu, Bing Wang, Min Chen, Jin Zhao Wang, and Wei Min Zheng. Optimization techniques for parallel forcedecomposition algorithm in molecular dynamic simulations. Computer Physics Communications, 154(2):121–130, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550300290X.

Swendsen:2002:CSP

[Swe02] Robert H. Swendsen. Computational statistical physics in the 21st century. Computer Physics Communications, 146(1):135– 136, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004745.

Strandlie:2000:PTF

[SWFL00] A. Strandlie, J. Wroldsen, R. Frühwirth, and B. Lillekjendlie. Particle tracks fitted on the Riemann sphere. Computer Physics Communications, 131(1–2):95–108, September 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000862.

Swift:2004:UHC

[Swi04] Daniel W. Swift. Use of a hybrid code for global-scale simulation of the Earth’s magnetosphere. Computer Physics Communications, 164(1–3):370–376, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003054.

Seaton:2009:WLS

[SWL09] D. T. Seaton, T. Wüst, and D. P. Landau. A Wang– Landau study of the phase transitions in a flexible homopolymer. Computer Physics Communications, 180(4):587–589, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003986.

Sun:2003:EAS

[SWP03] H. M. Sun, J. B. Wang, and J. Pan. An effective algorithm for simulating acoustical wave propagation. Computer Physics Communications, 151(2):241–249, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007002.

Shimobaba:2012:CWO

[SWS^{+}12] Tomoyoshi Shimobaba, Jiantong Weng, Takahiro Sakurai, Naohisa Okada, Takashi Nishitsuji, Naoki Takada, Atsushi Shiraki, Nobuyuki Masuda, and Tomoyoshi Ito. Computational wave optics library for C++: CWO++ library. Computer Physics Communications, 183(5):1124–1138, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465512000082.

Siu:2001:IFP

[SWY01] Y. L. Siu, Jones T. K. Wan, and K. W. Yu. Interparticle force in polydisperse electrorheological fluids: Beyond the dipole approximation. Computer Physics Communications, 142(1–3):446–452, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003861.

Stefanou:2000:MNV

[SYM00] N. Stefanou, V. Yannopapas, and A. Modinos. MULTEM 2: a new version of the program for transmission and bandstructure calculations of photonic crystals. Computer Physics Communications, 132(1–2):189–196, October 15, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001314.

Shima:2001:QTL

[SYN01] Hiroyuki Shima, Kousuke Yakubo, and Tsuneyoshi Nakayama. Quantum transport in long-range random magnetic fields. Computer Physics Communications, 142(1–3):424–428, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003824.

Schmidt:2000:STD

[SŽ00a] Burkhard Schmidt and Petra Žd̆ánská. Solution of the timedependent Schrödinger equation for highly symmetric potentials. Computer Physics Communications, 127(2–3):290–308, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046559900524X.

Sevastianov:2000:APM

[SZ00b] L. A. Sevastianov and E. P. Zhidkov. Analysis of problem in mathematical model for shadowed sputtering. Computer Physics Communications, 130(1–2):47–53, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000370.

Soloveva:2000:IMS

[SZ00c] T. M. Solov’eva and E. P. Zhidkov. Iteration method of solving the integral equation with nonlinear dependence on spectral parameter. Computer Physics Communications, 126(1–2):168– 177, April 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004786.

Subba:2004:MPW

[SZ04] F. Subba and R. Zanino. Modeling plasma-wall interactions in First Wall-Limiter geometry. Computer Physics Communications, 164(1–3):377–382, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003066.

Tafirout:2000:EEH

[TA00a] R. Tafirout and G. Azuelos. Erratum to: “EXOTIC — A heavy fermion and excited fermion Monte Carlo generator for e^{+}e^{−} physics” [Computer Physics Communication 126 (2000) 244– 260]. Computer Physics Communications, 133(1):136, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002137. See [TA00b].

Tafirout:2000:EHF

[TA00b] R. Tafirout and G. Azuelos. EXOTIC — A heavy fermion and excited fermion Monte Carlo generator for e^{+}e^{−} physics. Computer Physics Communications, 126(3):244–260, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004646. See [TA00a].

Takaishi:2000:CIH

[Tak00] Tetsuya Takaishi. Choice of integrator in the hybrid Monte Carlo algorithm. Computer Physics Communications, 133 (1):6–17, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001612.

Takahashi:2003:EIP

[Tak03] Daisuke Takahashi. Efficient implementation of parallel three-dimensional FFT on clusters of PCs. Computer Physics Communications, 152(2):144–150, May 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008184.

Takashima:2002:NAM

[TAKN02] Hajime Takashima, Takashi Amisaki, Kunihiro Kitamura, and Umpei Nagashima. Numerical accuracy on F_{m}(z) for molecular integral calculations. Computer Physics Communications, 148(2):182–187, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502005519.

Talman:2009:NSC

[Tal09] J. D. Talman. NumSBT: a subroutine for calculating spherical Bessel transforms numerically. Computer Physics Communications, 180(2):332–338, February 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003329.

Tamine:2003:MIM

[Tam03] M. Tamine. Mean interface magnon dispersion curves obtained by means of the matching procedure. Computer Physics Communications, 155(3):209–220, November 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003667.

Taguchi:2004:SHE

[TAM04] T. Taguchi, T. M. Antonsen, Jr., and K. Mima. Study of hot electron beam transport in high density plasma using 3D hybrid-Darwin code. Computer Physics Communications, 164(1–3):269–278, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002917.

Tablero:2001:GFI

[TAP01] César Tablero, Alfredo Aguado, and Miguel Paniagua. Global fit of ab initio potential energy surfaces: II.2. Tetratomic systems A_{2}B_{2} and ABC_{2}. Computer Physics Communications, 140(3):412–417, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002892.

Tatekawa:2007:PTD

[Tat07] Takayuki Tatekawa. Phase transition in d-dimensional longrange interacting systems. Computer Physics Communications, 177(1–2):190, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465507000707.

Tallinen:2009:DES

[TÅT09] T. Tallinen, J. A. Åström, and J. Timonen. Discrete element simulations of crumpling of thin sheets. Computer Physics Communications, 180(4):512–516, April 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004554.

Thompson:1985:CCF

[TB85] I. J. Thompson and A. R. Barnett. COULCC: a continuedfraction algorithm for Coulomb functions of complex order with complex arguments. Computer Physics Communications, 36(4):363–372, June 1985. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0010465585900256.

See [Tho04b].

Thompson:1987:MBF

[TB87] I. J. Thompson and A. R. Barnett. Modified Bessel functions I_{ν}(z) and K_{ν}(z) of real order and complex argument, to selected accuracy. Computer Physics Communications, 47(2–3):245–257, November/December 1987. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0010465587901111. See erratum [Tho04a].

Tsai:2002:DCB

[TBL02] Shan-Ho Tsai, Alex Bunker, and D. P. Landau. Dynamic critical behavior of the classical anisotropic BCC Heisenberg antiferromagnet. Computer Physics Communications, 147(1–2):97–100, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002242.

Tsuzuki:2007:SCD

[TBR07] Helio Tsuzuki, Paulo S. Branicio, and José P. Rino. Structural characterization of deformed crystals by analysis of common atomic neighborhood. Computer Physics Communications, 177(6):518–523, September 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002536.

Tarantola:2012:CTS

[TBZ12] Stefano Tarantola, William Becker, and Dirk Zeitz. A comparison of two sampling methods for global sensitivity analysis. Computer Physics Communications, 183(5):1061–1072, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465511004036.

Thacker:2006:PAC

[TC06] Robert J. Thacker and H. M. P. Couchman. A parallel adaptive P^{3}M code with hierarchical particle reordering. Computer Physics Communications, 174(7):540–554, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505006090.

Turner:2007:MDF

[TC07] M. M. Turner and P. Chabert. Modelling of dual-frequency capacitive discharges. Computer Physics Communications, 177(1–2):88–92, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550700121X.

Teixeira:2000:MST

[TCF00] Chris Teixeira, Hudong Chen, and David M. Freed. Multispeed thermal lattice Boltzmann method stabilization via equilibrium under-relaxation. Computer Physics Communications, 129(1–3):207–226, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001089.

Tsunoda:2000:NST

[TCO00] Shingo Tsunoda, Yu Chen, and Hirotada Ohashi. A new surface tracking algorithm for lattice gas automata. Computer Physics Communications, 129(1–3):138–144, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001016.

Tao:2008:SAP

[TCY^{+}08] WeiFeng Tao, DongSheng Cai, XiaoYang Yan, Nishikawa Kenichi, and Bertrand Lembege. Scalability analysis of parallel Particle-In-Cell codes on computational grids. Computer Physics Communications, 179(12):855–864, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002518.

Toffoli:2003:LSB

[TD03] D. Toffoli and P. Decleva. Least squares B-spline solutions of the radial Dirac equation in the continuum. Computer Physics Communications, 152(2):151–164, May 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008202.

Tonge:2004:TDS

[TDD04] J. Tonge, D. E. Dauger, and V. K. Decyk. Two-dimensional semiclassical Particle-In-Cell code for simulation of quantum plasmas. Computer Physics Communications, 164(1–3): 279–285, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465504002929.

Tsutahara:2000:SSF

[TdFK00] Michihisa Tsutahara, Shi de Feng, and Takeshi Kataoka. Simulation of the stratified flows using the two-component lattice Boltzmann method. Computer Physics Communications, 129(1–3):131–137, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001004.

Timbo:2009:CCC

[TdRGD09] C. Timbó, L. A. R. da Rosa, M. Gonçalves, and S. B. Duarte. Computational cancer cells identification by fractal dimension analysis. Computer Physics Communications, 180(6):850– 853, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004190.

Toth:2002:MNL

[TDY02] Géza Tóth, Colin Denniston, and J. M. Yeomans. Modeling nematohydrodynamics in liquid crystal devices. Computer Physics Communications, 147(1–2):7–12, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502001960.

Thieulot:2005:NID

[TE05] Cedric Thieulot and Pep Español. Non-isothermal diffusion in a binary mixture with smoothed particle hydrodynamics. Computer Physics Communications, 169(1–3):172–176, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001360.

Tehrani:2001:BPR

[Teh01] Francesco Safai Tehrani. The BaBar Prompt Reconstruction manager: a real life example of a constructive approach to software development. Computer Physics Communications, 140(1–2):56–64, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501002557.

Tilson:2000:PSO

[TEP00] J. L. Tilson, W. C. Ermler, and R. M. Pitzer. Parallel spin-orbit coupled configuration interaction. Computer Physics Communications, 128(1–2):128–138, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000618.

Tentyukov:2000:FDA

[TF00] M. Tentyukov and J. Fleischer. A Feynman diagram analyzer DIANA. Computer Physics Communications, 132(1– 2):124–141, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500001478.

Tentyukov:2004:PCF

[TF04] M. Tentyukov and J. Fleischer. Parallel computation of Feynman diagrams with DIANA. Computer Physics Communications, 160(3):167–186, July 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504002048.

Tartakovsky:2009:LPM

[TFM09] Alexandre M. Tartakovsky, Kim F. Ferris, and Paul Meakin. Lagrangian particle model for multiphase flows. Computer Physics Communications, 180(10):1874–1881, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001829.

Teuler:2000:DPI

[TG00] J. M. Teuler and J. C. Gimel. A direct parallel implementation of the Hoshen–Kopelman algorithm for distributed memory architectures. Computer Physics Communications, 130(1–2):118– 129, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000461.

Trail:2001:EDM

[TGB01] J. R. Trail, M. C. Graham, and D. M. Bird. Electronic damping of molecular motion at metal surfaces. Computer Physics Communications, 137(1):163–173, June 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001771.

Tsoulos:2006:GTF

[TGD06] Ioannis G. Tsoulos, Dimitris Gavrilis, and Evangelos Dermatas. GDF: a tool for function estimation through grammatical evolution. Computer Physics Communications, 174(7):555–559, April 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505005874.

Tsoulos:2007:GVE

[TGD07] Ioannis G. Tsoulos, Dimitris Gavrilis, and Evangelos Dermatas. GDF v2.0, an enhanced version of GDF. Computer Physics Communications, 177(12):976–977, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507003761.

Torras:2007:PSA

[THC^{+}07] Juan Torras, Yao He, Chao Cao, Krishna Muralidharan, E. Deumens, H.-P. Cheng, and S. B. Trickey. PUPIL: a systematic approach to software integration in multi-scale simulations. Computer Physics Communications, 177(3):265–279, August 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001828.

Theodorou:2005:HMA

[The05] Doros N. Theodorou. Hierarchical modeling of amorphous polymers. Computer Physics Communications, 169(1–3):82– 88, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001177.

Tackett:2001:PAW

[THM01] A. R. Tackett, N. A. W. Holzwarth, and G. E. Matthews. A Projector Augmented Wave (PAW) code for electronic structure calculations, Part II: pwpaw for periodic solids in a plane wave basis. Computer Physics Communications, 135(3):348–376, April 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500002411.

Thompson:2001:CSF

[Tho01] William J. Thompson. Computing special functions of mathematical physics. Computer Physics Communications, 142 (1–3):1–6, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003010.

Thompson:2004:EBB

[Tho04a] I. J. Thompson. Erratum to Modified Bessel functions I nu(z) and K_{ν}(z) of real order and complex argument [Comput. Phys. Commun. 47 (1987) 245–257]. Computer Physics Communications, 159(3):243–244, June 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001067. See [TB87].

Thompson:2004:ECC

[Tho04b] I. J. Thompson. Erratum to “COULCC: a continued-fraction algorithm for Coulomb functions of complex order with complex arguments”: [Comput. Phys. Commun. 36 (1985) 363– 372]. Computer Physics Communications, 159(3):241–242, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001079. See [TB85].

Tobimatsu:2001:PFR

[TI01] Keijiro Tobimatsu and Masataka Igarashi. Precise formula for the radiative Bhabha scattering. Computer Physics Communications, 136(1–2):105–116, May 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001473.

Tordella:2007:SSL

[TIM07] Daniela Tordella, Michele Iovieno, and Silvano Massaglia. Small scale localization in turbulent flows. A priori tests applied to a possible Large Eddy Simulation of compressible turbulent flows. Computer Physics Communications, 176(8):539–549, April 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000288. See erra-

tum [TIM08].

Tordella:2008:EBS

[TIM08] D. Tordella, M. Iovieno, and S. Massaglia. Erratum to: Small scale localization in turbulent flows. A priori tests applied to a possible Large Eddy Simulation of compressible turbulent flows [Comput. Phys. Comm. 176 (2007) 539–549]. Computer Physics Communications, 178(11):883–884, June 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000672. See [TIM07].

Tanabe:2009:SHG

[TIN^{+}09] Noriyuki Tanabe, Yasuyuki Ichihashi, Hirotaka Nakayama, Nobuyuki Masuda, and Tomoyoshi Ito. Speed-up of hologram generation using ClearSpeed Accelerator board. Computer Physics Communications, 180(10):1870–1873, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001805.

Tasaki:2001:PRP

[TiTD01] Hiroyasu Tasaki, Jun ichi Takimoto, and Masao Doi. Prediction of the rheological properties of polymers using a stochastic simulation. Computer Physics Communications, 142(1–3):136–139, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501003526.

Tabakin:2009:QPE

[TJD09] Frank Tabakin and Bruno Juliá-Díaz. QCMPI: a parallel environment for quantum computing. Computer Physics Communications, 180(6):948–964, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508004141.

Torrieri:2006:SFC

[TJLR06] G. Torrieri, S. Jeon, J. Letessier, and J. Rafelski. SHAREv2: fluctuations and a comprehensive treatment of decay feeddown. Computer Physics Communications, 175(10):635– 649, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506002918.

Trasca:2008:TON

[TK08] R. Andreea Trasca and Sabine H. L. Klapp. Translational order in a nanoconfined dipolar fluid. Computer Physics Communications, 179(1–3):66–70, July 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508000179.

Timoshenko:2009:WDA

[TK09] J. Timoshenko and A. Kuzmin. Wavelet data analysis of EX-AFS spectra. Computer Physics Communications, 180(6):920– 925, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004372.

Tennyson:2004:DPS

[TKB^{+}04] Jonathan Tennyson, Maxim A. Kostin, Paolo Barletta, Gregory J. Harris, Oleg L. Polyansky, Jayesh Ramanlal, and Nikolai F. Zobov. DVR3D: a program suite for the calculation of rotation-vibration spectra of triatomic molecules. Computer Physics Communications, 163(2):85–116, November 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003741.

Tsuno:2006:GEG

[TKK^{+}06] S. Tsuno, T. Kaneko, Y. Kurihara, S. Odaka, and K. Kato. GR@PPA 2.7 event generator for pp/pp̄ collisions. Computer Physics Communications, 175(10):665–677, November 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550600292X.

Tran:2008:FCO

[TKN^{+}08] Fabien Tran, Jan Kunes, Pavel Novák, Peter Blaha, Laurence D. Marks, and Karlheinz Schwarz. Force calculation for orbital-dependent potentials with FP-(L)APW + lo basis sets. Computer Physics Communications, 179(11):784–790, December 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550800235X.

Tokar:2006:NST

[TKP06] Mikhail Z. Tokar, Denis Kalupin, and Denis Pilipenko. Numerical solution of transport equations for plasmas with transport barriers. Computer Physics Communications, 175(1):30– 35, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600107X.

Tamura:2001:VRS

[TKS^{+}01] Yuichi Tamura, Akira Kageyama, Tetsuya Sato, Susumu Fujiwara, and Hiroaki Nakamura. Virtual reality system to visualize and auralize numerical simulation data. Computer Physics Communications, 142(1–3):227–230, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003137.

Tolke:2000:DBE

[TKSR00] Jonas Tölke, Manfred Krafczyk, Manuel Schulz, and Ernst Rank. Discretization of the Boltzmann equation in velocity space using a Galerkin approach. Computer Physics Communications, 129(1–3):91–99, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000965.

Tsoulos:2006:GGM

[TL06a] Ioannis G. Tsoulos and Isaac E. Lagaris. GenAnneal: Genetically modified Simulated Annealing. Computer Physics Communications, 174(10):846–851, May 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000142.

Tsoulos:2006:GCR

[TL06b] Ioannis G. Tsoulos and Isaac E. Lagaris. Genetically controlled random search: a global optimization method for continuous multidimensional functions. Computer Physics Communications, 174(2):152–159, January 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550500514X.

Tsoulos:2006:MLA

[TL06c] Ioannis G. Tsoulos and Isaac E. Lagaris. MinFinder: Locating all the local minima of a function. Computer Physics Communications, 174(2):166–179, January 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005680.

Tsoulos:2008:GEG

[TL08a] Ioannis G. Tsoulos and I. E. Lagaris. GenMin: An enhanced genetic algorithm for global optimization. Computer Physics Communications, 178(11):843–851, June 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550800043X.

Tsoulos:2008:MVI

[TL08b] Ioannis G. Tsoulos and Isaac E. Lagaris. Minfinder v2.0: An improved version of minfinder. Computer Physics Communications, 179(8):614–615, October 15, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508001781.

Tsai:2009:PDT

[TL09] Shan-Ho Tsai and D. P. Landau. Phase diagram of a twodimensional large-Q Potts model in an external field. Computer Physics Communications, 180(4):485–487, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003950.

Taccogna:2004:SPT

[TLCS04] Francesco Taccogna, Savino Longo, Mario Capitelli, and Ralf Schneider. Stationary plasma thruster simulation. Computer Physics Communications, 164(1–3):160–170, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002759.

Toeda:2003:CDS

[TLDM03] T. Toeda, M. Lamanna, V. Duic, and A. Manara. Conditions database system of the COMPASS experiment. Computer Physics Communications, 152(2):135–143, May 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008172.

Theos:2004:PSP

[TLP04] F. V. Theos, I. E. Lagaris, and D. G. Papageorgiou. PAN-MIN: sequential and parallel global optimization procedures with a variety of options for the local search strategy. Computer Physics Communications, 159(1):63–69, May 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504000773.

Tashiro:2001:IPS

[TMN01] Junichi Tashiro, Tetsuya Morishita, and Shuichi Nosé. Influence of k-point sampling on stability of structures obtained by ab initio MD simulations. Computer Physics Communications, 142(1–3):281–284, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003460.

Takada:2000:NST

[TMTF00] Naoki Takada, Masaki Misawa, Akio Tomiyama, and Shinya Fujiwara. Numerical simulation of two- and three-dimensional two-phase fluid motion by lattice Boltzmann method. Computer Physics Communications, 129(1–3):233–246, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001107.

Tolic-Norrelykke:2004:MPP

[TNBSF04] Iva Marija Tolić-Nørrelykke, Kirstine Berg-Sørensen, and Henrik Flyvbjerg. Matlab program for precision calibration of optical tweezers. Computer Physics Communications, 159(3):225– 240, June 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001043.

Tong:2000:MCC

[TNCG00] Longyu Tong, Michael Nolan, Tongwei Cheng, and J. C. Greer. A Monte Carlo configuration generation computer program for the calculation of electronic states of atoms, molecules, and quantum dots. Computer Physics Communications, 131(1–2):142–163, September 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001193.

Thompson:2004:FTT

[TND04] I. J. Thompson, F. M. Nunes, and B. V. Danilin. FaCE: a tool for three body Faddeev calculations with core excitation. Computer Physics Communications, 161(1–2):87–107, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002140. See erra-

tum [TND05].

Thompson:2005:EFT

[TND05] I. J. Thompson, F. M. Nunes, and B. V. Danilin. Erratum to “FaCE: a tool for three body Faddeev calculations with core excitation” [Comput. Phys. Commun. 161 (2004) 87–107]. Computer Physics Communications, 170(3):296–297, August 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505003371. See [TND04].

Taniguchi:2007:HCP

[TNI^{+}07] S. Taniguchi, M. Nakamura, Masaharu Isobe, N. Zhao, and M. Sugiyama. Heat conduction problem in a one-dimensional hard-point gas: Molecular dynamics and extended thermodynamics. Computer Physics Communications, 177(1–2):164– 165, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000835.

Tanaka:2000:CEC

[TNY00] R. Tanaka, T. Nakamura, and T. Yabe. Constructing exactly conservative scheme in a non-conservative form. Computer Physics Communications, 126(3):232–243, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465599004737.

Todd:2001:CSS

[Tod01] B. D. Todd. Computer simulation of simple and complex atomistic fluids by nonequilibrium molecular dynamics techniques. Computer Physics Communications, 142(1–3):14–21, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003046.

Toldra:2002:CCS

[Tol02] Ramon Toldrà. A C++ code to solve the DGLAP equations applied to Ultra High Energy Cosmic Rays. Computer Physics Communications, 143(3):287–304, March 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004581.

Tomasik:2009:DMC

[Tom09] Boris Tomásik. DRAGON: Monte Carlo generator of particle production from a fragmented fireball in ultrarelativistic nuclear collisions. Computer Physics Communications, 180(9):1642–1653, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000745.

Tonzani:2007:FFE

[Ton07] Stefano Tonzani. FERM3D: a finite element R-matrix electron molecule scattering code. Computer Physics Communications, 176(2):146–156, January 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465506003225.

Torok:2000:PTA

[Tör00] Csaba Török. 4-point transforms and approximation. Computer Physics Communications, 125(1–3):154–166, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046559900483X.

Tosiek:2008:FPM

[Tos08] Jaromir Tosiek. The Fedosov ∗-product in Mathematica. Computer Physics Communications, 179(12):924–930, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508002671.

Toth:2006:FER

[Tót06] Gábor Tóth. Flexible, efficient and robust algorithm for parallel execution and coupling of components in a framework. Computer Physics Communications, 174(10):793–802, May 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600018X.

Toth:2008:QVP

[Tót08] Géza Tóth. QUBIT4MATLAB V3.0: a program package for quantum information science and quantum optics for Matlab. Computer Physics Communications, 179(6):430–437, September 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001367.

Tews:2001:SRC

[TP01] Michael G. Tews and Warren F. Perger. Standalone relativistic continuum wavefunction solver. Computer Physics Communications, 141(1):205–216, November 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002089.

Theory:2004:DPW

[TPBE04] Plasma Theory, Simulation Group (PTSG), Charles K. (Ned) Birdsall, and Edison T. Estacio. Discoveries in plasmas while teaching simulation. Computer Physics Communications, 164(1–3):189–194, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002796.

Thibert-Plante:2003:SSA

[TPYV03] X. Thibert-Plante, D. A. Yuen, and A. P. Vincent. A simple spectral algorithm for solving large-scale Poisson equation in 2D. Computer Physics Communications, 154(2):89–97, August 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503002832.

Tian:2003:ESS

[TQ03]  Yi-Min Tian and Meng-Zhao Qin.  Explicit symplec-

tic schemes for investigating the evolution of vortices in a rotating Bose–Einstein condensate. Computer Physics Communications, 155(2):132–143, October 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503003527.

Tian:2008:MNM

[TQZM08] YiMin Tian, MengZhao Qin, YongMing Zhang, and Tao Ma. The multisymplectic numerical method for Gross–Pitaevskii equation. Computer Physics Communications, 178(6):449–458, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004638.

Tome:2009:SNP

[TRAdO09] Tânia Tomé, Áttila L. Rodrigues, Everaldo Arashiro, and Mário J. de Oliveira. The stochastic nature of predator-prey cycles. Computer Physics Communications, 180(4):536–539, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509000198.

Tabik:2008:MTD

[TRGR08] S. Tabik, L. F. Romero, E. M. Garzón, and J. I. Ramos. On a model of three-dimensional bursting and its parallel implementation. Computer Physics Communications, 178(7):471–485, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004626.

Tripiccione:2001:LSA

[Tri01] R. Tripiccione. LGT simulations on APE machines. Computer Physics Communications, 139(1):55–63, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002302.

Tripiccione:2005:SDC

[Tri05] Raffaele Tripiccione. Strategies for dedicated computing for lattice gauge theories. Computer Physics Communications, 169(1–3):442–448, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001955.

Troster:2008:MCS

[Trö08] Andreas Tröster. Monte Carlo simulation in Fourier space. Computer Physics Communications, 179(1–3):30–33, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000118.

Tiwari:2006:BSB

[TS06] Rakesh Prabhat Tiwari and Alok Shukla. A basis-set based Fortran program to solve the Gross–Pitaevskii equation for dilute Bose gases in harmonic and anharmonic traps. Computer Physics Communications, 174(12):966–982, June 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506000993.

Teodorescu:2008:HEP

[TS08] Liliana Teodorescu and Daniel Sherwood. High energy physics event selection with gene expression programming. Computer Physics Communications, 178(6):409–419, March 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507004481.

Tsai:2002:FPM

[Tsa02] M.-H. Tsai. First-principle molecular-dynamics simulations of the sticking of Ga and N gas-phase atoms on wurtzite GaN surfaces. Computer Physics Communications, 147(1–2):130–133, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200231X.

Tsuno:2003:GFB

[TSA^{+}03] S. Tsuno, K. Sato, T. Abe, J. Fujimoto, T. Ishikawa, Y. Kurihara, S. Odaka, and Y. Takaiwa. GR@PPA 4b: a four bottom quark production event generator for pp/ppν collisions. Computer Physics Communications, 151(2):216–240, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502007932.

Torrieri:2005:SSH

[TSB^{+}05] G. Torrieri, S. Steinke, W. Broniowski, W. Florkowski, J. Letessier, and J. Rafelski. SHARE: Statistical hadronization with resonances. Computer Physics Communications, 167(3):229–251, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505000755.

Tanaka:2002:RRD

[TSI02] Toshijiro Tanaka, Junko Shibata, and Masayoshi Inoue. On reevaluation rate in discrete time Hogg–Huberman model. Computer Physics Communications, 146(1):118–121, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004435.

Taioli:2006:WWP

[TT06] Simone Taioli and Jonathan Tennyson. WATERWAVES: wave particles dynamics on a complex triatomic potential. Computer Physics Communications, 175(1):41–51, July 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001032.

Tentyukov:2007:EFS

[TV07] M. Tentyukov and J. A. M. Vermaseren. Extension of the functionality of the symbolic program FORM by external software. Computer Physics Communications, 176(6):385–405, March 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465506004152.

Tong:2009:RUH

[TWY09] Xiaojun Tong, Zhu Wang, and Haining Yu. A research using hybrid RBF/Elman neural networks for intrusion detection system secure model. Computer Physics Communications, 180(10):1795–1801, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001519.

Tomiya:2001:NAL

[TY01] Mitsuyoshi Tomiya and Naotaka Yoshinaga. Numerical analysis of level statistical properties of two- and threedimensional coupled quartic oscillators. Computer Physics Communications, 142(1–3):82–87, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003289.

Takizawa:2002:MDS

[TYN02] K. Takizawa, T. Yabe, and T. Nakamura. Multi-dimensional semi-Lagrangian scheme that guarantees exact conservation. Computer Physics Communications, 148(2):137–159, October 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004721.

Tuckerman:2000:EML

[TYS^{+}00] Mark E. Tuckerman, D. A. Yarne, Shane O. Samuelson, Adam L. Hughes, and Glenn J. Martyna. Exploiting multiple levels of parallelism in molecular dynamics based calculations via modern techniques and software paradigms on distributed memory computers. Computer Physics Communications, 128(1–2):333–376, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500000771.

Tomiya:2005:LNH

[TYSH05] Mitsuyoshi Tomiya, Naotaka Yoshinaga, Shoichi Sakamoto, and Aki Hirai. A large number of higher-energy eigenvalues of a huge dimensional matrix for a quantum chaotic study of a quartic potential. Computer Physics Communications, 169(1–3):313–316, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001670.

Tang:2006:LST

[TZZ06] Ke Tang, Huaiwu Zhang, and Zhiyong Zhong. Lekner summation for two-dimensional magnetic dipolar interaction energy. Computer Physics Communications, 174(4):270–272, February 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505005916.

Ukolov:2005:RPN

[UCG^{+}05] Yu. A. Ukolov, N. A. Chekanov, A. A. Gusev, V. A. Rostovtsev, S. I. Vinitsky, and Y. Uwano. A REDUCE program for the normalization of polynomial Hamiltonians. Computer Physics Communications, 166(1):66–80, February 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504005284.

Uhlherr:2003:PMC

[Uhl03] Alfred Uhlherr. Parallel Monte Carlo simulations by asynchronous domain decomposition. Computer Physics Communications, 155(1):31–41, September 1–15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002893.

Urbach:2006:HAM

[UJSW06] C. Urbach, K. Jansen, A. Shindler, and U. Wenger. HMC algorithm with multiple time scale integration and mass preconditioning. Computer Physics Communications, 174(2):87–98, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004753.

Utsumi:2002:ANM

[UK02a] Takayuki Utsumi and James Koga. Accurate numerical method for the solutions of the Schrödinger equation and the radial integrals based on the CIP method. Computer Physics Communications, 148(3):267–280, November 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200557X.

Utsumi:2002:NNM

[UK02b] Takayuki Utsumi and James Koga. New numerical method for the solutions of the MCDF equations based on the CIP method. Computer Physics Communications, 148(3):281–290, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502005970.

Uhlherr:2002:LSA

[ULA^{+}02] Alfred Uhlherr, Stephen J. Leak, Nadia E. Adam, Per E. Nyberg, Manolis Doxastakis, Vlasis G. Mavrantzas, and Doros N. Theodorou. Large scale atomistic polymer simulations using Monte Carlo methods for parallel vector processors. Computer Physics Communications, 144(1):1–22, March 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501004647.

Umeda:2012:NOC

[UNK12] Takayuki Umeda, Yasuhiro Nariyuki, and Daichi Kariya. A non-oscillatory and conservative semi-Lagrangian scheme with fourth-degree polynomial interpolation for solving the Vlasov equation. Computer Physics Communications, 183(5):1094– 1100, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465512000264.

Umeda:2001:IMM

[UOM01] Takayuki Umeda, Yoshiharu Omura, and Hiroshi Matsumoto. An improved masking method for absorbing boundaries in electromagnetic particle simulations. Computer Physics Communications, 137(2):286–299, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001825.

Umeda:2003:NCC

[UOTM03] T. Umeda, Y. Omura, T. Tominaga, and H. Matsumoto. A new charge conservation method in electromagnetic particlein-cell simulations. Computer Physics Communications, 156 (1):73–85, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004375.

Ursescu:2005:SAC

[UTKF05] D. Ursescu, M. Tomaselli, T. Kuehl, and S. Fritzsche. Symbolic algorithms for the computation of Moshinsky brackets and nuclear matrix elements. Computer Physics Communications, 173(3):140–161, December 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505004273.

Umeda:2009:TDF

[UTO09] Takayuki Umeda, Kentaro Togano, and Tatsuki Ogino. Twodimensional full-electromagnetic Vlasov code with conservative scheme and its application to magnetic reconnection. Computer Physics Communications, 180(3):365–374, March 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003767.

Urbina-Villalba:2009:CSR

[UVLRRC09] German Urbina-Villalba, Aileen Lozsán, Kareem Rahn, and Manuel S. Romero-Cano. Calculation of the stability ratio of suspensions using Emulsion Stability Simulations. Computer Physics Communications, 180(11):2129–2139, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900201X.

Umansky:2009:SVE

[UXD^{+}09] M. V. Umansky, X. Q. Xu, B. Dudson, L. L. LoDestro, and J. R. Myra. Status and verification of edge plasma turbulence code BOUT. Computer Physics Communications, 180(6):887– 903, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508004293.

Utsumi:2004:ABS

[UYK^{+}04] Takayuki Utsumi, Takashi Yabe, James Koga, Takayuki Aoki, and Masatoshi Sekine. Accurate basis set by the CIP method for the solutions of the Schrödinger equation. Computer Physics Communications, 157(2):121–138, February 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550300496X.

Vay:2004:APA

[VAH04] J.-L. Vay, J.-C. Adam, and A. Héron. Asymmetric PML for the absorption of waves. Application to mesh refinement in electromagnetic Particle-In-Cell plasma simulations. Computer Physics Communications, 164(1–3):171–177, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002760.

Vakhidov:2000:CPA

[Vak00] Akmal A. Vakhidov. Construction of polynomial approximation for Hansen coefficients. Computer Physics Communications, 124(1):40–48, January 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004312.

Valuev:2005:RPM

[Val05] Ilya Valuev. Reactive potentials for molecular dynamics with cluster multi-range interpolation. Computer Physics Communications, 169(1–3):60–63, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550500113X.

Vigo-Aguiar:2008:EFB

[VAMVR08] J. Vigo-Aguiar, J. Martín-Vaquero, and H. Ramos. Exponential fitting BDF–Runge–Kutta algorithms. Computer Physics Communications, 178(1):15–34, January 1, 2008. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507003669.

VandeVyver:2005:CSS

[Van05a] Hans Van de Vyver. Comparison of some special optimized fourth-order Runge–Kutta methods for the numerical solution of the Schrödinger equation. Computer Physics Communications, 166(2):109–122, March 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005272.

VandeVyver:2005:RKN

[Van05b] Hans Van de Vyver. A Runge–Kutta–Nyström pair for the numerical integration of perturbed oscillators. Computer Physics Communications, 167(2):129–142, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000664.

VandeVyver:2005:SPL

[Van05c] Hans Van de Vyver. Stability and phase-lag analysis of explicit Runge–Kutta methods with variable coefficients for oscillatory problems. Computer Physics Communications, 173(3):115–130, December 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505004327.

VandeVyver:2006:FOS

[Van06] Hans Van de Vyver. A fourth-order symplectic exponentially fitted integrator. Computer Physics Communications, 174(4):255–262, February 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005795.

Varnik:2002:SPS

[Var02] Fathollah Varnik. Simulation of planar systems at constant normal pressure: Is it also possible to keep the plate separation constant? Computer Physics Communications, 149(2):61–70, December 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502006173.

Varga:2008:SFB

[Var08] Kálmán Varga. Solution of few-body problems with the stochastic variational method II: Two-dimensional systems. Computer Physics Communications, 179(8):591–596, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001926.

Vu:2007:SSP

[VBC07] K. T. Vu, J. Butcher, and J. Carminati. Similarity solutions of partial differential equations using DESOLV. Computer Physics Communications, 176(11–12):682–693, June 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002019.

Vlad:2001:GFS

[VBFD01] G. Vlad, S. Briguglio, G. Fogaccia, and B. Di Martino. Gridless finite-size-particle plasma simulation. Computer Physics Communications, 134(1):58–77, February 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001910.

VanDyck:2005:YAD

[VBFM05] D. Van Dyck, G. Brinkmann, V. Fack, and B. D. McKay. To be or not to be Yutsis: Algorithms for the decision problem. Computer Physics Communications, 173(1–2):61–70, December 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505004303.

VanDeun:2008:IPB

[VC08] Joris Van Deun and Ronald Cools. Integrating products of Bessel functions with an additional exponential or rational factor. Computer Physics Communications, 178(8):578–590, April 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004705.

Vaiana:2005:AMM

[VCCS05] A. C. Vaiana, Z. Cournia, I. B. Costescu, and J. C. Smith. AFMM: a molecular mechanics force field vibrational parametrization program. Computer Physics Communications, 167(1):34–42, April 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005624.

Vay:2004:IMR

[VCF^{+}04] J.-L. Vay, P. Colella, A. Friedman, D. P. Grote, P. McCorquodale, and D. B. Serafini. Implementations of mesh refinement schemes for Particle-In-Cell plasma simulations. Computer Physics Communications, 164(1–3):297–305, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002954.

vandenBroek:2008:VVI

[vdB08] Korneel van den Broek. Vscape V1.1.0. An interactive tool for metastable vacua. Computer Physics Communications, 178(1):52–72, January 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507003736.

vandenEshof:2002:NMQ

[vdEFL^{+}02] J. van den Eshof, A. Frommer, Th. Lippert, K. Schilling, and H. A. van der Vorst. Numerical methods for the QCDd overlap operator. I. Sign-function and error bounds. Computer Physics Communications, 146(2):203–224, July 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004551.

vanDyk:2009:HCL

[vDGM^{+}09] Danny van Dyk, Markus Geveler, Sven Mallach, Dirk Ribbrock, Dominik Göddeke, and Carsten Gutwenger. HONEI: a collection of libraries for numerical computations targeting multiple processor architectures. Computer Physics Communications, 180(12):2534–2543, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001295.

vanderHorst:2002:ECP

[vdHBP^{+}02] J.-W. van der Horst, P. A. Bobbert, W. F. Pasveer, M. A. J. Michels, G. Brocks, and P. J. Kelly. Excitons in conjugated polymers from first principles. Computer Physics Communications, 147(1–2):331–334, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002977.

vanderHolst:2008:MGA

[vdHKM08] B. van der Holst, R. Keppens, and Z. Meliani. A multidimensional grid-adaptive relativistic magnetofluid code. Computer Physics Communications, 179(9):617–627, November 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508001975.

vanderSman:2008:EDD

[vdSvdG08] R. G. M. van der Sman and S. van der Graaf. Emulsion droplet deformation and breakup with lattice Boltzmann model. Computer Physics Communications, 178(7):492–504, April 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004675.

vanErp:2008:EPS

[vE08] Titus S. van Erp. Efficient path sampling on multiple reaction channels. Computer Physics Communications, 179(1–3): 34–40, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550800012X.

Vegh:2004:SER

[Vég04] János Végh. A simple “embedded” reasoning inference engine with application example in the X-ray photoelectron spectroscopy. Computer Physics Communications, 160(1):8–22, June 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504001080.

VanLier:2008:ADC

[VEG08] G. Van Lier, C. P. Ewels, and P. Geerlings. Automated determination of chemical functionalisation addition routes based on magnetic susceptibility and nucleus independent chemical shifts. Computer Physics Communications, 179(1–3):165– 170, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000350.

Verges:2000:BCB

[Ver00] J. A. Vergés. A brief comparison between two programs that compute the static conductance of a disordered twodimensional tight-binding system. Computer Physics Communications, 127(2–3):268–270, May 10, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599005123.

Verboncoeur:2004:AEF

[Ver04] J. P. Verboncoeur. Aliasing of electromagnetic fields in stair step boundaries. Computer Physics Communications, 164(1–3):344–352, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504003029.

VanDyck:2003:GHB

[VF03a] D. Van Dyck and V. Fack. GYutsis: heuristic based calculation of general recoupling coefficients. Computer Physics Communications, 154(3):219–232, August 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002807.

VanDyck:2003:NHA

[VF03b] D. Van Dyck and V. Fack. New heuristic approach to the calculation of general recoupling coefficients. Computer Physics Communications, 151(3):354–368, April 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502007336.

vonHippel:2006:TAO

[vH06] G. M. von Hippel. TaylUR, an arbitrary-order diagonal automatic differentiation package for Fortran 95. Computer Physics Communications, 174(7):569–576, April 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000166.

vonHippel:2007:NVA

[vH07] G. M. von Hippel. New version announcement for TaylUR, an arbitrary-order diagonal automatic differentiation package for Fortran 95. Computer Physics Communications, 176(11–12):710–711, June 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S001046550700197X.

vanHameren:2000:FAG

[vHK00] André van Hameren and Ronald Kleiss. A fast algorithm for generating a uniform distribution inside a high-dimensional polytope. Computer Physics Communications, 133(1):1–5, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001570.

vonHippel:2008:EFM

[vHLP08] Georg M. von Hippel, Randy Lewis, and Robert G. Petry. Evolutionary fitting methods for the extraction of mass spectra in lattice field theory. Computer Physics Communications, 178(10):713–723, May 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465507005139.

Voglis:2009:NDL

[VHLP09] C. Voglis, P. E. Hadjidoukas, I. E. Lagaris, and D. G. Papageorgiou. A numerical differentiation library exploiting parallel architectures. Computer Physics Communications, 180(8):1404–1415, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509000484.

Viehland:2001:IAI

[Vie01] Larry A. Viehland. Ion-atom interaction potentials and transport properties. Computer Physics Communications, 142 (1–3):7–13, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501003022.

Viola:2004:BRB

[Vio04] Lorenza Viola. Book review: Quantum Computing: Mika Hirvensalo, (2001) Springer-Verlag, Berlin ISBN 3-540-66783-0. Computer Physics Communications, 163(1):65–66, October 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504004114. See [?] and

review [?].

VandenBerghe:2001:OIE

[VIV01] G. Vanden Berghe, L. Gr. Ixaru, and M. Van Daele. Optimal implicit exponentially-fitted Runge–Kutta methods. Computer Physics Communications, 140(3):346–357, November 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100279X.

Vlachos:2009:QGS

[VK09a] Panagiotis Vlachos and Ioannis G. Karafyllidis. Quantum game simulator, using the circuit model of quantum computation. Computer Physics Communications, 180(10):1990–1998, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550900160X.

Vlachos:2009:SQK

[VK09b] Panagiotis Vlachos and Ioannis G. Karafyllidis. Simulation of quantum key expansion using quantum cellular automata. Computer Physics Communications, 180(2):251–255, February 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508003317.

VandeVondele:2005:QFA

[VKM^{+}05] Joost VandeVondele, Matthias Krack, Fawzi Mohamed, Michele Parrinello, Thomas Chassaing, and Jürg Hutter. Quickstep: Fast and accurate density functional calculations using a mixed Gaussian and plane waves approach. Computer Physics Communications, 167(2):103–128, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505000615.

Vashishta:2007:MAS

[VKN07] Priya Vashishta, Rajiv K. Kalia, and Aiichiro Nakano. Multimillion atom simulations of dynamics of wing cracks and nanoscale damage in glass, and hypervelocity impact damage in ceramics. Computer Physics Communications, 177(1–2):202– 205, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000744.

Vanaverbeke:2009:GPS

[VKPB09] S. Vanaverbeke, R. Keppens, S. Poedts, and H. Boffin. GRADSPH: a parallel smoothed particle hydrodynamics code for self-gravitating astrophysical fluid dynamics. Computer Physics Communications, 180(7):1164–1182, July 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000071.

Virnau:2002:PDH

[VMMB02] P. Virnau, M. Müller, L. González MacDowell, and K. Binder. Phase diagrams of hexadecane-CO_{2} mixtures from histogramreweighting Monte Carlo. Computer Physics Communications, 147(1–2):378–381, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502003090.

Vogt:2005:EEU

[Vog05] A. Vogt. Efficient evolution of unpolarized and polarized parton distributions with QCD- pegasus. Computer Physics Communications, 170(1):65–92, July 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002584.

Voigt:2002:QDD

[Voi02] A. Voigt. Quantum disorder due to singlet formation: The plaquette lattice. Computer Physics Communications, 146(1):125– 128, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004459. See adden-

dum [Voi03].

Voigt:2003:AQD

[Voi03] A. Voigt. Addendum to: “Quantum disorder due to singlet formation: The Plaquette lattice”: [Comput. Phys. Commun. 146 (2002) 125–128]. Computer Physics Communications, 154(3):233, August 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503002789.

See [Voi02].

Vormoor:2002:LSF

[Vor02] Oliver Vormoor. Large scale fractal aggregates using the tunable dimension cluster-cluster aggregation. Computer Physics Communications, 144(2):121–129, April 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550200142X.

Voss:2006:NCE

[Vos06] Heinrich Voss. Numerical calculation of the electronic structure for three-dimensional quantum dots. Computer Physics Communications, 174(6):441–446, March 15, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000051.

Vadlamani:2004:PCM

[VPCK04] Srinath Vadlamani, Scott E. Parker, Yang Chen, and Charlson Kim. The particle-continuum method: an algorithmic unification of particle-in-cell and continuum methods. Computer Physics Communications, 164(1–3):209–213, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002826.

Vrahatis:2001:ACB

[VPK^{+}01] M. N. Vrahatis, A. E. Perdiou, V. S. Kalantonis, E. A. Perdios, K. Papadakis, R. Prosmiti, and S. C. Farantos. Application of the characteristic bisection method for locating and computing periodic orbits in molecular systems. Computer Physics Communications, 138(1):53–68, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001904.

Viel:2002:ISR

[VPNW02] Alexandra Viel, Mehul V. Patel, Parhat Niyaz, and K. Birgitta Whaley. Importance sampling in rigid body diffusion Monte Carlo. Computer Physics Communications, 145(1):24– 47, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001455.

Voglis:2012:MGO

[VPP^{+}12] C. Voglis, K. E. Parsopoulos, D. G. Papageorgiou, I. E. Lagaris, and M. N. Vrahatis. MEMPSODE: a global optimization software based on hybridization of populationbased algorithms and local searches. Computer Physics Communications, 183(5):1139–1154, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465512000136.

Virtanen:2001:CMS

[VS01] S. M. M. Virtanen and M. M. Salomaa. Computing microscopic structures of inhomogeneities in superconductors. Computer Physics Communications, 142(1–3):391–395, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550100368X.

Vlachos:2006:PPC

[VS06] D. S. Vlachos and T. E. Simos. PDSW: a program for the calculation of photon energy distribution resulting from radioactive elements in seawater. Computer Physics Communications, 174(5):391–395, March 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005813.

Vu:2000:LFL

[VSBD00] H. X. Vu, K. Y. Sanbonmatsu, B. Bezzerides, and D. F. DuBois. Laying a foundation for laser-plasma modeling for the national ignition facility. Computer Physics Communications, 127(1):71–90, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S001046550000117X.

Vladuca:2000:SCC

[VT00a] G. Vladuca and Anabella Tudora. SPECTRUM — A computer code for prompt fission neutron spectrum and prompt neutron multiplicity calculation. Computer Physics Communications, 125(1–3):221–238, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465599004580.

Volobuev:2000:MSQ

[VT00b] Yuri L. Volobuev and Donald G. Truhlar. An MIMD strategy for quantum mechanical reactive scattering calculations. Computer Physics Communications, 128(1–2):465–476, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000539.

Volobuev:2000:SMQ

[VT00c] Yuri L. Volobuev and Donald G. Truhlar. Stabilization methods for quantum mechanical resonance states of four-body systems. Computer Physics Communications, 128(1–2):516–526, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000606.

Vulcanov:2003:CDE

[Vul03] Dumitru N. Vulcanov. Calculation of the Dirac equation in curved spacetimes with possible torsion using Maple and RE-DUCE. Computer Physics Communications, 154(3):205–218, August 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503003448.

Vollinga:2005:NEM

[VW05] Jens Vollinga and Stefan Weinzierl. Numerical evaluation of multiple polylogarithms. Computer Physics Communications, 167(3):177–194, May 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505000706.

Velichko:2002:MCS

[VYK02] Yuri S. Velichko, Kenichi Yoshikawa, and Alexei R. Khokhlov. Monte Carlo simulation of circular double-stranded polymers. Computer Physics Communications, 146(1):122–124, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004447.

Wolff:2007:EMC

[WA07] Ulli Wolff and ALPHA Collaboration. Erratum to “Monte Carlo errors with less errors” [Comput. Phys. Comm. 156 (2004) 143–153]. Computer Physics Communications, 176 (5):383, March 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465506004322.

See [?].

Waldeer:2003:DSM

[Wal03] K. T. Waldeer. The direct simulation Monte Carlo method applied to a boltzmann-like vehicular traffic flow model. Computer Physics Communications, 156(1):1–12, December 1–15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503003680.

Wang:2000:BRB

[Wan00] Lin-Wang Wang. Book review: Quantum Wells, Wires and Dots: by Paul Harrison; Wiley (UK), 480 pp.; ISBN 0-471-98495-7. Computer Physics Communications, 133(1):137–138, December 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001727.

Wander:2001:NML

[Wan01] A. Wander. A new modular low energy electron diffraction package — DL LEED. Computer Physics Communications, 137(1):4–11, June 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465501001680.

Wang:2005:NEA

[Wan05a] Zhongcheng Wang. A new effective algorithm for the resonant state of a Schrödinger equation. Computer Physics Communications, 167(1):1–6, April 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504005636.

Wang:2005:PSL

[Wan05b] Zhongcheng Wang. P-stable linear symmetric multistep methods for periodic initial-value problems. Computer Physics Communications, 171(3):162–174, October 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002948.

Wang:2006:OOS

[Wan06a] Zhongcheng Wang. Obrechkoff one-step method fitted with Fourier spectrum for undamped Duffing equation. Computer Physics Communications, 175(11–12):692–699, December 1– 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506002955.

Wang:2006:TFMb

[Wan06b] Zhongcheng Wang. Trigonometrically-fitted method for a periodic initial value problem with two frequencies. Computer Physics Communications, 175(4):241–249, August 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465506001317.

Wang:2006:TFMa

[Wan06c] Zhongcheng Wang. Trigonometrically-fitted method with the Fourier frequency spectrum for undamped Duffing equation. Computer Physics Communications, 174(2):109–118, January 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005084.

Wang:2009:MNM

[Wan09a] Jian Wang. Multisymplectic numerical method for the Zakharov system. Computer Physics Communications, 180(7): 1063–1071, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004451.

Wang:2009:DMN

[Wan09b] Zhen Wang. Discrete tanh method for nonlinear differencedifferential equations. Computer Physics Communications, 180(7):1104–1108, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509000253.

Wittmer:2007:PMI

[WBC^{+}07] J. P. Wittmer, P. Beckrich, F. Crevel, C. C. Huang, A. Cavallo, T. Kreer, and H. Meyer. Are polymer melts “ideal”? Computer Physics Communications, 177(1–2):146–149, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507000914.

Wright:2004:URS

[WBDB04] J. C. Wright, P. T. Bonoli, E. D’Azevedo, and M. Brambilla. Ultrahigh resolution simulations of mode converted ion cyclotron waves and lower hybrid waves. Computer Physics Communications, 164(1–3):330–335, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002991.

Watson:2000:LSP

[WC00] Stuart C. Watson and Emily A. Carter. Linear-scaling parallel algorithms for the first principles treatment of metals. Computer Physics Communications, 128(1–2):67–92, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000643.

Wang:2005:TFO

[WC05] Zhongcheng Wang and Qimang Chen. A trigonometricallyfitted one-step method with multi-derivative for the numerical solution to the one-dimensional Schrödinger equation. Computer Physics Communications, 170(1):49–64, July 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505002559.

Woo:2005:EAD

[WCBN05] Teresa K. C. Woo, Vincent K. W. Cheng, Christopher D. Beling, and Michael K. P. Ng. An evaluation of algorithms for the deconvolution of Doppler broadening positron annihilation radiation spectroscopy spectra. Computer Physics Communications, 168(3):177–186, June 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002353.

Wheelock:2004:IIB

[WCG04] Adrian Wheelock, David L. Cooke, and Nikolaos A. Gatsonis. Investigation of ion beam neutralization processes with 2D and 3D PIC simulations. Computer Physics Communications, 164(1–3):336–343, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003005.

Warner:2000:SIE

[WCGL00] Simeon Warner, Simon Catterall, Eric Gregory, and Edward Lipson. SimScience: Interactive educational modules based on large simulations. Computer Physics Communications, 127(1): 1–5, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550000028X.

Wheaton:2009:TTM

[WCH09] S. Wheaton, J. Cleymans, and M. Hauer. THERMUS — a thermal model package for ROOT. Computer Physics Communications, 180(1):84–106, January 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508002750.

Weber:2004:ETE

[WD04] Valéry Weber and Claude Daul. Evaluation of two-electron integrals including the factors r_{12}^{k} exp(−γr_{12}^{2}) over Cartesian Gaussian functions. Computer Physics Communications, 158(1):1–11, March 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004971.

Weber:2004:RNI

[WDB04] Valéry Weber, Claude Daul, and Richard Baltensperger. Radial numerical integrations based on the sinc function. Computer Physics Communications, 163(3):133–142, November 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504004321.

Werth:2002:ACS

[WDF^{+}02] J. H. Werth, S. M. Dammer, Z. Farkas, H. Hinrichsen, and D. E. Wolf. Agglomeration in charged suspensions. Computer Physics Communications, 147(1–2):259–262, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502002850.

Wensch:2004:SSO

[WDHE04] J. Wensch, M. Däne, W. Hergert, and A. Ernst. The solution of stationary ODE problems in quantum mechanics by Magnus methods with stepsize control. Computer Physics Communications, 160(2):129–139, July 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001638.

Wei:2002:SCA

[Wei02a] G. W. Wei. Shock capturing by anisotropic diffusion oscillation reduction. Computer Physics Communications, 144 (3):317–342, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502002734.

Weinzierl:2002:FEP

[Wei02b] Stefan Weinzierl. Fast evolution of parton distributions. Computer Physics Communications, 148(3):314–326, November 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502005842.

Weinzierl:2002:SET

[Wei02c] Stefan Weinzierl. Symbolic expansion of transcendental functions. Computer Physics Communications, 145(3):357–370, June 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002618.

Weinzierl:2004:GFC

[Wei04] Stefan Weinzierl. gTybalt — a free computer algebra system. Computer Physics Communications, 156(2):180–198, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465503004685.

Wellisch:2001:HSM

[Wel01] J. P. Wellisch. Hadronic shower models in Geant 4 — the frameworks. Computer Physics Communications, 140(1–2):65–75, October 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501002569.

Weniger:2001:IID

[Wen01] Ernst Joachim Weniger. Irregular input data in convergence acceleration and summation processes: General considerations and some special Gaussian hypergeometric series as model problems. Computer Physics Communications, 133(2–3):202–228, January 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001752.

Wessel:2007:SAG

[Wes07] Stefan Wessel. Simulations of atomic gases on frustrated optical lattices. Computer Physics Communications, 177(1–2):166– 167, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507000847.

Wilson:2001:CMG

[WG01] K. F. Wilson and D. J. Goossens. A computational model of gastro-intestinal motility. Computer Physics Communications, 142(1–3):105–108, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003393.

Wang:2004:MPT

[WGDZ04] Zhongcheng Wang, Yonghua Ge, Yongming Dai, and Deyin Zhao. A Mathematica program for the two-step twelfthorder method with multi-derivative for the numerical solution of a one-dimensional Schrödinger equation. Computer Physics Communications, 160(1):23–45, June 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001092.

Wang:2006:NWB

[WGL06] Limin Wang, Wei Ge, and Jinghai Li. A new wall boundary condition in particle methods. Computer Physics Communications, 174(5):386–390, March 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005862.

Wagner:2000:SCB

[WGS00] A. J. Wagner, L. Giraud, and C. E. Scott. Simulation of a cusped bubble rising in a viscoelastic fluid with a new numerical method. Computer Physics Communications, 129(1–3):227– 232, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500001090.

Wan:2001:NEA

[WGY01] Jones T. K. Wan, G. Q. Gu, and K. W. Yu. Nonlinear ER effects in an ac applied field. Computer Physics Communications, 142(1–3):457–463, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003903.

Wolf:2000:ASC

[WH00] S. Wolf and Th. Henning. Accelerated self-consistent radiative transfer based on the Monte Carlo method. Computer Physics Communications, 132(1–2):166–188, October 15, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500001442.

Wolf:2005:DDR

[WH05] S. Wolf and L. A. Hillenbrand. Debris disk radiative transfer simulation tool (DDS). Computer Physics Communications, 171(3):208–218, October 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505002936.

Wu:2006:PUM

[WH06] Pei-Chi Wu and Kuo-Chan Huang. Parallel use of multiplicative congruential random number generators. Computer Physics Communications, 175(1):25–29, July 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506001007.

Wang:2007:AOM

[WHCL07] Di-Bao Wang, Fei-Bin Hsiao, Cheng-Hsin Chuang, and Yung-Chun Lee. Algorithm optimization in molecular dynamics simulation. Computer Physics Communications, 177(7):551–559, October 1, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507002482.

White:2000:VFE

[Whi00] Daniel A. White. Vector finite element modeling of optical tweezers. Computer Physics Communications, 128(3):558– 564, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000072.

Wang:2006:SOF

[WHJ06] Weichung Wang, Tsung-Min Hwang, and Jia-Chuan Jang. A second-order finite volume scheme for three dimensional truncated pyramidal quantum dot. Computer Physics Communications, 174(5):371–385, March 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505005849.

Wu:2005:PIM

[WHL05] Jong-Shinn Wu, Yu-Lin Hsu, and Yun-Min Lee. Parallel implementation of molecular dynamics simulation for shortranged interaction. Computer Physics Communications, 170 (2):175–185, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465505002663.

Wu:2007:DPE

[WHL^{+}07] J.-S. Wu, K.-H. Hsu, F.-L. Li, C.-T. Hung, and S.-Y. Jou. Development of a parallelized 3D electrostatic PIC-FEM code and its applications. Computer Physics Communications, 177(1–2): 98–101, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507001208.

Wu:2002:BPR

[WHO02] Pei-Chi Wu, Kuo-Chan Huang, and Shih-Ting Ouyang. Bitparallel random number generation for discrete uniform distributions. Computer Physics Communications, 144(3):252–260, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502001431.

Wilding:2002:NSA

[Wil02] Nigel B. Wilding. A new simulation approach to the freezing transition. Computer Physics Communications, 146(1):99–106, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200440X.

Williams:2009:NOO

[Wil09] M. Williams. Numerical object oriented quantum field theory calculations. Computer Physics Communications, 180 (10):1847–1852, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001659.

Wang:2009:POE

[WJW09] Hong-Yu Wang, Wei Jiang, and You-Nian Wang. Parallelization and optimization of electrostatic Particle-in-Cell/ Monte-Carlo coupled codes as applied to RF discharges. Computer Physics Communications, 180(8):1305–1314, August 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509000666.

Wenckebach:2002:OAI

[WK02] W. Th. Wenckebach and P. Kinsler. An optimized algorithm for ionized impurity scattering in Monte Carlo simulations. Computer Physics Communications, 143(2):136–141, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004404.

Waghmare:2001:HEM

[WKP^{+}01] U. V. Waghmare, Hanchul Kim, I. J. Park, Normand Modine, P. Maragakis, and Efthimios Kaxiras. HARES: an efficient method for first-principles electronic structure calculations of complex systems. Computer Physics Communications, 137(3):341–360, July 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465501001527.

Wang:2000:MCA

[WL00] Jian-Sheng Wang and Lik Wee Lee. Monte Carlo algorithms based on the number of potential moves. Computer Physics Communications, 127(1):131–136, May 1, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000163.

Wust:2008:HMP

[WL08] T. Wüst and D. P. Landau. The HP model of protein folding: a challenging testing ground for Wang–Landau sampling. Computer Physics Communications, 179(1–3):124–127, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508000271.

Wust:2009:MCS

[WLGX09] T. Wüst, D. P. Landau, C. Gervais, and Y. Xu. Monte Carlo simulations of systems with complex energy landscapes. Computer Physics Communications, 180(4):475–479, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508004475.

Wang:2000:ICV

[WLH00] K.-Y. Wang, D. J. Lary, and S. M. Hall. Improvement of a 3D CTM and a 4D variational data assimilation on a vector machine CRAY J90 through a multitasking strategy. Computer Physics Communications, 125(1–3):142–153, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004695.

Wensink:2008:LTS

[WLR^{+}08] H. H. Wensink, H. Löwen, M. Rex, C. N. Likos, and S. van Teeffelen. Long-time self-diffusion for Brownian gaussian-core particles. Computer Physics Communications, 179(1–3):77– 81, July 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000192.

Wong:2004:RCK

[WLW04] W. K. Wong, L. P. Lee, and K. W. Wong. Reply to the comment “Keystream cryptanalysis of a chaotic cryptographic method”. Computer Physics Communications, 156(2):208, January 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465503004338. See [kWpLwW01,

ÁMRP04].

Winstead:2000:PCS

[WM00] Carl Winstead and Vincent McKoy. Parallel computational studies of electron-molecule collisions. Computer Physics Communications, 128(1–2):386–398, June 9, 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500000667.

Wang:2009:PPI

[WMK09] Manhui Wang, Andrew J. May, and Peter J. Knowles. Parallel programming interface for distributed data. Computer Physics Communications, 180(12):2673–2679, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001611.

Weber:2005:TSC

[WML^{+}05] S. Weber, P.-H. Maire, R. Loubère, G. Riazuelo, P. Michel, V. Tikhonchuk, and J. Ovadia. A transport simulation code for inertial confinement fusion relevant laser-plasma interaction. Computer Physics Communications, 168(3):141–158, June 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002316.

Wei:2009:HML

[WMNS09] Tao Wei, Shengjing Mu, Aiichiro Nakano, and Katherine Shing. A hybrid multi-loop genetic-algorithm/simplex/spatial-grid method for locating the optimum orientation of an adsorbed protein on a solid surface. Computer Physics Communications, 180(5):669–674, May 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508003809.

Wada:2001:IRF

[WN01] Tatsuaki Wada and Tomotoshi Nishino. Interaction-round-a-face density-matrix renormalization-group method. Computer Physics Communications, 142(1–3):164–167, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003058.

Wolf:2003:MCR

[Wol03] S. Wolf. MC3D-3D continuum radiative transfer, version 2. Computer Physics Communications, 150(2):99–115, February 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006756.

Ward:2000:PCP

[WP00] A. J. Ward and J. B. Pendry. A program for calculating photonic band structures, Green’s functions and transmission/reflection coefficients using a non-orthogonal FDTD method. Computer Physics Communications, 128(3):590–621, June 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599005433.

Wang:2006:AWP

[WP06] J. B. Wang and J. Pan. Acoustical wave propagator with modified Chebyshev expansion. Computer Physics Communications, 174(3):187–190, February 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505005692.

Wang:2002:FTE

[WPL02] Z. J. Wang, A. J. Przekwas, and Yen Liu. A FV-TD electromagnetic solver using adaptive Cartesian grids. Computer Physics Communications, 148(1):17–29, October 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004642.

Wittkopf:2001:FDE

[WR01] Allan D. Wittkopf and Gregory J. Reid. Fast differential elimination in C: The CDiffElim environment. Computer Physics Communications, 139(2):192–217, September 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501001928.

Welch:2004:INI

[WRC^{+}04] D. R. Welch, D. V. Rose, R. E. Clark, T. C. Genoni, and T. P. Hughes. Implementation of an non-iterative implicit electromagnetic field solver for dense plasma simulation. Computer Physics Communications, 164(1–3):183–188, December 1–15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465504002784.

Winkler:2005:SCF

[WRMG05] R. G. Winkler, M. Ripoll, K. Mussawisade, and G. Gompper. Simulation of complex fluids by multi-particle-collision dynamics. Computer Physics Communications, 169(1–3):326– 330, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001700.

White:2001:CEI

[WRN01] R. D. White, R. E. Robson, and K. F. Ness. Computation of electron and ion transport properties in gases. Computer Physics Communications, 142(1–3):349–355, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003733.

Wrobel:2008:DHR

[Wro08] F. Wrobel. Detailed history of recoiling ions induced by nucleons. Computer Physics Communications, 178(2):88–104, January 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550700375X.

Wilding:2002:WSB

[WS02] N. B. Wilding and F. Schmid. Wetting of a symmetrical binary fluid mixture on a wall. Computer Physics Communications, 147(1–2):149–153, August 1, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502002345.

Wang:2009:NKD

[WS09a] Zhongcheng Wang and Hezhu Shao. A new kind of discretization scheme for solving a two-dimensional time-independent Schrödinger equation. Computer Physics Communications, 180(6):842–849, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003998.

Weyrauch:2009:CBC

[WS09b] Michael Weyrauch and Daniel Scholz. Computing the Baker– Campbell–Hausdorff series and the Zassenhaus product. Computer Physics Communications, 180(9):1558–1565, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001222.

Warrier:2004:SSP

[WSB04] Manoj Warrier, Ralf Schneider, and Xavier Bonnin. Subroutines for some plasma surface interaction processes: physical sputtering, chemical erosion, radiation enhanced sublimation, backscattering and thermal evaporation. Computer Physics Communications, 160(1):46–68, June 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504001110.

Welser-Sherrill:2009:ITD

[WSCW09] L. Welser-Sherrill, J. H. Cooley, and D. C. Wilson. Interactive tools designed to study mix in inertial confinement fusion implosions. Computer Physics Communications, 180(6):835– 841, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508003974.

Watanabe:2001:EAT

[WT01] Naoki Watanabe and Masaru Tsukada. Efficient algorithm for TD–Schrödinger equation and TD–Kohn–Sham equation. Computer Physics Communications, 142(1–3):255–258, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501003356.

Wang:2004:GPS

[WTH^{+}04] W. X. Wang, W. M. Tang, F. L. Hinton, L. E. Zakharov, R. B. White, and J. Manickam. Global δf particle simulation of neoclassical transport and ambipolar electric field in general geometry. Computer Physics Communications, 164(1–3):178–182, December 1–15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002772.

Wu:2004:PTD

[WTW04] Jong-Shinn Wu, Kun-Chang Tseng, and Fu-Yuan Wu. Parallel three-dimensional DSMC method using mesh refinement and variable time-step scheme. Computer Physics Communications, 162(3):166–187, October 1, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504003716.

Wu:2010:EER

[Wu10] Xinyuan Wu. Erratum to “Extended RKN-type methods for numerical integration of perturbed oscillators” [Computer Physics Communications 180 (2009) 1777–1794]. Computer Physics Communications, 181(9):1658, September 2010. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465510001876. See [YWYF09].

Wolf:2004:MSE

[WV04] S. Wolf and N. V. Voshchinnikov. Mie scattering by ensembles of particles with very large size parameters. Computer Physics Communications, 162(2):113–123, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003285.

Willatzen:2005:NIE

[WV05] M. Willatzen and L. C. Lew Yan Voon. Numerical implementation of the ellipsoidal wave equation and application to ellipsoidal quantum dots. Computer Physics Communications, 171(1):1–18, September 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002833.

Wang:2005:NKH

[WW05] Zhongcheng Wang and Yuan Wang. A new kind of highefficient and high-accurate P-stable Obrechkoff three-step method for periodic initial-value problems. Computer Physics Communications, 171(2):79–92, September 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002912.

Wu:2006:MPA

[WW06] Dongmei Wu and Zhongcheng Wang. A Mathematica program for the approximate analytical solution to a nonlinear undamped Duffing equation by a new approximate approach. Computer Physics Communications, 174(6):447–463, March 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505005138.

Wang:2008:IAL

[WWF08] Minghui Wang, Musheng Wei, and Yan Feng. An iterative algorithm for least squares problem in quaternionic quantum theory. Computer Physics Communications, 179(4):203–207, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508000945.

Wu:2009:NDO

[WYL09] Xinyuan Wu, Xiong You, and Jiyong Li. Note on derivation of order conditions for ARKN methods for perturbed oscillators. Computer Physics Communications, 180(9):1545–1549, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001209.

Wu:2009:OCA

[WYX09] Xinyuan Wu, Xiong You, and Jianlin Xia. Order conditions for ARKN methods solving oscillatory systems. Computer Physics Communications, 180(11):2250–2257, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509002318.

Wu:2006:SEH

[WZHZ06] X. Wu, J. F. Zhu, J. Z. He, and H. Zhang. On stabilization of energy for Hamiltonian systems. Computer Physics Communications, 175(1):15–24, July 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506000737.

Xie:2003:AMP

[XC03] Fuding Xie and Yong Chen. An algorithmic method in Painlevé analysis of PDE. Computer Physics Communications, 154(3):197–204, August 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503003436.

Xue:2008:NVI

[XD08] Changfeng Xue and Shaozhong Deng. New versions of image approximations to the ionic solvent induced reaction field. Computer Physics Communications, 178(3):171–185, February 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003943.

Xiao:2001:IMF

[Xia01] Feng Xiao. Implementations of multi-fluid hydrodynamic simulations on distributed memory computer with a fully parallelizable preconditioned Bi-CGSTAB method. Computer Physics Communications, 137(2):274–285, June 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501001813.

Xu:2008:PPS

[XON08] J. Xu, P. N. Ostroumov, and J. Nolen. A parallel 3D Poisson solver for space charge simulation in cylindrical coordinates. Computer Physics Communications, 178(4):290–300, February 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507004249.

Xu:2009:SVS

[XSC09] Dong Xu, Jernej Stare, and Andrew L. Cooksy. Solving the vibrational Schrödinger equation on an arbitrary multidimensional potential energy surface by the finite element method. Computer Physics Communications, 180(11):2079– 2094, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001908.

Xu:2012:ADI

[XZ12] Yiqiang Xu and Luming Zhang. Alternating direction implicit method for solving two-dimensional cubic nonlinear Schrödinger equation. Computer Physics Communications, 183(5):1082–1093, May 2012. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465512000094.

Yakubo:2001:TMA

[Yak01] K. Yakubo. Transfer-matrix approach to electron transport in inhomogeneous magnetic fields. Computer Physics Communications, 142(1–3):429–435, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003836.

Yamamoto:2000:EIM

[Yam00] Susumu Yamamoto. Effective implementations of multidimensional radix-2 FFT. Computer Physics Communications, 125(1–3):1–7, March 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004567.

Yan:2002:EJE

[Yan02] Zhenya Yan. Extended Jacobian elliptic function algorithm with symbolic computation to construct new doublyperiodic solutions of nonlinear differential equations. Computer Physics Communications, 148(1):30–42, October 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502004654.

Yan:2003:MNDa

[Yan03a] Zhenya Yan. Modified nonlinearly dispersive mK(m, n, k) equations: I. New compacton solutions and solitary pattern solutions. Computer Physics Communications, 152(1):25–33, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502007944.

Yan:2003:MNDb

[Yan03b] Zhenya Yan. Modified nonlinearly dispersive mK ( m, n, k ) equations: II. Jacobi elliptic function solutions. Computer Physics Communications, 153(1):1–16, June 1, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502008512.

Yan:2003:NEJ

[Yan03c] Zhenya Yan. The new extended Jacobian elliptic function expansion algorithm and its applications in nonlinear mathematical physics equations. Computer Physics Communications, 153(2):145–154, June 15, 2003. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503002078.

Yan:2003:REV

[Yan03d] Zhenya Yan. The Riccati equation with variable coefficients expansion algorithm to find more exact solutions of nonlinear differential equations. Computer Physics Communications, 152(1):1–8, April 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502007567.

Yang:2009:FAM

[Yan09] Hong Wei Yang. A FDTD analysis on magnetized plasma of Epstein distribution and reflection calculation. Computer Physics Communications, 180(1):55–60, January 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508002889. See addendum and com-

ment [Ram10].

Yao:2009:SDE

[Yao09] Huanmin Yao. Solution of the Duffing equation involving both integral and non-integral forcing terms. Computer Physics Communications, 180(9):1481–1488, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001088.

Yan:2002:NCS

[YB02a] Zhenya Yan and George Bluman. New compacton soliton solutions and solitary patterns solutions of nonlinearly dispersive Boussinesq equations. Computer Physics Communications, 149(1):11–18, November 15, 2002. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502005878.

Yepez:2002:EAQ

[YB02b] Jeffrey Yepez and Bruce Boghosian. An efficient and accurate quantum lattice-gas model for the many-body Schrödinger wave equation. Computer Physics Communications, 146(3): 280–294, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465502004198.

Yip:2007:MCM

[YC07] Man Ho Yip and M. J. Carvalho. A Monte–Carlo maplet for the study of the optical properties of biological tissues. Computer Physics Communications, 177(12):965–975, December 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465507003712.

Yasar:2006:SHC

[YD06] F. Yasar and K. Demir. The study of helix-coil transition of polyalanine with single histogram method. Computer Physics Communications, 175(9):604–611, November 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550600275X.

Yasar:2007:SSG

[YD07] F. Yasar and M. Dilaver. The simulation of spin glass model of neural networks by the Wang–Landau algorithm. Computer Physics Communications, 177(8):625–630, October 15, 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507002974.

Yepez:2002:QCP

[Yep02] Jeffrey Yepez. Quantum computation for physical modeling. Computer Physics Communications, 146(3):277–279, July 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502004186.

Yamamoto:2009:HTH

[YFM09] Sumie Yamamoto, Keisuke Fujii, and Akiya Miyamoto. A handy tool for history keeping of Geant4 tracks — J4HistoryKeeper. Computer Physics Communications, 180 (6):827–834, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508003962.

Yuan:2009:LRA

[YG09] Zhe Yuan and Shiwu Gao. Linear response approach to collective electronic excitations of solids and surfaces. Computer Physics Communications, 180(3):466–473, March 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508003779.

Yamazaki:2002:MPT

[YGT^{+}02] Yoshitake Yamazaki, Herbert Gleiter, Junji Tani, Yutaka Abe, Keiichiro Nakamura, Haruyasu Mizoguchi, Tomohiro Irie, Masahiro Watanabe, and Dai Ito. Mesoscopic phase transitions and their critical behavior of nanostructured materials. Computer Physics Communications, 147(1–2):267–271, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502002862.

Yamasaki:2002:NAA

[YH02] Junko Yamasaki and Naomichi Hatano. A new algorithm of analyzing the metal-insulator transition of the Anderson model. Computer Physics Communications, 147(1–2):263–266, August 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550200259X.

Yun:2007:ISP

[YKK07] Younsuk Yun, Sang Chul Kwon, and Whung Whoe Kim. Ab initio study of planar faults in bcc Fe and Mo. Computer Physics Communications, 177(1–2):49, July 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550700135X.

Yousif:2003:CBF

[YM03] Hashim A. Yousif and Richard Melka. Computing Bessel functions of the second kind in extreme parameter regimes. Computer Physics Communications, 151(1):25–34, March 1, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006975.

Ma:2001:NSF

[yMS01] Chun yu Ma and Danny Summers. A numerical scheme for Fokker–Planck wave-particle interaction equations in plasmas. Computer Physics Communications, 135(2):133–143, April 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500002241.

Yamamoto:2005:FII

[YN05a] Shigeyoshi Yamamoto and Umpei Nagashima. Four-index integral transformation exploiting symmetry. Computer Physics Communications, 166(1):58–65, February 15, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504005089.

Yang:2005:GSP

[YN05b] Xiaobao Yang and Jun Ni. Ground states of potassium adsorbate on single-walled carbon nanotubes. Computer Physics Communications, 169(1–3):20–23, July 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505001037.

Yamamoto:2005:MRH

[YNK05] Ryoichi Yamamoto, Yasuya Nakayama, and Kang Kim. A method to resolve hydrodynamic interactions in colloidal dispersions. Computer Physics Communications, 169(1–3):301– 304, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505001645.

Yokota:2009:FMM

[YNS^{+}09] R. Yokota, T. Narumi, R. Sakamaki, S. Kameoka, S. Obi, and K. Yasuoka. Fast multipole methods on a cluster of GPUs for the meshless simulation of turbulence. Computer Physics Communications, 180(11):2066–2078, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001891.

Yang:2009:IET

[YNZ^{+}09] Hua Yang, Yuchao Nie, Hongbin Zhang, Zengru Di, and Ying Fan. Insight to the express transport network. Computer Physics Communications, 180(9):1511–1515, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001143.

Yokoi:2009:VAM

[Yok09] Kensuke Yokoi. A variational approach to multi-phase motion of gas, liquid and solid based on the level set method. Computer Physics Communications, 180(7):1145–1149, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046550900037X.

Yoshida:2000:NPM

[Yos00] Mototsugu Yoshida. A national project on mesoscopic simulation in japan. Computer Physics Communications, 127(1):100– 104, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000692.

Yoshie:2001:RRM

[Yos01] Tomoteru Yoshié. Recent results from massively parallel computers dedicated to lattice QCD. Computer Physics Communications, 142(1–3):36–42, December 15, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003101.

Yoshiki:2003:CSO

[Yos03] H. Yoshiki. The cross sections for one phonon emission and absorption by slow neutrons in superfluid liquid helium. Computer Physics Communications, 151(2):141–148, March 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502008196. See addendum

[Yos07].

Yoshiki:2007:ACS

[Yos07] H. Yoshiki. Addendum to: “The cross sections for one phonon emission and absorption by slow neutrons in superfluid liquid helium” [Comput. Phys. Comm. 151 (2) (2003) 141–148]: Angular distribution of UCN. Computer Physics Communications, 177(11):877–878, December 1, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465507002962. See [Yos03].

Yoshida:2009:SSS

[Yos09] Naoki Yoshida. Supercomputer simulations of structure formation in the Universe. Computer Physics Communications, 180(4):625–627, April 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004487.

Young:2002:SGC

[You02] A. P. Young. Spin glasses: a computational challenge for the 21st century. Computer Physics Communications, 146(1):107– 112, June 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465502004411.

Young:2005:RNR

[You05] Peter Young. Recent numerical results on spin glasses. Computer Physics Communications, 169(1–3):144–148, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505001311.

Younes:2009:GME

[You09] W. Younes. Gaussian matrix elements in a cylindrical harmonic oscillator basis. Computer Physics Communications, 180(7):1013–1040, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508004396.

Yoo:2007:PSS

[YRR07] Jongsoo Yoo, Tongnyeol Rhee, and Chang-Mo Ryu. 1D PIC simulation study of nonlinear beam plasma interaction. Computer Physics Communications, 177(1–2):93–94, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507001087.

Yagi:2009:DMS

[YSM09] Manabu Yagi, Kanako Seki, and Yosuke Matsumoto. Development of a magnetohydrodynamic simulation code satisfying the solenoidal magnetic field condition. Computer Physics Communications, 180(9):1550–1557, September 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001210.

Yanchitsky:2001:DSG

[YT01a] B. Z. Yanchitsky and A. N. Timoshevskii. Determination of the space group and unit cell for a periodic solid. Computer Physics Communications, 139(2):235–242, September 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002120.

Yoshinaga:2001:NAL

[YT01b] Naotaka Yoshinaga and Mitsuyoshi Tomiya. New approach to level statistics of coupled quartic oscillators. Computer Physics Communications, 142(1–3):88–94, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501003307.

Yurtsever:2002:MCR

[Yur02] Ersin Yurtsever. Measuring chaos in rotating clusters. Computer Physics Communications, 145(1):194–202, May 1, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465502001546.

Yu:2005:APM

[YvG05] Haibo Yu and Wilfred F. van Gunsteren. Accounting for polarization in molecular simulation. Computer Physics Communications, 172(2):69–85, November 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505003772.

Yu:2000:IFP

[YW00] K. W. Yu and Jones T. K. Wan. Interparticle force in polydisperse electrorheological fluids. Computer Physics Communications, 129(1–3):177–184, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465500001053.

Yu:2001:LFD

[YW01] K. W. Yu and Jones T. K. Wan. Local field distribution near corrugated interfaces: Green’s function formulation. Computer Physics Communications, 142(1–3):368–373, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501003812.

Yao:2004:INL

[YWLC04] Zhenhua Yao, Jian-Sheng Wang, Gui-Rong Liu, and Min Cheng. Improved neighbor list algorithm in molecular simulations using cell decomposition and data sorting method. Computer Physics Communications, 161(1–2):27–35, August 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504002097.

Yang:2009:ERT

[YWYF09] Hongli Yang, Xinyuan Wu, Xiong You, and Yonglei Fang. Extended RKN-type methods for numerical integration of perturbed oscillators. Computer Physics Communications, 180 (10):1777–1794, October 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465509001520.

See erratum [Wu10].

Yang:2007:SPC

[YZD^{+}07] R. Y. Yang, R. P. Zou, K. J. Dong, X. Z. An, and Aibing Yu. Simulation of the packing of cohesive particles. Computer Physics Communications, 177(1–2):206–209, July 2007. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465507000756.

Yang:2002:CDS

[YZW02] S. Y. Yang, Y. C. Zhou, and G. W. Wei. Comparison of the discrete singular convolution algorithm and the Fourier pseudospectral method for solving partial differential equations. Computer Physics Communications, 143(2):113–135, February 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465501004271.

Zimmermann:2001:OPL

[ZA01] Karl-Heinz Zimmermann and Wolfgang Achtziger. Optimal piecewise linear schedules for LSGP- and LPGS-decomposed array processors via quadratic programming. Computer Physics Communications, 139(1):64–89, September 1, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002314.

Zaharioudakis:2000:CST

[Zah00] D. Zaharioudakis. Calculation of shape-truncation functions for cubic cells. Computer Physics Communications, 130(1–2): 22–40, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000138.

Zaharioudakis:2001:CST

[Zah01] Dimitris Zaharioudakis. Calculation of shape-truncation functions for hexagonal cells. Computer Physics Communications, 140(3):323–345, November 1, 2001. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465501002491.

Zaharioudakis:2004:TMB

[Zah04] Dimitris Zaharioudakis. Tetrahedron methods for Brillouin zone integration. Computer Physics Communications, 157 (1):17–31, February 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465503004892.

Zaharioudakis:2005:QCT

[Zah05] Dimitris Zaharioudakis. Quadratic and cubic tetrahedron methods for Brillouin zone integration. Computer Physics Communications, 167(2):85–89, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505000603.

Zaki:2000:SWI

[Zak00a] S. I. Zaki. Solitary wave interactions for the modified equal width equation. Computer Physics Communications, 126 (3):219–231, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465599004713.

Zaki:2000:SWK

[Zak00b] S. I. Zaki. Solitary waves of the Korteweg–de Vries–Burgers’ equation. Computer Physics Communications, 126(3):207–218, April 11, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465599004701.

Zaki:2001:SWS

[Zak01] S. I. Zaki. Solitary waves of the splitted RLW equation. Computer Physics Communications, 138(1):80–91, July 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465501002004.

Zakrzewski:2006:HPS

[Zak06] A. J. Zakrzewski. Highly precise solutions of the onedimensional Schrödinger equation with an arbitrary potential. Computer Physics Communications, 175(6):397–403, September 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550600213X.

Zatsarinny:2006:BBS

[Zat06] Oleg Zatsarinny. BSR: B-spline atomic R-matrix codes. Computer Physics Communications, 174(4):273–356, February 15, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505005783.

Zhang:2006:CFS

[ZBB^{+}06] Cheng Zhang, Bhupesh Bansal, Paulo S. Branicio, Rajiv K. Kalia, Aiichiro Nakano, Ashish Sharma, and Priya Vashishta. Collision-free spatial hash functions for structural analysis of billion-vertex chemical bond networks. Computer Physics Communications, 175(5):339–347, September 1, 2006. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506002311.

Zasada:2009:VAS

[ZC09] S. J. Zasada and P. V. Coveney. Virtualizing access to scientific applications with the Application Hosting Environment. Computer Physics Communications, 180(12):2513–2525, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465509001830.

Zhou:2005:LLR

[ZDKG05] J. Zhou, A. Deyhim, S. Krueger, and S. K. Gregurick. LORES: Low resolution shape program for the calculation of small angle scattering profiles for biological macromolecules in solution. Computer Physics Communications, 170(2):186–204, August 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505002535.

Zaloj:2000:PCM

[ZE00] Veaceslav Zaloj and Ron Elber. Parallel computations of molecular dynamics trajectories using the stochastic path approach. Computer Physics Communications, 128(1–2):118–127, June 9, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000382.

Zatsarinny:2000:GPC

[ZF00] Oleg Zatsarinny and Charlotte Froese Fischer. A general program for computing angular integrals of the Breit–Pauli Hamiltonian with non-orthogonal orbitals. Computer Physics Communications, 124(2–3):247–289, February 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465599004415.

Zatsarinny:2009:ASC

[ZF09] Oleg Zatsarinny and Charlotte Froese Fischer. Atomic structure calculations using MCHF and BSR. Computer Physics Communications, 180(11):2041–2065, November 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S001046550900188X.

Zhang:2000:FTC

[Zha00] Shiwei Zhang. Finite-temperature calculations of correlated electron systems. Computer Physics Communications, 127 (1):150–155, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465500000291.

Zhang:2001:ESD

[Zha01] C. Zhang. Electronic states and dielectric response of an electron gas under a quantizing magnetic field and an intense laser. Computer Physics Communications, 142(1–3):374–381, December 15, 2001. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550100385X.

Zhan:2008:PIW

[Zha08] Lixin Zhan. A parallel implementation of the Wang–Landau algorithm. Computer Physics Communications, 179(5):339– 344, September 1, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0010465508001392.

Zhang:2000:IST

[ZHC00] Raoyang Zhang, Xiaoyi He, and Shiyi Chen. Interface and surface tension in incompressible lattice Boltzmann multiphase model. Computer Physics Communications, 129(1–3):121– 130, July 2000. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465500000990.

Ziegler:2004:ABA

[Zie04] Udo Ziegler. An ADI-based adaptive mesh Poisson solver for the MHD code NIRVANA. Computer Physics Communications, 157(3):207–216, March 1, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465503005447.

Ziegler:2005:SAC

[Zie05] U. Ziegler. A solution-adaptive central-constraint transport scheme for magnetohydrodynamics. Computer Physics Communications, 170(2):153–174, August 1, 2005. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465505002675.

Ziegler:2008:NCP

[Zie08] U. Ziegler. The NIRVANA code: Parallel computational MHD with adaptive mesh refinement. Computer Physics Communications, 179(4):227–244, August 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465508000969.

Zimmermann:2002:EDS

[Zim02] Karl-Heinz Zimmermann. Efficient DNA sticker algorithms for NP-complete graph problems. Computer Physics Communications, 144(3):297–309, April 15, 2002. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465502002709.

Zimmermann:2005:SCC

[Zim05] Karl-Heinz Zimmermann. Solving constrained combinatorial optimization problems via importance sampling in the grand canonical ensemble. Computer Physics Communications, 165(3):243–259, February 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465504004977.

Zitko:2009:ALD

[Zit09] Rok Zitko. Adaptive logarithmic discretization for numerical renormalization group methods. Computer Physics Communications, 180(8):1271–1276, August 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509000630.

Zhang:2005:QSR

[ZKASS05] Shiwei Zhang, Henry Krakauer, Wissam A. Al-Saidi, and Malliga Suewattana. Quantum simulations of realistic systems by auxiliary fields. Computer Physics Communications, 169(1–3):394–399, July 1, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465505001840.

Zhao:2009:MWA

[ZLL09] Yin-Long Zhao, Yin-Ping Liu, and Zhi-Bin Li. A modified WTC algorithm for the Painlevé test of nonlinear variable-coefficient PDEs. Computer Physics Communications, 180(11):2123–2128, November 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0010465509001957.

Zuccaro:2004:PCS

[ZLM04] G. Zuccaro, G. Lapenta, and G. Maizza. Particle in cell simulation of combustion synthesis of TiC nanoparticles. Computer Physics Communications, 162(2):89–101, September 15, 2004. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465504003297.

Zhang:2000:SDC

[ZM00] Kai Zhang and Charles W. Manke. Simulation of diblock copolymer melts by dissipative particle dynamics. Computer Physics Communications, 129(1–3):275–281, July 2000. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465500001144.

Zerbetto:2009:SES

[ZPB09] Mirco Zerbetto, Antonino Polimeno, and Vincenzo Barone. Simulation of electron spin resonance spectroscopy in diverse environments: An integrated approach. Computer Physics Communications, 180(12):2680–2697, December 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465509001982.

Zakynthinaki:2003:SOT

[ZS03] M. S. Zakynthinaki and Y. G. Saridakis. Stochastic optimization for a tip-tilt adaptive correcting system. Computer Physics Communications, 150(3):274–292, February 15, 2003. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465502006811.

Zakynthinaki:2007:SOM

[ZS07] M. S. Zakynthinaki and J. R. Stirling. Stochastic optimization for modeling physiological time series: application to the heart rate response to exercise. Computer Physics Communications, 176(2):98–108, January 15, 2007. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465506003213.

Zakynthinaki:2008:SOCb

[ZS08] M. S. Zakynthinaki and J. R. Stirling. Stochastic optimization for the calculation of the time dependency of the physiological demand during exercise and recovery. Computer Physics Communications, 179(12):888–894, December 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465508002695.

Zakynthinaki:2008:SOCa

[ZSdD^{+}08] M. S. Zakynthinaki, J. R. Stirling, A. López Díaz de Durana, C. A. Cordente Martínez, M. Sillero Quintana, and J. Sampedro Molinuevo. Stochastic optimization for the calculation of the optimal critical curve from experimental data in a model of the process of regaining balance after perturbation from quiet stance. Computer Physics Communications, 179(8):562–568, October 15, 2008. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465508001811.

Zaitsev:2004:NSS

[ZSK^{+}04] F. S. Zaitsev, A. G. Shishkin, D. P. Kostomarov, M. R. O’Brien, R. J. Akers, M. Gryaznevich, A. B. Trefilov, and A. S. Yelchaninov. The numerical solution of the selfconsistent evolution of plasma equilibria. Computer Physics Communications, 157(2):107–120, February 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465503004958.

Zhou:2005:CEB

[ZSM05] Yunkai Zhou, Ron Shepard, and Michael Minkoff. Computing eigenvalue bounds for iterative subspace matrix methods. Computer Physics Communications, 167(2):90–102, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465505000627.

Zia:2000:CBE

[ZSSA00] R. K. P. Zia, L. B. Shaw, B. Schmittmann, and R. J. Astalos. Contrasts between equilibrium and non-equilibrium steady states: computer aided discoveries in simple lattice gases. Computer Physics Communications, 127(1):23–31, May 1, 2000. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465500000229.

Zhao:2005:IFO

[ZWD05] Deyin Zhao, Zhongcheng Wang, and Yongming Dai. Importance of the first-order derivative formula in the Obrechkoff method. Computer Physics Communications, 167(2):65–75, April 15, 2005. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0010465505000585.

Zhang:2004:CNA

[ZWY04] Wei Zhang, Zhiming Wu, and Genke Yang. Chaotic network attractor in packet traffic series. Computer Physics Communications, 161(3):129–142, August 15, 2004. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465504002309.

Zhang:2009:LBM

[ZY09] Jianying Zhang and Guangwu Yan. A lattice Boltzmann model for the Korteweg–de Vries equation with two conservation laws. Computer Physics Communications, 180(7):1054– 1062, July 2009. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S001046550800444X.

Zeng:2009:ECV

[ZZ09] Rong Zeng and Yanzhi Zhang. Efficiently computing vortex lattices in rapid rotating Bose–Einstein condensates. Computer Physics Communications, 180(6):854–860, June 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465508004207.

Zhao:2009:NTS

[ZZH09] Qingjie Zhao, Baojun Zhang, and Huosheng Hu. Novel twostep filtering scheme for a logging-while-drilling system. Computer Physics Communications, 180(9):1566–1571, September 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-

2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0010465509001234.